### v1

# Exploring Holomorphic Retracts

Abstract.
The purpose of this article is towards systematically characterizing holomorphic retracts of domains of holomorphy; to begin with, bounded balanced pseudoconvex domains $B\subset\mathbb{C}^{N}$. Specifically, we show that every (holomorphic) retract of $B$ passing through its center (i.e., the origin), is the graph of a holomorphic map over a linear subspace of $B$. If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics; the necessity of the boundary points to be extreme points is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well. This revisit of Vesentini’s result leads to a small modification that is useful to provide interesting examples. As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ which is submersive at its center, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$. To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of strictly convex balls, with respect to various norms and obtain a complete description of retracts passing through the origin, generalizing that of the polydisc due to Heath and Suffrdige and giving a simpler proof as well. The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$. Along the way, we discuss examples of non-convex but pseudoconvex bounded balanced domains in which most one-dimensional linear subspaces fail to be holomorphic retracts, in good contrast to the convex case. Indeed, for the simplest possible case here namely the ‘$\ell^{q}$-ball’ for $10<q<$ 1, we obtain a complete list of all retracts through the origin; this enables an illustration of applying retracts to establishing biholomorphic inequivalences. To go beyond balanced domains, we then first obtain a complete characterization of retracts of the Hartogs triangle and ‘analytic complements’ thereof. Thereafter, we drop the assumption of boundedness as well, to attain similar characterization results for domains which are neither bounded nor topologically trivial. We conclude by reporting some results on the retracts of $\mathbb{C}^{2}$.
Key words and phrases:
Union problem, Kobayashi hyperbolic, Kobayashi corank one, balanced pseudoconvex domains
1991 Mathematics Subject Classification:
Primary: 32F45; Secondary: 32H15, 32H20
1. Introduction
We begin with remarks for motivation and background, possibly a bit elaborate; this owes to the unavailability of a survey or expository article on holomorphic retracts to refer the reader to. This will also simultaneously help us set up notations and fix terminologies. By a retraction we shall mean an idempotent map, whereas by a retract, we mean the image of such a map. More precisely, we shall be interested in holomorphic retracts i.e., the image of a holomorphic retraction which by definition, is a holomorphic self-map $\rho$ of a domain $D\subset\mathbb{C}^{N}$ (or more generally, a complex manifold) satisfying the idempotent condition $\rho^{\circ 2}:=\rho\circ\rho=\rho$. This means that $\rho$ acts as the identity map on $Z=\rho(D)$ which shall be our notation for a retract throughout this article; a trivia worth noting is that $Z$ is precisely the set of points of $D$ fixed by $\rho$. A pair of extreme or trivial examples of retractions (resp. retracts) of any complex manifold are the identity map and constant maps (resp. $D$ and its points); obviously, we would not like to talk of them any further and thereby shall be referred to as trivial retracts henceforth. For instance, with this terminology and understanding, we shall say that in dimension $N=1$, there are no non-trivial retracts. This means that for any domain $D$ in $\mathbb{C}$ or any Riemann surface, any holomorphic self-map $\rho$ satisfying $\rho^{\circ 2}=\rho$ is either a constant map or else, the identity map on $D$. Indeed, this follows from the open mapping theorem together with, the idempotency condition which implies that $\rho(D)$ is a closed subset of $D$. In other words, holomorphic retracts are interesting only when the dimension $N$ gets bigger than one. Now in multidimensional complex analysis, a lot of effort has gone into the study of biholomorphic and more generally, proper (holomorphic [^1] Unless otherwise mentioned, all maps in this article are to be understood to be holomorphic.) maps, as is natural. To mention just one or two of the first basic facts about proper maps between a pair of domains in $\mathbb{C}^{N}$, they are known to be always surjective and for many a class of domains $D$ with ‘nice’ boundaries, proper self-maps turn out to be automorphisms of $D$. Though absolutely trivial to check, let us mention for contrast with the just mentioned facts about proper maps, that: non-trivial retractions are never surjective (and never are they, automorphisms!); as one another sharp contrast, proper maps never decrease dimension, whereas non-trivial retractions always decrease dimension. In other words, retractions could be viewed as being, loosely speaking, at the ‘opposite end of the spectrum’ when compared to proper maps.
While much work has been carried out on proper holomorphic maps and their images, far less has investigation has been done about retracts. Infact, in addition to (the foregoing remarks and) their role of in certain fundamental mapping problems above, retracts are of significance in complex dynamics as well. To just be specific by providing a standard result in this regard, we mention the following theorem which shows how retracts arise in the study of iterates of a holomorphic self-map of taut manifolds. Needless to say, this theorem has been applied to the study of the special case of proper holomorphic self-maps. In light of the the stark contrasts between proper maps and retraction maps mentioned above, the connection amongst these maps, is thereby all the more interesting.
Theorem 1.1 (cf. Bedford, [Bdfrd]; Abate, [Abate_Horosphre]).
Let $X$ be a taut manifold, $f$ a holomorphic self-map of $X$. Suppose that the sequence $\left\{f^{\circ k}\right\}$ of iterates of $f$ is not compactly divergent. Then there exist a submanifold $M$ of $X$ and a holomorphic retraction $\rho:X\rightarrow M$ such that every limit point $h\in\operatorname{Hol}(X,X)$ of $\left\{f^{\circ k}\right\}$ is of the form
$h=\gamma\circ\rho$
where $\gamma$ is an automorphism of $M$. Infact, $\rho$ is itself a limit point of the sequence $\left\{f^{\circ k}\right\}$ and $f$ acts as an automorphism of $M$.
[…]
Let us also mention here for clarity that our focus will be to determine the form of the retracts $Z=F(X)$ rather than retractions mappings $F$. This owes to the fact that for a given retract $Z$, there are in general plenty of retractions all of whose images are the same $Z$. This is manifestly brought out already in the simplest of cases for $X$ namely, when $X$ is the bidisc $\Delta^{2}$, a description of whose retracts was attained (indeed for $\Delta^{N}$) in 1981 by Heath and Suffridge in [Heath_Suff]. The case of the Euclidean ball dealt with in 1974 by Suffridge in [Suffridge-Ball] is exposited well in Rudin’s book [Rud_book_Fn_theory_unitball] (we shall not detail the timeline of the results, beyond the most basic cases of the ball and polydisc). As highlighted in the introduction of this article by Heath – Suffridge, there is a wide range of holomorphic retractions of the bidisc with the same image, which we recall here for some later significance for our article. Let $10<t<$ 1 and pick an arbitrary holomorphic function $f:\Delta^{2}\to\mathbb{C}$ which is bounded above in modulus by $t(1-t)/2$. Then the mapping on $\Delta^{2}$, defined by
$(z,w)\mapsto\left[(1-t)z+te^{i\theta}w+\left(z-e^{i\theta}w\right)^{2}f(z,w)\right]\left(1,e^{-i\theta}\right),$
for any fixed $\theta\in\mathbb{R}$, is a holomorphic retraction of $\Delta^{2}$ onto $\left\{z\left(1,e^{-i\theta}\right):|z|<1\right\}$, as is checked by a straightforward computation which we skip here, to get ahead with building a perspective in this introductory section. Note that though such a retraction may well be highly non-linear and have a complicated expression as just exemplified, its image (regardless of what $f$ is), is just a linear copy of the disc. This finishes the demonstration that for a given retract $Z$, there may very well be plenty of (as remarked in [Heath_Suff] ‘quite complicated’!) retractions all of whose images are the same $Z$. At this juncture, it may tempting to ask as to whether this phenomenon of the existence of rather complicated holomorphic retractions with the same image, owes to the polydisc being a domain that reduces to a product of one-dimensional discs, thereby in-particular, giving rise to a plethora of analytic discs in the boundary. That this is not the case is borne out by the fact that the same is true (in-fact, in some sense all the more! – see below) for Euclidean balls, as is recorded in chapter-8 of Rudin’s book [Rud_book_Fn_theory_unitball] wherein complicated retraction maps of the Euclidean ball $\mathbb{B}^{2}\subset\mathbb{C}^{2}$, all with the same image, are exemplified. Indeed to recall this briefly here, pick any $f\in H^{\infty}(\mathbb{B}^{2})$ of the form
$f(z,w)=z+w^{2}g_{1}(z,w)+w^{4}g_{2}(z,w)+w^{6}g_{3}(z,w)+\ldots$
where the $g_{j}$’s are arbitrary elements of $H^{\infty}(\mathbb{B}^{2})$; then the associated self-map $F$ of $\mathbb{B}^{2}$ given by $F(z,w)=\left(f(z,w),0\right)$ is a retraction map, as simply follows from the fact that $f(z,0)\equiv z$. Its image is $\pi_{1}(\mathbb{B}^{2})=\Delta$, where $\pi_{1}$ is the linear projection onto the first factor. Hence again, while $f$ and thereby $F$ may well be be very complicated, the retract here is just a disc. We are therefore motivated to focus on retracts first more than retraction mappings.
Let us then begin with the obvious observation that retracts are always fixed point sets of some holomorphic map; but the converse need not be true in general. But then again, an easy but noteworthy consequence of theorem 1.1 is that the fixed point set of every holomorphic self-map of any taut manifold $X$, is contained within a non-trivial retract. Furthermore, there are results for an exact equality rather than just containment, if so sought for: as shown in [Mazet_Vigue_fxdpnt], [Thai_Fxdpnt_cnvx] among others, the fixed point set of any holomorphic self map of any hyperbolic convex domain in $\mathbb{C}^{N}$ – whenever it is non-empty – is also a retract of the domain. Whether or not fixed point sets of a holomorphic self map is the retract of possibly some other holomorphic self map of $X$, a definitive basic feature that all retracts have in common is that every retract $Z$ has the extension property in the sense that every holomorphic function on $Z$ admits (just by composition with the retraction mapping $X$ onto $Z$) a holomorphic extension to $X$. When the given function on $Z$ is bounded, we also have (by the same means of composition with the retraction map), a holomorphic extension on $X$ with the same sup-norm and thus retracts form the simplest class of subvarieties of $X$ with the norm-preserving extension property to $X$. Whether retracts are the only subvarieties with this property was examined and shown to be so, in many special cases by Kosiński and McCarthy in the relatively recent work [Kosnski] wherein counterexamples have also been given. All this is only towards indicating the significance and motivation for the study of holomorphic retracts, in general. Our first goal for this article is rather specific and it is: to obtain a description of the holomorphic retracts of pseudoconvex bounded balanced domains passing through its center, namely the origin in $\mathbb{C}^{N}$. This itself is noteworthy; if not for anything else, this suffices for an important application to attain (employing the above theorem 1.1), the fairly recent result of [JJ] together with [Pranav_JJ] that: every proper holomorphic self-map of any smoothly bounded balanced pseudoconvex domain of finite type is an automorphism. In-fact our methods help simplify a part of the proof of this result which we shall not however elaborate here.
As mentioned in the abstract, we intend a systematic presentation; so, to not be amiss about addressing the simplest of questions in the aforementioned first goal then, we recall that in [Vesntn], E. Vesentini already proved the following fundamental result about the form of the retracts of certain convex domains. More important than the various hypotheses of this result, as pointed out by J. P. Vigue in [Vigue_fxdpnt_cnvx], Vesentini’s approach makes ‘heavy’ use of complex geodesics. We shall to begin with, give a relatively elementary proof of this result without using complex geodesics. It may be noted that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal and thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$ but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal while it is indeed $\mathbb{C}$-extremal. The notions of $\mathbb{R}$-extremality, $\mathbb{C}$-extremality and the relationship between them is recalled in the following section 2 about preliminaries.
Theorem 1.2 (Vesentini).
Let $D$ be a bounded balanced convex domain in $\mathbb{C}^{N}$ with $\mathbb{C}$-extremal boundary. Then every retract of $D$ passing through origin is a linear subspace of $D$.
Here and throughout the article, by a linear subspace $D_{L}$ of a balanced domain $D\subset\mathbb{C}^{N}$, we mean that $D_{L}$ is of the form $D_{L}=D\cap L$, where $L$ is a linear subspace of $\mathbb{C}^{N}$. Likewise, by an affine linear subspace $A^{\prime}$ of $D$, we mean $A^{\prime}=A\cap D$ where $A$ is an affine linear subspace of $\mathbb{C}^{N}$. To recall a relevant fact in this parlance, every retract of a ball (centered at the origin, as always in this article) linearly equivalent to the standard Euclidean ball $\mathbb{B}^{N}$ is affine linear. We postpone rendering clarity about a potential confusion about the notion of linear retracts to section 3; likewise, for definitions of a couple of terms as well, in order to focus more in this introductory section, on the narrative of our results and work about retracts in perspective of what is already known.
Now, the condition of convexity on $D$, which entails a certain condition to be satisfied by the entire boundary $\partial D$, can be dropped in various ways. To begin with here, let us only remark that we can do so in the following manner, which will actually be useful in dealing with certain concrete (counter-)examples in the sequel. Specifically, we shall show the following (without using ‘complex geodesics’).
Theorem 1.3.
(A small modification of Vesentini’s result): Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a retract through the origin with the property that $D\cap T_{0}Z$ is a convex domain in $T_{0}Z$ with $\mathbb{C}$-extremal boundary, then $Z$ is a linear subspace of $D$. In particular, every $1$-dimensional retract through the origin of a bounded balanced pseudoconvex domain $D$ with $\mathbb{C}$-extremal boundary, is linear.
We emphasize that this result is not really new and our inclusion of this is towards a relatively self-contained exposition, and more importantly, in the proof of proposition 5.1 below. Indeed, we would not like to be amiss to observe associated facts which are not explicitly recorded in the literature but can be attained by reading proofs of results carefully (and such observations are actually useful for us much later in the sequel): for instance, the apparently strong condition (when $\dim(Z)\geq 2$) on the convexity of $D\cap T_{0}Z$ in the above theorem can be dropped while attaining the same conclusion; specifically, we have the following.
Proposition 1.4.
(A small modification of proposition $4.5$ of [JJ]). Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a retract through the origin with the property that the boundary points of $D:=D\cap T_{0}Z$ viewed as a domain in $T_{0}Z$ are holomorphically extreme. Then, $Z$ is linear.
Amongst other results about retracts, this helps in establishing biholomorphic inequivalences, particularly with product domains. Stated differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., biholomorphic to a product domain. Let us only indicate the key idea here namely, the fundamental common feature amongst all product domains that: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed. Let us just state precisely here the simplest result to this effect to indicate what this means; leaving more general formulations to be laid out in section 7 wherein we shall have other remarks to tie up with.
Proposition 1.5.
Let $D_{1},D_{2}$ be domains in $\mathbb{C}$ with one of them bounded, say $D_{1}$. Then there exists $\epsilon>0$ such that: for every vector $v_{2}\in T_{p_{2}}D_{2}\simeq\mathbb{C}$ with $|v_{2}|<\epsilon$, there exists a retract $Z$ of $D_{1}\times D_{2}$ passing through $p=(p_{1},p_{2})$ with $(1,v_{2})\in T_{p}Z$.
As already alluded to, this paves the way for establishing the following.
Proposition 1.6.
Let $D$ be any bounded balanced domain of holomorphy in $\mathbb{C}^{2}$ (whose Minkowski functional we denote by $h=h_{D}$), which is not convex at an open dense subset of $\partial D$ (that there exist plenty of domains will be discussed later); here by not ‘convex at a boundary point’ $p\in\partial D$, means that there are no ‘supporting linear functionals’ for $p$ i.e., there is no $\mathbb{C}$-linear functional $\ell:\mathbb{C}^{2}\to\mathbb{C}$ with $|\ell|\leq h$ and $|\ell(p)|=h(p)$. Let $A$ be any proper analytic subset of $D$ not passing through the origin. Then, $D\setminus A$ is not biholomorphic to a product domain.
An extension of this to higher dimensions is briefly discussed in section 7. We also hasten to mention that there are better techniques to detect the occurrence of a product structure for domains far beyond the class of balanced domains owing particularly to the very recent work [Bh-Bo-Su], which however requires some quantitative knowledge about the squeezing function. However we believe that for balanced domains, the above proposition gives a much simpler approach in practice, since a balanced domain is most often given by its Minkowski functional. Needless to say, there are plenty of balanced domains of holomorphy satisfying the non-convexity condition in the above proposition, as will also be discussed in section 7.
Getting back to retracts, let us start by noting that while Vesentini’s theorem says that all retracts of $D$ (as in that theorem) passing through the origin, are linear, it does not say that all linear subspaces are retracts. Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]). Indeed, the question of precisely which linear subspaces are realizable as retracts of $D$ as in the above theorem, is quite an interesting question and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections (when we measure the norm of the projection with respect to the operator norm induced by $\mu$)? The answer turns out to vary even among the $l^{p}$ norms i.e., depend upon $p$ when we take $\mu=\|\cdot\|_{l^{p}}$ even if we exclude the case $p=2$, which may not be surprising at all for this case is distinguished as the one that arises from an inner product; specifically, the case $p=\infty$ is also different from the other cases. Before getting to such differences, let’s mention a unified result. As our interest lies in retracts, we shall phrase in it in the language of retracts rather than norm-one projections. Specifically, the characterization of linear retracts of the standard $l^{p}$ ball ($1\leq p<\infty,~{}p\neq 2$) can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
Theorem 1.7.
(theorem 2.a.4 of [Lindstraus]): Let $Y$ be a linear subspace of $\mathbb{C}^{N}$ equipped with $\|\cdot\|_{l^{p}}$ where $1\leq p<\infty,p\neq 2$. Then $Y\cap B_{\ell^{p}}$ is a linear retract of $B_{\ell^{p}}$ iff there exist vectors $\{u_{j}\}_{j=1}^{m}$ of norm 1 in $\mathbb{C}^{N}$ of the form
$u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i},~{}1\leq j\leq m,~{}with~{}\sigma_{j}\cap\sigma_{k}=\phi~{}for~{}k\neq j$
such that $Y=span\{u_{j}\}_{j=1}^{m}$, where $\sigma_{j}:=supp(u_{j})=\{1\leq i\leq N:u_{j}(i)\neq 0\}$ (and $\lambda_{i}$’s are some complex numbers).
The case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, in which all linear subspaces are retracts. We would rather like to draw attention to the more important fact that this result is not true for $p=\infty$ to demonstrate which (by a counterexample as simple as possible), we include an example in section 3. This naturally leads to the question as to whether it is possible at all to pin down precisely which linear subspaces of the polydisc $\Delta^{N}$ can be realized as linear retracts (recall that the polydisc is the exceptional case corresponding to $p=\infty$ which is left out by theorem 1.7). Towards this, we first note the following characterization of those linear subspaces of $\Delta^{N}$ which are its retracts. Indeed, this can readily derived from [Heath_Suff], though it is not stated in these explicit terms; contained within our treatment will be a simpler proof of the main result of [Heath_Suff] about the characterization of all retracts of the polydisc as well as a deduction of the following special one about its linear retracts.
Proposition 1.8.
A linear subspace $L$ of the polydisc $\Delta^{N}$ is a holomorphic retract of $\Delta^{N}$ if and only if $L=Y\cap\Delta^{N}$, where $Y$ is a linear subspace of $\mathbb{C}^{N}$ which can be expressed as the span of vectors $v_{1},\ldots,v_{r}$ (for some integer $r$ between $1$ and $N$) all from the boundary of the polydisc, satisfying the following conditions:
  1. (i)
    there exists a subset of indices $J=\{j_{1},\ldots j_{r}\}\subset\{1,2,\ldots,N\}$, such that for each $k=1,\ldots,r$, the vectors $v_{k}$ may be expressed in terms of the standard basis vectors $\{e_{i}\}_{i=1}^{N}$ of $\mathbb{C}^{N}$, as:
    $v_{k}=\left(\sum_{i\notin J}c_{i,j_{k}}e_{i}\right)+e_{j_{k}}$
  2. (ii)
    For each $m\in\{1,2,\ldots,N\}\setminus J$, the scalars $c_{m,j}$, satisfy $\sum_{j\in J}|c_{m,j}|\leq 1$.
In particular, not every two dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the $\mathbb{B}^{N}$. We must not be amiss to note however that every one dimensional linear subspace of a bounded balanced convex domain in $\mathbb{C}^{N}$ is a retract – this follows by the Hahn–Banach theorem. It is natural to ask whether this holds if we relax the convexity condition on the domain to pseudoconvexity. The answer is in the negative, as was noted by Lempert in [Lempert-FundArt] by indicating an example to complement his passing observation therein that the existence of $2$-dimensional retracts is rather restricted and to the effect that when convexity is relaxed, even $1$-dimensional retracts can fail to exist in abundance. The only rigorous record of the just-mentioned Lempert’s example is in [Jrncki_invrnt_dst] (specifically in example 11.3.10 therein). While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimension as well; moreover this can be achieved entirely through elementary means. In-fact, our example is as simple as the $l_{q}$-‘ball’ for $q<1$ and we record this in the following proposition. Let us also note that while the above-mentioned example of Lempert is a pseudoconvex complete Reinhardt domain much as ours, we believe recording the aforementioned $\ell^{q}$-‘ball’ is worthwhile due to the fact that the literature about retracts of dimension higher than one, is scant at-least compared to complex geodesics and one-dimensional retracts.
Theorem 1.9.
[…]
Consequently, the only $r$-dimensional linear retracts of $D_{q}$ are those spanned by an $r$-element subset of the standard basis $\{e_{1},\ldots,e_{N}\}\subset\mathbb{C}^{N}$. Thus, the holomorphic retracts through the origin of $D_{q}$ are precisely those linear subspaces spanned by some subset of the standard basis vectors.
Next, we get back to the discussion about Vesentini’s result. Let us remark a word or two, about concrete examples of domains satisfying the $\mathbb{C}$-extremality condition in his result. Firsly, all $\ell^{q}$-balls for $1\leq q<\infty$ satisfy this condition, whereas a product of any two of them fails to, much as the polydisc. While the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary, as will be verified in proposition 3.1 in section 3. Vesentini’s result applies to pin-down the retracts through the center of such a ‘ball of intersection’.
Next, we may want to go beyond the class of domains in Vesentini’s result. As the most natural class of domains in several complex variables are pseudoconvex domains, it is only natural to ask if convexity in the above theorem of Vesentini can be relaxed to pseudoconvexity and the extremality condition dropped. Towards this, we have the following result wherein we assume boundedness but it must be noted that it may be replaced by hyperbolicity (allowing unboundedness), which however is not significant here, owing to an old result of Kodama in [Kodama] according to which any hyperbolic balanced domain is equivalent to a bounded one.
Theorem 1.10.
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\longrightarrow Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a (single) holomorphic function over $D_{L}$.
Relegating some relevant observations here to section 3, we record here a corollary for an important class of domains. Indeed, when $D$ as in the above theorem (without any extremality condition about any boundary point), is homogeneous i.e., ${\rm Aut}(D)$ acts transitively on $D$, then of-course we may glean a characterization of all retracts of such a domain i.e., retracts not passing through the origin as well.
Corollary 1.11.
Let $D$ be any homogeneous bounded balanced domain in $\mathbb{C}^{N}$. Then, every non-trivial retract $Z$ (not necessarily passing through the origin in $D$) is biholomorphic via a member of ${\rm Aut}(D)$ to the graph of a holomorphic map over a linear retract $D_{L}$ (which as usual is a ball with respect to some norm in the lower dimensional complex Euclidean space $L=T_{0}Z$), of $D$.
Owing to the assumption of being balanced, we recall that the domains as in this corollary are the much studied bounded symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV alongwith two exceptional cases, all of whose automorphism groups are well-known (cf. Hua [Hua]). This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball (as in theorem 1.7 for the $l^{p}$-ball for $p<\infty$ and theorem 1.8 for the $p=\infty$ case much earlier above), is in general quite difficult. Moreover, it must be remembered that an orthogonal projection of such $l^{p}$-balls onto a linear subspace in general is not a retract, except for the case $p=2$. As these are classical facts, the significance of linear retracts from the viewpoint of recent research work may need a remark; let us only mention that it was shown by Mok in his fairly recent work [Mok], that every linear totally geodesic submanifold through the origin of any bounded symmetric domain $\Omega$ with respect to a ${\rm Aut}(\Omega)$-invariant Kähler metric (for instance, the Bergman metric) is realizable as a retract just by orthogonal projection. As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem, wherein which we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.10. It must first be noted that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifeslty seen in the simplest case of polydiscs. This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in simple cases such as domains decomposable into a product of balls (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc. Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains. This theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like. We shall address this here, to the extent we can.
Theorem 1.12.
Let $B_{1},B_{2}$ be a pair of bounded balanced strictly convex domains possibly in different complex Euclidean spaces. Then, every non-trivial retract of $D=B_{1}\times B_{2}$ through the origin is a holomorphic graph over such a linear subspace which itself is a retract of $D$. Infact, every non-linear retract of $D$ is the graph in $D$, either of a $B_{2}$-valued holomorphic map over a (complex) linear subspace of $B_{1}$ or, of a $B_{1}$-valued holomorphic map over a linear subspace of $B_{2}$ which itself is a linear retract of the respective ball.
More specifically and precisely, let $B_{1}\subset\mathbb{C}^{N_{1}}$ and $B_{2}\subset\mathbb{C}^{N_{2}}$ be as above i.e., bounded balanced strictly convex domains. Let $D=B_{1}\times B_{2}$ and suppose without loss of generality that $\min\{N_{1},N_{2}\}=N_{1}$. Let $\mu_{1},\mu_{2}$ denote the Minkowski functionals of $B_{1},B_{2}$ respectively, whose convexity ensures that $\mu_{1},\mu_{2}$ are norms on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively. Let $Z$ be any non-trivial retract of $D$, $L=T_{0}Z$, the tangent space to $Z$ at the origin, $D_{L}=L\cap D$ and $\partial D_{L}=L\cap\partial D$. Then:
  • (i)
    if $\partial D_{L}\subset\partial B_{1}\times\partial B_{2}$, then $Z$ is a complex linear subspace of $D$ which can be realized as the graph of a complex linear map over the linear subspace $\pi_{1}(D_{L})$ of $B_{1}$ (the ball whose dimension is $\min\{N_{1},N_{2}\}$):
    $Z=\{\big{(}w,\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}\;:\;w\in\pi_{1}(D_{L})\}$
    where the map
    $w\longmapsto\big{(}\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}$
    is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$, isometrically into $B_{2}$; $\pi_{1}(D_{L})$ itself is a linear retract of $B_{1}$,
  • (ii)
    if $\partial D_{L}\subset\partial B_{1}\times\overline{B}_{2}$, then $Z$ is the graph of a $B_{2}$-valued holomorphic map of a linear subspace $(B_{1})_{L_{1}}$ of $B_{1}$, with $(B_{1})_{L_{1}}$ being a linear retract of $B_{1}$,
  • (iii)
    if $\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$, then $Z$ is the graph of a $B_{1}$-valued holomorphic map of a linear subspace $(B_{2})_{L_{2}}$ of $B_{2}$, with $(B_{2})_{L_{2}}$ being a linear retract of $B_{2}$,
  • (iv)
    if $\partial D_{L}$ intersects both $\partial B_{1}\times B_{2}$ and $B_{1}\times\partial B_{2}$ – which are disjoint open pieces of the boundary of $D$, indeed connected components of $\partial D\setminus(\partial B_{1}\times\partial B_{2})$ – then $Z$ is a linear retract of $D$ of complex dimension $\geq 2$, which is neither contained in $B_{1}$ nor in $B_{2}$.
We cannot trivially extend this theorem to the case of a product $D=B_{1}\times B_{2}\times\ldots\times B_{n}$ – where $B_{j}$’s are all bounded balanced convex domains each of whose boundaries satisfy the strict convexity condition – just by replacing one of the factors in the above theorem by a product of $(n-1)$ of these balls i.e., writing for instance $D=(B_{1}\times B_{2}\times\ldots\times B_{n-1})\times B_{n}$, appealing to the above theorem and using an inductive argument, owing to the fact that $B_{1}\times B_{2}\times\ldots\times B_{n-1}$ fails to satisfy the strict convexity condition stipulated in the above theorem as soon as $n>2$. However, such an extension is indeed possible and not difficult either; provided only that we repeat the same arguments multiple times by projecting onto the individual factors $B_{j}$ to harness the strict convexity condition satisfied by each of these factor balls of the product $D$. As the complete formulation of the end result of this, turns out to be lengthy, we shall just lay it down in section 3. Here we record an application to the union problem for a nice class of polyballs which is interesting owing to being a case with some degeneracy. Recall that the union problem comprises the following: let $M$ be a complex manifold which is the union of an increasing sequence of open subsets $M_{j}$ each of which is biholomorphic to a fixed domain $\Omega\subset\mathbb{C}^{N}$; describe $M$ in terms of $\Omega$.
Theorem 1.13.
Take $\Omega$ in the union problem to be any homogeneous bounded balanced domain. If $M$ is non-hyperbolic and the corank of its (infinitesimal) Kobayashi metric is one, then $M$ is biholomorphic to $\Omega_{L}\times\mathbb{C}$ where $\Omega_{L}$ is a linear retract of $\Omega$.
Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is all the more significant. Towards this, we report that the following result.
Theorem 1.14.
Let $D$ be a bounded balanced strictly convex domain in $\mathbb{C}^{N}$. Let $Z$ be a retract of $D$ (not necessarily passing through origin) and suppose at-least one of the retraction maps for it say $\rho$, is a submersion at the origin. Then $Z=\rho(D)$ is the graph of a holomorphic function over some linear subspace $D_{L}$ of $D$. Indeed, $D_{L}=L\cap D$ where $L=T_{p}Z$ and $p=\rho(0)$.
Now, we have for the most part focussed attention on balanced domains which being contractible may be regarded as being topologically trivial. To turn next to topologically non-trivial cases, one simple case to begin with in this direction is that of analytic complements of domains with a Stein neighbourhood basis; specifically, see lemma 7.1. Next, to consider the simplest domain lacking a Stein neighbourhood basis then, we are led to considering the standard domain in this context namely the Hartogs triangle given by
$\Omega_{H}=\{(z,w)\in\mathbb{C}^{2}\;:\;|w|<|z|<1\}.$
We also contend that this is perhaps the simplest topologically non-trivial case to consider because it is biholomorphic to the product $\Delta\times\left(\Delta\setminus\{0\}\right)$; indeed the map $\phi(z,w)=(w/z,z)$, effects the aforementioned biholomorphic correspondence. Thus it suffices to determine the retracts of $\Delta\times\Delta^{*}$, which we do in the following theorem which leads to a complete characterization the retracts of the Hartogs triangle as well as its ‘analytic complements’.
Theorem 1.15.
Every non-trivial retract of $\Delta\times\Delta^{*}$ is a holomorphic graph over one of the factors; more precisely, either of the form $\{(z,f(z)):z\in\Delta\}$, where $f$ is a holomorphic map from $\Delta$ to $\Delta^{*}$, or of the form $\{(g(w),w):w\in\Delta^{*}\}$ for some holomorphic map self map $g$ of $\Delta$. Moreover, upto conjugation by an automorphism of the bidisc $\Delta^{2}$, every non-trivial retraction map is of the form
$\left(z,w\right)\rightarrow\left(z,f\left(z\right)\right)$
for some holomorphic $f:\Delta\to\Delta^{*}$, or of the form
$\left(z,w\right)\rightarrow\left[tz+(1-t)e^{-i\alpha}w+\left(e^{-i\alpha}w-z\right)^{2}h\left(z,w\right)\right]\left(1,e^{i\alpha}\right)$
for some real $\alpha$, $t\in(0,1)$ and holomorphic function $h:\Delta^{2}\to\Delta$.
Finally, retracts of $\Omega_{H}^{A}:=\Omega_{H}\setminus A$ where $A$ is any analytic subset of $\Omega_{H}$ are precisely given by $Z\cap\Omega_{H}$ where $Z$ is a retract of $\Omega_{H}$.
While retracts of $\Omega_{H}$ as in the above theorem, will be deduced as an immediate corollary of lemma 7.1 mentioned above, what does not follow so immediately are retracts of its analytic complements $\Omega_{H}^{A}$. This is detailed in section 8.
Next, it may be noted that an overarching hypothesis that we have imposed on all domains considered thus far, is their boundedness or equivalently (in our context of balanced domains here), hyperbolicity. To deal with the simplest case of unbounded and non-hyperbolic domains then, we first consider a product domain in which one of the factors is still a bounded domain but the other unbounded. We thought it better to separately record the result for the simplest case of this kind, namely of the form ×$\mathbb{C}\times D$ where $D$ is a bounded planar domain; the characterization of its retracts is laid down in the following.
Theorem 1.16.
If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of ×$\mathbb{C}\times D$ is either of the form ×$\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$.
It must be noted that this also addresses the other implicit assumption in most of our previous considerations, namely that of topological triviality. Indeed, the above theorem allows for domains with a variety of fundamental groups; more such variety is facilitated by the next set of results.
Proposition 1.17.
Let $X=\mathbb{C}\setminus A$ where $A$ is any discrete set (allowed to be empty as well), and $D$ any bounded domain in $\mathbb{C}^{N}$. Then, every retract $\widetilde{Z}$ of $X\times D$ is of the form
$\widetilde{Z}=\{(f(w),w):w\in Z\},$
where $Z$ is a retract of $D$ and $f$ is a holomorphic function from $Z$ to $\mathbb{C}\setminus A$, or of the form $X\times Z$ where again $Z$ is a retract of $D$.
While this article is focussed on characterization of retracts rather than retraction mappings, we thought it noteworthy to record in passing, a complete characterization of all retraction maps on the most special domain among the domains in the above theorem namely, ×$\mathbb{C}\times\Delta$.
Theorem 1.18.
Every non-trivial retraction map $F$ of ×$\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
While this article also focusses attention for the most part on domains in $\mathbb{C}^{N}$, we would like to interject here a result about retracts of manifolds, as it is a by-product of the essential ideas in the proof of the foregoing theorem 1.16 and more importantly, includes cases with far greater topological complexity.
Theorem 1.19.
Consider the following two cases for a product manifold $X\times Y$.
  • (i)
    $X$ is any compact Riemann surface and $Y$ is a non-compact Riemann surface,
  • (ii)
    $X$ is a non-hyperbolic Riemann surface and $Y$ is a (Kobayashi) hyperbolic complex manifold.
Then, every holomorphic retract $\widetilde{Z}$ of $X\times Y$ is either of the form
$\widetilde{Z}=\{(f(w),w):w\in Z\},$
where $Z$ is a retract of $Y$ and $f$ is a holomorphic map from $Z$ to $X$, or of the form $X\times Z$ where again $Z$ is a retract of $Y$.
Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{2}$ – indeed, this is the simplest of balanced domains of holomorphy or for that matter, the simplest of domains with non-trivial retracts and thereby too special to be remiss of us, if we do not remark about this atleast – and ask for a complete characterization of all of its holomorphic retracts (which seems unknown). Despite the apparent simplicity of the domain, this seems rather non-trivial; indeed, while graphs (in $\mathbb{C}^{2}$) of entire functions on $\mathbb{C}$ are retracts and the fact that retracts of products of planar domains in the foregoing, have literally turned out to be graphs of holomorphic maps over one of the factors, it is not the case that: every retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ is the graph of some entire function on either of its factors, as we shall show by providing a concrete example in the last section 9. On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this is attained by the following theorem.
Theorem 1.20.
If $R$ is a non-trivial polynomial retract of $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ on $\mathbb{C}^{2}$ such that $\psi(R)=\{(\zeta,0)\in\mathbb{C}^{2}\;:\;\zeta\in\mathbb{C}\}$.
Acknowledgements
This article forms the thesis work of the second author and we thank Prof. Kaushal Verma for giving us this problem of determining holomorphic retracts; as well as for helpful discussions, for which we also thank Dr. J. Jaikrishnan.
2. Preliminaries
We recall certain basic and general facts about retracts which would needless to say, apply to our specific case and put it in good perspective, as well. Let us first recall that a self-map $F$ of a set $X$ is called a retraction if $F\circ F=F$; note that this idempotent condition on $F$ means that $F$ acts as the identity map on its range $Z=F(X)$. So, for any Hausdorff topological space $X$, its retracts, which are the images of continuous retractions, are closed subsets of $X$. If $X$ is connected, then so is any of its retracts $Z$. When $X$ is furthermore contractible i.e., the space $X$ deformation retracts to a point then by composing any such deformation retraction with the retraction map $F$, gives a homotopy within $Z$ that reduces $Z=F(X)$ to a point, meaning that if $X$ is contractible then so are all of its retracts. If $X$ even happens to be a convex domain in some real topological vector space, then the retracts of $X$ are also strong deformation retracts. We are ofcourse interested in holomorphic retracts: if $X$ is a complex manifold and $F$ is a holomorphic self-map of $X$ which is idempotent (i.e., $F\circ F=F$) then, we say that $F$ is a holomorphic retraction and its image $Z=F(X)$ a holomorphic retract; as we shall only be talking about holomorphic retracts, we shall most often drop the adjective ‘holomorphic’. Moreover as mentioned in the introduction, our focus will be to determine the form of the retracts $Z=F(X)$ rather than the retractions $F$. A first basic and general fact about holomorphic retracts that is helpful to have in the background for the discussion in the sequel is laid down in the following well-known theorem which says that they are not just analytic varieties but actually (connected) manifolds i.e., are smooth and have the same dimension throughout.
Theorem 2.1.
Let $X$ be any complex manifold. Then any of its retracts is a (connected) complex submanifold of $X$ which is closed in $X$.
An exposition by Abate of the proof due to H. Cartan of this theorem, can be found in the monograph [Metrc_dynmcl_aspct] (specifically, lemma 2.1.3 therein). Abate calls this proof ‘clever’ and interestingly, this rather short and slick proof was given by Cartan only in 1986 in [Cartan-on-retracts]; Abate also notes that the theorem was proven for the first time by Rossi in [Rossi]. We sketch here, a flawed proof – the only place in this article where we deliberately do so – based on standard ideas in this context, for the sole purpose that it will serve in introducing some of the ideas (and our notations) to be used later while outlining the reasons underlying the necessity of a more careful approach as in the aforementioned proofs by Rossi and Cartan, to appreciate them better. We shall of-course clearly point out the flaw at the appropriate juncture, so as to be able to utilize the remaining (correct!) facts detailed in what follows. Indeed then, let $\rho:X\to Z$ be a retraction of $X$ onto $Z$. Rewrite the defining equation of $\rho$ being a retraction map namely, $\rho(\rho(x))=\rho(x)$ for all $x\in X$, in its infinitesimal form by differentiating using the chain-rule, as
$D\rho(\rho(x))D\rho(x)=D\rho(x),$
which holds for all $x\in X$. For $x=z\in Z$, $\rho(x)=x$ and so this simplifies to say that for every $z\in Z$, $D\rho(z)$ is an idempotent linear operator on $T_{z}X$ (the tangent space to $X$ at $z$); although fairly trivial, this will be of much use in the sequel, so we record this as a lemma for later reference.
Lemma 2.2.
Let $X$ be a complex manifold of dimension $n$, $Z$ a retract of $X$ and $\rho:X\to Z$ a retraction map. Then, for any $p\in Z$, the derivative $D\rho(p)$ is a linear projection (thereby a retraction) of $T_{p}X$ (the tangent space to $X$ at $p$) onto some linear subspace $L$ of $T_{p}X$.
As the assertion is essentially ‘local’, we may fix an arbitrary point $p\in Z$ and work in a chart about $p$. In particular, consider the mapping on a chart neighbourhood $U_{p}$ of $p$, given by
[…]
Denoting $D\rho(p)$ by $\pi_{p}$ we have
$DF_{p}(x)=\pi_{p}\circ D\rho_{p}=\pi_{p}^{\circ 2}=\pi_{p}$
implying that $F_{p}$ is of full rank. Note that $F_{p}$ maps $U_{p}$ into $L$.
Continuing with the proof theorem 2.1, we shall now show that in a neighbourhood of any given $p\in Z$, $Z$ is the level set of a holomorphic map whose derivative has constant rank near $p$; the implicit mapping theorem will then render the claim of the theorem. To this end then, we fix an arbitrary $p\in Z$ and passing to a holomorphic chart of the complex manifold $X$ near $p$, we may as well assume (as already mentioned), that we are in the complex Euclidean space $\mathbb{C}^{N}$, which allows us to define $\beta(x)=x-\rho(x)$ for $x$ near $p$, whose zero set is precisely $Z$ (all in the charts near $p$, to be precise). To show that the constancy of the rank of the derivative $D\beta(x)$ as mentioned, we first consider the derivative of $\rho$. The idempotence of the linear operator $D\rho(p)$ given by the above lemma, guarantees its diagonalizability as well. So, after a linear change of coordinates, $D\rho(p)$ is a diagonal matrix with diagonal entries either $1$ or 00. Let $r$ be the number of $1$’s i.e., $r={\rm rank}(D\rho(p))$. We may as well assume by a permutation of coordinates if needed, that the first principal $r\times r$ submatrix of $D\rho(p)$ is the identity matrix of order $r$, which by continuity of the determinant function implies that the determinant of the same principal minor of the nearby derivative matrices $D\rho(x)$ for all $x$ in some neighbourhood (contained in the chart) about $p$, is non-zero as well. This just means the well-known lower semi-continuity of the rank function, which when applied to $x\mapsto{\rm rank}(D\rho(x))$ implies that the rank of $D\rho(x)$ is atleast $r$ for all $x$ in some neighbourhood of $p$. Now, observe that the same arguments may be applied to $D\beta(x)$ as well. Firstly, note that $D\beta(p)=I-D\rho(p)$ where $I$ is the identity matrix of size $n$, is also an idempotent matrix, indeed diagonal with diagonal entries only $1$ or 00 and with the number of zeros being exactly $n-r$. In particular, ${\rm rank}(D\beta(p))=n-r$ and the lower semicontinuity of the rank applied to $x\mapsto{\rm rank}(D\beta(x))$ implies that the rank of $D\beta(x)$ is atleast $n-r$ in a neighbourhood of $p$. Now, we have a similar lower bound on the rank of $D\rho(x)$ as already noted, which is tempting to be translated this into an bound in-terms of $\beta$ instead of $\rho$; specifically, the upper bound ${\rm rank}(D\beta(x))\leq n-r$ in the some neighbourhood of $p$. While this is correct for $x=p$ and even along $Z$, it is not at all apriori clear as to why this bound continued to hold in an open neighbourhood of $p$. The reason for detailing this flawed argument is that it is then easy to finish the proof, within standard elementary ideas of calculus on differential manifolds. Indeed firstly, the foregoing pair of bounds when put together, only means that the rank of $D\beta(x)$ is constant in some neighbourhood of $p$, leading to the conclusion by the constant rank theorem that $Z$ is a complex manifold of dimension $r$ throughout, as we already know $Z=\rho(X)$ is connected but alas, this part of the argument is flawed, as pinpointed. This finishes the discussion of the proof of 2.1. Let us round this by noting a couple of very easy but noteworthy facts. Retracts of compact complex manifolds are compact complex submanifolds. Retracts of Stein manifolds are Stein and likewise, the property of a complex manifold being an Oka manifold, is preserved under passage to a retract thereof, as well.
Another general fact to keep in the background is the following; we include its proof, despite its obviousness, owing to its use for the rest of this article.
Proposition 2.3.
Suppose $X_{1}$ and $X_{2}$ are biholomorphically equivalent complex manifolds, with $F:X_{1}\to X_{2}$ being a biholomorphism. Then there is a one-to-one correspondence between retracts of $X_{1}$ and $X_{2}$, rendered via conjugation by $F$. In-particular, the image of a retract of a complex manifold $X$ under any member of ${\rm Aut}(X)$ is also a retract of $X$.
Proof.
Let $R_{1},~{}R_{2}$ be retracts of $X_{1},~{}X_{2}$ respectively; so, there exist holomorphic retraction maps $\rho_{1}:X_{1}\to X_{1}$ and $\rho_{2}:X_{2}\to X_{2}$ with $\rho_{1}(X_{1})=R_{1}$ and $\rho_{2}(X_{2})=R_{2}$. Let $Z_{1}:=F(R_{1}),~{}Z_{2}:=F^{-1}(R_{2})$. Consider the map $G:X_{2}\to X_{2}$ defined by $G(z)=F\circ\rho_{1}\circ F^{-1}(z)$. The idempotence of $\rho_{1}$ implies the same about $G$; so, $G$ is a holomorphic retraction on $X_{2}$ and ${\rm image}(G)=Z_{1}$. Thereby, $Z_{1}$ is a retract of $X_{1}$.
In other words, if $R_{1}$ is a retract of $X_{1}$ then $F(R_{1})$ is a retract of $X_{2}$. Similar arguments show that if $R_{2}$ is a retract of $X_{2}$ then $F^{-1}(R_{2})$ is a retract of $X_{1}$. ∎
As mentioned in the introduction, one may deduce out of the above elementary fact, the well-known characterization of retracts of $\Delta^{N}$, yet another way of demonstrating the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis, as soon as the dimension $N$ gets any bigger than one.
Corollary 2.4.
Let $D$ be any bounded balanced convex domain with $\mathbb{C}$-extremal boundary then $\Delta^{N}$ and $D$ are biholomophically inequivalent.
Proof.
If possible assume that $\Delta^{N}$ and $D$ are biholomorphically equivalent. Since automorphism group of $\Delta^{N}$ acts transitively, there exist a biholomorphism $F$ from $\Delta^{N}$ to $D$ with $F(0)=0$. By Cartan’s theorem (theorem 2.1.3 in [Rud_book_Fn_theory_unitball]), $F$ is a linear transformation. By theorem 1.2, every retract of $D$ passing through origin is a linear subspace of $D$. If $D_{L}:=D\cap L$ is a linear subspace of $D$, then $F^{-1}(D_{L})$ is also linear. By proposition 2.3, this leads to the conclusion that all retracts of $\Delta^{N}$ passing through origin are linear subspaces, which contradicts the fact that we have highlighted several times namely, the existence in plenty, of non-linear retracts passing through the origin for the polydisc. ∎
In passing, we may ofcourse note that by using the well-known theorem of Kaup–Upmeier [Kaup-Upm], we may replace the polydisc in the above by various other product domains which unlike polydiscs are neither convex nor homogeneous. Specifically, we may take a product $B_{1}\times\ldots\times B_{k}$ (with $k\geq 2$) of bounded balanced pseudoconvex domains with the property that one of the factors admits a non-linear holomorphic map into another, to ensure the existence of non-linear retracts, thereby the proof of the above corollary to go through. There are of-course better proofs and our reason for discussing this proof is to simultaneously make a few general remarks about non-linear retracts and about some simple ways of generating ample examples of retracts. We note that there are plenty of product domains of the kind just mentioned (satisfying the property about the existence of non-linear holomorphic maps, as mentioned above). To indicate a way of generating examples, let $E_{2m}:=\{(z,w)\;:\;|z|^{2m}+|w|^{2}<1\}$. For an arbitrary point $(z,w)\in E_{2m}$ note that $|z|<1$, thereby that, $|z^{2l}|+|w|^{2}<|z|^{2m}+|w|^{2}$ for every $l\geq m$. Consequently, $(z^{2l},w)$ lies in $E_{2l}$ for every $l\geq m$, implying that $(z,w)\mapsto(z^{2l},w)$, gives an example of a non-linear (in-fat proper) holomorphic mapping as desired and more importantly thereby, a non-linear retract passing through the origin in the polyball $E_{2m}\times E_{2l}$ for every $l\geq m$. Needless to say, similar constructions can be done with $l_{p}$-balls and more; we shall not digress further along this line.
2.1. Linear Retracts
Definition 2.5.
Let $D$ be any bounded balanced domain in $\mathbb{C}^{N}$ and $L$ a linear subspace of $\mathbb{C}^{N}$. We say that $D_{L}:=D\cap L$ is a linear retract of $D$, provided there exists an idempotent linear operator on $\mathbb{C}^{N}$ which maps $D$ onto $D_{L}$ (and thereby $\mathbb{C}^{N}$ onto $L$).
Remark 2.6.
A natural question arising for this definition, answering which will also settle some ambiguity that may otherwise arise in the sequel is the following. If it so happens that for a linear subspace $L$ of $\mathbb{C}^{N}$, $D_{L}:=D\cap L$ turns out to be a holomorphic retract i.e., the image of a possibly non-linear retraction $\rho$ of $D$, could it be said that $D_{L}$ is a linear retract in the sense of the above definition? We shall first answer this question under the assumption of convexity on $D$ as our domain of interest $D$ is convex; indeed then, the answer is in the affirmative and the reason is as follows. The image of $\rho$ being a linear subspace of $D$, namely $D_{L}$, it is in particular itself a bounded balanced convex domain in the Banach space $(L,\mu)$ where $\mu$ is the Minkowski functional of $D$ which is a norm on $\mathbb{C}^{N}$ by virtue of the convexity of $D$. This enables us to apply proposition C of Simha’s article [Simh] (exposited in theorem 8.1.2 of [Rud_book_Fn_theory_unitball] as well) which when applied to the pair of Banach spaces $(\mathbb{C}^{N},\mu)$, $(L,\mu)$ and the map $\rho$ between their unit balls, shows that the norm of the linear operator $D\rho(0)$ is atmost one. As there are points of norm one (with respect to the norm $\mu$) in $\overline{D}$ which are mapped to points of norm one in $L$ by $D\rho(0)$ namely, points of $\partial D_{L}$ itself, we immediately see that the norm of $D\rho(0)$ is exactly one. This leads to the desired conclusion that $D\rho(0)$ is a linear retraction of $D$ retracting $D$ onto $D_{L}$, as is brought out by the next general remark.
Remark 2.7.
Let $D$ be any bounded balanced convex domain in $\mathbb{C}^{N}$ and $\mu$ its Minkowski functional; as $D$ is convex, $\mu$ is a norm and $D$ is the unit ball with respect to $\mu$. Let $L$ be any complex linear subspace of $\mathbb{C}^{N}$ considered now as a Banach space with the norm $\mu$. Suppose there exists a linear projection $\pi_{L}:\mathbb{C}^{N}\to L$ of norm one, where we equip $\mathbb{C}^{N}$ and $L$ both with the norm $\mu$. This means that $\pi_{L}$ maps $D$ onto $D_{L}:=L\cap D$, the unit ball in $L$ with respect to the same norm $\mu$ restricted to $L$, which is fixed pointwise by $\pi_{L}$ (after all, $\pi_{L}$ itself is a retraction of $\mathbb{C}^{N}$ onto $L$); thus, $\pi_{L}$ is a linear retraction of $D$ onto $D_{L}$. Thus, a linear subspace of $D$ which must be of the form $D_{L}=D\cap L$ for some linear subspace $L$ of $\mathbb{C}^{N}$, is a linear retract of $D$ if and only if there exists a linear projection of the Banach space $(\mathbb{C}^{N},\mu)$ of norm one, mapping $D$ onto $D_{L}$. Finally, let us just remark here once for clarity that while this projection $\pi_{L}$ is linear, it may well not be (and most often not) an orthogonal projection.
Let us now address the basic question of generating (examples of) new retracts out of old. To begin with a first rudimentary step towards this, we put down the following transitivity result.
Lemma 2.8.
Let $X$ be any complex manifold. If $Z$ is a retract of $X$ and $W$ is a retract of $Z$ then $W$ is a retract of $X$.
Proof.
Since $Z$ is a retract of $X$ and $W$ is a retract of $Z$, there exist a pair of retraction maps $\rho_{1}:X\to Z$ and $\rho_{2}:Z\to W$. Note that if $x\in Z$ then $\rho_{1}\circ\rho_{2}(x)=\rho_{2}(x)$. Note that $\rho_{2}\circ\rho_{1}$ is a map from $X$ to $W$, and also that
$\rho_{2}\circ\rho_{1}\circ\rho_{2}\circ\rho_{1}(x)=\rho_{2}\circ\left(\rho_{1}\circ\rho_{2}(\rho_{1}(x))\right)=\rho_{2}\circ\rho_{2}(\rho_{1}(x))=\rho_{2}\circ\rho_{1}(x).$
So, $\rho_{2}\circ\rho_{1}$ is a retraction map on $X$ verifying that $W$ is a retract of $X$. ∎
While the above elementary lemma provides a manner of obtaining retracts of lower dimension, we may ask for a result with the converse effect. This will be addressed in the forthcoming sections at its appropriate place. Let us only mention here the answer to a special case of this question in the realm of balanced domains and linear retracts, namely the following. Suppose $D_{L_{1}}$ and $D_{L_{2}}$ are linear retracts of a bounded balanced convex domain $D$; then, is it true that $D_{L_{1}\oplus L_{2}}$ a linear retract of $D$? The following proposition shows that except essentially for the Euclidean ball, it can never happen that for every pair of linear retracts such a sum of (linear) retracts is again a retract.
Proposition 2.9.
Let $D$ be a bounded balanced convex domain in $\mathbb{C}^{N}$. If $D$ is not linearly isometric to Euclidean ball $\mathbb{B}$, then there exist one-dimensional linear subspaces $L_{1},L_{2}$ such that $D_{L_{1}\oplus L_{2}}$ is not a linear retract of $D$.
Proof.
[…]
For $D$ as in the above theorem, we have in background the observation that the convexity of $D$, implies by the Hahn-Banach theorem that every one dimensional subspace of $D$ is a retract of $D$. One may want to ascertain the converse, which is addressed in the following.
Proposition 2.10.
Suppose every one-dimensional linear subspace of a bounded balanced domain $D$ is indeed a retract, then domain $D$ must be convex.
Proof.
Assume to get a contradiction that there exists a pair of distinct points $p,q\in\partial D$, with the property that their mid-point, fails to lie within the closure of the domain i.e., $m:=(p+q)/2\;\not\in\overline{D}$. Let $T$ a linear retraction map (i.e., linear projection on $\mathbb{C}^{N}$) which maps $D$ onto $L_{m}$, the one-dimensional linear subspace spanned by this (mid-)point $m$ – the existence of such a map is guaranteed by hypothesis. Then, note first of all that: $T$ fixes $L_{m}$ pointwise; so $T(m)=m$, thereby $T(m)$ lies outside $\overline{D}$. On the other hand, $T(p)$ and $T(q)$ are a pair of points on $L_{m}\cap\overline{D}$, as $T$ maps $\overline{D}$ into $\overline{D}\cap L$. And, the mid-point of these image-points must lie in $L_{m}\cap\overline{D}$: by balancedness of $\overline{D}$, the line-segment joining $T(p)$ and $T(q)$ has got to be contained within $\overline{D}$ – note that this does not require convexity of $\overline{D}$! But then we have
$\frac{T(p)+T(q)}{2}\;=\;T\left(\frac{p+q}{2}\right)\;=\;T(m)$
It follows that $T(m)$ lies within $\overline{D}$. However, by earlier observations leading to $T(m)=m$, we have that the same point $T(m)$ lies outside $\overline{D}$! This absurdity only means that the assumption that $D$ is convex cannot hold, finishing the proof. ∎
2.2. Notions of extremality and strict convexity
Definition 2.11.
Let $K$ be the field $\mathbb{R}$ or $\mathbb{C}$. Let $\Omega\subset\mathbb{C}^{N}$ be an arbitrary domain and $a\in\partial{\Omega}$. We introduce a notion that captures the set of all directions along which one can move from $a$, at-least a little bit while staying within the boundary $\partial\Omega$. Towards this, define the face of $\partial\Omega$ at $a$ by
$F_{K}(a,\partial{\Omega})=\{u\in\mathbb{C}^{N}:\exists\epsilon>0,a+tu\in\partial{\Omega},\text{ for all }t\in K~{}with~{}|t|<\epsilon\}$
reduces to $\{0\}$. It is immediate that $F_{K}$ is closed under scalar multiplication. But a stronger claim that $F_{K}$ is a linear subspace of $\mathbb{C}^{N}$ requires more assumptions on $\Omega$. Indeed, if $\Omega$ is convex, then $F_{K}(a,\partial{\Omega})$ is $K$-linear subspace. A boundary point $a\in\partial{\Omega}$ is said to be a $K$-extremal boundary point for $\Omega$, if $F_{K}(a,\partial{\Omega})=0$; if the face reduces to $\{0\}$ for every $a\in\partial{\Omega}$, then we say that the domain $\Omega$ is $K$-extremal.
We would now like to make a couple of clarifications about the case $K=\mathbb{C}$ versus the case $K=\mathbb{R}$. To begin with, note that we obviously have
$\vec{F}_{\mathbb{C}}(a,X)=\vec{F}_{\mathbb{R}}(a,X)\cap i\vec{F}_{\mathbb{R}}(a,X).$
It follows that if $\Omega$ is $\mathbb{R}$-extremal, then $\Omega$ is $\mathbb{C}$-extremal. But the converse is not true. To see this, consider $B_{\ell^{1}}:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<1\}$; we shall presently proceed to showing that this is $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal. Perhaps even more basically, we must remark that every domain in $\mathbb{C}$ is $\mathbb{C}$-extremal. But only strict convex domains are $\mathbb{R}$- extremal.
Lemma 2.12.
Let $\rho:\Delta\to\overline{B_{\ell^{1}}}$ be a holomorphic map. If image of $\rho$ is contained in $\partial B_{\ell^{1}}$ then $\rho$ is a constant map.
Proof.
Write $\rho(z)=(\rho_{1}(z),\rho_{2}(z))$. Note that $\rho_{1},\rho_{2}$ are holomorphic maps from $\Delta$ to $\overline{\Delta}$. Write in polar-coordinates: $\rho_{1}(0)=r_{1}e^{i\theta_{1}}$ and $\rho_{2}(0)=r_{2}e^{i\theta_{2}}$. Define $g:\Delta\to\overline{\Delta}$ by
$g(z)=\rho_{1}(z)e^{-i\theta_{1}}+\rho_{2}(z)e^{-i\theta_{2}}.$
Note that $g$ is holomorphic and $g(0)=1$. By maximum modulus principle, $g$ is constant, indeed $g\equiv 1$. This when written out in terms of the $\rho_{j}^{\prime}s$ reads
$\rho_{1}(z)e^{-i\theta_{1}}+\rho_{2}(z)e^{-i\theta_{2}}=|\rho_{1}(z)e^{-i\theta_{1}}|+|\rho_{2}(z)e^{-i\theta_{2}}|,\text{ both being equal to }1.$
This in-particular, means that equality holds in triangle inequality. This implies that $\rho_{1}(z)e^{-i\theta_{1}}=c\rho_{2}(z)e^{-i\theta_{2}}$, for some $c\in\mathbb{C}$. Hence, $|\rho_{1}(z)|=k|\rho_{2}(z)|$, where $k=|c|\geq 0$. As $|\rho_{1}(z)|+|\rho_{2}(z)|=1$, this leads to
$k|\rho_{2}(z)|+|\rho_{2}(z)|=1$
This implies that $|\rho_{2}(z)|=1/(1+k).$ So $\rho_{2}$ is constant by maximum modulus principle, thereby $\rho_{1}$ is constant as well, finishing the proof. ∎
Proposition 2.13.
$B_{\ell^{1}}:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<1\}$ is $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal.
Proof.
We first show that $B_{\ell^{1}}$ is $\mathbb{C}$-extremal. To do this, let $a\in\partial B_{\ell^{1}}$. Suppose there exists $\epsilon_{0}>0,a\in\partial B_{\ell^{1}}$ and $u_{0}\in\mathbb{C}^{N}$ such that $a+tu_{0}\in\partial B_{\ell^{1}}$, for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$.
Without loss of generality, we may assume after an appropriate scaling that $\epsilon_{0}=1$; this ensures that the affine-linear (holomorphic) map defined by $\rho(t)=a+tu_{0}$ maps $\Delta$ into $\partial B_{\ell^{1}}$. By the above lemma, $\rho$ is constant, which is equal to $a$. This implies $u_{0}=0$, finishing the proof of $\mathbb{C}$-extremality.
We next show that $B_{\ell^{1}}$ is not $\mathbb{R}$-extremal. To do this, first let $a=(1/2,1/2)\in\partial B_{\ell^{1}}$ and consider the face
$F_{\mathbb{R}}(a,\partial B_{\ell^{1}})=\{u\in\mathbb{C}^{2}:\exists\epsilon>0,~{}a+tu\in\partial B_{\ell^{1}}~{},\forall t\in\mathbb{R}~{}\text{with}~{}|t|<\epsilon\}.$
Note then that $(1,-1)\in F_{\mathbb{R}}(a,\partial B_{\ell^{1}})$. Hence $B_{\ell^{1}}$ is not $\mathbb{R}$-extremal. ∎
Definition 2.14.
[…]
First we prove that the notion of rotundity and strictly convex are equivalent in a bounded balanced convex domain $\Omega$.
Theorem 2.15.
Let $\Omega$ be a bounded balanced convex domain in $\mathbb{C}^{N}$. Then $\Omega$ is rotund if and only if $\Omega$ is $\mathbb{R}$-extremal.
Proof.
Assume $\Omega$ is rotund. Let $a\in\partial\Omega$ and $(a+tu)\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. Note that $\frac{1}{2}(a+(a+tu))=a+\frac{t}{2}u\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. By proposition 5.1.2 of [Meggn], $a=a+tu$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$ and this implies that $u=0$. Therefore, $\Omega$ is $\mathbb{R}$extremal.
Conversely assume that $\Omega$ is $\mathbb{R}$extremal. Let $z_{1},z_{2}\in\partial\Omega$.
$(1-t)z_{1}+tz_{2}=z_{1}+t(z_{2}-z_{1})\in\partial\Omega\quad\text{for ~{}all}\ t\in[0,1].$
$\frac{z_{1}+z_{2}}{2}+t(z_{2}-z_{1})=(1/2-t)z_{1}+(1/2+t)z_{2}\in\partial{\Omega}~{}\text{for ~{}all}\ t\ \text{with}\ |t|<1/2$
Therefore $z_{2}-z_{1}\in F_{\mathbb{R}}(\frac{z_{1}+z_{2}}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$- extremal, $z_{2}=z_{1}$ and therefore $\Omega$ is rotund.
As the relation between the pair of notions gotten by taking $K=\mathbb{C}$ versus $K=\mathbb{R}$ will be basic importance in the sequel, we discuss this below. But one would perhaps like to know first the relation between strict convexity and extremality introduced above; as we do not find a quick reference for this, we include the details for this here.
Theorem 2.16.
Let $\Omega$ be a bounded balanced convex domain in $\mathbb{C}^{N}$. Then $\Omega$ is strictly $\mathbb{R}$-convex if and only if $\Omega$ is $\mathbb{R}$-extremal.
Proof.
Suppose that $\Omega$ is strictly $\mathbb{R}$-convex. If possible, assume that $\Omega$ is not $\mathbb{R}$- extremal. Then by theorem 2.15, $\Omega$ is not rotund. By corollary 5.1.16 of [Meggn], there exists $\mathbb{R}$- linear functional $f$ supports more than one point. This implies that there exists $z_{0},w_{0}\in\mathbb{C}^{N}$ such that $f(z_{0})=f(w_{0})=\|f\|$. Take $\alpha:=\|f\|$. Define a $\mathbb{R}$- linear functional $g$ defined by $g(z)=\frac{1}{\alpha}f(z)$. Note that $\|g|\ =1$ and $g(z_{0})=g(w_{0})=1$. This contradicts the strict $\mathbb{R}$-convexity of $\Omega$. Therefore, $\Omega$ is $\mathbb{R}$- extremal.
Conversely assume that $\Omega$ is $\mathbb{R}$- extremal. Let $T$ be a $\mathbb{R}-$ linear map with $\|T\|=1$. Assume that there exists $z,w\in\overline{\Omega}$ such that $Tz=1=Tw$. Then $z,w\in\partial{\Omega}$, otherwise $T\left(\frac{z}{\|z\|}\right)=\frac{1}{\|z\|}>1$. Note that $T((1-t)z+tw)=1$. Then $(1-t)z+tw\in\partial{\Omega}$ for all $t\in[0,1]$. In particular, if $t=1/2$, we get $\frac{1}{2}(z+w)\in\partial{\Omega}$. Note then that
$\frac{z+w}{2}+t(w-z)=(1/2-t)z+(1/2+t)w\in\partial{\Omega}~{}\ \text{for all t}\in\mathbb{R}\ \text{with}\ |t|<1/2.$
Then $w-z\in F_{\mathbb{R}}(\frac{z+w}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$-extremal, then $w=z$. Therefore $\Omega$ is strictly $\mathbb{R}$-convex. ∎
Theorem 2.17.
Let $\Omega$ be a bounded balanced convex domain in $\mathbb{C}^{N}$. Then $\Omega$ is strictly $\mathbb{R}$-convex if and only if $\Omega$ is strictly $\mathbb{C}$-convex.
Proof.
[…]
Conversely, assume that $\Omega$ is strictly $\mathbb{C}$-convex. We have to prove that $\Omega$ is strictly $\mathbb{R}$-convex. By above theorem, it suffices to show that $\Omega$ is $\mathbb{R}$-extremal. If possible assume that there exists $u\in\mathbb{C}^{N},~{}a\in\partial\Omega$ and an $\epsilon>0$ such that $a+tu\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. Let $M=span_{\mathbb{R}}\{u\}$.
Claim-1: $a\notin M$
Proof of Claim-1: If possible assume that $a\in M$. Choose $s_{0},~{}t_{0}\in\mathbb{R}$ such that $|s_{0}|<\epsilon$ and $|s_{0}t_{0}|<\epsilon$. Since $a\in M$, we get $a+s_{0}u\in M$. Hence we get $a+s_{0}u=ca$ for some $c\in\mathbb{R}$. Taking norm on both sides we get $|c|=1$.
$\|a+s_{0}t_{0}u\|^{2}=\|(1-t_{0})a+t_{0}(a+s_{0}u)\|^{2}=(1-t_{0})^{2}+2t_{0}(1-t_{0})c+c^{2}t_{0}^{2}<(1-t_{0})^{2}+2t_{0}(1-t_{0})+t_{0}^{2}=(1-t_{0}+t_{0})^{2}=1$.
Hence we get $\|(1-t_{0})a+t_{0}(a+s_{0}u)\|^{2}<1$. This implies that $\|a+s_{0}t_{0}u\|<1$. Hence, we get $a+s_{0}t_{0}u\notin\partial\Omega$. This contradicts our assumption, proving the claim.
By Hahn–Banach theorem, there exists a $\mathbb{C}$- linear functional $f$ with $\|f\|=1$ such that $f(z)=0$ for all $z\in M$ and $f(a)=dist(a,M)$.
Claim-2: ${\rm dist}(a,M)=1$.
Proof of Claim-2: Let us begin by recalling that $\rm dist(a,M)=\inf\{\|a+su\|:s\in\mathbb{R}\}$ and note firstly, that ${\rm dist}(a,M)\leq 1$. Take an arbitrary $s\in\mathbb{R}$. Choose $t_{0}\in[0,1)$ such that $|(1-t_{0})s|<\epsilon$ and note that
$\|(1-t_{0})(a+su)+t_{0}a\|\leq(1-t_{0})\|a+su\|+t_{0}\|a\|=(1-t_{0})\|a+su\|+t_{0}.$
We simplify the above inequality to write it as
(2.1) $\|a+(1-t_{0})su\|\leq(1-t_{0})\|a+su\|+t_{0}.$
Since $|(1-t_{0})s|<\epsilon$, we get from (2.1) that $\|a+(1-t_{0})su\|=1$. This implies $\|a+su\|\geq 1$ for all $s\in\mathbb{R}$. Hence, $dist(a,M)=1$, finishing the proof of claim-2.
Note that $f(a+tu)=f(a)=1$. Since $\Omega$ is strictly $\mathbb{C}$-convex, we have $a=a+tu$. This implies that $u=0$. So $\Omega$ is $\mathbb{R}$-convex. ∎
Remark 2.18.
$\mathbb{C}$-extremal domains need not be strictly $\mathbb{C}$-convex. To give the simplest example, we consider again the $\ell^{1}$ ball $B_{\ell^{1}}$. Consider $T:\mathbb{C}^{2}\longrightarrow\mathbb{C}$ defined by $T(z,w)=z+w$. We note that $T$ is linear and $\|T\|_{1}=1$, but $T(1,0)=T(0,1)=1$. Thus $B_{\ell^{1}}$ is not strictly $\mathbb{C}$-convex, but is $\mathbb{C}$-extremal.
We conclude with one other notion of extremality, whose definition is more restrictive than that of $\mathbb{C}$-extremality.
Definition 2.19.
Let $D\subset\mathbb{C}^{N}$ be a bounded domain. We say that a point $a\in\partial D$ is holomorphically extreme if there is no non-constant holomorphic mapping $\phi:\Delta\rightarrow\overline{D}$ such that $\phi(0)=a$.
The purpose of recalling this notion in this article about retracts is the following result; being a result about retracts, we prefer to include the details of the proof.
Theorem 2.20.
(A slight modification of proposition $4.5$ in [JJ]). Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a non-trivial retract through the origin with the property that the boundary points of $D_{L}:=D\cap T_{0}Z$ viewed as a domain in $T_{0}Z$ are holomorphically extreme. Then, $Z$ is linear.
Proof.
Let $\mathfrak{h}_{D_{L}}$ denote the Minkowski function on $D_{L}$. Since $Z$ is a retract of $D$, there exists a holomorphic retraction map $\rho:D\to D$ such that $\rho(D)=Z$. Note that $\rho|_{D_{L}}$ is a holomorphic map from $D_{L}$ to $D$ and $Z=image(\rho|_{D_{L}})$. Observe that $Z$ is a connected complex submanifold and complex tangent space at 00 is $T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}|_{D_{L}}(0)v=v\}$. Let $v\in D_{L},v\neq 0$. By our assumption on $\partial D_{L}$, it follows that the mapping
$\phi:\Delta\ni\lambda\longmapsto\lambda v/\mathfrak{h}_{D_{L}}(v)\in D_{L}.$
is the unique (modulo $Aut(\Delta)$) $K_{D_{L}}$ geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$. Since $D_{L}\subset D$ and $L$ is a linear subspace of $\mathbb{C}^{N}$, $\mathfrak{h}_{D_{L}}(v)=\mathfrak{h}_{D_{L}}(v)$. So $\phi$ is the unique (modulo $Aut(\Delta)$) $K_{D}$ geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$. Note that
$K_{D}(\rho|_{D_{L}}\circ\phi(0);(\rho|_{D_{L}}\circ\phi)^{\prime}(0))=K_{D}(0,v/\mathfrak{h}_{D_{L}}(v))=K_{D}(\phi(0);\phi^{\prime}(0)).$
Hence $\rho|_{D_{L}}\circ\phi$ is a geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$. By uniqueness, it follows that $\rho|_{D_{L}}\circ\phi=\phi\circ\psi$, where $\psi\in Aut(\Delta)$. Substituting $\psi^{-1}(\mathfrak{h}_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho|_{D_{L}}$. Hence $D_{L}\subset Z$. Now $T_{0}Z$ and $Z$ have the same dimension. Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$. Hence $Z$ is linear. ∎
2.3. Geodesics and totally geodesic subsets
We recall the notion of totally geodesic subsets of a general domain (or a complex manifold) $D$ from [Kosnski]. Let $p_{1},p_{2}$ be (distinct) points in $D$; the pair $\delta=(p_{1},p_{2})$ shall be referred to as a ‘datum’. A Kobayashi extremal for $\delta$ is a holomorphic map $k:\Delta\to D$ such that both $p_{1}$ and $p_{2}$ are in the range of $k$, and the pseudo-hyperbolic distance between the preimages of $p_{1}$ and $p_{2}$ is minimized at $k$, over all holomorphic maps from $\Delta$ to $D$. We shall call the range of a Kobayashi extremal with datum $\delta$, a geodesic through $\delta$. If a set $V\subset D$ has the property that for any $p_{1},p_{2}\in V$, a geodesic through $(p_{1},p_{2})$ is contained in $V$, we shall say that $V$ is totally geodesic.
The simple fact that the Kobayashi distance $k_{D}(p,q)$ between a pair of points $p,q$ in a domain $D$ remains invariant (i.e., $k_{D}(p,q)=k_{Z}(p,q)$) under passage to a retract $Z\subset D$, was recorded with proof explicitly as lemma $1$ in [HanPeters-Zeager]. As the proof is very simple and uses only the contractive feature of the Kobayashi distance under holomorphic maps, we therefore only note here that the same holds much more generally for any contractive metric or distance on any complex manifold.
Lemma 2.21.
Let $X$ be any complex manifold and $Z\subset X$, a retract. For a contractive pseudodistance $d$ and contractive (infinitesimal) metric $\delta$, we have
$d_{Z}(p,q)=d_{X}(p,q),\;\;\text{ and }\;\;\delta_{Z}(p,v)=\delta_{X}(p,v)$
for every pair of points $p,q\in Z$ and $v\in T_{p}X$.
It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic, a notion that shall be made use of, while recalling the work [Mok] of Mok.
2.4. One dimensional Retracts
There are plenty of one dimensional retracts which can be obtained from Caratheodory complex geodesics (referred to henceforth as $C$-geodesic) whose definition we now recall.
Definition 2.22.
We say that a holomorphic map $\varphi:\Delta\rightarrow D$ is a complex $C$-geodesic of $D$ if for all $\zeta_{1},\zeta_{2}\in\Delta$, we have $c_{D}\left(\varphi\left(\zeta_{1}\right),\varphi\left(\zeta_{2}\right)\right)=p\left(\zeta_{1},\zeta_{2}\right)$.
Similarly we can define complex $K$-geodesic (or complex Kobayashi geodesic) if we replace $c_{D}$ by $K_{D}$ in the above definition. By Lemma 2.6.1 (ii) i [Abate_Itrtn], every complex $K$-geodesic is a complex $C$-geodesic. But the converse is not true because there are domains in $\mathbb{C}^{N}$ which are Kobayashi hyperbolic but not Caratheodory hyperbolic. Owing to the direct relationship between $C$-geodesics and retracts, we shall mostly work with $C$-geodesics rather than $K$-geodesics.
E. Vesentini [Vesentini_1], gives the following characterization of the complex geodesics of $D$.
Proposition 2.23.
Let $\varphi:\Delta\rightarrow D$ be a holomorphic map. For $\varphi$ to be a complex geodesic of $D$, it is necessary and sufficient that $\varphi$ satisfies one of the following conditions:
  1. 1.
    there are two distinct points $\zeta_{1}$ and $\zeta_{2}$ in $\Delta$ such that $c_{D}\left(\varphi\left(\zeta_{1}\right),\varphi\left(\zeta_{2}\right)\right)=c_{\Delta}\left(\zeta_{1},\zeta_{2}\right)$.
  2. 2.
    there exists $\zeta_{1}\in\Delta$ and a nonzero vector $v\in\mathbb{C}$ such that $\gamma_{D}\left(\varphi\left(\zeta_{1}\right),\varphi^{\prime}\left(\zeta_{1}\right)\cdot v\right)=\gamma_{\Delta}\left(\zeta_{1},v\right)$.
The following proposition relates complex geodesics with retracts.
Proposition 2.24.
Let $D$ be a domain in $\mathbb{C}^{N}$. If $\varphi:\Delta\to D$ is a complex $C$-geodesic, then $\varphi(\Delta)$ is a one dimensional retract of $D$.
Proof.
First we want to prove that $\varphi$ has a holomorphic left inverse. For all $\zeta_{1},\zeta_{2}\in\Delta$, we have
$c_{D}\left(\varphi\left(\zeta_{1}\right),\varphi\left(\zeta_{2}\right)\right)=p(\zeta_{1},\zeta_{2})$
By definition of $c_{D}$, for all $\zeta_{1},\zeta_{2}\in\Delta$ we have
$\sup_{f\in\mathcal{O}(D,\Delta)}p\left((f\circ\varphi)(\zeta_{1}),(f\circ\varphi)(\zeta_{2})\right)=p(\zeta_{1},\zeta_{2}).$
There exists $f_{n}\in\mathcal{O}(D,\Delta)$ such that
$p\left((f_{n}\circ\varphi)(\zeta_{1}),(f_{n}\circ\varphi)(\zeta_{2})\right)\to p(\zeta_{1},\zeta_{2})\quad\text{as}~{}n\to\infty.$
By Montel’s theorem, there exists a subsequence $f_{n_{k}}$ converging to $g$ uniformly on compact subsets of $D$. So $p\left((g\circ\varphi)(\zeta_{1}),(g\circ\varphi)(\zeta_{2})\right)=p(\zeta_{1},\zeta_{2})$ for all $\zeta_{1},\zeta_{2}\in\Delta$. By Schwarz Pick lemma, we have $g\circ\varphi\in Aut(\Delta)$. This implies that $\varphi$ has a holomorphic left inverse.
Let $\psi\in\mathcal{O}(D,\Delta)$ such that $\psi\circ\varphi=id_{\Delta}$. If we set $\rho=\varphi\circ\psi$, then $\rho$ is a holomorphic self map on $D$ and
$\rho\circ\rho=(\varphi\circ\psi)\circ(\varphi\circ\psi)=\varphi\circ(\psi\circ\varphi)\circ\psi=\varphi\circ\psi=\rho.$
This proves that $\rho$ is a holomorphic retraction map on $D$ and $\varphi(\Delta)$ is a one dimensional retract of $D$. ∎
Theorem 2.25.
If $D$ is a simply connected Kobayashi hyperbolic domain in $\mathbb{C}^{N}$, then $1$-dimensional retracts of $D$ are precisely complex $C$-geodesics.
Proof.
By proposition 2.24, we get that every complex $C$-geodesic is a retract. We have only to prove that every one-dimensional retract is a complex $C$-geodesic. Let $Z$ be a $1$-dimensional retract of $D$. Note that $Z$ is a non-compact simply connected Riemann surface. By the uniformization theorem, $Z$ is biholomorphic to one of the Riemann surfaces: unit disk $\Delta,~{}\hat{\mathbb{C}},~{}\mathbb{C}$. Since $Z$ is non-compact and hyperbolic, $\hat{\mathbb{C}}$ and $\mathbb{C}$ are ruled out. Hence $Z$ is biholomorphic to $\Delta$. Let $\phi$ be a biholomorphic map from $\Delta$ to $Z$. By proposition 1.2 (the equivalence of parts a) and d) therein ) from [Dineen], we obtain that $\phi$ is a complex geodesic. So $Z$ is an image of complex $C$-geodesic. ∎
From proposition 2.10, we can conclude that not all linear subspaces of bounded balanced domain $D$ are a retract of $D$ unless $D$ is convex. One natural question arises: which linear subspaces of bounded balanced domain $D$ does not arises as a retract of $D$. The following corollary gives more general result. Before going to the statement, first we recall the notion of convexity at a boundary point.
Definition 2.26.
Let $D=\{z\in\mathbb{C}^{N}:\mathfrak{h}(z)<1\}$ be a balanced domain in $\mathbb{C}^{N}$ ($\mathfrak{h}$ is the Minkowski function on $D$) and $a\in D$. $D$ is said to be convex at $\frac{a}{\mathfrak{h}(a)}$ if there exists a $\mathbb{C}$- linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ with $|L|\leq\mathfrak{h}$ and $|L(a)|=\mathfrak{h}(a)$.
Corollary 2.27.
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that D is not convex at $a/\mathfrak{h}(a)$. Then there does not exists one dimensional retract of $D$ passing through 00 and $a$.
Proof.
Assume the contrary, there exists a one dimensional retract $R$ of $D$ passing through 00 and $a$. By theorem 2.25, $R$ is image of some complex C-geodesic $\varphi:\Delta\to D$. But by lemma 11.1.2 in [Jrncki_invrnt_dst] (see also remark 11.1.3(b) therein), there is no complex $C$- geodesic in $D$. This leads to a contradiction. ∎
Combining theorem 2.25 and theorem 3.2 in [Vigue_Cnvx], we obtain the following result:
Theorem 2.28.
Let $D$ be a bounded convex domain of $\mathbb{C}^{N}$,
  1. (1)
    Given two points $x$ and $y$ of $D$, there exists one dimensional retract $R$ of $D$ passing through $x$ and $y$.
  2. (2)
    Given $x\in D$ and $v\in T_{x}(D)$, there exists one dimensional retract $R$ of $D$ passing through $x$ with tangent vector $v$.
2.5. Fixed point sets and Retracts
As we already saw in the introduction, fixed point sets need not be realizable as retracts in bounded non-convex domains. However, there is still a connection between fixed point sets and retracts under mild hypotheses, which we formulate in the following proposition:
Proposition 2.29.
[…]
Proof.
Consider the sequence of holomorphic maps on $D$ given by $f_{n}:=f^{\circ n}$ which denotes as usual the $n$-fold composition of $f$ with itself. Since $\text{Fix}(f)\neq\phi$, $(f_{n})$ is not compactly divergent. By theorem 2.1.29 in [Abate_Itrtn], there exists a retract $Z$ of $D$ and a holomorphic retraction $\rho:D\to Z$ such that $\rho$ is a limit point of the sequence $\{f_{n}\}$. Note that $\rm Fix(f)\subset\rm Z$. $\rm dim(Z)\neq 0$, because $Z$ is a connected complex manifold and #$\#(Fix(f))>1$. By Corollary 2.1.30 in [Abate_Itrtn], $\rm dim(Z)$ is equal to the number of eigenvalues of $df_{z_{0}}$ contained in $\partial\Delta$. If all eigenvalues of $df_{z_{0}}$ are in $\partial\Delta$, then by theorem 2.1.21 (v) in [Abate_Itrtn], we get $f$ is an automorphism. This contradicts the hypothesis. Hence we can conclude that $\rm dim(Z)<N$. This proves that $Z$ is a non-trivial retract of $D$ and $\rm Fix(f)\subset Z$. ∎
3. Proofs of theorems 1.3, 1.8, 1.10, 1.12
We would like to admit at the outset that the proof of theorem 1.12 on retracts of polyballs, closely follows the recent work [BBMV-union]; but we give here a thorough treatment even if, at the cost of repetition of essential ideas, for it allows for a better exposition of the proof of both this theorem, as well as perhaps more importantly: proofs of theorems prior to it as in the title of this section, which are fundamental to this article.
Our goal here is to characterize the holomorphic retracts of the Cartesian product $B_{1}\times\ldots\times B_{N}$ (which will be referred to as polyballs for short), where the $B_{j}$’s are bounded balanced strictly convex domains or equivalently, the unit ball with respect to some norm in $\mathbb{C}^{n_{j}}$ for some $n_{j}\in\mathbb{N}$, for each $j=1,\ldots,N$; as noted in the previous section, this is equivalent to their boundaries being $\mathbb{R}$-extremal. As it turns out that the essential arguments for the case of a product of a pair of such balls extend to the more general case, we shall for simplicity of exposition, focus attention for the most part on the case of a product of $2$ such balls. So, let $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ (where $N_{1},N_{2}\in\mathbb{N}$ with say, $\min\{N_{1},N_{2}\}=N_{1}$) denote a pair of bounded balanced convex domains with $\mathbb{R}$-extremal boundaries; in-particular, $B_{1},B_{2}$ are balls with respect to some norms (on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively) and needless to say are domains of holomorphy. The simplest example of such a domain is the standard unit disc in $\mathbb{C}$ which we denote by $\Delta$. However, many nice properties of the disc may well not hold for such domains (i.e., bounded balanced convex domains). To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where []delimited-[][$\cdot$][ ⋅ ] denotes the greatest integer function. They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extermality, in the following.
Proposition 3.1.
If $D_{1},D_{2}$ are domains in $\mathbb{C}^{N}$ with $\mathbb{C}$-extremal boundary, then $D:=D_{1}\cap D_{2}$ is also a domain with $\mathbb{C}$-extremal boundary.
Proof.
Let $a$ be an arbitrary point in $\partial D$. We want to prove that $a$ is $\mathbb{C}$-extremal point in $D$. If possible assume that $a$ is not $\mathbb{C}$-extremal in $D$. That is, there exists non-zero $u\in\mathbb{C}^{N}$ and there exists $\epsilon>0$ such that $a+tu\in\partial D~{}\text{for~{} all }~{}t\in\mathbb{C}~{}\text{with}~{}|t|<\epsilon$. In particular, $a\in\partial D$ and we shall split the proof into cases depending upon which ‘piece’ of the boundary $a$ lies; herein, piece refers to one of the components in the decomposition of the boundary of $D$ as
$\partial{(D_{1}\cap D_{2})}=(\partial D_{1}\cap D_{2})\cup(D_{1}\cap\partial D_{2})\cup(\partial D_{1}\cap\partial D_{2})$
Without loss of generality assume that $\epsilon=1$. Consider a map $\alpha:\Delta\to\mathbb{C}$ defined by $\alpha(t)=a+tu$. Since $u\neq 0$, $\alpha$ is a non-constant holomorphic map. Note that $\alpha$ is a homeomorphism from $\Delta$ onto $\alpha(\Delta)$. Consider the sets $S_{1}:=\{t\in\Delta:\alpha(t)\in\partial{D_{1}}\}$ and $S_{2}:=\{t\in\Delta:\alpha(t)\in\partial{D_{2}}\}$. By above decomposition, $S_{1}\cup S_{2}=\Delta$ and note that $S_{1}=\alpha^{-1}(\partial D_{1})$ and $S_{2}=\alpha^{-1}(\partial D_{2})$. Again by continuity of $\alpha$, $S_{1}$ and $S_{2}$ are closed sets in $\Delta$.
We want to prove that $S_{1}$ or $S_{2}$ has atleast one interior point. Suppose $S_{1}$ has no interior points in $\Delta$. Let $\rm int(S_{i})$ denotes the set of all interior points of $S_{i}$ for $i=1,2$. Note that
$S_{2}=\overline{S_{2}}=\overline{\Delta\setminus S_{1}}=\Delta\setminus int(S_{1})=\Delta$
So we get $S_{2}=\Delta$. Hence we can conclude that $S_{1}$ or $S_{2}$ has atleast one interior point.
Without loss of generality assume that $S_{1}$ has an interior point. Let $t_{0}\in\Delta$ be an interior point of $S_{1}$. This means that there exists $\delta>0$ such that $A:=\{t\in\mathbb{C}:|t-t_{0}|<\delta\}\subset S_{1}$. Choose any $s\in\mathbb{C}$ with $|s|<\delta$. Let $b:=a+t_{0}u$. Note that $b\in\partial D_{1}$, because $t_{0}\in S_{1}$. We can write $b+su=a+(t_{0}+s)u$ and observe that $t_{0}+s\in A$. Hence $t_{0}+s\in S_{1}$ for all $s\in\mathbb{C}$ with $|s|<\delta$. By definition of $S_{1}$, we obtain that $b+su\in\partial D_{1}$ for all $s\in\mathbb{C}$ with $|s|<\delta$. Since $D_{1}$ is $\mathbb{C}$-extremal, we have $u=0$. This contradicts our assumption that $a$ is not a $\mathbb{C}$-extremal point of $D$. Hence we conclude that $a$ is $\mathbb{C}$-extremal. ∎
Thus an intersection of finitely many balls with $K$-extremal boundaries (for either case $K=\mathbb{R}$ or $\mathbb{C}$) is again a convex domain with $K$-extremal boundary. In-particular, for instance, the intersection of the $l^{1}$-ball with a suitable dilate of another $l^{p}$-ball for $p\in(1,\infty)$ can be used to get hold of a concrete instance of a domain with $\mathbb{C}$-extremal (but not smooth) boundary. As an example which is not derived out of balls with respect to the isotropic $l^{p}$-norms, we may consider the ‘anisotropic’ ball (also called a complex ellipsoid) :=$\mathcal{E}:=\{(z,w)\in\mathbb{C}^{2}\;:\;|z|^{2}+|w|^{4}<1\}$. So, there is a good variety of bounded balanced convex domains with $\mathbb{R}$-extremal and $\mathbb{C}$-extremal boundaries. The form of all holomorphic retracts of such domains passing through the point with respect to which they are balanced namely the origin, follows from the theorem due to Vesentini, already stated in the introductory section; though not in as straightforward a manner, for their products. The product $B_{1}\times B_{2}$, though bounded convex and balanced, certainly fails to satisfy the condition that its boundary points be $\mathbb{R}$-extremal, even when both factors $B_{1},B_{2}$ satisfy this condition individually. Indeed even in the simplest case when $B_{1}=B_{2}=\Delta$, the only points on the boundary of $B_{1}\times B_{2}$ which are $\mathbb{R}$-extremal are the points of its distinguished boundary, which is a rather thin subset of the boundary. As mentioned at the start, our goal is to first describe the retracts of such a product $B_{1}\times B_{2}\subset\mathbb{C}^{N_{1}+N_{2}}$ which as already noted is a bounded balanced convex domain in $\mathbb{C}^{N_{1}+N_{2}}$, indeed the unit ball with respect to the norm $\parallel(w,z)\parallel=\max\{|w|_{\mu_{1}},|z|_{{\mu_{2}}}\}$ where $\mu_{1},\mu_{2}$ denote the norms in which $B_{1},B_{2}$ are the standard (open) unit balls. As mentioned in the introduction, this theme of retracts of polyballs was initiated earlier in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, what the retracts of polyballs look like; the purpose here is to address this as systematically as we can.
Thus in particular, by the above theorem 2.1, all retracts of the domain of our interest here, namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, retracts of our convex domain $D$ are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^2] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract. Infact, there are contractible spaces which do not admit a strong deformation retraction to any of its points, though obviously they admit a deformation retraction to each of its points. during the homotopy and for each point of time in the homotopy, the mapping is holomorphic. While this means that all retracts of $D$ are topologically trivial thereby placing strong restrictions on which complex submanifolds of $D$ can happen to be a retract, we shall see on the other hand that there is some variety as well: for instance if the factors $B_{1},B_{2}$ are either a polydisc or a Euclidean ball, there are retracts for every possible dimension which thereby have got to be holomorphically distinct. Furthermore again, not every contractible complex submanifold of $D$ passing through the origin can qualify to be a retract, for as we shall show, retracts through the origin of $B_{1}\times B_{2}$ are submanifolds in normal form in the sense that they are all graphs of holomorphic maps; as is very well-known, graphs of holomorphic functions are submanifolds which are simultaneously in parametric form as well as in level set form and the implicit function theorem allows us to put every submanifold locally in such a normal form (as graphs). In view of the foregoing theorem (and the implicit function theorem) then, retracts are obviously graphs of holomorphic maps (over some set of axes) locally; however in the specific case of retracts of $D=B_{1}\times B_{2}$ passing through the origin, we shall show that they are actually so globally i.e., each retract of the product domain $D$ is the graph of a single holomorphic map over some linear submanifold of $D$. As we shall encounter retracts in ‘graphical form’ in several cases in the sequel, we make a general definition for clarity.
Definition 3.2.
Let $X,Y$ be a pair of complex manifolds, $S$ a submanifold of $X$ and $f$ a $Y$-valued holomorphic map on $S$ i.e., $f$ maps $S$ into $Y$. Then, the graph of $f$ over $S$ is the subset of $X\times Y$ given by $\tilde{S}=\{(s,f(s))\;:\;s\in S\}$ (i.e., our definition views the graph as a subset of $X\times Y$ rather than $S\times Y$). Needless to say, $\tilde{S}$ is then also a complex manifold, indeed a complex submanifold of $X\times Y$.
Another general fact to be noted at the outset is that for any product space $B_{1}\times B_{2}$ where $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ are some pair of domains (or for that matter, complex manifolds), the product of a retract of $B_{1}$ with one of $B_{2}$, gives a retract of $B_{1}\times B_{2}$ and the product of retractions is a retraction. In particular, each of the factors $B_{j}$ (for $j=1,2$) are retracts of $B_{1}\times B_{2}$ as the canonical projections $\pi_{j}:B_{1}\times B_{2}\to B_{j}$ for $j=1,2$ are also retractions. A further general observation that must not be missed is that while the product of an inclusion of one factor with a retract of the other is a retract of the product domain as we just noted, arbitrary holomorphic maps between the factor domains $B_{j}$ give rise to a retract of the product: if for instance, $f:B_{1}\to B_{2}$ is any holomorphic map, then the self-map of $D=B_{1}\times B_{2}$ given by
$(z,w)\to(z,f(z))$
is actually a retraction of $D$ whose image (which is nothing but the graph of $f$) is thereby a retract thereof. More generally, graphs over retracts in $B_{1}$ or in $B_{2}$ also give rise to retracts of $D$; indeed for instance, let $\rho:B_{1}\to S\subset B_{1}$ be a retraction, $f:S\to B_{2}$ be any holomorphic map, then the self-map of $D$ given by
(3.1) $(z,w)\to(\rho(z),f(\rho(z))$
is a retraction of $D$ onto the graph of $f$ over $S$ in $D$. These preliminary observations lead us to another important one for the sequel, namely a converse: projections of those retracts  [^3] Projections of retracts need not be retracts as is demonstrated for example, by any complex line segment $l$ in the bidisc $\Delta^{2}$ whose boundary $\partial l$ is contained in $\partial\Delta\times\Delta$; note then that in such a case, the projection $\pi_{2}$ onto the second factor, when restricted to $l$ is not proper and more importantly, $\pi_{2}(l)$ is compactly contained in $\Delta$ thereby showing that it is not a retract of $\Delta$ and providing the required counterexample. Note in passing that this does not contradict lemma 2.8. of $D=B_{1}\times B_{2}$ which are graphs in $D$ over some submanifold of $B_{1}$ (resp. $B_{2}$), are retracts of $B_{1}$ (resp. $B_{2}$) as well. In light of all this, our already mentioned goal thus amounts to showing that these exhaust all possibilities for the retracts of our special case of interest $D=B_{1}\times B_{2}$, passing through the origin, any of whose associated retractions we shall denote as usual by $\rho$; its range $\rho(D)$ will be denoted by $Z$ and our goal is to realize every retract $Z$ as a graph over some linear submanifold of $D$ which we shall simply refer to as a linear subspace of $D$, a terminology that we recall here a bit more than in the first section, to introduce further associated terms.
Definition 3.3.
By a linear subspace of a balanced domain $D$ in $\mathbb{C}^{N}$ (such as our $B_{j}$’s or their product $D$ as above), we mean the intercept $D_{L}=L\cap D$ where $L$ is any complex linear subspace of $\mathbb{C}^{N}$. By the boundary of such a linear subspace which we shall denote by $\partial D_{L}$, we mean the closed set $L\cap\partial D$.
[…]
Theorem 3.4.
(Theorem 3.6 of [Mazet_Rextrm]) Let $(F,\left\|\cdot\right\|)$ be a complex Banach space, $B_{F}$ its standard open unit ball centered at the origin $B_{F}=\{v\in F\;:\;\left\|v\right\|<1\}$ and $f$ a holomorphic curve in $B_{F}$ i.e., a holomorphic map $f:\Delta\to B_{F}$. Suppose $\left\|df(0)\right\|=1$ and that $df(0)$ is $\mathbb{R}$-extremal in $\overline{B}_{F}$. Then, $f$ must be a linear map: $f(z)=df(0)z$.
It must be noted that the condition $\parallel df(0)\parallel=1$ essentially means that $df(0)$ is an isometry from $\mathbb{C}$ into $F$, a condition which underlies and entails linearity of certain maps to be considered later. If we drop the $\mathbb{R}$-extremal condition in the above theorem, then the conclusion about the linearity of $f$ may well fail, as shown by the simplest non-linear holomorphic map $f(z)=(z,z^{2})$ of the disc into the bidisc $\Delta^{2}$; indeed, note that $df(z)=(1,2z)$, so $df(0)=(1,0)$ thereby $\left\|df(0)\right\|_{l^{\infty}}=1$ verifying all other conditions except for the $\mathbb{R}$-extremal condition which we dropped.
Though we shall continue making general observations for use in the sequel, we will once and for all now, specialize our ambient space from general complex manifolds or arbitrary Banach spaces to $\mathbb{C}^{N}$ equipped possibly with a norm different from the usual $l^{2}$-norm, so as to finish noting the implications of general facts for our specific case, without much digressions from this purpose. Getting to the retract $Z$ of our product domain $D=B_{1}\times B_{2}$, let $L=T_{0}Z$ which ofcourse is a complex linear subspace of $\mathbb{C}^{N_{1}+N_{2}}$; as we’ve already mentioned, we shall not be talking of trivial retracts and hence the possible dimensions of $L$ (over $\mathbb{C}$) are $11,2,\ldots,N_{1}+N_{2}-$ 1 which are also the possible dimensions of $Z$ as a complex submanifold; which of these dimensions are attained and what would be their forms then, are matters more subtle than the general features mentioned at the outset and as we shall see, actually depend on specific features of the factor domains $B_{1},B_{2}$. However, just the assumption of $\mathbb{R}$-extremality on the boundaries of $B_{1},B_{2}$ will enable us to say far more about $Z$ than just the already mentioned fact at the outset that it is a contractible submanifold of $D$. A first reason for this (as to why we shall be able to say more at all) is that we are in the setting of a complex Euclidean space which renders all the tangent spaces of $D$ to be canonically identified and for our case, parts of tangent spaces are contained within the domain $D$ itself; more precisely and in particular, there is a part of $L$ that is contained within our domain namely, the intercept $D_{L}=L\cap D$. This means that $\rho$ can be applied to this intercept (of the tangent space to $Z$) itself, which naturally gives rise to the question of the relationship between $\rho(D_{L})$ and $Z$. As to this, the version of the identity principle for complex analytic sets (as in section 5.6 of the first chapter of the book [Chirka] by Chirka, for instance) applies to ascertain that $\rho(D_{L})=Z$ where $D_{L}=L\cap D$, provided we know apriori that $\rho(D_{L})$ is a complex submanifold – obviously, $\rho(D_{L})\subset Z$ always; the role of the identity principle is to ensure the equality $\rho(D_{L})=Z$ which will be used several times in the sequel. Actually we can do better than doing this apriori verification that $\rho(D_{L})$ is a submanifold of $Z$ in each of the multiple cases that we shall be considering for the analysis of $Z$; while splitting up into multiple cases here would be seen to naturally arise owing to the difference in the forms of $Z$ that cases result in, the result $\rho(D_{L})=Z$ does not depend on any such case we may be in and can be shown to hold rather generally. As this also fits in well with our plan of first noting general facts following from from already known results, we now proceed to establish a general result about this as in theorem 3.7 below, which holds for any bounded balanced pseudoconvex domain $D$ – no $\mathbb{R}$-extremal hypotheses is required. While the proof of this will be done by borrowing an idea [Abate_Isometry], implementation of this idea requires preparatory work which itself yields another fundamental fact namely, that $D_{L}=D\cap L$ where $L=T_{0}Z$, itself is a retract of $D$, though not every linear subspace of $D$ can be its retract, unless $D$ is biholomorphic to the $l^{2}$-ball. Infact, $D_{L}$ is a linear retract of $D$ in the sense of definition 2.5.
We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.6 and 2.7, allowing us to skip details here.
Lemma 3.5.
Let $D$ be any bounded balanced convex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin and $\mu$ the Minkowski functional of $D$. Then, $D\rho(0)$ is a norm one projection of $(\mathbb{C}^{N},\mu)$ onto $(L,\mu)$, thus mapping $D$ onto $D_{L}:=D\cap L$ and showing in particular that $D_{L}$ is a linear retract of $D$.
If we relax the convexity condition on $D$ to pseudoconvexity, it is still true that the Minkowski functional of $D$ and the infinitesimal Kobayashi metric of $D$ agree at the origin (cf. proposition 3.5.3 of [Jrncki_invrnt_dst] wherein counterexamples to show the failure of the equality when pseudoconvexity is dropped are also given), thereby implying that the Kobayashi indicatrix of $D$ at the origin is $D$ itself; likewise the Kobayashi indicatrix of $D_{L}$ now also a bounded balanced pseudoconvex domain in $L$ is also (a copy of) $D_{L}$ itself. The general property of contractivity of the Kobayashi metric under holomorphic mappings when applied to $\rho$ viewed as a map from $D$ to $D_{L}$, then yields that $D\rho(0)(D)\subset D_{L}$. But then as $D\rho(0)$ fixes $L\supset D_{L}$ pointwise, it follows that $D\rho(0)(D)=D_{L}$, showing that the above lemma holds in this more general case which we record separately for some convenience.
Lemma 3.6.
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin and $D_{L}:=D\cap L$. Then, the linear map $D\rho(0)$ is a projection which retracts $D$ onto $D_{L}$.
In case $Z$ is already a linear subspace of $D$ to begin with, the above lemma still applies to show that a linear subspace of pseudoconvex bounded balanced domain $D$, which is apriori only known to be attainable as the image of a holomorphic (i.e., possibly non-linear) retraction map on $D$, can indeed be obtained as the image of a linear retraction map as well i.e., it is valid to say that $D_{L}$ as in the above lemma is a linear retract of $D$ without any further qualifications. We are now fully prepared to prove the highlighted general result about the equality $\rho(D_{L})=Z$ whose statement is formulated with slightly different notations than the one that we had adopted for our specific domain $D$. This is indeed the content of theorem 1.10, whose statement we lay down again here for convenience, and then prove it inspired by an idea in [Abate_Isometry].
Theorem 3.7.
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then, $\rho_{|D_{L}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a (single) holomorphic function over $D_{L}$.
Proof.
We first apply the foregoing lemma which realizes $D_{L}=D\cap L$ as a linear retract of $D$ by the linear projection $\pi_{L}:=D\rho(0)$. We then consider the map $F:=\pi_{L}\circ\rho_{|_{D_{L}}}$, which is a map from $D_{L}$ into itself. Note that $F$ fixes the origin and that $DF(0)$ is just the identity map as $D\rho(0)$ is the identity map on $L$ (since $L$ is a tangent space to $Z$ which is fixed pointwise by $\rho$) and $\pi_{L}$ is actually a linear retraction onto $L$ and so in particular is the identity on $L$. As $D_{L}$ is a bounded subdomain of the complex linear space $L$ containing the origin which is fixed by $F$, we may very well apply the classical uniqueness theorem of H. Cartan to deduce that $F$ itself is just the identity map. It follows that $\rho$ restricted to $D_{L}$ and $\pi_{L}$ restricted to $Z$ are inverses of each other and being holomorphic, finishes the proof of the first statement. It only remains to observe that as the inverse of $\rho_{|_{D_{L}}}$ actually happens to be the linear projection $\pi_{L}$ onto the linear subspace $D_{L}$, this means that $Z=\rho(D_{L})$ is realized essentially as the graph of a single holomorphic function over $D_{L}$ (the only linear change of coordinates that may need to be made is to orthogonalize the projection $\pi_{L}$ which may be ‘oblique’ to render $Z$ as a standard graph over $D_{L}$). ∎
Remark 3.8.
Under the hypotheses that the boundary $\partial B$ is holomorphically extreme, it was shown in [JJ] that $Z$ is actually linear. While this hypothesis is the best possible for obtaining such a strong conclusion, it is restrictive as well for, there are plenty of balanced pseudoconvex domains whose boundaries fail to satisfy such a strong extremality condition. To give just one example whose boundary does not contain any flat pieces, which indeed satisfies the weaker $\mathbb{C}$-extremality but not the holomorphic extermality condition, consider the bounded balanced pseudoconvex domain obtained by intersecting the domain $H=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/2\}$ with the standard Euclidean ball $\mathbb{B}$ in $\mathbb{C}^{2}$. This indicates how we may construct plenty of such examples.
Remark 3.9.
Being a graph, $Z$ is easily seen to be biholomorphic to $D_{L}$ as complex manifolds, in the setting of the above theorem. Thereby, it immediately follows that any pair of retracts through the origin in $D$, of the same dimension are diffeomorphic. Indeed, being star-shaped domains in $L$, they are diffeomorphic to $L$ – no pseudoconvexity is required here; the general fact being used here is that any star-shaped domain in $\mathbb{R}^{N}$ is $C^{\infty}$-diffeomorphic to $\mathbb{R}^{N}$ (cf. exercise $7$ of chapter $8$ of [Brocker-Janich-difftop-bk]). It is therefore of interest then (as is in Several Complex Variables), to note that such a pair of retracts (through the origin of the same balanced domain $D$ and) of the same dimension despite being mutually diffeomorphic need not be biholomorphic (indeed, they are rarely so). To see this first by what is perhaps the simplest example, consider $D=\Delta^{2}\times\mathbb{B}^{2}$; let $Z_{1}$ denote the projection of $D$ onto the span of the first pair of standard basis vectors of $\mathbb{C}^{4}$ and $Z_{2}$, the projection of $D$ onto the span of the remaining last pair of standard basis vectors of $\mathbb{C}^{4}$. Noting that $Z_{1}\simeq\Delta^{2}$, $Z_{2}\simeq\mathbb{B}^{2}$ and recalling that the ball and polydisc are biholomorphically inequivalent, we deduce the same about $Z_{1},Z_{2}$. To give another example, again as simple as possible but this time of a ball $D$ which is not biholomorphic to a product domain, in-fact with boundary both smooth as well as holomorphically extreme and, which contains pairs of linear retracts (thereby through the origin), of the same dimension that are biholomorphically inequivalent, consider
$D=\{z\in\mathbb{C}^{3}\;:\;|z_{1}|^{2}+|z_{2}|^{2}+|z_{3}|^{4}<1\}.$
Note that the projection $Z_{1}$ of $D$ onto the $(z_{1},z_{2})$-plane is $\mathbb{B}^{2}\subset\mathbb{C}^{2}\hookrightarrow\mathbb{C}^{3}$; whereas the projection $Z_{2}$ of $D$ onto the $(z_{2},z_{3})$-plane is a copy of the ‘egg domain’. Specifically,
$Z_{2}\simeq\{(z_{2},z_{3})\in\mathbb{C}^{2}\;:\;|z_{2}|^{2}+|z_{3}|^{4}<1\}$
which is not biholomorphically equivalent to $\mathbb{B}^{2}$; thus, $Z_{1}\not\simeq Z_{2}$. A noteworthy little observation that shall be discussed later is that while all linear projections are retraction maps on $\mathbb{C}^{N}$, their restrictions to balls $D$ (centered at the origin, with respect to some norm on $\mathbb{C}^{N}$) need not induce retraction mappings on $D$; in-fact, most often they do not. However, orthogonal projections onto spans of coordinate vectors, in the foregoing as well as forthcoming examples do (so, $Z_{1},Z_{2}$ are indeed retracts of $D$ in the above examples). Further such examples of irreducible smoothly bounded balanced convex domains in any dimension $N\geq 3$, can be constructed by considering the domain
:=$\mathcal{E}:=\{z\in\mathbb{C}^{N}\;:\;|z_{1}|^{p_{1}}+\ldots+|z_{N}|^{p_{N}}<1\},$
where the exponents $p_{j}$’s are taken to be bigger than one (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$). Then, for every $1<k<N$, there exist a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent. This is easily seen for instance, by projecting $\mathcal{E}$ onto the span of different subsets of $k$-many of the standard basis vectors of $\mathbb{C}^{N}$ following the foregoing concrete examples.
Remark 3.10.
We shall assume until further notice that $D$ is convex. The existence in abundance of one dimensional retracts of strictly convex domains (in $\mathbb{C}^{N}$) is well-known due to the work of Lempert. Another well-known work which obtains retracts of higher dimensions under certain conditions is [Ian_Grhm] of Graham; as noted therein, the condition in the above theorem that $D_{L}=D\cap L$ be a linear retract of $D$ which is equivalent to the existence of a norm one projection onto $L$ (denoted by $\pi_{L}$ in the proof), is always met if $L$ is one dimensional – this follows by the Hahn–Banach theorem which may be used to first get a linear functional $f$ on $(\mathbb{C}^{N},\mu)$ of norm one which takes the value $1$ at a boundary point of $D_{L}$ (one may view $f$ as mapping $D$ onto $D_{L}\subset L\simeq\mathbb{C}$); the desired projection is then obtained by taking the linear projection of $\mathbb{C}^{N}$ onto $L$ along $\ker(f)$. On the other hand (also noted therein), if there exists a norm one projection of the Banach space $(\mathbb{C}^{N},\mu)$ onto a every linear subspace, then $(\mathbb{C}^{N},\mu)$ is actually a Hilbert space i.e., $\mu$ is actually induced by an inner product owing to classical results mentioned in the introduction namely that of Kakutani in [Kakutni] for the real case and Bohnenblust in [Bhnblst] for the complex case. Thus it must be noted that not every linear subspace intersected with $D$ need be a retract of $D$. But then again, we shall show in the next theorem that the retracts through the origin are all linear.
While the foregoing theorem 3.7, already realizes $Z=\rho(D_{L})$ as the graph of a holomorphic function over $L$, it must be noted that this does not mean that the graph of any holomorphic function over any given linear retract of $D$, is a retract: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well. The proof of this, atleast as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true for every bounded balanced convex domain with $\mathbb{C}$-extremal boundary. This is of-course the result of Vesentini which we extend and give a more elementary proof as already mentioned in the introduction i.e., we now lay down the proof of theorem 1.3.
Proof of theorem 1.3: Let $\rho:D\longrightarrow D$ be a retraction and $\rho(0)=0$. By lemma 3.6, $D\rho|_{0}$ is a projection of norm $1$ from $D$ onto $D_{L}=D\cap L$, where $L=T_{0}Z$. Note that $(D\rho|_{0})_{|_{D_{L}}}:D_{L}\longrightarrow D_{L}$ is the identity map on $D_{L}$ and that $D_{L}$ is a bounded balanced convex domain in $\mathbb{C}^{N}$. It follows from the proof of Krein-Milman theorem (the reader is referred to the book [Meggn] for the standard proof we have in mind here), that $D_{L}$ must have atleast one boundary point $\mathbb{R}$-extremal point, say $z_{0}$. As $D\rho(0)$ acts as the identity map on $D_{L}$ (thereby in particular $D\rho(0)(z_{0})=z_{0}$), it follows that there is atleast one $\mathbb{R}$-extremal point in $D\rho(0)(D_{L})$ namely $z_{0}$ itself! Theorem 3.7 part-(iii) of Mazet’s article [Mazet_Rextrm] applies to the map $\rho_{|_{D_{L}}}$ finishes the proof that $Z$ is linear.
To not be amiss to discuss concrete examples and the simplest ones at that, recall that the first standard choices for $D$, namely the $l^{p}$-balls was already mentioned in the introduction, specifically theorem 1.7. Recall also that we had promised an example to show the failure of this theorem for $p=\infty$ i.e., when $D$ is the polydisc.
Example: The linear map $T:\mathbb{C}^{3}\longrightarrow\mathbb{C}^{3}$ defined by $T(z_{1},z_{2},z_{3})=\left(z_{1},z_{2},(z_{1}+z_{2})/2\right)$ is a projection of norm $1$ in $(\mathbb{C}^{3},\|\cdot\|_{\infty})$. But its image is not spanned by mutually disjoint supported vectors. To see this, assume to obtain a contradiction that there exists a basis $\{v_{1},v_{2}\}$ whose supports are disjoint. Write $v_{1}=(\zeta_{1},\zeta_{2},\zeta_{3}),~{}v_{2}=(\eta_{1},\eta_{2},\eta_{3})$. Consider the case $Supp(v_{1})=\{1,2\}$ and $Supp(v_{2})=\{3\}$. Then $\zeta_{1},\zeta_{2}\neq 0,\zeta_{3}=0$ and $\eta_{1},\eta_{2}=0,\eta_{3}\neq 0$. Note that $image(T)$ is the subspace $w_{1}+w_{2}-2w_{3}=0$. Since $v_{1},v_{2}\in image(T)$, we have $\zeta_{1}+\zeta_{2}=2\zeta_{3}$ and $\eta_{1}+\eta_{2}=2\eta_{3}$ which leads to $\zeta_{1}+\zeta_{2}=0$ and $\eta_{3}=0$. It then follows that $v_{2}=0$, a contradiction. Similar arguments, show that the other cases here also do not arise.∎
We now finish off the discussion of this $p=\infty$ by proving the characterization of its linear retracts as stated in proposition 1.8.
[…]
$L=\Delta^{N}\cap Y=\{(F_{1}(z),\ldots,F_{N}(z)):z\in M\}$
Observe that the $F_{j}$’s are holomorphic functions from $\Delta^{N}$ to $\Delta$ and $F_{j}(z)=G_{j}(z_{j_{1}},\ldots,z_{j_{r}})$ for some holomorphic function $G_{j}$, which is bounded by $1$. Without loss of generality assume that each $F_{i}$’s are linear, since $F(z)=(F_{1}(z),\ldots,F_{N}(z))$ restricted to $L$ is linear. So we get $F_{i}(z)=z_{i}$, if $i\in J$ and $F_{i}(z)=c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}}$, if $i\notin J$ for some scalars $c_{i,j},~{}j\in J$.
Fix $m\notin J$ and for $1\leq k\leq N$, define
$w_{k}:=\begin{cases}\frac{|c_{m,j_{r}}|}{c_{m,j_{r}}},\ \text{if}\ k=j_{r}~{}\text{for some}~{}r\ \text{and}\ c_{m,j_{r}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}$
We know that $|F_{m}(z)|\leq 1$ for all $z\in\Delta^{N}$. If $w:=(w_{1},\ldots,w_{N})$ then
$|F_{m}(w)|=\sum_{j\in J}|c_{m,j}|\leq 1.$
Hence we get $\sum_{j\in J}|c_{m,j}|\leq 1$.
Note that $M$ is spanned by $\{e_{j_{1}},\ldots,e_{j_{r}}\}$. Since $F$ is determined by the variables $z_{j_{1}},\ldots,z_{j_{r}}$, we get $Y$ is spanned by $\{F(e_{j_{1}}),\ldots,F(e_{j_{r}})\}$. By definition of $F_{i}$, we get
$F_{i}(e_{j_{k}})=\begin{cases}1,\quad\qquad\text{if}~{}i\in J\\ c_{i,j_{k}},\qquad\text{if}~{}i\notin J,\\ \end{cases}$
By definition of $F$, we will get
$F(e_{j_{k}})=(F_{1}(e_{j_{k}}),\ldots,F_{N}(e_{j_{k}}))=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$
This shows that $Y=span\{v_{k}\}_{k=1}^{r}$ where $v_{k}=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$.
To attain the converse implication, assume that $L=Y\cap\Delta^{N}$ for some subspace $Y=span\{v_{1},\ldots,v_{r}\}$ and also satisfies the conditions (i) and (ii). Define $J:=\{j_{1},\ldots,j_{r}\}$ and
$F_{i}(z)=\begin{cases}z_{i},\quad\qquad\qquad\qquad\qquad\text{if}~{}i\in J\\ c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}},\quad\text{if}~{}i\notin J,\\ \end{cases}$
Note that $F_{i}(z)$ is a holomorphic function from $\Delta^{N}$ to $\Delta$, because by condition (ii) of the theorem. Next we define $F(z):=(F_{1}(z),\ldots,F_{N}(z))$. $F$ is a holomorphic retraction map on $\Delta^{N}$ whose image is $span\{v_{k}\}_{k=1}^{r}\cap\Delta^{N}$, where $v_{k}=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$, which completes the proof.
We already saw in the introduction section, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors. The converse implication however is true and can now be derived as a corollary.
Corollary 3.11.
If $Y$ is a linear subspace of $\mathbb{C}^{N}$ spanned by disjointly supported vectors $\{v_{1},\ldots,v_{r}\}$, then $Z=Y\cap\Delta^{N}$ is a linear retract of $\Delta^{N}$.
Proof.
Without loss of generality assume that $\|v_{i}\|_{\infty}=1$ for $i=1,\dots,r$. There exists $J=\{j_{1},\ldots,j_{r}\}$ such that $\pi_{j_{i}}(v_{i})=1$, where $\pi_{i}$ denotes the projection to the $i^{th}$ component. We can write $v_{k}=e_{j_{k}}+w_{k}$. Note that $w_{k}$’s are disjointly supported vectors. If $m\in\{1,2,\ldots,N\}\setminus J$, then we define $c_{m,{j_{k}}}:=\pi_{m}(w_{k})$. Since $\text{supp}(w_{i})$’s are disjoint, we get $\pi_{m}(w_{k})\pi_{m}(w_{l})=0$ for $k\neq l$. Hence $\sum_{j\in J}|c_{m,j}|\leq 1$. By theorem 1.8, $Y=span\{v_{1},\ldots,v_{k}\}$ is a linear retract of $\Delta^{N}$. ∎
Turning now to the proof of theorem 1.12 and proceeding forth with it, as already remarked, the boundary of our product domain $D=B_{1}\times B_{2}$ does not enjoy the $\mathbb{R}$-extremality property. Nevertheless, we shall show that $Z$ is linear (indeed, $\rho(D_{L})=D_{L}=Z$) in several cases to be dealt with in the sequel and in all cases can be realized as holomorphic graphs over some subspace contained within the coordinate spaces $\mathbb{C}^{N_{1}}$ or $\mathbb{C}^{N_{2}}$ intercepted by $B_{1}$ or $B_{2}$ respectively. That is to say, while we already know that $\rho(D_{L})=Z$ is posited in $D$ as a holomorphic graph over $D_{L}$, what we shall be more precisely showing is that $Z$ can always be realized a holomorphic graph over some retract of $B_{1}$ or $B_{2}$ which in view of the last theorem is a linear subspace of the respective ball ($B_{1}$ or $B_{2}$). Indeed, we have finished our general observations and we shall now progress into the analysis of our specific $D$. First, a few words about notations about retracts $\rho$ of our specific product domain $D\subset\mathbb{C}^{N_{1}+N_{2}}$. When written out explicitly, $\rho$ has $N_{1}+N_{2}$ components: $(\rho_{1},\rho_{2},\ldots,\rho_{N_{1}},\rho_{N_{1}+1},\ldots,\rho_{N_{1}+N_{2}})$ where each of the $\rho_{j}^{\prime}s$ for $j=1,\ldots,N_{1}+N_{2}$ can apriori depend on all of the variables $w_{1},\ldots,w_{N_{1}},z_{1},\ldots,z_{N_{2}}$; so, we use $w=(w_{1},\ldots,w_{N_{1}})$ as our notation for the coordinates on $\mathbb{C}^{N_{1}}\supset B_{1}$ and $z=(z_{1},\ldots,z_{N_{2}})$ for that on $\mathbb{C}^{N_{2}}\supset B_{2}$.
Next, note that as our non-trivial retract $Z$ is necessarily a non-compact subset of $B_{1}\times B_{2}$, it will therefore meet the boundary of $D$ i.e., $\overline{Z}\cap\partial(B_{1}\times B_{2})\neq\emptyset$ where it may noted that
$\partial(B_{1}\times B_{2})=(\partial B_{1}\times B_{2})\cup(B_{1}\times\partial B_{2})\cup(\partial B_{1}\times\partial B_{2}).$
However, rather than this disjoint decomposition of the boundary, we shall for the most part work with a slightly different decomposition in which the pieces are not disjoint but are closed (in $\mathbb{C}^{N_{1}+N_{2}}$); essentially,
(3.2) $\partial(B_{1}\times B_{2})=(\partial B_{1}\times\overline{B}_{2})\cup(\overline{B}_{1}\times\partial B_{2}).$
Moreover, we shall break up our discussion depending on where $L$ (rather than $Z$) meets the boundary of $D:=B_{1}\times B_{2}$ i.e., let $\partial D_{L}:=L\cap\partial D$ and consider the various cases:
  • (I)
    $\partial D_{L}\subset\partial B_{1}\times\partial B_{2}$.
  • (II)
    $\partial D_{L}\subset(\partial B_{1}\times B_{2})\cup(\partial B_{1}\times\partial B_{2})=\partial B_{1}\times\overline{B}_{2}$.
  • (III)
    $\partial D_{L}\subset(B_{1}\times\partial B_{2})\cup(\partial B_{1}\times\partial B_{2})=\overline{B}_{1}\times\partial B_{2}$.
  • (IV)
    $\partial D_{L}$ intersects both $\partial B_{1}\times B_{2}$ and $B_{1}\times\partial B_{2}$.
Observe that while these cases are not mutually exclusive they are indeed exhaustive; to see this clearly, first note the two obviously mutually exclusive cases:
  • (i)
    $\partial D_{L}$ is contained in one of $\partial B_{1}\times\overline{B}_{2}$ or $\overline{B}_{1}\times\partial B_{2}$, or,
  • (ii)
    case-(i) does not hold i.e., $\partial D_{L}$ is not fully contained within one of the pieces mentioned in (i).
As $\partial D_{L}$ is contained in $\partial(B_{1}\times B_{2})$ by definition, case-(ii) means by equation (3.2), that $\partial D_{L}$ must intersect both the pieces of case-(i) and conversely. Now, if $\partial D_{L}$ intersects both pieces mentioned in case-(i), then noting the subtle differences in the closure signs, we observe that either case-(IV) or case-(I) must happen in such an event. This verifies that the cases (I) through (IV) suffice to exhaust all possibilities; we note in passing that since the last case (IV), does not preclude the possibility that $L$ intersects $\partial B_{1}\times\partial B_{2}$ (and hence subsumes case-(I)), the possibility that $L$ intersects each of $\partial B_{1}\times B_{2}$, $B_{1}\times\partial B_{2}$ and $\partial B_{1}\times\partial B_{2}$ is taken care of, as well. We now take up each case in turn. We rigorously argue out these cases trying to minimize repetition at the same time, for we will provide no details for the extended product $B_{1}\times B_{2}\times\ldots\times B_{N}$.
Case-(I)
In this case, both the projections $\pi_{1}:D_{L}\to B_{1}$ and $\pi_{2}:D_{L}\to B_{2}$ are proper. Indeed, if $\{(w^{k},z^{k}):k\in\mathbb{N}\}$ is a sequence in $D_{L}$ accumulating only on its boundary, then by definition of the case-(I) that we are in now, $w^{k}\to\partial B_{1}$ and $z^{k}\to\partial B_{2}$ i.e., $\pi_{1}(w^{k},z^{k})\to\partial B_{1}$ and $\pi_{2}(w^{k},z^{k})\to\partial B_{2}$.
Next, note that ${\rm dim}(D_{L})={\rm dim}(L)\leq\min\{N_{1},N_{2}\}=N_{1}$ since proper holomorphic maps cannot decrease dimension and $\pi_{1}$ is a proper holomorphic map carrying $D_{L}$ into the $N_{1}$-dimensional $B_{1}$. Infact, as $D_{L}$ is just a linear subvariety of $D$, the properness of $\pi_{1}$ means also that the fibres of $\pi_{1}$ are actually singletons; so, $\pi_{1}$ maps $D_{L}$ biholomorphically onto its range in $B_{1}$ and the inverse of this map can be used to parametrize $D_{L}$ as the graph of a linear map over the linear subspace $\pi_{1}(D_{L})\subset B_{1}$ as:
(3.3) $D_{L}=\{\big{(}w,\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}\;:\;w\in\pi_{1}(D_{L})\}.$
for some $\mathbb{C}$-linear functionals $\beta_{1},\ldots,\beta_{N_{2}}$ on the complex linear subspace $\pi_{1}(L)\subset\mathbb{C}^{N_{1}}$. To tie this up with our goal of describing $Z$, we shall now show that $Z=D_{L}$! Indeed to this end, pick any point $(w^{0},z^{0})\in\partial D_{L}\subset\partial B_{1}\times\partial B_{2}$ and consider the complex line segment $l$ through it i.e., $l$ is the image of the map $\varphi(t)$ defined for $t\in\Delta$, by
$\varphi(t)=(w^{0}t,z^{0}t)=(w^{0},z^{0})t\;=t(w^{0},z^{0}),$
where we have written out $\varphi(t)$ in various notations only to indicate once of the alternative usages (that may be more natural in certain contexts to come later). Next compose this parametrization of $l$ with our retraction $\rho$:
$t\longmapsto\big{(}\rho_{1}(tw^{0},tz^{0}),\ldots,\rho_{N_{1}+N_{2}}(tw^{0},tz^{0})\big{)}.$
This is ofcourse the restriction of the retraction $\rho$ to $l$ and we shall now analyze it componentwise. More precisely, let
$g_{1}(t)=(\pi_{1}\circ\rho\circ\varphi)(t)=\big{(}\rho_{1}(tw^{0},tz^{0}),\ldots,\rho_{N_{1}}(tw^{0},tz^{0})\big{)}$
and
$g_{2}(t)=(\pi_{2}\circ\rho\circ\varphi)(t)=\Big{(}\rho_{N_{1}+1}(tw^{0},tz^{0}),\ldots,\rho_{N_{1}+N_{2}}(tw^{0},tz^{0})\Big{)}.$
So $g_{1},g_{2}$ are holomorphic maps from the unit disc into the balls $B_{1},B_{2}$ respectively, mapping $0\in\Delta$ to the origin in the respective balls. We are thus essentially in the setting of the Schwarz lemma of Mazet recalled above as theorem 3.4, provided only that we verify that the derivatives of these maps at the origin satisfy the norm condition as in theorem 3.4. To do this, let us begin with the map $g_{1}$ whose derivative we compute using the chain-rule as:
$dg_{1}(0)=d\pi_{1}\circ d\rho_{|_{\varphi(0)}}\circ d\varphi(0).$
Then, using the facts that $d\pi_{1}\equiv\pi_{1}$ (as the projection $\pi_{1}$ is just linear), $\varphi$ maps $0\in\Delta$ to the origin in $B_{1}$ which inturn is fixed by $\rho$, and that $d\rho|_{0}$ is just the identity map on $T_{0}Z$ and that $(w^{0},z^{0})=d\varphi(0)\in T_{0}Z$, we get $dg_{1}(0)=w^{0}\in\partial B_{1}$, verifying the norm condition $\mu(w^{0})=1$ (and the $\mathbb{R}$-extremality of $dg_{1}(0)$, which is now trivial here owing to the hyporthesis that all boundary points of $B_{1}$ are $\mathbb{R}$-extremal). Hence, the aforementioned Schwarz lemma of Mazet, applied here to $g_{1}$ shows that that $\pi_{1}\circ\rho(tw_{0},tz^{0})=tw_{0}$, which means that $\rho^{1}:=\pi_{1}\circ\rho$ acts like the projection onto the $w$-axes on $D_{L}$:
${\rho^{1}}_{|_{D_{L}}}(w,z)=w.$
The same considerations then applied to the other component $g_{2}$ gives likewise for all $t\in\Delta$ that $\rho^{2}(tw_{0},tz^{0})=tz^{0}$ where $\rho^{2}=\pi_{2}\circ\rho$. In other words, $\rho^{2}$ also acts as a projection, this time onto the $z$-space:
${\rho^{2}}_{|_{D_{L}}}(w,z)=z.$
Putting this together with the similar conclusion made about $\rho^{1}$, this just means that $\rho$ acts as the identity mapping on $l$. But this argument may be carried out for every one-dimensional linear subspace of $D_{L}$, as we are in the case that $\partial D_{L}$ is completely contained within $\partial B_{1}\times\partial B_{2}$. It therefore follows that $\rho$ must act as the identity mapping on $D_{L}$ leading to the conclusion that $D_{L}=Z$; thus, $Z$ in the present case (Case-(I)) must be of the form
$Z=\{\big{(}w,\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}\;:\;w\in\pi_{1}(D_{L})\}.$
Finally, what the fact that the boundary of $Z=D_{L}$ (what we have denoted by $\partial D_{L}$), lies in $\partial B_{1}\times\partial B_{2}$ means for the functionals $\beta_{j}$’s appearing in the above description of $Z$, is that the linear map represented by their collective (whose graph is $Z$), namely
$w\longmapsto\big{(}\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}$
is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$. Hence, we see that this places strong restrictions about when this case-(I) can happen. For instance, this case cannot occur if $N_{1}=N_{2}=N>1$ and $B_{1}=\Delta^{N}$ is the standard unit polydisc and $B_{2}=\mathbb{B}$, the standard ball with respect to the $l^{2}$-norm in $\mathbb{C}^{N}$ and ${\rm dim}(D_{L})=N$; or more generally, case-(I) cannot arise as soon as $\pi_{1}(D_{L})$ is not biholomorphic to any of the linear sections of $B_{2}$. But then again, one must not be amiss to note that case-(I) does happen atleast once regardless of what particular domains $B_{1},B_{2}$ are, namely when ${\rm dim}(D_{L})=1$; indeed, the complex line segment spanned by every point in $\partial B_{1}\times\partial B_{2}$ is a retract (as is more generally true of the $\mathbb{C}$-line segment spanned by any boundary point).
Case-(II):
$\partial L\subset\partial B_{1}\times\overline{B}_{2}$. In this case, note that the projection $\pi_{1}:D_{L}\to B_{1}$ is proper and again the general fact that proper holomorphic maps cannot decrease dimension gives ${\rm dim}D_{L}\leq{N_{1}}$. Noting that the arguments in case-(I) leading to (3.3), relied only on the properness of $\pi_{1}$ and not that of $\pi_{2}$, we see that the same arguments apply in this case as well to show again that $D_{L}$ must be of the form
$D_{L}=\{\big{(}w,\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}\;:\;w\in\pi_{1}(D_{L})\},$
for some $\mathbb{C}$-linear functionals $\beta_{1},\ldots,\beta_{N_{2}}$ on the complex linear subspace $\pi(L)\subset\mathbb{C}^{N_{1}}$; again, the fact that the boundary of $D_{L}$ lies in $B_{1}\times\partial B_{2}$ implies that the linear transformation given by
$T_{\beta}(w)=\big{(}\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}$
is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$. Picking a point $(w^{0},z^{0})\in\partial D_{L}$, considering the complex line segment joining the origin to this point and the restriction of the retraction of $\rho$ to this segment, we get by the arguments as in Case-(I) by applying the Schwarz lemma of Mazet, that ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ where $\rho^{1}=\pi_{1}\circ\rho$. As points of $D_{L}$ are of the form $(w,T_{\beta}(w))$ for $w\in\pi_{1}(D_{L})$, we have for all $w\in\pi_{1}(D_{L})$ that
(3.4) $\rho(w,T_{\beta}(w))=\big{(}w,\rho^{2}(w,T_{\beta}(w))\big{)}$
where $\rho^{2}=\pi_{2}\circ\rho$; thereby, it follows that $\rho(D_{L})$ is of the form
$\rho(D_{L})=\{\big{(}w,\rho^{2}(w,T_{\beta}(w))\big{)}\;:\;w\in\pi_{1}(D_{L})\}.$
This means that $\rho(D_{L})$ is the graph of a $\mathbb{C}^{N_{2}}$-valued holomorphic function over $B_{1}^{L}:=\pi_{1}(D_{L})$, more precisely the graph of the $B_{2}$-valued holomorphic map given by
$w\to\big{(}\rho_{N_{1}+1}(w,T_{\beta}(w)),\ldots,\rho_{N_{1}+N_{2}}(w,T_{\beta}(w))\big{)}.$
for $w$ varying through $B_{1}^{L}$, the unit ball of $\pi_{1}(L)$ with respect to the norm $\mu_{1}$. Recalling that theorem 3.7 gives $Z=\rho(D_{L})$ whose form we just determined, finishes this case.
Case-(III)
$\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$. In this case, the projection onto the second-factor namely, $\pi_{2}:D_{L}\to B_{2}$ is proper and we deduce as usual at the outset, that the general fact that proper holomorphic maps do not decrease dimension implies that ${\rm dim}D_{L}\leq N_{2}$.
We then start as in Case-(I) by picking a boundary point $(w^{0},z^{0})\in\partial D_{L}$, which we may assume to not be from the part of $D_{L}$ that lies (if it does at all) in $\partial B_{1}\times\partial B_{2}$ for this has already been dealt with in case-(I); the arguments that follow do not depend on such an assumption and we therefore do not make any assumption about which part of the boundary of $D_{L}$ the point $(w^{0},z^{0})$ comes from. Denote by $l^{\prime}$, the complex line spanned by this point and consider the restriction of our retraction map $\rho$ to the intercept $l^{\prime}\cap D$ of this line with our domain $D=B_{1}\times B_{2}$, which we call $l$. Indeed, compose the standard parametrization of $l$ namely $\varphi(t)=t(w^{0},z^{0})$ with $\rho$ followed by $\pi_{2}$ to get the holomorphic curve in $B_{2}$ given by the parametrization $g_{2}(t):=(\pi_{2}\circ\rho\circ\varphi)(t)$ for $t\in\Delta$. Noting that $g_{2}$ sends $0\in\Delta$ to the origin in $B_{2}$, computing by the chain-rule that $dg_{2}(0)=z^{0}\in\partial B_{2}$ and recalling that all boundary points of $B_{2}$ are $\mathbb{R}$-extremal, we see that all conditions required by theorem 3.4 are met, which when applied to this map $g_{2}$ yields that $g_{2}$ is a linear map: $g_{2}(t)=z^{0}t=tz^{0}$. To see the implications of this, let us write this more elaborately using the original definition of $g_{2}$ as $\pi_{2}(\rho\circ\varphi(t))=tz^{0}$, which unravels explicitly as
$\big{(}\rho_{N_{1}+1}(\varphi(t)),\ldots,\rho_{N_{1}+N_{2}}(\varphi(t))\big{)}=(tz^{0}_{1},\ldots,tz^{0}_{N_{2}}).$
This means that for all $j=1,2,\ldots,N_{2}$, the restriction of $\rho_{N_{1}+j}$ to $l$ just returns the $j$-th variable:
${\rho_{N_{1}+j}}_{|_{l}}(w,z_{1},\ldots,z_{N_{2}})=z_{j}.$
This just means that ${\rho^{2}}_{|_{l}}(w,z)=z$. Now, we may extend this independence of $\rho^{2}$ on $w$ throughout $l$, by recalling that $l$ was actually an arbitrarily chosen complex line segment from $D_{L}$. Indeed, recall that $l$ is the line spanned by $(w^{0},z^{0})$ which was just an arbitrary point of $\partial D_{L}$. Observing that every one dimensional complex linear subspace of $L$ can be obtained as the $\mathbb{C}$-span of some point of $\partial D_{L}$ (and that the foregoing arguments to reach the last equation works equally well for every complex line segment through the origin in $D_{L}$), we deduce the persistence of the foregoing equation throughout $D_{L}$. That is to say, $\rho^{2}$ restricted to $D_{L}$ does not depend on $w$ and infact,
(3.5) ${\rho^{2}}_{|_{D_{L}}}(w,z)=z$
Though we cannot claim this form for $\rho^{1}$ as well, observe that the fact that the linear projection $\pi_{2}:D_{L}\to B_{2}$ is proper, means that $\pi_{2}$ maps $D_{L}$ biholomorphically (indeed, as a linear isomorphism) onto its image $\pi_{2}(D_{L})$, a linear subspace of $B_{2}$. In particular therefore, $\pi_{2}$ is invertible over $\pi_{2}(D_{L})$. Denoting the $w$-component of the inverse of this map by $T^{\beta}$, we see (analogously to Case-(II)) that this means that points of $D_{L}$ are of the form $(T^{\beta}(z),z)$ for some (holomorphic) linear map $T^{\beta}$ of $z=(z_{1},\ldots,z_{N_{2}})$ varying in $\pi_{2}(D_{L})$. If we the compose this map with our retraction $\rho$, we deduce as before that points of $\rho(D_{L})$ are precisely of the form
(3.6) $\rho\big{(}\beta(z),z\big{)}=\big{(}\rho^{1}(T^{\beta}(z),z),z\big{)}\;\;\text{for all}\;z\in\pi_{2}(D_{L}),$
meaning that $\rho(D_{L})$ is a $k$-dimensional holomorphic graph over the linear subspace $\pi_{2}(D_{L})$ of $B_{2}$, where $k={\rm dim}(D_{L})={\rm dim}(Z)$. As this verifies in particular that $\rho(D_{L})$ is a submanifold of $Z$, we may now apply the identity principle for analytic sets or just recall theorem 3.7, which allows us to say that the whole of $Z$ is obtained by applying $\rho$ just once to $D_{L}$ i.e., $\rho(D_{L})=Z$; as we just deciphered the form of $\rho(D_{L})$, this finishes determining $Z$. To conclude this case, note that equation (3.6) gives $(\pi_{2}\circ\rho)(0,z)=z$ for all $z\in\pi_{2}(D_{L})$. Moreover, $(\pi_{2}\circ\rho)(D)=\pi_{2}(Z)=\pi_{2}(D_{L})$, since $Z$ is a graph over $\pi_{2}(D_{L})$ i.e., $\pi_{2}\circ\rho$ is a retraction of $D$ with $\pi_{2}(D_{L})$ being the corresponding retract. Restricting $\pi_{2}\circ\rho$ to $B_{2}$ also shows that $\pi_{2}(D_{L})$ is a retract of $B_{2}$ itself. Similar arguments apply to the previous and forthcoming cases as well, which we skip for some conciseness.
Remark 3.12.
The reason we are able to apply the reasoning of the first case in this case, ‘line-by-line’ to each complex line in $D_{L}$ but not be confronted with the possibility that the projection $\pi_{2}(Z)$ is a ‘ruled-surface/ruled-submanifold’ but actually a linear subspace is because of the first made observation by the argument based on theorem 3.4 that $\pi_{2}(\rho(tw_{0},tz^{0}))=tz^{0}$ which is stronger than just saying that the image of $D_{L}$ under $\rho$ when projected onto the second factor is a line – $\rho$ actually preserves the parametrization in the $z$-variables.
Case-(IV)
This is the final case when $\partial D_{L}$ intersects both $\Gamma_{1}:=\partial B_{1}\times B_{2}$ as well as $\Gamma_{2}:=B_{1}\times\partial B_{2}$. While this does not preclude the possibility that $L$ does not intersect $\Gamma_{0}:=\partial B_{1}\times\partial B_{2}$, certainly any one dimensional complex linear subspace $l^{\prime}$ which intersects $\Gamma_{j}$ for $j=1$ or $2$, intersects the boundary of $D$ only in $\Gamma_{j}$ (note firstly that the $\Gamma_{j}$’s are disjoint). That is to say, if $l=l\cap D$ then $\partial D_{L}$ is fully contained in $\Gamma_{j}$ (for $j=1$ or $2$), as soon as it contains a point from $\Gamma_{j}$. Indeed, if $(w^{0},z^{0})\in\Gamma_{j}$ then the complex line $l$ is the two dimensional real linear subspace spanned by $(w^{0},z^{0})$ and $i\cdot(w^{0},z^{0})=(iw_{0},iz^{0})$. Noting that $w^{0}$ (resp. $z^{0}$) lies in the interior of $\overline{B}_{1}$ (resp. $\overline{B}_{2}$) if and only if $iw^{0}$ (resp. $iz^{0}$) does, verifies the claim that $\partial D_{L}$ is fully contained within $\Gamma_{j}$ (for $j=1$ or $2$) as soon as it intersects $\Gamma_{j}$. As $L$ intersects both $\Gamma_{j}$ for $j=1,2$ we have in particular that $L=T_{0}Z$ cannot equal $l^{\prime}$ and ${\rm dim}(D_{L})\geq 2$. This also means that for each of $\Gamma_{1},\Gamma_{2}$ there are complex lines in $L$ which intersect it. Moreover, if $l$ is a complex line in $D_{L}$ which intersects $\Gamma_{j}$ (for $j=1$ or $2$) then lines obtained by perturbing the coefficients of linear equations which define $L$, slightly, also intersect the open piece $\Gamma_{j}$ (for $j=1$ or $2$) of the boundary of $D$ and by our forgoing remarks, the boundary of the intercept of each such line (near $D_{L}$) with $D$ is fully contained in $\Gamma_{j}$. Observe that the union of these lines contains an open subset $U_{L}$ of $D_{L}$. Next, we apply the Schwarz lemma arguments in Case-(II) (resp. Case-(III)) to each line that intersects $\Gamma_{1}$ (resp. $\Gamma_{2}$). As there are lines in $D_{L}$ which intersect $\Gamma_{1}$ as well as (other) lines in $D_{L}$ which intersect $\Gamma_{2}$, we have both the conclusions of these cases about the major pair of components of $\rho$ namely, $\rho^{j}:=\pi_{j}\circ\rho$ for $j=1,2$, which are that $\rho^{1}$ just returns the first set of variables on $D_{L}$: ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ and likewise for the second component ${\rho^{2}}_{|_{D_{L}}}(w,z)=z$. Putting together, in short we see that $\rho$ acts just as the identity map, first atleast on $U_{L}$, and thereby on $L$ as well, by the identity principle applied to both the holomorphic maps $\rho^{1},\rho^{2}$. As ${\rm dim}(D_{L})={\rm dim}(Z)$ obviously, this means that $D_{L}=Z$. Therefore, we conclude in this case that $Z$ is a linear subspace of $D$ (of complex dimension $\geq 2$), which is neither contained in $B_{1}$ nor $B_{2}$.
Conclusion
The analysis of the foregoing cases can all be now summed up and it leads precisely to theorem 1.12. Certain apparent redundancies are better clarified by the following remarks.
Remark 3.13.
Though case-(i) is a subcase of both case-(ii) and case-(iii) in the statement of theorem 1.12, we have stated it separately, for we can provide finer information in this case where the boundary $\partial D_{L}$ is contained in the distinguished boundary of $D$, namely $\partial B_{1}\times\partial B_{2}$. Indeed in case-(i), the linear retract $Z$ can be viewed as a graph of a linear map over a linear subspace in $B_{2}$ as well; this cannot be done in case-(iv) in general: consider for instance, the product of a one-dimensional retract of $B_{1}$ with one of $B_{2}$ which obviously cannot be expressed as a graph of a function neither over a linear subspace in $B_{1}$ nor over one in $B_{2}$ (unless the dimension $N=2$). Nevertheless, one may broadly say as in the opening sentence of the theorem, that retracts in all cases are indeed graphs of holomorphic maps over some linear subspace in the (full) domain $D$ just by theorem 1.10 itself. However, what cannot be derived from that theorem but is readily implied by the above theorem is that certain graphs are ruled out from being retracts – for instance, graphs of non-linear holomorphic maps which intersect $\partial B_{1}\times B_{2}$ and $B_{1}\times\partial B_{2}$.
If we drop case-(i), viewing it as a subcase of case (ii) or (iii), then it may be noted that cases (ii) through (iv) are both mutually exclusive and exhaustive as well. Next, it must be noted that each of the possibilities expressed by cases (i) through (iv) are actually realized and can be used to construct the retracts. However, we note that though this theorem seems to provide a complete characterization of the retracts of $D$, we cannot pinpoint which of the linear subspaces of $D$ can happen to be its retract much as we cannot say which linear subspaces of $B_{1}$ or $B_{2}$ are their (linear) retracts as well – this depends on the specific nature of the balls $B_{j}$ i.e., the geometric features of their boundaries, for it is known for certain that not every linear subspace of $D$ can be its retract in general, for instance even in the simplest case $B_{1}=B_{2}=\Delta$, in which case however we can indeed so pinpoint, as was precisely the purpose of our theorem 1.8 above. Finally, we note that the maximum dimension of the nonlinear retracts of $D$ is $\max\{N_{1},N_{2}\}$ far less, in general, than the dimension of $D$ namely, $N_{1}+N_{2}$.
More importantly perhaps is the following observation.
Remark 3.14.
Although $Z$ is biholomorphic to $D_{L_{1}}$, it is unreasonable to expect at-least in general, that there is an automorphism of $D$ which maps $Z$ biholomorphic $D_{L_{1}}$. This fails already for the simplest polyball namely, the polydisc, wherein such a failure can be seen as follows. All the automorphisms of the polydisc are linear fractional transformations, thereby each one of them maps a complex linear subspaces to at-most a quadratic complex surface. Hence, as soon as $Z$ is ‘sufficiently’ non-linear (there are plenty of retracts in the polydisc which are graphs of monomial maps of degree $>2$ for instance, which are herein regarded as ‘sufficiently’ non-linear), it is certainly clear that there does not exist any automorphism of the polydisc mapping $Z$ onto $D_{L_{1}}$. Theorem 1.12 may therefore be viewed as providing an answer to the problem of recasting retracts in a domain into a simple ‘normal’ form (which cannot be reduced to a further simple one in general), by applying some appropriate automorphism of the domain.
Remark 3.15.
The condition of strict convexity is equivalent to the notion of $\mathbb{R}$-extremality (as will be discussed in the next section), which is obtained just by changing the ground field from $\mathbb{C}$ to the reals in that of $\mathbb{C}$-extremality already used in Vesentini’s theorem. The proof of the foregoing theorem will be phrased in a manner to make it clear that as a corollary to its proof it will follow that: we may relax the strict convexity condition on the pair of balls $B_{1},B_{2}$ to $\mathbb{C}$-extremality, to get the same conclusions for those linear retracts $D_{L}$ of $D=B_{1}\times B_{2}$ with the property that some point of $\partial D_{L}$ is an $\mathbb{R}$-extremal point of $\partial D$. This relaxation allows the applicability of this result to the case of balls which are $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal, for which simple examples exist: indeed, the ball in $\mathbb{C}^{N}$ with respect to the $l^{1}$-norm is one such (though balls in all other $l^{p}$-norms for $1<p<\infty$ satisfy the stronger condition of $\mathbb{R}$-extremality.
As mentioned in the introduction, we shall skip the details of the extension of the arguments above to the case of the product of an arbitrary number of balls and only lay down the end result.
Theorem 3.16.
Let $B_{1},B_{2},\ldots,B_{N}$ be a collection of bounded balanced strictly convex domains possibly in different complex Euclidean spaces (say, $B_{j}\subset\mathbb{C}^{n_{j}}$ for $j=1,\ldots,N$). Then, every non-trivial retract through the origin of $D=B_{1}\times B_{2}\times\ldots\times B_{N}$ is a holomorphic graph over a linear subspace which itself is a retract of $D$; if the retract is not linear, this linear subspace is infact a subspace of the product of a subcollection of these balls $B_{j}$’s with values in the product of the remaining balls in this collection.
More specifically and precisely, let $L=T_{0}Z$, $D_{L}=L\cap D$ and $\partial D_{L}=L\cap\partial D$. For each $p\in\{1,2,\ldots,N\}$, denote by $\Sigma_{p}$, the set of all $p$-tuples $(j_{1},j_{2},\ldots,j_{p})$ where $\{j_{1},j_{2},\ldots,j_{p}\}\subset\{1,2,\ldots,N\}$ with $j_{1}<j_{2}<\ldots<j_{N}$ (the cardinality of $\Sigma_{p}$ is ($N\choose p$) and note that there exists any given $Z,D_{L}$ as above, $p\in\{1,2,\ldots,N\}$ such that $\partial D_{L}$ is contained in that part of the boundary of the convex body $\overline{D}=\overline{B}_{1}\times\overline{B}_{2}\times\ldots\times\overline{B}_{N}$, obtained by replacing the factor $\overline{B}_{l}$ herein by $\partial B_{l}$ for each $l=j_{1},j_{2},\ldots,j_{p}$ where $\sigma=(j_{1},j_{2},\ldots,j_{p})\in\Sigma_{p}$. Then, $Z$ is the graph of a $B^{\sigma_{c}}$-valued holomorphic map over a $q$-dimensional linear retract of $B_{j_{1}}\times B_{j_{2}}\times\ldots\times B_{j_{p}}$ (which is included in the canonical manner into our product domain $D=B_{1}\times B_{2}\times\ldots\times B_{N}$), where:
  • (i)
    $q\leq n_{j_{1}}+\ldots+n_{j_{p}}$ and $(j_{1},j_{2},\ldots,j_{p})\in\Sigma_{p}$ and,
  • (ii)
    $B^{\sigma_{c}}$ is the product of balls given by $B_{k_{1}}\times B_{k_{2}}\times\ldots\times B_{k_{N-p}}$ where $\sigma_{c}\in\Sigma_{N-p}$ is the $(N-p)$-tuple complementary to $\sigma$ in the sense that $\sigma^{c}=(k_{1},\ldots,k_{N-p})\subset\{1,\ldots,N\}\setminus\{j_{1},\ldots,j_{p}\}$ and $k_{1}<k_{2}<\ldots<k_{N-p}$.
Let $m=\min\{n_{1},n_{2},\ldots,n_{N}\}$, $\hat{\mu}_{m}=\max\{\mu_{1},\mu_{2},\ldots,\mu_{m-1},\mu_{m+1},\ldots,\mu_{N}\}$, where for all $j=1,2,\ldots,N$ we denote by $\mu_{j}$, the Minkowski functional of $B_{j}$. Then, in the subcase when $\partial L$ is contained within the distinguished boundary of $D$ namely, $\partial B_{e}:=\partial B_{1}\times\partial B_{2}\times\ldots\partial B_{N}$, the retract $Z$ is actually a linear subspace of $D$ occurring as the graph of a (holomorphic) linear isometry of the Banach space $(\pi_{m}(L),\mu_{m})$ into the Banach space $(\mathbb{C}^{n_{1}}\times\ldots\times\mathbb{C}^{n_{m-1}}\times\mathbb{C}^{n_{m+1}}\times\mathbb{C}^{n_{N}},\hat{\mu}_{m})$, and therefore maps $B_{m}^{L}:=\pi_{m}(D_{L})$ which is the unit ball in $\pi_{m}(D_{L})$ with respect to $\mu_{m}$, isometrically into $B_{1}\times\ldots B_{m-1}\times B_{m+1}\times\ldots\times B_{N}$.
In the pending case when $\partial D_{L}$ intersects more than one connected component of $\partial D\setminus\partial B_{e}$, the retract $Z$ is again a linear subspace of $D$ of complex dimension equal to the number of connected components that it intersects.
The simplest case for $D$ as in the above theorem is obtained by taking all the $B_{j}$’s therein to be just the unit disc, in which case ofcourse $D$ is the standard unit polydisc $\Delta^{N}$, whose retracts were determined originally by Heath and Suffridge [Heath_Suff]. To recover their result from the above theorem which only characterizes the retracts through the origin, we only have to note that the automorphism group of $\Delta^{N}$ acts transitively thereby implying that all retracts of $\Delta^{N}$ are obtained as biholomorphic images of those passing through the origin under some automorphism of $\Delta^{N}$; recalling that all automorphisms of $\Delta^{N}$ are actually fractional linear maps, we also note that the main essential feature of retracts through the origin described in the above theorem namely, as holomorphic graphs over some linear subspace is inherited by all retracts of $\Delta^{N}$, as fractional linear maps carry linear subspaces to affine-linear subspaces.
Similar considerations apply whenever the factors $B_{j}$ are homogeneous domains; recall that a domain is said to be homogeneous if its automorphism group acts transitively and examples of bounded balanced convex domains which are also homogeneous include apart from $\Delta^{N}$, the standard Euclidean ball $\mathbb{B}^{N}$ and more generally, products of Cartan domains of various types (better known as bounded balanced symmetric domains). In such a case their product $D=B_{1}\times\ldots B_{N}$ is homogeneous as well and all its retracts are obtained as images under some element of ${\rm Aut}(D)$, of those retracts of $D$ passing through the origin which have been determined by the above theorem. However, we may not be able to say that every such retract (unless it passes through the origin), is a holomorphic graph over a linear subspace since elements of ${\rm Aut}(D)$ may well not be so simple (as linear fractional transformations) to preserve linearity as in the aforementioned instance of the polydisc. To give another concrete instance, where the main features of the retracts through the origin are indeed preserved, let us take each $B_{j}$ in the above theorem to be the standard unit ball with respect to the $l^{2}$-norm; so, let $\mathbb{B}_{1},\ldots,\mathbb{B}_{N}$ be standard Euclidean balls in complex Euclidean spaces of possibly different dimensions. Then as ${\rm Aut}(D)={\rm Aut}(\mathbb{B}_{1})\times\ldots\times{\rm Aut}(\mathbb{B}_{N})$, we see that the product domain $D=\mathbb{B}_{1}\times\ldots\times\mathbb{B}_{N}$ is also homogeneous and more importantly, all of its automorphisms are fractional-linear maps. However, despite all retracts of each of these $l^{2}$-balls being affine-linear, this does not mean that every retract of $D$ is affine-linear because it must be kept in mind that the operation of taking Cartesian product leads to new retracts in the form of holomorphic graphs (even through the origin), which are not linear.
The foregoing remarks for a product of Euclidean balls apply as well to the case of products of balls with polydiscs such as $\Delta^{k}\times\mathbb{B}$ (or a larger collection of Euclidean balls and polydiscs in any order) which are also homogeneous domains with all of is automorphisms being fractional-linear maps. We conclude by noting that in these cases the characterization of retracts provided by theorems 1.12 and 3.16, is complete.
4. Proof of theorem 1.13
Let us begin by noting that it follows from theorem 1.10 proved in the foregoing section that every retract of a bounded balanced homogeneous domain is biholomorphic to the graph of some holomorphic function over a linear subspace of $D$.
Since $\Omega$ is homogenous, $\Omega/Aut(\Omega)$ is singleton and hence compact. By the Main theorem in [Fost], there exists a closed complex submanifold $Z$ of $\Omega$ such that M is biholomorphic to a locally trivial holomorphic fibre bundle with fibre $\mathbb{C}$ over $Z$. As mentioned in the begining, by theorem 1.10 we get that Z is biholomorphic to a linear retract $\Omega_{L}:=\Omega\cap L$ for some linear subspace $L$ of $\mathbb{C}^{N}$. We can cover $M$ by coordinate charts biholomorphic to $\{U_{\alpha}\times\mathbb{C}\}$ where $U_{\alpha}$ is an open cover of $Z$. The coordinate transformations from the $\alpha$ to the $\beta$- coordinates are of the form
$\varphi_{\alpha\beta}(z,w)=(z,h_{\alpha\beta}(z)w+k_{\alpha\beta}(z))$
for some holomorphic functions $k_{\alpha\beta},h_{\alpha\beta}$ on $U_{\alpha}\cap U_{\beta},h_{\alpha\beta}\neq 0$. The data $h_{\alpha\beta}$ determine a multiplicative Cousin problem on $Z$. Since $Z$ is homeomorphic to polydisk and Oka’s theorem in [Oka](exposited in theorem 3.9 of [oka_0] as well), multiplicative Cousin problem for $Z$ can be solved. Hence we might have assumed already above that $h_{\alpha\beta}\equiv 1$ for all $\alpha,\beta$. However in that case, the data $k_{\alpha\beta}$ determine an additive Cousin problem. By Oka’s theorem in [Oka_1] (exposited in theorem 3.6 of [oka_0]), additive Cousin problem can be solved. Hence $M$ is biholomorphic to $Z\times\mathbb{C}$.
Remark 4.1.
We recall for one last time, that the class of homogeneous bounded balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (cf. the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein). It follows that a domain $\Omega$ as in the above theorem is a polyball (and convex), in-particular.
5. Proof of Theorem 1.9
It must first be noted that by Theorem 1.3, every holomorphic retract through the origin is actually linear i.e., $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Of-course, we are only interested in non-trivial retracts, so $Y_{q}$ is of the form $Y_{q}=D_{q}\cap Y$ for some $r$-dimensional $\mathbb{C}$-linear subspace $Y$ of $\mathbb{C}^{N}$. We only need to show that if $Y$ cannot be of the form $\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r},~{}r+1\leq i\leq N\}$ of $\mathbb{C}^{N}$, where not all $(a_{i,j})$ are zero.
To prove this proposition by contradiction, we assume that $Y_{q}$ is a linear retract of $D_{q}$ and a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Let $\{e_{i}:1\leq i\leq N\}$ be the standard basis of $\mathbb{C}^{N}$. Define another basis $\{v_{i}:1\leq i\leq N\}$ of $\mathbb{C}^{N}$, as follows:
$v_{j}$ $=(\underbrace{0,\ldots,0,t_{j},0,\ldots,0}_{r},a_{j,r+1}t_{j},\ldots,a_{j,N}t_{j})\quad\text{for}~{}1\leq j\leq r,$
$v_{j}$ $=e_{j}\qquad\text{for}~{}r+1\leq j\leq N$
where
$t_{j}=\frac{1}{\left(1+\sum_{k=r+1}^{N}|a_{j,k}|^{q}\right)^{1/q}}$
For $1\leq j\leq r$, write
$e_{j}=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}v_{l}\right)$
After applying $P$ on both sides, we get
$P(e_{j})=\frac{1}{t_{j}}P(v_{j})-\left(\sum_{l=r+1}^{N}a_{j,l}P(v_{l})\right)$
Since $P$ maps $\mathbb{C}^{N}$ into $Y$, $P(v_{j})\in Y$. Write
(5.1) $P(v_{j})=\left(u_{1,j},\ldots,u_{r,j},\sum_{k=1}^{r}a_{k,r+1}u_{k,j},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,j}\right)~{}\text{for}~{}r+1\leq j\leq N$
Since $P$ is a projection linear map from $\mathbb{C}^{N}$ onto $Y$ and $v_{j}\in Y$ for $1\leq j\leq r$, $P(v_{j})=v_{j}$ for $1\leq j\leq r$. Hence for $1\leq j\leq r$, we have
$P(e_{j})=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}\left(u_{1,l},\ldots,u_{r,l},\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)\right)\\ =\left(-\sum_{l=r+1}^{N}a_{j,l}u_{1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{j-1,l},1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l},-\sum_{l=r+1}^{N}a_{j,l}u_{j+1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{r,l},\right.\\ \left.a_{j,r+1}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,a_{j,N}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)$
Note that $P(v_{j})\in\overline{D_{q}}$ for all $j$ So, we get the following inequality for $1\leq j\leq r$,
(5.2) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}\left|a_{j,s}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,s}u_{k,l}\right|^{q}\leq 1$
The following inequality holds for $r+1\leq j\leq N$:
(5.3) $\sum_{k=1}^{r}|u_{k,j}|^{q}+\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}~{}u_{k,j}\right|^{q}\leq 1$
By reverse triangle inequality for square roots: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.2) reduces to
$-\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0$
Combining the above inequality with (5.3), we obtain the following inequality
$-\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,j}\right|^{q}\right)+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0$
Simplifing above inequality, we get
$\sum_{l=r+1}^{N}|a_{j,l}|^{q}\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}\leq 0$
From this inequality, we obtain
(5.4) $a_{j,l}u_{k,l}=0~{}~{}\quad r+1\leq l\leq N~{},1\leq j,k\leq r~{}\text{with}~{}j\neq k.$
From inequality (5.2), we get
(5.5) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|\leq t_{j},\quad 1\leq j\leq r$
Since not all $(a_{i,j})$ are zero, atleast one of $a_{i,j}$ is non-zero. If $a_{j_{0},l_{0}}\neq 0$ for some $j_{0},l_{0}$ then from (5.4) we obtain that $u_{j,l_{0}}=0$ for all $j$ with $j\neq j_{0}$. From (5.3), we get $|u_{j_{0},l_{0}}|\leq t_{j_{0}}$. We can conclude that for all $(j,l)$, we have $|u_{j,l}|\leq t_{j}$. Using (5.5), we get
$|1|$ $=\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}+\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|$
$\leq\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|+\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|$
$\leq t_{j}+t_{j}(\sum_{l=r+1}^{N}|a_{j,l}|)<1$
This is obviously a contradiction, which completes the proof. ∎
To make some noteworthy remarks here, let us also record a special case that characterises the non-trivial retracts through the origin of such a domain in $\mathbb{C}^{2}$ where this result is not new (though our proof is direct and more elementary).
Proposition 5.1.
Let $q$ be an arbitrary positive number less than one. Consider the $\ell^{q}$-‘ball’ given by $D_{q}=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<1\}$ and the subspace $Y=\{(z,w)\in\mathbb{C}^{2}:z-mw=0\}$ (where $0\neq m\in\mathbb{C}$) of $\mathbb{C}^{2}$. Then the one dimensional subspace $Y_{q}=D_{q}\cap Y$ is not a retract of $D_{q}$. In-fact, the holomorphic retracts through the origin are linear and more precisely, the intercepts of the coordinate axes with $D_{q}$.
Owing to remarks made in the introduction (in greater detail, the work of section 7 below), it follows as an immediate corollary that $D_{q}$ is not biholomorphic to any product domain; it must also be noted here that retracts can be used to immediately show inequivalence with convex domains – for instance, the non-existence of retracts of $D_{q}$ in certain directions suffices to prove its biholomorphic distinctness from bounded convex domains (the unbounded case is ruled out simply by hyperbolicity of $D_{q}$). A more important relevance of this example is noted next.
Remark 5.2.
We recall from [Vigue] that fixed point sets of holomorphic self-maps of a any given bounded (not necessarily balanced) convex domain $D$ are precisely the same as its retracts. The example of the above proposition proves that this coincidence fails as soon as we drop convexity. Indeed, for the non-convex (but pseudoconvex) balanced domain $D_{q}$ in $\mathbb{C}^{2}$ as above, the fixed point set of the holomorphic self-map of $D_{q}$ as simple as the flip $(z,w)\to(w,z)$, namely the linear subspace $\{(z,w)\in\mathbb{C}^{2}\;:\;w=z\}$, is not realizable as a holomorphic retract of $D_{q}$, which is part of the content of the proposition 5.1 and its higher dimensional analogue in theorem 1.9 above.
6. Proof of theorem 1.14
Let $\pi:=D\rho|_{\rho(0)}$ and $L:=T_{p}Z$. Since $\rho$ is a submersion at the origin, $D\rho_{|_{0}}:\mathbb{C}^{N}\longrightarrow T_{p}Z$ is onto. Next, write
$\mathbb{C}^{N}=Ker(D\rho_{|_{0}})\oplus V_{p},$
where $V_{p}$ is some complemented linear subspace of $Ker(D\rho_{|_{0}})$. Now, the submersivity condition means that $D\rho_{|_{0}}$ when restricted to $V_{p}$ maps $V_{p}$ onto $T_{p}Z$. Since ${\rm dim}(V_{p})=\rm dim(T_{p}Z)$, it follows that $D\rho_{|_{0}}:V_{p}\longrightarrow T_{p}Z$ is a bijective linear map i.e., a linear isomorphism. Now, consider the map
$G=(D\rho_{|_{0}})^{-1}\circ\pi\circ\rho:D_{L}\longrightarrow V_{p}.$
Note then that
$DG_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}\circ\pi\circ D\rho_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}\circ\pi(D\rho_{|_{0}}(v)).$
Further, note that $\pi_{|_{T_{p}Z}}$ is just the identity map. So we get $DG_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}(D\rho_{|_{0}}(v))=v$. We then get by Mazet’s Schwarz lemma that $G$ acts as just the identity map on $D_{L}$. Note also that
$(D\rho_{|_{0}})^{-1}\circ\pi\circ\rho\equiv I_{D_{L}}.$
Therefore, $Z={\rm image}(\rho)$ is the graph of holomorphic function over some linear subspace $D_{L}$ of $D$, finishing the proof.
We ought to make a few remarks to settle any first questions that may arise as to whether the result is vacuous(!) i.e., whether there exist retracts not passing through the origin at all. For this let us only recall Lempert’s results which guarantee the existence of one-dimensional retracts $Z$ through any given pair of points $p,q$ which we take to be linearly independent; if $Z$ passes through origin, then the strict convexity hypotheses here leads by Vesentini’s theorem to the conclusion that $Z$ is linear. But then the linear independence of $p,q$ means that ${\rm dim}(Z)>1$; this contradiction finishes the verification about the existence of many retracts not passing through the origin, the matter of the above theorem. This discussion also answers another basic question: are there examples of a bounded balanced domain of holomorphy $D$, wherein there exist pairs of one-dimensional retracts which are not ${\rm Aut}(D)$-equivalent i.e., there is no automorphism carrying one onto the other? In light of the above discussion which actually only requires $\mathbb{C}$-extremality, we may readily cite $D$ to be the $\ell^{1}$-ball in which the origin is fixed by every of its automorphisms; a retract through any pair of linearly independent points in $D$ is therefore not ${\rm Aut}(D)$-equivalent to any of the retracts passing through the origin.
To address other hypotheses of the foregoing theorem briefly, let us remark that a non-trivial holomorphic retraction map is always a submersion except perhaps on a set of measure zero (cf. page $74$ of the book [Hrsch]). Nevertheless, we admit that it would be more significant to ascertain whether the origin falls outside such a measure zero set, for some retraction map $\rho$ as in the above theorem; or, if all retraction maps of such a class of domains is a submersion throughout the domain, making the submersive hypotheses redundant.
Perhaps a more useful remark would be that as a corollary of the proof of this theorem, it will be seen that the strict convexity can be relaxed to $\mathbb{C}$-extremality to obtain the same description of $Z$, as soon as it ascertained that $L\cap\partial D$ has at-least one point which is $\mathbb{R}$-extremal in $\partial D$.
7. Proof of proposition 1.6
Before going into the proof, we need to make some general remarks; to indicate as to why there are enough instances of domains verifying the conditions stipluated in proposition 1.6. First let us indicate as to why it may perhaps be of mild interest atleast, to note in passing that one may deduce out of the general facts about retracts and the already mentioned well-known characterization of retracts of polydiscs, yet another proof of the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis namely, the biholomorphic inequivalence of two of the simplest bounded simply connected domains the ball $\mathbb{B}$ and $\Delta^{N}$ (despite their diffeomeorphic equivalence), as soon as the dimension $N$ gets any bigger than one; and other such inequivalences using retracts. We have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (and so that proof is not repeated here). However here, we would indeed like to hasten to mention that this does not obtain for such $D$, results as good as those, which already well-known techniques in this realm can! – so, this leads to the question as to why one may want to even consider this proof even if the problem of biholomorphic inequivalence and its special case of establishing irreducibility of a domain $D$ i.e., showing that $D$ is not biholomorphic to a product of lower dimensional domains, is undoubtedly fundamental. To address this briefly here, let us recall that while the well-known techniques alluded to above, more explicitly such as that of Remmert – Stein (exposited well in [Encyclo-Vol6]; more specifically, see theorem $4.1$ therein alongwith the subsequent theorem $4.2$ due to Rischel) for showing irreducibility of domains, work for far more general domains $D$ beyond balanced domains of holomorphy, they impose the condition that $D$ not have any non-trivial analytic varieties in the boundary. We shall presently show that the route via retracts can very well dispense with such conditions, if $D$ is a balanced domain of holomorphy. To exposit this better by comparing and contrasting with the Remmert – Stein techniques, let us mention that the aforementioned proof by retracts for $\mathbb{B}\not\simeq\Delta^{N}$, shares (unlike a proof based on comparison of automorphism groups for instance), with the best possible techniques in this context, the essential idea about the difference in the lack or presence of analytic discs in the boundary of $D$ as being responsible for such inequivalences, as it has a bearing on the geometry of retracts through the origin in $D$. Moreover, this provides a shorter way to establish inequivalences without having to worry about the boundary extendibility (or boundary behaviour) of biholomorphic mappings since the boundary geometry is ‘internalized’ as features about the retracts through points in the interior; we believe this will be useful for establishing new inequivalences between pairs of bounded balanced domains in both of whose boundaries there reside non-trivial analytic discs. To illustrate this point with a concrete and simple example, consider the question of whether the analytic polyhedron
$D_{h}=\{(z,w)\in\mathbb{C}^{2}:|w^{2}-z^{2}|<1\}\cap\{(z,w)\in\mathbb{C}^{2}:|zw|<2\}$
is biholomorphic to a product domain  [^4] A motivation for the specific choices here, is that the simplest examples of bounded product domains namely polydiscs are defined as sublevel sets of a pair of (homogeneous) holomorphic linear polynomials and we are interested in bounded domains whose definition bears some semblance to that of polydiscs, but are nevertheless irreducible.. To begin with analysing this example first, we note that while an open set cut-out by such polynomial inequalities in general may fail to be connected, such issues do not arise here: the homogeneity of the defining pair of quadratic polynomials leads to $D_{h}$ being balanced, thereby (immediately seen to be connected and) contractible. Let us note in passing that $D_{h}$ is not Reinhardt; more importantly, that $D_{h}$ is indeed bounded: to verify this, note that the defining pair of inequalities together leads to
(7.1) $|z+iw|^{2}=\left|(z^{2}-w^{2})+2izw\right|\leq|z^{2}-w^{2}|+2|zw|<5.$
The reverse triangle inequality
$\left||z|-|w|\right|\leq|z+iw|$
when squared and combined with 7.1, leads to
$|z|^{2}+|w|^{2}<5+2|zw|.$
Looking at the second inequality in the definition of $D_{h}$, we arrive at the conclusion that $|z|^{2}+|w|^{2}<9$, meaning $D_{h}\subset 3\cdot\mathbb{B}$ finishing the desired verification. Now, the simply connectedness of $D_{h}$ (unfortunately tempts) an application of the Riemann mapping theorem to reduce the problem of establishing irreducibility of $D_{h}$ directly to that of proving the inequivalence of the hyperbolic domain $D_{h}$ with $\Delta^{2}$. The balancedness of the special model domain here namely $\Delta^{2}$, together with its homogeneity, enables a reduction of the problem by an application of Cartan’s theorem to linear equivalence; this, in-turn is immediately settled as linear maps preserve convexity whereas $D_{h}$ is non-convex.
However, we would like to show this and other modified versions now by circumventing use of the Riemann mapping theorem which may not work in dealing with more complex cases. Indeed, to spoil the above proof which is based on the convenient availability of a model domain in the equivalence class of balanced domains which are biholomorphic to products domains as nice as the bidisc, let us remove an analytic hypersurface $A$ not passing through the origin [^5] Not removing the origin thus, helps in rendering a short illustration of our use of retracts, owing to scarcity of results that we can readily cite and use. It may be noted to begin with here that this modified domain only has more analytic varieties in the boundary, that is to say: the modification to spoil the foregoing proof, has not been done at the cost of the very purpose of the previous example!, and consider $D_{h}^{A}=D_{h}\setminus\{A\}$. While the removal of $A$ from $D_{h}$ never disconnects $D_{h}$, it always destroys the simply connectedness of $D_{h}$, as is always the case when proper analytic sets are subtracted from any domain. Though not required by the statement of proposition 1.6, the case $D=D_{h}$ also serves as a concrete example of a (hyperbolic) balanced domain of holomorphy with analytic discs in its boundary (without purposefully removing analytic sets!), which feature as already mentioned is the case of greatest interest in contributing to any new results or routes about establishing non-existence of a product structure in a domain.
To establish the irreducibility of such $D_{h}^{A}$, let us start with the most natural attempt of assuming to obtain a contradiction, the existence of a biholomorphism $F$ from $D_{h}^{A}$ onto a product domain $D$; no less standard then, is to analyze the boundary behaviour of $F$. However, this is rather cumbersome owing to potential complexity of the boundary of the factors $B_{1},B_{2}$. It is here where consideration of retracts provides a far more convenient approach, as already alluded to in the above.
However, before we discuss the proof of propositions about the irreducibility of analytic complements as in proposition 1.6, it is convenient to formulate part of its proof as the following more general lemma.
Lemma 7.1.
Let $D$ be any bounded domain in $\mathbb{C}^{N}$ with a Stein neighbourhood basis. Then any holomorphic retraction map on $D\setminus A$, where $A$ is a proper analytic subset of $D$ can be extended to a holomorphic retraction map on $D$. Consequently, every retract $R^{A}$ of $D\setminus A$ is of the form $R^{A}=R\cap(D\setminus A)$ for some retract $R$ of $D$.
Proof.
Firstly recall that removing an analytic set from a domain does not disconnect it; by definition $A$ is closed in $D$, thereby $D\setminus A$ is a domain in $\mathbb{C}^{N}$. Let $\rho:D\setminus A\to D\setminus A$ be a holomorphic retraction map. By applying the Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each of the component functions of $\rho$ we get a (uniquely determined) holomorphic extension of $\rho$ which we denote by $\tilde{\rho}$. The fact that $\rho$ maps $D\setminus A$ into itself implies firstly that $\tilde{\rho}$ maps $D$ into its closure $\overline{D}$. However, it is easy to see that $\tilde{\rho}$ actually maps $D$ into itself. Indeed, pick any $a\in A$ and enquire if it is possible for $b=\tilde{\rho}(a)$ to lie in $\partial D$. Recalling the local structure of analytic sets, we know that: there exists a complex line segment $\ell\subset D$ through $a$ which intersects $A$ only at $a$. If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $D$. Since $D$ has a Stein neighbourhood basis, there exists an exhaustion function $\psi$ for $D$ which is plurisubharmonic on an open set containing $D$ in particular. Considering then the composition $\psi\circ\tilde{\rho}\circ\varphi$, we see that the maximum principle for this plurisubharmonic function is violated, under our assumption that $b\in\partial D$. Thus $\rho(D)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
We finish off by verifying that $\tilde{\rho}$ is a retraction on $D$. Indeed, we need only verify the required idempotence equation along $A$, as outside of $A$, $\tilde{\rho}$ is the given retraction map $\rho$ on $D\setminus A$. To do this, pick an arbitrary $a\in A$ and a sequence $(z_{n})\subset D\setminus A$ with $z_{n}$ converging to $a$. Then just by the idempotence of $\rho$ on $D\setminus A$, we have: $\rho\left(\rho(z_{n})\right)=\rho(z_{n})$, wherein by passing to limit using the continuity of $\tilde{\rho}$ twice, we get $\tilde{\rho}\left(\tilde{\rho}(a)\right)=\tilde{\rho}(a)$. As noted earlier, we may now conclude that the same equation holds not just on $A$ but throughout $D$ i.e., $\tilde{\rho}\circ\tilde{\rho}\equiv\tilde{\rho}$; denoting $\tilde{\rho}(D),\rho(D\setminus A)$ by $R^{A}$ and $R$ respectively, we arrive at the statement of the lemma. ∎
We now detail the proof of proposition 1.5 and lay down some of its generalizations as well. This will be used in the sequel.
Proof of proposition 1.5 . Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ and $v_{2}\in T_{p_{2}}{D_{2}}$. Since $D_{1}$ is bounded, there exists $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$. Since $D_{2}$ is open, there exists $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$. Consider the domain $D:=D_{1}\times D_{2}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $|v_{2}|<\frac{r_{2}}{r_{1}}$. Consider a holomorphic map $F:D_{1}\to\mathbb{C}$ defined by $F(z)=p_{2}+(z-p_{1})v_{2}$. Note that $F(p_{1})=p_{2}$ and $F^{\prime}(p_{1})=v_{2}$. Take $\epsilon=\frac{r_{2}}{r_{1}}$. If $|v_{2}|<\epsilon$, then $F$ maps $D_{1}$ into $D_{2}$. Consider a holomorphic retraction map $G:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $G(z,w)=(z,F(z))$. Hence $Z:=\{(z,F(z)):z\in D_{1}\}$ is a 1 dimensional retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(1,v_{2})\in T_{p}Z$. ∎
Proposition 7.2.
Suppose $D_{1},D_{2}$ be domains in $\mathbb{C}^{N},\mathbb{C}^{M}$ respectively and atleast one of domains is Caratheodory hyperbolic, say $D_{1}$; let $p=(p_{1},p_{2})$ be any given point in the product domain $D=D_{1}\times D_{2}$. Let $v_{1}\in T_{p_{1}}D_{1}$ be any non-zero vector. Then there exists $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, and there exists a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ such that $(w_{1},v_{2})\in T_{p}Z$, for some non-zero vector $w_{1}$ with $w_{1}\in span_{\mathbb{C}}\{v_{1}\}$.
Proof.
Since $D_{2}$ is open in $\mathbb{C}^{M}$, there exists a polydisk $\Delta^{M}(p_{2},r_{2})\subset D_{2}$. Since $D_{1}$ is Caratheodory hyperbolic, we have $\gamma_{D_{1}}(p_{1},v_{1})\neq 0$. This means that there exists a non-constant holomorphic map $g:D_{1}\to\Delta$ such that $g(p_{1})=0$ and $g^{\prime}(p_{1})(v_{1})\neq 0$. So $g^{\prime}(p_{1})$ is a non-zero linear map from $\mathbb{C}^{M}$ onto $\mathbb{C}$. Hence there exists $w_{1}\in T_{p_{1}}D_{1}$ with $w_{1}=\lambda v_{1}$ for some $\lambda\in\mathbb{C}$ and $g^{\prime}(p_{1})(w_{1})=1$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $v_{2}\in\Delta^{M}(0,r)$ Consider a holomorphic map $\phi:D_{1}\to\mathbb{C}^{M}$ defined by $\phi(z)=p_{2}+v_{2}g(z)$. Note that
$\|\phi(z)-p_{2}\|_{\infty}=\|v_{2}g(z)\|=|g(z)|\|v_{2}\|_{\infty}\leq 1\cdot\|v_{2}\|_{\infty}$
Write $r_{2}=(r_{2}^{(1)},\ldots,r_{2}^{(M)})$. Choose $\epsilon:=\min\{r_{2}^{(1)},\ldots,r_{2}^{(M)}\}$. If $\|v_{2}\|_{\infty}<\epsilon$, then $\phi$ maps $D_{1}$ into $D_{2}$. Note that $\phi(p_{1})=p_{2}$ and $\phi^{\prime}(p_{1})(w_{1})=v_{2}g^{\prime}(p_{1})(w_{1})=v_{2}$. Define a holomorphic map $F:D_{1}\times D_{2}\to D_{1}\times D_{2}$ by $F(z,w)=(z,\phi(z))$. Again note that $F(p_{1},w)=(p_{1},p_{2})$ and $F^{\prime}(p)(w_{1},w_{2})=(w_{1},\phi^{\prime}(p_{1})(w_{1}))=(w_{1},v_{2})$. Hence $Z:=\{(z,\phi(z)):z\in D_{1}\}$ is a retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(w_{1},v_{2})\in T_{p}Z$, where $w_{1}\in span_{\mathbb{C}}<v_{1}>$. ∎
Next we shall prove proposition 1.6.
Proof of proposition 1.6. Let us begin by considering the set of boundary points of the form $q/h(q)$ for $q$ varying within $D\setminus A$. Denote the subset of such boundary points satisfying the additional condition that $\partial D$ is non-convex at $q/h(q)$, by $S^{h}$. It follows from the hypotheses that $S^{h}$ is a dense subset of $\partial D$ (as is not difficult to see). Next, we observe that it follows from the above lemma 7.1 together with corollary 2.27, there is no one-dimensional retract through 00 in the direction of $p\in S^{h}$. Note that lemma 7.1 is applicable here owing to the fact that bounded balanced pseudoconvex domains are guaranteed to have a Stein neighbourhood basis by theorem 1.17.16 in [Jrnck_Frst_stps].
With these observations, we are ready to prove the proposition by contradiction. Indeed, if $F$ were a biholomorphism mapping $D\setminus A$ onto a product domain $D$, then through $F(0)$ we have a retract along those directions given by proposition 7.2 – note that the hypothesis about Caratheodory hyperbolicity therein is verified owing to the fact that $D\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D\setminus A$ must be Caratheodory hyperbolic as well. In particular, there always exists a non-trivial retract (necessarily of dimension one here) of $D$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$. By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exists a (non-trivial) retract of $D$ through the origin with its tangent direction at 00 belonging to $S^{\prime}$. But this contradicts the hypotheses that $D$ is not convex at an open dense subset of $\partial D$ which implies that for each non-convex boundary point $a$ of $D$, there are no retracts of $D\setminus A$ passing through the origin in the direction of $a$. This finishes the proof. ∎
Remark 7.3.
Though we had assumed in the statement of proposition 1.6 that the dimension $n=2$, the above proof goes through for arbitrary dimension to show atleast that $D$ is a balanced domain of holomorphy in $\mathbb{C}^{N}$, $A$ an analytic subset not passing through origin, then $D\setminus A$ is not biholomorphic to a product domain with a one-dimensional factor; in-particular, it is not completely reducible meaning that $D\setminus A$ is not biholomorphic to a product of one-dimensional domains.
8. Retracts of Unbounded and topologically non-trivial domains – Proofs of theorems 1.151.19
As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allows topologically non-triviality, the boundedness assumption prevailed. To go beyond, to more general settings, we require the following lemma:
Lemma 8.1.
Let $X$ and $Y$ be complex manifolds. If $F:X\times Y\to X\times Y$ is a holomorphic retraction map from $X\times Y$ onto $Z\times\{c\}$ for some $c\in Y$, then $Z$ is a retract of $X$.
Proof.
Write $F(x,y)=(f(x,y),c)$. By comparing the first components of $F\circ F=F$ (the defining equation of a retraction map), we get
(8.1) $f(f(x,y),c)=f(x,y)$
Define $G:X\to Z$ by $G(x)=f(x,c)$. Note that $G(G(x))=G(f(x,c))=f\left(f(x,c),c)\right)$. From (8.1), we get $G(G(x))=f(x,c)=G(x)$. So, $G\circ G=G$ i.e., $G$ is a retraction map from $X$ onto $Z$, thereby $Z$ is a retract of $X$. ∎
We are now ready to detail the proofs of various theorems dealing with unbounded and topologically non-trivial domains (and manifolds) as stated in the introduction, starting with that of theorem (1.16).
Proof of theorem 1.16.
Let $Z$ be a retract of ×$\mathbb{C}\times D$ passing through $p:=(a,c)$. Let $F:\mathbb{C}\times D\to\mathbb{C}\times D$ be a holomorphic retraction map such that ${\rm image}(F)=Z$. Write $F(z,w)=\left(f_{1}(z,w),f_{2}(z,w)\right)$.
We start by considering the second component of $F$. For each fixed $w\in D$, $f^{2}_{w}:\mathbb{C}\to D$ defined by $f^{2}_{w}(z)=f_{2}(z,w)$ is a bounded holomorphic map on $\mathbb{C}$. By Liouville’s theorem, $f^{2}_{w}$ is a constant function and this implies that $f_{2}(z,w)$ is independent of $z$. So $f_{2}(z,w)=g(w)$ for some holomorphic map $g$ on $D$. Note that $g$ is a retraction map on $D$ such that $g(c)=c$. So either $g\equiv c$, or $g\equiv I$, the identity map on $D$ and accordingly, we are led to breaking into $2$ cases as dealt with below.
Case-1: $g\equiv c$. In this case, $F(z,w)=(f_{1}(z,w),c)$. By comparing the first component of $F\circ F=F$, we get $f_{1}(f_{1}(z,w),c)=f_{1}(z,w)$. Note that $F$ is a holomorphic retraction map on ×$\mathbb{C}\times D$ onto $Z_{1}\times\{c\}$, where $Z_{1}:={\rm image}(f_{1})$ on ×$\mathbb{C}\times D$. By lemma 8.1, $Z_{1}$ is a retract of $\mathbb{C}$ passing through $a$. Hence either $Z_{1}=\mathbb{C}$, or $Z_{1}=\{a\}$. It follows that $Z=\mathbb{C}\times\{c\}$ or $Z=\{p\}$.
Case-2: $g\equiv I$, the identity map on $D$. In this case, $F(z,w)=(f_{1}(z,w),w)$. For each fixed $w\in D$, we then consider the holomorphic function $f^{1}_{w}:\mathbb{C}\to\mathbb{C}$ defined by $f^{1}_{w}(z)=f_{1}(z,w)$. The fact that $F$ is a retraction map i.e., $F\circ F=F$ implies in particular that $f_{1}\left(f_{1}(z,w),w\right)=f_{1}(z,w)$. This translates to the following for the function $f^{1}_{w}$:
$f^{1}_{w}\circ f^{1}_{w}(z)=f^{1}_{w}(f_{1}(z,w))=f_{1}(f_{1}(z,w),w)=f_{1}(z,w)=f^{1}_{w}(z).$
In short: $f^{1}_{w}\circ f^{1}_{w}=f^{1}_{w}$, thereby $f^{1}_{w}$ is a retraction map on $\mathbb{C}$. Consequently, either $f^{1}_{w}\equiv I$, the identity map on $\mathbb{C}$, or $f^{1}_{w}$ is a constant function. This means that either $f_{1}(z,w)=z$ or $f_{1}(z,w)$ is a function of $w$ alone, i.e. $f_{1}(z,w)=f(w)$ for some holomorphic function $f$ on $D$. So, we conclude that $Z=\mathbb{C}\times D$, or $Z=\{(f(w),w):w\in D\}$. ∎
Next we shall detail the proof of retraction mappings of ×$\mathbb{C}\times\Delta$.
Proof of theorem 1.18:.
Let $F:\mathbb{C}\times\Delta\to\mathbb{C}\times\Delta$ be a retraction map such that ${\rm image}(F)=Z$ and $F(0)=0$. Write $F(z,w)=(f(z,w),g(z,w))$. For each fixed $w$, $g_{w}:\mathbb{C}\to\Delta$ defined by $g_{w}(z)=g(z,w)$ is a bounded holomorphic map on $\mathbb{C}$. By Liouville’s theorem, $g_{w}$ is a constant function and this implies that $g(z,w)$ is independent of $z$. So $g(z,w)=\tilde{g}(w)$, where $\tilde{g}$ is a holomorphic function on $\Delta$. Note that $\tilde{g}$ is a retraction map on $\Delta$ such that $\tilde{g}(0)=0$. So $\tilde{g}\equiv 0$ or $\tilde{g}\equiv I$, the identity map on $\Delta$.
Case-1: $\tilde{g}\equiv 0$. In this case, $F(z,w)=(f(z,w),0)$. The fact that $F$ is a retraction map i.e. $F\circ F=F$ implies that $f(f(z,w),0)=f(z,w)$. Since $f$ has a single power series representation centred at origin, we can write
$f(z,w)=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots,$
where $P_{i}(z,w)$ is a homogenous polynomial in $z,w$ of degree $i$.
Since $f(f(z,w),0)=f(z,w)$, we get
$f(P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots,0)=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots.$
(8.2) $P_{1}(P_{1}+P_{2}+P_{3}+\ldots,0)+P_{2}(P_{1}+P_{2}+P_{3}+\ldots,0)+P_{3}(P_{1}+P_{2}+P_{3}+\ldots,0)\\ +\ldots=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots.$
Equating degree one terms on both sides, we get $P_{1}(P_{1},0)=P_{1}(z,w)$. Either $P_{1}\equiv 0$ or $P_{1}(z,w)=z+bw$ for some $b\in\mathbb{C}$ and accordingly split into two subcases.
Subcase-(1a): $P_{1}\equiv 0$. From (8.2), we get
$P_{2}(P_{2}+P_{3}+\ldots,0)+P_{3}(P_{2}+P_{3}+\ldots,0)+\ldots=P_{2}(z,w)+P_{3}(z,w)+\ldots.$
The least degree term on left side is $P_{2}(P_{2},0)$, whose degree is $4$ and the least degree term on right side is $P_{2}(z,w)$, whose degree is $2$. So we conclude that $P_{2}\equiv 0$. Continuing in this manner, we get $P_{i}\equiv 0$ for all $i=1,2,\ldots$. Thereby we conclude in this subcase that $F\equiv 0$. Hence $Z=\{(0,0)\}$.
Subcase-(1b): $P_{1}(z,w)=z+bw$. From (8.2), we get
$P_{2}(P_{1}+P_{2}+P_{3}+\ldots,0)+P_{3}(P_{1}+P_{2}+P_{3}+\ldots,0)+\ldots=0.$
Equating degree $2$ terms on both sides, we get $P_{2}(P_{1},0)=0$. This implies that $P_{2}(z,w)=w(Q_{1}(z,w))$, where $Q_{1}(z,w)$ is a polynomial in $z$ and $w$. Substituting $P_{2}$ in 8.2, we get
$P_{3}(P_{1}+P_{2}+P_{3}+\ldots,0)+P_{4}(P_{1}+P_{2}+P_{3}+\ldots)+\ldots=0$
Next, equating degree $3$ terms on both sides, we get $P_{3}(P_{1},0)=0$. This implies that $P_{3}(z,w)=w(Q_{2}(z,w))$, where $Q_{2}(z,w)$ is a polynomial in $z$ and $w$. Continuing in this way, we get $P_{i}(z,w)=w(Q_{i-1}(z,w))$, where $Q_{i-1}$ is a polynomial in $z$ and $w$. Finally $f(z,w)=z+w(h(z,w))$, where $h$ is some holomorphic map on ×$\mathbb{C}\times\Delta$. In this subcase, $F(z,w)=(z+w(h(z,w)),0)$. Hence $Z=\mathbb{C}\times\{0\}$.
Case-2: $\tilde{g}\equiv I$. In this case, $F(z,w)=(f(z,w),w)$. Since $F\circ F=F$, we get $f(f(z,w),w)=f(z,w)$.
Since $f$ has a single power series expansion centred at origin, we can write
$f(z,w)=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots,$
where $P_{i}(z,w)$ is a homogenous polynomial in $z,w$ of degree $i$
Since $f(f(z,w),w)=f(z,w)$, we get
$f(P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots,w)=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots$
(8.3) $P_{1}(P_{1}+P_{2}+P_{3}+\ldots,w)+P_{2}(P_{1}+P_{2}+P_{3}+\ldots,w)+P_{3}(P_{1}+P_{2}+P_{3}+\ldots,w)\\ +\ldots=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots$
Comparing the smallest degree term on both sides, we get $P_{1}(P_{1}(z,w),w)=P_{1}(z,w)$. Write $P_{1}(z,w)=az+bw$ for some $a,b\in\mathbb{C}$. From the above relation, we get $P_{1}(az+bw,w)=az+bw$. This implies that $a(az+bw)+bw=az+bw$. After rearranging we get $a^{2}z+b(a+1)w=az+bw$. Equating the coefficient of $z$ and $w$, we get $a^{2}=a$ and $b(a+1)=b$. From this we can conclude either $a=0$ or $1$. So we divide into two cases.
Subcase-(2a): $a=1$. From the relation $b(a+1)=b$, we get $b=0$. So $P_{1}(z,w)=z$.
From (8.3), we get
$(z+P_{2}+P_{3}+\ldots)+P_{2}(z+P_{2}+P_{3}+\ldots,w)+P_{3}(z+P_{2}+P_{3}+\ldots,w)+\ldots=z+P_{2}+P_{3}+\ldots$
After simplifying, above equation reduces to
$P_{2}(z+P_{2}+P_{3}+\ldots,w)+P_{3}(z+P_{2}+P_{3}+\ldots,w)+\ldots=0.$
Again equating smallest degree on both sides, we get $P_{2}\equiv 0$.
Continuing in this manner, we get $P_{i}\equiv 0$ for $i=1,2,\ldots$. Hence $f_{1}(z,w)=z$.
In this subcase, $F(z,w)=(z,w)$. Hence we get $Z=\mathbb{C}\times\Delta$.
Subcase-(2b): $a=0$. In this subcase, we get $P_{1}(z,w)=bw$.
From (8.3), we get
$P_{2}(bw+P_{2}+P_{3}+\ldots,w)+P_{3}(bw+P_{2}+P_{3}+\ldots,w)+\ldots=P_{2}+P_{3}+\ldots$
Comparing the smallest degree terms, we get $P_{2}(bw,w)=P_{2}(z,w)$. This implies $P_{2}$ is a function of $w$ alone.
(8.3) reduces to
$P_{3}(bw+P_{2}+P_{3}+\ldots,w)+P_{4}(bw+P_{2}+P_{3}+\ldots,w)\ldots=P_{3}+P_{4}+\ldots$
After equating the smallest degree, we get $P_{3}$ is a function of $w$ alone.
Continuing this manner, we get $P_{i}$ is a function of $w$ alone. So $f(z,w)=h(w)$, where $h(w)=P_{1}(w)+P_{2}(w)+P_{3}(w)+\ldots$
In this subcase, $F(z,w)=\left(h(w),w\right)$. Hence we get $Z=\{(h(w),w):w\in\Delta\}$.
Since ×$\mathbb{C}\times\Delta$ is homogenous, upto a conjugation of automorphism, every non-trivial retraction map is of the form $F(z,w)=\left(z+wh(z,w),0\right)$, where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$ or $F(z,w)=(h(w),w)$, where $h\in\mathcal{O}(\Delta)$
Proof of theorem 1.17.
Let $\widetilde{Z}$ be a retract of $(\mathbb{C}\setminus A)\times D$ passing through an arbitrary point $(a,b)\in(\mathbb{C}\setminus A)\times D$. Let $F:(\mathbb{C}\setminus A)\times D\to(\mathbb{C}\setminus A)\times D$ be a retraction map such that ${\rm image}(F)=\widetilde{Z}$. Write $F(z,w)=(f(z,w),g(z,w))$. For each fixed $w\in D$, $g_{w}:(\mathbb{C}\setminus A)\to D$ defined by $g_{w}(z)=g(z,w)$ is a bounded holomorphic map on $(\mathbb{C}\setminus A)$. Applying Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each component of $g_{w}$, $g_{w}$ can extend holomorphically as a map $g_{w}^{\prime}:\mathbb{C}\to\overline{D}$. By Liouville’s theorem, $g_{w}^{\prime}$ is a constant function. In particular, $g_{w}$ is constant function and thus implies that $g(z,w)$ is independent of $z$. So $g(z,w)=\tilde{g}(w)$ for some holomorphic map $\tilde{g}$ on $D$. From the definition of retraction map, we get $F\circ F=F$. This implies in particular that $f\left(f(z,w),\tilde{g}(w)\right)=f(z,w)$ and $\tilde{g}(\tilde{g}(w))=\tilde{g}$. Since $F(a,b)=(a,b)$, we get $\tilde{g}(b)=b$. Note that $\tilde{g}$ is a retraction map on $D$ such that $\tilde{g}(b)=b$. For each fixed $w\in D$, we get a holomorphic map $f_{w}:(\mathbb{C}\setminus A)\to(\mathbb{C}\setminus A)$ defined by $f_{w}(z)=f(z,\tilde{g}(w))$. Note that
$f_{w}\circ f_{w}(z)=f_{w}\left(f(z,\tilde{g}(w))\right)=f\left(f(z,\tilde{g}(w)),\tilde{g}(\tilde{g}(w)\right)=f(z,\tilde{g}(w))=f_{w}(z).$
This shows that $f_{w}$ is a retraction map on $(\mathbb{C}\setminus A)$. Since every domain in $\mathbb{C}$, the only retraction maps are identity and constant function, either $f_{w}\equiv I$, the identity map on $(\mathbb{C}\setminus A)$ or $f_{w}\equiv$ constant function (depends on $w$), hence imples that $f_{w}$ is independent of $z$. Hence for all $z,w\in(\mathbb{C}\setminus A)\times D$, we have
(8.4) $f(z,\tilde{g}(w))=H(\tilde{g}(w))\quad\text{for some holomorphic function}~{}H~{}\text{on}~{}D$
or
(8.5) $f(z,\tilde{g}(w))=z.$
Define $Z:={\rm image}(\tilde{g})$. According to (8.4) and (8.5), we divide into $2$ cases.
Case-1: $f(z,\tilde{g}(w))=H(\tilde{g}(w))$ for all $(z,w)\in(\mathbb{C}\setminus A)\times D$
We claim that $\widetilde{Z}=\{(H(w),w):w\in Z\}$. To prove this claim, observe that for each $w\in Z$, we have $\tilde{g}(w)=w$ and $f(H(w),w)=f(H(w),\tilde{g}(w))=H(w)$ This implies that $\{(H(w),w):w\in Z\}\subset\widetilde{Z}$. For attaining converse implication, let $(z_{0},w_{0})$ be an arbitrarily point in $\widetilde{Z}$. Since $\tilde{g}(w_{0})=w_{0}$ and $f(z_{0},w_{0})=z_{0}$ Hence we get $z_{0}=f(z_{0},w_{0})=f(z_{0},\tilde{g}(w_{0}))=H(\tilde{g}(w_{0}))=H(w_{0})$. This implies that $(z_{0},w_{0})=\left(H(w_{0}),w_{0}\right)$. Hence $\widetilde{Z}\subset\{\left(H(w),w\right):w\in Z\}$, and this proves the claim.
Case-2: $f(z,\tilde{g}(w))=z$ for all $(z,w)\in(\mathbb{C}\setminus A)\times D.$
In this case, we claim that $\widetilde{Z}=(\mathbb{C}\setminus A)\times Z$. To prove this claim, note that each $(z_{0},w_{0})\in(\mathbb{C}\setminus A)\times Z$, $\tilde{g}(w_{0})=w_{0}$ and $f(z_{0},\tilde{g}(w_{0}))=f(z_{0},w_{0})=z_{0}$. Hence $(\mathbb{C}\setminus A)\times Z\subset\widetilde{Z}$. Conversely assume that $(z_{0},w_{0})$ be an arbitrary point in $\widetilde{Z}$. Note that $\widetilde{g}(w_{0})=w_{0}\in Z$ and $f(z_{0},w_{0})=z_{0}$. This implies that $(z_{0},w_{0})\in(\mathbb{C}\setminus A)\times Z$, which proves our claim.
Finally we get that every retract $\widetilde{Z}$ is of the form
$\widetilde{Z}=\{(f(w),w):w\in Z\}$
where $Z$ is a retract of $D$ and $f$ is a holomorphic map on $Z$ or, of the form $(\mathbb{C}\setminus A)\times Z$.
We shall next record the proof of theorem 1.16, whose verifying details will be shortened as most of its arguments are contained in the prequel. To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of ×$\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$.
Proof of theorem 1.19:.
(i)  All arguments used in the proof of theorem 1.16 holds for this case too except for proving $g_{w}$ is independent of $z$. So, we have only to prove that $g_{w}$ is independent of $z$.
Let $Z$ be a retract of $X\times Y$ passing through $p:=(a,c)$, where $(a,c)\in X\times Y$. Let $F:X\times Y\to X\times Y$ be a holomorphic retraction map such that ${\rm image}(F)=Z$. Write $F(z,w)=(f(z,w),g(z,w))$ in standard component notations.
For each fixed $w\in Y$, $g_{w}:X\to Y$ defined by $g_{w}(z)=g(z,w)$ is a holomorphic map from a compact Riemann surface $X$ to non-compact Riemann surface $Y$. If $g_{w}$ is non-constant map, then by open mapping theorem, $g_{w}(X)$ is open and by continuity of $g_{w}$, $g_{w}(X)$ is compact, hence closed. Since, $X$ is connected, $g_{w}$ is surjective. This does not holds because $Y$ is non-compact. Hence we conclude that $g_{w}$ is a constant function and this implies that $g(z,w)$ is independent of $z$. So $g(z,w)=\tilde{g}(w)$ for some holomorphic map $\tilde{g}$ on $Y$.
(ii)  The proof for this case follows similar line of arguments used in the proof of proposition 1.17. So, we have only to prove that $g_{w}$ is independent of $z$.
Let $\widetilde{Z}$ be a retract of $X\times Y$ passing through $p:=(a,c)$, where $(a,c)\in X\times Y$. Let $F:X\times Y\to X\times Y$ be a holomorphic retraction map such that ${\rm image}(F)=\widetilde{Z}$. Write $F(z,w)=(f(z,w),g(z,w))$. For each fixed $w\in Y$, consider then the map $g_{w}:X\to Y$ defined by $g_{w}(z)=g(z,w)$, which is a holomorphic map from a non-hyperbolic Riemann surface $X$ to hyperbolic complex manifold $Y$. Let $d_{X},d_{Y}$ denotes Kobayashi hyperbolic pseudo-distance on $X$ and $Y$ respectively. By the contractive property of hyperbolic pseudodistance under holomorphic maps, we get $d_{Y}(g_{w}(x_{1}),g_{w}(x_{2}))\leq d_{X}(x_{1},x_{2})$. Since $X$ is non-hyperbolic Riemann surface, $d_{X}\equiv 0$. By hyperbolicity of $Y$, we get $g_{w}(x_{1})=g_{w}(x_{2})$. Since this is true for every $x_{1},x_{2}\in X$, we get $g_{w}$ is a constant function for each fixed $w\in Y$. This implies that $g(z,w)$ is independent of $z$. ∎
We shall now detail the proof of theorem 1.15 about retracts of the Hartogs triangle and its analytic complements. Making the identification $\Omega_{H}\simeq\Delta\times\Delta^{*}$, retracts of the Hartogs triangle can be deduced as a special corollary of lemma 7.1 taking $D$ to be just the bidisc $\Delta^{2}$. However, for analytic complements $\Omega_{H}\setminus A$ this cannot be done merely by citing lemma 7.1 (at the very least, directly!), as the Hartogs triangle $\Omega_{H}$ does not admit a Stein neighbourhood basis. We therefore prefer to give a direct proof as follows.
Proof of theorem 1.15.
Let $F$ be a non-constant holomorphic retraction map on $(\Delta\times\Delta^{*})\setminus A$ such that $Z={\rm image}(F)$. Write $F(z,w)=\left(F_{1}(z,w),F_{2}(z,w)\right)$, where $F_{1}:(\Delta\times\Delta^{*})\setminus A\to\Delta$ and $F_{2}:(\Delta\times\Delta^{*})\setminus A\to\Delta^{*}$ are holomorphic functions. Since $A$ is a proper analytic subset of $\Delta\times\Delta^{*}$ and by Riemann removable singularities theorem, $F_{1},F_{2}$ can be extended holomorphically as a map $\widetilde{F}_{1},\widetilde{F}_{2}:\Delta\times\Delta^{*}\to\overline{\Delta}$ respectively. Write $\Delta\times\Delta^{*}=\Delta^{2}\setminus A_{0}$, where $A_{0}=\{(z,w)\in\Delta^{2}:w=0\}$. Note that $A_{0}$ is a proper analytic subset of $\Delta^{2}$. Applying the same argument, $\widetilde{F}_{1},\widetilde{F}_{2}$ can be extended to a holomorphic map from $\Delta^{2}$ to $\overline{\Delta}$. By abuse of notation, $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorphic maps from $\Delta^{2}$ to $\Delta$. We now observe that the images of these maps actually does not contain any point from $\partial\Delta$ i.e., they map into $\Delta$; this follows by the open mapping theorem (non-constant holomorphic functions from domains in $\mathbb{C}^{N}$ are open maps into $\mathbb{C}$;). Note that atleast one of $\widetilde{F}_{1}$ or $\widetilde{F}_{2}$ are non-constant! So, $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorpic maps from $\Delta^{2}$ to $\Delta$. Thereby, $F$ extends holomorphically as a map $\widetilde{F}:\Delta^{2}\to\Delta^{2}$; elaborately for the sequel, $\widetilde{F}(z,w)=\left(\widetilde{F}_{1}(z,w),\widetilde{F}_{2}(z,w)\right)$ to $\Delta^{2}$.
Next we want to prove that $\widetilde{F}$ is a retraction map on $\Delta^{2}$. By hypothesis, $F\circ F\equiv F$ on $(\Delta\times\Delta^{*})\setminus A$. Let $(z_{0},w_{0})\in A$ be an arbitrary point. Pick any sequence $(z_{n},w_{n})\in(\Delta\times\Delta^{*})\setminus A$ converging to $(z_{0},w_{0})$. So, $z_{n}\to z_{0}$ and $w_{n}\to w_{0}$. Since $(z_{n},w_{n})\in(\Delta\times\Delta^{*})\setminus A$ and $\widetilde{F}\equiv F$ on $(\Delta\times\Delta^{*})\setminus A$, we have $\widetilde{F}\circ\widetilde{F}(z_{n},w_{n})=\widetilde{F}(z_{n},w_{n})$. By continuity of $\widetilde{F}$, as $n\to\infty$ we get $\widetilde{F}\circ\widetilde{F}(z_{0},w_{0})=\widetilde{F}(z_{0},w_{0})$. Since $(z_{0},w_{0})$ is an arbitrary point, we get the relation $\widetilde{F}\circ\widetilde{F}\equiv\widetilde{F}$ on $\Delta^{2}$. This proves that $\widetilde{F}$ is a holomorphic retraction on $\Delta^{2}$.
Let $\widetilde{Z}:={\rm image}(\widetilde{F})$. By theorem 3 in [Heath_Suff], $\widetilde{Z}$ is either of the form $\{(z,f(z)):z\in\Delta\}$ for some holomorphic self map $f$ on $\Delta$ or $\{(g(w),w):w\in\Delta\}$ for some holomorphic self map $g$ on $\Delta$. Since $Z\subset(\Delta\times\Delta^{*})\setminus A$ and $Z\subset\widetilde{Z}$ ($\widetilde{F}$ is a holomorphic extension of $F$), we get $Z$ is either of the form $\{(z,f(z)):z\in\Delta\}\cap((\Delta\times\Delta^{*})\setminus A)$ for some holomorphic map $f:\Delta\to\Delta$ or $\{(g(w),w):w\in\Delta^{*}\}\cap((\Delta\times\Delta^{*})\setminus A)$ for some holomorphic map $g:\Delta\to\Delta$. ∎
9. Retracts of $\mathbb{C}^{2}$; proof of theorem 1.20
We first give an example of a holomorphic retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ which can neither be realized as the graph over the first factor nor the second. This is discussed in the following
Proposition 9.1.
The one-dimensional complex submanifold of $\mathbb{C}^{2}$ given by
$Z:=\{(e^{z},ze^{-z}):z\in\mathbb{C}\}$
is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},zw~{}e^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
Proof.
It is easily first noted that $Z$ cannot be written as $\{(z,g(z)):z\in\mathbb{C}\}$ for some entire function $g$ on $\mathbb{C}$ because the first projection $\pi_{1}(z,g(z))=z$ maps onto $\mathbb{C}$ whereas $\pi_{1}(e^{z},ze^{-z})=e^{z}$ is not onto. Next, we shall establish that $Z$ cannot be written in the form $\{(g(w),w):w\in\mathbb{C}\}$ either. To prove this by contradiction, suppose $Z=\{(g(w),w):w\in\mathbb{C}\}$ for some entire function $g$. This implies that $g$ satisfies $e^{wg(w)}=g(w)$. Hence we get $(e^{g(w)})^{w}=g(w)$. Therefore, if we define $h:\mathbb{C}\to\mathbb{C}^{*}$ by $h(w)=e^{g(w)}$, it follows that: $h$ satisfies $(h(w))^{w}=g(w)$. Note that $h~{}\text{and}~{}g$ are surjective holomorphic functions from $\mathbb{C}$ to $\mathbb{C}^{*}$.
Let $L_{1}$ be an arbitrary ray emanating from origin. For each $z\in\mathbb{C}$, there exists an open disk $D$ such that $D\subset\mathbb{C}\setminus L_{1}$. Let $\ell_{1}$ denotes the branch of logarithm on $\mathbb{C}\setminus L_{1}$. Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ non-vacuously. $L_{2}$ divides $D$ into two regions namely $D^{+}~{}\text{and}~{}D^{-}$. Let $\ell_{2}$ denotes the branch of logarithm on $\mathbb{C}\setminus L_{2}$.
Since $h(w)^{w}=g(w)$ for all $w\in\mathbb{C}$, $w\ell_{1}(h(w))=\ell_{1}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$ and this implies that $h(w)=e^{\frac{\ell_{1}(g(w))}{w}}$. Similarly, we get $h(w)=e^{\frac{\ell_{2}(g(w))}{w}}~{}\text{for all}~{}w\in D^{+}\cup D^{-}$.
On $D^{+}\cup D^{-}$,
$e^{\frac{\ell_{1}(g(w))}{w}}=e^{\frac{\ell_{2}(g(w))}{w}}$
This implies that for all $w\in D^{+}\cup D^{-}$, we have
(9.1) $\ell_{1}(g(w))-\ell_{2}(g(w))=2n\pi iw~{}$
But $\ell_{1}(g(w))-\ell_{2}(g(w))$ is an integer multiple of $2\pi i$ for any $w\in D^{+}\cup D^{-}$. So, (9.1) holds only for $n=0$. Hence we get $\ell_{1}(g(w))=\ell_{2}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$. This implies that $\ell_{2}$ is a branch of logarithm on $\mathbb{C}^{*}$. This contradicts the fact that there does not exists a branch of logarithm on $\mathbb{C}^{*}$. ∎
We conclude by reporting some results about polynomial retracts of $\mathbb{C}^{2}$, the simplest of domains admitting non-trivial retracts. It seems difficult to get a complete characterization of holomorphic retracts of $\mathbb{C}^{2}$. However it is possible to deduce a complete characterization of polynomial retracts on $\mathbb{C}^{2}$ from known results about the polynomial ring in complex two variables. Even this deduction is not entirely trivial and needs some work which we record here. In order to lay this work out properly, we begin with some definitions.
Definition 9.2.
A holomorphic map $\beta:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is said to be a polynomial retraction if both components of $\beta$ are polynomials and satisfy $\beta\circ\beta=\beta$.
Definition 9.3 ([Shplrn]).
Let $K[x,y]$ be the polynomial algebra in two variables over a field $K$ of characteristic 00. A subalgebra $R$ of $K[x,y]$ is called a retract if it satisfies any of the following equivalent conditions:
  1. (R1)
    There is an idempotent homomorphism (a retraction, or projection) $\varphi:K[x,y]\rightarrow$ $K[x,y]$ such that $\varphi(K[x,y])=R$.
  2. (R2)
    There is a homomorphism $\varphi:K[x,y]\rightarrow R$ that fixes every element of $R$.
  3. (R3)
    $K[x,y]=R\oplus I$ for some ideal $I$ of the algebra $K[x,y]$.
  4. (R4)
    $K[x,y]$ is a projective extension of $R$ in the category of $K$-algebras. In other words, there is a splitting exact sequence $1\rightarrow I\rightarrow K[x,y]\rightarrow R\rightarrow 1$, where $I$ is the same ideal as in (R3) above.
To provide just one concrete example to illustrate the above definitions, take $p=x+x^{2}y$; then $K[p]$ is a retract of $K[x,y]$ (the retraction map here, is given by $x\mapsto p$ , $y\mapsto 0$).
In [Cost], Costa proved that every proper retract of $K[x,y]$ has the form $K[p]$ for some polynomial $p\in K[x,y]$.
Next we shall show how polynomial retraction maps on $\mathbb{C}^{2}$ transforms to a retraction map on the polynomial ring $\mathbb{C}[z,w]$. We can achieve this using the following lemma.
Lemma 9.4.
A holomorphic map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\rho(z,w)=(p(z,w),q(z,w))$, is a polynomial retraction if and only if there exists a retraction map $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ whose components are polynomials $p(z,w)$ and $q(z,w)$.
Proof.
Assume that $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is a retraction map on $\mathbb{C}^{2}$ such that whose components are polynomials namely $p(z,w)$ and $q(z,w)$. Define $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ defined by $z\mapsto p(z,w),~{}w\mapsto q(z,w)$. Since $\rho\circ\rho=\rho$, we get $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$. This imply that $\tilde{\rho}\circ\tilde{\rho}=\tilde{\rho}$. So, we get that $\tilde{\rho}$ is a retraction on $\mathbb{C}[z,w]$ whose image is precisely $R$, the subalgebra generated by $p,q\in\mathbb{C}[z,w]$.
Conversely, assume $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ is a retraction map. Then there exists $p(z,w),\\ q(z,w)\in\mathbb{C}[z,w]$ such that $\rho(z)=p(z,w)$ and $\rho(w)=q(z,w)$. Note that the retract $R$ is the subalgebra generated by $p(z,w),~{}q(z,w)$.
Define $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\rho(z,w)=(p(z,w),q(z,w))$. Since $p,q\in R$, $\tilde{\rho}(p)=p$ and $\tilde{\rho}(q)=q$. This implies that $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$. Hence we get that $\rho$ is a polynomial retraction on $\mathbb{C}^{2}$.
9.1. Proof of theorem 1.20
Let $Z$ be a non-trivial polynomial retract of $\mathbb{C}^{2}$. There exists a retraction map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is of the form $\rho(z,w)=(p(z,w),q(z,w))$, where $p,q\in\mathbb{C}[z,w]$ such that $\rho(\mathbb{C}^{2})=Z$. By lemma 9.4, $\rho$ induces a retraction map $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ such that $\tilde{\rho(z)}=p(z,w)$ and $\tilde{\rho}(w)=q(z,w))$, where $p,q\in\mathbb{C}[z,w]$. Note that ${\rm image}(\tilde{\rho})=\mathbb{C}[p,q]$, the subalgebra generated by $p,q$.
By Proposition 3.1 in [Cost], $\mathbb{C}[p,q]=\mathbb{C}[r]$, where $r=r(z,w)\in\mathbb{C}[z,w]$. Write $p=q_{1}(r),~{}q=q_{2}(r)$, where $q_{1},q_{2}\in\mathbb{C}[z]$ So $\rho(z,w)=(q_{1}(r(z,w)),q_{2}(r(z,w)))$.
Suppose $q_{1}$ and $q_{2}$ have degree $n\geq 1$ and $m\geq 1$ respectively. By Abhyankar - Moh theorem in [Abynkr], either $n$ divides $m$ or $m$ divides $n$. Suppose $deg(q_{1})=k\cdot deg(q_{2})$ for some integer $k\geq 1$.
Consider an automorphism $\psi:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\psi(z,w)=(z-cw^{k},w)$, where $c$ is chosen so that leading term of polynomial $q_{1}-cq_{2}^{k}$ cancel out.
$\psi\circ\rho\circ\psi^{-1}(z,w)$ $=$ $\psi\circ\rho(z+cw^{k},w)=\psi(q_{1}(r(z+cw^{k},w)),q_{2}(r(z+cw^{k},w)))$
$=$ $(q_{1}(r(z+cw^{k},w))-cq_{2}^{k}(r(z+cw^{k},w)),q_{2}(r(z+cw^{k}.w)))$
$=$ $(\tilde{q_{1}}(\tilde{r}(z,w)),\tilde{q_{2}}(\tilde{r}(z,w)))$
where $\tilde{q_{1}}=q_{1}-cq_{2}^{k},~{}\tilde{q_{2}}=q_{2},~{}\tilde{r}(z,w)=r(z+cw^{k},w)$. Note that $\tilde{q_{1}}(\tilde{r}),~{}\tilde{q_{2}}(\tilde{r})$ generates $\mathbb{C}[r]$.
Continuing this process, we eventually arrive at a pair of polynomials. One of which is a constant polynomial $c$, and the other polynomial which is denoted by $h$, then we must have $\mathbb{C}[h]=\mathbb{C}[r]$ (i.e. $h=c_{1}r+c_{2}$). Finally we get $\phi\in Aut(\mathbb{C}^{2})$ such that
$\phi\circ\rho\circ\phi^{-1}(z,w)=(c_{1}r(z,w)+c_{2},c).$
After applying conjugation of affine automorphism, we get $\phi(z,w)$ is of the form $(s(z,w),0)$, where $s(z,w)\in\mathbb{C}[z,w]$ and $s(0,0)=0$. Since $\phi\circ\rho\circ\phi^{-1}$ is a retraction map, we will get $s(s(z,w),0)=s(z,w)$. Write $s(z,0)=a_{1}z+a_{2}z^{2}+\ldots+a_{k}z^{k}$. Note that
$s(s(z,w),0)=a_{1}s(z,w)+\ldots+a_{k}s(z,w)^{k}=s(z,w).$
This implies $a_{1}=1$ and $a_{2}=...=a_{k}=0$. So $s(z,w)=z+w\cdot l(z,w)$, where $l(z,w)\in\mathbb{C}[z,w]$. Hence $Z$ is of the form $\{(\zeta,0):\zeta\in\mathbb{C}\}$.

### v6

@@
 If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics; the necessity of the boundary points to be extreme points is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well.
-This revisit of Vesentini’s result leads to a small modification that is useful to provide interesting examples.
+This revisit of Vesentini’s result leads to a small modification that is useful to provide interesting examples; specifically, in completely determining the retracts through the origin of certain balanced domains with analytic varieties in their boundaries.
+The converse of Vesentini’s theorem is established and general results characterizing the non-linear retracts through the origin are attained (for fairly general domains in this context) when the boundary fails to have extreme points.
 As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ which is submersive at its center, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$.
-To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of strictly convex balls, with respect to various norms and obtain a complete description of retracts passing through the origin, generalizing that of the polydisc due to Heath and Suffrdige and giving a simpler proof as well.
+To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of strictly convex balls, with respect to various norms and obtain a complete description of retracts passing through the origin, generalizing that of the polydisc due to Heath and Suffridge and giving a simpler proof as well.
 The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$.
 Along the way, we discuss examples of non-convex but pseudoconvex bounded balanced domains in which most one-dimensional linear subspaces fail to be holomorphic retracts, in good contrast to the convex case.
-Indeed, for the simplest possible case here namely the ‘$\ell^{q}$-ball’ for $10<q<$ 1, we obtain a complete list of all retracts through the origin; this enables an illustration of applying retracts to establishing biholomorphic inequivalences.
+Indeed, for the simplest possible case here namely the ‘$\ell^{q}$-ball’ for $10<q<$ 1, we obtain a complete list of all retracts through the origin of each possible dimension; this enables an illustration of applying retracts to establishing biholomorphic inequivalences.
 To go beyond balanced domains, we then first obtain a complete characterization of retracts of the Hartogs triangle and ‘analytic complements’ thereof.
 Thereafter, we drop the assumption of boundedness as well, to attain similar characterization results for domains which are neither bounded nor topologically trivial.
-We conclude by reporting some results on the retracts of $\mathbb{C}^{2}$.
+We conclude with some expositions about retracts of $\mathbb{C}^{2}$.
 Key words and phrases:
@@
 This will also simultaneously help us set up notations and fix terminologies.
+We emphasize that this article is a survey in parts, striving to give references for the known results (to the best of our knowledge).
 By a retraction we shall mean an idempotent map, whereas by a retract, we mean the image of such a map.
@@
 In other words, retractions could be viewed as being, loosely speaking, at the ‘opposite end of the spectrum’ when compared to proper maps.
-While much work has been carried out on proper holomorphic maps and their images, far less has investigation has been done about retracts.
+While much work has been carried out on proper holomorphic maps and their images, far less investigation has been done about retracts.
 Infact, in addition to (the foregoing remarks and) their role of in certain fundamental mapping problems above, retracts are of significance in complex dynamics as well.
@@
 In light of the the stark contrasts between proper maps and retraction maps mentioned above, the connection amongst these maps, is thereby all the more interesting.
-Theorem 1.1 (cf. Bedford, [Bdfrd]; Abate, [Abate_Horosphre]).
+Theorem 1.1 (Bedford, [Bdfrd]; Abate, [Abate_Horosphre]).
 Let $X$ be a taut manifold, $f$ a holomorphic self-map of $X$.
@@
 This itself is noteworthy; if not for anything else, this suffices for an important application to attain (employing the above theorem 1.1), the fairly recent result of [JJ] together with [Pranav_JJ] that: every proper holomorphic self-map of any smoothly bounded balanced pseudoconvex domain of finite type is an automorphism.
-In-fact our methods help simplify a part of the proof of this result which we shall not however elaborate here.
+We shall not however elaborate upon this here.
 As mentioned in the abstract, we intend a systematic presentation; so, to not be amiss about addressing the simplest of questions in the aforementioned first goal then, we recall that in [Vesntn], E.
@@
 We shall to begin with, give a relatively elementary proof of this result without using complex geodesics.
-It may be noted that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal and thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$ but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal while it is indeed $\mathbb{C}$-extremal.
+It may be noted that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal; thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$, but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal but indeed $\mathbb{C}$-extremal.
 The notions of $\mathbb{R}$-extremality, $\mathbb{C}$-extremality and the relationship between them is recalled in the following section 2 about preliminaries.
+The necessity of such extremality hypotheses is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well.
 Theorem 1.2 (Vesentini).
@@
 Now, the condition of convexity on $D$, which entails a certain condition to be satisfied by the entire boundary $\partial D$, can be dropped in various ways.
-To begin with here, let us only remark that we can do so in the following manner, which will actually be useful in dealing with certain concrete (counter-)examples in the sequel.
-Specifically, we shall show the following (without using ‘complex geodesics’).
+To begin with here, let us only remark that we can do so in the following manner; this will actually be useful in dealing with certain concrete examples in the sequel (example 8.7 in section 8).
+This is one another major reason for in revisiting this result of Vesentini.
+Specifically, we shall show the following without using complex geodesics as alluded to earlier.
 Theorem 1.3.
 (A small modification of Vesentini’s result): Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
-Suppose $Z$ is a retract through the origin with the property that $D\cap T_{0}Z$ is a convex domain in $T_{0}Z$ with $\mathbb{C}$-extremal boundary, then $Z$ is a linear subspace of $D$.
+Suppose $Z$ is a retract through the origin with the property that: $D\cap T_{0}Z$ is convex with points of $\partial D\cap T_{0}Z$ being $\mathbb{C}$-extremal boundary points of $D$; then $Z$ is a linear subspace of $D$.
 In particular, every $1$-dimensional retract through the origin of a bounded balanced pseudoconvex domain $D$ with $\mathbb{C}$-extremal boundary, is linear.
-We emphasize that this result is not really new and our inclusion of this is towards a relatively self-contained exposition, and more importantly, in the proof of proposition 5.1 below.
-Indeed, we would not like to be amiss to observe associated facts which are not explicitly recorded in the literature but can be attained by reading proofs of results carefully (and such observations are actually useful for us much later in the sequel): for instance, the apparently strong condition (when $\dim(Z)\geq 2$) on the convexity of $D\cap T_{0}Z$ in the above theorem can be dropped while attaining the same conclusion; specifically, we have the following.
+Our inclusion of this result first towards a relatively self-contained exposition, has other more important motivations than just a more elementary proof: to furnish examples of balanced non-convex (but pseudoconvex) domains possessing non-trivial analytic varieties in the boundary, wherein the retracts through the center can be completely determined.
+This is accomplished in section 8 wherein the above modified version of Vesentini’s result has also a role to play in other propositions there.
+Also in general, we would not like to be amiss to observe associated facts which are not explicitly recorded in the literature but can be attained by reading proofs of results carefully – these are mostly exposited in the ‘preliminaries’ section; more importantly, such observations are actually useful for us much later in the sequel.
+While we relegate such observations to later, one that we think must be mentioned right away here, is the affirmative answer to the question about the very existence of atleast one retract through the origin of an arbitrary bounded balanced domain of holomorphy without any further assumptions.
+This is verified in proposition 2.35.
+Getting back to the above theorem let us now mention for instance, the apparently strong condition (when $\dim(Z)\geq 2$) on the convexity of $D\cap T_{0}Z$ in the above theorem can be dropped while attaining the same conclusion; specifically, we have the following.
 Proposition 1.4.
-(A small modification of proposition $4.5$ of [JJ]).
-Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
-Suppose $Z$ is a retract through the origin with the property that the boundary points of $D:=D\cap T_{0}Z$ viewed as a domain in $T_{0}Z$ are holomorphically extreme.
+(A small modification of proposition $4.5$ of [JJ]): Let $D$ be any balanced pseudoconvex domain in $\mathbb{C}^{N}$.
+Suppose $D\cap T_{0}Z$ is bounded and $Z$ is a retract through the origin with the property that all points of $\partial D\cap T_{0}Z$ are holomorphically extreme boundary points of $D$.
 Then, $Z$ is linear.
-Amongst other results about retracts, this helps in establishing biholomorphic inequivalences, particularly with product domains.
-Stated differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., biholomorphic to a product domain.
+Given this, it may seem questionable as to why we have stated theorem 1.3 in this introductory section; in this regard let us highlight the existence of domains as already alluded to, whose boundaries are not holomorphically extreme but are $\mathbb{C}$-extremal.
+Also, $\mathbb{C}$-extremality is far easier to verify than holomorphic extremality which apparently is a more restrictive condition.
+Indeed, a concrete example of a domain of the kind just mentioned is recorded in proposition 8.8 wherein the domain also serves to show that the converse of the above proposition is false.
+As for the significance of the above proposition 1.4 itself, let us only remark that it will be employed in determining retracts through the origin of all possible dimensions in $\ell_{q}$-‘balls’ as in theorem 1.9 below.
+It also helps in establishing biholomorphic inequivalences, particularly with product domains; for instance, a domain as in the above proposition with holomorphically extreme boundary and the polydisc – this is only to give a quick example here.
+To get new results, we must consider domains with greater complexity, to proceed systematically towards which we deal first with a simple class of domains which allow analytic varieties in their boundaries.
+Stating the essence here differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., biholomorphic to a product domain.
 Let us only indicate the key idea here namely, the fundamental common feature amongst all product domains that: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed.
@@
 Then, $D\setminus A$ is not biholomorphic to a product domain.
+Domains as in the above proposition may be regarded as being maximally non-convex in the realm of bounded balanced domains, remembering that it is impossible to have a bounded domain which is non-convex throughout the boundary i.e., at every boundary point.
+While convex domains is a much studied class, the other extreme has by far less been; this proposition contributes towards that.
 An extension of this to higher dimensions is briefly discussed in section 8.
@@
 Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]).
-Indeed, the question of precisely which linear subspaces are realizable as retracts of $D$ as in the above theorem, is quite an interesting question and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections (when we measure the norm of the projection with respect to the operator norm induced by $\mu$)?
+The question of precisely which linear subspaces are realizable as retracts of $D$ as in the above theorem is significant for us.
+For instance: while every retract of the polydisc $\Delta^{N}$ is a graph, not every graph in $\Delta^{N}$ is a retract; only those which are graphs on linear retracts of $\Delta^{N}$ are, making it important to precisely pin down which linear subspaces are retracts.
+Indeed, this is quite an interesting question in its own right and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections (when we measure the norm of the projection with respect to the operator norm induced by $\mu$)?
 The answer turns out to vary even among the $l^{p}$ norms i.e., depend upon $p$ when we take $\mu=\|\cdot\|_{l^{p}}$ even if we exclude the case $p=2$, which may not be surprising at all for this case is distinguished as the one that arises from an inner product; specifically, the case $p=\infty$ is also different from the other cases.
@@
 As our interest lies in retracts, we shall phrase in it in the language of retracts rather than norm-one projections.
-Specifically, the characterization of linear retracts of the standard $l^{p}$ ball ($1\leq p<\infty,~{}p\neq 2$) can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
+Specifically, the characterization of linear (thereby holomorphic, owing to Vesentini’s theorem) retracts of the standard $l^{p}$ ball ($1\leq p<\infty,~{}p\neq 2$) can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
 Theorem 1.7.
 (Theorem 2.a.4 of [Lindstraus]): Let $Y$ be a linear subspace of $\mathbb{C}^{N}$ equipped with $\|\cdot\|_{l^{p}}$ where $1\leq p<\infty,p\neq 2$.
-Then $Y\cap B_{\ell^{p}}$ is a linear retract of $B_{\ell^{p}}$ iff there exist vectors $\{u_{j}\}_{j=1}^{m}$ of norm 1 in $\mathbb{C}^{N}$ of the form
+Then $Y\cap B_{\ell^{p}}$ is a linear retract of $B_{\ell^{p}}$ if and only if there exist vectors $\{u_{j}\}_{j=1}^{m}$ of $\ell^{p}$-norm 1 in $\mathbb{C}^{N}$ of the form
 $u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i},~{}1\leq j\leq m,~{}with~{}\sigma_{j}\cap\sigma_{k}=\phi~{}for~{}k\neq j$
-such that $Y=span\{u_{j}\}_{j=1}^{m}$, where $\sigma_{j}:=supp(u_{j})=\{1\leq i\leq N:u_{j}(i)\neq 0\}$ (and $\lambda_{i}$’s are some complex numbers).
-The case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, in which all linear subspaces are retracts.
+such that $Y=span\{u_{j}\}_{j=1}^{m}$, where $\sigma_{j}:=supp(u_{j})=\{1\leq i\leq N:u_{j}(i)\neq 0\}$ (and $\lambda_{i}$’s are some complex numbers such that the $\ell^{p}$-norm of $u_{j}$ are equal to $1$).
+Here $e_{i}$ denotes the standard basis of $\mathbb{C}^{N}$.
+We shall discuss alternative formulations of the above result which are more geometrically revealing in section 5.
+In-particular, retracts of $\ell^{p}$-balls are isometric copies of lower dimensional standard $\ell^{p}$-balls.
+Next, we note that the case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, in which all linear subspaces are retracts.
 We would rather like to draw attention to the more important fact that this result is not true for $p=\infty$ to demonstrate which (by a counterexample as simple as possible), we include an example in section 3.
@@
     For each $m\in\{1,2,\ldots,N\}\setminus J$, the scalars $c_{m,j}$, satisfy $\sum_{j\in J}|c_{m,j}|\leq 1$.
-In particular, not every two dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the $\mathbb{B}^{N}$.
+It is immediate that through any given point of the polydisc, there exist retracts of all possible dimensions ($k=0,1,\ldots,N$), as its automorphism group acts transitively; a little contemplation shows that the transitive action is unnecessary to confirm the existence of such retracts.
+Indeed, for $1\leq p<\infty$, the $\ell_{p}$-ball fails to be homogeneous unlike the polydisc as soon as $p\neq 2$; but it is easy to derive as a corollary of theorem 1.7 that through each point of the $\ell_{p}$-ball, there exist retracts of all possible dimensions, as the retracts through the origin serve to confirm this.
+It is too much to as for retracts along any given set of directions because as the above proposition shows: not every two dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the ball with respect to the $l^{2}$-norm, $\mathbb{B}^{N}$.
 We must not be amiss to note however that every one dimensional linear subspace of a bounded balanced convex domain in $\mathbb{C}^{N}$ is a retract – this follows by the Hahn–Banach theorem.
@@
 While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimension as well; moreover this can be achieved entirely through elementary means.
-In-fact, our example is as simple as the $l_{q}$-‘ball’ for $q<1$ and we record this in the following proposition.
+In-fact, our example is as simple as the standard $\ell^{q}$-‘ball’ for $q<1$ and we record this in the following proposition; an interesting feature of the $\ell_{q}$-‘ball’ in $\mathbb{C}^{N}$ is that apart from failing to be convex, it fails to be concave unlike its real counter-part.
 Let us also note that while the above-mentioned example of Lempert is a pseudoconvex complete Reinhardt domain much as ours, we believe recording the aforementioned $\ell^{q}$-‘ball’ is worthwhile due to the fact that the literature about retracts of dimension higher than one, is scant at-least compared to complex geodesics and one-dimensional retracts.
+In that regard, we would like to mention the very recent article of [GZ] which appeared on the arXiv, a couple of months after the earlier versions of our article.
+Ghosh – Zwonek attain in this article [GZ], a complete characterization of the two dimensional retracts (and thereby all retracts) of the three dimensional Lie ball and the tetrablock; this article also has noteworthy general results about retracts of Lempert domains.
+We now get to the special case of the $\ell^{q}$-‘ball’ for $q<1$ in arbitrary (finite) dimension, which is amongst the simplest examples of (pseudoconvex) domains which fail to be a Lempert domain (see proposition 5.8 below as well); indeed, we deal with large collections of domains that fail to be Lempert domains (however, it goes without saying that the case of Lempert domains is indeed of interest).
 Theorem 1.9.
-[…]
-Consequently, the only $r$-dimensional linear retracts of $D_{q}$ are those spanned by an $r$-element subset of the standard basis $\{e_{1},\ldots,e_{N}\}\subset\mathbb{C}^{N}$.
+Let $q$ be an arbitrary positive number less than one.
+Consider the $\ell^{q}$-‘ball’ given by $D_{q}=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;|z_{1}|^{q}+\ldots+|z_{N}|^{q}<1\}$ and the $r$-dimensional subspace $Y=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r}~{}\text{for}~{}r+1\leq i\leq N\}$ of $\mathbb{C}^{N}$, where for each fixed $i$, not all $a_{i,j}$’s are zero.
+Then the $r$-dimensional subspace $Y_{q}=D_{q}\cap Y$ is not a retract of $D_{q}$.
+In-fact, the only $r$-dimensional linear retracts of $D_{q}$ are those spanned by an $r$-element subset of the standard basis $\{e_{1},\ldots,e_{N}\}\subset\mathbb{C}^{N}$.
 Thus, the holomorphic retracts through the origin of $D_{q}$ are precisely those linear subspaces spanned by some subset of the standard basis vectors.
-Next, we get back to the discussion about Vesentini’s result.
-Let us remark a word or two, about concrete examples of domains satisfying the $\mathbb{C}$-extremality condition in his result.
-Firsly, all $\ell^{q}$-balls for $1\leq q<\infty$ satisfy this condition, whereas a product of any two of them fails to, much as the polydisc.
-While the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary, as will be verified in proposition 3.1 in section 3.
-Vesentini’s result applies to pin-down the retracts through the center of such a ‘ball of intersection’.
-Next, we may want to go beyond the class of domains in Vesentini’s result.
-As the most natural class of domains in several complex variables are pseudoconvex domains, it is only natural to ask if convexity in the above theorem of Vesentini can be relaxed to pseudoconvexity and the extremality condition dropped.
-Towards this, we have the following result wherein we assume boundedness but it must be noted that it may be replaced by hyperbolicity (allowing unboundedness), which however is not significant here, owing to an old result of Kodama in [Kodama] according to which any hyperbolic balanced domain is equivalent to a bounded one.
+Note that the boundaries of these domains are holomorphically extreme; indeed $\mathbb{C}$-extremal, as is the case for $\ell^{p}$-balls for $1\leq p<\infty$ as well, in which case Vesentini’s theorem reduces considerations from holomorphic retracts to linear ones.
+We also furnish examples (of balanced non-convex pseudoconvex domains) for the case where such extremality condition fails in the section where we detail the proof of the above theorem.
+Indeed, we shall show in such examples that retracts of all possible dimension exists through each point of the domain other than the origin as well.
+That retracts cannot also be expected in any given direction for the non-convex case, is substantiated by $D_{q}$ as in the above theorem, owing to the spectacular failure even at the origin (of $1$-dimensional retracts for instance, along any direction other than the coordinate directions).
+Getting back to analysing a bit further, examples arising out of $\ell^{p}$-balls, let us remark a word or two, about concrete examples of domains satisfying the $\mathbb{C}$-extremality condition in his result before we pass out of it.
+Firstly, while the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary; Vesentini’s theorem applies to pin-down the retracts through the center of such a ‘ball of intersection’.
+Finally and more importantly, to bring out the role of the extremality condition both in Vesentini’s theorem and allied results both in the foregoing as well as the sequel (such as theorem 1.15 involving $\mathbb{R}$-extremality), let us mention that the converse of Vesentini’s theorem is also true: if all retracts through the origin of a bounded balanced convex domain are linear, then the domain is $\mathbb{C}$-extremal.
+In-fact, it is enough to just have linearity of the one-dimensional retracts and furthermore, a general and local version can be formulated as follows.
+Theorem 1.10.
+Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $p\in\partial D$.
+Assume that $\partial D$ is convex near $p$ or more generally, convex at $p$ with the Minkowski functional of $D$ being continuous near $p$.
+Suppose the only one-dimensional retract through the origin in the direction of $p$ is the linear one, then $p$ is $\mathbb{C}$-extremal i.e., there exist non-linear retracts, as soon as $p$ fails to be a $\mathbb{C}$-extremal boundary point.
+Remark 1.11.
+  • (i)
+    That the hypotheses that $\partial D$ is convex near $p$ implies convexity of $\partial D$ at $p$ together-with the continuity of the Minkowski functional near $p$; this will be discussed during the course of the proof.
+It may happen that $\partial D$ is convex at some boundary point but not in any open piece of the boundary about it; to see a simple example, consider $l^{q}$-‘balls’ for $q<1$ in $\mathbb{C}^{2}$ and examine points $p$ on the coordinate axes.
+This gives an example wherein convexity near $p$ fails but the weaker condition of convexity at $p$ together-with the continuity of the Minkowski functional near $p$ indeed holds.
+Finally here let us recall that if $p$ is a non-convex boundary point then there are no one-dimensional retracts through the origin in the direction of $p$.
+  • (ii)
+    More importantly, a precise description of all the non-linear retracts through the origin in a bounded balanced pseudoconvex domain along the direction of a boundary point $p$ at which $\partial D$ is convex but fails to be $\mathbb{C}$-extremal – with no assumptions about the continuity of the Minkowski functional near $p$ – is attained in proposition 6.4.
+We record the essence of this in the following theorem.
+Theorem 1.12.
+Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
+Suppose that $\partial D$ is convex near $p\in\partial D$ but with the boundary not being $\mathbb{C}$-extremal at $p$.
+Let $Z$ be an arbitrary one-dimensional retract of $D$ through the origin along the direction of $p$.
+Then $Z$ can be realized as the graph of a holomorphic map on the linear retract ${\rm span}(p)\cap D$, taking values in a bounded balanced pseudoconvex domain within the face $F_{\mathbb{C}}(p,\partial D)$ where $F_{\mathbb{C}}(p,\partial D)$ is the complex face of $\overline{D}$ at the boundary point $p$.
+Next, we wish to go beyond the purview of Vesentini’s result.
+To do this by starting with the simplest step, we retain the assumption of balancedness; also, as the most natural class of domains in several complex variables are pseudoconvex domains, we ask if convexity in the above theorem of Vesentini can be relaxed to pseudoconvexity with no further assumptions about convexity as in the above theorem and extremality conditions dropped as well.
+Towards this, we have the following result wherein we assume boundedness.
+It must be noted that dropping boundedness leads to the failure of the result whereas replacing boundedness by the apparently more general condition of Kobayashi hyperbolicity is insignificant here, owing to the already mentioned result of Kodama in [Kodama], according to which any hyperbolic balanced domain is bounded.
 Theorem 1.13.
@@
 Let $L$ denote the tangent space to $Z$ at the origin.
-Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a (single) holomorphic function over $D_{L}$.
+Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$.
 Relegating some relevant observations here to section 3, we record here a corollary for an important class of domains.
@@
 Then, every non-trivial retract $Z$ (not necessarily passing through the origin in $D$) is biholomorphic via a member of ${\rm Aut}(D)$ to the graph of a holomorphic map over a linear retract $D_{L}$ (which as usual is a ball with respect to some norm in the lower dimensional complex Euclidean space $L=T_{0}Z$), of $D$.
-Owing to the assumption of being balanced, we recall that the domains as in this corollary are the much studied bounded symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV alongwith two exceptional cases, all of whose automorphism groups are well-known (cf.
-Hua [Hua]).
-This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball (as in theorem 1.7 for the $l^{p}$-ball for $p<\infty$ and theorem 1.8 for the $p=\infty$ case much earlier above), is in general quite difficult.
+Owing to the assumption of being balanced, we recall that the domains as in this corollary are the much studied bounded symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV alongwith two exceptional cases, all of whose automorphism groups are well-known (Hua [Hua]).
+This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball – as in theorem 1.7 for the $l^{p}$-ball for $p<\infty$ and theorem 1.8 for the $p=\infty$ case much earlier above; and few other relevant results as in the already mentioned very recent work [GZ] which deals with among other things, Cartan domains of type IV – is in general quite difficult.
 Moreover, it must be remembered that an orthogonal projection of such $l^{p}$-balls onto a linear subspace in general is not a retract, except for the case $p=2$.
@@
 As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem, wherein which we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.13.
-It must first be noted that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifeslty seen in the simplest case of polydiscs.
-This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in simple cases such as domains decomposable into a product of balls (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc.
-Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains.
-This theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like.
+It must first be noted that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifestly seen in the simplest case of polydiscs.
+This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in the simplest of test cases such as domains decomposable into a product of balls (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc.
+Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains and with a simpler proof following the recent work [BBMV-union].
+As indicated above, this theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like.
 We shall address this here, to the extent we can.
@@
 As the complete formulation of the end result of this, turns out to be lengthy, we shall just lay it down in section 3.
-Here we record an application to the union problem for a nice class of polyballs which is interesting owing to being a case with some degeneracy.
+Here we record an application to the union problem for a nice class of polyballs which is interesting owing to it being a case with some degeneracy; this is based on the ideas of Fornaess – Sibony in their article [Fost].
 Recall that the union problem comprises the following: let $M$ be a complex manifold which is the union of an increasing sequence of open subsets $M_{j}$ each of which is biholomorphic to a fixed domain $\Omega\subset\mathbb{C}^{N}$; describe $M$ in terms of $\Omega$.
@@
 Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is all the more significant.
-Towards this, we report that the following result.
+Towards this, we report the following result.
 Theorem 1.17.
@@
 for some real $\alpha$, $t\in(0,1)$ and holomorphic function $h:\Delta^{2}\to\Delta$.
-Finally, retracts of $\Omega_{H}^{A}:=\Omega_{H}\setminus A$ where $A$ is any analytic subset of $\Omega_{H}$ are precisely given by $Z\cap\Omega_{H}$ where $Z$ is a retract of $\Omega_{H}$.
+Finally, retracts of $\Omega_{H}^{A}:=\Omega_{H}\setminus A$ where $A$ is any analytic subset of $\Omega_{H}$ are precisely given by $Z\cap\Omega_{H}^{A}$ where $Z$ is a retract of $\Omega_{H}$.
 While retracts of $\Omega_{H}$ as in the above theorem, will be deduced as an immediate corollary of lemma 8.1 mentioned above, what does not follow so immediately are retracts of its analytic complements $\Omega_{H}^{A}$.
@@
 Theorem 1.19.
-If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of ×$\mathbb{C}\times D$ is either of the form ×$\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$.
+If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of ×$\mathbb{C}\times D$ is either of the form ×$\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$; and, conversely every such graph is a retract.
 It must be noted that this also addresses the other implicit assumption in most of our previous considerations, namely that of topological triviality.
@@
 Let $X=\mathbb{C}\setminus A$ where $A$ is any discrete set (allowed to be empty as well), and $D$ any bounded domain in $\mathbb{C}^{N}$.
-Then, every retract $\widetilde{Z}$ of $X\times D$ is of the form
+Then, every retract $\widetilde{Z}$ of $X\times D$ is precisely of the form
 $\widetilde{Z}=\{(f(w),w):w\in Z\},$
@@
 Theorem 1.21.
-Every non-trivial retraction map $F$ of ×$\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
+Every non-trivial retraction map $F$ of ×$\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is precisely: either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
 While this article also focusses attention for the most part on domains in $\mathbb{C}^{N}$, we would like to interject here a result about retracts of manifolds, as it is a by-product of the essential ideas in the proof of the foregoing theorem 1.19 and more importantly, includes cases with far greater topological complexity.
@@
     $X$ is a non-hyperbolic Riemann surface and $Y$ is a (Kobayashi) hyperbolic complex manifold.
-Then, every holomorphic retract $\widetilde{Z}$ of $X\times Y$ is either of the form
+Then, every holomorphic retract $\widetilde{Z}$ of $X\times Y$ is precisely: either of the form
 $\widetilde{Z}=\{(f(w),w):w\in Z\},$
-where $Z$ is a retract of $Y$ and $f$ is a holomorphic map from $Z$ to $X$, or of the form $X\times Z$ where again $Z$ is a retract of $Y$.
-Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{2}$ – indeed, this is the simplest of balanced domains of holomorphy or for that matter, the simplest of domains with non-trivial retracts and thereby too special to be remiss of us, if we do not remark about this atleast – and ask for a complete characterization of all of its holomorphic retracts (which seems unknown).
+where $Z$ is a retract of $Y$ and $f$ is a holomorphic map from $Z$ to $X$ or, of the form $X\times Z$ where again $Z$ is a retract of $Y$.
+Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{N}$ itself.
+While it is rather hard to precisely characterize all retracts of $\mathbb{C}^{N}$ and we shall show by actual examples that it is possible for non-hyperbolic domains to possess hyperbolic retracts, it is easy to observe a qualitative general feature common to all retracts of $\mathbb{C}^{N}$ apart from their topological triviality namely, that they are non-hyperbolic to the same (maximal) extent as $\mathbb{C}^{N}$ i.e., their Kobayashi metric vanishes identically.
+It is also easy to see that the one-dimensional retracts of $\mathbb{C}^{N}$ are biholomorphic to $\mathbb{C}$.
+For more precise results, we consider $\mathbb{C}^{2}$ – indeed, this is the simplest of balanced domains of holomorphy or for that matter, the simplest of domains with non-trivial retracts – and ask for a complete characterization of all of its holomorphic retracts (which seems unknown).
 Despite the apparent simplicity of the domain, this seems rather non-trivial; indeed, while graphs (in $\mathbb{C}^{2}$) of entire functions on $\mathbb{C}$ are retracts and the fact that retracts of products of planar domains in the foregoing, have literally turned out to be graphs of holomorphic maps over one of the factors, it is not the case that: every retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ is the graph of some entire function on either of its factors, as we shall show by providing a concrete example in the last section 10.
-On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this is attained by the following theorem.
-Theorem 1.20.
-If $R$ is a non-trivial polynomial retract of $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ on $\mathbb{C}^{2}$ such that $\psi(R)=\{(\zeta,0)\in\mathbb{C}^{2}\;:\;\zeta\in\mathbb{C}\}$.
-Acknowledgements
-This article forms the thesis work of the second author and we thank Prof.
-Kaushal Verma for giving us this problem of determining holomorphic retracts; as well as for helpful discussions, for which we also thank Dr.
-J.
-Jaikrishnan.
+On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this in essence is attained by Shpilrain and Yu in [Shpilrain--Yu] wherein it is deduced from other known results about polynomial rings and their retracts; more about this in the final section.
+Theorem 1.23 (Shpilrain – Yu).
+If $Z$ is a non-trivial polynomial retract of $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\Phi$ on $\mathbb{C}^{2}$ such that $\Phi(Z)=\{(\zeta,0)\in\mathbb{C}^{2}\;:\;\zeta\in\mathbb{C}\}$.
+Such a ‘rectifiability’ assertion cannot in general be made for all holomorphic retracts of $\mathbb{C}^{N}$ for any $N\geq 2$, as follows by a simple analysis of the proof of the main result in the work [Forstn] due to Forstneric–Globevnik–Rosay.
+Acknowledgments: We thank Gautam Bharali, J.
+Jaikrishnan and Kaushal Verma.
+The second named author acknowledges the support of the CSIR-JRF fellowship provided by the Council of Scientific and Industrial Research, India.
 2. Preliminaries
-We recall certain basic and general facts about retracts which would needless to say, apply to our specific case and put it in good perspective, as well.
-Let us first recall that a self-map $F$ of a set $X$ is called a retraction if $F\circ F=F$; note that this idempotent condition on $F$ means that $F$ acts as the identity map on its range $Z=F(X)$.
-So, for any Hausdorff topological space $X$, its retracts, which are the images of continuous retractions, are closed subsets of $X$.
-If $X$ is connected, then so is any of its retracts $Z$.
-When $X$ is furthermore contractible i.e., the space $X$ deformation retracts to a point then by composing any such deformation retraction with the retraction map $F$, gives a homotopy within $Z$ that reduces $Z=F(X)$ to a point, meaning that if $X$ is contractible then so are all of its retracts.
-If $X$ even happens to be a convex domain in some real topological vector space, then the retracts of $X$ are also strong deformation retracts.
-We are ofcourse interested in holomorphic retracts: if $X$ is a complex manifold and $F$ is a holomorphic self-map of $X$ which is idempotent (i.e., $F\circ F=F$) then, we say that $F$ is a holomorphic retraction and its image $Z=F(X)$ a holomorphic retract; as we shall only be talking about holomorphic retracts, we shall most often drop the adjective ‘holomorphic’.
-Moreover as mentioned in the introduction, our focus will be to determine the form of the retracts $Z=F(X)$ rather than the retractions $F$.
+We recall certain basic and general facts about retracts which would needless to say, apply to our specific cases and put it in good perspective, as well.
+While we would like to minimize repetition, we prefer to put the details on record to ensure a rigorous exposition.
+But here we repeat a few generalities from the introductory section for one last time, to cast this section systematically.
+Let us begin with the most general definition relevant to us: a self-map $\rho$ of a set $X$ is called a retraction if $\rho\circ\rho=\rho$; note that this idempotent condition on $\rho$ means that $\rho$ acts as the identity map on its range $Z=\rho(X)$.
+So, for any Hausdorff topological space $X$, its retracts, which are the images of continuous retractions, are closed subsets of $X$; henceforth Hausdorffness on $X$ will be in force.
+We regard $\rho$ as being non-trivial, if it is neither a constant map nor the identity.
+In addition to the contrasts between proper maps and retractions mentioned in the introduction, we note that non-trivial retraction mappings may very well fail to be closed maps notwithstanding their ranges being closed subsets always; indeed, as soon as $X$ is a non-compact complex manifold – the general setting of interest in this article – they always fail to be closed maps, despite the general fact that $\rho(X)$ is always closed as a subset of $X$, for every retraction map $\rho$ on $X$ as just mentioned (proofs detailed later below).
+However, the feature of being a quotient map is common to both retraction maps as well as proper maps.
+Getting back to the general case of a Hausdorff topological space $X$, if we add the assumption that $X$ is connected, then so are every of its retracts $Z$.
+When $X$ is furthermore contractible i.e., the space $X$ deformation retracts to a point then by composing any such deformation retraction with the retraction map $\rho$, gives a homotopy within $Z$ that reduces $Z=\rho(X)$ to a point, meaning that if $X$ is contractible then so are all of its retracts.
+If $X$ even happens to be a convex domain in some real topological vector space, then the retracts of $X$ are also strong deformation retracts; indeed, the homotopy $H(t,x)=(1-t)x+t\rho(x)$ substantiates this.
+We are ofcourse interested in holomorphic retracts: if $X$ is a complex manifold and $\rho$ is a holomorphic self-map of $X$ which is idempotent then, we say that $\rho$ is a holomorphic retraction and its image $Z=\rho(X)$ a holomorphic retract; as we shall only be talking about holomorphic retracts, we shall most often drop the adjective ‘holomorphic’.
+Moreover as mentioned in the introduction, our focus will be to determine the form of the retracts $Z=\rho(X)$ rather than the retractions $\rho$.
 A first basic and general fact about holomorphic retracts that is helpful to have in the background for the discussion in the sequel is laid down in the following well-known theorem which says that they are not just analytic varieties but actually (connected) manifolds i.e., are smooth and have the same dimension throughout.
@@
 implying that $F_{p}$ is of full rank. Note that $F_{p}$ maps $U_{p}$ into $L$.
-Continuing with the proof theorem 2.1, we shall now show that in a neighbourhood of any given $p\in Z$, $Z$ is the level set of a holomorphic map whose derivative has constant rank near $p$; the implicit mapping theorem will then render the claim of the theorem.
+Continuing with the proof of theorem 2.1, we shall now show that in a neighbourhood of any given $p\in Z$, $Z$ is the level set of a holomorphic map whose derivative has constant rank near $p$; the implicit mapping theorem will then render the claim of the theorem.
 To this end then, we fix an arbitrary $p\in Z$ and passing to a holomorphic chart of the complex manifold $X$ near $p$, we may as well assume (as already mentioned), that we are in the complex Euclidean space $\mathbb{C}^{N}$, which allows us to define $\beta(x)=x-\rho(x)$ for $x$ near $p$, whose zero set is precisely $Z$ (all in the charts near $p$, to be precise).
@@
 Retracts of Stein manifolds are Stein and likewise, the property of a complex manifold being an Oka manifold, is preserved under passage to a retract thereof, as well.
-Another general fact to keep in the background is the following; we include its proof, despite its obviousness, owing to its use for the rest of this article.
+Proposition 2.3.
+Let $X$ be a non-compact complex manifold with $Z$ being any non-trivial retract of $X$. Then none of the retraction maps $\rho:X\to Z$ is a closed map.
+Proof.
+Choose any convergent sequence $(z_{n})$ in $Z$ of distinct points; as we already know $Z$ is closed, thereby the limit $z_{0}$ of the chosen sequence lies in $Z$.
+Our strategy will be to pick points $p_{n}$ in the fibers over $z_{n}$ such that the $p_{n}$’s form a discrete, thereby a closed subset $C$  of $X$.
+If we let $K$ denote the (infinite) compact set $K=\{z_{n}\}\cup\{z_{0}\}$ and note that $\rho(C)$ equals $K$ without its limit point, we would have proven the proposition.
+The possibility of choosing a sequence $p_{n}$ as desired above really requires the non-compactness in hypothesis.
+Indeed, consider the sequence of fibers of $\rho$ given by $F_{n}:=\rho^{-1}(z_{n})$.
+Note firstly that being non-trivial analytic sets, each of the $F_{n}$’s is non-compact and $\rho^{-1}(K\setminus\{z_{0}\})=\cup_{n=1}^{\infty}F_{n}$.
+We then just use the fact from general topology that: being second countable, locally compact and Hausdorff, $X$ admits an exhaustion by compact subsets $K_{n}$.
+In particular, since $X$ is not compact, the complements $K_{n}^{c}=X\setminus K_{n}$ is non-empty for each $n\in\mathbb{N}$, ensuring that we can choose a point $p_{n}\in F_{n}\cap K_{n}^{c}$.
+To prove that $p_{n}$’s form a discrete subset of $X$ as needed by our strategy, assume to get a contradiction that there exists a subsequence $(p_{n_{k}})$ converging to $p$.
+This implies that there exists $N\in\mathbb{N}$ such that the $N$-th compact captures our sequence eventually: $p_{n_{k}}\in K_{N}$ for all $n_{k}$; but that contradicts the choice of $p_{n}$, finishing the proof that $p_{n}$ has no limit point in $X$.
+Thus $C:=\{p_{n}:n\in\mathbb{N}\}$ is a closed subset of $X$ whose image $\rho(C)=\{z_{n}:n\in\mathbb{N}\}$ fails to be closed as its limit point $z_{0}$ lies outside it, establishing that $\rho$ is not a closed map.
+∎
+We turn to questions which arise when we have more than one retraction map and some initial remarks about the basic question about how do we generate new retraction maps out of old.
+Algebraic or other operations would require special assumptions about the ambient space; but whatever be the ambient space $X$, viewing retraction maps as self-maps of $X$, we may always compose any two such maps and ask when does it result in a retraction?
+As the first answer, we have: if two retraction maps $\rho_{1},\rho_{2}$ on $X$ commute, then $\rho:=\rho_{1}\circ\rho_{2}=\rho_{2}\circ\rho_{1}$ is also a retraction map on $X$ whose image is the retract given by the intersection of the retracts $\rho_{1}(X)$ and $\rho_{2}(X)$.
+If we are allowed to change the space, then there are even more trivial ways such as the following.
+Given any retraction with notations as above, we may pick an arbitrary subset $A$ of the retract $Z$ and consider the restriction of $\rho$ to $X\setminus\rho^{-1}(A)$; this renders $Z\setminus A$ as a retract of $X\setminus\rho^{-1}(A)$.
+In other words, we may get retracts of $X$ as intersections of retracts of a space containing $X$ with $X$.
+The question of whether all retracts of $X\setminus A$ can be realized as being obtained this way is interesting only when $X$ and $A$ satisfy certain conditions i.e., the converse of the last statement indeed fails and there are nice counterexamples towards this.
+To put down the simplest of examples in the complex analytic category, consider $X=\mathbb{C}^{2}\setminus\{\rm w-axis\}$ where the standard coordinates are herein and henceforth in this article, denoted by $(z,w)$.
+We then observe that the retract $Z$ described by the equation $zw=1$ which is the image of the retraction $(z,w)\mapsto(z,1/z)$, cannot be realized as mentioned just earlier – the proof of this is recorded in remark 8.2 following a lemma to answer the aforementioned question in the positive direction.
+In passing, we put down a rudimentary fact to keep in the background about comparing retracts of a pair of related spaces; we include its proof alongwith statement despite its obviousness (as with certain other facts), owing to its use for the rest of this article.
 Proposition 2.4.
@@
 To indicate a way of generating examples, let $E_{2m}:=\{(z,w)\;:\;|z|^{2m}+|w|^{2}<1\}$.
-For an arbitrary point $(z,w)\in E_{2m}$ note that $|z|<1$, thereby that, $|z^{2l}|+|w|^{2}<|z|^{2m}+|w|^{2}$ for every $l\geq m$.
-Consequently, $(z^{2l},w)$ lies in $E_{2l}$ for every $l\geq m$, implying that $(z,w)\mapsto(z^{2l},w)$, gives an example of a non-linear (in-fat proper) holomorphic mapping as desired and more importantly thereby, a non-linear retract passing through the origin in the polyball $E_{2m}\times E_{2l}$ for every $l\geq m$.
+For an arbitrary point $(z,w)\in E_{2m}$ note that $|z|<1$, thereby that, $|z^{n}|^{2l}+|w|^{2}<|z|^{2m}+|w|^{2}$ for every $l\geq m$.
+Consequently, $(z^{n},w)$ lies in $E_{2l}$ for every $l\geq m$, implying that $(z,w)\mapsto(z^{n},w)$, gives an example of a non-linear holomorphic mapping as soon as $n\geq 2$ as desired and more importantly thereby, a non-linear retract passing through the origin in the polyball $E_{2m}\times E_{2l}$ for every $l\geq m$.
 Needless to say, similar constructions can be done with $l_{p}$-balls and more; we shall not digress further along this line.
@@
 Let $K$ be the field $\mathbb{R}$ or $\mathbb{C}$.
-Let $\Omega\subset\mathbb{C}^{N}$ be an arbitrary domain and $a\in\partial{\Omega}$.
-We introduce a notion that captures the set of all directions along which one can move from $a$, at-least a little bit while staying within the boundary $\partial\Omega$.
-Towards this, define the face of $\partial\Omega$ at $a$ by
-$F_{K}(a,\partial{\Omega})=\{u\in\mathbb{C}^{N}:\exists\epsilon>0,a+tu\in\partial{\Omega},\text{ for all }t\in K~{}with~{}|t|<\epsilon\}$
+Let $A\subset\mathbb{C}^{N}$ be an arbitrary set and $a\in\partial{A}$.
+We introduce a notion that captures the set of all directions along which one can move from $a$, at-least a little bit while staying within the boundary $\partial A$.
+Towards this, define the face of $\partial A$ at $a$ by
+$F_{K}(a,\partial{A})=\{u\in\mathbb{C}^{N}:\exists\epsilon>0,a+tu\in\partial{A},\text{ for all }t\in K~{}with~{}|t|<\epsilon\}$
 reduces to $\{0\}$.
 It is immediate that $F_{K}$ is closed under scalar multiplication.
-But a stronger claim that $F_{K}$ is a linear subspace of $\mathbb{C}^{N}$ requires more assumptions on $\Omega$.
-Indeed, if $\Omega$ is convex, then $F_{K}(a,\partial{\Omega})$ is $K$-linear subspace.
-A boundary point $a\in\partial{\Omega}$ is said to be a $K$-extremal boundary point for $\Omega$, if $F_{K}(a,\partial{\Omega})=0$; if the face reduces to $\{0\}$ for every $a\in\partial{\Omega}$, then we say that the domain $\Omega$ is $K$-extremal.
+But a stronger claim that $F_{K}$ is a linear subspace of $\mathbb{C}^{N}$ requires more assumptions on $A$.
+Indeed, if $A$ is convex set, then $F_{K}(a,\partial{A})$ is $K$-linear subspace.
+We shall mostly deal with the case when $A=\Omega$ is a domain.
+In such a case, by abuse of notation we shall write $F(a,\partial\Omega)$ or $F(a,\overline{\Omega})$ in place of $F(a,\Omega)$.
+A boundary point $a\in\partial{\Omega}$ is said to be a weakly  $K$-extremal boundary point for $\Omega$, if $F_{K}(a,\partial{\Omega})=0$; if the face reduces to $\{0\}$ for every $a\in\partial{\Omega}$, then we say that the domain $\Omega$ is weakly $K$-extremal.
+A boundary point $a\in\partial{\Omega}$ is said to be a strongly  $K$-extremal boundary point for $\Omega$, if the face
+$F_{K}(a,\overline{\Omega})=\{u\in\mathbb{C}^{N}:\exists\epsilon>0,a+tu\in\overline{\Omega},\text{ for all }t\in K~{}with~{}|t|<\epsilon\}$
+reduces to $\{0\}$.
+As is usual, if all boundary points are strongly $K$-extremal, then we say that the domain itself is strongly $K$-extremal.
+In general, if $\Omega$ is strongly $K$-extremal, then we shall drop the adjective ‘strongly’ and call $\Omega$ simply as a $K$-extremal domain.
+We now prove that in the case of bounded balanced pseudoconvex domains, strongly $\mathbb{C}$-extremal and weakly $\mathbb{C}$-extremal are equivalent, thereby we can forget about the qualifications ‘weak’ and ‘strong’ essentially for the rest of this article.
+We leave the easier implication about strong extremality implying weak extremality aside and detail only the relatively non-trivial reverse implication.
+Proposition 2.13.
+Suppose $D$ is a bounded balanced pseudoconvex domain and $p\in\partial D$ is a weakly $\mathbb{C}$-extremal boundary point. Then it is also a strongly $\mathbb{C}$-extremal boundary point.
+Proof.
+Assume that $p$ is not strongly $\mathbb{C}$-extremal.
+Then there exists $\epsilon_{0}>0$ and non-zero $u\in\mathbb{C}^{N}$ such that $p+tu\in\overline{D}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$.
+After a rescaling, we may assume that $\epsilon_{0}=1$.
+Then the affine (holomorphic) map $f$ defined by $f(t)=p+tu$ maps $\Delta$ into $\overline{D}$; note that $f(0)=p$.
+Let $h$ denote the Minkowski functional of $D$.
+Consider the map $\varphi:\Delta\to\mathbb{R}$ defined by $\varphi(z)=h\circ f(z)$.
+Note that $\varphi$ is a plurisubharmonic function whose modulus attains maximum at $z=0$.
+By maximum principle of plurisubharmonic functions, $\varphi$ is identically constant, which is equal to $1$.
+This means that $h\circ f\equiv 1$.
+So ${\rm image}(f)\subset\partial D$.
+Hence $p$ is not weakly $\mathbb{C}$-extreme boundary point.
+This finishes the proof by contradiction.
+∎
+Remark 2.14.
+If $K=\mathbb{R}$, then in the case of bounded balanced convex domains, the notion of weak $\mathbb{R}$-extremality and strong $\mathbb{R}$-extremality are equivalent.
 We would now like to make a couple of clarifications about the case $K=\mathbb{C}$ versus the case $K=\mathbb{R}$. To begin with, note that we obviously have
-$\vec{F}_{\mathbb{C}}(a,X)=\vec{F}_{\mathbb{R}}(a,X)\cap i\vec{F}_{\mathbb{R}}(a,X).$
-It follows that if $\Omega$ is $\mathbb{R}$-extremal, then $\Omega$ is $\mathbb{C}$-extremal.
+$\vec{F}_{\mathbb{C}}(a,\overline{\Omega})=\vec{F}_{\mathbb{R}}(a,\overline{\Omega})\cap i\vec{F}_{\mathbb{R}}(a,\overline{\Omega}).$
+It follows that if $\Omega$ is weakly (resp.
+strongly) $\mathbb{R}$-extremal, then $\Omega$ is weakly (resp.
+strongly) $\mathbb{C}$-extremal.
 But the converse is not true.
 To see this, consider $B_{\ell^{1}}:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<1\}$; we shall presently proceed to showing that this is $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal.
-Perhaps even more basically, we must remark that every domain in $\mathbb{C}$ is $\mathbb{C}$-extremal.
-But only strict convex domains are $\mathbb{R}$- extremal.
+Perhaps even more basically, we must remark that (vacuously) every domain in $\mathbb{C}$ is $\mathbb{C}$-extremal whereas only the strictly convex ones are $\mathbb{R}$-extremal.
+To discuss a ‘non-vacuous’ example, we discuss the well-known $\ell^{1}$-ball as mentioned above whose proof is included here owing to its basic importance for us.
 Lemma 2.15.
@@
 To do this, let $a\in\partial B_{\ell^{1}}$.
-Suppose there exists $\epsilon_{0}>0,a\in\partial B_{\ell^{1}}$ and $u_{0}\in\mathbb{C}^{N}$ such that $a+tu_{0}\in\partial B_{\ell^{1}}$, for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$.
+Suppose there exists $\epsilon_{0}>0$ and $u_{0}\in\mathbb{C}^{N}$ such that $a+tu_{0}\in\partial B_{\ell^{1}}$, for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$.
 Without loss of generality, we may assume after an appropriate scaling that $\epsilon_{0}=1$; this ensures that the affine-linear (holomorphic) map defined by $\rho(t)=a+tu_{0}$ maps $\Delta$ into $\partial B_{\ell^{1}}$.
@@
 Note then that $(1,-1)\in F_{\mathbb{R}}(a,\partial B_{\ell^{1}})$. Hence $B_{\ell^{1}}$ is not $\mathbb{R}$-extremal. ∎
+We shall now discuss the relationship between the above extremality notions and that of strict convexity.
 Definition 2.17.
 […]
-First we prove that the notion of rotundity and strictly convex are equivalent in a bounded balanced convex domain $\Omega$.
+In order to establish equivalence of $\mathbb{R}$-extremality with strict convexity, it is convenient to do this, the intermediate concept namely of rotundity, which is defined as follows:
+Definition 2.18.
+[…]
+First we prove that the notion of rotundity and strict convexity are equivalent in the realm of bounded balanced convex domains.
 Theorem 2.19.
@@
 Assume $\Omega$ is rotund.
-Let $a\in\partial\Omega$ and $(a+tu)\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$.
+Let $a\in\partial\Omega$ and assume that there exists $u\in\mathbb{C}^{N}$ such that $a+tu\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$.
 Note that $\frac{1}{2}(a+(a+tu))=a+\frac{t}{2}u\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$.
 By proposition 5.1.2 of [Meggn], $a=a+tu$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$ and this implies that $u=0$.
-Therefore, $\Omega$ is $\mathbb{R}$extremal.
-Conversely assume that $\Omega$ is $\mathbb{R}$extremal. Let $z_{1},z_{2}\in\partial\Omega$.
+Therefore, $\Omega$ is $\mathbb{R}$-extremal.
+Conversely assume that $\Omega$ is $\mathbb{R}$-extremal. Let $z_{1},z_{2}\in\partial\Omega$.
 $(1-t)z_{1}+tz_{2}=z_{1}+t(z_{2}-z_{1})\in\partial\Omega\quad\text{for ~{}all}\ t\in[0,1].$
-$\frac{z_{1}+z_{2}}{2}+t(z_{2}-z_{1})=(1/2-t)z_{1}+(1/2+t)z_{2}\in\partial{\Omega}~{}\text{for ~{}all}\ t\ \text{with}\ |t|<1/2$
-Therefore $z_{2}-z_{1}\in F_{\mathbb{R}}(\frac{z_{1}+z_{2}}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$- extremal, $z_{2}=z_{1}$ and therefore $\Omega$ is rotund.
+It follows that when $|t|\leq 1/2$, we have
+$\frac{z_{1}+z_{2}}{2}+t(z_{2}-z_{1})=(1/2-t)z_{1}+(1/2+t)z_{2}\in\partial{\Omega}.$
+Therefore $z_{2}-z_{1}\in F_{\mathbb{R}}(\frac{z_{1}+z_{2}}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$- extremal, $z_{2}=z_{1}$ and therefore $\Omega$ is rotund. ∎
 As the relation between the pair of notions gotten by taking $K=\mathbb{C}$ versus $K=\mathbb{R}$ will be basic importance in the sequel, we discuss this below.
-But one would perhaps like to know first the relation between strict convexity and extremality introduced above; as we do not find a quick reference for this, we include the details for this here.
+But one would perhaps like to ascertain first the relation between strict convexity and extremality introduced above.
+This follows from known facts (see for instance [Meggn]), but owing to its central importance for us, we include the details.
 Theorem 2.20.
@@
 By corollary 5.1.16 of [Meggn], there exists $\mathbb{R}$-linear functional $f$ supports more than one point.
-This implies that there exists $z_{0},w_{0}\in\mathbb{C}^{N}$ such that $f(z_{0})=f(w_{0})=\|f\|$.
+This implies that there exists $z_{0},w_{0}\in\overline{\Omega}$ such that $f(z_{0})=f(w_{0})=\|f\|$.
 Take $\alpha:=\|f\|$.
-Define a $\mathbb{R}$- linear functional $g$ defined by $g(z)=\frac{1}{\alpha}f(z)$.
+Consider the $\mathbb{R}$-linear functional $g$ defined by $g(z)=\frac{1}{\alpha}f(z)$.
 Note that $\|g|\ =1$ and $g(z_{0})=g(w_{0})=1$.
@@
 Therefore, $\Omega$ is $\mathbb{R}$- extremal.
-Conversely assume that $\Omega$ is $\mathbb{R}$- extremal.
+For the converse, assume that $\Omega$ is $\mathbb{R}$-extremal.
 Let $T$ be a $\mathbb{R}-$ linear map with $\|T\|=1$.
-Assume that there exists $z,w\in\overline{\Omega}$ such that $Tz=1=Tw$.
-Then $z,w\in\partial{\Omega}$, otherwise $T\left(\frac{z}{\|z\|}\right)=\frac{1}{\|z\|}>1$.
+Suppose that there exist $z,w\in\overline{\Omega}$ such that $Tz=1=Tw$.
+Then $z,w\in\partial{\Omega}$, otherwise $T\left(z/\|z\|\right)=1/\|z\|>1$.
 Note that $T((1-t)z+tw)=1$.
@@
 $\frac{z+w}{2}+t(w-z)=(1/2-t)z+(1/2+t)w\in\partial{\Omega}~{}\ \text{for all t}\in\mathbb{R}\ \text{with}\ |t|<1/2.$
-Then $w-z\in F_{\mathbb{R}}(\frac{z+w}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$-extremal, then $w=z$. Therefore $\Omega$ is strictly $\mathbb{R}$-convex. ∎
+Then $w-z\in F_{\mathbb{R}}(\frac{z+w}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$-extremal, it follows that $w=z$. Therefore $\Omega$ is strictly $\mathbb{R}$-convex. ∎
 Theorem 2.21.
@@
 If possible assume that there exists $u\in\mathbb{C}^{N},~{}a\in\partial\Omega$ and an $\epsilon>0$ such that $a+tu\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$.
-Let $M=span_{\mathbb{R}}\{u\}$.
+Let $M=span_{\mathbb{C}}\{u\}$.
+We split the proof now into two steps, formulated as ‘claims’.
 Claim-1: $a\notin M$
 Proof of Claim-1: If possible assume that $a\in M$.
-Choose $s_{0},~{}t_{0}\in\mathbb{R}$ such that $|s_{0}|<\epsilon$ and $|s_{0}t_{0}|<\epsilon$.
-Since $a\in M$, we get $a+s_{0}u\in M$.
-Hence we get $a+s_{0}u=ca$ for some $c\in\mathbb{R}$.
-Taking norm on both sides we get $|c|=1$.
-$\|a+s_{0}t_{0}u\|^{2}=\|(1-t_{0})a+t_{0}(a+s_{0}u)\|^{2}=(1-t_{0})^{2}+2t_{0}(1-t_{0})c+c^{2}t_{0}^{2}<(1-t_{0})^{2}+2t_{0}(1-t_{0})+t_{0}^{2}=(1-t_{0}+t_{0})^{2}=1$.
-Hence we get $\|(1-t_{0})a+t_{0}(a+s_{0}u)\|^{2}<1$. This implies that $\|a+s_{0}t_{0}u\|<1$. Hence, we get $a+s_{0}t_{0}u\notin\partial\Omega$. This contradicts our assumption, proving the claim.
-By Hahn–Banach theorem, there exists a $\mathbb{C}$- linear functional $f$ with $\|f\|=1$ such that $f(z)=0$ for all $z\in M$ and $f(a)=dist(a,M)$.
-Claim-2: ${\rm dist}(a,M)=1$.
-Proof of Claim-2: Let us begin by recalling that $\rm dist(a,M)=\inf\{\|a+su\|:s\in\mathbb{R}\}$ and note firstly, that ${\rm dist}(a,M)\leq 1$.
-Take an arbitrary $s\in\mathbb{R}$.
-Choose $t_{0}\in[0,1)$ such that $|(1-t_{0})s|<\epsilon$ and note that
-$\|(1-t_{0})(a+su)+t_{0}a\|\leq(1-t_{0})\|a+su\|+t_{0}\|a\|=(1-t_{0})\|a+su\|+t_{0}.$
-We simplify the above inequality to write it as
-(2.1) $\|a+(1-t_{0})su\|\leq(1-t_{0})\|a+su\|+t_{0}.$
-Since $|(1-t_{0})s|<\epsilon$, we get from (2.1) that $\|a+(1-t_{0})su\|=1$. This implies $\|a+su\|\geq 1$ for all $s\in\mathbb{R}$. Hence, $dist(a,M)=1$, finishing the proof of claim-2.
-Note that $f(a+tu)=f(a)=1$. Since $\Omega$ is strictly $\mathbb{C}$-convex, we have $a=a+tu$. This implies that $u=0$. So $\Omega$ is $\mathbb{R}$-convex. ∎
+Since $a\in\partial\Omega$, there exists non-zero $\lambda\in\mathbb{C}$ such that $a=\lambda u$.
+Choose $t_{0}$ be a non-zero real number such that $|t_{0}|<\epsilon$ and $|\lambda+t_{0}|\neq|\lambda|$.
+Taking norm on both sides of the equation $a=\lambda u$, we get $|\lambda|=1/\|u\|$.
+From the assumption $a+tu\in\partial\Omega$, we obtain
+$1=\|a+t_{0}u\|=\|\lambda u+t_{0}u\|=\|u\|\cdot|\lambda+t_{0}|=\frac{|\lambda+t_{0}|}{|\lambda|}.$
+The above equations tells that $|\lambda+t_{0}|=|\lambda|$. This contradicts our choice of $t_{0}$ and finishes the proof of the claim.
+[…]
+Proof of Claim-2: Let us begin by recalling that ${\rm dist}(a,M)=\inf\{\|a-su\|:s\in\mathbb{C}\}$ and note firstly (by considering $s=0$) that ${\rm dist}(a,M)\leq 1$.
+To prove this claim then, it suffices to show that $\rm dist(a,M)\geq 1$.
+To this end, it is convenient to rewrite the definition of the aforementioned distance as ${\rm dist}(a,M)=\inf\{\|a+su\|:s\in\mathbb{C}\}$ – this is not a mere sign change, it is a changing over to the ‘dual’ viewpoint that the aforementioned distance is also the infimum of distances of the origin to points on the affine line through $a$ in the direction of $u$.
+Since $D$ is convex, there exists a supporting hyperplane for every $p\in\partial D$.
+In particular, there exists a supporting hyperplane $\widetilde{H}=\{z\in\mathbb{C}^{N}:g(z)=g(a)\}$ (for some $\mathbb{R}$-linear functional $g$) at the point $a$.
+Let $l:=\{a+tu:t\in\mathbb{R}\}$.
+Let us first verify that $l\subset\tilde{H}$.
+To prove this by contradiction, assume the contrary that $l\not\subset\tilde{H}$.
+Consider the linear subspace $H:=\{z\in\mathbb{C}^{N}\;:\;g(z)=0\}$.
+By the definition of supporting hyperplane, $a\notin H$.
+So, $span\{a\}\cap H=\{0\}$.
+We now show that $span\{u\}\cap H=\{0\}$ as well.
+Indeed, if $\lambda u\in H$ for some $\lambda\in\mathbb{C}$, then $g(\lambda u)=\lambda g(u)=0$.
+Since $l\not\subset\widetilde{H}$, there exists a non-zero real number $t_{0}$ such that $a+t_{0}u\notin\widetilde{H}$.
+This means that $u\notin\ker(g)$.
+From $\lambda g(u)=0$, we obtain that $\lambda=0$.
+Hence $span\{u\}\cap H=\{0\}$.
+Now by Claim-1, $a$ and $u$ are linearly independent.
+Since $span\{a\}\cap H=\{0\}$ and $span\{u\}\cap H=\{0\}$, it follows that $dim(H)<N-1$.
+This contradicts the fact that $H$ is a hyperplane.
+Hence $a+su\in\widetilde{H}$ for all $s\in\mathbb{R}$.
+Therefore $\|a+su\|\geq 1$.
+This implies that $\rm dist(a,M)\geq 1$.
+This completes the proof of the claim.
+Completing the proof of the theorem consists in now observing that: by the Hahn–Banach theorem, there exists a $\mathbb{C}$-linear functional $f$ with $\|f\|=1$ such that $f(z)=0$ for all $z\in M$ and $f(a)=\rm dist(a,M)$.
+Note then that $f(a+tu)=f(a)=1$.
+Since $\Omega$ is strictly $\mathbb{C}$-convex, this means $a=a+tu$, which in turn implies $u=0$.
+So $\Omega$ is $\mathbb{R}$-convex, finishing the proof in all.
+∎
+Remark 2.22.
+We shall henceforth not explicitly mention the field and simply refer to a strictly $K$-convex domain as strictly convex domain.
 Remark 2.23.
@@
 We say that a point $a\in\partial D$ is holomorphically extreme if there is no non-constant holomorphic mapping $\phi:\Delta\rightarrow\overline{D}$ such that $\phi(0)=a$.
-The purpose of recalling this notion in this article about retracts is the following result; being a result about retracts, we prefer to include the details of the proof.
-Theorem 2.20.
-(A slight modification of proposition $4.5$ in [JJ]).
-Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
-Suppose $Z$ is a non-trivial retract through the origin with the property that the boundary points of $D_{L}:=D\cap T_{0}Z$ viewed as a domain in $T_{0}Z$ are holomorphically extreme.
-Then, $Z$ is linear.
-Proof.
-Let $\mathfrak{h}_{D_{L}}$ denote the Minkowski function on $D_{L}$.
-Since $Z$ is a retract of $D$, there exists a holomorphic retraction map $\rho:D\to D$ such that $\rho(D)=Z$.
-Note that $\rho|_{D_{L}}$ is a holomorphic map from $D_{L}$ to $D$ and $Z=image(\rho|_{D_{L}})$.
-Observe that $Z$ is a connected complex submanifold and complex tangent space at 00 is $T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}|_{D_{L}}(0)v=v\}$.
-Let $v\in D_{L},v\neq 0$.
-By our assumption on $\partial D_{L}$, it follows that the mapping
-$\phi:\Delta\ni\lambda\longmapsto\lambda v/\mathfrak{h}_{D_{L}}(v)\in D_{L}.$
-is the unique (modulo $Aut(\Delta)$) $K_{D_{L}}$ geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$.
-Since $D_{L}\subset D$ and $L$ is a linear subspace of $\mathbb{C}^{N}$, $\mathfrak{h}_{D_{L}}(v)=\mathfrak{h}_{D_{L}}(v)$.
-So $\phi$ is the unique (modulo $Aut(\Delta)$) $K_{D}$ geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$.
-Note that
-$K_{D}(\rho|_{D_{L}}\circ\phi(0);(\rho|_{D_{L}}\circ\phi)^{\prime}(0))=K_{D}(0,v/\mathfrak{h}_{D_{L}}(v))=K_{D}(\phi(0);\phi^{\prime}(0)).$
-Hence $\rho|_{D_{L}}\circ\phi$ is a geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$.
-By uniqueness, it follows that $\rho|_{D_{L}}\circ\phi=\phi\circ\psi$, where $\psi\in Aut(\Delta)$.
-Substituting $\psi^{-1}(\mathfrak{h}_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho|_{D_{L}}$.
-Hence $D_{L}\subset Z$.
-Now $T_{0}Z$ and $Z$ have the same dimension.
-Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$.
-Hence $Z$ is linear.
-∎
+The purpose of recalling this notion in this article about retracts is to record the details for proposition 1.4 in the introductory section; which will be done in the next section.
 2.3. Geodesics and totally geodesic subsets
@@
 It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic, a notion that shall be made use of, while recalling the work [Mok] of Mok.
+This lemma also shows that if $X$ is hyperbolic then every retract of $X$ is hyperbolic.
 2.4. One dimensional Retracts
@@
 We say that a holomorphic map $\varphi:\Delta\rightarrow D$ is a complex $C$-geodesic of $D$ if for all $\zeta_{1},\zeta_{2}\in\Delta$, we have $c_{D}\left(\varphi\left(\zeta_{1}\right),\varphi\left(\zeta_{2}\right)\right)=p\left(\zeta_{1},\zeta_{2}\right)$.
-Similarly we can define complex $K$-geodesic (or complex Kobayashi geodesic) if we replace $c_{D}$ by $K_{D}$ in the above definition.
-By Lemma 2.6.1 (ii) i [Abate_Itrtn], every complex $K$-geodesic is a complex $C$-geodesic.
+Here $c_{D}$ denotes the Caratheodory distance; as we shall be working with a variant of this as well, we recall both of them and their notations:
+$c_{D}(z,w)=\sup\{p(f(z),f(w)):f\in\mathcal{O}(G,\Delta)\}$
+The variant of $c_{D}$ denoted $c_{D}^{*}$ which will also be useful later for us, is defined as follows.
+$c_{D}^{*}(z,w)=\sup\{m(f(z),f(w)):f\in\mathcal{O}(G,\Delta)\}$
+‘Dualizing’ this definition i.e., by swapping the roles of $G$ and $\Delta$, considering $\mathcal{O}(\Delta,G)$ in place of $\mathcal{O}(G,\Delta)$, we have what is known as the Lempert function of $D$ denoted $l_{D}$.
+For the precise definition of this invariant function as well as other invariant metrics and distances alongwith various of their fundamental properties particularly of the Kobayashi metric, we refer to the best known book on the subject [Jrncki_invrnt_dst]; here we shall only recall to the extent of relevance to retracts and convenience of later use in this article.
+Thus, we only mention that we can similarly define complex $K$-geodesic (i.e., complex Kobayashi geodesic) if we replace $c_{D}$ by $K_{D}$ in the above definition.
+By Lemma 2.6.1 (ii) in [Abate_Itrtn], every complex $C$-geodesic is a complex $K$-geodesic.
 But the converse is not true because there are domains in $\mathbb{C}^{N}$ which are Kobayashi hyperbolic but not Caratheodory hyperbolic.
 Owing to the direct relationship between $C$-geodesics and retracts, we shall mostly work with $C$-geodesics rather than $K$-geodesics.
-E. Vesentini [Vesentini_1], gives the following characterization of the complex geodesics of $D$.
+Vesentini [Vesentini_1], gives the following characterization of the complex geodesics of $D$.
 Proposition 2.27.
@@
 ∎
-From proposition 2.10, we can conclude that not all linear subspaces of bounded balanced domain $D$ are a retract of $D$ unless $D$ is convex.
-One natural question arises: which linear subspaces of bounded balanced domain $D$ does not arises as a retract of $D$.
-The following corollary gives more general result.
-Before going to the statement, first we recall the notion of convexity at a boundary point.
+Recall that by proposition 2.11, not all one-dimensional linear subspaces of bounded balanced domain $D$ are retracts of $D$, unless $D$ is convex.
+This naturally leads to the question about which are those one-dimensional linear subspaces of bounded balanced non-convex domain $D$ which fail to be a retract of $D$?
+We next lay down a corollary that answers this; its statement requires the notion of convexity at a boundary point, whose definition we recall from [Jrncki_invrnt_dst].
 Definition 2.30.
-Let $D=\{z\in\mathbb{C}^{N}:\mathfrak{h}(z)<1\}$ be a balanced domain in $\mathbb{C}^{N}$ ($\mathfrak{h}$ is the Minkowski function on $D$) and $a\in D$.
-$D$ is said to be convex at $\frac{a}{\mathfrak{h}(a)}$ if there exists a $\mathbb{C}$- linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ with $|L|\leq\mathfrak{h}$ and $|L(a)|=\mathfrak{h}(a)$.
+Let $D=\{z\in\mathbb{C}^{N}:h(z)<1\}$ be a balanced domain in $\mathbb{C}^{N}$ where as usual $h$ denotes the Minkowski functional of $D$.
+Let $a$ be any point of $D$, other than the origin such that $D$ is bounded in the direction of $a$; note that this forces $h(a)$ to be non-zero.
+The domain $D$ is said to be convex at the boundary point $a/h(a)$, if there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ with $|L|\leq h$ and $|L(a)|=h(a)$.
+Note that the latter (equality) condition may be written as $|L(p)|=1$ where $p=a/h(a)\in\partial D$.
+When the aforementioned pair of conditions hold, we also say that $\partial D$ is convex at the boundary point $p$ (or that the balanced domain $D$ is convex near the boundary point closest in the direction of $a$).
+The boundary $\partial D$ (or the domain $D$ itself) is said to be convex near $a/h(a)$ if there exists an open ball $B$ around $a/h(a)$ such that $B\cap\partial D$ is convex at every $q\in B\cap\partial D$.
+One may want to relate the above notion of convexity with the more standard geometric one whose requirement does not involve its boundary points explicitly.
+Namely, we recall that $D$ is said to be geometrically convex near the boundary point $p$ if there exists an open set $U$ about $p$ such that $U\cap D$ is convex in the usual sense that for every pair of points $p_{1},p_{2}\in U\cap D$, we have $(1-t)p_{1}+tp_{2}\in U\cap D$ for all $t\in[0,1]$.
+The relation between the above pair of notions of convexity will be discussed in section 6.
+Here we shall clarify their connection with another standard one namely, the analogue of the above definition 2.30 obtained by restricting to real scalars instead of complex ones.
+Definition 2.31.
+Let $D=\{z\in\mathbb{C}^{N}:h(z)<1\}$ be a balanced domain in $\mathbb{C}^{N}$ where as before $h$ denotes the Minkowski functional of $D$.
+Let $a$ be any point of $D$, other than the origin such that $D$ is bounded in the direction of $a$ i.e., $h(a)\neq 0$.
+The domain $D$ is said to be convex at the boundary point $a/h(a)$, if there exists a $\mathbb{R}$-linear functional $f:\mathbb{C}^{N}\to\mathbb{R}$ with $|f|\leq h$ and $|f(a)|=h(a)$.
+We note that this definition can be made for any real vector space, of-course $\mathbb{R}^{n}$ being the case of our interest.
+Proposition 2.32.
+Let $D$ be a balanced domain in $\mathbb{C}^{N}$. The above two definitions 2.30 and 2.31 are equivalent.
+Proof.
+Assume that $D$ is convex at $a/h(a)$ in the sense of definition 2.30.
+Then there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(a)|=h(a)$.
+Consider the real linear functional $:\mathbb{C}^{N}\to\mathbb{R}$ defined by $T(z)=Re(\alpha L(z))$, where $\alpha=|L(a)|/L(a)$.
+Note that $|\alpha|=1$ and we obtain that
+$|T(z)|=|Re(\alpha L(z))|\leq|\alpha||L(z)|=|L(z)|\leq h(z).$
+Hence we obtain that $|T(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|T(a)|=Re(\alpha L(a))=|L(a)|=h(a)$. Therefore $D$ is convex at $a/h(a)$ according to definition 2.31.
+Conversely, assume that $D$ is convex at $a/h(a)$ as per definition 2.31.
+Then there exists a $\mathbb{R}$-linear functional $f:\mathbb{C}^{N}\to\mathbb{R}$ with $|f(z)|\leq h(z)$ and $|f(a)|=h(a)$.
+Define a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ by $L(z)=f(z)-if(iz)$.
+Observe that for each $z\in\mathbb{C}^{N}$, there exist $\lambda_{z}\in\mathbb{C}$ with $|\lambda_{z}|=1$ such that $\lambda_{z}L(z)=|L(z)|$.
+Note that
+$|L(z)|=|Re(\lambda_{z}L(z))|=|Re(L(\lambda_{z}z))|=|f(\lambda_{z}z)|\leq h(\lambda_{z}z)=|\lambda_{z}|h(z)=h(z)$
+Hence we get that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$.
+In-particular, $|L(a)|\leq h(a)$.
+But then the very definition of $L$ shows that $|L(a)|\geq h(a)$ and we thereby get that $|L(a)|=h(a)$ (and $ia$ lies in the kernel of $f$ though this observation is not necessary for the proof per se).
+Hence we have proved that $D$ is convex at $a/h(a)$ according to definition 2.30.
+∎
+The next pair of results show that the existence (or non-existence) of retracts of a balanced pseudoconvex domain through the origin in any given direction is in essence, completely determined by the convexity of the nearest boundary point in that direction, as clarified by the foregoing definitions.
 Corollary 2.33.
-Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that D is not convex at $a/\mathfrak{h}(a)$.
-Then there does not exists one dimensional retract of $D$ passing through 00 and $a$.
+Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is not convex at $a/h(a)$.
+Then there does not exist any one-dimensional retract of $D$ passing through 00 and $a$.
 Proof.
-Assume the contrary, there exists a one dimensional retract $R$ of $D$ passing through 00 and $a$.
+Assume the contrary, there exists a one-dimensional retract $R$ of $D$ passing through 00 and $a$.
 By theorem 2.29, $R$ is image of some complex C-geodesic $\varphi:\Delta\to D$.
@@
 This leads to a contradiction.
+∎
+Proposition 2.34.
+Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is convex at $a/h(a)$.
+Then $\varphi:\Delta\to D$ defined by $\varphi(\lambda)=\lambda a/h(a)$ is a complex $C$-geodesic on $D$.
+In particular, there always exists a one dimensional linear retract of $D$ passing through 00 and $a$.
+Proof.
+Let $\lambda$ be any non-zero element in $\Delta$. Since $D$ is convex at $a/h(a)$ and by proposition 2.3.1(b) in [Jrncki_invrnt_dst], we get
+$c_{D}^{*}(\varphi(0),\varphi(\lambda))=c_{D}^{*}(0,\lambda a/h(a))=h(\lambda a/h(a))=|\lambda|$
+Note that $c_{D}=\tanh^{-1}c_{D}^{*}$. So
+$c_{D}(\varphi(0),\varphi(\lambda))=\tanh^{-1}\left(c_{D}^{*}(\varphi(0),\varphi(\lambda))\right)=\tanh^{-1}(|\lambda|)=p(0,\lambda)$
+By proposition 11.1.4 in [Jrncki_invrnt_dst], $\varphi$ is a complex $C$-geodesic. ∎
+We spell out the following immediate consequence as it settles a first basic question about the very existence of any non-trivial retract in bounded balanced domains of holomorphy of central interest of this article.
+Proposition 2.35.
+If $D$ is an arbitrary bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, then there exists atleast one (non-trivial) retract passing through origin.
+Proof.
+Let $a\in\partial D$ be the farthest point from the origin.
+Note that $D$ is convex at $a$.
+By the above proposition, there always exists a one-dimensional linear retract of $D$ passing through $a$.
+Alternatively, we may observe that the orthogonal projection onto ${\rm span}(a)$ maps $D$ onto it and thereby a retraction map on $D$.
+The aforementioned retract is the image of this retraction map.
 ∎
@@
 Proposition 2.37.
-[…]
+Let $D$ be a bounded taut domain in $\mathbb{C}^{N}$.
+Assume that $f:D\to D$ be a holomorphic map which is not an automorphism and $\text{Fix}(f)$ is a non-empty subset of $D$ which consists of more than one point.
+Then there exists a non-trivial retract $Z$ of $D$ such that ${\rm Fix}(f)\subset Z$
 Proof.
@@
 ∎
-3. Proofs of theorems 1.3, 1.8, 1.10, 1.12
+3. Proofs of theorems 1.3, 1.4, 1.8, 1.13, 1.15
 We would like to admit at the outset that the proof of theorem 1.15 on retracts of polyballs, closely follows the recent work [BBMV-union]; but we give here a thorough treatment even if, at the cost of repetition of essential ideas, for it allows for a better exposition of the proof of both this theorem, as well as perhaps more importantly: proofs of theorems prior to it as in the title of this section, which are fundamental to this article.
@@
 To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where []delimited-[][$\cdot$][ ⋅ ] denotes the greatest integer function.
-They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extermality, in the following.
+They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extremality, in the following.
 Proposition 3.1.
-If $D_{1},D_{2}$ are domains in $\mathbb{C}^{N}$ with $\mathbb{C}$-extremal boundary, then $D:=D_{1}\cap D_{2}$ is also a domain with $\mathbb{C}$-extremal boundary.
+If $D_{1},D_{2}$ are any pair of domains in $\mathbb{C}^{N}$ with weakly (resp. strongly) $\mathbb{C}$-extremal boundary, then so is every connected component of $D:=D_{1}\cap D_{2}$.
 Proof.
-Let $a$ be an arbitrary point in $\partial D$.
-We want to prove that $a$ is $\mathbb{C}$-extremal point in $D$.
-If possible assume that $a$ is not $\mathbb{C}$-extremal in $D$.
+Let $D_{1},D_{2}$ be pair of domains with weakly $\mathbb{C}$-extremal boundary.
+To prove this result by contradiction, assume $a\in\partial D$ is not weakly $\mathbb{C}$-extremal.
 That is, there exists non-zero $u\in\mathbb{C}^{N}$ and there exists $\epsilon>0$ such that $a+tu\in\partial D~{}\text{for~{} all }~{}t\in\mathbb{C}~{}\text{with}~{}|t|<\epsilon$.
-In particular, $a\in\partial D$ and we shall split the proof into cases depending upon which ‘piece’ of the boundary $a$ lies; herein, piece refers to one of the components in the decomposition of the boundary of $D$ as
+We shall split the proof into cases depending upon which ‘piece’ of the boundary $a$ lies; herein, piece refers to one of the components in the decomposition of the boundary of $D$ as
 $\partial{(D_{1}\cap D_{2})}=(\partial D_{1}\cap D_{2})\cup(D_{1}\cap\partial D_{2})\cup(\partial D_{1}\cap\partial D_{2})$
@@
 Since $u\neq 0$, $\alpha$ is a non-constant holomorphic map.
-Note that $\alpha$ is a homeomorphism from $\Delta$ onto $\alpha(\Delta)$.
+Note that $\alpha$ is a homeomorphism from $\Delta$ onto $\alpha(\Delta)$, which is in $\partial D$.
 Consider the sets $S_{1}:=\{t\in\Delta:\alpha(t)\in\partial{D_{1}}\}$ and $S_{2}:=\{t\in\Delta:\alpha(t)\in\partial{D_{2}}\}$.
 By above decomposition, $S_{1}\cup S_{2}=\Delta$ and note that $S_{1}=\alpha^{-1}(\partial D_{1})$ and $S_{2}=\alpha^{-1}(\partial D_{2})$.
-Again by continuity of $\alpha$, $S_{1}$ and $S_{2}$ are closed sets in $\Delta$.
+Note that $\partial D_{i}\cap\partial D$ is a closed subset of $\partial D$ for $i=1,2$.
+By continuity of $\alpha$, $S_{1}$ and $S_{2}$ are closed sets in $\Delta$.
 We want to prove that $S_{1}$ or $S_{2}$ has atleast one interior point.
@@
 $S_{2}=\overline{S_{2}}=\overline{\Delta\setminus S_{1}}=\Delta\setminus int(S_{1})=\Delta$
-So we get $S_{2}=\Delta$. Hence we can conclude that $S_{1}$ or $S_{2}$ has atleast one interior point.
+So, we get $S_{2}=\Delta$. Hence we can conclude that $S_{1}$ or $S_{2}$ has atleast one interior point.
 Without loss of generality assume that $S_{1}$ has an interior point.
@@
 By definition of $S_{1}$, we obtain that $b+su\in\partial D_{1}$ for all $s\in\mathbb{C}$ with $|s|<\delta$.
-Since $D_{1}$ is $\mathbb{C}$-extremal, we have $u=0$.
-This contradicts our assumption that $a$ is not a $\mathbb{C}$-extremal point of $D$.
-Hence we conclude that $a$ is $\mathbb{C}$-extremal.
+Since $D_{1}$ is weakly $\mathbb{C}$-extremal, we have $u=0$.
+This contradicts our assumption that $a$ is not a weakly $\mathbb{C}$-extremal point of $\partial D$.
+Hence we conclude that $a$ is weakly $\mathbb{C}$-extremal.
+Next, assume that $D_{1},D_{2}$ are domains in $\mathbb{C}^{N}$ with strongly $\mathbb{C}$-extremal boundary.
+Let $a\in\partial(D_{1}\cap D_{2})$.
+If possible assume that there exists $\epsilon>0$ and $u\in\mathbb{C}^{N}$ such that $a+tu\in\partial(D_{1}\cap D_{2})$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$.
+Note that $\partial D_{1}\cap\partial D_{2}\subset\partial(D_{1}\cap D_{2})$.
+Hence $\overline{D_{1}\cap D_{2}}=\overline{D_{1}}\cap\overline{D_{2}}$.
+If $a+tu\in\overline{D_{1}\cap D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$, then $a+tu\in\overline{D_{1}}\cap\overline{D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$.
+Since $D_{1}$ and $D_{2}$ are domains with strongly $\mathbb{C}$-extremal boundary, we have $u=0$.
+Therefore, $D_{1}\cap D_{2}$ is a domain with strongly $\mathbb{C}$-extremal boundary.
 ∎
@@
 As mentioned in the introduction, this theme of retracts of polyballs was initiated earlier in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, what the retracts of polyballs look like; the purpose here is to address this as systematically as we can.
-Thus in particular, by the above theorem 2.1, all retracts of the domain of our interest here, namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, retracts of our convex domain $D$ are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^2] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract.
+Thus in particular, by theorem 2.1 all retracts of the domain of our interest here, namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, retracts of our convex domain $D$ are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^3] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract.
 Infact, there are contractible spaces which do not admit a strong deformation retraction to any of its points, though obviously they admit a deformation retraction to each of its points.
@@
 If we drop the $\mathbb{R}$-extremal condition in the above theorem, then the conclusion about the linearity of $f$ may well fail, as shown by the simplest non-linear holomorphic map $f(z)=(z,z^{2})$ of the disc into the bidisc $\Delta^{2}$; indeed, note that $df(z)=(1,2z)$, so $df(0)=(1,0)$ thereby $\left\|df(0)\right\|_{l^{\infty}}=1$ verifying all other conditions except for the $\mathbb{R}$-extremal condition which we dropped.
+The following theorem is the generalization of above Mazet theorem, which is also useful in this sequel.
+Theorem 3.5.
+(Theorem 3.7 of [Mazet_Rextrm]) Let $E$ and $F$ be Banach spaces and $B_{E},B_{F}$ be the corresponding unit ball. Consider a holomorphic map $f:B_{E}\to B_{F}$ be a holomorphic map. Then we have:
+  1. i)
+    If there exists $u$ of norm 1 in $E$ such that $f^{\prime}(0)\cdot u$ be $\mathbb{R}$-extremal in $\overline{B_{F}}$ then $f(0)=0$.
+  2. ii)
+    If, for any $u$ of norm 1 in $E,f^{\prime}(0)\cdot u$ is $\mathbb{C}$-extremal in $B_{F}$ then $f$ is a polynomial of degree at most 2.
+  3. iii)
+    If the two previous hypotheses are simultaneously verified, then $f$ is a linear map, $f(u)=$ $f^{\prime}(0)\cdot u$
 Though we shall continue making general observations for use in the sequel, we will once and for all now, specialize our ambient space from general complex manifolds or arbitrary Banach spaces to $\mathbb{C}^{N}$ equipped possibly with a norm different from the usual $l^{2}$-norm, so as to finish noting the implications of general facts for our specific case, without much digressions from this purpose.
@@
 Infact, $D_{L}$ is a linear retract of $D$ in the sense of definition 2.6.
-We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.6 and 2.7, allowing us to skip details here.
+We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.7 and 2.8, allowing us to skip details here.
 Lemma 3.6.
@@
 Then, $D\rho(0)$ is a norm one projection of $(\mathbb{C}^{N},\mu)$ onto $(L,\mu)$, thus mapping $D$ onto $D_{L}:=D\cap L$ and showing in particular that $D_{L}$ is a linear retract of $D$.
-If we relax the convexity condition on $D$ to pseudoconvexity, it is still true that the Minkowski functional of $D$ and the infinitesimal Kobayashi metric of $D$ agree at the origin (cf.
-proposition 3.5.3 of [Jrncki_invrnt_dst] wherein counterexamples to show the failure of the equality when pseudoconvexity is dropped are also given), thereby implying that the Kobayashi indicatrix of $D$ at the origin is $D$ itself; likewise the Kobayashi indicatrix of $D_{L}$ now also a bounded balanced pseudoconvex domain in $L$ is also (a copy of) $D_{L}$ itself.
+If we relax the convexity condition on $D$ to pseudoconvexity, it is still true that the Minkowski functional of $D$ and the infinitesimal Kobayashi metric of $D$ agree at the origin (proposition 3.5.3 of [Jrncki_invrnt_dst] wherein counterexamples to show the failure of the equality when pseudoconvexity is dropped are also given), thereby implying that the Kobayashi indicatrix of $D$ at the origin is $D$ itself; likewise the Kobayashi indicatrix of $D_{L}$ now also a balanced pseudoconvex domain in $L$ is also (a copy of) $D_{L}$ itself.
 The general property of contractivity of the Kobayashi metric under holomorphic mappings when applied to $\rho$ viewed as a map from $D$ to $D_{L}$, then yields that $D\rho(0)(D)\subset D_{L}$.
@@
 Lemma 3.7.
-Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map.
+Let $D$ be any balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map.
 Let $L$ denote the tangent space to $Z$ at the origin and $D_{L}:=D\cap L$.
 Then, the linear map $D\rho(0)$ is a projection which retracts $D$ onto $D_{L}$.
-In case $Z$ is already a linear subspace of $D$ to begin with, the above lemma still applies to show that a linear subspace of pseudoconvex bounded balanced domain $D$, which is apriori only known to be attainable as the image of a holomorphic (i.e., possibly non-linear) retraction map on $D$, can indeed be obtained as the image of a linear retraction map as well i.e., it is valid to say that $D_{L}$ as in the above lemma is a linear retract of $D$ without any further qualifications.
+As a corollary to this lemma, we not only get an alternative proof of corollary 2.33 – which was only for one-dimensional retracts – but also an improvement wherein there is no restriction on the dimension of the retract.
+Corollary 3.8.
+Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$.
+If there exists a point $p\in\partial D$ such that $D$ is not convex at $p$, then there does not exist a retract of $D$ passing through the origin in the direction of $p$.
+Proof.
+Assume to obtain a contradiction that there exists a retract $Z$ of $D$ passing through the origin in the direction of $p$.
+Let $\rho$ be any one of the retraction maps on $D$ such that ${\rm image}(\rho)=Z$.
+By the previous lemma 3.7, the linear projection map $\pi:=D\rho|_{0}\;:\;\mathbb{C}^{N}\to\mathbb{C}^{N}$ maps $D$ onto $D_{L}$, where $L=T_{0}Z$.
+Consider the affine subspace $K:=p+Ker(\pi)$ of $\mathbb{C}^{N}$.
+Since $D$ is not convex at $p$, there exists a point $q\in K$ such that $q\in D\cap K$.
+But $\pi(q)=p$, which contradicts the fact that $\pi$ maps $D$ into $D$.
+∎
+We return to the preceding lemma for a little clarification that saves some confusion once and for all.
+In case $Z$ is already a linear subspace of $D$ to begin with, the above lemma 3.7, still applies to show that a linear subspace of pseudoconvex bounded balanced domain $D$, which is apriori only known to be attainable as the image of a holomorphic (i.e., possibly non-linear) retraction map on $D$, can indeed be obtained as the image of a linear retraction map as well i.e., it is valid to say that $D_{L}$ as in the above lemma is a linear retract of $D$ without any further qualifications.
 We are now fully prepared to prove the highlighted general result about the equality $\rho(D_{L})=Z$ whose statement is formulated with slightly different notations than the one that we had adopted for our specific domain $D$.
-This is indeed the content of theorem 1.10, whose statement we lay down again here for convenience, and then prove it inspired by an idea in [Abate_Isometry].
+This is indeed the content of theorem 1.13, whose statement we lay down again here for convenience, and then prove it inspired and based on an idea in [Abate_Isometry].
 Theorem 3.9.
@@
 Let $L$ denote the tangent space to $Z$ at the origin.
-Then, $\rho_{|D_{L}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a (single) holomorphic function over $D_{L}$.
+Then, $\rho_{|D_{L}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$ taking values in $M:=\ker\left(D\rho(0)\right)$.
 Proof.
@@
 ∎
+While the above theorem allows retracts to be non-linear and we shall indeed furnish several examples of non-product (balanced, pseudoconvex) domains possessing non-linear retracts through origin in the form of graphs, we recall that such a phenomenon is precluded by Vesentini’s theorem in the convex case.
+Extremality boundary assumptions as in Vesentini’s theorem are indeed necessary as we shall see while establishing a converse to his theorem later.
+However, the hypothesis that the entire boundary be holomorphically extreme as in the extension of Vesetini’s theorem given by proposition 1.4 in [JJ] is inconvenient for discussing interesting intermediate examples in the sequel.
+This is addressed by proposition 1.4 which we now establish.
+Though this is in essence an imitation of the proof of the aforementioned proposition in [JJ], we prefer to include some details as this is a result about retracts that we shall need later in this article wherein the aforementioned relaxation will actually be used.
+And, also because it provides an answer towards the natural query as to whether an extension of Vesentini’s theorem holds if we add only an extremality hypothesis just as much as is needed about the boundary, but none about convexity in the setting of the above theorem 3.9.
+Proof of proposition 1.4: Choose any holomorphic retraction map $\rho:D\to D$ with $\rho(D)=Z$.
+We may observe by the above theorem 3.9, that the restriction $\rho|_{D_{L}}$ is a holomorphic map on $D_{L}$ whose image is precisely equal to $Z$; but we shall not need this for the rest of this proof here.
+Recall that $Z$ is a connected complex submanifold and the tangent space to $Z$ at the origin, is denoted by $L$; so, $L=T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}|_{D_{L}}(0)v=v\}$.
+Let $v\in D_{L},v\neq 0$.
+By our assumption on $\partial D_{L}$, it follows that the mapping
+$\phi:\Delta\ni\lambda\longmapsto\lambda v/h_{D_{L}}(v)\in D_{L}.$
+is the unique (modulo $Aut(\Delta)$) $K_{D_{L}}$-geodesic for $(0,v/h_{D_{L}}(v))$; here of-course $h_{D_{L}}$ denotes the Minkowski functional of the balanced domain $D_{L}\subset L$.
+Since $D_{L}$ is a linear subspace of $D$, $h_{D_{L}}(v)=h_{D}(v)$.
+So $\phi$ is the unique (modulo $Aut(\Delta)$) $K_{D}$-geodesic for the datum $(0,v/h_{D_{L}}(v))$.
+Note that
+$K_{D}\left(\rho|_{D_{L}}\circ\phi(0);(\rho|_{D_{L}}\circ\phi)^{\prime}(0)\right)=K_{D}\left(0,v/h_{D_{L}}(v)\right)=K_{D}\left(\phi(0);\phi^{\prime}(0)\right).$
+Hence $\rho|_{D_{L}}\circ\phi$ is a $K_{D}$-geodesic for $(0,v/h_{D_{L}}(v))$.
+Because of our hypothesis, we have that $v/h_{D_{L}}(v)$ is a holomorphic extreme point of $D$.
+By uniqueness of complex geodesics in $D$ (exposited as proposition 11.3.5 in [Jrncki_invrnt_dst]), it follows that $\rho|_{D_{L}}\circ\phi=\phi\circ\psi$, where $\psi\in{\rm Aut}(\Delta)$.
+Substituting $\psi^{-1}(h_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho|_{D_{L}}$.
+Hence $D_{L}\subset Z$.
+Now $T_{0}Z$ and $Z$ have the same dimension.
+Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$.
+Hence $Z$ is linear.
+∎
+Some remarks about the foregoing pair of results are in order.
 Remark 3.10.
-Under the hypotheses that the boundary $\partial B$ is holomorphically extreme, it was shown in [JJ] that $Z$ is actually linear.
-While this hypothesis is the best possible for obtaining such a strong conclusion, it is restrictive as well for, there are plenty of balanced pseudoconvex domains whose boundaries fail to satisfy such a strong extremality condition.
+To indicate just a few concrete examples to illustrate where proposition 1.4 may be applied (but not the other results in this subject directly at-least), we first deal with domains where the holomorphic extremality condition is neither violated maximally nor minimally; examples for ‘maximal’ violation is also considered but in section 8.
+Among the simplest of domains satisfying the hypotheses of proposition 1.4 in this intermediate fashion, we consider the domain $\Omega_{1}:=\Delta^{2}\cap D_{q}$, where $D_{q}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<2^{q}\}$ for $10<q<$ 1 is a dilated copy of the $\ell_{q}$-ball.
+Note that $\Omega_{1}$ is a bounded balanced pseudoconvex domain and $\partial\Omega_{1}$ fails to be holomorphically extreme (in its entirety); though of-course, there exists an open piece of $\partial\Omega_{1}$, which is holomorphically extreme.
+A combined application of theorem 1.10, theorem 1.13 and corollary 2.33, proves that every retract passing through origin is either a linear subspace of $\Omega_{1}$ or a graph of a non-linear holomorphic function over some linear subspace of $\Omega_{1}$.
+The above example was non-convex; so, next we consider a convex domain just by taking its limiting case i.e., letting $q\to 1$ in the above namely, $\Omega_{2}:=\Delta^{2}\cap D_{1}$, where $D_{1}:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<3/2\}$ is some concrete dilated copy of the $\ell_{1}$-ball.
+Note that $\partial\Omega_{2}$ is not strictly convex, indeed not holomorphically extreme.
+Though no little piece of the boundary is $\mathbb{R}$-extremal, there exists an open piece of the boundary which satisfies the $\mathbb{C}$-extremality condition.
+By proposition 1.4 whose applications we are after, we obtain that every retract through the origin is either a linear subspace of $\Omega_{2}$ or a graph of a non-linear holomorphic function over some linear subspace of $\Omega_{2}$.
+Remark 3.11.
+While it is clear that without any extremality assumptions, linearity of retracts is unreasonable to be expected (polydisc being the simplest ‘counterexample’) and, holomorphic extremality seems the best possible condition for obtaining such a strong conclusion as in the proposition 1.4 just discussed, it is restrictive as well.
+For, there are plenty of balanced pseudoconvex domains whose boundaries fail to satisfy such a strong extremality condition.
 To give just one example whose boundary does not contain any flat pieces, which indeed satisfies the weaker $\mathbb{C}$-extremality but not the holomorphic extermality condition, consider the bounded balanced pseudoconvex domain obtained by intersecting the domain $H=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/2\}$ with the standard Euclidean ball $\mathbb{B}$ in $\mathbb{C}^{2}$.
-This indicates how we may construct plenty of such examples.
+This indicates how we may construct plenty of such examples which also we shall deal later in the sequel at appropriate places; in-particular, subsequent to the next remark.
 Remark 3.12.
-Being a graph, $Z$ is easily seen to be biholomorphic to $D_{L}$ as complex manifolds, in the setting of the above theorem.
+In the setting of the foregoing proposition 1.4, being a graph, $Z$ is easily seen to be biholomorphic to $D_{L}$ as complex manifolds, in the setting of the above theorem.
 Thereby, it immediately follows that any pair of retracts through the origin in $D$, of the same dimension are diffeomorphic.
-Indeed, being star-shaped domains in $L$, they are diffeomorphic to $L$ – no pseudoconvexity is required here; the general fact being used here is that any star-shaped domain in $\mathbb{R}^{N}$ is $C^{\infty}$-diffeomorphic to $\mathbb{R}^{N}$ (cf.
-exercise $7$ of chapter $8$ of [Brocker-Janich-difftop-bk]).
+Indeed, being star-shaped domains in $L$, they are diffeomorphic to $L$ – no pseudoconvexity is required here; the general fact being used here is that any star-shaped domain in $\mathbb{R}^{N}$ is $C^{\infty}$-diffeomorphic to $\mathbb{R}^{N}$ (exercise $7$ of chapter $8$ of [Brocker-Janich-difftop-bk]).
 It is therefore of interest then (as is in Several Complex Variables), to note that such a pair of retracts (through the origin of the same balanced domain $D$ and) of the same dimension despite being mutually diffeomorphic need not be biholomorphic (indeed, they are rarely so).
-To see this first by what is perhaps the simplest example, consider $D=\Delta^{2}\times\mathbb{B}^{2}$; let $Z_{1}$ denote the projection of $D$ onto the span of the first pair of standard basis vectors of $\mathbb{C}^{4}$ and $Z_{2}$, the projection of $D$ onto the span of the remaining last pair of standard basis vectors of $\mathbb{C}^{4}$.
-Noting that $Z_{1}\simeq\Delta^{2}$, $Z_{2}\simeq\mathbb{B}^{2}$ and recalling that the ball and polydisc are biholomorphically inequivalent, we deduce the same about $Z_{1},Z_{2}$.
+To see this first by what is perhaps the simplest example, consider $D=\Delta^{2}\times\mathbb{B}^{2}$ as a subdomain of $\mathbb{C}^{4}$; let $Z_{1}$ denote the projection of $D$ onto the span of the first pair of standard basis vectors of $\mathbb{C}^{4}$ and $Z_{2}$, the projection of $D$ onto the span of the remaining last pair of standard basis vectors of $\mathbb{C}^{4}$.
+By the well-known biholomorphic inequivalence of the ball and polydisc, $Z_{1}\not\simeq Z_{2}$ as $Z_{1},Z_{2}$ are just copies of the bidisc $\Delta^{2}$ and the ball $\mathbb{B}^{2}$ respectively.
 To give another example, again as simple as possible but this time of a ball $D$ which is not biholomorphic to a product domain, in-fact with boundary both smooth as well as holomorphically extreme and, which contains pairs of linear retracts (thereby through the origin), of the same dimension that are biholomorphically inequivalent, consider
@@
 :=$\mathcal{E}:=\{z\in\mathbb{C}^{N}\;:\;|z_{1}|^{p_{1}}+\ldots+|z_{N}|^{p_{N}}<1\},$
-where the exponents $p_{j}$’s are taken to be bigger than one (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$).
+where the exponents $p_{j}$’s are taken to be bigger than one and distinct (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$).
 Then, for every $1<k<N$, there exist a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent.
@@
 But then again, we shall show in the next theorem that the retracts through the origin are all linear.
-While the foregoing theorem 3.7, already realizes $Z=\rho(D_{L})$ as the graph of a holomorphic function over $L$, it must be noted that this does not mean that the graph of any holomorphic function over any given linear retract of $D$, is a retract: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well.
+The converse of above theorem 3.9 is also true.
+Proposition 3.14.
+Let $D$ be a bounded balanced pseudoconvex domain and $D_{L}:=D\cap L$ be a linear retract of $D$.
+Let $\varphi$ be any holomorphic map from $D_{L}$ to $M$, where $M$ is a linear subspace complementary to $L$ i.e., $\mathbb{C}^{N}=L\oplus M$, is a holomorphic map with the property that $(w,\varphi(w))\in D$, whenever $w\in D_{L}$.
+Then ${\rm Graph}(\varphi)$, the graph of $\varphi$ over $D_{L}$ in $D$, is a retract of $D$.
+Proof.
+Since $D_{L}$ is a linear retract of $D$, there exists a linear projection map $P$ such that $image(P)=D_{L}$.
+By idempotency of $P$, we can write $\mathbb{C}^{N}=L\oplus M$, where $M=\ker(P)$ and $L={\rm image}(I-P)$.
+Consider the map $\psi:D\to D$ defined by $\psi(z)=(T(z),\varphi(T(z)))$; this can be verified to be idempotent by a direct computation, thereby showing that $\psi$ is a holomorphic retraction from $D$ onto ${\rm Graph}(\varphi)$ over $D_{L}$, finishing the proof.
+∎
+It must be noted that the above does not mean that non-linear retracts automatically come into existence by taking the graph of any holomorphic function over any given linear retract of $D$ for, the graph can fail to be contained within $D$; that this is the only obstruction to generating non-linear retracts by graphs is the content of the above proposition.
+However, the obstruction is substantial: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well.
 The proof of this, atleast as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true for every bounded balanced convex domain with $\mathbb{C}$-extremal boundary.
@@
 Note that $(D\rho|_{0})_{|_{D_{L}}}:D_{L}\longrightarrow D_{L}$ is the identity map on $D_{L}$ and that $D_{L}$ is a bounded balanced convex domain in $\mathbb{C}^{N}$.
-It follows from the proof of Krein-Milman theorem (the reader is referred to the book [Meggn] for the standard proof we have in mind here), that $D_{L}$ must have atleast one boundary point $\mathbb{R}$-extremal point, say $z_{0}$.
+It follows from the proof of Krein-Milman theorem (the reader is referred to the book [Meggn] for the standard proof we have in mind here), that $D_{L}$ must have atleast one $\mathbb{R}$-extremal boundary point, say $z_{0}$.
 As $D\rho(0)$ acts as the identity map on $D_{L}$ (thereby in particular $D\rho(0)(z_{0})=z_{0}$), it follows that there is atleast one $\mathbb{R}$-extremal point in $D\rho(0)(D_{L})$ namely $z_{0}$ itself!
-Theorem 3.7 part-(iii) of Mazet’s article [Mazet_Rextrm] applies to the map $\rho_{|_{D_{L}}}$ finishes the proof that $Z$ is linear.
+To be precise $z_{0}$ is an $\mathbb{R}$-extremal point of $D$ not just $D_{L}$.
+Applying theorem 3.5 to the map $\rho_{|_{D_{L}}}$ viewed as a map from $D_{L}$ into the (larger) domain $D$, finishes the proof that $Z$ is linear.
 To not be amiss to discuss concrete examples and the simplest ones at that, recall that the first standard choices for $D$, namely the $l^{p}$-balls was already mentioned in the introduction, specifically theorem 1.7.
@@
 […]
-$L=\Delta^{N}\cap Y=\{(F_{1}(z),\ldots,F_{N}(z)):z\in M\}$
+$L=\Delta^{N}\cap Y=\{(F_{1}(z),\ldots,F_{N}(z)):z\in M\}.$
 Observe that the $F_{j}$’s are holomorphic functions from $\Delta^{N}$ to $\Delta$ and $F_{j}(z)=G_{j}(z_{j_{1}},\ldots,z_{j_{r}})$ for some holomorphic function $G_{j}$, which is bounded by $1$.
 Without loss of generality assume that each $F_{i}$’s are linear, since $F(z)=(F_{1}(z),\ldots,F_{N}(z))$ restricted to $L$ is linear.
-So we get $F_{i}(z)=z_{i}$, if $i\in J$ and $F_{i}(z)=c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}}$, if $i\notin J$ for some scalars $c_{i,j},~{}j\in J$.
+So, we get
+$F_{i}(z)=\begin{cases}z_{i},\quad\qquad\qquad\qquad\qquad\text{if}~{}i\in J\\ c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}},\quad\text{if}~{}i\notin J,\\ \end{cases}$
+for some scalars $c_{i,j},~{}j\in J$.
 Fix $m\notin J$ and for $1\leq k\leq N$, define
-$w_{k}:=\begin{cases}\frac{|c_{m,j_{r}}|}{c_{m,j_{r}}},\ \text{if}\ k=j_{r}~{}\text{for some}~{}r\ \text{and}\ c_{m,j_{r}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}$
+$w_{k}:=\begin{cases}\frac{|c_{m,j_{r}}|}{c_{m,j_{r}}},\ \text{if}\ k=j_{r}~{}\text{for some}~{}r\ \text{and}\ c_{m,j_{r}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}.$
 We know that $|F_{m}(z)|\leq 1$ for all $z\in\Delta^{N}$. If $w:=(w_{1},\ldots,w_{N})$ then
@@
 Hence we get $\sum_{j\in J}|c_{m,j}|\leq 1$.
-Note that $M$ is spanned by $\{e_{j_{1}},\ldots,e_{j_{r}}\}$.
+Recall that $M$ is spanned by $\{e_{j_{1}},\ldots,e_{j_{r}}\}$.
 Since $F$ is determined by the variables $z_{j_{1}},\ldots,z_{j_{r}}$, we get $Y$ is spanned by $\{F(e_{j_{1}}),\ldots,F(e_{j_{r}})\}$.
 By definition of $F_{i}$, we get
-$F_{i}(e_{j_{k}})=\begin{cases}1,\quad\qquad\text{if}~{}i\in J\\ c_{i,j_{k}},\qquad\text{if}~{}i\notin J,\\ \end{cases}$
-By definition of $F$, we will get
-$F(e_{j_{k}})=(F_{1}(e_{j_{k}}),\ldots,F_{N}(e_{j_{k}}))=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$
+$F_{i}(e_{j_{k}})=\begin{cases}\delta_{i,j_{k}},\quad\qquad\text{if}~{}i\in J\\ c_{i,j_{k}},\qquad\text{if}~{}i\notin J,\\ \end{cases}$
+where $\delta_{i,j}$ denotes the Kronecker delta function.
+By definition of $F$, we get
+$F(e_{j_{k}})=(F_{1}(e_{j_{k}}),\ldots,F_{N}(e_{j_{k}}))=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}.$
 This shows that $Y=span\{v_{k}\}_{k=1}^{r}$ where $v_{k}=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$.
@@
 Note that $F_{i}(z)$ is a holomorphic function from $\Delta^{N}$ to $\Delta$, because by condition (ii) of the theorem.
-Next we define $F(z):=(F_{1}(z),\ldots,F_{N}(z))$.
-$F$ is a holomorphic retraction map on $\Delta^{N}$ whose image is $span\{v_{k}\}_{k=1}^{r}\cap\Delta^{N}$, where $v_{k}=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$, which completes the proof.
-We already saw in the introduction section, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors.
+Consider the map $F$ defined in terms of these $F_{i}$’s by $F(z):=(F_{1}(z),\ldots,F_{N}(z))$.
+It can be readily checked that $F$ is a holomorphic retraction map on $\Delta^{N}$ whose image is $span\{v_{k}\}_{k=1}^{r}\cap\Delta^{N}$, where $v_{k}=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$, which completes the proof.
+As noted above, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors.
 The converse implication however is true and can now be derived as a corollary.
@@
 The foregoing remarks for a product of Euclidean balls apply as well to the case of products of balls with polydiscs such as $\Delta^{k}\times\mathbb{B}$ (or a larger collection of Euclidean balls and polydiscs in any order) which are also homogeneous domains with all of is automorphisms being fractional-linear maps.
-We conclude by noting that in these cases the characterization of retracts provided by theorems 1.12 and 3.16, is complete.
+We conclude by noting that in these cases the characterization of retracts provided by theorems 1.15 and 3.20, is complete.
 4. Proof of theorem 1.16
+The proof follows ideas from the artice [Fost] of Fornaess–Sibony.
 Let us begin by noting that it follows from theorem 1.13 proved in the foregoing section that every retract of a bounded balanced homogeneous domain is biholomorphic to the graph of some holomorphic function over a linear subspace of $D$.
-Since $\Omega$ is homogenous, $\Omega/Aut(\Omega)$ is singleton and hence compact.
+Since $\Omega$ is homogeneous, $\Omega/{\rm Aut}(\Omega)$ is singleton and hence compact.
 By the Main theorem in [Fost], there exists a closed complex submanifold $Z$ of $\Omega$ such that M is biholomorphic to a locally trivial holomorphic fibre bundle with fibre $\mathbb{C}$ over $Z$.
-As mentioned in the begining, by theorem 1.10 we get that Z is biholomorphic to a linear retract $\Omega_{L}:=\Omega\cap L$ for some linear subspace $L$ of $\mathbb{C}^{N}$.
+As mentioned in the beginning, by theorem 1.13 we get that Z is biholomorphic to a linear retract $\Omega_{L}:=\Omega\cap L$ for some linear subspace $L$ of $\mathbb{C}^{N}$.
 We can cover $M$ by coordinate charts biholomorphic to $\{U_{\alpha}\times\mathbb{C}\}$ where $U_{\alpha}$ is an open cover of $Z$.
@@
 The data $h_{\alpha\beta}$ determine a multiplicative Cousin problem on $Z$.
-Since $Z$ is homeomorphic to polydisk and Oka’s theorem in [Oka](exposited in theorem 3.9 of [oka_0] as well), multiplicative Cousin problem for $Z$ can be solved.
+Since $Z$ is homeomorphic to a polydisc, it follows by Oka’s theorem (theorem 3.9 of [oka_0] as well) that the multiplicative Cousin problem for $Z$ can be solved.
 Hence we might have assumed already above that $h_{\alpha\beta}\equiv 1$ for all $\alpha,\beta$.
 However in that case, the data $k_{\alpha\beta}$ determine an additive Cousin problem.
-By Oka’s theorem in [Oka_1] (exposited in theorem 3.6 of [oka_0]), additive Cousin problem can be solved.
+By Oka’s theorem in (theorem 3.6 of [oka_0]), the additive Cousin problem on $Z$ can also be solved.
 Hence $M$ is biholomorphic to $Z\times\mathbb{C}$.
 Remark 4.1.
-We recall for one last time, that the class of homogeneous bounded balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (cf.
-the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein).
+We recall for one last time, that the class of homogeneous bounded balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein).
 It follows that a domain $\Omega$ as in the above theorem is a polyball (and convex), in-particular.
 5. Proof of Theorem 1.9
-It must first be noted that by Theorem 1.3, every holomorphic retract through the origin is actually linear i.e., $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$.
+Before discussing our result as in the title of this section on the retracts of $\ell_{q}$ balls for $10<q<$ 1 we first review the case $q\geq 1$.
+The main theorem about this case was already mentioned in theorem 1.7.
+However that formulation is highly coordinate dependent.
+Here we first re-formulate it in-terms of commutavity with standard projections; the proof is based on an idea from the work [Baronti-Pappini] of Baronti – Pappini.
+Theorem 5.1.
+$R$ is a $k$-dimensional retract of the $\ell^{p}$-ball in $\mathbb{C}^{N}$ if and only if one of the following holds:
+  1. (1)
+    There exists a norm-one projection which maps $\mathbb{B}_{p}$ onto $R$ which commutes with the standard orthogonal projection onto one of the $(N-k)$ or $(N-(k-1))$-dimensional subspace spanned by co-ordinate vectors.
+  2. (2)
+    There exist a finite sequence of projections $P_{k},P_{k-1},\ldots,P_{1}$ such that for each $r=1,2,\ldots,k$, $P_{r}$ is a projection of norm one which maps $\mathbb{C}^{N}$ onto an $r$-dimensional subspace and $P_{k}(\mathbb{B}_{p})=R$.
+Proof.
+It is easy to check that one implication of the theorem holds by using the definition of norm one projection.
+So we have only to prove the other implication.
+Assume that $R$ is a $k$-dimensional linear retract of $\mathbb{B}_{p}$.
+By theorem 2.a.4 in [Lindstraus], we obtain that $R=L\cap\mathbb{B}_{p}$ where $L$ is a $k$-dimensional linear subspace of $\mathbb{C}^{N}$ spanned by mutually disjointly supported vectors.
+Hence $L$ can be written as the solution set of $(N-k)$ linear equations, where each linear equation depends on atmost two variables.
+After applying Gaussian elimination and renaming variables, there is no loss of generality assume that above mentioned system of equations can be written as $AZ=0$, where $A$ is a $(N-k)\times N$ matrix $[B~{}B^{\prime}~{}B^{\prime\prime}]^{T}$ and $Z=[z_{1}~{}\ldots~{}z_{N}]^{T}$.
+The submatrices $B,B^{\prime},B^{\prime\prime}$ are of the following form:
+  1. (i)
+    $B=[B_{0}~{}B_{1}~{}\ldots B_{r-1}]^{T}$, where for each $i=0,1,\ldots,r-1$, $B_{i}=[b_{k_{i}+1}\ldots b_{k_{i+1}}]^{T}$ and $0=k_{0}<k_{1}<\ldots<k_{r}$.
+Futhermore for each $i$, $B_{i}$ satisfies the following properties:
+    1. (a)
+      For $j=k_{i}+1$ to $k_{i+1}-1$, $supp(b_{j})\cap supp(b_{j+1})\neq\phi$ and $|supp(b_{j})\cap supp(b_{j+1})|=1$.
+    2. (b)
+      For $i=0,1,\ldots,r-1$, if we define $supp(B_{i}):=\bigcup_{j=k_{i}+1}^{k_{i+1}}supp(b_{j})$ then $supp(B_{i})\cap supp(B_{j})=\phi$ for $i\neq j$.
+  2. (ii)
+    $B^{\prime}$ is of the form $[b_{1}^{\prime}~{}\ldots~{}b_{l}^{\prime}]^{T}$, where $supp(b_{i}^{\prime})\cap supp(b_{j}^{\prime})=\phi$ for $i,j=1,2,\ldots,l$ with $i\neq j$.
+  3. (iii)
+    $B^{\prime\prime}$ is of the form $[e_{s_{0}+1}~{}\ldots e_{s_{1}}]^{T}$. where $s_{0}=k_{r}+r+2l,s_{1}=N-k+r+l$ and $e_{i}$ denotes the standard basis of $\mathbb{C}^{N}$.
+Note that $|supp(B)|=|\bigcup_{i=0}^{r-1}supp(B_{i})|=k_{r}+r$, $|supp(B^{\prime})|=2l$ and $|supp(B^{\prime\prime})|=s_{1}-s_{0}$.
+Since the number of equation is $(N-k)$, we have $k_{r}+l+(s_{1}-s_{0})=N-k$.
+Because of this equation, we take $s_{1}=N-k+r+l$.
+Hence, we obtain the total number of dependent variables in the system of equations $AZ=0$ is equal to
+$|supp(B)|+|supp(B^{\prime})|+|supp(B^{\prime\prime})|$ $=$ $k_{r}+r+2l+(s_{1}-s_{0})$
+$=$ $k_{r}+r+2l+(N-k)-(k_{r}+l)$
+$=$ $N-k+r+l.$
+First we want to construct mutually disjoint supported vectors $(v_{i})_{i=1}^{k}$ such that $(v_{i})$ forms a basis for $L$.
+For each $i=1,\ldots,r$, consider the system of equations $B_{i}Z_{i}=0$, where $Z_{i}=[z_{k_{i-1}+1}~{}\ldots~{}z_{k_{i}}]^{T}$.
+This solution space is spanned by a non-zero vector $w_{i}=(\mu_{k_{i-1}+1},\ldots\mu_{k_{i}})\in\mathbb{C}^{k_{i}-k_{i-1}}$.
+Since every two co-ordinates of $w_{i}$ satisfies a linear equation in two variables, every component of $w_{i}$ is non-zero.
+For each $i=1,\ldots,r$, define $v_{i}:=\mu_{k_{i-1}+1}e_{k_{i-1}+1}+\ldots+\mu_{k_{i}}e_{k_{i}}$, where $e_{i}$ denote the standard basis in $\mathbb{C}^{N}$.
+Note that $v_{i}\in L$ and $supp(v_{i})=k_{i}-k_{i-1}$.
+For $i=r+1$ to $r+l$, choose a non-zero vector $v_{i}$ which is orthogonal to $b^{\prime}_{i-r}$ and $supp(v_{i})=supp(b^{\prime}_{i-r})$.
+For $i=r+l+1$ to $k$, $v_{i}=e_{i+t}$, where $t=N-k$.
+Define $u_{i}:=\frac{v_{i}}{\|v_{i}\|_{p}}$ for $i=1,2,\ldots,k$.
+Note that $(u_{i})_{i=1}^{k}$ forms a mutually disjointly supported vectors and a basis for $L$.
+Observe that $|supp(u_{i})|=2$ for $i=r+1$ to $r+l$ and $|supp(u_{i})|=1$ for $i=r+l+1$ to $k$.
+Using the idea in the proof of theorem 2.a.4 in [Lindstraus], we can construct projection $P$ of norm $1$ from given vectors $u_{1},\ldots,u_{k}$.
+Let $\sigma_{j}:=supp(u_{j})$ for $j=1,2,\ldots,k$.
+If we write $u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i}$, then we define the linear functional $u_{j}^{*}=\sum_{i\in\sigma_{j}}\frac{|\lambda_{i}|^{p}}{\lambda_{i}}e_{i}^{*}$.
+Observe that for $i=1,2,\ldots,r$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{i-1}+1},\ldots,z_{k_{i}}$, for $i=r+1,\ldots,r+l$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{r}+2i-r-1}$, $z_{k_{r}+2i-r}$ and for each $i\neq j$, $u_{j}^{*}$ does not depends on the variables $z_{k_{r}+2i-r-1}$ and $z_{k_{r}+2i-r}$.
+For $i=r+l+1$ to $k$, $u_{i}^{*}(z)=z_{i+t}$.
+Define a projection $P:\mathbb{C}^{N}\to\mathbb{C}^{N}$ by
+$P(z)=\sum_{j=1}^{k}u_{j}^{*}(z)u_{j}.$
+Note that $P$ is a projection of norm $1$ onto $L$. Next we are looking for standard projections which commutes with $P$. For $i=1,\ldots,r$, consider the standard projection,
+$\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
+Note that
+$P\circ\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
+For each $i=1,\ldots,r$, note that $k_{i-1}+1,\ldots,k_{i}$ components of $P$ only depends on $z_{k_{i-1}+1},\ldots,z_{k_{i}}$.
+So, $\pi_{k_{i-1}+1,\ldots,k_{i}}\circ P(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$.
+Let $m_{i}:=k_{r}+2i-r-1$. For $i=r+1,\ldots,r+l$, consider another projection
+$\pi_{m_{i},m_{i}+1}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{m_{i}-1},0,0,z_{m_{i+2}},\ldots,z_{N})$
+Using similar arguments as mentioned above, we can conclude that $P\circ\pi_{m_{i},m_{i}+1}=\pi_{s_{i},s_{i}+1}\circ P$.
+As we noted before, the total number of dependent variables in the system of equation $AZ=0$ is $N-k+r+l$.
+So, we consider another projection $\pi$ defined by $\pi(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{s_{1}},0,\ldots,0)$.
+Note that $P$ commutes with $\pi$ and $rank(\pi)=N-k+r+l$.
+If $r+l\geq 1$, then rank of $\pi$ is atleast $N-(k-1)$.
+If $r+l=0$, then $rank(\pi)$ is atleast $(N-k)$.
+This finishes the proof of (i).
+To prove (ii), take $P_{k}:=P$, which is a projection of norm one defined in the proof of (i).
+From the proof of (i), we obtain a standard projection $\pi_{k-1}$ of rank $(N-1)$ which commute with $P$.
+Let $P_{k-1}:=\pi_{k-1}\circ P$.
+Observe that $P_{k-1}$ is a norm one projection and $\rm image(P_{k-1})$ is a $k$-dimensional is a $(k-1)$-dimensional subspace of $\mathbb{C}^{N}$.
+Again applying the same argument to $P_{k-1}$, we get another projection $\pi_{k-2}$ of rank $(N-1)$ which commutes with $P$.
+Define $P_{k-2}:=\pi_{k-2}\circ P_{k-1}=\pi_{k-2}\circ\pi_{k-1}\circ P$.
+Hence we get that $P_{k-2}$ is a projection of norm one and it’s image is a $(k-2)$ dimensional subspace.
+Continuing this process, until we will obtain a projection $P_{1}$ of norm one such that $\rm image(P_{1})$ is a one dimensional subspace of $\mathbb{C}^{N}$.
+∎
+An apparently even more geometric formulation of the retracts of $\ell_{p}$ balls for $1\leq p<\infty,p\neq 2$ is that the retracts are precisely isometric copies of lower dimensional $\ell_{p}$-balls; see theorem 6.3 of [Beata-survey].
+This is a feature shared by the $\ell_{p}$-‘balls’ even for positive values of the parameter $p<1$.
+Indeed, we now move onto this case dealt with by theorem 1.9.
+Before going into its proof 1.9, we need to make a couple of observations about the geometry of the $\ell_{q}$ ball $D_{q}$ in $\mathbb{C}^{N}$ (for $q<1$).
+Firstly, although its absolute image of is concave (easily and vividly seen in the two-dimensional case), $D_{q}$ is actually not concave (nor convex) domain in $\mathbb{C}^{N}$; not even locally, near any of the smooth boundary points.
+We may establish this by computing the real Hessian at any smooth boundary point with positive real coordinates as $D_{q}$ is multicircular; we show this in the two-dimensional setting for simplicity, while it holds in higher dimensions as well by straight-forward extension of the arguments and computations.
+Rewrite the defining function for $D_{q}$ as
+$\varphi(z,w)=z^{q/2}\overline{z}^{q/2}+w^{q/2}\overline{w}^{q/2}-1$
+To compute the real Hessian, note the following.
+$H$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}\frac{q^{2}}{4}|z|^{q-2}&0\\ 0&\frac{q^{2}}{4}|w|^{q-2},\end{bmatrix}$
+$\tilde{H}$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z^{2}}&\frac{\partial^{2}\varphi}{\partial z\partial w}\\ \frac{\partial^{2}\varphi}{\partial w\partial z}&\frac{\partial^{2}\varphi}{\partial w^{2}}\end{bmatrix}=\begin{bmatrix}\frac{q}{2}(\frac{q}{2}-1)|z|^{q}z^{-2}&0\\ 0&\frac{q}{2}(\frac{q}{2}-1)|w|^{q}w^{-2}\end{bmatrix}$
+Before proceeding, we remark that we shall utilize the above to compute the real Hessian only at the (smooth) absolute boundary points of $D_{q}$, such points being precisely those points $(x,y)$ of $\partial D_{q}$ both of whose coordinates are positive; while this suffices as $D_{q}$ is Reinhardt, it is important to note that the (tangent) vectors along which the Hessian form is evaluated need to be allowed to have non-real (complex) coordinates.
+Let $v=[v_{1}~{}v_{2}]$ and $\overline{v}=[\overline{v_{1}}~{}\overline{v_{2}}]$.
+Then the real Hessian at the point $(x,y)$ evaluated along the direction $v$ is
+(5.1) $vH\overline{v}^{T}|_{(x,y)}+Re(vHv^{T})_{(x,y)}\\ =\frac{q^{2}}{4}\left(x^{q-2}|v_{1}|^{2}+y^{q-2}|v_{2}|^{2}\right)+\frac{q}{2}(\frac{q}{2}-1)Re\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right).$
+If $v_{1},v_{2}$ are purely imaginary, then the above expression (5.1) is strictly positive. If $v_{1}$ and $v_{2}$ are real, then the above expression (5.1) becomes
+$\frac{q^{2}-q}{2}\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right)$
+which is negative as $q<1$.
+We conclude therefore that $D_{q}$ is neither convex nor concave (though of-course pseudoconvex) at $(x,y)\in\partial D_{q}$.
+Owing to already made remarks this holds at each point of $\partial D_{q}$ other than the negligible set of non-smooth boundary points of $D_{q}$.
+Proceeding now towards the retracts of $D_{q}$ (where we no longer assume $N=2$ for that would be an oversimplification unlike the matter in the preceding paragraph), it must first be noted that by proposition 1.4, every holomorphic retract of $\ell_{q}$-ball through the origin is actually linear even when $q<1$; for $1$-dimensional retracts, this follows from theorem 1.3 but as highlighted in the introduction, the thrust here is on retracts of arbitrary dimension.
+So, if $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$.
 Of-course, we are only interested in non-trivial retracts, so $Y_{q}$ is of the form $Y_{q}=D_{q}\cap Y$ for some $r$-dimensional $\mathbb{C}$-linear subspace $Y$ of $\mathbb{C}^{N}$.
-We only need to show that if $Y$ cannot be of the form $\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r},~{}r+1\leq i\leq N\}$ of $\mathbb{C}^{N}$, where not all $(a_{i,j})$ are zero.
-To prove this proposition by contradiction, we assume that $Y_{q}$ is a linear retract of $D_{q}$ and a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$.
+We only need to show that $Y$ cannot be of the form
+(5.2) $\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r}~{}\text{for}~{}r+1\leq i\leq N\},$
+where not all $(a_{i,j})$ are zero.
+To prove this theorem 1.9 by contradiction as already indicated above, assume that $Y_{q}$ is a linear retract of $D_{q}$ and a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$.
 Let $\{e_{i}:1\leq i\leq N\}$ be the standard basis of $\mathbb{C}^{N}$.
-Define another basis $\{v_{i}:1\leq i\leq N\}$ of $\mathbb{C}^{N}$, as follows:
+With notations as in (5.2), define another basis $\{v_{i}:1\leq i\leq N\}$ of $\mathbb{C}^{N}$, as follows:
 $v_{j}$ $=(\underbrace{0,\ldots,0,t_{j},0,\ldots,0}_{r},a_{j,r+1}t_{j},\ldots,a_{j,N}t_{j})\quad\text{for}~{}1\leq j\leq r,$
@@
 where
-$t_{j}=\frac{1}{\left(1+\sum_{k=r+1}^{N}|a_{j,k}|^{q}\right)^{1/q}}$
+$t_{j}=\frac{1}{\left(1+\sum_{k=r+1}^{N}|a_{j,k}|^{q}\right)^{1/q}}.$
 For $1\leq j\leq r$, write
-$e_{j}=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}v_{l}\right)$
-After applying $P$ on both sides, we get
-$P(e_{j})=\frac{1}{t_{j}}P(v_{j})-\left(\sum_{l=r+1}^{N}a_{j,l}P(v_{l})\right)$
+$e_{j}=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}v_{l}\right).$
+Applying $P$ on both sides, we get
+$P(e_{j})=\frac{1}{t_{j}}P(v_{j})-\left(\sum_{l=r+1}^{N}a_{j,l}P(v_{l})\right).$
 Since $P$ maps $\mathbb{C}^{N}$ into $Y$, $P(v_{j})\in Y$. Write
-(5.1) $P(v_{j})=\left(u_{1,j},\ldots,u_{r,j},\sum_{k=1}^{r}a_{k,r+1}u_{k,j},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,j}\right)~{}\text{for}~{}r+1\leq j\leq N$
+(5.3) $P(v_{j})=\left(u_{1,j},\ldots,u_{r,j},\sum_{k=1}^{r}a_{k,r+1}u_{k,j},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,j}\right)~{}\text{for}~{}r+1\leq j\leq N.$
 Since $P$ is a projection linear map from $\mathbb{C}^{N}$ onto $Y$ and $v_{j}\in Y$ for $1\leq j\leq r$, $P(v_{j})=v_{j}$ for $1\leq j\leq r$. Hence for $1\leq j\leq r$, we have
-$P(e_{j})=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}\left(u_{1,l},\ldots,u_{r,l},\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)\right)\\ =\left(-\sum_{l=r+1}^{N}a_{j,l}u_{1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{j-1,l},1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l},-\sum_{l=r+1}^{N}a_{j,l}u_{j+1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{r,l},\right.\\ \left.a_{j,r+1}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,a_{j,N}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)$
+$P(e_{j})=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}\left(u_{1,l},\ldots,u_{r,l},\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)\right)\\ =\left(-\sum_{l=r+1}^{N}a_{j,l}u_{1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{j-1,l},1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l},-\sum_{l=r+1}^{N}a_{j,l}u_{j+1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{r,l},\right.\\ \left.a_{j,r+1}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,a_{j,N}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,N}u_{k,l}\right).$
 Note that $P(v_{j})\in\overline{D_{q}}$ for all $j$ So, we get the following inequality for $1\leq j\leq r$,
-(5.2) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}\left|a_{j,s}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,s}u_{k,l}\right|^{q}\leq 1$
-The following inequality holds for $r+1\leq j\leq N$:
-(5.3) $\sum_{k=1}^{r}|u_{k,j}|^{q}+\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}~{}u_{k,j}\right|^{q}\leq 1$
-By reverse triangle inequality for square roots: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.2) reduces to
-$-\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0$
+(5.4) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}\left|a_{j,s}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,s}u_{k,l}\right|^{q}\leq 1.$
+Therefore, following inequality holds for $r+1\leq j\leq N$:
+(5.5) $\sum_{k=1}^{r}|u_{k,j}|^{q}+\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}~{}u_{k,j}\right|^{q}\leq 1.$
+By reverse triangle inequality for $q^{\text{th}}$ power where $10<q<$ 1: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.4) becomes
+$-\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
 Combining the above inequality with (5.5), we obtain the following inequality
-$-\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,j}\right|^{q}\right)+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0$
-Simplifing above inequality, we get
-$\sum_{l=r+1}^{N}|a_{j,l}|^{q}\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}\leq 0$
-From this inequality, we obtain
+$-\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,j}\right|^{q}\right)+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
+Simplifying this inequality, we get
+$\sum_{l=r+1}^{N}|a_{j,l}|^{q}\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}\leq 0.$
+As all quantities on the left are obviously non-negative, it follows that
 (5.6) $a_{j,l}u_{k,l}=0~{}~{}\quad r+1\leq l\leq N~{},1\leq j,k\leq r~{}\text{with}~{}j\neq k.$
-From inequality (5.2), we get
-(5.5) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|\leq t_{j},\quad 1\leq j\leq r$
-Since not all $(a_{i,j})$ are zero, atleast one of $a_{i,j}$ is non-zero.
+Combining this result with inequality (5.4), we get
+(5.7) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|\leq t_{j},\quad\text{ for each }j\in\{1,2,\ldots,r\}.$
+By our hypothesis about the $a_{i,j}$’s, atleast one of $a_{i,j}$ is non-zero.
 If $a_{j_{0},l_{0}}\neq 0$ for some $j_{0},l_{0}$ then from (5.6) we obtain that $u_{j,l_{0}}=0$ for all $j$ with $j\neq j_{0}$.
 From (5.5), we get $|u_{j_{0},l_{0}}|\leq t_{j_{0}}$.
-We can conclude that for all $(j,l)$, we have $|u_{j,l}|\leq t_{j}$.
+Therefore, we conclude that for all $(j,l)$, we have $|u_{j,l}|\leq t_{j}$.
 Using (5.7), we get
@@
 $\leq\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|+\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|$
-$\leq t_{j}+t_{j}(\sum_{l=r+1}^{N}|a_{j,l}|)<1$
+$\leq t_{j}+t_{j}(\sum_{l=r+1}^{N}|a_{j,l}|)<1.$
 This is obviously a contradiction, which completes the proof. ∎
-To make some noteworthy remarks here, let us also record a special case that characterises the non-trivial retracts through the origin of such a domain in $\mathbb{C}^{2}$ where this result is not new (though our proof is direct and more elementary).
+To make some noteworthy remarks here, let us also record a special case about all non-trivial retracts through the origin of such a domain in $\mathbb{C}^{2}$ where this result is not new (though our proof is direct and more elementary).
 Proposition 5.2.
@@
 Indeed, for the non-convex (but pseudoconvex) balanced domain $D_{q}$ in $\mathbb{C}^{2}$ as above, the fixed point set of the holomorphic self-map of $D_{q}$ as simple as the flip $(z,w)\to(w,z)$, namely the linear subspace $\{(z,w)\in\mathbb{C}^{2}\;:\;w=z\}$, is not realizable as a holomorphic retract of $D_{q}$, which is part of the content of the proposition 5.2 and its higher dimensional analogue in theorem 1.9 above.
+The $\ell_{q}$-balls dealt with above are bounded by holomorphically extreme real hypersurfaces.
+We next discuss examples for the ‘other extreme’ as well as some intermediate cases.
+We begin with an example where we can show existence of higher dimensional retracts; the purpose being: to record atleast one simple example to affirmatively answer the question as to whether there exists non-homogeneous pseudoconvex domains in $\mathbb{C}^{N}$ (i.e., when $N$ is not restricted to be $2$ or the domain being convex) with the property that there are retracts through every point of the domain and all possible dimensions; given Lempert’s apt remarks about the non-existence of retracts of dimension any bigger than one, in general.
+As our purpose is in recording the simplest cases first, the examples may well not be impressive.
+Example: Consider the unbounded balanced non-convex domain given by $H^{N}:=\{(z_{1},\cdots,z_{N})\in\mathbb{C}^{N}:|z_{1}\ldots z_{N}|<1\}$.
+Since none of the boundary points of $H^{N}$ are convex, there are no one-dimensional retracts passing through the origin except for the lines which are contained within one of the coordinate ‘planes’ i.e., lines spanned by vectors one of whose coordinates is zero.
+Such lines being completely contained within the domain, it follows that $H^{N}$ is a non-hyperbolic domain, in particular.
+Of-course $H^{N}$ is not biholomoprhic to $\mathbb{C}^{N}$ or any such domain which is ‘maximally non-hyperbolic’ wherein the Kobayashi metric vanishes identically; thereby $H^{N}$ seems to serve as a nice first example of an intermediate case with respect to hyperbolicity and in-fact, with respect to convexity as well.
+We now get to showing the existence of retracts of all possible dimensions through each point in this domain.
+To this end, pick any $c=(c_{1},c_{2},\ldots,c_{N})\in H^{N}$ and any positive integer $k$ less than $N$; we shall show that there exists a $k$-dimensional retract of $H^{N}$ passing through $c$.
+Consider the affine subspace $A=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{j}=c_{j}~{}\text{for}~{}k+1\leq j\leq N\}$, defined by setting an appropriate number of components equal to that of the fixed point $c$.
+It suffices to prove that the $k$-dimensional affine subspace $H^{N}\cap A$ is a retract of $H$.
+First we do this for $k=N-1$ and the rest follows recursively as noted below.
+To this end, we let $L_{c_{N}}:=\{(z_{1},\ldots,z_{N-1},z_{N})\in\mathbb{C}^{N}:z_{N}=c_{N}\}$ and show that the $(N-1)$-dimensional affine subspace $H^{N}\cap L_{c_{N}}$ of $H^{N}$ is a retract of $H^{N}$.
+We split into $2$ cases depending on $c_{N}$ being zero or not.
+If $c_{N}=0$, then the linear projection map $\pi(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{N-1},0)$ suffices to render ${\rm image}(\pi)=H^{N}\cap L_{c_{N}}$ as a retract of $H^{N}$.
+If $c_{N}\neq 0$, then we consider the holomorphic map $\rho$ on $H^{N}$ defined by
+$\rho(z_{1},\ldots,z_{N})=\left(\frac{1}{c_{N}}z_{1}z_{N},z_{2},\ldots,z_{N-1},c_{N}\right).$
+Let $\rho_{1},\ldots,\rho_{N}$ be the components of $\rho$. Note that
+$|\rho_{1}\cdots\rho_{N}|=\left|\frac{1}{c_{N}}z_{1}z_{N}z_{2}\cdots z_{N-1}c_{N}\right|=|z_{1}\cdots z_{N}|<1$
+So $\rho$ maps $H^{N}$ into $H^{N}$ and
+$\rho(\rho(z))=\rho(\frac{1}{c_{N}}\rho_{1}\rho_{N},\rho_{2},\ldots,\rho_{N-1},c_{N})=\left(\frac{1}{c_{N}}(\frac{1}{c_{N}}z_{1}z_{N}c_{N}),z_{2},\ldots,z_{N-1},c_{N}\right)=\rho(z)$
+verifying that $\rho$ is a retraction map onto $H^{N}\cap L_{c_{N}}$ which is biholomorphically – indeed linearly – equivalent to $H^{N-1}_{1/c_{N}}$ which in-turn is linearly equivalent to $H^{N-1}$ (in-case $c_{N}=0$, note that $H^{N}\cap L_{c_{N}}$ is just a copy of $\mathbb{C}^{N-1}$).
+The same arguments as above, show that $H^{N-1}\cap L_{c_{N-1}}$ is either a holomorphic retract of $H^{N-1}$ or a holomorphic retract of $\mathbb{C}^{N-1}$ depending on whether $c_{N-1}$ is non-zero or not.
+By lemma 2.9, we obtain that $\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{N-1}=c_{N-1},z_{N}=c_{N}\}\cap H^{N}$ is a retract of $H^{N}$.
+Continuing this process, we establish in $k$ steps, that $H^{N}\cap A$ is a $k$-dimensional retract of $H^{N}$.∎
+At this juncture, we can settle a question about whether holomorphic retraction mappings are automatically submersions so as (for instance) to render the hypotheses of theorem 1.17 redundant; to answer this, it suffices to consider the two-dimensional version of the above domain.
+Theorem 5.4.
+Consider the retract $Z=\{(\zeta,1):\zeta\in\Delta\}$ of $H^{2}$ in $\mathbb{C}^{2}$. Then there does not exist any retraction map from $H^{2}$ onto $Z$, which is submersive at origin.
+Proof.
+Let $\rho=(\rho_{1},\rho_{2})$ be any of the retractions which map $H^{2}$ onto $Z$.
+Since second component of any point in $Z$ is $1$, we have $\rho_{2}\equiv 1$.
+No such drastic conclusion cannot be made for the first component of $\rho$ namely $\rho_{1}$, as we only have to begin with that it is a holomorphic function which maps $H^{2}$ onto $\Delta$; nevertheless we may deduce about its form as a consequence of certain features of $H^{2}$.
+To do so, consider first its restriction: $f(w):=\rho_{1}(0,w)$.
+Being an entire function which maps into $\Delta$, Liouville’s theorem implies that $f$ is a constant function.
+As $f(1)=0$, that constant is zero, thereby $f(w)=\rho_{1}(0,w)=0$ for all $w\in\mathbb{C}$.
+As $H^{2}$ is a logarithmically convex complete Reinhardt domain, $\rho_{1}$ can be expressed as a single power series expansion on $H^{2}$.
+This means that there exist scalars $c_{j,k}$ such that for all $(z,w)\in H^{2}$,
+$\rho_{1}(z,w)=\sum_{j,k=0}^{\infty}c_{j,k}z^{j}w^{k}$
+As the series above converges absolutely on $H^{2}$, we may rewrite by grouping together the terms as indicated next:
+(5.8) $\rho_{1}(z,w)=P_{1}(z)+P_{2}(w)+P_{3}(z,w)$
+where
+$P_{1}(z)=\sum_{j=1}^{\infty}c_{j,0}z^{j},~{}P_{2}(w)=\sum_{k=0}^{\infty}c_{0,k}w^{k},~{}P_{3}(z,w)=\sum_{j,k=1}^{\infty}c_{j,k}z^{j}w^{k}.$
+Since $\rho_{1}(0,w)=0$ for all $w\in\mathbb{C}$, we obtain that $P_{2}\equiv 0$.
+Similarly we can consider another entire function $g:\mathbb{C}\to\Delta$ defined by $g(z)=\rho_{1}(z,0)$.
+By the same arguments, we obtain that $g$ for all $z\in\mathbb{C}$ and hence $P_{1}\equiv 0$.
+Therefore from the above equation (5.8), we arrive at our aforementioned deduction about the form of the first component namely, $\rho_{1}(z,w)=zwh(z,w)$ for some holomorphic function $h$ on $H^{2}$.
+We may now prove that $\rho$ is not submersive at origin by directly writing down the expression for its Jacobian matrix, which at any point $(z,w)$ is if the form
+$\begin{bmatrix}w(z\frac{\partial h_{2}}{\partial z}+h_{2}(z,w))&z(w\frac{\partial h_{2}}{\partial w}+h_{2}(z,w))\\ 0&0\end{bmatrix}$
+It is immediate that for $(z,w)=(0,0)$ this is the zero matrix, finishing the proof that none of the retraction maps onto $Z$ is not submersive at origin. ∎
+Remark 5.5.
+While the above theorem with a very similar proof holds for any of the hyperbolic retracts $Z^{h}$ of $H^{2}$ (to give that none of the retraction mappings onto $Z^{h}$ is a submersion at the origin), it is worth noting down a concrete example $Z$ as in the theorem for an amusing point.
+Namely, while linear subspaces (i.e., through origin) of a balanced domain which are obtained as images of non-linear retraction mappings are indeed obtainable as images of linear maps, this is not the case with affine-linear retracts in general.
+Indeed, this can be seen as a corollary of the above theorem.
+While the retract $Z=\{(\zeta,1):\zeta\in\Delta\}$ of $H^{2}$ is an affine-linear subspace of $H^{2}$, there exists no affine linear projection mapping the domain onto $Z$ since such a map would be a retraction which is submersive everywhere (ruled out by the above theorem).
+We also note that $Z$ gives a concrete example of a hyperbolic retract (being a copy of the disc) of the non-hyperbolic domain $H$ which therefore serves as a good intermediate between completely non-hyperbolic domains such as $\mathbb{C}^{N}$ which do not admit any hyperbolic retracts on the one hand, and, hyperbolic domains which admit only hyperbolic ones on the other.
+A more important point is noted next.
+A direct consequence of the above theorem is that unless we are in very special settings even within the class of balanced pseudoconvex domains, the hypothesis about submersivity of retractions as in theorem 1.17 cannot be dropped.
+For a quick (counter-)example, observe that the retract $Z$ in the foregoing theorem 5.4 cannot be realized as the graph over any linear retract of $H^{2}$ as in the conclusion of theorem 1.17.
+Indeed, the only (non-trivial) linear retracts of $H^{2}$ are the pair of coordinate axes, on neither of which $Z$ can be expressed as a graph, for then $Z$ would be biholomorphic to $\mathbb{C}$, manifestly false.
+It is also possible to give a non-affine example in this simple domain itself to show that there are retracts which are impossible to be realized as a graph over any affine linear subspace of $H$, thereby showing (again) that the conclusion of theorem 1.17 can fail as soon as its hypotheses, particularly about submersivity of the retractions fail which is the purpose of the following corollary.
+Corollary 5.6.
+Consider the retract $\tilde{Z}=\{(\zeta e^{-\zeta},e^{\zeta}):\zeta\in\Delta\}$ of $H^{2}$ in $\mathbb{C}^{2}$.
+Then there does not exist a retraction map from $H^{2}$ onto $\tilde{Z}$, which is submersive at the origin.
+Proof.
+Suppose to arrive at a contradiction that $\tilde{\rho}:H^{2}\to H^{2}$ is a retraction map with ${\rm image}(\tilde{\rho})=\tilde{Z}$ and such that it is a submersion at the origin.
+Consider the automorphism $\psi:H^{2}\to H^{2}$ defined by $\psi(z,w)=(ze^{-zw},we^{zw})$.
+Conjugating $\rho$ by $\psi$, we then have the holomorphic map $\rho:H^{2}\to H^{2}$ given by $\rho(z,w)=\psi^{-1}\circ\tilde{\rho}\circ\psi(z,w)$.
+Note that $\rho$ is a retraction map on $H^{2}$ and ${\rm image}(\rho)=Z=\{(\zeta,1):\zeta\in\Delta\}$ and we may apply theorem 5.4 to $\rho$.
+But then the submersitivity of $\tilde{\rho}$ at the origin is equivalent to that of $\rho$ at the origin, as $\psi$ is not just a conjugating automorphism but also fixes at the origin; this contradicts theorem 5.4 finishing the proof.
+∎
+We next record a pair of bounded modifications of $H^{2}$ in the above examples wherein the some of the results hold as is, but some arguments cease to.
+Example: Consider the domain $\Omega:=\Delta^{2}\cap H^{2}_{1/4}$, where $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}:|zw|<1/4\}$ (and $\Delta^{2}$ the standard bidisc, as usual).
+We demonstrate below that for each point $(a,b)\in\Omega$, there exists a non-trivial retract of $\Omega$ passing through it.
+Let $L$ denote the affine complex line $\{(z,w)\in\mathbb{C}^{2}:w=b\}$.
+Observe that to attain our goal, it suffices to show that $L\cap\Omega$ is a retract of $\Omega$ for which purpose, we first decompose the boundary of our domain $\Omega$ as:
+$\partial\Omega=(\partial\Delta^{2}\cap H^{2}_{1/4})\cup(\Delta^{2}\cap\partial H^{2}_{1/4})\cup(\partial\Delta^{2}\cap\partial H^{2}_{1/4}).$
+Note that $L\cap\Omega$ can be viewed as an open subset of $L$ and viewed thus, its boundary satisfies $\partial(L\cap\Omega)=L\cap\partial\Omega$.
+We split-up the rest of the proof into claims and cases.
+Claim-1: Either $\partial(L\cap\Omega)\subset\partial\Delta^{2}$ or $\partial(L\cap\Omega)\subset\partial H^{2}_{1/4}$.
+Proof: Assume to get a contradiction that $\partial(L\cap\Omega)\nsubseteq\partial\Delta^{2}$ and $\partial(L\cap\Omega)\nsubseteq\partial H^{2}_{1/4}$.
+Then there are points $(a_{1},b)$ and $(b_{1},b)$ in $\partial(L\cap\Omega)$ such that $(a_{1},b)\notin\partial\Delta^{2}$ and $(b_{1},b)\notin\partial H^{2}_{1/4}$.
+From the above decomposition of $\partial\Omega$ and $(a_{1},b),(b_{1},b)\in\partial\Omega$, we obtain that $(a_{1},b)\in\Delta^{2}\cap\partial H^{2}_{1/4}$ and $(b_{1},b)\in\partial\Delta^{2}\cap H^{2}_{1/4}$.
+This implies that $a_{1},b_{1},b$ satisfies the following inequalities:
+$|a_{1}b|=1/4,~{}\max\{|a_{1}|,|b|\}<1$
+$|b_{1}b|<1/4,~{}\max\{|b_{1}|,|b|\}=1.$
+Note that $|b|,|a_{1}|<1$ because $\max\{|a_{1}|,|b|\}<1$.
+Using the inequality $|b|<1$ and the equality $\max\{|b_{1}|,|b|\}=1$, we obtain $|b_{1}|=1$.
+Since $|b_{1}b|<1/4$, we have $|b|<1/4$.
+Hence we get that
+$\frac{1}{4}=|a_{1}b|<\frac{1}{4}|a_{1}|$
+This implies that $|a_{1}|>1$, which contradicts the fact that $|a_{1}|<1$, finishing the proof of claim-$1$.
+We are now in a position to bifurcate the rest of the arguments into two cases depending on whether $\partial(L\cap\Omega)\subset\partial\Delta^{2}$ or $\partial(L\cap\Omega)\subset\partial H^{2}_{1/4}$.
+The arguments in either of these cases proceeds by fixing a point $(a^{\prime},b)$ in $\partial(L\cap\Omega)$; it is also introduce the notations: $Z:=L\cap\Delta^{2}$, $W:=L\cap H^{2}_{1/4}$.
+Note that $Z$ is bounded whereas $W$ could be unbounded, depending upon where $L$ intersects $\partial\Delta^{2}$.
+We first deal with the easier case.
+Case-1: $\partial(L\cap\Omega)\subset\partial\Delta^{2}$.
+In this case, by the aforementioned decomposition of $\partial\Omega$, we first note that the boundary point $(a^{\prime},b)$ lies in $\partial\Delta^{2}\cap\overline{H^{2}_{1/4}}$.
+The map $\rho$ on $\Delta^{2}$ defined by $\rho(z,w)=(z,b)$ suffices to give a holomorphic retraction map on $\Delta^{2}$ with ${\rm image}(\rho)=Z$.
+However, we must check that $Z$ is a retract of $\Omega$, not just of $\Delta^{2}$.
+To do this of-course we need only consider the restriction $\rho|_{\Omega}$, which maps $\Omega$ into $Z$ and that would finish the proof provided we verify the following.
+Claim-2: $Z=L\cap\Omega$.
+Proof: Firstly note that if $(z,b)\in Z$ then $|z|<|a^{\prime}|=1$ because $(a^{\prime},b)\in\partial\Delta^{2}$ and $(a,b)\in\Delta^{2}$.
+If $(z,b)\in Z$, then $|zb|<|a^{\prime}b|$.
+Since $(a^{\prime},b)\in\overline{H^{2}_{1/4}}$, we get that $|zb|<1/4$.
+Hence $Z\subset L\cap\Omega$.
+Since $\Omega\subset\Delta^{2}$, we have $L\cap\Omega\subset Z$.
+This finishes the proof of the claim.
+It now follows immediately that $\rho$ maps $\Omega$ into itself and ${\rm image}(\rho|_{\Omega})\subset L\cap\Omega=Z$; indeed, as we have for every $z\in Z=L\cap\Omega$ that $\rho(z)=z$, we have ${\rm image}(\rho|_{\Omega})=L\cap\Omega$.
+This finishes the proof completely that $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$ in the above case.
+Case-2: $\partial(L\cap\Omega)\subset\partial H^{2}_{1/4}$.
+In this case, again by the aforementioned decomposition of $\partial\Omega$, we start by noting that the boundary point $(a^{\prime},b)$ lies in $\overline{\Delta^{2}}\cap\partial H^{2}_{1/4}$.
+Note that $b\neq 0$ as $(a^{\prime},b)\in\partial H^{2}_{1/4}$.
+Now in contrast to the previous case, no linear choices work (the proof of which is easy but we shall not digress into that here), thereby leading us to consider the non-linear map $\varphi:H^{2}_{1/4}\to H^{2}_{1/4}$ defined by $\varphi(z,w)=(\frac{1}{b}zw,b)$.
+which is of-course a holomorphic retraction map on $H^{2}_{1/4}$ with ${\rm image}(\varphi)=W$.
+It remains to realize $W$ as a retraction of our smaller bounded $\Omega$, rather than the unbounded $H^{2}_{1/4}$.
+Again we need only consider the restriction $\varphi|_{\Omega}$ (which maps $\Omega$ into $W$), provided we verify the following.
+Claim-3: $W=L\cap\Omega$.
+Proof: If $(z,b)\in W$ then $|z|<|a^{\prime}|$, because if $|z|\geq|a^{\prime}|$ then $|zb|\geq|a^{\prime}b|=1/4$, which contradicts the fact that $(z,b)\in W$.
+Since $(a^{\prime},b)\in\overline{\Delta^{2}}$, we have $|z|<1$.
+Hence $(z,b)\in L\cap\Delta^{2}$ and therefore $W\subset L\cap\Omega$.
+As $\Omega\subset H^{2}_{1/4}$, we have $L\cap\Omega\subset W$, finishing the proof of the claim.
+So, $\varphi$ maps $\Omega$ into $\Omega$ and the proof that ${\rm image}(\varphi)=L\cap\Omega$ is a non-trivial retract of $\Omega$ passing through $(a,b)$ is now completed in all cases.
+Remark 5.7.
+Consider the retract $Z=\{(\zeta,b):|\zeta|<\frac{1}{4|b|}\}$ of $\Omega$ as above.
+We observe that $Z$ can be expressed as a graph over a linear retract (through the origin) unlike the situation in the unbounded version, as follows.
+Fix any $z_{0}$ with $|z_{0}|\leq 1/4$ and let $L=\{\lambda(z_{0},1):\lambda\in\mathbb{C}\}$, the linear subspace of $\mathbb{C}^{2}$ spanned by the boundary point $(z_{0},1)$.
+If we let $\phi$ be the the holomorphic map given by $\phi(\lambda)=(\frac{1}{4b}\lambda^{2},b)$, then it is straightforward to check that $\phi$ maps $\Delta$ into $H^{2}$, indeed onto $Z$, owing to the chosen constraint on $z_{0}$.
+If we view $\mathbb{C}^{2}$ as the direct sum of $L$ with the span of $(0,b)$ (indeed, the $z$-axis!), and define $\varphi$ on $L$ by $\varphi\left(\lambda(z_{0},1)\right)=\phi(\lambda)$, we observe then that: $Z$ is in-fact the graph of $\varphi$ over the linear subspace $\Omega_{L}=\Omega\cap L$.
+An interesting trivia to note here that when we choose $z_{0}=0$, we obtain the ‘horizontal’ retract $Z$ as the graph over the ‘vertical’ subspace obtained by intersecting the $w$-axis with the domain i.e., the same $L$ when $z_{0}=0$.
+At the other extreme i.e., when $|z_{0}|=1/4$, we observe that $\varphi$ is just the restriction of the retraction map $f(z,w)=(zw/b,b)$.
+In any case, we can neither choose $f$ nor $\varphi$ to be affine-linear owing to theorem 5.4.
+As another difference from the unbounded $H^{2}$, thereby a purpose of the above $\Omega$, is that the sets: $S_{1}$ of unit vectors representing directions along which there are linear retracts in $\Omega$, as well as, $S_{2}$ of those unit vectors along which there are no non-trivial retracts through the origin, are both non-trivial subsets of the unit sphere neither of which have measure zero (and with both having non-empty interior in the sphere).
+We may ask for variants of the above $H^{2}_{1/4}\cap\Delta^{2}$ and more; we shall only record the case when the polydisc $\Delta^{2}$ is replaced by the Euclidean ball $\mathbb{B}$.
+Example: Consider the domain $\Omega:=\mathbb{B}\cap H^{2}_{1/4}$, where $\mathbb{B}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{2}+|w|^{2}<1\}$ be the euclidean ball centred at 00 of radius $1$ and $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}:|zw|<1/4\}$
+We want to prove that for each $(a,b)\in\Omega$, there exist a retract (of dimension 1) of $\Omega$ passing through $(a,b)$.
+Let $L$ denotes the hyperplane $\{(z,w)\in\mathbb{C}^{2}:w=b\}$ Firstly note that
+$\partial\Omega=(\partial\mathbb{B}\cap H^{2}_{1/4})\cup(\mathbb{B}\cap\partial H^{2}_{1/4})\cup(\partial\mathbb{B}\cap\partial H^{2}_{1/4})$
+Note that $\partial(L\cap\Omega)=L\cap\partial\Omega$ First we prove the following claim:
+Claim-1: Either $\partial(L\cap\Omega)\subset\partial\mathbb{B}$ or $\partial(L\cap\Omega)\subset\partial H^{2}_{1/4}$ for $i=1,2$.
+Proof of the claim-1: If possible assume that $\partial(L\cap\Omega)\nsubseteq\partial\mathbb{B}$ and $\partial(L\cap\Omega)\nsubseteq\partial H^{2}_{1/4}$ Then there exist points $(a_{1},b)$ and $(b_{1},b)$ in $\partial(L\cap\Omega)$ such that $(a_{1},b)\notin\partial\mathbb{B}$ and $(b_{1},b)\notin\partial H^{2}_{1/4}$.
+From the above decomposition of $\partial\Omega$ and $(a_{1},b),(b_{1},b)\in\partial\Omega$, we obtain that $(a_{1},b)\in\mathbb{B}\cap\partial H^{2}_{1/4}$ and $(b_{1},b)\in\partial\mathbb{B}\cap H^{2}_{1/4}$.
+This implies that $a_{1},b_{1},b$ satisfies the following equations:
+$|a_{1}|^{2}+|b|^{2}<1,~{}|a_{1}b|=1/4$
+$|b_{1}|^{2}+|b|^{2}=1,~{}|b_{1}b|<1/4$
+Dividing $|a_{1}b|$ (which is equal to $1/4$) on both sides of the equation $|b_{1}b|<1/4$, we get $|b_{1}|<|a_{1}|$.
+Squaring and adding $|b|^{2}$ on both sides of the equation $|b_{1}|<|a_{1}|$, we get that $11=|b_{1}|^{2}+|b|^{2}<|a_{1}|^{2}+|b|^{2}<$ 1, which is vividly a contradiction.
+This finishes the proof of the claim.
+So we can divide into two cases depending on whether $\partial(L\cap\Omega)\subset\partial\mathbb{B}$ or $\partial(L\cap\Omega)\subset\partial H^{2}_{1/4}$.
+Fix a point $(a^{\prime},b)$ in $\partial(L\cap\Omega)$ and let $Z:=L\cap\mathbb{B}$, $W:=L\cap H^{2}_{1/4}$.
+Case-1: $\partial(L\cap\Omega)\subset\partial\mathbb{B}$.
+Since $(a^{\prime},b)\in\partial\Omega$ and the above decomposition of $\partial\Omega$, we obtain that $(a^{\prime},b)\in\partial\mathbb{B}\cap\overline{H^{2}_{1/4}}$.
+We know that $Z$ is an retract of $\mathbb{B}$.
+Then there exist a holomorphic retraction map $\rho$ on $\mathbb{B}$ such that $\rho(\mathbb{B})=Z$.
+Now we prove that $Z$ is a retract of $\Omega$.
+To prove this, consider the map $\rho|_{\Omega}$, which maps $\Omega$ into $Z$.
+Next we prove the following claim:
+Claim -2: $Z=L\cap\Omega$.
+Proof of claim-2: Note that $|a^{\prime}|\leq\frac{1}{4|b|}$, because if $|a^{\prime}|>\frac{1}{4|b|}$ then $|a^{\prime}b|>1/4$ and thereby implies that $(a^{\prime},b)\notin\overline{H^{2}_{1/4}}$, which contradicts our assumption that $(a^{\prime},b)\in\partial(L\cap\Omega)\subset\partial\mathbb{B}$.
+If $(z,b)\in Z$ then $|z|<|a^{\prime}|$ because if $|z|\geq|a^{\prime}|$, then $|z^{2}|+|b|^{2}\geq|a^{\prime}|^{2}+|b|^{2}=1$, which contradict the fact that $(z,b)\in\mathbb{B}$.
+Note that $|zb|<|a^{\prime}b|\leq 1/4$.
+This implies that $(z,b)\in L\cap\Omega$ and hence $Z\subset L\cap\Omega$.
+Since $\Omega\subset\mathbb{B}$, we have $L\cap\Omega\subset Z$, .
+This finishes the proof of the claim.
+Using the above claim-2, $\rho$ maps $\Omega$ into itself and $image(\rho|_{\Omega})\subset L\cap\Omega=Z$.
+If $z\in Z=L\cap\Omega$, then $\rho(z)=z$.
+So $image(\rho|_{\Omega})=L\cap\Omega$.
+Hence $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$.
+Case-2: $\partial(L\cap\Omega)\subset\partial H^{2}_{1/4}$.
+Firstly note that $b\neq 0$ because $(a^{\prime},b)\in\partial H^{2}_{1/4}$. Consider the holomorphic map $\varphi:H^{2}_{1/4}\to H^{2}_{1/4}$ defined by $\varphi(z,w)=(\frac{1}{b}zw,b)$.
+Note that $\varphi$ is a retraction map on $H^{2}_{1/4}$. Now we consider the map $\varphi|_{\Omega}$, which maps $\Omega$ into $W$. Next we prove the following claim:
+Claim-3: $W=L\cap\Omega$.
+Proof of claim-3: First we prove that $|a^{\prime}|\geq\frac{1}{4|b|}$.
+If $|a^{\prime}|<\frac{1}{4|b|}$ then $|a^{\prime}b|<1/4$ and thus implies that $(a^{\prime},b)\in H^{2}_{1/4}$, which contradicts the fact that $(a^{\prime},b)\in\partial H^{2}_{1/4}$ Hence we proved that $|a^{\prime}|\geq\frac{1}{4|b|}$
+If $(z,b)\in W$, then $|z|<\frac{1}{4|b|}$. Note that
+$|z|^{2}+|b|^{2}<\left(\frac{1}{4|b|}\right)^{2}+|b|^{2}\leq|a^{\prime}|^{2}+|b|^{2}.$
+Since $(a^{\prime},b)\in(L\cap\partial\Omega)\subset\partial H^{2}_{1/4}$ and the above decomposition of $\partial\Omega$, we obtain that $(a^{\prime},b)\in\overline{\mathbb{B}}$.
+Hence $|z|^{2}+|b|^{2}<1$ and therefore we get that $W\subset L\cap\Omega$.
+Since $\Omega\subset D$, we have $L\cap\Omega\subset W$.
+This finishes the proof of the claim.
+Hence $\varphi$ maps $\Omega$ into $\Omega$. Therefore $image(\varphi)=L\cap\Omega$ is a one-dimensional retract of $\Omega$ passing through $(a,b)$.
+We conclude this section by noting that the $\ell_{q}$ ball in $\mathbb{C}^{N}$ is not a Lempert domain, the class of domains considered in the very recent article [GZ], since a balanced pseudoconvex domain is a Lempert domain if and only if it is convex, as verified below.
+Indeed, recall that a taut domain $D\subset\mathbb{C}^{N}$ is a Lempert domain if the identity $c_{D}=\ell_{D}$ holds.
+Proposition 5.8.
+Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$. Then $D$ is a Lempert domain if and only if $D$ is convex.
+Proof.
+Assuming first that $D$ is a Lempert domain, we have $c_{D}=\ell_{D}$. By proposition 3.1.11 in [Jrncki_invrnt_dst], we have $p(0,h(z))=\ell_{D}(0,z)=\tanh^{-1}(h(z))$. Thereby,
+$c_{D}(0,z)=\tanh^{-1}(c_{D}^{*}(0,z))=\tanh^{-1}(h(z))=\ell_{D}(0,z).$
+This implies that $c_{D}^{*}(0,z)=h(z)$ for all $z\in D$. By proposition 2.3.1 in [Jrncki_invrnt_dst], $D$ is convex.
+The converse is the famous theorem of Lempert (exposited as theorem 11.2.1 in [Jrncki_invrnt_dst]). ∎
+6. Proof of theorem 1.10
+We begin by connecting the notions of convexity as used in the statement of the theorem, usual geometric convexity and continuity of the Minkowski functional – throughout this section $D$ will stand for a bounded balanced domain of holomorphy in $\mathbb{C}^{N}$ without further mention; we may only mention additional assumptions if made.
+This will be done in the statement and in course of the proof of the following lemma.
+Though this is not really a new result, we include the proof not only to keep it self-contained but also because it is crucial for what follows.
+Lemma 6.1.
+If $\partial D$ is convex near $p$ then the Minkowski functional $h$ of $D$ is continuous near $p$.
+Proof.
+Let us firstly quickly note that if $D$ is convex near $p$ is the sense of definition 2.30, then $D$ is geometrically convex near $p$ i.e., there exist an open ball $U$ in $\mathbb{C}^{N}$ centered at $p$ such that $U\cap D$ is convex in the usual sense of being closed under the formation of line segments.
+Indeed, this follows by considering the closed half-spaces (containing the domain) given by the linear functionals as per definition 2.30 due to the points of $\partial D\cap U$ being convex boundary points; supporting hyperplanes (and likewise half-spaces containing $D\cap U$) at other points of $\partial(D\cap U)$ exist just by convexity of the ball $U$.
+As the intersection of these closed half-spaces precisely equals $\overline{U\cap D}$, the convexity of $U\cap D$ follows.
+Next, we show the subadditivity of $h$.
+Let $z,w\in U\cap D$.
+By definition of the Minkowski functional, there exist positive reals $t,s$ such that $z/t,w/s\in U\cap D$ and $t<h(z)+\epsilon$, $s<h(w)+\epsilon$ for some $\epsilon>0$.
+Writing $(z+w)/(s+t)$ as the convex combination:
+$\frac{z+w}{s+t}=\frac{t}{s+t}\cdot\frac{z}{t}+\frac{s}{s+t}\cdot\frac{w}{s},$
+it follows that by the convexity of $U\cap D$, that $(z+w)/(s+t)\in U\cap D$.
+Hence $h(z+w)\leq s+t<h(z)+h(w)+2\epsilon$.
+As $\epsilon$ is arbitrary, we get $h(z+w)\leq h(z)+h(w)$ for all $z,w\in U\cap D$, finishing the verification of the required subadditivity.
+We now proceed to the original goal about continuity of $h$.
+Let $q$ be arbitrary point in $U\cap\partial D$ and $(z_{n})$ be a sequence converging to $q\in U\cap\partial D$.
+(without loss of generality, we assume that $z_{n}\in U\cap D$ for all $n$).
+By subadditivity of $h$ and the well-known basic fact that the Minkowski functional is dominated by some scaling of the Euclidean norm, we have $|h(z_{n})-h(q)|\leq h(z_{n}-q)\leq c\|z_{n}-q\|$ for some $c>0$.
+The convergence of $z_{n}$ to $q$, thereby readily implies that $h(z_{n})$ converges to $h(q)$.
+Hence $h$ is continuous near $p$, finishing the proof owing to what was noted earlier.
+∎
+The converse of the above lemma is false.
+Indeed, there are plenty of (counter-) examples, the simplest of which was given in the remark 1.11 of the introduction; another explicit example is obtained be examining the non-smooth boundary points of the domain called $D_{h}$ in section 8 below; it is not difficult to construct families of further examples in plenty.
+Proof of theorem 1.10: We shall work with the more general and weaker assumption that $\partial D$ is convex at the point $p$ together-with the continuity of the Minkowski functional $h$ of $D$; the other assumption that $p$ is not a $\mathbb{C}$-extremal boundary point as in the statement of theorem 1.10 remains in force.
+Just by the hypotheses that $D$ is balanced, we may describe $D$ as $D=\{z\in\mathbb{C}^{N}:h(z)<1\}$.
+As $p\in\partial D$ is not $\mathbb{C}$-extremal, there exist $\epsilon_{0}>0$ and a non-zero $u\in\mathbb{C}^{N}$ such that $p+tu\in\partial D$ for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$.
+After a rescaling of $u$, we may assume $\epsilon_{0}=1$ and also (by the hypotheses) that $h$ is continuous at $p+tu$ for all $t\in\Delta$.
+Consider the holomorphic map $\varphi:\Delta\to\mathbb{C}^{N}$ defined by $\varphi(t)=(p+tu)t$.
+As $p+tu\in\partial D$ for all $t\in\Delta$ and $h$ is continuous at $p+tu$ for all $t\in\Delta$, we have by a well-known fundamental fact (see for instance Remark 2.2.1 (h) of [Jrncki_invrnt_dst]) that $h(p+tu)=1$ for all $t\in\Delta$.
+Hence
+$h(\varphi(t))=h((p+tu)t)=|t|h(p+tu)=|t|<1.$
+So, $\varphi$ maps $\Delta$ into $D$.
+Next we want to prove that $\varphi$ is a complex geodesic in $D$.
+First we claim that there exist $t_{0}\in\Delta^{*}:=\Delta\setminus\{0\}$ such that $D$ is convex at $p+t_{0}u$.
+By the definition of convexity of $\partial D$ at $p$, there exist a $\mathbb{C}$-linear functional $L$ such that
+$|L(p)|=h(p)=1~{}~{}~{}\text{and}~{}~{}|L(z)|\leq h(z)\;\;\text{ for all }z\in\mathbb{C}^{N}$
+To prove the aforementioned claim, it suffices to show that there exist $t_{0}\in\Delta^{*}$ such that
+(6.1) $|L(z)|\leq h(z)~{}~{}\text{and}~{}~{}|L(p+t_{0}u)|=h(p+t_{0}u)=1.$
+Since $\mathbb{C}^{N}=Ker(L)\oplus span\{p\}$, we can write $u=u_{1}+u_{2}$, where $u_{1}\in Ker(L)$ and $u_{2}\in span\{p\}$. Write $u_{2}=\lambda p$ for some $\lambda\in\mathbb{C}$. So
+$|L(p+tu)|=|L(p)+tL(u)|=|L(p)+t\left(L(u_{1})+L(u_{2})\right)|=|1+\lambda t||L(p)|=|1+\lambda t|.$
+We need only prove that there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$, which we do now.
+If $\lambda=0$, then $u\in Ker(L)$.
+Hence (6.1) holds for any $t_{0}\in\Delta^{*}$.
+Assume that $\lambda\neq 0$.
+Write $\lambda t=re^{i\theta}$ where $\theta\in[-\pi,\pi)$.
+From this we get $t=\frac{1}{\lambda}(re^{i\theta})$.
+The relations $|1+\lambda t|=1$ and $|t|<1$ becomes $|1+re^{i\theta}|=1$ and $|re^{i\theta}|<|\lambda|$.
+We want to prove that there exist $(r,\theta)$ such that
+$|re^{i\theta}+1|=1~{}~{}\text{ and }~{}~{}~{}r<|\lambda|.$
+Geometrically above equation represents the intersection of circle centred at $-1$ with radius $1$ and the disc centred at 00 of radius $|\lambda|$.
+So there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$.
+This proves the claim.
+Remark 6.2.
+The arguments above can actually be used to show that $\partial D$ is convex at all points near $p$ in the complex line segment through $p$ in the direction of $u$.
+To proceed further with the proof that $\varphi$ is a complex $C$-geodesic, let us now note that as $D$ is convex at the boundary point $p+t_{0}u$ by the foregoing, we have by proposition 2.3.1(b) in [Jrncki_invrnt_dst], with the same notations therein, that
+$c_{D}^{*}(0,(p+t_{0}u)t_{0})=h((p+t_{0}u)t_{0})=|t_{0}|h(p+t_{0}u)=|t_{0}|;$
+for the definition of $c_{D}^{*}$ refer to section 2.4. Recall also the definition of the Caratheodory distance $c_{D}$ and its relationship to $c_{D}^{*}$: $c_{D}=\tanh^{-1}c_{D}^{*}$. So
+$c_{D}(0,(p+t_{0}u)t_{0})=\tanh^{-1}\left(c_{D}^{*}(0,(p+t_{0}u)t_{0})\right)=\tanh^{-1}|t_{0}|$
+For the term at the left note that: $c_{D}(\varphi(0),\varphi(t_{0}))=c_{D}(0,(p+t_{0}u)t_{0})$; the term on the right can just be realized as the Poincare distance between the $t_{0}$ and the origin in the disc $\Delta$.
+Indeed if $m(a,b)$ denotes the Mobius distance between two points $a,b\in\Delta$, recall that $p(a,b)=\tanh^{-1}(m(a,b))$.
+Therefore, $c_{D}(\varphi(0),\varphi(t_{0}))=p(0,t_{0})$.
+By proposition 11.1.4 in [Jrncki_invrnt_dst], we have $c_{D}(\varphi(\lambda),\varphi(\mu))=p(\lambda,\mu)$ for all $\lambda,\mu\in\Delta$.
+Hence $\varphi$ is a complex $C$-geodesic on $D$.
+It follows that $Z={\rm image}(\varphi)$ is a one-dimensional retract of $D$.
+To finish, let us check that $Z$ is indeed non-linear.
+Assume the contrary that $Z$ is linear.
+Let $b$ be a non-zero element in $Z$; it has got to be of the form: $b=(p+t_{0}u)t_{0}\in Z$ for some $t_{0}\in\Delta^{*}$.
+As $Z$ is linear, $\frac{1}{2}(p+t_{0}u)t_{0}\in Z$.
+This implies that
+(6.2) $\frac{1}{2}(p+t_{0}u)t_{0}=(p+s_{0}u)s_{0},$
+for some $s_{0}\in\Delta^{*}$.
+So, $h(\frac{1}{2}(p+t_{0}u)t_{0})=h(s_{0}(p+s_{0}u))$ and this implies that $|t_{0}|h(p+t_{0}u)=2|s_{0}|h(p+s_{0}u)$.
+Since $p+tu\in\partial D$ for all $t\in\Delta$ and $D$ is convex, we have $|t_{0}|=2|s_{0}|$.
+We can write $s_{0}=(t_{0}/2)e^{i\theta}$ for some $\theta\in[-\pi,\pi)$.
+From (6.2), we get
+$(p+t_{0}u)t_{0}=t_{0}e^{i\theta}(p+\frac{t_{0}}{2}ue^{i\theta}).$
+Since $t_{0}\neq 0$, the above equation becomes
+(6.3) $p+t_{0}u=e^{i\theta}p+\frac{e^{i2\theta}t_{0}}{2}u.$
+We claim $p,u$ are linearly independent. To prove this, assume the contrary: $p=\lambda u$ for some $\lambda\in\mathbb{C}$. Applying $h$ on both sides, we get
+$h(p)=|\lambda|h(u).$
+By the hypotheses that $D$ is bounded and from remark 2.2.1 (l) in [Jrncki_invrnt_dst], we obtain that $h(u)\neq 0$.
+Hence we get $|\lambda|=1/h(u)$.
+We now split into two cases depend upon whether $h(u)>1$ or not.
+If $h(u)\leq 1$, then $p+\frac{1}{2\overline{\lambda}}u\in\partial D$. Since $p=\lambda u$, we have
+(6.4) $p+\frac{1}{2\overline{\lambda}}u=\lambda u+\frac{1}{2\overline{\lambda}}u=\left(\frac{2|\lambda|^{2}+1}{2\overline{\lambda}}\right)u.$
+Applying $h$ on the above equation (6.4), we get
+$1=h\left(p+\frac{1}{2\overline{\lambda}}u\right)=\left(\frac{2|\lambda|^{2}+1}{2|\lambda|}\right)h(u)=\frac{2+(h(u))^{2}}{2}$
+Comparing the left-most and the right-most terms in this equation, we obtain $h(u)=0$ which is a contradiction.
+If $h(u)>1$, then $p+\lambda u\in\partial D$. From $p=\lambda u$, we get that $p+\lambda u=2\lambda u$. Applying $h$ on both sides, we get
+$1=h(p+\lambda u)=2|\lambda|h(u)=2,$
+vividly a contradiction.
+This establishes the linear independence of $p$ and $u$.
+Now we get back to finishing the proof that $Z$ is linear, to which end, we get back to the equation (6.3).
+Since $p,u$ are linearly independent, we obtain in-particular, $e^{i2\theta}/2=1$.
+As such a $\theta$ does not exist, this concludes our proof by contradiction to establish that $Z$ is a non-linear retract of $D$.
+∎
+Remark 6.3.
+  1. (i)
+    If $N=2$, then we can replace convex near $p$ by convexity at $p$ because if $D$ is convex at $p$ then $D$ is convex at $p+tu$ for all $t\in\Delta$.
+  2. (ii)
+    The above proof gives an explicit (quadratic) retract.
+Ofcourse, retracts of higher degree as well as other non-linear retracts, may be obtained by taking $\varphi(t)=tp+t\tilde{\psi}(t)u$, where $\tilde{\psi}$ is an arbitrary holomorphic self-map of $\Delta$ fixing the origin.
+We note that after factoring out $t$, the resulting map $t+\tilde{\psi}u$ takes values in $\partial D$ i.e., gives an analytic disc in the boundary $\partial D$.
+The next proposition establishes that this exhausts all possible examples of non-linear (one-dimensional) retracts through the origin in the direction of a boundary point that fails to be $\mathbb{C}$-extremal.
+Indeed, this proposition is an expanded version of theorem 1.12 whose proof will thereby get finished here.
+Before we get to the details, it is very helpful to introduce a notation for brevity: if $a=(a_{1},\ldots,a_{k})$ is any point in $\mathbb{C}^{k}$ (for some $k\in\mathbb{N}$) and each of $v_{1},v_{2},\ldots,v_{k}$ are vectors in $\mathbb{C}^{N}$, then we define $a\cdot v=a_{1}v_{1}+a_{2}v_{2}+\ldots+a_{k}v_{k}$; indeed this is the ‘$a$-linear’ combination of the vectors $v_{1},\ldots,v_{k}$.
+Proposition 6.4.
+Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
+Suppose that $\partial D$ is convex at $p\in\partial D$ but not $\mathbb{C}$-extremal at $p$.
+Let $Z$ be an arbitrary one dimensional retract of $D$ passing through 00 in the direction of $p$; let $\rho:D\to Z$ be a retraction map.
+Then, $Z$ can be be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\tilde{\psi}(t)\cdot v$ where $\tilde{\psi}:\Delta\to\mathbb{C}^{N-1}$ is an analytic disc with $\tilde{\psi}(0)=0$ and where $v$ stands for a list of basis vectors $(v_{1},\ldots,v_{N})$ of $\ker\left(D\rho(0)\right)$.
+Also, after factoring out $t$, the resulting map $t\mapsto p+\tilde{\psi}(t)\cdot v$ gives an analytic disc in the boundary $\partial D$; indeed, this analytic disc is the radial projection of $Z$ onto the boundary of our balanced domain $D$.
+We may also view $Z$ as the graph of the map on ${\rm span}(p)\cap D$ given by $tp\to t\tilde{\psi}(t)\cdot v$ taking values in $\ker\left(D\rho(0)\right)$.
+Suppose next that $\partial D$ is convex near $p$.
+In-case $p$ is a smooth boundary point, we have: $\ker\left(D\rho(0)\right)=H_{p,\mathbb{C}}^{0}$, where $H_{p,\mathbb{C}}^{0}$ denotes the translate of the complex supporting hyperplane to $\overline{D}$ at $p$, to the origin; if $p$ is not a smooth boundary point, $\ker\left(D\rho(0)\right)$ coincides with one of the translates of the supporting hyperplanes at $p$, to the origin.
+In either case, $Z$ may be viewed as the graph of a holomorphic map on the linear retract ${\rm span}(p)\cap D$ taking values in a bounded balanced pseudoconvex domain in the face $F_{\mathbb{C}}(p,\partial D)$, indeed in a linear projection of $D$ into $F_{\mathbb{C}}(p,\partial D)$.
+Proof.
+Let $Z$ be a one dimensional retract of $D$ and $\rho:D\to Z$ be one of the retraction maps on $D$ whose image is $Z$; it suffices to address and focus on the case when $Z$ is non-linear, which we assume for the rest of the proof – by theorem 1.10 and its proof above, this case is non-vacuous and the thrust of the claim is about non-linear retracts.
+By theorem 1.13, $Z$ is a graph of a holomorphic map $\psi$ over $D_{L}:=D\cap L$, where $L$ as usual is the tangent space $T_{0}Z$.
+Note that since $Z$ passes through origin in the direction of $p$, $L=T_{0}Z$ is the subspace spanned by $p$ and $Z$ may be expressed as $Z=\{tp+\psi(t)\cdot v\}$, where
+$\psi(t)\cdot v\;=\;\psi_{1}(t)v_{1}+\ldots+\psi_{N-1}(t)v_{N-1}$
+with $v_{1},\ldots,v_{N-1}$ being a basis for the $(N-1)$-dimensional subspace $M:=\ker\left(D\rho(0)\right)$.
+Note that $\psi(0)=0$ and $\psi^{\prime}(0)=0$ because $Z$ passes through 00 in the direction of $p$ and $p,v_{1},v_{2},...,v_{N-1}$ are linearly independent.
+Note that we can write $\psi(t)=t\widetilde{\psi}(t)$, for some holomorphic map $\widetilde{\psi}$ with $\widetilde{\psi}(0)=0$.
+Now, let $f(t)=h\circ\varphi(t)$, where $h$ as usual is the Minkowski functional and $\varphi(t)=tp+\psi(t)\cdot v$ for $t\in\Delta$; the last mentioned factorization of $\psi$ leads to a similar one about $\varphi$: we may write $\varphi(t)=t\widetilde{\varphi}(t)$ where $\widetilde{\varphi}(t)=p+\widetilde{\psi}(t)\cdot v$.
+Note that pseudoconvexity of $D$ implies $h$ and thereby $f$ is plurisubharmonic.
+By maximum principle of plurisubharmonic functions, we get that for all $t\in\Delta$,
+$|f(t)|\leq\lim\sup_{\begin{subarray}{c}|\lambda|\to 1\\ \lambda\in\Delta\end{subarray}}|f(\lambda)|$
+As $|f(t)|<1$ for all $t\in\Delta$, this limsup is also atmost $1$. Therefore $h(\varphi(t))\leq 1$ for all $t\in\Delta$. It follows that $|t|h(p+\widetilde{\psi}(t)\cdot v)\leq 1$. This implies that
+$\lim\sup_{\begin{subarray}{c}|t|\to 1\\ t\in\Delta\end{subarray}}h(p+\widetilde{\psi}(t)\cdot v)\leq 1$
+Hence $h(p+\widetilde{\psi}(t)\cdot v)\leq 1$ for all $t\in\Delta$.
+Note that $h(p+\widetilde{\psi}(0)\cdot v)=h(p)=1$.
+By another application of the maximum principle, $h(p+\widetilde{\psi}(t)\cdot v)=1$ for all $t\in\Delta$.
+As the set of points $z$ where $h(z)=1$, is always part of $\partial D$, it follows that $\widetilde{\varphi}(t)=p+\widetilde{\psi}(t)\cdot v\in\partial D$ for all $t\in\Delta$.
+As $\varphi$ parametrizes our retract $Z$, it follows that $\widetilde{\varphi}$ gives the radial projection of $Z$ onto the boundary $\partial D$ (equivalently, one may view $Z$ as lying within the ‘surface’ spanned by points in the image of $\widetilde{\varphi}$ i.e., the ‘surface’ obtained by joining points of this analytic disc to the origin).
+We next proceed to relate $\ker\left(D\rho(0)\right)$ with the supporting hyperplane(s) at the boundary point $p$ near which $\partial D$ is convex, thereby $D$ is geometrically convex near $p$ as well.
+For ease of exposition, we first consider the simpler case when $p$ happens to be a smooth boundary point.
+In such a case, the complex supporting hyperplane at $p$ to $\overline{D}$ (near $p$ to be specific as $D$ need not be convex, our assumption is only that $D$ is convex near $p$ and this will be implicitly understood for the rest of this proof) is uniquely determined.
+Denote by $H_{p,\mathbb{C}}^{0}$, the translate of this complex supporting hyperplane to the origin.
+We claim that the sought-for relationship in general first is that of the containment: $\ker\left(D\rho(0)\right)\subset H_{p,\mathbb{C}}^{0}$ and in the setting of this proposition being proved, these subspaces are actually the same.
+The proof of this claim is as follows.
+First we want to show that $\ker(D\rho(0))\subset H_{p,\mathbb{R}}^{0}:=H_{p,\mathbb{R}}-p$, where $H_{p,\mathbb{R}}$ denotes the real supporting hyperplane to $\overline{D}$ in a neighbourhood of $p$.
+Let $T:=D\rho(0)$.
+Suppose there exist $v\in\ker(T)$ such that $v\notin H_{p,\mathbb{R}}^{0}$.
+By definition of $H_{p,\mathbb{R}}$ and $v\notin H_{p,\mathbb{R}}^{0}$, we get $(p+{\rm span}(v))\cap D$ is non-empty.
+Let $w\in(p+{\rm span}(v))\cap D$.
+Writing $w=\lambda v+p$ for some $\lambda\in\mathbb{C}$, we see by the linearity of the map $T$ that
+$T(w)=T(\lambda v+p)=\lambda T(v)+T(p)=p.$
+This contradicts the fact that $T$ maps points of $D$ to (interior) points of $D$.
+This establishes $\ker(T)\subset H_{p,\mathbb{R}}^{0}$.
+Now, remembering that $\ker(T)$ is a $\mathbb{C}$-linear subspace, we conclude that it is contained within the maximal complex subspace of $H_{p,\mathbb{R}}^{0}$ given by the ortho-complement of ${\rm span}(ip)$ in $H_{p,\mathbb{R}}^{0}$.
+As $\ker(T)$ is also of complex co-dimension one, the aforementioned containment is actually as equality: $\ker(T)=H_{p,\mathbb{C}}^{0}$.
+Next we consider the general case that $p$ is (possibly) a non-smooth boundary point.
+Now we claim that $\ker(T)$ is a translate of one of the complex supporting hyperplanes at $p$.
+It suffices to prove that $(p+\ker(T))\cap D=\phi$.
+Suppose there exist $v\in\ker(T)$ such that $p+v\in D$.
+It follows by linearity that $T(p+v)=p$ which as before gives the contradiction that $T$ maps $D$ into $D$.
+This establishes $(p+\ker(T))\cap D=\phi$, thereby that $p+\ker(T)$ is a complex supporting hyperplane at $p$.
+Next, we prove $image(\widetilde{\varphi})\subset p+F_{\mathbb{C}}(p,\partial D)$. We divide the proof into the following steps labelled as claims.
+Claim-1: $image(\widetilde{\varphi})\subset\partial D\cap H_{p,\mathbb{R}}$.
+We already proved that $image(\widetilde{\varphi})\subset\partial D$.
+It remains to show that $image(\widetilde{\varphi})\subset H_{p,\mathbb{R}}$.
+Note that $H_{p,\mathbb{R}}$ can be described by a single linear equation with real coefficients, say
+$\sum_{j=1}^{N}a_{j}{\rm Re}(z_{j})+\sum_{j=1}^{N}b_{j}{\rm Im}(z_{j})=\alpha,$
+for some constants $a_{j},b_{j},\alpha\in\mathbb{R}$. Using the rudimentary identity ${\rm Re}(i\zeta)=-{\rm Im}(\zeta)$ for any complex number $\zeta$, we rewrite the above equation
+${\rm Re}\left(\sum_{j=1}^{N}((a_{j}-ib_{j})z_{j}-\alpha)\right)=0.$
+We shall use below the fact that the left hand side is a pluriharmonic function of $z=(z_{1},\ldots,z_{N})$, which we call $\Phi$ i.e., $\Phi:\mathbb{C}^{N}\to\mathbb{R}$ by
+$\Phi(z)={\rm Re}\left(\sum_{j=1}^{N}((a_{j}-ib_{j})z_{j}-\alpha)\right).$
+Since $\partial D$ is convex near $p$ and $\widetilde{\varphi}(0)=p$, there exist a smaller subdisc $\Delta^{\prime}$ of $\Delta$ such that $\partial D$ is convex near $\widetilde{\varphi}(t)$ for all $t\in\Delta^{\prime}$.
+Observe that that $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}$ is a harmonic function on $\Delta^{\prime}$ and $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}\leq 0$ on $\Delta^{\prime}$.
+Since $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}(0)=0$, by the maximum principle we obtain $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}\equiv 0$ on $\Delta^{\prime}$ Consider the holomorphic function $F:\Delta\to\mathbb{C}$ defined by
+$F(t)=\sum_{j=1}^{N}\left((a_{j}-ib_{j})\widetilde{\varphi}_{j}(t)-\alpha\right),$
+where $\widetilde{\varphi}_{j}$ denotes the $j^{\text{th}}$ component of $\widetilde{\varphi}$.
+Since $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}$ vanishes on $\Delta^{\prime}$, ${\rm Re}(F)$ vanishes on $\Delta^{\prime}$.
+By the open mapping theorem applied to $F$, we deduce that $F$ vanishes on $\Delta^{\prime}$.
+It follows then by the identity principle that $F\equiv 0$ on $\Delta$.
+In particular, ${\rm Re}F=\Phi\circ\widetilde{\varphi}$ vanishes on $\Delta$.
+Hence we conclude that $image(\widetilde{\varphi})\subset H_{p,\mathbb{R}}$.
+This finishes the proof of claim-1.
+To proceed to the next claim, let $V$ be the real affine hull of $\partial D\cap H_{p,\mathbb{R}}$.
+Claim-2: $V=p+F_{\mathbb{R}}(p,\partial D)$, where $F_{\mathbb{R}}(p,\partial D)$ denotes the real face of $\overline{D}$ at $p$.
+First we prove the inclusion $V\subset p+F_{\mathbb{R}}(p,\partial D)$ If $u$ belongs to $V$, then there exist non-zero vectors $v_{i}\in H_{p,\mathbb{R}}^{0}$ and non-zero real $t_{i}$ for $1\leq i\leq k$ such that $u=p+(t_{1}v_{1}+\ldots+t_{k}v_{k})$ and $p+v_{i}\in\partial D$.
+Since $H_{p,\mathbb{R}}^{0}$ is a subspace over $\mathbb{R}$, $p+sv_{i}\in H_{p,\mathbb{R}}$ for all $s\in\mathbb{R}$.
+Recall from the proof of lemma 6.1, we get that $\partial D$ is convex near $p$ implies the sub-additivity of Minkowski functional $h$ and thereby for all $s\in[0,1]$ we have
+$h(p+(1/2-s)v_{i})=h((1/2+s)p+(1/2-s)(p+v_{i}))\leq(1/2+s)h(p)+(1/2-s)h(p+v_{i})=1$
+This means that $p+tv_{i}\in\overline{D}$ for all $t\in\mathbb{R}$ with $|t|<1/2$.
+Using the fact that $H_{p,\mathbb{R}}\cap D=\phi$ and $p+tv_{i}\in H_{p,\mathbb{R}}$, we get that $p+tv_{i}\in\partial D$ for all $t\in\mathbb{R}$ with $|t|<1/2$.
+So $v_{i}\in F_{\mathbb{R}}(p,\partial D)$ for $1\leq i\leq k$.
+Hence $u=p+(t_{1}v_{1}+\ldots+t_{k}v_{k})\in p+F_{\mathbb{R}}(p,\partial D)$.
+To establish the reverse inclusion, let $u\in F_{\mathbb{R}}(p,\partial D)$.
+This means that there exist $\epsilon>0$ such that $p+tu\in\partial D$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$.
+Recall from the proof of theorem 1.10, we get that each supporting hyperplane at $p$ acts as a supporting hyperplane at $p+tu$ for all $t\in\mathbb{R}$ with $|t|<<1$.
+This implies that there exist $\epsilon^{\prime}>0$ such that $p+tu\in\partial D\cap H_{p,\mathbb{R}}$ for all $t\in\mathbb{R}$ with $|t|<\epsilon^{\prime}$.
+This finishes the proof of claim-2.
+Remark 6.5.
+In this claim, we obtain that $F_{\mathbb{C}}(p,\partial D)+p$ is contained in every complex supporting hyperplane to $\partial D$ at $p$ i.e., every complex supporting hyperplane to the boundary at the point $p$ has got to contain the copy of the face $F_{\mathbb{C}}(p,\partial D)$ translated to $p$.
+Continuing with the proof of the proposition 6.4, write $\widetilde{\varphi}(t)=p+\widetilde{\phi}(t)$, where $\widetilde{\phi}:\Delta\to\mathbb{C}^{N}$ is the holomorphic map given by $\widetilde{\phi}(t)=\widetilde{\psi}(t)\cdot v$.
+Let $L_{p}$ denote the smallest $\mathbb{R}$-linear subspace of $\mathbb{C}^{N}$ containing $image({\widetilde{\phi}})$ ($L_{p}$ is of-course uniquely determined as the intersection of all such $\mathbb{R}$-linear subspaces).
+We now prove that $L_{p}$ is a complex vector space.
+Note that $L_{p}$ is the solution of some (finite) system of linear equations with real coefficients, where each linear equation can be written in the following form
+$\sum_{j=1}^{N}a_{j}{\rm Re}(z_{j})+\sum_{j=1}^{N}b_{j}{\rm Im}(z_{j})=0.$
+Since $image(\widetilde{\phi})$ is contained in $L_{p}$, we have
+$\sum_{j=1}^{N}a_{j}{\rm Re}(\widetilde{\phi}_{j})+\sum_{j=1}^{N}b_{j}{\rm Im}(\widetilde{\phi}_{j})=0,$
+where $\widetilde{\phi}=(\widetilde{\phi}_{1},\ldots,\widetilde{\phi}_{N})$.
+Recalling the identity ${\rm Re}(i\zeta)=-{\rm Im}(\zeta)$, the above equation can be rewritten as ${\rm Re}\left(\sum_{j=1}^{N}(a_{j}-ib_{j})\widetilde{\phi}_{j}\right)=0$.
+Owing to the Cauchy–Riemann equations, $\sum_{j=1}^{N}(a_{j}-ib_{j})\widetilde{\phi}_{j}=0$.
+In particular, ${\rm Im}\left(\sum_{j=1}^{N}(a_{j}-ib_{j})\widetilde{\phi}_{j}\right)=0$ where-from: ${\rm Re}\left(i\left(\sum_{j=1}^{N}(a_{j}-ib_{j})\widetilde{\phi}_{j}\right)\right)=0$.
+Application of this observation to each of the linear equations which define $L_{p}$, we infer that $L_{p}$ is closed under multiplication by $i$.
+Hence $L_{p}$ is a complex linear subspace.
+Recalling the results established in ‘Claim-1’ and ‘Claim-2’, we then get that $image(\widetilde{\phi})\subset F_{\mathbb{R}}(p,\partial D)$.
+Since $L_{p}$ is a complex vector space, $image(\widetilde{\phi})\subset iF_{\mathbb{R}}(p,\partial D)$.
+Remembering the fact that $F_{\mathbb{C}}(p,\partial D)=F_{\mathbb{R}}(p,\partial D)\cap iF_{\mathbb{R}}(p,\partial D)$, we get that $image(\widetilde{\phi})\subset F_{\mathbb{C}}(p,\partial D)$.
+Hence $image(\widetilde{\varphi})\subset p+F_{\mathbb{C}}(p,\partial D)$ and thereby, $image(\widetilde{\varphi})\subset(p+F_{\mathbb{C}})\cap\partial D$.
+The last remaining part of the proposition comprises only in recasting $Z$ – realized as the image of $\widetilde{\varphi}$ above – as a graph on ${\rm span}(p)\cap D$ which as we know is a linear retract owing to $p$ being a convex boundary point of $D$.
+To be brief, let us only mention that this is facilitated by the decomposition:
+$\mathbb{C}^{N}\;=\;{\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)\oplus\tilde{V},$
+where $\tilde{V}$ is a complementary subspace to ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$.
+If $\pi$ denotes the projection of $\mathbb{C}^{N}$ onto ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$, then note that $\pi(Z)=Z$.
+Indeed, this is just a re-statement of the fact that $Z$ is contained within ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$ i.e., has no $\tilde{V}$-component.
+From the fact that $Z=image(\widetilde{\varphi})$ and the expression of $\widetilde{\varphi}(t)$, it is now easy to see that $Z$ is the graph of the holomorphic map on ${\rm span}(p)$ given by $tp\to t\tilde{\psi}\cdot v$; where $\tilde{\psi}(t)\cdot v$ is a holomorphic map from $\Delta$ into the face $F_{\mathbb{C}}(p,\partial D)$ with the property that $p+\tilde{\psi}(t)\cdot v$ is contained within the connected component of
+$\left(p+F_{\mathbb{C}}(p,\partial D)\right)\cap\partial D;$
+indeed in its interior, as soon as $\tilde{\psi}\not\equiv 0$.
+Noting that this connected component is contained within the affine linear projection of $D$ into $p+F_{\mathbb{C}}(p,\partial D)$, we are led to the last statement of the proposition; translating this projection back to the origin, gives a bounded balanced pseudoconvex domain which is a projection of $D$ into the lower dimensional complex linear subspace given by the face $F_{\mathbb{C}}(p,\partial D)$.
+This finishes the proof in all, both of theorem 1.12 as well as the assertions of this proposition 6.4.
+∎
+Remark 6.6.
+When $N=2$, we may rephrase the conclusion above more precisely as follows: every non-linear retract $Z$ passing through the origin in the direction of $p$ can be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\widetilde{\psi}(t)u$, where $u$ belongs to the face $F_{\mathbb{C}}(p,\partial D)$ and $\widetilde{\psi}$ maps $\Delta$ holomorphically onto bounded domain in $\mathbb{C}$ containing the origin, with $\psi(0)=0$ and $\psi^{\prime}(0)=0$.
+The proof of this is as follows.
+First observe that from the proof of above proposition, we obtain there exist non-zero $v\in\mathbb{C}^{N}$ such that $p+\widetilde{\psi}(t)v\in\partial D$ for all $t\in\Delta$.
+Since $Z$ is non-linear, $\widetilde{\psi}$ is not identically zero.
+By continuity of $\widetilde{\psi}$ and $\widetilde{\psi}(0)=0$, we get $v\in F_{\mathbb{C}}(p,\partial D)$.
+It follows from the connectedness of image of $\widetilde{\psi}$, that $\widetilde{\psi}$ maps $\Delta$ into $\widetilde{\Delta}$.
+By the open mapping theorem, $\widetilde{\psi}$ maps $\Delta$ onto an open subset $\widetilde{\psi}(\Delta)$ within ${\rm int}(\widetilde{\Delta})$.
+Remark 6.7.
+Though there may well be plenty of retraction maps $\rho$ on $D$ such that $image(\rho)=Z$, we observe that for every retraction map $\rho$, we have $\ker D\rho(0)\supset F_{\mathbb{C}}(p,\partial\Omega)$.
+This can be seen by an argument by contradiction as follows.
+If possible assume that there exist $u\in F_{\mathbb{C}}(p,\partial\Omega)$ such that $D\rho(0)(u)\neq 0$.
+Let $u^{\prime}:=D\rho(0)(u)$.
+Since $D\rho(0)$ maps $\mathbb{C}^{N}$ onto ${\rm span}(p)$, $u^{\prime}=\lambda p$ for some $\lambda\in\mathbb{C}$.
+Since $D$ is convex near $p$, we get by recalling the proof of theorem 1.10, that $p$ and $u$ are linearly independent.
+It follows that $u^{\prime}$ and $u$ are linearly independent.
+Hence $\dim(\ker\left(D\rho(0\right))<N-1$, contradicting the fact that $\ker\left(D\rho(0)\right)$ is of complex co-dimension one.
+Now, a few words about examples of domains satisfying the hypotheses of the foregoing results just discussed.
+The simplest examples of domains as in proposition 6.4 and theorem 1.10 are obtained by taking products of bounded balanced domains of holomorphy with some convex boundary points.
+Other examples (which are not products) can be obtained by taking an intersection of sublevel sets of moduli of linear functionals with suitable balanced domains of holomorphy; to just mention one concrete instance of this kind, consider
+$D:=\{(z,w)\in\mathbb{C}^{2}\;:\;|z+w|<c\}\cap B.$
+where $B$ is a strictly convex ball such as $l^{p}$ ball with $p\neq 1,\infty$ and $c=1$.
+In-fact, we can also let $B$ be a non-convex balanced domain of holomorphy such as the $\ell_{q}$-‘ball’ for $10<q<$ 1 (in which case we must choose $c$ appropriately for instance $c=1/2$) to obtain more concrete examples.
+Let us now conclude this section with a couple of remarks.
+We formulate the first one as the following corollary to highlight that: as a consequence of lemma 6.1 and our proof of theorem 1.10 immediately verifies the truth of that theorem in an important special case.
+Corollary 6.8.
+Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $p\in\partial D$.
+If $\partial D$ is convex near $p$ and the only one-dimensional retract through the origin in the direction of $p$ is the linear one, then $p$ is a $\mathbb{C}$-extremal boundary point.
+Remark 6.9.
+If $D$ fails to be convex near $p$, then the following two cases arise:
+  1. (i)
+    $D$ is convex at $p$ but not near $p$.
+In this case, $span\{p\}\cap D$ is a linear retract of $D$.
+Non-linear retracts of $D$ may or may not exist i.e., no definite statement can be made in general as it depends on finer features of the geometry of $\partial D$ near $p$.
+  2. (ii)
+    $D$ is neither convex at $p$ nor near $p$. In this case, there are no retracts of $D$ passing through origin in the direction of $p$.
 7. Proof of theorem 1.17
@@
 In light of the above discussion which actually only requires $\mathbb{C}$-extremality, we may readily cite $D$ to be the $\ell^{1}$-ball in which the origin is fixed by every of its automorphisms; a retract through any pair of linearly independent points in $D$ is therefore not ${\rm Aut}(D)$-equivalent to any of the retracts passing through the origin.
-To address other hypotheses of the foregoing theorem briefly, let us remark that a non-trivial holomorphic retraction map is always a submersion except perhaps on a set of measure zero (cf.
-page $74$ of the book [Hrsch]).
+To address other hypotheses of the foregoing theorem briefly, let us remark that a non-trivial holomorphic retraction map is always a submersion except perhaps on a set of measure zero (page $74$ of the book [Hrsch]).
 Nevertheless, we admit that it would be more significant to ascertain whether the origin falls outside such a measure zero set, for some retraction map $\rho$ as in the above theorem; or, if all retraction maps of such a class of domains is a submersion throughout the domain, making the submersive hypotheses redundant.
@@
 8. Proof of proposition 1.6
-Before going into the proof, we need to make some general remarks; to indicate as to why there are enough instances of domains verifying the conditions stipluated in proposition 1.6.
-First let us indicate as to why it may perhaps be of mild interest atleast, to note in passing that one may deduce out of the general facts about retracts and the already mentioned well-known characterization of retracts of polydiscs, yet another proof of the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis namely, the biholomorphic inequivalence of two of the simplest bounded simply connected domains the ball $\mathbb{B}$ and $\Delta^{N}$ (despite their diffeomeorphic equivalence), as soon as the dimension $N$ gets any bigger than one; and other such inequivalences using retracts.
+Before we begin to work out the proof, we need to make some general remarks; to indicate as to why there are enough instances of domains verifying the conditions stipulated in proposition 1.6.
+Indeed, even before this, let us indicate as to why it may perhaps be of mild interest atleast, to note in passing (as we did in the preliminaries section) that: one may deduce out of the general facts about retracts and the already mentioned well-known characterization of retracts of polydiscs, yet another proof of the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis which is, the biholomorphic inequivalence of two of the simplest bounded simply connected domains namely, the ball $\mathbb{B}$ and $\Delta^{N}$ (despite their diffeomeorphic equivalence), as soon as the dimension $N$ gets any bigger than one; and other such inequivalences using retracts.
 We have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (and so that proof is not repeated here).
@@
 $\left||z|-|w|\right|\leq|z+iw|$
-when squared and combined with 7.1, leads to
+when squared and combined with 8.1, leads to
 $|z|^{2}+|w|^{2}<5+2|zw|.$
@@
 Lemma 8.1.
-Let $D$ be any bounded domain in $\mathbb{C}^{N}$ with a Stein neighbourhood basis.
+Let $D$ be any bounded pseudoconvex domain in $\mathbb{C}^{N}$ either with a Stein neighbourhood basis or with a bounded plurisubharmonic defining function.
 Then any holomorphic retraction map on $D\setminus A$, where $A$ is a proper analytic subset of $D$ can be extended to a holomorphic retraction map on $D$.
@@
 Indeed, pick any $a\in A$ and enquire if it is possible for $b=\tilde{\rho}(a)$ to lie in $\partial D$.
-Recalling the local structure of analytic sets, we know that: there exists a complex line segment $\ell\subset D$ through $a$ which intersects $A$ only at $a$.
-If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $D$.
-Since $D$ has a Stein neighbourhood basis, there exists an exhaustion function $\psi$ for $D$ which is plurisubharmonic on an open set containing $D$ in particular.
-Considering then the composition $\psi\circ\tilde{\rho}\circ\varphi$, we see that the maximum principle for this plurisubharmonic function is violated, under our assumption that $b\in\partial D$.
-Thus $\rho(D)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
+Recalling the local structure of analytic sets, we know that: there exist a complex line segment $\ell\subset D$ through $a$ which intersects $A$ only at $a$.
+If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $a$.
+If $D$ admits a bounded plurisubharmonic defining function $\psi$, then we consider the composition $\phi:=\psi\circ\tilde{\rho}\circ\varphi$, which is a subharmonic function on $\Delta^{*}=\Delta\setminus\{0\}$; as it is also bounded (by the boundedness of $\psi$), it follows by that the origin is a removable singularity for this subharmonic function, thereby $\phi$ extends to a subharmonic function on $\Delta$ which we continue to denote by $\phi$.
+But then, the maximum principle for this subharmonic function is violated, under our assumption that $b\in\partial D$.
+We deduce therefore that $b$ cannot lie in $\partial D$ and it follows that $\tilde{\rho}(D)\subset D$.
+Even if the existence of such a $\psi$ is not known, it is indeed true that $\tilde{\rho}(D)$ lies within $D$ (the above argument argument was given to show a simple route under the aforementioned assumption about bounded plurisubharmonic defining functions).
+This follows from the last result in [Abe], owing to our hypothesis that $D$ has a Stein neighbourhood basis; as the work [Abe] has some non-major errors, we also cite the Masters thesis of Persson [Persson], wherein the errors are set right with a better exposition.
+Thus as $\tilde{\rho}(a)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
 We finish off by verifying that $\tilde{\rho}$ is a retraction on $D$.
-Indeed, we need only verify the required idempotence equation along $A$, as outside of $A$, $\tilde{\rho}$ is the given retraction map $\rho$ on $D\setminus A$.
+Indeed, we need only verify the required idempotence equation along $A$ (as outside of $A$: $\tilde{\rho}$ is the given retraction map $\rho$ on $D\setminus A$).
 To do this, pick an arbitrary $a\in A$ and a sequence $(z_{n})\subset D\setminus A$ with $z_{n}$ converging to $a$.
@@
 ∎
-We now detail the proof of proposition 1.5 and lay down some of its generalizations as well. This will be used in the sequel.
+Remark 8.2.
+One may want to know to what extent we may weaken the hypotheses of the above lemma.
+We do not pursue this here but draw attention to the fact that rather simple counterexamples exist to show that boundedness cannot be dropped easily.
+Indeed, consider the unbounded domain $D=\mathbb{C}^{2}\setminus A$, where $A=\{z=0\}$, the $w$-axis.
+Notice then that $R^{A}=\{(z,1/z):z\in\mathbb{C}^{*}\}$ is a retract of $D$.
+We observe that there exists no retract $R$ of $\mathbb{C}^{2}$ such that $R^{A}=R\cap(D\setminus A)$.
+Indeed, this follows from the fact that the (smooth) analytic variety $R^{A}$ described by the equation $zw-1=0$, is irreducible, together with the principle of analytic continuation for complex analytic sets.
+Notwithstanding the example in the above remark, we may certainly remove analytic varieties of higher codimension to attain the analogue of the foregoing lemma for the case when the domain is the entire complex Euclidean space itself: to adapt the essential ideas of its proof to deal with the unboundedness now, we only need to employ (theorem 5B in [Whitney]) in place of the standard Riemann removable singularities theorem.
+We shall not repeat the details and be content with stating it as follows.
+Corollary 8.3.
+(Corollary to the proof of lemma 8.1) Let $A$ be any analytic variety of codimension at-least two in $\mathbb{C}^{N}$.
+Then every retract $R^{A}$ of $\mathbb{C}^{N}\setminus A$ is of the form $R^{A}=R\cap(\mathbb{C}^{N}\setminus A)$ for some retract $R$ of $\mathbb{C}^{N}$.
+Without further ado now, we proceed to detailing the proof of proposition 1.5 and lay down some of its generalizations as well. This will be used in the sequel.
 Proof of proposition 1.5 .
-Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ and $v_{2}\in T_{p_{2}}{D_{2}}$.
-Since $D_{1}$ is bounded, there exists $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$.
-Since $D_{2}$ is open, there exists $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$.
+Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ Since $D_{1}$ is bounded, there exist $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$.
+Since $D_{2}$ is open, there exist $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$.
 Consider the domain $D:=D_{1}\times D_{2}$.
@@
 Consider a holomorphic retraction map $G:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $G(z,w)=(z,F(z))$.
-Hence $Z:=\{(z,F(z)):z\in D_{1}\}$ is a 1 dimensional retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(1,v_{2})\in T_{p}Z$.
+Hence $Z:=\{(z,F(z)):z\in D_{1}\}$ is a one-dimensional retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(1,v_{2})\in T_{p}Z$.
 ∎
@@
 Let $v_{1}\in T_{p_{1}}D_{1}$ be any non-zero vector.
-Then there exists $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, and there exists a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ such that $(w_{1},v_{2})\in T_{p}Z$, for some non-zero vector $w_{1}$ with $w_{1}\in span_{\mathbb{C}}\{v_{1}\}$.
+Then there exist $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, there is a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ such that $(w_{1},v_{2})\in T_{p}Z$, for some non-zero vector $w_{1}$ with $w_{1}\in span_{\mathbb{C}}\{v_{1}\}$.
 Proof.
-Since $D_{2}$ is open in $\mathbb{C}^{M}$, there exists a polydisk $\Delta^{M}(p_{2},r_{2})\subset D_{2}$.
+Since $D_{2}$ is open in $\mathbb{C}^{M}$, there exist a polydisk $\Delta^{M}(p_{2},r_{2})\subset D_{2}$.
 Since $D_{1}$ is Caratheodory hyperbolic, we have $\gamma_{D_{1}}(p_{1},v_{1})\neq 0$.
-This means that there exists a non-constant holomorphic map $g:D_{1}\to\Delta$ such that $g(p_{1})=0$ and $g^{\prime}(p_{1})(v_{1})\neq 0$.
+This means that there exist a non-constant holomorphic map $g:D_{1}\to\Delta$ such that $g(p_{1})=0$ and $g^{\prime}(p_{1})(v_{1})\neq 0$.
 So $g^{\prime}(p_{1})$ is a non-zero linear map from $\mathbb{C}^{M}$ onto $\mathbb{C}$.
-Hence there exists $w_{1}\in T_{p_{1}}D_{1}$ with $w_{1}=\lambda v_{1}$ for some $\lambda\in\mathbb{C}$ and $g^{\prime}(p_{1})(w_{1})=1$.
-Choose $v_{2}\in T_{p_{2}}D_{2}$ with $v_{2}\in\Delta^{M}(0,r)$ Consider a holomorphic map $\phi:D_{1}\to\mathbb{C}^{M}$ defined by $\phi(z)=p_{2}+v_{2}g(z)$.
+Hence there exist $w_{1}\in T_{p_{1}}D_{1}$ with $w_{1}=\lambda v_{1}$ for some $\lambda\in\mathbb{C}$ and $g^{\prime}(p_{1})(w_{1})=1$.
+For $v_{2}\in T_{p_{2}}D_{2}$, consider the holomorphic map $\phi:D_{1}\to\mathbb{C}^{M}$ defined by $\phi(z)=p_{2}+v_{2}g(z)$.
 Note that
@@
 Write $r_{2}=(r_{2}^{(1)},\ldots,r_{2}^{(M)})$.
-Choose $\epsilon:=\min\{r_{2}^{(1)},\ldots,r_{2}^{(M)}\}$.
-If $\|v_{2}\|_{\infty}<\epsilon$, then $\phi$ maps $D_{1}$ into $D_{2}$.
+Let $\epsilon:=\min\{r_{2}^{(1)},\ldots,r_{2}^{(M)}\}$.
+Choose $v_{2}\in T_{p_{2}}D_{2}$ with $\|v_{2}\|_{\infty}<\epsilon$.
+Hence $\phi$ maps $D_{1}$ into $D_{2}$.
 Note that $\phi(p_{1})=p_{2}$ and $\phi^{\prime}(p_{1})(w_{1})=v_{2}g^{\prime}(p_{1})(w_{1})=v_{2}$.
-Define a holomorphic map $F:D_{1}\times D_{2}\to D_{1}\times D_{2}$ by $F(z,w)=(z,\phi(z))$.
+Consider the holomorphic map $F:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $F(z,w)=(z,\phi(z))$.
 Again note that $F(p_{1},w)=(p_{1},p_{2})$ and $F^{\prime}(p)(w_{1},w_{2})=(w_{1},\phi^{\prime}(p_{1})(w_{1}))=(w_{1},v_{2})$.
@@
 Indeed, if $F$ were a biholomorphism mapping $D\setminus A$ onto a product domain $D$, then through $F(0)$ we have a retract along those directions given by proposition 8.4 – note that the hypothesis about Caratheodory hyperbolicity therein is verified owing to the fact that $D\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D\setminus A$ must be Caratheodory hyperbolic as well.
-In particular, there always exists a non-trivial retract (necessarily of dimension one here) of $D$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$.
-By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exists a (non-trivial) retract of $D$ through the origin with its tangent direction at 00 belonging to $S^{\prime}$.
+In particular, there always exist a non-trivial retract (necessarily of dimension one here) of $D$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$.
+By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exist a (non-trivial) retract of $D$ through the origin with its tangent direction at 00 belonging to $S^{\prime}$.
 But this contradicts the hypotheses that $D$ is not convex at an open dense subset of $\partial D$ which implies that for each non-convex boundary point $a$ of $D$, there are no retracts of $D\setminus A$ passing through the origin in the direction of $a$.
@@
 Remark 8.5.
-Though we had assumed in the statement of proposition 1.6 that the dimension $n=2$, the above proof goes through for arbitrary dimension to show atleast that $D$ is a balanced domain of holomorphy in $\mathbb{C}^{N}$, $A$ an analytic subset not passing through origin, then $D\setminus A$ is not biholomorphic to a product domain with a one-dimensional factor; in-particular, it is not completely reducible meaning that $D\setminus A$ is not biholomorphic to a product of one-dimensional domains.
+Though we had assumed in the statement of proposition 1.6 that the dimension $n=2$, the above proof goes through for arbitrary dimension to show atleast that if $D$ is a balanced domain of holomorphy in $\mathbb{C}^{N}$ with the same assumptions as in that proposition (and $A$ an analytic subset not passing through origin), then $D\setminus A$ is not biholomorphic to a product domain with a one-dimensional factor; in-particular, it is not completely reducible meaning that $D\setminus A$ is not biholomorphic to a product of one-dimensional domains.
+How about not complete reducibility of the domain but an intermediate one?
+Establishing such a complete non-existence of a product structure for a large class of domains is rather challenging.
+Let us only remark that as a direct application of proposition 8.4, we may derive the following corollary from theorem 1.9 proved in an earlier section.
+Corollary 8.6.
+Let $D_{q}$ denote the $\ell_{q}$-‘ball’ for $10<q<$ 1 and $A$ any analytic subset of $D_{q}$ not containing the origin.
+Then, neither $D_{q}$ nor $D_{q}\setminus A$, is biholomorphic to a product domain.
+We conclude this section with a complete determination of retracts of $D_{h}\setminus A$ through the origin, where $D_{h}$ is the polynomial polyhedron discussed around the beginning of this section.
+As above, $A$ is an analytic subset of $D_{h}$ not passing through origin, which is allowed to be empty but not $D_{h}$ itself.
+This of-course is tantamount to determining retracts of $D_{h}$ by lemma 8.1 and we may therefore forget about $A$.
+Example 8.7.
+The point here is to record at-least one case of a (non-convex but pseudoconvex) domain with $\mathbb{C}$-extremal but not holomorphically extreme boundary; thereby an example where proposition 1.4 does not apply but the modified form of Vesentini’s result as in theorem 1.3 does.
+Indeed, $D_{h}$ serves this purpose as we now discuss.
+First, note that almost every point of $\partial D_{h}$ fails to be holomorphically extreme; specifically, every point of $\partial D_{h}$ which is not in the intersection of the real analytic varieties $H_{1}$ given by $|z^{2}-w^{2}|=1$ and $H_{2}$ given by $|zw|=2$.
+However, every point of $\partial D_{h}$ is a $\mathbb{C}$-extremal boundary point; this is substantiated by realizing $D_{h}$ as the intersection of the pair of balanced domains $D_{1},D_{2}$ given respectively as the sub-level sets $|z^{2}-w^{2}|<1$ and $|zw|<2$ both of which have (weakly) $\mathbb{C}$-extremal boundaries and then using proposition 3.1.
+Now, we contend that for certain special points $p$ (which we shall precisely determine in due course) of $\partial D_{h}$ in the intersection $H_{1}\cap H_{2}$, there exists precisely one retract through the origin in the direction of $p$ which is given by the linear subspace spanned by $p$; along directions given by any other boundary point there is no retract, completing the characterization of retracts through the center of $D_{h}$.
+We shall presently put forth the proof of this claim.
+Let us begin by settling at the outset, the non-existence of retracts along directions given by $D_{h}\setminus(H_{1}\cap H_{2})$: this follows by noting that every point of this open piece of the boundary is a non-convex boundary point of $D_{h}$ and invoking corollary 2.33.
+With this we may restrict ourselves to points of $H_{1}\cap H_{2}\subset\partial D_{h}$ and proceed to discussing retracts through such points alone; it also being borne in mind that by retracts we mean the linear ones for, by theorem 1.3 there are no non-linear holomorphic retracts through the origin in $D_{h}$.
+To proceed forth indeed then, pick any point $p=(p_{1},p_{2})$ in the intersection $H_{1}\cap H_{2}\subset\partial D_{h}$; so, $|p_{2}^{2}-p_{1}^{2}|=1$ and $|p_{1}p_{2}|=2$.
+Consider the linear projection $L$ from $\mathbb{C}^{2}$ onto the one-dimensional complex subspace spanned by $p$ denoted $\langle p\rangle$, defined by $L(z,w)=\lambda(z,w)p$, where
+$\lambda(z,w)=\frac{\overline{p_{1}}z+\overline{p_{2}}w}{|p_{1}|^{2}+|p_{2}|^{2}}.$
+To prove our claim, it suffices to check that $L$ maps $D_{h}$ into itself.
+Towards this, write $L$ in standard component notations $L=(L_{1},L_{2})$; it needs to be verified that the pair of inequalities
+(8.2) $|\left(L_{2}(z,w)\right)^{2}-\left(L_{1}(z,w)\right)^{2}|<1\;\;\;\text{and}\;\;\;|L_{1}(z,w)L_{2}(z,w)|<2,$
+hold for all $(z,w)\in D_{h}$. We shall now deal with the quantities on the left of each of these inequalities; we begin with the first one. Note that
+$|\left(L_{2}(z,w)\right)^{2}-\left(L_{1}(z,w)\right)^{2}|$ $=$ $|\lambda(z,w)|^{2}|p_{2}^{2}-p_{1}^{2}|=|\lambda(z,w)|^{2}$
+It therefore follows that for our task of verifying the first inequality at (8.2), it suffices to find the maximum modulus of $\varphi(z,w)=\overline{p_{1}}z+\overline{p_{2}}w$ on $D_{h}$ or equivalently, on its closure $\overline{D_{h}}$.
+While the compactness of $\overline{D_{h}}$ ensures this maximum being attained, holomorphicity of $\varphi$ allows us to restrict attention on $\partial D_{h}$ for determining the maximum.
+In-fact as $\varphi$ is a linear functional, we may further restrict attention to the extreme points on the boundary of the convex hull of $\overline{D_{h}}$.
+However, rather than rigorously determining this convex hull and its extreme points, it is convenient here to directly proceed to solving our optimization problem on $\overline{D_{h}}$.
+Indeed, let just first use the holomorphicity of $\varphi$ to deduce that the aforementioned maximum – call it $m$ – cannot be attained in the open piece $X_{1}\cup X_{2}$ within the boundary $\partial D_{h}$, where
+$X_{1}$ =$=\{|z^{2}-w^{2}|=1\}\cap\{|zw|<2\},$
+$X_{2}$ =$=\{|z^{2}-w^{2}|<1\}\cap\{|zw|=2\}.$
+To see this, suppose $q=(q_{1},q_{2})$ is a point in $\partial D_{h}$ where $m$ is attained by $|\varphi|$.
+Let us only argue as to why $q$ cannot lie in $X_{1}$; the arguments for $q\not\in X_{2}$ are similar, therefore omitted.
+To argue by contradiction supposing $q\in X_{1}$, we would firstly have $q_{1}^{2}-q_{2}^{2}=e^{i\theta}$ for some $\theta\in(-\pi,\pi]$.
+We collect together all points in $X_{1}$ with this feature i.e., consider $\tilde{X}_{1}$ given by
+$\tilde{X}_{1}=\{(z,w)\in X_{1}\;:\;z^{2}-w^{2}=e^{i\theta}\}.$
+An immediate application of the holomorphic implicit function theorem ensures that this is smooth i.e., $\tilde{X}_{1}$ is a Riemann surface (noting that the gradient of $z^{2}-w^{2}$ is nowhere vanishing on this level set).
+Thus, we are assured of an analytic disc $\psi:\Delta\to\partial D_{h}$ whose image is actually contained in $\tilde{X}_{1}\ni q$; we may assume $\psi(0)=q$.
+Recalling what we have supposed about $q$, we note that the holomorphic function $h=\varphi\circ\psi$ on $\Delta$ assumes its maximum modulus at the origin, thereby reduces to a constant.
+This means that $\varphi$ is constant on $\psi(\Delta)$, thereby on $\tilde{X}_{1}$.
+But then it is not difficult to check that the projection of $\tilde{X}_{1}$ onto $\langle p\rangle$ is a non-trivial open subset of $\langle p\rangle$.
+This means, owing to its linearity, that $\varphi$ must be constant on all of $\mathbb{C}^{2}$ which is manifestly false.
+This concludes the proof of the observation that the maximum of $|\varphi|$ on $\partial D_{h}$ is attained within $H_{1}\cap H_{2}$, which as we shall presently see is a real analytic submanifold of $\mathbb{C}^{2}$.
+In-particular therefore, one may apply the method of Lagrange multipliers to determine this maximum (of $|\varphi|$ on this submanifold).
+However, for our purposes the exact maximum is not necessary and moreover, we shall adopt a more elementary approach, not because it is elementary but owing rather to its relative quickness in directly leading to what is required for our goal here.
+To go about further solving this optimization problem, we rephrase it for convenience by ‘squaring it’, as follows: maximize
+$f(z,w)=|\varphi|^{2}=|\overline{p_{1}}z+\overline{p_{2}}w|^{2}$
+subject to the simultaneous constraints:
+[…]
+Despite the squares, this is of-course equivalent to the previous maximization problem.
+A first convenience of the squares here is that the (real) gradients $\nabla g_{1}$ and $\nabla g_{2}$ are easily checked to be nowhere vanishing along the level sets of $g_{1},g_{2}$ respectively as mentioned above.
+Consequently, it follows by the implicit function theorem that $H_{1}\cap H_{2}$ is a smooth (real analytic) submanifold.
+Now, getting to the aforementioned optimization problem, we shall work in polar coordinates and so write the variables $z=re^{i\alpha},w=se^{i\beta}$ where we may let $\alpha,\beta$ vary in $\mathbb{R}$ rather than on a $2\pi$-length interval such as $(-\pi,\pi]$, for convenience in the sequel; let us also note in passing that $r,s$ are also variables, taking values in non-negative reals, but each of the constraints at (8.3) curtail the possible values of all these variables, as we shall detail below.
+Also write $\overline{p_{1}}=r_{1}e^{i\theta_{1}},\overline{p_{2}}=r_{2}e^{i\theta_{2}}$; here ofcourse, $\theta_{1},\theta_{2}\in(-\pi,\pi]$ and $r_{1},r_{2}$ are fixed numbers which also satisfy both the above constraints at (8.3).
+The first advantage of polar coordinates is of-course the dropping of absolute values in the constraint $|zw|=2$ to write it as $rs=2$, which we use below in the form $r^{2}s^{2}=4$.
+The other constraint (8.3) however, does not yield to such simple re-expressions; but nevertheless is good enough for our problem at hand to be solved.
+To this end, we write it in polar form: $|r^{2}e^{2i\alpha}-s^{2}e^{2i\beta}|^{2}=1$.
+As $r^{2}s^{2}=4$, we may replace the second radial parameter $s$ by the first one $r$, to write it in various forms as:
+$|r^{2}e^{2i\alpha}-\frac{4}{r^{2}}e^{2i\beta}|^{2}=1,$
+$\Longleftrightarrow(r^{2})^{2}+(\frac{4}{r^{2}})^{2}-8Re(e^{2i(\beta-\alpha)})=1,$
+$\Longleftrightarrow\left(r^{2}+\frac{4}{r^{2}}\right)^{2}-8-8\cos(2(\beta-\alpha))=1.$
+Let us record the last equation as:
+(8.4) $\left(r^{2}+\frac{4}{r^{2}}\right)^{2}=1+8(1+\cos(2(\beta-\alpha))).$
+This only finishes expressing the constraints in polar coordinates. To get to the main objective here of maximizing $f$, we likewise express it also in polar form:
+$f(z,w)=|\overline{p_{1}}z+\overline{p_{2}}w|^{2}=|r_{1}e^{i\theta_{1}}re^{i\alpha}+r_{2}e^{i\theta_{2}}se^{i\beta}|^{2}.$
+Since $rs=2$ and likewise $r_{1}r_{2}=2$ as well, the above equation can be written as
+$\left|r_{1}re^{i(\alpha+\theta_{1})}+(\frac{2}{r_{1}})(\frac{2}{r})e^{i(\theta_{2}+\beta)}\right|^{2}$ $=$ $r_{1}^{2}r^{2}+\frac{16}{r_{1}^{2}r^{2}}+8Re(e^{i(\alpha+\theta_{1}-\theta_{2}-\beta)})$
+$=$ $r_{1}^{2}r^{2}+\frac{16}{r_{1}^{2}r^{2}}+8\cos(\alpha+\theta_{1}-\theta_{2}-\beta)$
+$=$ $(r_{1}^{2}+\frac{4}{r_{1}^{2}})(r^{2}+\frac{4}{r^{2}})-4\left(\frac{r^{2}}{r_{1}^{2}}+\frac{r_{1}^{2}}{r^{2}}\right)+$
+$8\cos(\alpha+\theta_{1}-\theta_{2}-\beta)$
+$=$ $(r_{1}^{2}+\frac{4}{r_{1}^{2}})(\sqrt{1+8(1+\cos(2(\beta-\alpha)))})-4\left(\frac{r^{2}}{r_{1}^{2}}+\frac{r_{1}^{2}}{r^{2}}\right)$
+$+8\cos((\beta-\alpha)-(\theta_{1}-\theta_{2})).$
+If we introduce the notations: $\theta=\beta-\alpha$ and $\theta_{0}=\theta_{1}-\theta_{2}$, then our optimization problem now fully expressed in polar coordinates, consists of maximizing
+(8.5) $\tilde{f}(r,\theta)=r_{1}^{2}r^{2}+\frac{16}{r_{1}^{2}r^{2}}+8\cos(\theta-\theta_{0})$
+subject to the constraint
+$\left(r^{2}+\frac{4}{r^{2}}\right)^{2}=1+8(1+cos(2\theta)).$
+Thus our original double constraints problem has now got reduced to a single constraint one.
+To proceed further, let $Q(\theta):=\sqrt{1+8(1+\cos 2\theta)}$.
+We can write $Q(\theta)=\sqrt{1+16\cos^{2}\theta}$, whereby the just mentioned constraint can be rewritten as
+(8.6) $r^{2}+\frac{4}{r^{2}}=Q(\theta)=\sqrt{1+16\cos^{2}\theta}.$
+It must be noted that this contains within it implicitly, certain bounds on $\theta$: a simple application of the inequality about arithmetic and geometric means, the left hand side is bounded below by $4$.
+It follows thereby that
+(8.7) $\cos^{2}\theta\geq\frac{4^{2}-1}{16}\;\Longleftrightarrow\;\sin^{2}\theta\leq\frac{1}{16}.$
+Let us note in passing that these bounds are indeed satisfied for $\theta=\theta_{0}$, due to the trivial reason that $p$ is already given to lie in $H_{1}\cap H_{2}$ i.e., the coordinates of $p$ satisfy the constraints (8.3) and thereby all its consequences derived above.
+Another helpful remark to keep in the background for the sequel is the description of $H:=H_{1}\cap H_{2}$ in polar coordinates as:
+$H:=H_{1}\cap H_{2}=\Big{\{}\left((r,\alpha),(s,\beta)\right)\;:\;rs=2,\;r^{2}+s^{2}=Q(\beta-\alpha)\Big{\}}.$
+Before proceeding ahead, we would also like to record a special case of the foregoing equation (8.6), for later reference; namely taking $\theta=\theta_{0}$ and $r=r_{1}$, we have
+(8.8) $r_{1}^{2}+\frac{4}{r_{1}^{2}}=Q(\theta_{0}).$
+Next we determine $r_{1}$ and more generally $r$ as a function of $\theta$, when $r,\theta$ are to vary in manner satisfying the constraints at (8.3).
+To do this efficiently from the above data, it is convenient to consider the quadratic equation whose roots are $r^{2}$ and $4/r^{2}$ namely, $x^{2}-Q(\theta)x+4=0$ – this follows immediately from the fact that the sum and product of its roots is what is given above.
+On the other hand by the standard formula for the roots of a quadratic, we may also express the roots as $x=(Q\pm\sqrt{Q^{2}-16})/2$.
+From (8.6), we get
+$x=\frac{Q\pm\sqrt{1+16\cos^{2}\theta-16}}{2}=\frac{Q(\theta)\pm\sqrt{1-16\sin^{2}\theta}}{2}.$
+But then we already know that $r^{2}$ and $\frac{4}{r^{2}}$ are the roots of the aforementioned quadratic.
+The desired expression of $r$ as a function of $\theta$, thereby follows; we record in a form suited for direct use in the subsequent analysis.
+(8.9) $r^{2}=\frac{1}{2}\left[\sqrt{1+16\cos^{2}\theta}+\sqrt{1-16\sin^{2}\theta}\right],~{}~{}\frac{4}{r^{2}}=\frac{1}{2}\left[\sqrt{1+16\cos^{2}\theta}-\sqrt{1-16\sin^{2}\theta}\right].$
+Let us remark in passing here that the factor $\sqrt{1-16\sin^{2}\theta}$ is a well-defined real number owing to the bounds on $\theta$ as noted in (8.7).
+We also need to record for later reference a few particular relationships following from the foregoing pair of expressions; firstly the difference:
+$r^{2}-\frac{4}{r^{2}}=\sqrt{1-16\sin^{2}\theta}.$
+Further in particular then, this equation holds for $\theta=\theta_{0}$ and $r=r_{1}$, in which case this reads:
+(8.10) $r_{1}^{2}-\frac{4}{r_{1}^{2}}=\sqrt{1-16\sin^{2}\theta_{0}}.$
+Next, substituting the expressions for $r^{2},\frac{4}{r^{2}}$ from (8.9) into (8.5), we note that $\tilde{f}(r,\theta)$ restricted to $H_{1}\cap H_{2}$ reduces to a function of $\theta$ alone, which we call $h$:
+(8.11) $h(\theta)\;=\tilde{f}_{|_{H_{1}\cap H_{2}}}(r,\theta)=\;\frac{1}{2}\Big{(}r_{1}^{2}(\sqrt{1+16\cos^{2}\theta}+\sqrt{1-16\sin^{2}\theta})\\ +\frac{4}{r_{1}^{2}}(\sqrt{1+16\cos^{2}\theta}-\sqrt{1-16\sin^{2}\theta})\Big{)}+8\cos(\theta-\theta_{0}).$
+After some simplifications, we get
+(8.12) $h(\theta)=\frac{1}{2}\left[\left(r_{1}^{2}+\frac{4}{r_{1}^{2}}\right)(\sqrt{1+16\cos^{2}\theta})+\left(r_{1}^{2}-\frac{4}{r_{1}^{2}}\right)(\sqrt{1-16\sin^{2}\theta})\right]+8\cos(\theta-\theta_{0}).$
+This is the simplified expression of the single variable function $h(\theta)$ that we shall presently put to use.
+As preliminary observations about $h$ that can be read-off from this expression would perhaps be that it is a periodic function, indeed $2\pi$-periodic and smooth (in-fact real analytic) on the entire real line but alas, this is not true: the occurrence of the term $1-16\sin^{2}\theta$ under the square root forbids it from being real-valued for all real $\theta$.
+Indeed, it must be remembered that it was for this reason that we have throughout the bound on $\theta$ stipulated in (8.7).
+Actually, it is possible to circumvent this issue simply by allowing $h$ to take complex values and fixing a branch of the square root on the simply connected domain $\mathbb{C}\setminus(i\mathbb{R}^{-})$ where $i\mathbb{R}^{-}$ is the negative imaginary axis, for instance.
+While this may ensures well-definedness and even continuity of $h$ on the real line (which will render boundedness owing to periodicity), it is worth noting that differentiability at those values of $\theta$ at which $\sin^{2}\theta=1/16$ persists to be invalid.
+Thus, we shall be content in having the codomain of $h$ restricted to real values; interest in the differentiability stems from the desire to apply standard calculus techniques of locating maxima of the function $h$ whose points of global maximum is what we wish to examine for our ultimate goal about retracts of $D_{h}$, as must not be lost sight of.
+We then have to restrict the domain of the function $h$, as mentioned above and we may do so by taking it to be its natural domain in the real line, namely the set of all those real numbers $\theta$ for which $h(\theta)$ also lies in the real line (and not a non-real complex number).
+Owing to the ‘essential’ periodicity of $h$ as indicated above, we may restrict the domain of $h$ to the union of three intervals $I=I^{-}\cup I_{0}\cup I^{+}$ where these intervals are defined as follows.
+First we define the middle one, namely the interval:
+$I_{0}=\text{ connected component containing the origin in $\mathbb{R}$ of its subset }\Big{\{}\theta\in\mathbb{R}\;:\;\sin^{2}\theta\leq\frac{1}{16}\Big{\}}.$
+More concisely put $I_{0}=[-\epsilon_{0},\epsilon_{0}]$ where $\epsilon_{0}$ stands for the smallest positive number satisfying $\sin^{2}\epsilon_{0}=1/16$.
+The other pair of intervals are defined as: $I^{-}=[-\pi,-\pi+\epsilon_{0}]$ and $I^{+}=[\pi-\epsilon_{0},\pi]$.
+The boundedness of $h$ then follows by the continuity of $h$ on the compact set $I$, as already alluded to; but it is also possible to lay down a bound directly from the explicit expression defining $h$ in (8.12) which estimates are of much use in the sequel.
+(8.13) $|h(\theta)|\leq\frac{1}{2}\left[\left(r_{1}^{2}+\frac{4}{r_{1}^{2}}\right)\sqrt{17}+\left(r_{1}^{2}-\frac{4}{r_{1}^{2}}\right)\right]+8.$
+This bound of-course depends on $r_{1}$ which depends on $p$ but is of-course independent of $\theta$ (so, suffices to finish the verification that $h$ is bounded by rendering an explicit bound).
+In this connection, it must more importantly be noted that the aforementioned dependence on $r_{1}$ means that the function $h$ depends the parameter $\theta_{0}$ and perhaps must be written $h_{\theta_{0}}$; of-course we drop the subscript to prevent overloading of notations, unless emphasis enhances clarity.
+Better estimates on the maximum value of $h$ will be derived but only to the extent that is needed for our goal.
+Indeed, let us recall that our goal does not really require us to determine the maximum but rather only to ascertain whether $\theta_{0}$ is a point at which the maximum is attained.
+This is ofcourse just a problem of single variable calculus if we could solve this just by applying the second derivative test to locate first the local maxima and then glean out of them the desired global maximum.
+To this end, we compute the derivative
+$h^{\prime}(\theta)$ $=$ $\frac{1}{2}\left[(r_{1}^{2}+\frac{4}{r_{1}^{2}})\left(\frac{-32\sin\theta\cos\theta}{2\sqrt{1+16\cos^{2}\theta}}\right)+(r_{1}^{2}-\frac{4}{r_{1}^{2}})\left(\frac{-32\sin\theta\cos\theta}{2\sqrt{1-16\sin^{2}\theta}}\right)\right]$
+$-8\sin(\theta-\theta_{0})$
+$=$ $-8\sin\theta\cos\theta\left[(r_{1}^{2}+\frac{4}{r_{1}^{2}})\left((\frac{1}{\sqrt{1+16\cos^{2}\theta}}\right)+(r_{1}^{2}-\frac{4}{r_{1}^{2}})\left(\frac{1}{\sqrt{1-16\sin^{2}\theta}}\right)\right]$
+$-8\sin(\theta-\theta_{0})$
+From (8.8) and (8.10),
+(8.14) $h^{\prime}(\theta_{0})=-16\sin\theta_{0}\cos\theta_{0}=-8\sin 2\theta_{0}.$
+We proceed next to compute the second order derivative:
+$\frac{-h^{\prime\prime}(\theta)}{8}=\sin\theta\cos\theta\left[\left(r_{1}^{2}+\frac{4}{r_{1}^{2}}\right)\left(\frac{32\sin\theta\cos\theta}{2\sqrt{1+16\cos^{2}\theta}(1+16\cos^{2}\theta)}\right)\right.\\ +\left.\left(r_{1}^{2}-\frac{4}{r_{1}^{2}}\right)\left(\frac{32\sin\theta\cos\theta}{2\sqrt{1-16\sin^{2}\theta}(1-16\sin^{2}\theta)}\right)\right]+\\ \left[\left(r_{1}^{2}+\frac{4}{r_{1}^{2}}\right)\left(\frac{1}{\sqrt{1+16\cos^{2}\theta}}\right)+\left(r_{1}^{2}-\frac{4}{r_{1}^{2}}\right)\left(\frac{1}{\sqrt{1-16\sin^{2}\theta}}\right)\right]\left(\cos^{2}\theta-\sin^{2}\theta\right)+\cos(\theta-\theta_{0}).$
+We simplify this a bit, as
+$\frac{-h^{\prime\prime}(\theta)}{8}=16\sin^{2}\theta\cos^{2}\theta\left[\left(r_{1}^{2}+\frac{4}{r_{1}^{2}}\right)\left(\frac{1}{\sqrt{1+16\cos^{2}\theta}(1+16\cos^{2}\theta)}\right)+\right.\\ \left.\left(r_{1}^{2}-\frac{4}{r_{1}^{2}}\right)\left(\frac{1}{\sqrt{1-16\sin^{2}\theta}(1-16\sin^{2}\theta)}\right)\right]+\\ \left[\left(r_{1}^{2}+\frac{4}{r_{1}^{2}}\right)\left(\frac{1}{\sqrt{1+16\cos^{2}\theta}}\right)+\left(r_{1}^{2}-\frac{4}{r_{1}^{2}}\right)\left(\frac{1}{\sqrt{1-16\sin^{2}\theta}}\right)\right]\left(\cos^{2}\theta-\sin^{2}\theta\right)+\cos(\theta-\theta_{0}).$
+Next we put $\theta=\theta_{0}$ in the above equation, in which case it simplifies considerably; more importantly, we do this only because $\theta_{0}$ is the point of focus for us, as already mentioned.
+From (8.8) and (8.10), we get
+$\frac{-h^{\prime\prime}(\theta_{0})}{8}=16\sin^{2}\theta_{0}\cos^{2}\theta_{0}\left[\frac{1}{1+16\cos^{2}\theta_{0}}+\frac{1}{1-16\sin^{2}\theta_{0}}\right]+2\cos 2\theta_{0}+1$
+(8.15) $\frac{-h^{\prime\prime}(\theta_{0})}{8}=4\sin^{2}2\theta_{0}\left[\frac{1}{1+16\cos^{2}\theta_{0}}+\frac{1}{1-16\sin^{2}\theta_{0}}\right]+2\cos 2\theta_{0}+1$
+to not lose sight of our main goal in the midst of computations, we keep this aside for a moment, and observe that our main goal translates here to checking only whether: $\theta_{0}$ is a point of global maximum for $h$ or not.
+This is due to the following observation: if at all $p$ is a a point of maximum for $f(z,w)=\tilde{f}(r,\theta)$, then $\theta_{0}$ must be a maximum point for $h$.
+Now, owing to its smoothness on the open interval $I^{0}$ (the interior of $I$), a point of global maximum for $h$ on $I^{0}$ if any, is not only a point of local maximum but also a point about which $h$ is smooth and so, the standard derivative tests of calculus are applicable.
+So, if at all $\theta_{0}$ is a point of maximum for $h$ on the compact interval $I$, we must have $h^{\prime}(\theta_{0})=0$ or that $\theta_{0}$ is one of the two boundary points of $I$.
+But then we infer from (8.14) that: $h^{\prime}(\theta_{0})=0$ implies that $\theta_{0}=0,\pm\pi/2\text{ or }\pi$.
+Thus as soon as $\theta_{0}$ is not one among the very few possibilities: 00, ±$\pm\pi/2$, $\pi$, or the remaining end-points of the intervals occurring in the definition of $I$, it cannot even be a point of local extremum for $h$.
+Moreover, the pair of possibilities $\theta_{0}=\pm\pi/2$ are ruled out simply because $h(\theta_{0})$ is not real and therefore not of relevance for our consideration of the function $h$.
+We presently proceed to dispose off the cases $\theta_{0}=\pm\epsilon_{0}$ where $\epsilon_{0}$ be the smallest positive number such that $\sin^{2}\epsilon_{0}=1/16$ which means that ±$\pm\epsilon_{0}$ are the end-points of the interval $I$; by showing that $h_{\theta_{0}}(\theta)$ does not attain its (global) maximum on $I$ at $\theta=\theta_{0}=\epsilon_{0}$ – the argument for the other case $-\theta_{0}$ being similar is skipped.
+Now when $\theta_{0}=\epsilon_{0}$, the expression of $h(\theta)$ simplifies a bit and reads:
+$h(\theta)=2\sqrt{1+16\cos^{2}\theta}+8\cos(\theta-\epsilon_{0}).$
+Note that $h(\epsilon_{0})=2\sqrt{1+16\cos^{2}\epsilon_{0}}+8$.
+Recalling that $\cos^{2}\epsilon_{0}=15/16$, we get $h(\epsilon_{0})=16$.
+To show that this is not the maximum value of $h$, we show that the value at $\epsilon_{0}/2$ is strictly bigger i.e., we claim $h(\frac{\epsilon_{0}}{2})>16$.
+To prove this, we write this claim in various equivalent forms in order to unravel it and reach an inequality that is readily checked to be true, beginning with explicitly writing $h(\frac{\epsilon_{0}}{2})>16$:
+$2\sqrt{1+16\cos^{2}(\frac{\epsilon_{0}}{2})}+8\cos\left(\frac{\epsilon_{0}}{2}\right)>16\iff\sqrt{1+16\cos^{2}(\frac{\epsilon_{0}}{2})}>4\left(2-\cos\left(\frac{\epsilon_{0}}{2}\right)\right)$
+which is equivalent to
+$1+16\cos^{2}\left(\frac{\epsilon_{0}}{2}\right)>16\left(2-\cos\left(\frac{\epsilon_{0}}{2}\right)\right)^{2}.$
+Expanding the square on the right, it should be noted that the term $16\cos^{2}\left(\frac{\epsilon_{0}}{2}\right)$ actually cancels, leading to a reduced but of-course equivalent form:
+$\cos\left(\frac{\epsilon_{0}}{2}\right)>1-\frac{1}{64},$
+which uses the fact that the quantity of the left is positive owing to the definition of $\epsilon_{0}$.
+Its precise value is not at all difficult to compute; indeed the half-angle formula $\cos^{2}\theta=(1+\cos 2\theta)/2$, leads to this value being determined as $\cos(\epsilon_{0}/2)=\sqrt{\left(1+\sqrt{15/16}\right)/2}$.
+So the last inequality (thereby all its equivalent forms prior to it) is equivalent to:
+$\frac{\sqrt{15}}{4}+1>2\left(1-\frac{1}{64}\right)^{2}\iff\sqrt{15}>\frac{1921}{8^{3}}\iff 8^{6}\times 15>(1921)^{2}.$
+As the last inequality is readily checked to be true indeed, this finishes the proof that $h(\epsilon_{0}/2)>h(\epsilon_{0})$, thereby that $h_{\theta_{0}}$ does not attain its global maximum at $\epsilon_{0}$, in the case when $\theta_{0}=\epsilon_{0}$.
+The case $\theta_{0}=\pi-\epsilon_{0}$ and $\theta_{0}=-\pi+\epsilon_{0}$ can also be disposed off by similar arguments – for example, it can be shown that $h(\pi-\epsilon_{0}/2)>h(\pi-\epsilon_{0})$ – which ascertain that $h$ does not attain (global) maximum at these ‘end’-points.
+We now note that the only possible cases left are: $\theta_{0}=0$ and $\theta_{0}=\pi$; in both cases, we see from equation (8.15) that $h^{\prime\prime}(\theta_{0})$ is strictly negative.
+Of-course, despite being the only left cases for $\theta_{0}$, this still only establishes that $\theta_{0}$ is a point of local maximum in such cases.
+We now prove that these cases correspond to points of global maximum.
+First suppose $\theta_{0}=0$: this case means that the arguments in the polar coordinates for $p=(p_{1},p_{2})$ are equal ($\theta_{1}=\theta_{2}$), in which case we may write: $p_{1}=r_{1}e^{-it_{0}},p_{2}=r_{2}e^{-it_{0}}$, where $t_{0}=\theta_{1}=\theta_{2}$, wherein $r_{1},r_{2}$ satisfy the polar analogue of the basic constraints (8.3) namely, (8.8) .
+Recall that this means that $r_{1},r_{2}$ satisfy $r_{1}^{2}-r_{2}^{2}=1$ and $r_{1}r_{2}=2$.
+From this, we deduce the constraint satisfied by $r_{1}$ namely, $r_{1}^{2}-\frac{4}{r_{1}^{2}}=1$ which may be rewritten as a quadratic equation in $r_{1}^{2}$: $r_{1}^{4}-r_{1}^{2}-4=0$.
+Solving this gives $r_{1}^{2}=(1+\sqrt{17})/2$ and therefore, $r_{2}^{2}=8/(1+\sqrt{17})$.
+So, we have
+(8.16) $|p_{1}|^{2}+|p_{2}|^{2}=r_{1}^{2}+r_{2}^{2}=\frac{1+\sqrt{17}}{2}+\frac{8}{1+\sqrt{17}}=\sqrt{17}.$
+We shall furnish the details of the conclusion of this case.
+But before that it is convenient to also deal with the case $\theta_{0}=\pi$ simultaneously now, as the arguments are similar to the above case.
+So, in-case $\theta_{0}=\pi$ we have $\theta_{2}=\theta_{1}-\pi$ just by definition of $\theta_{0}$.
+The coordinates of $p$ may be written, in such a case, as $p_{1}=r_{1}e^{-i\theta_{1}},p_{2}=-r_{2}e^{-i\theta_{1}}$.
+Note then that $Q(\theta_{0})=\sqrt{1+16\cos^{2}\theta_{0}}=\sqrt{17}$.
+Substituting $\theta_{0}=\pi$ in (8.8), we obtain $|p_{1}|^{2}+|p_{2}|^{2}=\sqrt{17}$ again.
+The significance of this number here is that regardless of the value of $\theta_{0}$, the bound $|h_{\theta_{0}}(\theta)|\leq 17$ always holds; this follows from (8.13), the estimate whose usefulness was promised there.
+Indeed, employing that estimate alongwith (8.16) we get $|h(\theta)|\leq\frac{1}{2}[\sqrt{17}\cdot\sqrt{17}+1]+8=17$.
+That is to say, $h$ is bounded in magnitude by $17$, which is equal to $(|p_{1}|^{2}+|p_{2}|^{2})^{2}$.
+Hence, $\theta_{0}=0$ and $\theta_{0}=\pi$ is a point of global maximum.
+What this translates to about our original function $f(z,w)=|\overline{p_{1}}z+\overline{p_{2}}w|^{2}$ from around (8.3) is that: its maximum value on $H_{1}\cap H_{2}$ and in-fact by our earlier arguments on the closure of our entire domain $D_{h}$ is $(|p_{1}|^{2}+|p_{2}|^{2})^{2}$.
+Indeed, recalling the discussions and observations around (8.3) and (8.2), this is equivalent to the $|\lambda(z,w)|<1$ where $\lambda$ is the function defined prior to (8.2) and involved in the two inequalities determining whether $L(z,w)$ lies in $D_{h}$.
+Indeed, the just-mentioned bound on $\lambda$ leads immediately to the first of these inequalities: $|(L_{2}(z,w))^{2}-(L_{1}(z,w))^{2}|<1$; the other inequality at (8.2) is also now immediate as
+$\left|L_{1}(z,w)L_{2}(z,w)\right|=|(\lambda(z,w))^{2}p_{1}p_{2}|=|\lambda(z,w)|^{2}<1.$
+Hence, we conclude that $L$ maps $D_{h}$ onto $D_{h}\cap{\rm span}\{p\}$, thereby finishing the proof that $D_{h}\cap{\rm span}\{p\}$ is indeed a linear retract of $D_{h}$ for points $p$ as above.
+As the details of the foregoing example have been rather lengthy, let us summarize and record it as a proposition.
+Proposition 8.8.
+Let $D_{h}$ be the domain in $\mathbb{C}^{2}$ given by
+$D_{h}=\{(z,w)\in\mathbb{C}^{2}\;:\;|z^{2}-w^{2}|<1,\;|zw|<2\}.$
+Then: firstly, every holomorphic retract through the origin is linear.
+More specifically and precisely, there are no retracts in the directions of $p\in\partial D_{h}\setminus H$, where $H$ is the real analytic manifold cut out by $|z^{2}-w^{2}|=1$ and $|zw|=2$.
+As for points $p\in H$, there is a holomorphic retract in the direction of $p$ if and only if it is of the form $(e^{it_{1}}r_{1},e^{it_{2}}r_{2})$ where $r_{1},r_{2}$ are positive real numbers satisfying $r_{1}^{2}-r_{2}^{2}=\pm 1$ and $r_{1}r_{2}=2$ while $t_{1},t_{2}\in\mathbb{R}$ with $|t_{2}-t_{1}|$ being either 00 or $\pi$ (i.e., the arguments of $p_{1},p_{2}$ are either same or differ by $\pi$); and every such retract is linear.
+Let us remark that there are of-course points $p$ in $H$ other than those of the form mentioned in the theorem and note more importantly, that those points are precisely the ones in the boundary of the domain $D_{h}$ which are farthest from the origin.
+Let us briefly justify these claims.
+For the former, note that for any $\alpha_{0}$ satisfying $160\leq\alpha_{0}\leq 1/$ 16, we can find $p=(p_{1},p_{2})=\left(re^{i\theta_{1}},(2/r)e^{i\theta_{2}}\right)\in H_{1}\cap H_{2}$ such that $\sin^{2}\theta_{0}=\alpha_{0}$, where $\theta_{0}=\theta_{1}-\theta_{2}$.
+Now choose $\theta_{0}\in I$ such that $\sin^{2}\theta_{0}=\alpha_{0}$.
+Write $\theta_{0}=\theta_{1}-\theta_{2}$ for some $\theta_{1},\theta_{2}\in[0,2\pi)$.
+Using this $\theta_{0}$, define
+$r^{2}:=\frac{1}{2}\left[\sqrt{1+16\cos^{2}\theta_{0}}+\sqrt{1-16\sin^{2}\theta_{0}}\right].$
+From (8.4), we get that $p=(re^{i\theta_{1}},\frac{2}{r}e^{i\theta_{2}})$ lies in the intersection of $H_{1}$ and $H_{2}$, finishing the proof of the first claim.
+As for the second claim, note that these points are farthest points from origin because we obtain the following relation from (8.4),
+$|p_{1}|^{2}+|p_{2}|^{2}=r^{2}+\frac{4}{r^{2}}=\sqrt{1+16\cos^{2}\theta_{0}}.$
+Remark 8.9.
+The foregoing arguments – again using theorem 1.3 at the outset to reduce considerations from the generality of holomorphic retracts to linear ones – can be adapted to determine precisely the retracts through the origin of other ‘simpler’ examples of non-convex (but pseudoconvex) bounded Reinhardt domains such as $\mathbb{B}\cap H^{2}_{1/4}$ or $B_{p}\cap H^{2}_{1/4}$, where $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/4\}$ and where $B_{p}$ denotes the standard $\ell^{p}$-ball with $1\leq p<\infty$.
+The case $p=\infty$ is different as in that case (i.e., for $\Delta^{2}\cap H^{2}_{1/4}$) there are non-linear retracts through the origin and therefore, discussed separately below.
+The reason we chose to deal with $D_{h}$ extensively is not because it fails to be Reinhardt (though that is an added virtue) but rather because it fails to be non-convex in a maximal manner: note that there does not exist any bounded domain which is non-convex at all boundary points, as follows by standard considerations involving farthest point(s) from a fixed point in the interior and appealing to the Hahn – Banach theorem; for a balanced domain the next best thing to ask for is one that is non-convex at an open dense subset of its boundary or, that the set of convexity points is of measure zero.
+Both of these conditions are fulfilled by $D_{h}$.
+Despite its simplicity, one can still ask natural questions about the choice of $D_{h}$; without elaborating on them all, let us just mention one: why not consider $D_{q}\cap H^{2}_{1/4}$ (where $D_{q}$ denotes the ‘$\ell_{q}$-ball’ for $10<q<$ 1 as before), as this is also a maximally non-convex domain?
+Our answer to this is that though this is easier to analyse owing to it being Reinhardt, it turns out that through every non-smooth point of $D_{q}\cap H^{2}_{1/4}$ there is a linear retract and is a convex boundary point whereas this is not the case for $D_{h}$: not every non-smooth point of $\partial D_{h}$ is a point of convexity, thereby is an even better example of a maximally non-convex domain.
+Moreover, through almost every point $\partial D_{h}$, there passes analytic discs which lie within $\partial D_{h}$ making it at the other extreme compared to convex domains of finite type which are by far, well-studied.
+Let us conclude this remark by finishing off what we began it with namely, retracts of $B_{p}\cap H^{2}_{1/4}$ for $1\leq p<\infty$ which provide simple examples of balanced domains which are neither convex nor maximally non-convex i.e., lie somewhere in-between these extremes (the role of intersection with $H^{2}_{1/4}$ is mainly to create this intermediate possibility by destroying the convexity of $B_{p}$ and could be replaced with other domains).
+The retracts through the origin of such domains are precisely those points of $\partial B_{p}$ which remain after the intersection with $H^{2}_{1/4}$ is performed.
+Let us also draw attention to the fact that the case $p=1$ is allowed whereas $p=\infty$ is not because there are plenty of non-linear retracts (through the origin) in the polydisc.
+Finally, let us mention in passing from this section to the next that while images of retracts under biholomorphic maps are retracts, this becomes false if we more generally consider proper holomorphic maps and expect a direct extension.
+That is to say, neither the backward nor forward images of retracts under proper (holomorphic) maps need be retracts.
+We may use the examples dealt with here to establish this.
+While $D_{h}$ does not admit any biholomorphic mappings to $\Delta^{2}$, it does admit proper holomorphic maps.
+Indeed, the map $F(z,w)=\left(z^{2}-w^{2},zw/2\right)$ renders a proper map of $D_{h}$ onto $\Delta^{2}$ fixing the origin; note that the inverse image of the linear subspace in $\Delta^{2}$ spanned by $(1,1)$ is not even a submanifold of $D_{h}$ leave alone a retract.
+As for forward images, this does not help; however, we can give even simpler examples!
+Consider the map $(z,w)\mapsto(z^{2},w)$; it effects a proper map from the ellipsoid $E_{4}=\{(z,w)\;:\;|z|^{4}+|w|^{2}<1\}$ to the Euclidean ball $\mathbb{B}$.
+Note that the linear subspace given by $w=z$ is retract of $E_{4}$ (as every one-dimensional linear subspace is).
+However, its image $\{z=w^{2}\}$ being the graph of a non-linear map is not a retract of $\mathbb{B}^{2}$.
 9. Retracts of Unbounded and topologically non-trivial domains – Proofs of theorems 1.181.22
-As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allows topologically non-triviality, the boundedness assumption prevailed.
+As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allows topological non-triviality, as we considered ‘analytic complements’ $D\setminus A$ at the beginning therein, the boundedness assumption prevailed.
 To go beyond, to more general settings, we require the following lemma:
@@
 ∎
-Next we shall detail the proof of retraction mappings of ×$\mathbb{C}\times\Delta$.
+Next we shall detail the complete determination of the retraction maps of ×$\mathbb{C}\times\Delta$.
 Proof of theorem 1.21:.
@@
 $P_{2}(P_{1}+P_{2}+P_{3}+\ldots,0)+P_{3}(P_{1}+P_{2}+P_{3}+\ldots,0)+\ldots=0.$
-Equating degree $2$ terms on both sides, we get $P_{2}(P_{1},0)=0$. This implies that $P_{2}(z,w)=w(Q_{1}(z,w))$, where $Q_{1}(z,w)$ is a polynomial in $z$ and $w$. Substituting $P_{2}$ in 8.2, we get
+Equating degree $2$ terms on both sides, we get $P_{2}(P_{1},0)=0$. This implies that $P_{2}(z,w)=w(Q_{1}(z,w))$, where $Q_{1}(z,w)$ is a polynomial in $z$ and $w$. Substituting $P_{2}$ in 9.2, we get
 $P_{3}(P_{1}+P_{2}+P_{3}+\ldots,0)+P_{4}(P_{1}+P_{2}+P_{3}+\ldots)+\ldots=0$
@@
 Since ×$\mathbb{C}\times\Delta$ is homogenous, upto a conjugation of automorphism, every non-trivial retraction map is of the form $F(z,w)=\left(z+wh(z,w),0\right)$, where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$ or $F(z,w)=(h(w),w)$, where $h\in\mathcal{O}(\Delta)$
-Proof of theorem 1.17.
+Next, we turn to the proofs of theorems dealing with domains or manifolds with much greater topological complexity.
+Proof of proposition 1.20.
 Let $\widetilde{Z}$ be a retract of $(\mathbb{C}\setminus A)\times D$ passing through an arbitrary point $(a,b)\in(\mathbb{C}\setminus A)\times D$.
@@
 To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of ×$\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$.
+Indeed as mentioned earlier, this class contains domains with a wide variety of fundamental (or for that matter higher homotopy and homology) groups.
 Proof of theorem 1.22:.
@@
 10. Retracts of $\mathbb{C}^{2}$; proof of theorem 1.23
-We first give an example of a holomorphic retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ which can neither be realized as the graph over the first factor nor the second.
+To begin with a few words in general about the non-trivial (holomorphic) retracts of the complex Euclidean space $\mathbb{C}^{N}$ of any given dimension $N$, note that all its retracts are contractible complex submanifolds which are non-compact as $\mathbb{C}^{N}$ does not admit any non-trivial compact submanifold.
+Apart from being Stein submanifolds, they are non-hyperbolic; indeed, they are ‘Liouville’ submanifolds as their Kobayashi metric vanishes identically (due to lemma 2.25).
+In-particular, we see as an immediate consequence of the uniformization theorem that any one-dimensional retract $Z$ of $\mathbb{C}^{N}$ is biholomorphic to $\mathbb{C}$; thereby, $Z\times\mathbb{C}^{N-1}$ is biholomorphic to $\mathbb{C}^{N}$.
+Thus, we may decompose $\mathbb{C}^{N}$ as $Z\times\mathbb{C}^{N-1}$ as well, though it may be asked as to why we may want such a decomposition.
+More importantly, our purpose in proving that retracts of the various kinds of bounded balanced domains earlier, are graphs of holomorphic maps, was to cast the retracts in as simple a ‘normal form’ as possible.
+It is natural to first examine if this as literally true for $\mathbb{C}^{N}$.
+Towards this, we first give an example of a holomorphic retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ which can neither be realized as the graph over the first factor nor the second; moreover, it is also not a product of retracts – taking graphs and products of retracts are the two canonical ways in which retracts of product domains arise and this provides an example that cannot be realized as arising directly from either of these ways.
 This is discussed in the following
@@
 $Z:=\{(e^{z},ze^{-z}):z\in\mathbb{C}\}$
-is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},zw~{}e^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
+is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},\;zwe^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
 Proof.
@@
 Note that $h~{}\text{and}~{}g$ are surjective holomorphic functions from $\mathbb{C}$ to $\mathbb{C}^{*}$.
-Let $L_{1}$ be an arbitrary ray emanating from origin.
-For each $z\in\mathbb{C}$, there exists an open disk $D$ such that $D\subset\mathbb{C}\setminus L_{1}$.
-Let $\ell_{1}$ denotes the branch of logarithm on $\mathbb{C}\setminus L_{1}$.
-Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ non-vacuously.
-$L_{2}$ divides $D$ into two regions namely $D^{+}~{}\text{and}~{}D^{-}$.
-Let $\ell_{2}$ denotes the branch of logarithm on $\mathbb{C}\setminus L_{2}$.
-Since $h(w)^{w}=g(w)$ for all $w\in\mathbb{C}$, $w\ell_{1}(h(w))=\ell_{1}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$ and this implies that $h(w)=e^{\frac{\ell_{1}(g(w))}{w}}$.
+Let $L_{1}$ be an arbitrary ray emanating from origin and $z$ an arbitrary point of $\mathbb{C}\setminus L_{1}$.
+Just by the closedness of $L_{1}$ in $\mathbb{C}$, we may pick an open disc $D\subset\mathbb{C}\setminus L_{1}$.
+As $\mathbb{C}\setminus L_{1}$ is simply connected, there exists atleast one (indeed infinitely many) branch of the (complex) logarithm on $\mathbb{C}\setminus L_{1}$; pick one of them and call it $\ell_{1}$.
+Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ Note that $L_{2}$ divides $D$ into two regions which shall be referred to by $D^{+}~{}\text{and}~{}D^{-}$.
+As with $\ell_{1}$, we likewise pick and denote by $\ell_{2}$, a branch of logarithm on $\mathbb{C}\setminus L_{2}$.
+Since $h(w)^{w}=g(w)$ for all $w\in\mathbb{C}$, we have $w\ell_{1}(h(w))=\ell_{1}(g(w))~{}\text{for all}~{}w\in D^{+}\cup D^{-}$ which implies that $h(w)=e^{\frac{\ell_{1}(g(w))}{w}}$.
 Similarly, we get $h(w)=e^{\frac{\ell_{2}(g(w))}{w}}~{}\text{for all}~{}w\in D^{+}\cup D^{-}$.
-On $D^{+}\cup D^{-}$,
-$e^{\frac{\ell_{1}(g(w))}{w}}=e^{\frac{\ell_{2}(g(w))}{w}}$
+It then follows that on $D^{+}\cup D^{-}$,
+$e^{\frac{\ell_{1}(g(w))}{w}}=e^{\frac{\ell_{2}(g(w))}{w}}.$
 This implies that for all $w\in D^{+}\cup D^{-}$, we have
-(9.1) $\ell_{1}(g(w))-\ell_{2}(g(w))=2n\pi iw~{}$
+(10.1) $\ell_{1}(g(w))-\ell_{2}(g(w))=2n\pi iw.$
 But $\ell_{1}(g(w))-\ell_{2}(g(w))$ is an integer multiple of $2\pi i$ for any $w\in D^{+}\cup D^{-}$.
@@
 This implies that $\ell_{2}$ is a branch of logarithm on $\mathbb{C}^{*}$.
-This contradicts the fact that there does not exists a branch of logarithm on $\mathbb{C}^{*}$.
+This contradicts the fact that there is no branch of logarithm on $\mathbb{C}^{*}$ and completes the proof.
 ∎
-We conclude by reporting some results about polynomial retracts of $\mathbb{C}^{2}$, the simplest of domains admitting non-trivial retracts.
-It seems difficult to get a complete characterization of holomorphic retracts of $\mathbb{C}^{2}$.
-However it is possible to deduce a complete characterization of polynomial retracts on $\mathbb{C}^{2}$ from known results about the polynomial ring in complex two variables.
-Even this deduction is not entirely trivial and needs some work which we record here.
+We note that the above retract is however, straightenable and thereby conjugate to a graph – we say that a complex submanifold $M$ of $\mathbb{C}^{N}$ of dimension $k$ is straightenable or rectifiable if we can apply an automorphism of $\mathbb{C}^{N}$ which maps $M$ onto $\mathbb{C}^{k}$.
+Getting examples of retracts which are not conjugate to graphs is more non-trivial but it is not difficult to derive it from the work of Forstneric – Globevnik – Rosay [Forstn].
+To briefly recall their main result, let us recall as they do in the introduction of [Forstn] that Abhyankar and Moh in [Abynkr] and M.Suzuki in [Suzuki] proved that if $P$ is a polynomial embedding of $\mathbb{C}$ into $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ of $\mathbb{C}^{2}$ such that $\psi\circ P(\mathbb{C})=\mathbb{C}\times\{0\}$.
+This result is not true for holomorphic embeddings of $\mathbb{C}$.
+By proposition $1$ in [Forstn], there exists a proper holomorphic embedding $H:\mathbb{C}\to\mathbb{C}^{N}$ such that for no automorphism $\psi$ of $\mathbb{C}^{N}$, $\psi\circ H(\mathbb{C})=\mathbb{C}\times\{0\}\subset\mathbb{C}^{N}$.
+We want to prove that ${\rm image}(H)$ is a retract of $\mathbb{C}^{N}$.
+Towards this the manner in which this proper holomorphic embedding $H$ is defined and attained in the proof of proposition $2$ of their paper [Forstn] is now very briefly recalled.
+Start with the embedding $H_{0}:\mathbb{C}\to\mathbb{C}^{N}$ defined by $H_{0}(\zeta)=(\zeta,0)$.
+For $j\geq 1$, $H_{j}:=\psi_{j}\circ H_{j-1}$, where $\psi_{j}$ is some automorphism of $\mathbb{C}^{N}$.
+These $(H_{j})$’s consitute a sequence of proper embeddings of $\mathbb{C}$ satisfying some properties.
+Taking their uniform limit on compacts $H:=limH_{j}$ yields the desired embedding.
+Consider now, the holomorphic retraction $\rho_{0}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{0}(z,w)=(z,0)$, where $z\in\mathbb{C}$ and $w\in\mathbb{C}^{N-1}$.
+Note that ${\rm image}(H_{0})={\rm image}(\rho_{0})$.
+For $j\geq 1$, let $\rho_{j}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{j}:=\psi_{j}\circ H_{j}\circ\psi_{j}^{-1}$, where $\psi_{j}$ is the automorphism mentioned in the above construction.
+Observe that each $\rho_{j}$ is a holomorphic retraction with ${\rm image}(\rho_{j})={\rm image}(H_{j})$.
+Note that the $\rho_{j}$’s converge uniformly on compacts.
+If we denote the limit map by $\rho$ then, $\rho$ is also a retraction – this follows by passing to the limit in the defining equation of a retraction wherein taking a double limit is justified owing to the uniform convergence.
+Finally, we note that the limit of this retraction map $\rho$ satisfies ${\rm image}(H)={\rm image}(\rho)$.
+Hence we have verfied the existence of a one-dimensional holomorphic retract $Z$ of $\mathbb{C}^{N}$ which is not straightenable, thereby not conjugate to a graph as well.
+We conclude by expositing on results about polynomial retracts of $\mathbb{C}^{2}$, the simplest of domains admitting non-trivial retracts.
+As mentioned in the foregoing, holomorphic retracts in $\mathbb{C}^{2}$ are not in general rectifiable.
+Also, it seems difficult to get a complete characterization of holomorphic retracts of $\mathbb{C}^{2}$.
+However it is possible to deduce a complete characterization of polynomial retracts on $\mathbb{C}^{2}$ from known results about the polynomial ring in two complex variables and this yileds in-particular, that all polynomial retracts are rectifiable.
+Even this deduction is not entirely trivial and needs some work which was done by Shpilrain and Yu in [Shpilrain--Yu], which we exposit upon.
 In order to lay this work out properly, we begin with some definitions.
@@
 Let $K[x,y]$ be the polynomial algebra in two variables over a field $K$ of characteristic 00.
-A subalgebra $R$ of $K[x,y]$ is called a retract if it satisfies any of the following equivalent conditions:
-  1. (R1)
-    There is an idempotent homomorphism (a retraction, or projection) $\varphi:K[x,y]\rightarrow$ $K[x,y]$ such that $\varphi(K[x,y])=R$.
-  2. (R2)
-    There is a homomorphism $\varphi:K[x,y]\rightarrow R$ that fixes every element of $R$.
-  3. (R3)
-    $K[x,y]=R\oplus I$ for some ideal $I$ of the algebra $K[x,y]$.
-  4. (R4)
-    $K[x,y]$ is a projective extension of $R$ in the category of $K$-algebras.
-In other words, there is a splitting exact sequence $1\rightarrow I\rightarrow K[x,y]\rightarrow R\rightarrow 1$, where $I$ is the same ideal as in (R3) above.
+An idempotent $K$-algebra homomorphism $\varphi$ from $K[x,y]$ into itself is called a retraction (or a projection) map on $K[x,y]$.
+The image of such a homomorphism is called a retract of (the polynomial ring/algebra) $K[x,y]$.
+Thus, a subalgebra $R$ of $K[x,y]$ is called a retract if there is an idempotent homomorphism $\varphi:K[x,y]\rightarrow$ $K[x,y]$ such that $\varphi(K[x,y])=R$.
 To provide just one concrete example to illustrate the above definitions, take $p=x+x^{2}y$; then $K[p]$ is a retract of $K[x,y]$ (the retraction map here, is given by $x\mapsto p$ , $y\mapsto 0$).
+Of-course as always, before proceeding to such examples, one can cite the trivial ones, namely $K[x,y]$ or $(0)$; a retract of $K[x,y]$ which is neither of these two extremes, is referred to as a proper retract in this context.
 In [Cost], Costa proved that every proper retract of $K[x,y]$ has the form $K[p]$ for some polynomial $p\in K[x,y]$.
-Next we shall show how polynomial retraction maps on $\mathbb{C}^{2}$ transforms to a retraction map on the polynomial ring $\mathbb{C}[z,w]$. We can achieve this using the following lemma.
+We shall be using this result in the sequel.
+It is helpful in all of what follows, to first record the elementary fact as to how polynomial retraction maps on $\mathbb{C}^{2}$ transform to (or induce) a retraction map on the polynomial ring $\mathbb{C}[z,w]$, as discussed next.
 Lemma 10.4.
-A holomorphic map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\rho(z,w)=(p(z,w),q(z,w))$, is a polynomial retraction if and only if there exists a retraction map $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ whose components are polynomials $p(z,w)$ and $q(z,w)$.
+A holomorphic map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ given by $\rho(z,w)=(p(z,w),q(z,w))$ where $p,q\in\mathbb{C}[z,w]$ is a polynomial retraction if and only if there exists a retraction map $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ which take the standard generators $z,w$ to the polynomials $p(z,w)$ and $q(z,w)$.
 Proof.
-Assume that $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is a retraction map on $\mathbb{C}^{2}$ such that whose components are polynomials namely $p(z,w)$ and $q(z,w)$.
-Define $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ defined by $z\mapsto p(z,w),~{}w\mapsto q(z,w)$.
-Since $\rho\circ\rho=\rho$, we get $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$.
-This imply that $\tilde{\rho}\circ\tilde{\rho}=\tilde{\rho}$.
-So, we get that $\tilde{\rho}$ is a retraction on $\mathbb{C}[z,w]$ whose image is precisely $R$, the subalgebra generated by $p,q\in\mathbb{C}[z,w]$.
-Conversely, assume $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ is a retraction map.
-Then there exists $p(z,w),\\ q(z,w)\in\mathbb{C}[z,w]$ such that $\rho(z)=p(z,w)$ and $\rho(w)=q(z,w)$.
-Note that the retract $R$ is the subalgebra generated by $p(z,w),~{}q(z,w)$.
-Define $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\rho(z,w)=(p(z,w),q(z,w))$.
-Since $p,q\in R$, $\tilde{\rho}(p)=p$ and $\tilde{\rho}(q)=q$.
+Assume that $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is a (holomorphic) retraction map on $\mathbb{C}^{2}$ such that its components namely $p(z,w)$ , $q(z,w)$ are polynomials.
+Define $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ by $z\mapsto p(z,w),~{}w\mapsto q(z,w)$.
+Since $\rho\circ\rho=\rho$, we have $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$.
+This implies that $\hat{\rho}\circ\hat{\rho}=\hat{\rho}$ i.e., $\hat{\rho}$ is a retraction map.
+Conversely, assume $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ is a retraction map.
+Then there exists $p(z,w),\\ q(z,w)\in\mathbb{C}[z,w]$ such that $\hat{\rho}(z)=p(z,w)$ and $\hat{\rho}(w)=q(z,w)$.
+Let the subalgebra generated by $p(z,w)$ and $q(z,w)$, which is a retract of the polynomial algebra $\mathbb{C}[z,w]$ be denoted by $R$.
+Define $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ by $\rho(z,w)=(p(z,w),q(z,w))$.
+Since $p,q\in R$, $\hat{\rho}(p)=p$ and $\hat{\rho}(q)=q$.
 This implies that $p(p(z,w),q(z,w))=p(z,w)$ and $q(p(z,w),q(z,w))=q(z,w)$.
-Hence we get that $\rho$ is a polynomial retraction on $\mathbb{C}^{2}$.
+But this just means that $\rho\circ\rho=\rho$ i.e., $\rho$ is a polynomial retraction mapping on $\mathbb{C}^{2}$.
 10.1. Proof of theorem 1.23
+This is essentially an exposition of the proof of theorem $1.1$ of [Shpilrain--Yu] with greater details.
 Let $Z$ be a non-trivial polynomial retract of $\mathbb{C}^{2}$.
-There exists a retraction map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is of the form $\rho(z,w)=(p(z,w),q(z,w))$, where $p,q\in\mathbb{C}[z,w]$ such that $\rho(\mathbb{C}^{2})=Z$.
-By lemma 9.4, $\rho$ induces a retraction map $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ such that $\tilde{\rho(z)}=p(z,w)$ and $\tilde{\rho}(w)=q(z,w))$, where $p,q\in\mathbb{C}[z,w]$.
-Note that ${\rm image}(\tilde{\rho})=\mathbb{C}[p,q]$, the subalgebra generated by $p,q$.
-By Proposition 3.1 in [Cost], $\mathbb{C}[p,q]=\mathbb{C}[r]$, where $r=r(z,w)\in\mathbb{C}[z,w]$.
-Write $p=q_{1}(r),~{}q=q_{2}(r)$, where $q_{1},q_{2}\in\mathbb{C}[z]$ So $\rho(z,w)=(q_{1}(r(z,w)),q_{2}(r(z,w)))$.
-Suppose $q_{1}$ and $q_{2}$ have degree $n\geq 1$ and $m\geq 1$ respectively.
-By Abhyankar - Moh theorem in [Abynkr], either $n$ divides $m$ or $m$ divides $n$.
-Suppose $deg(q_{1})=k\cdot deg(q_{2})$ for some integer $k\geq 1$.
-Consider an automorphism $\psi:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\psi(z,w)=(z-cw^{k},w)$, where $c$ is chosen so that leading term of polynomial $q_{1}-cq_{2}^{k}$ cancel out.
+There exists a retraction map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ of the form $\rho(z,w)=(p(z,w),q(z,w))$, where $p,q\in\mathbb{C}[z,w]$ such that $\rho(\mathbb{C}^{2})=Z$.
+By lemma 10.4, $\rho$ induces a retraction map $\hat{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ such that $\hat{\rho}(z)=p(z,w)$ and $\hat{\rho}(w)=q(z,w)$, the components of the retraction mapping $\rho$ on $\mathbb{C}^{2}$.
+Note that ${\rm image}(\hat{\rho})=\mathbb{C}[p,q]$, the subalgebra generated by $p,q$ which we shall denote by $R$.
+By theorem 3.5 in [Cost], $\mathbb{C}[p,q]=\mathbb{C}[r]$, for some $r=r(z,w)\in R\subset\mathbb{C}[z,w]$.
+Hence we may write $p=q_{1}(r),~{}q=q_{2}(r)$, for some $q_{1},q_{2}\in\mathbb{C}[z]$ i.e., $q_{1},q_{2}$ are polynomials of a single variable.
+Let us write out this factorization for our retraction mapping $\rho$ on $\mathbb{C}^{2}$, a bit elaborately for convenience in the computations that follow, as
+$\rho(z,w)=\left(q_{1}(r(z,w)),q_{2}(r(z,w))\right).$
+Suppose $q_{1}$ and $q_{2}$ have degree $n\geq 1$ and $m\geq 1$ respectively (the case where atleast one of them has degree zero i.e., is constant, is simpler and dealt with during the course of concluding the proof).
+It follows by the Abhyankar – Moh theorem (main theorem of [Abynkr]) that: either $n$ divides $m$ or $m$ divides $n$.
+Suppose without loss of generality that $\deg(q_{1})=k\cdot\deg(q_{2})$ for some integer $k\geq 1$.
+Consider then the automorphism $\psi:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\psi(z,w)=(z-cw^{k},w)$, where $c$ is chosen so that leading term of the polynomials $q_{1}$ and $cq_{2}^{k}$ in $q_{1}-cq_{2}^{k}$ cancel out.
+The point of considering this polynomial automorphism $\psi$ is due to its effect in rendering a strict reduction of degree of the polynomial $q_{1}$ when we pass from our retraction map $\rho$ to its conjugate by $\psi$, explained as follows:
 $\psi\circ\rho\circ\psi^{-1}(z,w)$ $=$ $\psi\circ\rho(z+cw^{k},w)=\psi(q_{1}(r(z+cw^{k},w)),q_{2}(r(z+cw^{k},w)))$
@@
 where $\tilde{q_{1}}=q_{1}-cq_{2}^{k},~{}\tilde{q_{2}}=q_{2},~{}\tilde{r}(z,w)=r(z+cw^{k},w)$. Note that $\tilde{q_{1}}(\tilde{r}),~{}\tilde{q_{2}}(\tilde{r})$ generates $\mathbb{C}[r]$.
-Continuing this process, we eventually arrive at a pair of polynomials.
-One of which is a constant polynomial $c$, and the other polynomial which is denoted by $h$, then we must have $\mathbb{C}[h]=\mathbb{C}[r]$ (i.e.
+Observe that the degree of the single variable polynomial $\tilde{q}_{1}$ is strictly less than that of $q_{1}$; we shall call the polynomials $\tilde{q}_{1},\tilde{q}_{2}$ now simply by $q_{1},q_{2}$ and repeat the above process of passing from one retraction to another through conjugation by the (new and appropriate) automorphism $\psi$.
+So, continuing this process , we eventually arrive within finitely many steps, at a pair of polynomials such that the following happens.
+One of them is a constant polynomial $c$, and the other polynomial, denoted by $h$, is such that $\mathbb{C}[h]=\mathbb{C}[r]$ (i.e.
 $h=c_{1}r+c_{2}$).
-Finally we get $\phi\in Aut(\mathbb{C}^{2})$ such that
-$\phi\circ\rho\circ\phi^{-1}(z,w)=(c_{1}r(z,w)+c_{2},c).$
-After applying conjugation of affine automorphism, we get $\phi(z,w)$ is of the form $(s(z,w),0)$, where $s(z,w)\in\mathbb{C}[z,w]$ and $s(0,0)=0$.
-Since $\phi\circ\rho\circ\phi^{-1}$ is a retraction map, we will get $s(s(z,w),0)=s(z,w)$.
-Write $s(z,0)=a_{1}z+a_{2}z^{2}+\ldots+a_{k}z^{k}$.
-Note that
-$s(s(z,w),0)=a_{1}s(z,w)+\ldots+a_{k}s(z,w)^{k}=s(z,w).$
-This implies $a_{1}=1$ and $a_{2}=...=a_{k}=0$. So $s(z,w)=z+w\cdot l(z,w)$, where $l(z,w)\in\mathbb{C}[z,w]$. Hence $Z$ is of the form $\{(\zeta,0):\zeta\in\mathbb{C}\}$.
+The composition of the various (finitely many) automorphisms involved in the aforementioned process, gives a polynomial automorphism $\phi$ of $\mathbb{C}^{2}$ such that
+$\left(\phi\circ\rho\circ\phi^{-1}\right)(z,w)\;=\;(c_{1}r(z,w)+c_{2},c).$
+Let $\tilde{\rho}:=\phi\circ\rho\circ\phi^{-1}$, which of-course is a retraction map on $\mathbb{C}^{2}$.
+After conjugating it by an affine automorphism namely $\tilde{\psi}(z,w)=(z,w-c)$, we can ensure that the retraction map – which we shall call $\tilde{\rho}$ again – has a particularly simple form wherein the second component is just zero i.e., $\tilde{\rho}(z,w)$ is of the form $(s(z,w),0)$ for some $s(z,w)\in\mathbb{C}[z,w]$.
+Since $\tilde{\psi}\circ\tilde{\rho}\circ\tilde{\psi}^{-1}$ is a retraction map, it follows that $s(s(z,w),0)=s(z,w)$.
+Writing $s(z,0)=a_{0}+a_{1}z+a_{2}z^{2}+\cdots+a_{k}z^{k}$, we then have:
+$s(s(z,w),0)=a_{0}+a_{1}s(z,w)+\cdots+a_{k}s(z,w)^{k}=s(z,w).$
+Comparing the highest degree monomials on both sides (of the second equality above), we first deduce that: $a_{2}=\cdots=a_{k}=0$ and $a_{1}=0$; it is then immediate that $a_{0}$ must also be zero.
+Thus, $s(z,0)=z$ meaning that among the monomials not divisible by $w$, there are no monomials other than $z$.
+Hence, we conclude that $s$ must be of the form: $s(z,w)=z+w\cdot l(z,w)$, for some $l(z,w)\in\mathbb{C}[z,w]$.
+Therefore in all, it follows from the foregoing arguments that there exists a polynomial automorphism $\Phi$ on $\mathbb{C}^{2}$ such that
+$\tilde{\rho}(z,w)\;=\;\left(\Phi\circ\rho\circ\Phi^{-1}\right)(z,w)\;=\;\left(s(z,w),0\right).$
+Thus, after applying the polynomial change of coordinates given by $\Phi$, the retract $Z$ is transformed into
+$\Phi(Z)=\{(\zeta,0):\zeta\in\mathbb{C}\},$
+finishing the proof of the theorem that all non-trivial polynomial retracts of $\mathbb{C}^{2}$ can be rectified. ∎

### v7

@@
 Specifically, we show that every (holomorphic) retract of $B$ passing through its center (i.e., the origin), is the graph of a holomorphic map over a linear subspace of $B$.
-If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics; the necessity of the boundary points to be extreme points is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well.
-This revisit of Vesentini’s result leads to a small modification that is useful to provide interesting examples; specifically, in completely determining the retracts through the origin of certain balanced domains with analytic varieties in their boundaries.
-The converse of Vesentini’s theorem is established and general results characterizing the non-linear retracts through the origin are attained (for fairly general domains in this context) when the boundary fails to have extreme points.
-As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ which is submersive at its center, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$.
-To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of strictly convex balls, with respect to various norms and obtain a complete description of retracts passing through the origin, generalizing that of the polydisc due to Heath and Suffridge and giving a simpler proof as well.
-The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$.
+If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics which works when convexity is dropped and $\mathbb{C}$-extremality is replaced by holomorphic extremality.
+The converse of Vesentini’s theorem is also established and general results characterizing the non-linear retracts through the origin are attained (for fairly general domains in this context) when the boundary fails to have extreme points.
+As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ whose derivative at the center is a norm-one projection, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$.
+To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of bounded balanced domains of holomorphy with holomorphic extreme boundaries and obtain a complete description of retracts passing through the origin.
+The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$ (a description of all of its retracts being attained by the aforementioned result).
 Along the way, we discuss examples of non-convex but pseudoconvex bounded balanced domains in which most one-dimensional linear subspaces fail to be holomorphic retracts, in good contrast to the convex case.
-Indeed, for the simplest possible case here namely the ‘$\ell^{q}$-ball’ for $10<q<$ 1, we obtain a complete list of all retracts through the origin of each possible dimension; this enables an illustration of applying retracts to establishing biholomorphic inequivalences.
+Indeed, for the simplest possible case here namely the $\ell^{q}$-ball’ for $0<q<1$ as well as its ‘anisotropic’ analogues, we obtain a complete list of all retracts through the origin of each possible dimension; whether $q$ is less than $1$ or not, retracts through the origin are shown to be biholomorphic copies of lower dimensional $\ell^{q}$-balls.
+The same is done also for a class of domains lying at the other extreme in their Levi geometry but nevertheless share with $\ell^{q}$-balls the feature of their boundaries being devoid of complex line segments; namely, balanced $\mathbb{C}$-extremal analytic polyhedra.
+In the course of achieving this, we prove a generalization of Mazet’s Schwarz lemma for pseudoconvex balanced domains.
+The above mentioned determination of retracts enables an illustration of applying retracts to establishing biholomorphic inequivalences.
 To go beyond balanced domains, we then first obtain a complete characterization of retracts of the Hartogs triangle and ‘analytic complements’ thereof.
 Thereafter, we drop the assumption of boundedness as well, to attain similar characterization results for domains which are neither bounded nor topologically trivial.
-We conclude with some expositions about retracts of $\mathbb{C}^{2}$.
+We conclude with remarks about retracts of $\mathbb{C}^{2}$ including an alternative proof about the characterization of its polynomial retracts.
 Key words and phrases:
-Union problem, Kobayashi hyperbolic, Kobayashi corank one, balanced pseudoconvex domains
+Holomorphic Retracts, Balanced pseudoconvex domains, Schwarz lemma, Biholomorphic inequivalences, Union Problem, Kobayashi hyperbolic, Caratheodory geodesics
 1991 Mathematics Subject Classification:
@@
 1. Introduction
-We begin with remarks for motivation and background, possibly a bit elaborate; this owes to the unavailability of a survey or expository article on holomorphic retracts to refer the reader to.
+We begin with remarks for motivation and background, possibly somewhat elaborate; this owes to the unavailability of a survey or expository article on holomorphic retracts to refer the reader to.
 This will also simultaneously help us set up notations and fix terminologies.
-We emphasize that this article is a survey in parts, striving to give references for the known results (to the best of our knowledge).
+While the main purpose of this article is to contribute original results, we wish to put our results systematically in connection with what is known for reasons already indicated.
+In this regard, we emphasize that this article is a survey in parts as well, striving to give references for the known results (to the best of our knowledge).
 By a retraction we shall mean an idempotent map, whereas by a retract, we mean the image of such a map.
 More precisely, we shall be interested in holomorphic retracts i.e., the image of a holomorphic retraction which by definition, is a holomorphic self-map $\rho$ of a domain $D\subset\mathbb{C}^{N}$ (or more generally, a complex manifold) satisfying the idempotent condition $\rho^{\circ 2}:=\rho\circ\rho=\rho$.
-This means that $\rho$ acts as the identity map on $Z=\rho(D)$ which shall be our notation for a retract throughout this article; a trivia worth noting is that $Z$ is precisely the set of points of $D$ fixed by $\rho$.
+This means that $\rho$ acts as the identity map on $Z=\rho(D)$ which shall be our notation for a retract throughout this article; a trivia worth noting is that $Z$ is precisely the set of points in $D$ fixed by $\rho$.
 A pair of extreme or trivial examples of retractions (resp.
@@
 This means that for any domain $D$ in $\mathbb{C}$ or any Riemann surface, any holomorphic self-map $\rho$ satisfying $\rho^{\circ 2}=\rho$ is either a constant map or else, the identity map on $D$.
-Indeed, this follows from the open mapping theorem together with, the idempotency condition which implies that $\rho(D)$ is a closed subset of $D$.
+Indeed, this follows from the open mapping theorem together with the idempotency condition which implies that $\rho(D)$ is a closed subset of $D$.
 In other words, holomorphic retracts are interesting only when the dimension $N$ gets bigger than one.
@@
 While much work has been carried out on proper holomorphic maps and their images, far less investigation has been done about retracts.
-Infact, in addition to (the foregoing remarks and) their role of in certain fundamental mapping problems above, retracts are of significance in complex dynamics as well.
+Infact, in addition to (the foregoing remarks and) their role in certain fundamental mapping problems indicated above, retracts are of significance in complex dynamics as well.
 To just be specific by providing a standard result in this regard, we mention the following theorem which shows how retracts arise in the study of iterates of a holomorphic self-map of taut manifolds.
@@
 where $\gamma$ is an automorphism of $M$. Infact, $\rho$ is itself a limit point of the sequence $\left\{f^{\circ k}\right\}$ and $f$ acts as an automorphism of $M$.
-[…]
+The manifold $M$ occurring in the above theorem – referred to as the limit manifold – is, as is evident, a holomorphic retract of $X$.
+Determining all holomorphic retracts of taut manifolds in general is way too difficult.
+We shall therefore be taking up first, the question of describing the holomorphic retracts in what may seem to be a rather small class, namely that of bounded balanced pseudoconvex domains $D$ in $\mathbb{C}^{N}$; not all such domains need be taut, but all bounded balanced domains which are taut, are contained within this class (this owes to results of Barth [Barth]).
+Whether or not $D$ is taut, a first general fact to be noted is that every retract of $D$ is a Stein manifold.
+This follows from the basic fact that retracts of (connected) complex manifolds are closed connected complex submanifolds and if the manifold is a Stein manifold, then so is every of its retracts; as our class of domains of interest is contained within that comprising domains of holomorphy in $\mathbb{C}^{N}$ which are the simplest instances of Stein manifolds, the retracts in this article, all have this property.
+Perhaps before the just mentioned inheritance of the property of being a Stein manifold, we must recall in a sentence here the elementary fact that various topological properties such as simply connectedness or more generally the triviality of any of the homotopy or homology groups (of the ambient space) is inherited under passage to a retract.
+Indeed in general, if $R\xhookrightarrow{\iota}X$ denotes a (continuous) retraction map on a topological space $X$, then the homomorphism induced by the inclusion $\iota$ at the level of each homotopy or homology group, is injective since $r_{*}\circ\iota_{*}\equiv{\rm identity}$, in standard notations; this may be paraphrased as the general observation that topological complexity cannot increase for a retract (and our results will establish certain finer analogues of this in several special cases in the complex analytic category).
+While this means that all such groups must be trivial for every retract of a contractible space $X$, we cannot deduce as readily out of this, the fact that contractibility is inherited as well.
+Indeed, this inheritance follows directly from the definition of contractibility and the only reason we are spending words here to spell this out is because we shall be dealing with balanced domains for the most part, and all such domains are evidently contractible.
+It may be worth remarking once here that as soon as we restrict our considerations to the category of smooth manifolds, we gain easily defined properties that can be used to distinguish a pair of retracts, namely for instance, the dimension of a retract i.e., as soon as we know that a given pair of retracts of a contractible manifold are of different dimensions, we are ascertained that they are diffeomorphically inequivalent.
+This is just to indicate that classifying retracts of a manifold which is topologically trivial (here, contractible) is not an intractable or meaningless exercise.
+We shall in-fact be focussing our attention in good measure on bounded balanced domains of holomorphy.
+Let us only remark here as to why we do not replace boundedness by the apparently more general assumption of Kobayashi hyperbolicity: a balanced domain is hyperbolic if and only if it is bounded ([Kodama]) whereby in particular, the class of hyperbolic balanced convex domains (in $\mathbb{C}^{N}$) is precisely the class of bounded balanced convex domains.
+Now, boundedness implies hyperbolicity and it is known that a convex domain is taut if and only if it is hyperbolic  [^2] In such a case the convex domain is biholomorphic to a bounded domain but not necessarily to a bounded convex domain, as shown by Zimmer [Zim].
+([Bracci-Saracco])).
+It follows that the foregoing theorem in conjunction with our results would be useful towards obtaining a description of the limit manifolds of (proper) holomorphic self-maps of those bounded balanced convex domains that we consider in this article, in particular.
+Even within this class this question of determining all retracts has indeed been of interest but has been rather a difficult one and a complete description of all retracts is available only for the Euclidean ball $\mathbb{B}$, the polydisc $\Delta^{N}$ and perhaps their Cartesian products (herein and henceforth, all balls are centered at the origin and of unit radius with respect to some norm on $\mathbb{C}^{N}$; when the norm is the $l^{2}$ norm, the corresponding ball referred to as the Euclidean ball is denoted $\mathbb{B}^{N}$ or $\mathbb{B}$ for short; $\Delta$ denotes the standard unit disc in $\mathbb{C}$).
 Let us also mention here for clarity that our focus will be to determine the form of the retracts $Z=F(X)$ rather than retractions mappings $F$.
@@
 This is manifestly brought out already in the simplest of cases for $X$ namely, when $X$ is the bidisc $\Delta^{2}$, a description of whose retracts was attained (indeed for $\Delta^{N}$) in 1981 by Heath and Suffridge in [Heath_Suff].
-The case of the Euclidean ball dealt with in 1974 by Suffridge in [Suffridge-Ball] is exposited well in Rudin’s book [Rud_book_Fn_theory_unitball] (we shall not detail the timeline of the results, beyond the most basic cases of the ball and polydisc).
+The case of the Euclidean ball dealt with in 1974 by Suffridge in [Suffridge-Ball] is exposited well in Rudin’s book [Rud_book_Fn_theory_unitball]; we shall not detail the timeline of the results, beyond the most basic cases of the ball and polydisc.
 As highlighted in the introduction of this article by Heath – Suffridge, there is a wide range of holomorphic retractions of the bidisc with the same image, which we recall here for some later significance for our article.
-Let $10<t<$ 1 and pick an arbitrary holomorphic function $f:\Delta^{2}\to\mathbb{C}$ which is bounded above in modulus by $t(1-t)/2$.
+Let $0<t<1$ and pick an arbitrary holomorphic function $f:\Delta^{2}\to\mathbb{C}$ which is bounded above in modulus by t(1t)/2t(1-t)/2.
 Then the mapping on $\Delta^{2}$, defined by
@@
 This finishes the demonstration that for a given retract $Z$, there may very well be plenty of (as remarked in [Heath_Suff] ‘quite complicated’!) retractions all of whose images are the same $Z$.
-At this juncture, it may tempting to ask as to whether this phenomenon of the existence of rather complicated holomorphic retractions with the same image, owes to the polydisc being a domain that reduces to a product of one-dimensional discs, thereby in-particular, giving rise to a plethora of analytic discs in the boundary.
+At this juncture, it may be tempting to ask as to whether this phenomenon of the existence of rather complicated holomorphic retractions with the same image, owes to the polydisc being a domain that reduces to a product of one-dimensional discs, thereby in-particular, giving rise to a plethora of analytic discs in the boundary.
 That this is not the case is borne out by the fact that the same is true (in-fact, in some sense all the more!
@@
 where the $g_{j}$’s are arbitrary elements of $H^{\infty}(\mathbb{B}^{2})$; then the associated self-map $F$ of $\mathbb{B}^{2}$ given by $F(z,w)=\left(f(z,w),0\right)$ is a retraction map, as simply follows from the fact that $f(z,0)\equiv z$.
-Its image is $\pi_{1}(\mathbb{B}^{2})=\Delta$, where $\pi_{1}$ is the linear projection onto the first factor.
+Its image is $\pi_{1}(\mathbb{B}^{2})=\Delta$, where $\pi_{1}$ is the standard linear projection onto the first factor.
 Hence again, while $f$ and thereby $F$ may well be be very complicated, the retract here is just a disc.
-We are therefore motivated to focus on retracts first more than retraction mappings.
-Let us then begin with the obvious observation that retracts are always fixed point sets of some holomorphic map; but the converse need not be true in general.
-But then again, an easy but noteworthy consequence of theorem 1.1 is that the fixed point set of every holomorphic self-map of any taut manifold $X$, is contained within a non-trivial retract.
-Furthermore, there are results for an exact equality rather than just containment, if so sought for: as shown in [Mazet_Vigue_fxdpnt], [Thai_Fxdpnt_cnvx] among others, the fixed point set of any holomorphic self map of any hyperbolic convex domain in $\mathbb{C}^{N}$ – whenever it is non-empty – is also a retract of the domain.
+We are therefore motivated to focus on retracts, first more than retraction mappings.
+Let us then begin with the obvious observation that retracts are always fixed point sets of some holomorphic map; the converse however is not true in general.
+But then again, an easy but noteworthy consequence of theorem 1.1 is that the fixed point set of every holomorphic self-map on any taut manifold $X$, is contained within a non-trivial retract333these little claims will be justified in the preliminaries and subsequent sections..
+Furthermore, there are results for an exact equality rather than just containment: as shown in [Mazet_Vigue_fxdpnt], [Thai_Fxdpnt_cnvx] among others, the fixed point set of any holomorphic self map of any hyperbolic convex domain in $\mathbb{C}^{N}$ – whenever it is non-empty – is also a retract of the domain.
 Whether or not fixed point sets of a holomorphic self map is the retract of possibly some other holomorphic self map of $X$, a definitive basic feature that all retracts have in common is that every retract $Z$ has the extension property in the sense that every holomorphic function on $Z$ admits (just by composition with the retraction mapping $X$ onto $Z$) a holomorphic extension to $X$.
@@
 Whether retracts are the only subvarieties with this property was examined and shown to be so, in many special cases by Kosiński and McCarthy in the relatively recent work [Kosnski] wherein counterexamples have also been given.
+This is the only sentence that we spend in recalling the recent history about the works on retracts in general, all other citations being actually used in one way or the other in the sequel.
 All this is only towards indicating the significance and motivation for the study of holomorphic retracts, in general.
@@
 As mentioned in the abstract, we intend a systematic presentation; so, to not be amiss about addressing the simplest of questions in the aforementioned first goal then, we recall that in [Vesntn], E.
-Vesentini already proved the following fundamental result about the form of the retracts of certain convex domains.
-More important than the various hypotheses of this result, as pointed out by J.
+Vesentini already attained a fundamental result on the form of retracts through the center of certain convex balls.
+As pointed out by J.
 P.
 Vigue in [Vigue_fxdpnt_cnvx], Vesentini’s approach makes ‘heavy’ use of complex geodesics.
-We shall to begin with, give a relatively elementary proof of this result without using complex geodesics.
-It may be noted that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal; thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$, but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal but indeed $\mathbb{C}$-extremal.
+We shall to begin with, give a relatively elementary proof of a generalization of this result without using complex geodesics.
+We however we emphasize that complex geodesics are conceptually significant here; they are particularly advantageous when the boundary satisfies extremality conditions, but in the presence of analytic varieties in the boundary, it seems fruitful to take a more direct approach.
+For instance, this approach led us among other things, to attain theorem 1.14 giving a fairly precise description of the retracts when extremality conditions are completely dropped in the setting of Vesentini’s theorem.
+To see what we mean by this further, we request the reader to see its proof in section 6 as well as that of proposition 1.3 below (theorem 1.17 is also based on this, though the details are longer).
+Perhaps more importantly, we first note that all Caratheodory geodesics (abbreviated complex $C$-geodesic) are retracts but not conversely  [^4] To mention a simple example consider the copy of $\Delta^{*}$ in $\Delta\times\Delta^{*}$, which is evidently a retract but not a complex $C$-geodesic.
+It may be noted in Vesentini’s theorem that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal; thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$, but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal but indeed $\mathbb{C}$-extremal.
 The notions of $\mathbb{R}$-extremality, $\mathbb{C}$-extremality and the relationship between them is recalled in the following section 2 about preliminaries.
@@
 Likewise, by an affine linear subspace $A^{\prime}$ of $D$, we mean $A^{\prime}=A\cap D$ where $A$ is an affine linear subspace of $\mathbb{C}^{N}$.
-To recall a relevant fact in this parlance, every retract of a ball (centered at the origin, as always in this article) linearly equivalent to the standard Euclidean ball $\mathbb{B}^{N}$ is affine linear.
-We postpone rendering clarity about a potential confusion about the notion of linear retracts to section 3; likewise, for definitions of a couple of terms as well, in order to focus more in this introductory section, on the narrative of our results and work about retracts in perspective of what is already known.
-Now, the condition of convexity on $D$, which entails a certain condition to be satisfied by the entire boundary $\partial D$, can be dropped in various ways.
-To begin with here, let us only remark that we can do so in the following manner; this will actually be useful in dealing with certain concrete examples in the sequel (example 8.7 in section 8).
-This is one another major reason for in revisiting this result of Vesentini.
-Specifically, we shall show the following without using complex geodesics as alluded to earlier.
-Theorem 1.3.
-(A small modification of Vesentini’s result): Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
-Suppose $Z$ is a retract through the origin with the property that: $D\cap T_{0}Z$ is convex with points of $\partial D\cap T_{0}Z$ being $\mathbb{C}$-extremal boundary points of $D$; then $Z$ is a linear subspace of $D$.
-In particular, every $1$-dimensional retract through the origin of a bounded balanced pseudoconvex domain $D$ with $\mathbb{C}$-extremal boundary, is linear.
-Our inclusion of this result first towards a relatively self-contained exposition, has other more important motivations than just a more elementary proof: to furnish examples of balanced non-convex (but pseudoconvex) domains possessing non-trivial analytic varieties in the boundary, wherein the retracts through the center can be completely determined.
-This is accomplished in section 8 wherein the above modified version of Vesentini’s result has also a role to play in other propositions there.
-Also in general, we would not like to be amiss to observe associated facts which are not explicitly recorded in the literature but can be attained by reading proofs of results carefully – these are mostly exposited in the ‘preliminaries’ section; more importantly, such observations are actually useful for us much later in the sequel.
-While we relegate such observations to later, one that we think must be mentioned right away here, is the affirmative answer to the question about the very existence of atleast one retract through the origin of an arbitrary bounded balanced domain of holomorphy without any further assumptions.
-This is verified in proposition 2.35.
-Getting back to the above theorem let us now mention for instance, the apparently strong condition (when $\dim(Z)\geq 2$) on the convexity of $D\cap T_{0}Z$ in the above theorem can be dropped while attaining the same conclusion; specifically, we have the following.
+To recall a relevant fact in this language, every retract of a ball (centered at the origin, as always in this article) linearly equivalent to the standard Euclidean ball $\mathbb{B}^{N}$, is an affine linear subspace.
+A question which arises here is whether linear retracts $D_{L}$ can fail to be realizable as the image of a linear (retraction) map.
+While it is easy to ascertain that this cannot happen, it may be interesting to note that this can happen if ‘linear’ is replaced by ‘affine-linear’ i.e., there exist affine-linear retracts for certain balanced domains wherein they cannot be realized as images of affine-linear retraction maps.
+We postpone rendering clarity about such ambiguities and questions to section 3; likewise, for definitions of a couple of other terms as well, in order to focus more in this introductory section, on the narrative of our results and work about retracts in perspective of what is already known.
+Now, generalizations of Vesentini’s theorem have certainly been obtained in the intervening years.
+Let us only mention a relatively recent, best result in this direction extended a bit more for our applications.
+In discussing this, we also attain a generalization of Mazet’s Schwarz lemma, namely theorem 3.6 of the sequel, that is a key tool for some of the main results about retracts as well as the following proposition.
 Proposition 1.3.
@@
 Then, $Z$ is linear.
-Given this, it may seem questionable as to why we have stated theorem 1.3 in this introductory section; in this regard let us highlight the existence of domains as already alluded to, whose boundaries are not holomorphically extreme but are $\mathbb{C}$-extremal.
-Also, $\mathbb{C}$-extremality is far easier to verify than holomorphic extremality which apparently is a more restrictive condition.
-Indeed, a concrete example of a domain of the kind just mentioned is recorded in proposition 8.8 wherein the domain also serves to show that the converse of the above proposition is false.
-As for the significance of the above proposition 1.4 itself, let us only remark that it will be employed in determining retracts through the origin of all possible dimensions in $\ell_{q}$-‘balls’ as in theorem 1.9 below.
-It also helps in establishing biholomorphic inequivalences, particularly with product domains; for instance, a domain as in the above proposition with holomorphically extreme boundary and the polydisc – this is only to give a quick example here.
-To get new results, we must consider domains with greater complexity, to proceed systematically towards which we deal first with a simple class of domains which allow analytic varieties in their boundaries.
-Stating the essence here differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., biholomorphic to a product domain.
-Let us only indicate the key idea here namely, the fundamental common feature amongst all product domains that: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed.
-Let us just state precisely here the simplest result to this effect to indicate what this means; leaving more general formulations to be laid out in section 8 wherein we shall have other remarks to tie up with.
+We shall employ this in determining retracts through the origin of all possible dimensions in $\ell_{q}$-‘balls’ as in theorem 1.8 below.
+The $\ell_{q}$-balls give the simplest domains of holomorphy with holomorphically extreme but non-convex boundaries, at-least for the case where global $C^{1}$-smoothness of boundary fails.
+Now, the above proposition also helps in establishing biholomorphic inequivalences, particularly with product domains; for instance, a bounded balanced pseudoconvex domain with holomorphically extreme boundary, and the polydisc – this is only to give a quick example here.
+We now discuss other possible results about biholomorphic inequivalences, but only briefly as the focus of this article is more on attaining descriptions of retracts.
+Towards this, we consider domains with greater complexity to proceed systematically towards which, we deal with the simplest class of domains generalizing the polydisc other than products, that are also rich in analytic varieties in their boundaries, namely analytic polyhedra.
+Now, stating the essence of the foregoing differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., not being biholomorphic to a product domain.
+The key idea here is the the following common feature amongst all product domains: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed.
+Let us just state here the simplest result to this effect to indicate what this means; leaving more general formulations to be laid out in section 8 wherein we shall have other remarks to tie up with and furnish applications.
 Proposition 1.4.
 Let $D_{1},D_{2}$ be domains in $\mathbb{C}$ with one of them bounded, say $D_{1}$.
-Then there exists $\epsilon>0$ such that: for every vector $v_{2}\in T_{p_{2}}D_{2}\simeq\mathbb{C}$ with $|v_{2}|<\epsilon$, there exists a retract $Z$ of $D_{1}\times D_{2}$ passing through $p=(p_{1},p_{2})$ with $(1,v_{2})\in T_{p}Z$.
-As already alluded to, this paves the way for establishing the following.
-Proposition 1.6.
-Let $D$ be any bounded balanced domain of holomorphy in $\mathbb{C}^{2}$ (whose Minkowski functional we denote by $h=h_{D}$), which is not convex at an open dense subset of $\partial D$ (that there exist plenty of domains will be discussed later); here by not ‘convex at a boundary point’ $p\in\partial D$, means that there are no ‘supporting linear functionals’ for $p$ i.e., there is no $\mathbb{C}$-linear functional $\ell:\mathbb{C}^{2}\to\mathbb{C}$ with $|\ell|\leq h$ and $|\ell(p)|=h(p)$.
-Let $A$ be any proper analytic subset of $D$ not passing through the origin.
-Then, $D\setminus A$ is not biholomorphic to a product domain.
-Domains as in the above proposition may be regarded as being maximally non-convex in the realm of bounded balanced domains, remembering that it is impossible to have a bounded domain which is non-convex throughout the boundary i.e., at every boundary point.
-While convex domains is a much studied class, the other extreme has by far less been; this proposition contributes towards that.
-An extension of this to higher dimensions is briefly discussed in section 8.
+Let $p$ be an arbitrary point of $D=D_{1}\times D_{2}$.
+Then there exists $\epsilon>0$ such that: for every vector $v_{2}\in T_{p_{2}}D_{2}\simeq\mathbb{C}$ with $|v_{2}|<\epsilon$, there exists a retract $Z$ of the product domain $D$ passing through $p=(p_{1},p_{2})$ with $(1,v_{2})\in T_{p}Z$.
+The version of this for arbitrary (finite) dimensions is formulated in proposition 8.5 which as already alluded to, paves the way for establishing the following.
+Theorem 1.5.
+Let $D$ be a (bounded) balanced $\mathbb{C}$-extremal analytic polyhedron.
+Let $G$ be any domain in $\mathbb{C}^{N}$ admitting retracts for some open piece of directions at each $w\in G$.
+Then $D$ is not biholomorphic to $G$.
+The same is true when $D$ is replaced by $D\setminus A$ as well, where $A$ is any non-trivial analytic set (possibly of codimension one) not containing the origin.
+In particular, $D$ and $D\setminus A$ are irreducible i.e., not biholomorphic to a product domain.
+The precise meaning of what it means for a domain $G$ to admit retracts through a point $w\in G$ for some open piece of directions (based at the origin in $T_{w}G$) is laid out in definition 5.9.
+The condition of $\mathbb{C}$-extremality means that there are no (non-trivial) complex line segments in the boundary.
+As the foregoing contains results involving the conditions of $\mathbb{C}$-extremality and holomorphic extremality, let us mention here that: while the two notions are equivalent in the presence of convexity, this is not so in general, as is manifestly seen in the class of analytic polyhedra.
+The inequivalence of analytic polyhedra with strictly pseudoconvex domains (one of the simplest classes of domains with holomorphically extreme boundaries) is well-known due to Henkin’s 1973 article [henkin].
+Product domains also share the same property: none of them is biholomorphic to any domain with holomorphically extreme boundary; the techniques of Remmert – Stein attribute such inequivalences to the lack of analytic varieties in the boundaries of domains in the latter class.
+The question about the irreducibility of analytic polyhedra $D$ and associated ‘co-analytic domains’ $D\setminus A$ for analytic sets $A\subset D$, particularly when the codimension of $A$ is one, is therefore interesting.
+One cannot hope for results without imposing conditions as there are many analytic polyhedra which are completely reducible, the polydisc being the simplest example.
+The above result shows that irreducibility of $D$ is guaranteed, as soon as the milder $\mathbb{C}$-extremality condition holds at-least within the class of balanced domains.
+We hasten to mention that while there may well be shorter proofs than the one via retracts to prove the irreducibility of those analytic polyhedra $D$ as in the above theorem, it gets somewhat more difficult firstly for the associated co-analytic domains $D\setminus A$ therein; and all the more difficult for more general cases.
+As our purpose in this article is to attain results characterizing retracts more than various applications, we do not strive to include beyond a couple of results of this kind.
 We also hasten to mention that there are better techniques to detect the occurrence of a product structure for domains far beyond the class of balanced domains owing particularly to the very recent work [Bh-Bo-Su], which however requires some quantitative knowledge about the squeezing function.
-However we believe that for balanced domains, the above proposition gives a much simpler approach in practice, since a balanced domain is most often given by its Minkowski functional.
-Needless to say, there are plenty of balanced domains of holomorphy satisfying the non-convexity condition in the above proposition, as will also be discussed in section 8.
-Getting back to retracts, let us start by noting that while Vesentini’s theorem says that all retracts of $D$ (as in that theorem) passing through the origin, are linear, it does not say that all linear subspaces are retracts.
-Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]).
-The question of precisely which linear subspaces are realizable as retracts of $D$ as in the above theorem is significant for us.
-For instance: while every retract of the polydisc $\Delta^{N}$ is a graph, not every graph in $\Delta^{N}$ is a retract; only those which are graphs on linear retracts of $\Delta^{N}$ are, making it important to precisely pin down which linear subspaces are retracts.
-Indeed, this is quite an interesting question in its own right and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections (when we measure the norm of the projection with respect to the operator norm induced by $\mu$)?
-The answer turns out to vary even among the $l^{p}$ norms i.e., depend upon $p$ when we take $\mu=\|\cdot\|_{l^{p}}$ even if we exclude the case $p=2$, which may not be surprising at all for this case is distinguished as the one that arises from an inner product; specifically, the case $p=\infty$ is also different from the other cases.
-Before getting to such differences, let’s mention a unified result.
-As our interest lies in retracts, we shall phrase in it in the language of retracts rather than norm-one projections.
-Specifically, the characterization of linear (thereby holomorphic, owing to Vesentini’s theorem) retracts of the standard $l^{p}$ ball ($1\leq p<\infty,~{}p\neq 2$) can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
+However, we believe that the approach via retracts is worth recording, as there are no universal techniques to establish biholomorphic inequivalences.
+We shall indeed show that there are plenty of balanced polyhedra as in the above theorem, in chapter 8.
+Getting back to retracts, let us note that while Vesentini’s theorem says that all retracts of $D$ as in that theorem, passing through the origin, are linear, it does not say that all linear subspaces are retracts.
+Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the main result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]).
+The question of precisely which linear subspaces are realizable as retracts of $D$ as in Vesentini’s theorem is significant for us.
+For instance: while every retract of the polydisc $\Delta^{N}$ is a graph, not every graph in $\Delta^{N}$ is a retract; only those which are graphs on linear retracts of $\Delta^{N}$ are, making it important to first precisely pin down which linear subspaces are retracts.
+Indeed, this is quite an interesting question in its own right and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections when we measure the norm of the projection on $\mathbb{C}^{N}$ with respect to the operator norm induced by $\mu$?
+There are differences in the answers for the cases $p=1,\infty$ even if we exclude $p=2$, the well-known case distinguished as arising from an inner product, thereby not surprising at all.
+Before getting to such differences, let us mention that the result is uniform for all finite $p$ other than two and just state this unified (known) result.
+As our interest lies in retracts, we shall phrase it in the language of retracts rather than norm-one projections.
+Specifically, the characterization of linear retracts (thereby those holomorphic retracts which pass through the origin, owing to Vesentini’s theorem) of the standard $l^{p}$ ball denoted $B_{\ell^{p}}$, can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
 Theorem 1.6.
@@
 Then $Y\cap B_{\ell^{p}}$ is a linear retract of $B_{\ell^{p}}$ if and only if there exist vectors $\{u_{j}\}_{j=1}^{m}$ of $\ell^{p}$-norm 1 in $\mathbb{C}^{N}$ of the form
-$u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i},~{}1\leq j\leq m,~{}with~{}\sigma_{j}\cap\sigma_{k}=\phi~{}for~{}k\neq j$
+(1.1) $u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i},~1\leq j\leq m,~with~\sigma_{j}\cap\sigma_{k}=\phi~for~k\neq j$
 such that $Y=span\{u_{j}\}_{j=1}^{m}$, where $\sigma_{j}:=supp(u_{j})=\{1\leq i\leq N:u_{j}(i)\neq 0\}$ (and $\lambda_{i}$’s are some complex numbers such that the $\ell^{p}$-norm of $u_{j}$ are equal to $1$).
 Here $e_{i}$ denotes the standard basis of $\mathbb{C}^{N}$.
-We shall discuss alternative formulations of the above result which are more geometrically revealing in section 5.
-In-particular, retracts of $\ell^{p}$-balls are isometric copies of lower dimensional standard $\ell^{p}$-balls.
-Next, we note that the case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, in which all linear subspaces are retracts.
+We shall derive in section 5, alternative formulations (some of which are known) of the above result, which are more geometrically revealing to show that the linear retracts of $\ell^{p}$-balls are precisely isometric copies of lower dimensional standard $\ell^{p}$-balls.
+Stated differently, only those linear subspaces of $\ell^{p}$-balls which are biholomorphic to lower dimensional $\ell^{p}$-balls, turn out to be their holomorphic retracts through the origin.
+Next, we note that the case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, wherein all linear subspaces are retracts.
 We would rather like to draw attention to the more important fact that this result is not true for $p=\infty$ to demonstrate which (by a counterexample as simple as possible), we include an example in section 3.
-This naturally leads to the question as to whether it is possible at all to pin down precisely which linear subspaces of the polydisc $\Delta^{N}$ can be realized as linear retracts (recall that the polydisc is the exceptional case corresponding to $p=\infty$ which is left out by theorem 1.7).
-Towards this, we first note the following characterization of those linear subspaces of $\Delta^{N}$ which are its retracts.
-Indeed, this can readily derived from [Heath_Suff], though it is not stated in these explicit terms; contained within our treatment will be a simpler proof of the main result of [Heath_Suff] about the characterization of all retracts of the polydisc as well as a deduction of the following special one about its linear retracts.
+Nevertheless, the aforementioned geometric characterization of the linear retracts of $\ell^{p}$-balls for finite $p$, as lower dimensional versions of themselves, holds true for the polydisc as well.
+Indeed, this can derived without much difficulty from [Heath_Suff], though it is not stated in these explicit terms.
 Proposition 1.7.
@@
   2. (ii)
-    For each $m\in\{1,2,\ldots,N\}\setminus J$, the scalars $c_{m,j}$, satisfy $\sum_{j\in J}|c_{m,j}|\leq 1$.
+    For each $m\in\{1,2,\ldots,N\}\setminus J$, the scalars $c_{m,j}$ satisfy $\sum_{j\in J}|c_{m,j}|\leq 1$.
+Consequently, non-trivial linear retracts of the polydisc $\Delta^{N}$ are precisely those linear subspaces which are isometric copies of lower dimensional polydiscs $\Delta^{k}$ (for $1\leq k\leq N$).
+It follows that all holomorphic retracts of $\Delta^{N}$ (not necessarily passing through the origin) are biholomorphic copies of lower dimensional polydiscs.
 It is immediate that through any given point of the polydisc, there exist retracts of all possible dimensions ($k=0,1,\ldots,N$), as its automorphism group acts transitively; a little contemplation shows that the transitive action is unnecessary to confirm the existence of such retracts.
-Indeed, for $1\leq p<\infty$, the $\ell_{p}$-ball fails to be homogeneous unlike the polydisc as soon as $p\neq 2$; but it is easy to derive as a corollary of theorem 1.7 that through each point of the $\ell_{p}$-ball, there exist retracts of all possible dimensions, as the retracts through the origin serve to confirm this.
-It is too much to as for retracts along any given set of directions because as the above proposition shows: not every two dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the ball with respect to the $l^{2}$-norm, $\mathbb{B}^{N}$.
+Indeed, for every $1\leq p<\infty$, the $\ell^{p}$-ball fails to be homogeneous unlike the polydisc as soon as $p\neq 2$; but it is easy to derive as a corollary of theorem 1.6 that through each point of the $\ell^{p}$-ball, there exist retracts of all possible dimensions, as the retracts through the origin serve to confirm this.
+It is too much to ask for retracts along any given set of directions because as the above proposition shows: not every two-dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the ball with respect to the $l^{2}$-norm, $\mathbb{B}^{N}$.
 We must not be amiss to note however that every one dimensional linear subspace of a bounded balanced convex domain in $\mathbb{C}^{N}$ is a retract – this follows by the Hahn–Banach theorem.
 It is natural to ask whether this holds if we relax the convexity condition on the domain to pseudoconvexity.
-The answer is in the negative, as was noted by Lempert in [Lempert-FundArt] by indicating an example to complement his passing observation therein that the existence of $2$-dimensional retracts is rather restricted and to the effect that when convexity is relaxed, even $1$-dimensional retracts can fail to exist in abundance.
+The answer is in the negative, as was noted by Lempert in [Lempert-FundArt] by indicating an example to complement his passing observation therein, that the existence of $2$-dimensional retracts is rather restricted i.e., to the effect that when convexity is relaxed, even $1$-dimensional retracts can fail to exist in abundance.
 The only rigorous record of the just-mentioned Lempert’s example is in [Jrncki_invrnt_dst] (specifically in example 11.3.10 therein).
-While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimension as well; moreover this can be achieved entirely through elementary means.
-In-fact, our example is as simple as the standard $\ell^{q}$-‘ball’ for $q<1$ and we record this in the following proposition; an interesting feature of the $\ell_{q}$-‘ball’ in $\mathbb{C}^{N}$ is that apart from failing to be convex, it fails to be concave unlike its real counter-part.
+While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimensions as well; moreover this can be achieved entirely through elementary means.
+In-fact, our example is as simple as the standard $\ell^{q}$-‘ball’ for $q<1$ and we record this in the following proposition; an interesting feature of the $\ell^{q}$-‘ball’ in $\mathbb{C}^{N}$ is that apart from failing to be convex, it fails to be concave unlike its real counter-part.
 Let us also note that while the above-mentioned example of Lempert is a pseudoconvex complete Reinhardt domain much as ours, we believe recording the aforementioned $\ell^{q}$-‘ball’ is worthwhile due to the fact that the literature about retracts of dimension higher than one, is scant at-least compared to complex geodesics and one-dimensional retracts.
-In that regard, we would like to mention the very recent article of [GZ] which appeared on the arXiv, a couple of months after the earlier versions of our article.
+In that regard, we would like to mention the very recent article [GZ]  [^5] This appeared on the arXiv, a couple of months after the earlier versions of our article.
 Ghosh – Zwonek attain in this article [GZ], a complete characterization of the two dimensional retracts (and thereby all retracts) of the three dimensional Lie ball and the tetrablock; this article also has noteworthy general results about retracts of Lempert domains.
-We now get to the special case of the $\ell^{q}$-‘ball’ for $q<1$ in arbitrary (finite) dimension, which is amongst the simplest examples of (pseudoconvex) domains which fail to be a Lempert domain (see proposition 5.8 below as well); indeed, we deal with large collections of domains that fail to be Lempert domains (however, it goes without saying that the case of Lempert domains is indeed of interest).
+We now get to the special case of the $\ell^{q}$-‘ball’ for $q<1$ in arbitrary (finite) dimension, which is amongst the simplest examples of (pseudoconvex) domains which fail to be a Lempert domain (see proposition 5.17 below as well); indeed, we deal with large collections of domains that fail to be Lempert domains.
 Theorem 1.8.
-Let $q$ be an arbitrary positive number less than one.
-Consider the $\ell^{q}$-‘ball’ given by $D_{q}=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;|z_{1}|^{q}+\ldots+|z_{N}|^{q}<1\}$ and the $r$-dimensional subspace $Y=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r}~{}\text{for}~{}r+1\leq i\leq N\}$ of $\mathbb{C}^{N}$, where for each fixed $i$, not all $a_{i,j}$’s are zero.
-Then the $r$-dimensional subspace $Y_{q}=D_{q}\cap Y$ is not a retract of $D_{q}$.
-In-fact, the only $r$-dimensional linear retracts of $D_{q}$ are those spanned by an $r$-element subset of the standard basis $\{e_{1},\ldots,e_{N}\}\subset\mathbb{C}^{N}$.
-Thus, the holomorphic retracts through the origin of $D_{q}$ are precisely those linear subspaces spanned by some subset of the standard basis vectors.
-Note that the boundaries of these domains are holomorphically extreme; indeed $\mathbb{C}$-extremal, as is the case for $\ell^{p}$-balls for $1\leq p<\infty$ as well, in which case Vesentini’s theorem reduces considerations from holomorphic retracts to linear ones.
-We also furnish examples (of balanced non-convex pseudoconvex domains) for the case where such extremality condition fails in the section where we detail the proof of the above theorem.
+Let $q=(q_{1},\ldots,q_{N})$, where $q_{j}$’s are arbitrary positive numbers, all less than one. Consider the $\ell^{q}$-‘ball’ given by
+(1.2) $D_{q}=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;|z_{1}|^{q_{1}}+\ldots+|z_{N}|^{q_{N}}<1\}.$
+The holomorphic retracts through the origin of $D_{q}$ are precisely the orthogonal projections of $D$ onto a span of a finite subset of the standard basis vectors of $\mathbb{C}^{N}$; thereby, the retracts are all copies of lower-dimensional $\ell^{q}$-balls.
+The techniques used to obtain this theorem can also be used to characterize retracts through the origin of some (bounded balanced pseudoconvex) non-decoupled domains as well (see corollary 5.11).
+Note that the boundaries of these domains are holomorphically extreme, in particular $\mathbb{C}$-extremal, as is the case for $\ell^{p}$-balls for $1\leq p<\infty$ as well; in such cases proposition 1.3 reduces considerations from holomorphic retracts to linear ones.
+We also include a couple of examples of certain balanced non-convex pseudoconvex domains to illustrate the case where extremality condition fails (in the same section 5 where we detail the proof of the above theorem).
 Indeed, we shall show in such examples that retracts of all possible dimension exists through each point of the domain other than the origin as well.
-That retracts cannot also be expected in any given direction for the non-convex case, is substantiated by $D_{q}$ as in the above theorem, owing to the spectacular failure even at the origin (of $1$-dimensional retracts for instance, along any direction other than the coordinate directions).
-Getting back to analysing a bit further, examples arising out of $\ell^{p}$-balls, let us remark a word or two, about concrete examples of domains satisfying the $\mathbb{C}$-extremality condition in his result before we pass out of it.
+That retracts cannot also be expected in any given direction for the non-convex case, is substantiated by $D_{q}$ as in the above theorem, owing to the spectacular failure even at the origin (of $1$-dimensional retracts, along any direction other than the coordinate directions).
+Getting back to analysing a bit further, examples arising out of $\ell^{p}$-balls (where $p\geq 1$), let us remark a word or two, about generating examples of domains satisfying the $\mathbb{C}$-extremality condition in Vesentini’s result, before we pass out of it.
 Firstly, while the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary; Vesentini’s theorem applies to pin-down the retracts through the center of such a ‘ball of intersection’.
-Finally and more importantly, to bring out the role of the extremality condition both in Vesentini’s theorem and allied results both in the foregoing as well as the sequel (such as theorem 1.15 involving $\mathbb{R}$-extremality), let us mention that the converse of Vesentini’s theorem is also true: if all retracts through the origin of a bounded balanced convex domain are linear, then the domain is $\mathbb{C}$-extremal.
-In-fact, it is enough to just have linearity of the one-dimensional retracts and furthermore, a general and local version can be formulated as follows.
+Now towards non-convex cases, Vesentini’s theorem cannot in general be modified easily so as to adapt to other classes of domains despite proposition 1.3; for instance, to deal with those analytic polyhedra whose boundaries are $\mathbb{C}$-extremal.
+Such domains can never be holomorphically extreme (so proposition 1.3 cannot be applied), as their boundaries are Levi flat on an open dense subset  [^6] Indeed, this is the maximum possible Levi flatness that can be expected in this context; for instance, while it is possible for a smoothly bounded domain $D$ to be holomorphically extreme all along $\partial D$, it is impossible for $\partial D$ to be Levi flat everywhere..
+However, we were able to attain the following result.
+Theorem 1.9.
+There do not exist any (non-trivial) retracts through the origin of (bounded) balanced $\mathbb{C}$-extremal analytic polyhedron $D$ in $\mathbb{C}^{N}$ along the direction of any of the points in the open faces of $\partial D$.
+To be more precise, suppose that $D$ can be described as
+$D=\{z\in U\;:\;|f_{k}(z)|<1,\text{ for all }1\leq k\leq m\},$
+for some holomorphic $f_{k}$’s in a neighbourhood $U$ of $\overline{D}$ (and that $D$ happens to be an open, connected, bounded, balanced subset of $\mathbb{C}^{N}$ with $\mathbb{C}$-extremal boundary).
+Suppose also that the system of defining functions $(f_{k})$ is minimal i.e., none of the functions $f_{k}$ can be removed without changing $D$.
+Let $p$ be a boundary point on any of the open faces of $\partial D$, i.e.
+$p\in\tilde{\sigma}_{j}$ for some $j$ with $1\leq j\leq m$, where
+$\tilde{\sigma}_{j}:=\{z\in\partial D:|f_{j}(z)|=1,~|f_{l}(z)|<1~\text{for}~1\leq l\leq m~\text{with}~l\neq j\}.$
+Then there does not exist any retract $Z$ of $D$ passing through the origin tangent to $p$ thereat i.e., with $p\in T_{0}Z$.
+We know that $T_{0}Z$ is itself a retract, indeed a linear retract.
+What the above theorem means for linear retracts of $D$ is that they cannot intersect the open faces $\tilde{\sigma}_{j}$.
+Thereby, every (non-trivial) linear retract is spanned by (at-most N1N-1) points from the lower dimensional ribs of the polyhedron.
+Consequently, owing to theorem 1.15 below, every (non-trivial holomorphic) retract through the origin is a biholomorphic copy of a lower dimensional linear slice of the same polyhedron spanned by vectors from its ribs.
+Existence of domains as in the above theorem in abundance will be shown in section 8.
+A complete determination of retracts (even through the origin) of such domains seems possible only when either $f_{j}$’s are explicitly given or when sufficient data about them is given.
+We shall do this exercise only for what is perhaps the simplest of such polyhedrons namely, $D_{h}:=\{(z,w)\in\mathbb{C}^{2}:|z^{2}-w^{2}|<1,\;|zw|<2\}$.
+It may be noted that this itself is interesting to start with, as its boundary is not entirely complex non-degenerate, (i.e.
+its ribs has singular points within it).
+The next main result of this paper, from which we deduce both the above theorem as well as theorem 1.8, is the following.
+Theorem 1.10.
+Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $C$ be the convex hull of $\overline{D}$.
+Let $L$ be any linear subspace which cannot be precisely spanned by vectors from $\partial D\cap\partial C$.
+Then $D_{L}:=D\cap L$ is not a linear retract of $D$.
+We now lay down the aforementioned application of theorem 1.8 and theorem 1.9 about retracts of $D$ to establish irreducibility, i.e., biholomorphic inequivalence with product domains.
+As a consequence of the foregoing pair of theorems, we see that a holomorphic retract $Z$ through the origin of an analytic polyhedron as in the above theorem 1.9 has the property that $T_{0}Z$ intersects $\partial D$ only in its closed nowhere dense subset comprising ribs (formed by intersections of at-least two of the faces) of $\partial D$.
+Thereby, such retracts are sparse: $D$ does not admit any retract along any open piece of directions at the origin.
+This leads to theorem 1.5; though this was already stated, the point here is to highlight that the proof is very similar for theorem 1.8 we well.
+Indeed, to minimize repetition, we prove both cases simultaneously in section 8.
+Corollary 1.11.
+Let $D$ be a non-convex decoupled egg domain $D$ as in (1.2) of theorem 1.8.
+Then $D$ is not biholomorphic to a product domain.
+If $A$ is a non-trivial analytic subset of $D$ not passing through the origin then the same conclusion holds for $D\setminus A$.
+Among others, the above corollary alongwith theorem 1.5 deals with three classes of domains: product domains, analytic polyhedra and ‘co-analytic’ domains  [^7] This is not standard terminology; however, this helps prevent repeating long phrases to describe a class of domains that will occur oftentimes in this article.
+We shall also refer to them as ‘analytic complements’ at times.
+i.e., domains of the form $D\setminus A$ wherein $A$ is a non-trivial analytic subset of $D$ (we only deal with domains of holomorphy $D$ assuming ${\dim A}>0$ needless to say).
+Members of all these classes, share the common feature that their boundaries have non-trivial analytic varieties in them.
+Corollary 1.11 is to the effect that nevertheless it is easy to get hold of mutually inequivalent members from these classes.
+In-fact, as we shall see along the way in section 8 (one of the major sections of this article), an analytic polyhedron is never equivalent to a co-analytic domain $D\setminus A$, whatever $D$ may be.
+As can be guessed, such inequivalences stem from the differences in the location of the analytic boundary pieces (for the former, these are not situated in the interior as is evidently the case with the latter).
+An analytic polyhedron can be equivalent to a product domain, the simplest example being $\Delta^{2}$; however, proposition 8.11 will confirm that this is the only bounded balanced domain in $\mathbb{C}^{2}$ for which such an equivalence occurs.
+Finally and more importantly, to bring out the role of the extremality condition both in Vesentini’s theorem and allied results (such as the foregoing theorem 1.10 and theorem 1.17 below), let us mention that the converse of Vesentini’s theorem is also true: if all retracts through the origin of a bounded balanced convex domain are linear, then the domain is $\mathbb{C}$-extremal.
+In-fact, it is enough to just have linearity of the one-dimensional retracts and furthermore, a general and local version can be formulated, namely as follows.
 Theorem 1.12.
@@
   • (i)
-    That the hypotheses that $\partial D$ is convex near $p$ implies convexity of $\partial D$ at $p$ together-with the continuity of the Minkowski functional near $p$; this will be discussed during the course of the proof.
-It may happen that $\partial D$ is convex at some boundary point but not in any open piece of the boundary about it; to see a simple example, consider $l^{q}$-‘balls’ for $q<1$ in $\mathbb{C}^{2}$ and examine points $p$ on the coordinate axes.
-This gives an example wherein convexity near $p$ fails but the weaker condition of convexity at $p$ together-with the continuity of the Minkowski functional near $p$ indeed holds.
+    We note that the hypotheses that $\partial D$ is convex near $p$ implies convexity of $\partial D$ at $p$ together-with the continuity of the Minkowski functional near $p$; this will be discussed during the course of the proof.
+When $\partial D$ is convex near $p$, it is enough to check for the existence of some holomorphic variety through $p$ rather than a linear one, in order to ascertain existence of non-linear retracts; this is due to the equivalence of $\mathbb{C}$-extremality with holomorphic extremality in the setting of the above theorem, which is discussed in proposition 2.37.
+It may happen that $\partial D$ is convex at some boundary point but not in any open piece of the boundary about it; to mention a simple example, we may consider $l^{q}$-‘balls’ for $q<1$ in $\mathbb{C}^{2}$ and examine points $p$ on the coordinate axes.
+This gives an example wherein convexity near $p$ fails but the weaker condition of convexity at $p$ together-with the continuity of the Minkowski functional near $p$ as in the statement of the above theorem, indeed holds.
 Finally here let us recall that if $p$ is a non-convex boundary point then there are no one-dimensional retracts through the origin in the direction of $p$.
   • (ii)
-    More importantly, a precise description of all the non-linear retracts through the origin in a bounded balanced pseudoconvex domain along the direction of a boundary point $p$ at which $\partial D$ is convex but fails to be $\mathbb{C}$-extremal – with no assumptions about the continuity of the Minkowski functional near $p$ – is attained in proposition 6.4.
-We record the essence of this in the following theorem.
+    More importantly, a precise description of all the one-dimensional non-linear retracts through the origin in a bounded balanced pseudoconvex domain along the direction of a boundary point $p$ at which $\partial D$ is convex but fails to be $\mathbb{C}$-extremal – with no assumptions about the continuity of the Minkowski functional near $p$ as in the statement of the above theorem – is attained in proposition 6.4.
+That leads to the following theorem about retracts of all dimensions.
 Theorem 1.14.
 Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
-Suppose that $\partial D$ is convex near $p\in\partial D$ but with the boundary not being $\mathbb{C}$-extremal at $p$.
-Let $Z$ be an arbitrary one-dimensional retract of $D$ through the origin along the direction of $p$.
-Then $Z$ can be realized as the graph of a holomorphic map on the linear retract ${\rm span}(p)\cap D$, taking values in a bounded balanced pseudoconvex domain within the face $F_{\mathbb{C}}(p,\partial D)$ where $F_{\mathbb{C}}(p,\partial D)$ is the complex face of $\overline{D}$ at the boundary point $p$.
-Next, we wish to go beyond the purview of Vesentini’s result.
-To do this by starting with the simplest step, we retain the assumption of balancedness; also, as the most natural class of domains in several complex variables are pseudoconvex domains, we ask if convexity in the above theorem of Vesentini can be relaxed to pseudoconvexity with no further assumptions about convexity as in the above theorem and extremality conditions dropped as well.
-Towards this, we have the following result wherein we assume boundedness.
-It must be noted that dropping boundedness leads to the failure of the result whereas replacing boundedness by the apparently more general condition of Kobayashi hyperbolicity is insignificant here, owing to the already mentioned result of Kodama in [Kodama], according to which any hyperbolic balanced domain is bounded.
+Let $Z$ be any (non-trivial) retract of $D$ through the origin with the property that $\partial D$ is convex near points where $L:=T_{0}Z$ intersects $\partial D$.
+Then $Z$ can be realized as the graph of a holomorphic map over the linear subspace $D_{L}$ (which itself is a retract of $D$), taking values in $\bigcup_{p\in\partial D_{L}}F_{\mathbb{C}}(p,\partial D)\cap D_{M}$ where $F_{\mathbb{C}}(p,\partial D)$ is the $\mathbb{C}$-linear face of $\overline{D}$ corresponding to the boundary point $p$ and $M$ is a certain linear subspace complementary to $L$ (and as usual $D_{M}:=D\cap M$).
+In proving this theorem, we first derive from an idea from [Abate_Isometry], a general result without any convexity hypothesis, thus going beyond the purview of Vesentini’s result.
+Namely, the following result wherein we assume boundedness.
+It must be noted that dropping boundedness leads to the failure 888see proposition 10.1 of section 10.
+of the result whereas replacing boundedness by the apparently more general condition of Kobayashi hyperbolicity is insignificant here, owing to the already mentioned result of Kodama in [Kodama], according to which any hyperbolic balanced domain is bounded.
 Theorem 1.15.
-Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\longrightarrow Z$, a retraction map.
+Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ any (non-trivial) retract passing through the origin with $\rho:D\longrightarrow Z$, a retraction map.
 Let $L$ denote the tangent space to $Z$ at the origin.
-Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$.
-Relegating some relevant observations here to section 3, we record here a corollary for an important class of domains.
-Indeed, when $D$ as in the above theorem (without any extremality condition about any boundary point), is homogeneous i.e., ${\rm Aut}(D)$ acts transitively on $D$, then of-course we may glean a characterization of all retracts of such a domain i.e., retracts not passing through the origin as well.
+Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$ taking values in $D_{M}:=\ker\left(D\rho(0)\right)$.
+It follows immediately from this theorem that $Z$ is biholomorphic to a (lower dimensional) the linear retract $D_{L}$.
+In the cases where $\partial D$ is holomorphically extreme, $Z=D_{L}$.
+Whatever be the case for $D$ and $Z$ as above, several inheritance properties extending the list mentioned earlier follow: if $D$ is $\mathbb{C}$-extremal analytic polyhedron or, strongly pseudoconvex or more generally (pseudoconvex and) of finite type, then $Z$ is biholomorphic to a domain of the same kind.
+While it is true that when $D$ is an $\ell^{p}$-ball, $Z$ is biholomorphic to a lower dimensional $\ell^{p}$-ball, this is nowhere close to being as trivial as the just mentioned inheritance properties.
+Next, we note that it does not follow that any two retracts of the same dimension through the origin, are biholomorphically equivalent; indeed, even a pair of linear retracts of the same dimension may fail to be biholomorphic.
+To mention a concrete example, we may consider $D=\mathbb{B}^{2}\times\Delta^{2}$ wherein both factors can be seen as two-dimensional retracts through the origin in $D$.
+Relegating further relevant observations here to section 3, we record here a corollary for an important class of domains.
+Indeed, when $D$ as in the above theorem (without any extremality condition about any boundary point), is homogeneous i.e., ${\rm Aut}(D)$ acts transitively on $D$, then of-course we may glean a characterization of all retracts of such a domain.
 Corollary 1.16.
@@
 Then, every non-trivial retract $Z$ (not necessarily passing through the origin in $D$) is biholomorphic via a member of ${\rm Aut}(D)$ to the graph of a holomorphic map over a linear retract $D_{L}$ (which as usual is a ball with respect to some norm in the lower dimensional complex Euclidean space $L=T_{0}Z$), of $D$.
-Owing to the assumption of being balanced, we recall that the domains as in this corollary are the much studied bounded symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV alongwith two exceptional cases, all of whose automorphism groups are well-known (Hua [Hua]).
-This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball – as in theorem 1.7 for the $l^{p}$-ball for $p<\infty$ and theorem 1.8 for the $p=\infty$ case much earlier above; and few other relevant results as in the already mentioned very recent work [GZ] which deals with among other things, Cartan domains of type IV – is in general quite difficult.
+Owing to the assumption of being balanced, we recall that the domains as in this corollary are same as the (much studied) bounded balanced symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV, all of whose automorphism groups are well-known (Hua [Hua]); this also means that domains as in the above theorem are convex.
+This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball – as in theorem 1.6 for the $l^{p}$-ball for $p<\infty$ and proposition 1.7 for the $p=\infty$ case much earlier above, and few other relevant results as in the already mentioned very recent work [GZ] which deals with among other things, Cartan domains of type IV – is in general quite difficult.
 Moreover, it must be remembered that an orthogonal projection of such $l^{p}$-balls onto a linear subspace in general is not a retract, except for the case $p=2$.
 As these are classical facts, the significance of linear retracts from the viewpoint of recent research work may need a remark; let us only mention that it was shown by Mok in his fairly recent work [Mok], that every linear totally geodesic submanifold through the origin of any bounded symmetric domain $\Omega$ with respect to a ${\rm Aut}(\Omega)$-invariant Kähler metric (for instance, the Bergman metric) is realizable as a retract just by orthogonal projection.
-As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem, wherein which we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.13.
-It must first be noted that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifestly seen in the simplest case of polydiscs.
-This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in the simplest of test cases such as domains decomposable into a product of balls (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc.
+As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem; wherein which, we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.15.
+Recall that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifestly seen in the simplest case of polydiscs.
+This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in the simplest of test cases such as domains decomposable into a product of balls with respect to some norms (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc.
 Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains and with a simpler proof following the recent work [BBMV-union].
-As indicated above, this theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like.
-We shall address this here, to the extent we can.
+As indicated above, this theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll]; it was shown therein that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like.
+We shall address this here, in the following next main result of this paper.
 Theorem 1.17.
-Let $B_{1},B_{2}$ be a pair of bounded balanced strictly convex domains possibly in different complex Euclidean spaces.
-Then, every non-trivial retract of $D=B_{1}\times B_{2}$ through the origin is a holomorphic graph over such a linear subspace which itself is a retract of $D$.
-Infact, every non-linear retract of $D$ is the graph in $D$, either of a $B_{2}$-valued holomorphic map over a (complex) linear subspace of $B_{1}$ or, of a $B_{1}$-valued holomorphic map over a linear subspace of $B_{2}$ which itself is a linear retract of the respective ball.
-More specifically and precisely, let $B_{1}\subset\mathbb{C}^{N_{1}}$ and $B_{2}\subset\mathbb{C}^{N_{2}}$ be as above i.e., bounded balanced strictly convex domains.
+  • (A)
+    Let $B_{1},B_{2}$ be a pair of bounded balanced domains of holomorphy with holomorphically extreme boundaries, possibly in different complex Euclidean spaces.
+Then, every non-trivial retract of $D=B_{1}\times B_{2}$ through the origin is the graph, either of a $B_{2}$-valued holomorphic map over a (complex) linear subspace of $B_{1}$ or, of a $B_{1}$-valued holomorphic map over a linear subspace of $B_{2}$.
+    More specifically and precisely, let $B_{1}\subset\mathbb{C}^{N_{1}}$ and $B_{2}\subset\mathbb{C}^{N_{2}}$ be as above i.e., not necessarily convex but pseudoconvex bounded balanced domains with holomorphically extreme boundaries.
 Let $D=B_{1}\times B_{2}$ and suppose without loss of generality that $\min\{N_{1},N_{2}\}=N_{1}$.
-Let $\mu_{1},\mu_{2}$ denote the Minkowski functionals of $B_{1},B_{2}$ respectively, whose convexity ensures that $\mu_{1},\mu_{2}$ are norms on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively.
+Let $\mu_{1},\mu_{2}$ denote the Minkowski functionals of $B_{1},B_{2}$ respectively.
 Let $Z$ be any non-trivial retract of $D$, $L=T_{0}Z$, the tangent space to $Z$ at the origin, $D_{L}=L\cap D$ and $\partial D_{L}=L\cap\partial D$.
@@
     • (i)
-    if $\partial D_{L}\subset\partial B_{1}\times\partial B_{2}$, then $Z$ is a complex linear subspace of $D$ which can be realized as the graph of a complex linear map over the linear subspace $\pi_{1}(D_{L})$ of $B_{1}$ (the ball whose dimension is $\min\{N_{1},N_{2}\}$):
+      if $\partial D_{L}\subset\partial B_{1}\times\partial B_{2}$, then $Z$ is a complex linear subspace of $D$ which can be realized as the graph of a complex linear map over the linear subspace $\pi_{1}(D_{L})$ of $B_{1}$:
       $Z=\{\big{(}w,\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}\;:\;w\in\pi_{1}(D_{L})\}$
-    where the map
+      where the mapping
       $w\longmapsto\big{(}\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}$
-    is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$, isometrically into $B_{2}$; $\pi_{1}(D_{L})$ itself is a linear retract of $B_{1}$,
+      maps $B_{1}^{L}:=\pi_{1}(D_{L})$ into $B_{2}$,
     • (ii)
-    if $\partial D_{L}\subset\partial B_{1}\times\overline{B}_{2}$, then $Z$ is the graph of a $B_{2}$-valued holomorphic map of a linear subspace $(B_{1})_{L_{1}}$ of $B_{1}$, with $(B_{1})_{L_{1}}$ being a linear retract of $B_{1}$,
+      if $\partial D_{L}\subset\partial B_{1}\times\overline{B}_{2}$, then $Z$ is the graph of a $B_{2}$-valued holomorphic map over a linear subspace $(B_{1})_{L_{1}}$ of $B_{1}$, given by $(B_{1})_{L_{1}}:=B_{1}\cap L_{1}$, where $L_{1}=\pi_{1}(L)$.
     • (iii)
-    if $\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$, then $Z$ is the graph of a $B_{1}$-valued holomorphic map of a linear subspace $(B_{2})_{L_{2}}$ of $B_{2}$, with $(B_{2})_{L_{2}}$ being a linear retract of $B_{2}$,
+      if $\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$, then $Z$ is the graph of a $B_{1}$-valued holomorphic map over a linear subspace $(B_{2})_{L_{2}}$ of $B_{2}$, given by $(B_{2})_{L_{2}}:=B_{2}\cap L_{2}$, where $L_{2}=\pi_{2}(L)$.
     • (iv)
       if $\partial D_{L}$ intersects both $\partial B_{1}\times B_{2}$ and $B_{1}\times\partial B_{2}$ – which are disjoint open pieces of the boundary of $D$, indeed connected components of $\partial D\setminus(\partial B_{1}\times\partial B_{2})$ – then $Z$ is a linear retract of $D$ of complex dimension $\geq 2$, which is neither contained in $B_{1}$ nor in $B_{2}$.
-We cannot trivially extend this theorem to the case of a product $D=B_{1}\times B_{2}\times\ldots\times B_{n}$ – where $B_{j}$’s are all bounded balanced convex domains each of whose boundaries satisfy the strict convexity condition – just by replacing one of the factors in the above theorem by a product of $(n-1)$ of these balls i.e., writing for instance $D=(B_{1}\times B_{2}\times\ldots\times B_{n-1})\times B_{n}$, appealing to the above theorem and using an inductive argument, owing to the fact that $B_{1}\times B_{2}\times\ldots\times B_{n-1}$ fails to satisfy the strict convexity condition stipulated in the above theorem as soon as $n>2$.
+  • (B)
+    Suppose now that $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ are bounded balanced domains of holomorphy (i.e., with the holomorphic extremal assumption on their boundaries dropped).
+Then statements (ii) and (iii) of (A) above, hold true.
+Statements (i) and (iv) hold true if $Z$ is replaced by $D_{L}$ (thus they hold when $Z$ is linear, in which case $Z=D_{L}$; if $Z$ is not linear, then $Z$ is the graph over $D_{L}$).
+As the first assertion about retracts through the origin is already contained in theorem 1.15, the real matter of here is the finer information in the statments (i) – (iv) of (A) which we note are not mutually exclusive; it may also be noted that (ii) and (iii) are actually exhaustive.
+We further note that it may very well happen that neither $\pi_{1}(Z)$ nor $\pi_{2}(Z)$ is a retract of the respective factors $B_{1},B_{2}$ (see the example in remark 3.20).
+Next, we note that we cannot trivially extend all of part-(A) to extended products $D=B_{1}\times B_{2}\times\ldots\times B_{N}$ in the following manner even in the special case wherein all the $B_{j}$’s to be bounded balanced convex domains each of whose boundaries satisfy the strict convexity condition.
+Namely: by replacing one of the factors in the above theorem by a product of (N1)(N-1) of these balls i.e., writing for instance $D=(B_{1}\times B_{2}\times\ldots\times B_{N-1})\times B_{N}$, appealing to the above theorem and using an inductive argument; this is due to the fact that $B_{1}\times B_{2}\times\ldots\times B_{N-1}$ fails to satisfy the extremality condition stipulated in part-(A) as soon as $N>2$.
 However, such an extension is indeed possible and not difficult either; provided only that we repeat the same arguments multiple times by projecting onto the individual factors $B_{j}$ to harness the strict convexity condition satisfied by each of these factor balls of the product $D$.
 As the complete formulation of the end result of this, turns out to be lengthy, we shall just lay it down in section 3.
-Here we record an application to the union problem for a nice class of polyballs which is interesting owing to it being a case with some degeneracy; this is based on the ideas of Fornaess – Sibony in their article [Fost].
+Here we record an application of part-(B) of the above theorem (in-fact, its essence namely theorem 1.15, suffices) to: the ‘union problem’ for a nice class of polyballs which is interesting owing to it being a case with some degeneracy; this is based on the ideas of Fornæss - Sibony in their article [Fost].
 Recall that the union problem comprises the following: let $M$ be a complex manifold which is the union of an increasing sequence of open subsets $M_{j}$ each of which is biholomorphic to a fixed domain $\Omega\subset\mathbb{C}^{N}$; describe $M$ in terms of $\Omega$.
 Theorem 1.18.
-Take $\Omega$ in the union problem to be any homogeneous bounded balanced domain.
-If $M$ is non-hyperbolic and the corank of its (infinitesimal) Kobayashi metric is one, then $M$ is biholomorphic to $\Omega_{L}\times\mathbb{C}$ where $\Omega_{L}$ is a linear retract of $\Omega$.
-Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is all the more significant.
+Take $\Omega$ in the union problem to be any homogeneous bounded balanced domain; as is well-known such a domain is a polyball.
+If $M$ is non-hyperbolic and the corank of its (infinitesimal) Kobayashi metric is one, then $M$ is biholomorphic to $\Omega_{L}\times\mathbb{C}$ where $\Omega_{L}$ is a linear retract of the polyball $\Omega$.
+Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is in general more difficult and significant.
 Towards this, we report the following result.
 Theorem 1.19.
-Let $D$ be a bounded balanced strictly convex domain in $\mathbb{C}^{N}$.
-Let $Z$ be a retract of $D$ (not necessarily passing through origin) and suppose at-least one of the retraction maps for it say $\rho$, is a submersion at the origin.
-Then $Z=\rho(D)$ is the graph of a holomorphic function over some linear subspace $D_{L}$ of $D$.
-Indeed, $D_{L}=L\cap D$ where $L=T_{p}Z$ and $p=\rho(0)$.
-Now, we have for the most part focussed attention on balanced domains which being contractible may be regarded as being topologically trivial.
-To turn next to topologically non-trivial cases, one simple case to begin with in this direction is that of analytic complements of domains with a Stein neighbourhood basis; specifically, see lemma 8.1.
-Next, to consider the simplest domain lacking a Stein neighbourhood basis then, we are led to considering the standard domain in this context namely the Hartogs triangle given by
+Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$, $\rho:D\to Z$ a (non-trivial) retraction map and $p$ the point in $Z$ which is the image of the origin under $\rho$; so, $Z$ is a retract not necessarily passing through the origin.
+Suppose that: (i) the slice $D_{L}:=D\cap L$ is a bounded convex domain in $L:=T_{p}Z$ with $\mathbb{C}$-extremal boundary and (ii) $D\rho_{|_{0}}$ is a projection which maps $D$ onto $D_{L}$, where $L=T_{p}Z$.
+Then $Z=\rho(D)$ is the graph of a holomorphic map over the linear subspace $D_{L}$ of $D$.
+Examples wherein all the hypotheses of the theorem are satisfied will be provided in the section where we lay down the proof.
+To move forward, we note that we have for the most part focussed attention on bounded balanced pseudoconvex domains which being contractible, may be regarded as being topologically trivial; moreover, they have other special properties such as Stein neighbourhood bases, and plurisubharmonic barriers all along their boundary (i.e., are hyperconvex) as soon as the Minkowski function is continuous.
+To turn next to topologically non-trivial cases failing to be hyperconvex, one simple case to begin with in this direction, is that of analytic complements of domains with a Stein neighbourhood basis; specifically, see the remarks following lemma 8.1.
+Next, to consider the simplest domain lacking a Stein neighbourhood basis and plurisubharmonic barriers throughout the boundary then, we are led to considering the standard domain in this context namely the Hartogs triangle given by
 $\Omega_{H}=\{(z,w)\in\mathbb{C}^{2}\;:\;|w|<|z|<1\}.$
 We also contend that this is perhaps the simplest topologically non-trivial case to consider because it is biholomorphic to the product $\Delta\times\left(\Delta\setminus\{0\}\right)$; indeed the map $\phi(z,w)=(w/z,z)$, effects the aforementioned biholomorphic correspondence.
-Thus it suffices to determine the retracts of $\Delta\times\Delta^{*}$, which we do in the following theorem which leads to a complete characterization the retracts of the Hartogs triangle as well as its ‘analytic complements’.
+Thus it suffices to determine the retracts of $\Delta\times\Delta^{*}$, which we do in the following theorem giving a complete characterization of all retracts of the Hartogs triangle as well as its ‘analytic complements’; this will be deduced as a consequence of a far more general lemma about retracts of ’co-analytic’ domains namely lemma 8.1.
 Theorem 1.20.
@@
 Finally, retracts of $\Omega_{H}^{A}:=\Omega_{H}\setminus A$ where $A$ is any analytic subset of $\Omega_{H}$ are precisely given by $Z\cap\Omega_{H}^{A}$ where $Z$ is a retract of $\Omega_{H}$.
-While retracts of $\Omega_{H}$ as in the above theorem, will be deduced as an immediate corollary of lemma 8.1 mentioned above, what does not follow so immediately are retracts of its analytic complements $\Omega_{H}^{A}$.
-This is detailed in section 9.
-Next, it may be noted that an overarching hypothesis that we have imposed on all domains considered thus far, is their boundedness or equivalently (in our context of balanced domains here), hyperbolicity.
+Next, it may be noted that an overarching hypothesis that we have imposed on all domains considered thus far, is their boundedness or, equivalently in our major context of balanced domains here, hyperbolicity.
 To deal with the simplest case of unbounded and non-hyperbolic domains then, we first consider a product domain in which one of the factors is still a bounded domain but the other unbounded.
-We thought it better to separately record the result for the simplest case of this kind, namely of the form ×$\mathbb{C}\times D$ where $D$ is a bounded planar domain; the characterization of its retracts is laid down in the following.
+We thought it better to separately record the result for the simplest case of this kind, namely of the form $\mathbb{C}\times D$ where $D$ is a bounded planar domain; the characterization of its retracts is laid down in the following.
 Theorem 1.21.
-If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of ×$\mathbb{C}\times D$ is either of the form ×$\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$; and, conversely every such graph is a retract.
-It must be noted that this also addresses the other implicit assumption in most of our previous considerations, namely that of topological triviality.
-Indeed, the above theorem allows for domains with a variety of fundamental groups; more such variety is facilitated by the next set of results.
+If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of $\mathbb{C}\times D$ is either of the form $\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$; and, conversely every such graph is a retract.
+The above theorem allows for domains with a variety of fundamental groups; more such variety is facilitated by the next set of results.
 Proposition 1.22.
@@
 where $Z$ is a retract of $D$ and $f$ is a holomorphic function from $Z$ to $\mathbb{C}\setminus A$, or of the form $X\times Z$ where again $Z$ is a retract of $D$.
-While this article is focussed on characterization of retracts rather than retraction mappings, we thought it noteworthy to record in passing, a complete characterization of all retraction maps on the most special domain among the domains in the above theorem namely, ×$\mathbb{C}\times\Delta$.
+While this article is focussed on characterization of retracts rather than retraction mappings, we thought it noteworthy to record in passing, a complete characterization of the retraction maps on the most special domain among the domains in the above theorem namely, $\mathbb{C}\times\Delta$.
 Theorem 1.23.
-Every non-trivial retraction map $F$ of ×$\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is precisely: either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
-While this article also focusses attention for the most part on domains in $\mathbb{C}^{N}$, we would like to interject here a result about retracts of manifolds, as it is a by-product of the essential ideas in the proof of the foregoing theorem 1.19 and more importantly, includes cases with far greater topological complexity.
+Every non-trivial retraction map $F$ of $\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is precisely: either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
+While this article also focusses attention for the most part on domains in $\mathbb{C}^{N}$, we would like to interject here a result about retracts of manifolds, as it is a by-product of the essential ideas in the proof of the foregoing theorem 1.21 and includes cases with far greater topological complexity.
+For instance, fundamental groups of planar domains are countably generated and in fact always freely generated, whereas for Riemann surfaces this is in general not the case.
 Theorem 1.24.
@@
 where $Z$ is a retract of $Y$ and $f$ is a holomorphic map from $Z$ to $X$ or, of the form $X\times Z$ where again $Z$ is a retract of $Y$.
+It must be noted that this theorem does not subsume proposition 1.22 above, for both factors of the product therein may very well be hyperbolic.
+We hasten to mention that there are indeed similar ideas in the proofs of the foregoing four results on products, which are therefore dealt with together in section 9 towards reducing repetition of details.
 Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{N}$ itself.
-While it is rather hard to precisely characterize all retracts of $\mathbb{C}^{N}$ and we shall show by actual examples that it is possible for non-hyperbolic domains to possess hyperbolic retracts, it is easy to observe a qualitative general feature common to all retracts of $\mathbb{C}^{N}$ apart from their topological triviality namely, that they are non-hyperbolic to the same (maximal) extent as $\mathbb{C}^{N}$ i.e., their Kobayashi metric vanishes identically.
+While it is rather hard to precisely characterize all retracts of $\mathbb{C}^{N}$, and we shall exemplify the possibility of non-hyperbolic domains possessing hyperbolic retracts, it is easy to observe a qualitative general feature common to all retracts of $\mathbb{C}^{N}$ apart from their topological triviality namely: they are non-hyperbolic to the same (maximal) extent as $\mathbb{C}^{N}$ i.e., their Kobayashi metric vanishes identically.
 It is also easy to see that the one-dimensional retracts of $\mathbb{C}^{N}$ are biholomorphic to $\mathbb{C}$.
@@
 Despite the apparent simplicity of the domain, this seems rather non-trivial; indeed, while graphs (in $\mathbb{C}^{2}$) of entire functions on $\mathbb{C}$ are retracts and the fact that retracts of products of planar domains in the foregoing, have literally turned out to be graphs of holomorphic maps over one of the factors, it is not the case that: every retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ is the graph of some entire function on either of its factors, as we shall show by providing a concrete example in the last section 10.
-On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this in essence is attained by Shpilrain and Yu in [Shpilrain--Yu] wherein it is deduced from other known results about polynomial rings and their retracts; more about this in the final section.
-Theorem 1.23 (Shpilrain – Yu).
+On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this in essence is attained by Shpilrain and Yu in [Shpilrain--Yu] wherein it is deduced from other known results about polynomial rings and their retracts, specifically, the Abhyankar – Moh theorem and a result due to Costa in [Cost].
+One may even wonder as to why is this result not a direct corollary of the Abhyankar – Moh theorem; we address this in the final section.
+We shall however give a self-contained and elementary proof not relying on such theorems, which does not seem to be recorded in the literature to the best of our knowledge; we believe such a proof will help clarify where exactly such theorems are actually necessary.
+We emphasize for one last time that this result is not new, and is due to various major contributors as clarified above.
+Theorem 1.25.
 If $Z$ is a non-trivial polynomial retract of $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\Phi$ on $\mathbb{C}^{2}$ such that $\Phi(Z)=\{(\zeta,0)\in\mathbb{C}^{2}\;:\;\zeta\in\mathbb{C}\}$.
 Such a ‘rectifiability’ assertion cannot in general be made for all holomorphic retracts of $\mathbb{C}^{N}$ for any $N\geq 2$, as follows by a simple analysis of the proof of the main result in the work [Forstn] due to Forstneric–Globevnik–Rosay.
-Acknowledgments: We thank Gautam Bharali, J.
+Acknowledgments: We thank Gautam Bharali, C.
+R.
+Jayanarayanan, Rohith Varma, J.
 Jaikrishnan and Kaushal Verma.
-The second named author acknowledges the support of the CSIR-JRF fellowship provided by the Council of Scientific and Industrial Research, India.
+The second named author would like to acknowledge the support of the CSIR-JRF fellowship provided by the Council of Scientific and Industrial Research, India.
 2. Preliminaries
@@
 While we would like to minimize repetition, we prefer to put the details on record to ensure a rigorous exposition.
-But here we repeat a few generalities from the introductory section for one last time, to cast this section systematically.
+We believe many results here (if not all) are probably folklore or even if new, are relatively straightforward but essential.
+However, we repeat here a few generalities from the introductory section for one last time, to cast this section and the rest, systematically.
 Let us begin with the most general definition relevant to us: a self-map $\rho$ of a set $X$ is called a retraction if $\rho\circ\rho=\rho$; note that this idempotent condition on $\rho$ means that $\rho$ acts as the identity map on its range $Z=\rho(X)$.
@@
 As the assertion is essentially ‘local’, we may fix an arbitrary point $p\in Z$ and work in a chart about $p$. In particular, consider the mapping on a chart neighbourhood $U_{p}$ of $p$, given by
-[…]
+$F_{p}(x):=\left(D\rho(p)\circ\rho\right)(x)\;:\;U_{p}\to T_{p}X=L\oplus E\;\text{ (say)}.$
 Denoting $D\rho(p)$ by $\pi_{p}$ we have
@@
 The idempotence of the linear operator $D\rho(p)$ given by the above lemma, guarantees its diagonalizability as well.
-So, after a linear change of coordinates, $D\rho(p)$ is a diagonal matrix with diagonal entries either $1$ or 00.
+So, after a linear change of coordinates, $D\rho(p)$ is a diagonal matrix with diagonal entries either $1$ or 0.
 Let $r$ be the number of $1$’s i.e., $r={\rm rank}(D\rho(p))$.
@@
 Now, observe that the same arguments may be applied to $D\beta(x)$ as well.
-Firstly, note that $D\beta(p)=I-D\rho(p)$ where $I$ is the identity matrix of size $n$, is also an idempotent matrix, indeed diagonal with diagonal entries only $1$ or 00 and with the number of zeros being exactly $n-r$.
-In particular, ${\rm rank}(D\beta(p))=n-r$ and the lower semicontinuity of the rank applied to $x\mapsto{\rm rank}(D\beta(x))$ implies that the rank of $D\beta(x)$ is atleast $n-r$ in a neighbourhood of $p$.
+Firstly, note that $D\beta(p)=I-D\rho(p)$ where $I$ is the identity matrix of size $n$, is also an idempotent matrix, indeed diagonal with diagonal entries only $1$ or 0 and with the number of zeros being exactly nrn-r.
+In particular, ${\rm rank}(D\beta(p))=n-r$ and the lower semicontinuity of the rank applied to $x\mapsto{\rm rank}(D\beta(x))$ implies that the rank of $D\beta(x)$ is atleast nrn-r in a neighbourhood of $p$.
 Now, we have a similar lower bound on the rank of $D\rho(x)$ as already noted, which is tempting to be translated this into an bound in-terms of $\beta$ instead of $\rho$; specifically, the upper bound ${\rm rank}(D\beta(x))\leq n-r$ in the some neighbourhood of $p$.
@@
 Retracts of Stein manifolds are Stein and likewise, the property of a complex manifold being an Oka manifold, is preserved under passage to a retract thereof, as well.
+In addition to the contrasts between retractions and proper mappings mentioned in the introduction, we have the following.
 Proposition 2.3.
@@
 Needless to say, similar constructions can be done with $l_{p}$-balls and more; we shall not digress further along this line.
+Since for a good part of the article, we deal with circular domains, we state one of the most famous results about mappings between such domains which will be used in numerous occasions in this article; we shall simply refer to this theorem due to H.Cartan as Cartan’s theorem.
+Theorem 2.6.
+(H.
+Cartan) Suppose $\Omega_{1},\Omega_{2}$ are circular domains in $\mathbb{C}^{N}$ with $0\in\Omega_{1}\cap\Omega_{2}$ and $\Omega_{1}$ is bounded.
+If $F$ is a biholomorphic map of $\Omega_{1}$ onto $\Omega_{2}$ with $F(0)=0$, then $F$ is a linear transformation.
+There shall also be occasions where another result of H.Cartan will be used, as stated next. We shall refer to it as the classical uniqueness theorem of Cartan.
+Theorem 2.7.
+Suppose $\Omega$ is a bounded domain in $\mathbb{C}^{N}$ and $F:\Omega\to\Omega$ is a holomorphic map. If for some $p\in\Omega,~F(p)=p$ and $F^{\prime}(p)=I$, then $F(z)=z$ for all $z\in\Omega$.
+2.1. Balanced Domains in $\mathbb{C}^{N}$
+In this subsection, we recall the basic properties of balanced domain and its associated Minkowski function, following [Jrncki_invrnt_dst].
+Definition 2.8.
+Let $D\subset\mathbb{C}^{N}$ be a balanced domain. Define its Minkowski function $h_{D}:\mathbb{C}^{N}\longrightarrow\mathbb{R}^{+}$,
+$h_{D}(z):=\inf\{t>0:z/t\in D\},\quad z\in\mathbb{C}^{N}.$
+Proposition 2.9.
+We have with the notations as in the above definition, the following properties of the Minkowski functional.
+  1. (a)
+    $h_{D}(\lambda z)=|\lambda|h_{D}(z)$, for all $z\in\mathbb{C}^{N},\lambda\in\mathbb{C}$.
+  2. (b)
+    $D=\left\{z\in\mathbb{C}^{N}:h_{D}(z)<1\right\}$.
+  3. (c)
+    If $h:\mathbb{C}^{N}\longrightarrow\mathbb{R}^{+}$is an upper semicontinuous function such that $h(\lambda z)=|\lambda|h(z),z\in\mathbb{C}^{N},\lambda\in\mathbb{C}$, then the set $D:=\left\{z\in\mathbb{C}^{N}:h(z)<1\right\}$ is a balanced domain and $h\equiv h_{D}$.
+  4. (d)
+    If $h_{D}$ is continuous, then $\partial D=\left\{z\in\mathbb{C}^{N}:h_{D}(z)=1\right\}$.
+  5. (e)
+    $D$ is convex if and only if $h_{D}$ is a seminorm, i.e., $h_{D}(z+w)\leq h_{D}(z)+h_{D}(w),z,w\in\mathbb{C}^{N}$.
+  6. (f)
+    If $D$ is bounded, then $h_{D}^{-1}(0)=\{0\}$.
+An important little observation that will be particularly useful to us is the following.
+Fix any norm $\|\cdot\|$ on $\mathbb{C}^{N}$.
+Then, for any $z\in\mathbb{C}^{N}$, the Minkowski functional of $D$ can be expressed as:
+$h_{D}(z)\;=\;\frac{\|z\|}{\|\pi(z)\|},$
+where $\pi(z)=\alpha_{0}z\in\partial D$ with $\alpha_{0}:=\sup\{\alpha>0:\alpha z\in D\}$.
 2.2. Linear Retracts
@@
 Proof.
-[…]
+It suffices to assume that $D_{L_{1}\oplus L_{2}}$ is a linear retract of $D$ for every pair of (complex) lines $L_{1},L_{2}$ through the origin $\mathbb{C}^{N}$ to derive a contradiction.
+Indeed with such an assumption, we would have that every $2$ dimensional subspace of $D$ is a linear retract of $D$.
+But then, this would imply by theorem-A in [Bhnblst] that $(\mathbb{C}^{N},\|\cdot\|)$ is a Hilbert space, where $\|\cdot\|$ denotes the Minkowski functional on $D$ which in turn implies the existence of a unitary map $T$ from $(\mathbb{C}^{N},\|\cdot\|)$ to $(\mathbb{C}^{N},\|\cdot\|_{\ell^{2}})$.
+Therefore $T(D)=\mathbb{B}$, contradicting the hypothesis that $D$ is not equivalent to $\mathbb{B}$.
+∎
 For $D$ as in the above theorem, we have in background the observation that the convexity of $D$, implies by the Hahn-Banach theorem that every one dimensional subspace of $D$ is a retract of $D$.
@@
 Definition 2.21.
-[…]
+Let $\Omega$ be a balanced bounded convex domain in $\mathbb{C}^{N}$; so, $\mathbb{C}^{N}$ may be regarded as a Banach space with respect to the norm $\|\cdot\|$ whose unit ball is $\Omega$.
+We say that $\Omega$ is strictly $K$-convex if for every $K$-linear functional $T$ with $\|T\|=1$, corresponds just one $z\in\overline{\Omega}$ such that $Tz=1$.
 In order to establish equivalence of $\mathbb{R}$-extremality with strict convexity, it is convenient to do this, the intermediate concept namely of rotundity, which is defined as follows:
 Definition 2.22.
-[…]
+Let $\Omega$ be a balanced bounded convex domain in $\mathbb{C}^{N}$; so, $\mathbb{C}^{N}$ may be regarded as a Banach space with respect to the norm $\|\cdot\|$ whose unit ball is $\Omega$.
+We say that $\Omega$ is rotund if $tx_{1}+(1-t)x_{2}\in\Omega$ whenever $x_{1}$ and $x_{2}$ are different points of $\partial\Omega$ and $0<t<1$.
 First we prove that the notion of rotundity and strict convexity are equivalent in the realm of bounded balanced convex domains.
@@
 Then $z,w\in\partial{\Omega}$, otherwise $T\left(z/\|z\|\right)=1/\|z\|>1$.
-Note that $T((1-t)z+tw)=1$.
+Note that T((1t)z+tw)=1T((1-t)z+tw)=1.
 Then $(1-t)z+tw\in\partial{\Omega}$ for all $t\in[0,1]$.
@@
 Proof.
-[…]
+Let $\Omega$ be a bounded balanced convex domain in $\mathbb{C}^{N}$.
+Note that $\Omega$ is the unit ball with respect to some norm $\|\cdot\|$ in $\mathbb{C}^{N}$.
+Assume $\Omega$ is strictly $\mathbb{R}$-convex.
+We want to prove that $\Omega$ is strictly $\mathbb{C}$-convex.
+Let $T$ be $\mathbb{C}$ -linear functional with $\|T\|=1$ such that $T(z)=T(w)=1$ for some $z,w\in\overline{\Omega}$.
+Every $\mathbb{C}$ -linear functional must be an $\mathbb{R}$-linear functional.
+Then $T$ is a $\mathbb{R}$-linear functional with $\|T\|=1$ and $T(z)=T(w)=1$.
+Since $\Omega$ is strictly $\mathbb{R}$-convex, we get $z=w$.
 Conversely, assume that $\Omega$ is strictly $\mathbb{C}$-convex.
@@
 The above equations tells that $|\lambda+t_{0}|=|\lambda|$. This contradicts our choice of $t_{0}$ and finishes the proof of the claim.
-[…]
+Claim-2: ${\rm dist}(a,M)=1$, where $\rm dist$} denotes the distance induced by the norm $\|\cdot\|$ with respect to which $\Omega$ is the unit ball.
 Proof of Claim-2: Let us begin by recalling that ${\rm dist}(a,M)=\inf\{\|a-su\|:s\in\mathbb{C}\}$ and note firstly (by considering $s=0$) that ${\rm dist}(a,M)\leq 1$.
@@
 Now by Claim-1, $a$ and $u$ are linearly independent.
-Since $span\{a\}\cap H=\{0\}$ and $span\{u\}\cap H=\{0\}$, it follows that $dim(H)<N-1$.
+Since $span\{a\}\cap H=\{0\}$ and $span\{u\}\cap H=\{0\}$, it follows that dim(H)<N1dim(H)<N-1.
 This contradicts the fact that $H$ is a hyperplane.
@@
 2.4. Geodesics and totally geodesic subsets
-We recall the notion of totally geodesic subsets of a general domain (or a complex manifold) $D$ from [Kosnski].
+We recall the notion of totally $K$-geodesic subsets of a general domain (or a complex manifold) $D$ from [Kosnski].
 Let $p_{1},p_{2}$ be (distinct) points in $D$; the pair $\delta=(p_{1},p_{2})$ shall be referred to as a ‘datum’.
-A Kobayashi extremal for $\delta$ is a holomorphic map $k:\Delta\to D$ such that both $p_{1}$ and $p_{2}$ are in the range of $k$, and the pseudo-hyperbolic distance between the preimages of $p_{1}$ and $p_{2}$ is minimized at $k$, over all holomorphic maps from $\Delta$ to $D$.
-We shall call the range of a Kobayashi extremal with datum $\delta$, a geodesic through $\delta$.
-If a set $V\subset D$ has the property that for any $p_{1},p_{2}\in V$, a geodesic through $(p_{1},p_{2})$ is contained in $V$, we shall say that $V$ is totally geodesic.
+A Kobayashi extremal for $\delta$ is a holomorphic map $k:\Delta\to D$ such that both $p_{1}$ and $p_{2}$ are in the range of $k$, and the pseudo-hyperbolic distance (also known as the Möbius distance on $\Delta$) between the preimages of $p_{1}$ and $p_{2}$ is minimized at $k$, over all holomorphic maps from $\Delta$ to $D$.
+We shall call the range of a Kobayashi extremal with datum $\delta$, a $K$-geodesic through $\delta$.
+If a set $V\subset D$ has the property that for any $p_{1},p_{2}\in V$, a $K$-geodesic through $(p_{1},p_{2})$ is contained in $V$, we shall say that $V$ is totally $K$-geodesic.
+A similar definition can be made with invariant metrics or distances, in particular the Caratheodory distance, but we do not pursue such matter here.
 The simple fact that the Kobayashi distance $k_{D}(p,q)$ between a pair of points $p,q$ in a domain $D$ remains invariant (i.e., $k_{D}(p,q)=k_{Z}(p,q)$) under passage to a retract $Z\subset D$, was recorded with proof explicitly as lemma $1$ in [HanPeters-Zeager].
@@
 Lemma 2.29.
-Let $X$ be any complex manifold and $Z\subset X$, a retract. For a contractive pseudodistance $d$ and contractive (infinitesimal) metric $\delta$, we have
+Let $X$ be any complex manifold and $Z\subset X$, a retract. For a contractive pseudo-distance $d$ and contractive (infinitesimal) pseudo-metric $\delta$, we have
 $d_{Z}(p,q)=d_{X}(p,q),\;\;\text{ and }\;\;\delta_{Z}(p,v)=\delta_{X}(p,v)$
 for every pair of points $p,q\in Z$ and $v\in T_{p}X$.
-It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic, a notion that shall be made use of, while recalling the work [Mok] of Mok.
-This lemma also shows that if $X$ is hyperbolic then every retract of $X$ is hyperbolic.
+It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic; the more standard definition of ‘totally geodesic’ shall also be made use of, while recalling the work [Mok] of Mok.
+Getting back to the above lemma, it also shows that if $X$ is hyperbolic then every retract of $X$ is hyperbolic.
 2.5. One dimensional Retracts
@@
 We say that a holomorphic map $\varphi:\Delta\rightarrow D$ is a complex $C$-geodesic of $D$ if for all $\zeta_{1},\zeta_{2}\in\Delta$, we have $c_{D}\left(\varphi\left(\zeta_{1}\right),\varphi\left(\zeta_{2}\right)\right)=p\left(\zeta_{1},\zeta_{2}\right)$.
+Occasionally, by abuse of language, we shall also refer to the image of $\varphi$ as a complex $C$-geodesic.
 Here $c_{D}$ denotes the Caratheodory distance; as we shall be working with a variant of this as well, we recall both of them and their notations:
@@
     there exists $\zeta_{1}\in\Delta$ and a nonzero vector $v\in\mathbb{C}$ such that $\gamma_{D}\left(\varphi\left(\zeta_{1}\right),\varphi^{\prime}\left(\zeta_{1}\right)\cdot v\right)=\gamma_{\Delta}\left(\zeta_{1},v\right)$.
-The following proposition relates complex geodesics with retracts.
+The following proposition relates complex $C$-geodesics with retracts.
+This seems well-known; it can be found stated in [Dineen].
+However, the proof details seem not to have been recorded anywhere which we therefore include here; particularly because it is a basic result about one-dimensional retracts.
 Proposition 2.32.
@@
 So $p\left((g\circ\varphi)(\zeta_{1}),(g\circ\varphi)(\zeta_{2})\right)=p(\zeta_{1},\zeta_{2})$ for all $\zeta_{1},\zeta_{2}\in\Delta$.
-By Schwarz Pick lemma, we have $g\circ\varphi\in Aut(\Delta)$.
+By Schwarz Pick lemma, we have $g\circ\varphi\in{\rm Aut}(\Delta)$.
 This implies that $\varphi$ has a holomorphic left inverse.
@@
 $\rho\circ\rho=(\varphi\circ\psi)\circ(\varphi\circ\psi)=\varphi\circ(\psi\circ\varphi)\circ\psi=\varphi\circ\psi=\rho.$
-This proves that $\rho$ is a holomorphic retraction map on $D$ and $\varphi(\Delta)$ is a one dimensional retract of $D$. ∎
+This proves that $\rho$ is a holomorphic retraction map on $D$ and $\varphi(\Delta)$ is a one-dimensional retract of $D$. ∎
+It may be surprising that the hypothesis has no condition about $D$.
+Indeed, it may happen that $c_{D}\equiv 0$, there are no complex $C$- geodesic and the result is vacuously true.
+The converse is true for the class of domains of much interest in this article namely, bounded balanced pseudoconvex domains.
+More generally, we have the following.
 Theorem 2.33.
-If $D$ is a simply connected Kobayashi hyperbolic domain in $\mathbb{C}^{N}$, then $1$-dimensional retracts of $D$ are precisely complex $C$-geodesics.
+If $D$ is a simply connected Kobayashi hyperbolic domain in $\mathbb{C}^{N}$, then the one-dimensional retracts of $D$ are precisely its complex $C$-geodesics.
 Proof.
@@
 Note that $Z$ is a non-compact simply connected Riemann surface.
-By the uniformization theorem, $Z$ is biholomorphic to one of the Riemann surfaces: unit disk $\Delta,~{}\hat{\mathbb{C}},~{}\mathbb{C}$.
+By the uniformization theorem, $Z$ is biholomorphic to one of the Riemann surfaces: unit disk $\Delta,~\hat{\mathbb{C}}$$\mathbb{C}$.
 Since $Z$ is non-compact and hyperbolic, $\hat{\mathbb{C}}$ and $\mathbb{C}$ are ruled out.
@@
 Let $\phi$ be a biholomorphic map from $\Delta$ to $Z$.
-By proposition 1.2 (the equivalence of parts a) and d) therein ) from [Dineen], we obtain that $\phi$ is a complex geodesic.
-So $Z$ is an image of complex $C$-geodesic.
+By proposition 1.2 (the equivalence of parts a) and d) therein ) from [Dineen] (proposition 11.1.7 in [Jrncki_invrnt_dst]), we obtain that $\phi$ is a complex $C$-geodesic.
+So $Z$ is an image of a complex $C$-geodesic.
 ∎
@@
 Proposition 2.36.
-Let $D$ be a balanced domain in $\mathbb{C}^{N}$. The above two definitions 2.30 and 2.31 are equivalent.
+Let $D$ be a balanced domain in $\mathbb{C}^{N}$. The above two definitions 2.34 and 2.35 are equivalent.
 Proof.
@@
 Then there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(a)|=h(a)$.
-Consider the real linear functional $:\mathbb{C}^{N}\to\mathbb{R}$ defined by $T(z)=Re(\alpha L(z))$, where $\alpha=|L(a)|/L(a)$.
+Consider the real linear functional $\mathbb{C}^{N}\to\mathbb{R}$ defined by $T(z)=Re(\alpha L(z))$, where $\alpha=|L(a)|/L(a)$.
 Note that $|\alpha|=1$ and we obtain that
@@
 Then there exists a $\mathbb{R}$-linear functional $f:\mathbb{C}^{N}\to\mathbb{R}$ with $|f(z)|\leq h(z)$ and $|f(a)|=h(a)$.
-Define a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ by $L(z)=f(z)-if(iz)$.
+Define a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ by L(z)=f(z)if(iz)L(z)=f(z)-if(iz).
 Observe that for each $z\in\mathbb{C}^{N}$, there exist $\lambda_{z}\in\mathbb{C}$ with $|\lambda_{z}|=1$ such that $\lambda_{z}L(z)=|L(z)|$.
@@
 $|L(z)|=|Re(\lambda_{z}L(z))|=|Re(L(\lambda_{z}z))|=|f(\lambda_{z}z)|\leq h(\lambda_{z}z)=|\lambda_{z}|h(z)=h(z)$
-Hence we get that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$.
+Hence, we get that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$.
 In-particular, $|L(a)|\leq h(a)$.
 But then the very definition of $L$ shows that $|L(a)|\geq h(a)$ and we thereby get that $|L(a)|=h(a)$ (and $ia$ lies in the kernel of $f$ though this observation is not necessary for the proof per se).
-Hence we have proved that $D$ is convex at $a/h(a)$ according to definition 2.30.
+Hence, we have proved that $D$ is convex at $a/h(a)$ according to definition 2.34.
+∎
+We also note that the notion of holomorphic extremality is equivalent to the notion of $\mathbb{C}$-extremality of the boundary in the presence of convexity, as clarified in the following proposition.
+Proposition 2.37.
+Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
+Assume that $D$ is convex near $p\in\partial D$.
+Then $D$ is $\mathbb{C}$-extremal at $p$ if and only if $D$ is holomorphic extreme at $p$.
+Proof.
+Since every holomorphic extreme boundary point of $D$ is $\mathbb{C}$- extremal, it suffices to show that if $D$ is $\mathbb{C}$-extremal at $p$ then $D$ is holomorphic extreme at $p$.
+Assume that $D$ is $\mathbb{C}$-extremal at $p$.
+Let $\text{co}(D)$ denote the convex hull of $D$.
+Firstly note that that $\text{co}(D)$ is a balanced convex domain.
+Next, we divide the proof into the following steps:
+Step-1: If $p\in\partial D$ and $D$ is convex at $p$ then $p\in\partial\text{co}(D)$.
+By proposition 2.3.1 (b) in [Jrncki_invrnt_dst], we get that $\gamma_{D}(0;p)=h(p)=1$.
+Applying proposition 2.3.1 (d), we obtain that $\hat{h}(p)=1$, where $\hat{h}$ is the Minkowski function of $\text{co}(D)$.
+This shows that $p\in\partial\text{co}(D)$.
+Step-2: $p$ is a $\mathbb{C}$-extremal boundary point of $\text{co}(D)$.
+To prove this by contradiction, assume that $p$ is not $\mathbb{C}$-extremal boundary point of $\text{co}(D)$.
+Then there exists non-zero $u\in\mathbb{C}^{N}$ such that $p+tu\in\text{co}(D)$ for all $t\in\Delta$.
+After applying proposition 2.3.1(a) in [Jrncki_invrnt_dst] we get that $\gamma_{D}(0,p+tu)=1$.
+Since $D$ is convex near $p$, we may assume that $D$ is convex at $(p+tu)/h(p+tu)$ for all $t\in\Delta$; the equivalence of (iv) and (v) in proposition 2.3.1 in [Jrncki_invrnt_dst] shows that $h(p+tu)=1$ for all $t\in\Delta$.
+This implies that $p+tu\in\partial D$ for all $t\in\Delta$, which contradicts the fact that $p$ is $\mathbb{C}$-extremal point of $D$.
+This finishes the proof of Step 2.
+Thereafter, applying lemma 11.3.7 in [Jrncki_invrnt_dst] to $\text{co}(D)$, we obtain that $p$ is a holomorphic extreme point of $\text{co}(D)$.
+Using arguments as in the proof of Step 2, we obtain that $p$ is a holomorphic extreme point of $D$.
+This finishes the proof of the proposition.
 ∎
@@
 Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is not convex at $a/h(a)$.
-Then there does not exist any one-dimensional retract of $D$ passing through 00 and $a$.
-Proof.
-Assume the contrary, there exists a one-dimensional retract $R$ of $D$ passing through 00 and $a$.
-By theorem 2.29, $R$ is image of some complex C-geodesic $\varphi:\Delta\to D$.
-But by lemma 11.1.2 in [Jrncki_invrnt_dst] (see also remark 11.1.3(b) therein), there is no complex $C$-geodesic in $D$.
-This leads to a contradiction.
-∎
+Then there does not exist any one-dimensional retract of $D$ passing through 0 and $a$.
+A proof of this will be provided later in the course of other things in the next section. We next get to a positive result.
 Proposition 2.39.
 Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is convex at $a/h(a)$.
-Then $\varphi:\Delta\to D$ defined by $\varphi(\lambda)=\lambda a/h(a)$ is a complex $C$-geodesic on $D$.
-In particular, there always exists a one dimensional linear retract of $D$ passing through 00 and $a$.
+Then there always exists a one dimensional linear retract of $D$ passing through 0 and $a$.
 Proof.
-Let $\lambda$ be any non-zero element in $\Delta$. Since $D$ is convex at $a/h(a)$ and by proposition 2.3.1(b) in [Jrncki_invrnt_dst], we get
-$c_{D}^{*}(\varphi(0),\varphi(\lambda))=c_{D}^{*}(0,\lambda a/h(a))=h(\lambda a/h(a))=|\lambda|$
-Note that $c_{D}=\tanh^{-1}c_{D}^{*}$. So
-$c_{D}(\varphi(0),\varphi(\lambda))=\tanh^{-1}\left(c_{D}^{*}(\varphi(0),\varphi(\lambda))\right)=\tanh^{-1}(|\lambda|)=p(0,\lambda)$
-By proposition 11.1.4 in [Jrncki_invrnt_dst], $\varphi$ is a complex $C$-geodesic. ∎
+Since $D$ is convex at $p$, there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(p)|=1$.
+By multiplying $L$ by a suitable unimodular constant, we may assume that $L(p)=1$.
+Consider the linear transformation $T:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $T(z)=L(z)p$.
+Note that $T$ maps $D$ into $D$ because
+$h(T(z))=h(L(z)p)=|L(z)|h(p)\leq h(z)<1.$
+Moreover, $T$ is a linear projection map:
+T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).
+This shows that $\text{span}(p)\cap D$ is a linear retract of $D$.
+Since $D$ is convex at $p$, there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(p)|=1$.
+By multiplying $L$ by a suitable unimodular constant, we may assume that $L(p)=1$.
+Consider the linear transformation $T:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $T(z)=L(z)p$.
+Note that $T$ maps $D$ into $D$ because
+$h(T(z))=h(L(z)p)=|L(z)|h(p)\leq h(z)<1.$
+Moreover, $T$ is a linear projection map:
+T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).
+This shows that $\text{span}(p)\cap D$ is a linear retract of $D$. ∎
+Remark 2.40.
+  1. (1)
+    In the case of bounded balanced convex domains, the geometric Hahn-Banach theorem ensures that every one-dimensional subspace of $D$ is a retract of $D$.
+  2. (2)
+    Every one dimensional retract $Z$ of $D$ passing through the origin and $a$ can be parametrized by $\lambda a/h(a)$ for $\lambda\in\Delta$.
 We spell out the following immediate consequence as it settles a first basic question about the very existence of any non-trivial retract in bounded balanced domains of holomorphy of central interest of this article.
@@
 ∎
-Combining theorem 2.29 and theorem 3.2 in [Vigue_Cnvx], we obtain the following result:
+Combining theorem 2.33 and theorem 3.2 in [Lemp_geodesc], the following result is obtained.
 Theorem 2.42.
@@
     Given $x\in D$ and $v\in T_{x}(D)$, there exists one dimensional retract $R$ of $D$ passing through $x$ with tangent vector $v$.
+Remark 2.43.
+Assume for this remark that $D$ is convex.
+The existence in abundance of one dimensional retracts of strictly convex domains (in $\mathbb{C}^{N}$) is well-known due to the work of Lempert.
+Another well-known work which obtains retracts of higher dimensions under certain conditions is [Ian_Grhm] of Graham; as noted therein, the condition in the above theorem that $D_{L}=D\cap L$ be a linear retract of $D$ which is equivalent to the existence of a norm one projection onto $L$ (denoted by $\pi_{L}$ in the proof), is always met if $L$ is one dimensional – this follows by the Hahn–Banach theorem which may be used to first get a linear functional $f$ on $(\mathbb{C}^{N},\mu)$ of norm one which takes the value $1$ at a boundary point of $D_{L}$ (one may view $f$ as mapping $D$ onto $D_{L}\subset L\simeq\mathbb{C}$); the desired projection is then obtained by taking the linear projection of $\mathbb{C}^{N}$ onto $L$ along $\ker(f)$.
+On the other hand (also noted therein), if there exists a norm one projection of the Banach space $(\mathbb{C}^{N},\mu)$ onto a every linear subspace, then $(\mathbb{C}^{N},\mu)$ is actually a Hilbert space i.e., $\mu$ is actually induced by an inner product owing to classical results mentioned in the introduction namely that of Kakutani in [Kakutni] for the real case and Bohnenblust in [Bhnblst] for the complex case.
+Thus, not every linear subspace intersected with $D$ need be a retract of $D$.
 2.6. Fixed point sets and Retracts
@@
 Assume that $f:D\to D$ be a holomorphic map which is not an automorphism and $\text{Fix}(f)$ is a non-empty subset of $D$ which consists of more than one point.
-Then there exists a non-trivial retract $Z$ of $D$ such that ${\rm Fix}(f)\subset Z$
+Then there exists a non-trivial retract $Z$ of $D$ such that ${\rm Fix}(f)\subset Z$.
 Proof.
@@
 Note that $\rm Fix(f)\subset\rm Z$.
-$\rm dim(Z)\neq 0$, because $Z$ is a connected complex manifold and #$\#(Fix(f))>1$.
+Note also that $\rm dim(Z)\neq 0$, because $Z$ is a connected complex manifold and #(Fix(f))>1\#(Fix(f))>1.
 By Corollary 2.1.30 in [Abate_Itrtn], $\rm dim(Z)$ is equal to the number of eigenvalues of $df_{z_{0}}$ contained in $\partial\Delta$.
@@
 This contradicts the hypothesis.
-Hence we can conclude that $\rm dim(Z)<N$.
+Hence, we can conclude that $\rm dim(Z)<N$.
 This proves that $Z$ is a non-trivial retract of $D$ and $\rm Fix(f)\subset Z$.
 ∎
-3. Proofs of theorems 1.3, 1.4, 1.8, 1.13, 1.15
-We would like to admit at the outset that the proof of theorem 1.15 on retracts of polyballs, closely follows the recent work [BBMV-union]; but we give here a thorough treatment even if, at the cost of repetition of essential ideas, for it allows for a better exposition of the proof of both this theorem, as well as perhaps more importantly: proofs of theorems prior to it as in the title of this section, which are fundamental to this article.
-Our goal here is to characterize the holomorphic retracts of the Cartesian product $B_{1}\times\ldots\times B_{N}$ (which will be referred to as polyballs for short), where the $B_{j}$’s are bounded balanced strictly convex domains or equivalently, the unit ball with respect to some norm in $\mathbb{C}^{n_{j}}$ for some $n_{j}\in\mathbb{N}$, for each $j=1,\ldots,N$; as noted in the previous section, this is equivalent to their boundaries being $\mathbb{R}$-extremal.
+2.7. Hyperconvex Domains
+In this subsection, we briefly recall a standard class of pseudoconvex domains containing convex domains, namely hyperconvex domains.
+Hyperconvexity is an intermediate notion between convexity and pseudoconvexity that is a generalization of a key concept (going by the name of barriers) characterizing solvability of the Dirichlet problem in dimension one.
+Owing to this, we recall the connection between hyperconvexity and psh (plurisubharmonic) barrier boundary points.
+To begin with, let us recall the definition of a psh barrier point from [barrier].
+Definition 2.45.
+A boundary point $a$ of an domain $D\subset\mathbb{C}^{N}$ is said to be a (global) psh barrier point for $D$ if there is a negative plurisubharmonic function $\varphi$ on $D$ with $\lim_{D\ni z\to a}\varphi(z)=0$.
+We call $a$ a local psh barrier point for $D$ if there is a (connected) open neighborhood $U$ of $a$ such that $a$ is a psh barrier point for $D\cap U$.
+We shall say that $D$ has psh barrier boundary (resp.
+local psh barrier boundary) if all of its boundary points are psh (resp.
+local psh) barrier points.
+Remark 2.46.
+Examples of domains whose boundary satisfying the local psh barrier condition, as in the definition, abound.
+All bounded strictly pseudoconvex domains, more generally, all bounded (weakly) pseudoconvex domains of finite type, serve as examples.
+More interestingly, non-examples are obtained by looking at domains whose boundaries contain one or more piece of analytic varieties.
+For instance, while all analytic polyhedra $D$ satisfy this condition; as soon as we remove any non-trivial analytic set $A$, $D\setminus A$ no longer satisfies this condition at any of the (boundary) points $a\in A$.
+Let us recall the definition of a hyperconvex manifold from [Hyp-cnvx].
+Definition 2.47.
+A complex manifold $M$ is said to be hyperconvex if there exists a bounded strictly plurisubharmonic (psh) exhaustion function $\varphi$ on $M$.
+Remark 2.48.
+  • (i)
+    If $F:M_{1}\to M_{2}$ is a proper holomorphic map between complex manifolds $M_{1}$ and $M_{2}$ with $M_{2}$ is hyperconvex, then any bounded strictly plurisubharmonic exhaustion function $\varphi$ on $M_{2}$ induces a bounded strictly plurisubharmonic exhaustion function on $M_{1}$ via the pullback $\psi:=\varphi\circ F$.
+Hence, $M_{1}$ is hyperconvex.
+In particular, this shows that hyperconvexity is a biholomorphic invariant.
+  • (ii)
+    Every hyperconvex domain in $\mathbb{C}^{N}$ is pseudoconvex (see theorem 5.2.1 and proposition 5.2.2 in [Kobyshi]).
+  • (iii)
+    In the case of bounded balanced pseudoconvex domains, hyperconvexity is equivalent to the continuity of their Minkowski functionals (see remark 3.2.3 a), c) in [Jrncki_invrnt_dst]).
+Proposition 2.49.
+Every retract of a hyperconvex complex manifold is hyperconvex.
+Proof.
+Let $Z$ be a retract of a hyperconvex complex manifold $X$ and let $\varphi:X\to[-\infty,c)$ be a bounded plurisubharmonic exhaustion on $X$.
+Note that $\varphi_{|_{Z}}$ serves as a bounded plurisubharmonic exhaustion function on $Z$.
+Hence, $Z$ is a hyperconvex complex manifold.
+∎
+The following theorem (see theorem 1.17.5 in [Jrnck_Frst_stps]) describes the relationship between hyperconvexity and the existence of psh barriers.
+Theorem 2.50.
+Let $D\subset\mathbb{C}^{N}$ be a domain.
+Then $D$ is hyperconvex if and only if for any boundary point $a\in\partial D$ (including $a=\infty$ when $D$ is unbounded), $a$ is a psh barrier point for $D$.
+Remark 2.51.
+If $D$ as in the above theorem is bounded, then by theorem 12.4.7 in [Jrncki_invrnt_dst], $D$ is hyperconvex iff every boundary point of $D$ admits a local psh barrier boundary.
+2.8. Growth Vectors and Tangent Cone
+In this subsection, we recall the definitions and related results concerning growth vectors and tangent cones from [Whitney].
+Definition 2.52.
+A set $C\subset\mathbb{C}^{m}$ is an open complex cone if $z\in C$ implies $\lambda z\in C$ for all $\lambda\in\mathbb{C}$.
+We may construct open complex cones by taking an arbitrary set $S\subset\mathbb{C}^{N}\setminus\{0\}$ and considering $C=C(S)=\{\lambda z|z\in S,~\lambda\in\Delta\}$.
+We then refer to $C$ as the cone generated by $S$.
+The associated punctured (complex) cone is given by $C^{*}=\{\lambda z|z\in S,~\lambda\in\Delta^{*}\}$.
+Note that if $S$ is convex, then $C(S)$ is a convex cone.
+Definition 2.53.
+Suppose $f$ is holomorphic near $a$, and ord $_{a}f\geqq v$. Then we may write
+(2.1) $f(x)=\sum_{|\lambda|=v}A_{\lambda}\left(x_{1}-a_{1}\right)^{\lambda_{1}}\cdots\left(x_{n}-a_{n}\right)^{\lambda_{n}}+\sum_{|\lambda|=v+1}F_{\lambda}(x)\left(x_{1}-a_{1}\right)^{\lambda_{1}}\cdots\left(x_{n}-a_{n}\right)^{\lambda_{n}}$
+where the $F_{\lambda}$ is holomorphic near $a$. The first sum on the right of (2.1) is the initial polynomial of $f$ at $a$; call it $f_{a}^{[*]}(x-a)$.
+As a consequence of (2.1), consider any vector $v\in\mathbb{C}^{n}$, use $x=a+zv,z\in\mathbb{C}$ for sufficiently small. Then we have the expansion
+$f(a+zv)=z^{v}\left[f_{a}^{[*]}(v)+zR(a,z,v)\right].$
+where $R$ is holomorphic for small $z$.
+Definition 2.54.
+Let $f$ be holomorphic and not identically zero near $a$, with $f(a)=0$.
+A non-zero vector $v$ is a growth vector for $f$ at $a$ if for the function $g(z):=f(a+zv)$ defined on sufficiently small disk $\Delta_{r}\subset\mathbb{C}$, we have $\operatorname{ord}_{0}g=\operatorname{ord}_{a}f$.
+Proposition 2.55.
+With $f$ as in the above definition 2.54, let $V_{f}$ be the set of non-growth vectors for $f$ at $a$.
+Then $V_{f}$ is the set of solutions of $f_{a}^{[*]}(v)=0$; hence it is an algebraic cone, and is closed and nowhere dense in $\mathbb{C}^{N}$.
+Remark 2.56.
+Note that $dim(V_{f})=2N-2$, because by above proposition, $V_{f}$ is the zero set of homogeneous polynomial $f_{a}^{[*]}$.
+Definition 2.57.
+Let $V$ be an analytic variety in $\mathbb{C}^{N}$.
+The tangent cone $C(V,p)$ is the set of all vectors $v$ with the property that for each $\varepsilon>0$ there is a point $q\in V$ and there is a complex number $\alpha\in\mathbb{C}$ such that
+$|q-p|<\varepsilon,\quad|\alpha(q-p)-v|<\varepsilon$
+Equivalently, $v\in C(V,p)$ if and only if there exist sequences $(q_{i})$ and $(\alpha_{i})$ such that
+$q_{i}\in V,\quad q_{i}\rightarrow p,\quad\alpha_{i}\left(q_{i}-p\right)\rightarrow v$
+Note that the tangent cone is a complex cone. Thus the direction of (non-zero) $v$ is the limit of directions of secants from $p$ to points of $V$ converging to $p$.
+The following theorem establishes the relationship between non-growth vectors of $f$ at $p$ and the tangent cone to the zero set of $f$ at $p$.
+Theorem 2.58.
+Let $f$ be holomorphic and not identically zero near $p$ with $Z$ as its zero set and $\operatorname{ord}_{p}f=m$. For a non-zero vector $v\in\mathbb{C}^{N}$, define the function
+$g_{v}(t)=f(p+tv),~t\in\Delta_{r},$
+where $\Delta_{r}$ is a disc of sufficiently small radius $r>0$. Then, $\operatorname{ord}_{0}g_{v}>m$ if and only if $v\in C(Z,p)$, the tangent cone to $Z$ at $p$.
+2.9. Analytic and Semi-analytic sets
+In this subsection, we shall recall the definitions and results regarding algebraic and analytic sets (varieties) over one of the fields $\mathbb{R}$ or $\mathbb{C}$), from [Singularity] and [Strat].
+Throughout this subsection, the term analytic refers to real analytic when $K=\mathbb{R}$, and complex analytic when $K=\mathbb{C}$.
+Let $M$ denote a connected $n$-dimensional (real or complex) analytic manifold and let $\mathcal{O}(M)$ denote the ring of analytic functions from $M$ to $K$.
+Definition 2.59.
+Let $A$ be an analytic subset of $M$.
+A point $p\in A$ is called smooth, of dimension $d$, if and only if there exists an open neighbourhood $W$ of $p$ in $M$ such that $W\cap A$ is an analytic submanifold of $W$ of dimension $d$ (over the field $K$ ).
+Thus, $p\in A$ is smooth of dimension $d$ if and only if there exists an open neighborhood $W$ of $p$ in $M$ and $f_{d+1},\ldots f_{n}\in\mathcal{O}(W)$ such that $W\cap A=Z\left(f_{d+1},\ldots,f_{n}\right)$ and, for all $x\in W,\nabla f_{d+1}(x),\ldots,\nabla f_{n}(x)$ are linearly independent.
+We denote the set of smooth points of $A$ by $\stackrel{{\scriptstyle\circ}}{{A}}$, and the set of smooth points of dimension $d$ by $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$.
+A smooth point of $A$ of highest dimension is called a regular point (this definition of regular is not followed by all authors); we denote the set of regular points of $A$ by $A^{\text{reg }}$.
+A point $p\in A$ which is not smooth is called a singular point (or, a singularity).
+We denote the set of singular points, the singular locus, of $A$ by $\Sigma A$.
+Theorem 2.60.
+Let $A$ be an analytic subset of $M$. Suppose that $0\leq d\leq n$. Then
+  1. (1)
+    $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$ is a d-dimensional analytic submanifold of $M$, and is an open subset of $A$;
+  2. (2)
+    the $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$ are disjoint for different $d,\stackrel{{\scriptstyle\circ}}{{A}}=\stackrel{{\scriptstyle\circ}}{{A}}^{(0)}\cup\ldots\stackrel{{\scriptstyle\circ}}{{A}}^{(n)},\stackrel{{\scriptstyle\circ}}{{A}}$ is an analytic submanifold of $M$, and $\stackrel{{\scriptstyle\circ}}{{A}}$ is open in $X$;
+  3. (3)
+    $\stackrel{{\scriptstyle\circ}}{{A}}$ is dense in $A$; and
+  4. (4)
+    $\Sigma A$ is a closed, nowhere dense subset of $A$.
+Definition 2.61.
+The dimension (over $K$), denoted $\operatorname{dim}A$, of an analytic set $A\subseteq M$ is the largest $d$ such that $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$ is non-empty.
+The codimension of $A$ is the number $\dim M-\dim A$.
+The dimension of $A$ at a point $p\in A$, denoted $\operatorname{dim}_{p}A$, is the largest $d$ such that $p$ is in the closure of $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$.
+Definition 2.62.
+A subset $V$ of $M$ is called a semivariety if locally at each point $x\in V$ it is a finite union of subsets defined by equations and inequalities
+$f_{1}=\cdots=f_{k}=0\quad\begin{cases}g_{1}\neq 0,\ldots,g_{l}\neq 0&\text{ (complex case) },\\ g_{1}>0,\ldots,g_{l}>0&\text{ (real case) },\end{cases}$
+where $f_{j}$ ’s and $g_{k}$ ’s are real (or complex) analytic (or algebraic) depending on the case under consideration.
+In the real algebraic case, semivarieties are commonly referred to as semialgebraic sets; in the complex algebraic case, they are known as constructible, and in either analytic case, they are called semianalytic sets.
+Definition 2.63.
+Let $V$ be a semianalytic subset of $M$. We then define smooth points, regular points and singular points exactly as in the above Definition 2.59 for analytic subsets, using the same notations.
+In the case of semivarieties, it is known that the smooth points of a semianalytic set are dense.
+Therefore, we define the dimension of $X$, and the dimension of $X$ at a point exactly as in the above Definition 2.61.
+3. Proofs of theorems 1.3, 1.7, 1.15, 1.17
+One of our main goals of this section is to characterize the holomorphic retracts of the Cartesian product $B_{1}\times\ldots\times B_{N}$ where the $B_{j}$’s are bounded balanced domains of holomorphy.
+The result seems to be new at-least to the extent of not being noted elsewhere; in-fact new even in the case when all the $B_{j}$’s are convex in which case the above product will be referred to as a polyball for short.
+However we hasten to admit at the outset here, that the proof of theorem 1.17 about the aforementioned goal, closely follows the recent work [BBMV-union]; with the new ingredient being theorem 3.6, an analogue of Mazet’s Schwarz lemma for the non-convex case, all discussed below.
+We give here a thorough treatment despite repetition of some of the essential ideas as it allows for a better exposition of the proof of both this theorem, as well as proofs of theorems prior to it as in the title of this section, which are fundamental to this article.
+Recall that theorem 1.17 has $2$ parts, wherein its part (A) requires the boundaries of all the factors of the product domain to satisfy an extremality condition.
+We discuss some examples for this now.
+Since the result is new even for the strictly convex case, we shall dwell upon this special case of interest first.
+Thus until further notice, suppose the $B_{j}$’s are bounded balanced strictly convex domains or equivalently, the unit ball with respect to some norm in $\mathbb{C}^{N_{j}}$ for some $N_{j}\in\mathbb{N}$, for each $j=1,\ldots,N$; as noted in the previous section, this is equivalent to their boundaries being $\mathbb{R}$-extremal and implies (owing to the temporary assumption of convexity) the holomorphic extremality condition required by 1.17 (A).
 As it turns out that the essential arguments for the case of a product of a pair of such balls extend to the more general case, we shall for simplicity of exposition, focus attention for the most part on the case of a product of $2$ such balls.
@@
 However, many nice properties of the disc may well not hold for such domains (i.e., bounded balanced convex domains).
-To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where []delimited-[][$\cdot$][ ⋅ ] denotes the greatest integer function.
+To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where $\cdot$] denotes the greatest integer function.
 They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extremality, in the following.
 Proposition 3.1.
-If $D_{1},D_{2}$ are any pair of domains in $\mathbb{C}^{N}$ with weakly (resp. strongly) $\mathbb{C}$-extremal boundary, then so is every connected component of $D:=D_{1}\cap D_{2}$.
+If $D_{1},D_{2}$ are any pair of (not necessarily convex) domains in $\mathbb{C}^{N}$ with weakly (resp.
+strongly) $\mathbb{C}$-extremal boundary, then so is every connected component of $D:=D_{1}\cap D_{2}$.
 Proof.
@@
 Without loss of generality assume that $\epsilon=1$.
-Consider a map $\alpha:\Delta\to\mathbb{C}$ defined by $\alpha(t)=a+tu$.
+Consider a map $\alpha:\Delta\to\mathbb{C}^{N}$ defined by $\alpha(t)=a+tu$.
 Since $u\neq 0$, $\alpha$ is a non-constant holomorphic map.
@@
 Let $a\in\partial(D_{1}\cap D_{2})$.
-If possible assume that there exists $\epsilon>0$ and $u\in\mathbb{C}^{N}$ such that $a+tu\in\partial(D_{1}\cap D_{2})$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$.
-Note that $\partial D_{1}\cap\partial D_{2}\subset\partial(D_{1}\cap D_{2})$.
-Hence $\overline{D_{1}\cap D_{2}}=\overline{D_{1}}\cap\overline{D_{2}}$.
+If possible assume that there exists $\epsilon>0$ and $u\in\mathbb{C}^{N}$ such that $a+tu\in\overline{D_{1}\cap D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$.
+Note that $\overline{D_{1}\cap D_{2}}=\overline{D_{1}}\cap\overline{D_{2}}$.
 If $a+tu\in\overline{D_{1}\cap D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$, then $a+tu\in\overline{D_{1}}\cap\overline{D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$.
@@
 In-particular, for instance, the intersection of the $l^{1}$-ball with a suitable dilate of another $l^{p}$-ball for $p\in(1,\infty)$ can be used to get hold of a concrete instance of a domain with $\mathbb{C}$-extremal (but not smooth) boundary.
-As an example which is not derived out of balls with respect to the isotropic $l^{p}$-norms, we may consider the ‘anisotropic’ ball (also called a complex ellipsoid) :=$\mathcal{E}:=\{(z,w)\in\mathbb{C}^{2}\;:\;|z|^{2}+|w|^{4}<1\}$.
+As an example which is not derived out of balls with respect to the $l^{p}$-norms (which are symmetric under a permutation of coordinates, despite not being strictly isotropic), we may for example, consider the ‘more’ anisotropic ball $\mathcal{E}:=\{(z,w)\in\mathbb{C}^{2}\;:\;|z|^{2}+|w|^{4}<1\}$ referred to as a ‘complex ellipsoid’ or an ‘egg’ domain .
 So, there is a good variety of bounded balanced convex domains with $\mathbb{R}$-extremal and $\mathbb{C}$-extremal boundaries.
@@
 Indeed even in the simplest case when $B_{1}=B_{2}=\Delta$, the only points on the boundary of $B_{1}\times B_{2}$ which are $\mathbb{R}$-extremal are the points of its distinguished boundary, which is a rather thin subset of the boundary.
-As mentioned at the start, our goal is to first describe the retracts of such a product $B_{1}\times B_{2}\subset\mathbb{C}^{N_{1}+N_{2}}$ which as already noted is a bounded balanced convex domain in $\mathbb{C}^{N_{1}+N_{2}}$, indeed the unit ball with respect to the norm $\parallel(w,z)\parallel=\max\{|w|_{\mu_{1}},|z|_{{\mu_{2}}}\}$ where $\mu_{1},\mu_{2}$ denote the norms in which $B_{1},B_{2}$ are the standard (open) unit balls.
-As mentioned in the introduction, this theme of retracts of polyballs was initiated earlier in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, what the retracts of polyballs look like; the purpose here is to address this as systematically as we can.
-Thus in particular, by theorem 2.1 all retracts of the domain of our interest here, namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, retracts of our convex domain $D$ are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^3] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract.
+As mentioned at the start, one of our goals is to first describe the retracts of such a product $B_{1}\times B_{2}\subset\mathbb{C}^{N_{1}+N_{2}}$, the unit ball with respect to the norm $\parallel(w,z)\parallel=\max\{|w|_{\mu_{1}},|z|_{{\mu_{2}}}\}$ where $\mu_{1},\mu_{2}$ denote the norms in which $B_{1},B_{2}$ are the standard (open) unit balls.
+Thus in particular, by theorem 2.1 all retracts of the domains as in theorem 1.17 namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, the retracts in the case $B_{1},B_{2}$ are convex, are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^9] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract.
 Infact, there are contractible spaces which do not admit a strong deformation retraction to any of its points, though obviously they admit a deformation retraction to each of its points.
@@
 By the boundary of such a linear subspace which we shall denote by $\partial D_{L}$, we mean the closed set $L\cap\partial D$.
-[…]
+Before delving into a detailed analysis of retracts through the origin of balanced domains $D$, one may like to see at this stage what can be said if at all, about such retracts of bounded balanced convex domains $D$ in general.
+Let us begin with the trivia that retractions $\rho$ associated to such retracts of $D$ fix the origin $\rho(0)=0$, and are self-maps of $D$; note also that $D$ being a bounded balanced convex domain, is the unit ball with respect to some norm $\left\|\cdot\right\|$ on $\mathbb{C}^{N}$, indeed its Minkowski functional.
+This already prompts us to apply an appropriate version of the Schwarz lemma on $\rho$.
+A befitting version for this context alongwith an elementary proof is provided by proposition C of the article [Simh] by Simha, by which we get $\left\|D\rho(0)\right\|\leq 1$.
+However, unlike the classical Schwarz lemma, equality herein need not always imply linearity of the map, to attain which, additional conditions need to be imposed.
+This brings us to the version of the Schwarz lemma due to Mazet mentioned at the outset, which we shall generalize below so as to make recurrent use of it in the sequel.
 Theorem 3.4.
@@
 If we drop the $\mathbb{R}$-extremal condition in the above theorem, then the conclusion about the linearity of $f$ may well fail, as shown by the simplest non-linear holomorphic map $f(z)=(z,z^{2})$ of the disc into the bidisc $\Delta^{2}$; indeed, note that $df(z)=(1,2z)$, so $df(0)=(1,0)$ thereby $\left\|df(0)\right\|_{l^{\infty}}=1$ verifying all other conditions except for the $\mathbb{R}$-extremal condition which we dropped.
-The following theorem is the generalization of above Mazet theorem, which is also useful in this sequel.
+The following extension also due to Mazet, which completely drops restrictions on the dimension of the source of the map is no less relevant in the sequel.
+We shall generalize this to some extent below for our needs.
 Theorem 3.5.
-(Theorem 3.7 of [Mazet_Rextrm]) Let $E$ and $F$ be Banach spaces and $B_{E},B_{F}$ be the corresponding unit ball. Consider a holomorphic map $f:B_{E}\to B_{F}$ be a holomorphic map. Then we have:
+(Theorem 3.7 of [Mazet_Rextrm]) Let $E,F$ be Banach spaces with their unit balls denoted $B_{E},B_{F}$ respectively. Let $f:B_{E}\to B_{F}$ be any holomorphic map. Then we have the following:
   1. i)
-    If there exists $u$ of norm 1 in $E$ such that $f^{\prime}(0)\cdot u$ be $\mathbb{R}$-extremal in $\overline{B_{F}}$ then $f(0)=0$.
+    If there exists $u$ of norm 1 in $E$ such that $f^{\prime}(0)\cdot u$ is $\mathbb{R}$-extremal in $\overline{B_{F}}$ then $f(0)=0$.
   2. ii)
-    If, for any $u$ of norm 1 in $E,f^{\prime}(0)\cdot u$ is $\mathbb{C}$-extremal in $B_{F}$ then $f$ is a polynomial of degree at most 2.
+    If, for every $u$ of norm $1$ in $E$, $f^{\prime}(0)\cdot u$ is $\mathbb{C}$-extremal in $B_{F}$ then $f$ is a polynomial of degree at most 2.
   3. iii)
-    If the two previous hypotheses are simultaneously verified, then $f$ is a linear map, $f(u)=$ $f^{\prime}(0)\cdot u$
-Though we shall continue making general observations for use in the sequel, we will once and for all now, specialize our ambient space from general complex manifolds or arbitrary Banach spaces to $\mathbb{C}^{N}$ equipped possibly with a norm different from the usual $l^{2}$-norm, so as to finish noting the implications of general facts for our specific case, without much digressions from this purpose.
-Getting to the retract $Z$ of our product domain $D=B_{1}\times B_{2}$, let $L=T_{0}Z$ which ofcourse is a complex linear subspace of $\mathbb{C}^{N_{1}+N_{2}}$; as we’ve already mentioned, we shall not be talking of trivial retracts and hence the possible dimensions of $L$ (over $\mathbb{C}$) are $11,2,\ldots,N_{1}+N_{2}-$ 1 which are also the possible dimensions of $Z$ as a complex submanifold; which of these dimensions are attained and what would be their forms then, are matters more subtle than the general features mentioned at the outset and as we shall see, actually depend on specific features of the factor domains $B_{1},B_{2}$.
+    If the two previous hypotheses are simultaneously satisfied, then $f$ is a linear map, $f(u)=$ $f^{\prime}(0)\cdot u$.
+An analogue of the above for the non-convex case can be derived from some standard ideas from chapter $11$ of [Jrncki_invrnt_dst], which we do now.
+This will quickly lead to a generalization of Vesentini’s theorem for the non-convex case, which will however be derived a bit later after theorem 3.10 where it is more relevant.
+Theorem 3.6.
+Let $D$ be a balanced pseudoconvex (not necessarily bounded) domain in $\mathbb{C}^{N}$ and $\varphi:\Delta\to D$ be a holomorphic map.
+If $\varphi(0)=0$ and $\varphi^{\prime}(0)$ happens to be a holomorphic extreme boundary point in $D$, then $\varphi$ is a linear map: $\varphi(t)=\varphi^{\prime}(0)t$.
+More generally, if $G$ is a bounded balanced pseudoconvex domain in $\mathbb{C}^{M}$ and $F:G\to D$ be a holomorphic map with the property that $F(0)=0$ and for all $u\in\partial G$, $DF_{|_{0}}(u)$ is a holomorphic extreme boundary point of $D$, then $F$ is linear.
+Proof.
+We start with the holomorphic map $\varphi:\Delta\to D$ as in the first part of the statement i.e., satisfying $\varphi(0)=0$ with $\varphi^{\prime}(0)=p$ being a holomorphic extreme boundary point of $D$.
+As $\varphi(0)=0$, we can write $\varphi(t)=t\tilde{\varphi}(t)$, where $\tilde{\varphi}\in\mathcal{O}(\Delta,\mathbb{C}^{N})$.
+Since $\varphi$ maps $\Delta$ into $D$, we have
+$h(\varphi(t))=|t|h\circ\tilde{\varphi}(t)<1.$
+Because the above inequality holds for all $t\in\Delta$, we obtain a similar inequality (but with the possibility of equality): $h\circ\tilde{\varphi}(t)\leq 1$ for all $t\in\Delta$. Note that
+$\tilde{\varphi}(0)=\varphi^{\prime}(0)=p.$
+This implies that $h\circ\tilde{\varphi}(0)=h(p)=1$.
+By the maximum principle applied to the subharmonic function $h\circ\tilde{\varphi}$, we get that $h\circ\tilde{\varphi}\equiv 1$.
+This shows that ${\rm image}(\tilde{\varphi})\subset\partial D$.
+By the hypothesis that $p\in\partial D$ is a holomorphic extreme boundary point of $D$, we obtain that $\tilde{\varphi}\equiv p$ and hence we conclude that $\varphi(t)=tp$ for all $t\in\Delta$.
+We shall now use the above to prove the general case expressed in-terms of the map $F$ as in the statement of the theorem.
+Let $p$ be an arbitrary point in $\partial G$.
+Consider the holomorphic map $\varphi:\Delta\to D$ defined by $\varphi(t)=F(tp)$.
+Note that $\varphi(0)=0$ and $\varphi^{\prime}(0)=DF_{|_{0}}(p)$ is a holomorphic extreme boundary point of $D$.
+By the special case $G=\Delta$ already dealt with, we obtain that $\varphi(t)=tDF_{|_{0}}(p)$.
+As $p$ was arbitrarily chosen, this means that $F(tp)=DF_{|_{0}}(tp)$ for all $p$ in the boundary of the bounded balanced domain $G$ (and all $t\in\Delta$).
+It follows thereby that $F(z)=DF_{|_{0}}(z)$, finishing the proof that $F$ is a linear map.
+∎
+We now get to the retract $Z$ of our product domain $D=B_{1}\times B_{2}$ and denote its tangent space at the origin $T_{0}Z$ by $L$, which ofcourse is a complex linear subspace of $\mathbb{C}^{N_{1}+N_{2}}$; as we have already mentioned, we shall not be talking of trivial retracts and hence the possible dimensions of $L$ (over $\mathbb{C}$) are $1,2,\ldots,N_{1}+N_{2}-1$ which are also the possible dimensions of $Z$ as a complex submanifold.
+Which of these dimensions are attained and what would be their forms then, are matters more subtle than the general features mentioned at the outset and as we shall see, actually depend on specific features of the factor domains $B_{1},B_{2}$.
 However, just the assumption of $\mathbb{R}$-extremality on the boundaries of $B_{1},B_{2}$ will enable us to say far more about $Z$ than just the already mentioned fact at the outset that it is a contractible submanifold of $D$.
@@
 Actually we can do better than doing this apriori verification that $\rho(D_{L})$ is a submanifold of $Z$ in each of the multiple cases that we shall be considering for the analysis of $Z$; while splitting up into multiple cases here would be seen to naturally arise owing to the difference in the forms of $Z$ that cases result in, the result $\rho(D_{L})=Z$ does not depend on any such case we may be in and can be shown to hold rather generally.
-As this also fits in well with our plan of first noting general facts following from from already known results, we now proceed to establish a general result about this as in theorem 3.9 below, which holds for any bounded balanced pseudoconvex domain $D$ – no $\mathbb{R}$-extremal hypotheses is required.
-While the proof of this will be done by borrowing an idea [Abate_Isometry], implementation of this idea requires preparatory work which itself yields another fundamental fact namely, that $D_{L}=D\cap L$ where $L=T_{0}Z$, itself is a retract of $D$, though not every linear subspace of $D$ can be its retract, unless $D$ is biholomorphic to the $l^{2}$-ball.
-Infact, $D_{L}$ is a linear retract of $D$ in the sense of definition 2.6.
-We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.7 and 2.8, allowing us to skip details here.
+As this also fits in well with our plan of first noting general facts following from already known results, we now proceed to establish a general result about this as in theorem 3.10 below, which holds for any bounded balanced pseudoconvex domain $D$ – no $\mathbb{R}$-extremal hypotheses is required.
+While the proof of this will be done by borrowing an idea from [Abate_Isometry], implementation of this idea requires preparatory work which itself yields another fundamental fact namely, that $D_{L}=D\cap L$ where $L=T_{0}Z$, itself is a retract of $D$, though not every linear subspace of $D$ can be its retract, unless $D$ is biholomorphic to the $l^{2}$-ball.
+Indeed, $D_{L}$ is a linear retract of $D$ in the sense of definition 2.10.
+We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.11 and 2.12, allowing us to skip details here.
 Lemma 3.7.
@@
 Then, the linear map $D\rho(0)$ is a projection which retracts $D$ onto $D_{L}$.
-As a corollary to this lemma, we not only get an alternative proof of corollary 2.33 – which was only for one-dimensional retracts – but also an improvement wherein there is no restriction on the dimension of the retract.
+As a corollary to this lemma, we give a proof of corollary 2.38 not because this proof avoids geodesics but because it brings out the reason directly.
 Corollary 3.9.
 Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$.
-If there exists a point $p\in\partial D$ such that $D$ is not convex at $p$, then there does not exist a retract of $D$ passing through the origin in the direction of $p$.
+If there exists a point $p\in\partial D$ such that $D$ is not convex at $p$, then there does not exist any one-dimensional retract of $D$ passing through the origin in the direction of $p$.
 Proof.
-Assume to obtain a contradiction that there exists a retract $Z$ of $D$ passing through the origin in the direction of $p$.
+Assume to obtain a contradiction that there exists a one-dimensional retract $Z$ of $D$ passing through the origin in the direction of $p$.
 Let $\rho$ be any one of the retraction maps on $D$ such that ${\rm image}(\rho)=Z$.
-By the previous lemma 3.7, the linear projection map $\pi:=D\rho|_{0}\;:\;\mathbb{C}^{N}\to\mathbb{C}^{N}$ maps $D$ onto $D_{L}$, where $L=T_{0}Z$.
-Consider the affine subspace $K:=p+Ker(\pi)$ of $\mathbb{C}^{N}$.
+By the previous lemma 3.8, the linear projection map $\pi:=D\rho_{|_{0}}\;:\;\mathbb{C}^{N}\to\mathbb{C}^{N}$ maps $D$ onto $D_{L}$, where $L=T_{0}Z$.
+Consider the affine subspace $K:=p+\ker(\pi)$ of $\mathbb{C}^{N}$.
+Note that $K$ is a complex hyperplane at $p$.
 Since $D$ is not convex at $p$, there exists a point $q\in K$ such that $q\in D\cap K$.
@@
 We are now fully prepared to prove the highlighted general result about the equality $\rho(D_{L})=Z$ whose statement is formulated with slightly different notations than the one that we had adopted for our specific domain $D$.
-This is indeed the content of theorem 1.13, whose statement we lay down again here for convenience, and then prove it inspired and based on an idea in [Abate_Isometry].
+This is indeed the content of theorem 1.15, whose statement we lay down again here for convenience, and then prove it based on an idea in [Abate_Isometry].
 Theorem 3.10.
@@
 It follows that $\rho$ restricted to $D_{L}$ and $\pi_{L}$ restricted to $Z$ are inverses of each other and being holomorphic, finishes the proof of the first statement.
-It only remains to observe that as the inverse of $\rho_{|_{D_{L}}}$ actually happens to be the linear projection $\pi_{L}$ onto the linear subspace $D_{L}$, this means that $Z=\rho(D_{L})$ is realized essentially as the graph of a single holomorphic function over $D_{L}$ (the only linear change of coordinates that may need to be made is to orthogonalize the projection $\pi_{L}$ which may be ‘oblique’ to render $Z$ as a standard graph over $D_{L}$).
+It only remains to observe that as the inverse of $\rho_{|_{D_{L}}}$ actually happens to be the linear projection $\pi_{L}$ onto the linear subspace $D_{L}$ along its kernel $M=\ker\left(D\rho(0)\right)$, this means that $Z=\rho(D_{L})$ is realized essentially as the graph of a (single) holomorphic map over $D_{L}$ taking values in $M=\ker\left(D\rho(0)\right)$.
 ∎
-While the above theorem allows retracts to be non-linear and we shall indeed furnish several examples of non-product (balanced, pseudoconvex) domains possessing non-linear retracts through origin in the form of graphs, we recall that such a phenomenon is precluded by Vesentini’s theorem in the convex case.
-Extremality boundary assumptions as in Vesentini’s theorem are indeed necessary as we shall see while establishing a converse to his theorem later.
-However, the hypothesis that the entire boundary be holomorphically extreme as in the extension of Vesetini’s theorem given by proposition 1.4 in [JJ] is inconvenient for discussing interesting intermediate examples in the sequel.
-This is addressed by proposition 1.4 which we now establish.
-Though this is in essence an imitation of the proof of the aforementioned proposition in [JJ], we prefer to include some details as this is a result about retracts that we shall need later in this article wherein the aforementioned relaxation will actually be used.
-And, also because it provides an answer towards the natural query as to whether an extension of Vesentini’s theorem holds if we add only an extremality hypothesis just as much as is needed about the boundary, but none about convexity in the setting of the above theorem 3.9.
-Proof of proposition 1.4: Choose any holomorphic retraction map $\rho:D\to D$ with $\rho(D)=Z$.
-We may observe by the above theorem 3.9, that the restriction $\rho|_{D_{L}}$ is a holomorphic map on $D_{L}$ whose image is precisely equal to $Z$; but we shall not need this for the rest of this proof here.
-Recall that $Z$ is a connected complex submanifold and the tangent space to $Z$ at the origin, is denoted by $L$; so, $L=T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}|_{D_{L}}(0)v=v\}$.
+While the above theorem allows retracts to be non-linear and we shall indeed furnish several examples of non-product (balanced, pseudoconvex) domains possessing non-linear retracts through origin in the form of graphs, we recall that such a phenomenon is precluded by Vesentini’s theorem in the convex case when the boundary is $\mathbb{C}$-extremal.
+That such extremal boundary conditions are indeed necessary will be shown in course of establishing a converse to Vesentini’s theorem later.
+However, the hypothesis that the entire boundary be holomorphically extreme as in the extension of Vesentini’s theorem given by proposition 1.4 in [JJ] is inconvenient for discussing interesting intermediate examples in the sequel.
+This is addressed by proposition 1.3 whose proof we now lay down.
+Though this can be done by an imitation of the proof of the aforementioned proposition in [JJ], we prefer to include some details as this is a result about retracts and further, an occassion will arise later in this article wherein the aforementioned relaxation will actually be used.
+It also provides an answer towards the natural query as to whether an extension of Vesentini’s theorem holds if we add only an extremality hypothesis just as much as is needed about the boundary, but none about convexity in the setting of the above theorem 3.10.
+Proof of proposition 1.3: This is one of the few occasions in this article wherein we discuss $2$ proofs, with a one sentence proof following this, to show a possible advantage of circumventing geodesics at times, as mentioned in the introduction.
+Choose any holomorphic retraction map $\rho:D\to D$ with $\rho(D)=Z$.
+We may observe by the above theorem 3.10, that the restriction $\rho_{|_{D_{L}}}$ is a holomorphic map on $D_{L}$ whose image is precisely equal to $Z$; but we shall not need this for the rest of this proof here.
+Recall that $Z$ is a connected complex submanifold and the tangent space to $Z$ at the origin, is denoted by $L$; so, $L=T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}_{|_{D_{L}}}(0)v=v\}$.
 Let $v\in D_{L},v\neq 0$.
@@
 Note that
-$K_{D}\left(\rho|_{D_{L}}\circ\phi(0);(\rho|_{D_{L}}\circ\phi)^{\prime}(0)\right)=K_{D}\left(0,v/h_{D_{L}}(v)\right)=K_{D}\left(\phi(0);\phi^{\prime}(0)\right).$
-Hence $\rho|_{D_{L}}\circ\phi$ is a $K_{D}$-geodesic for $(0,v/h_{D_{L}}(v))$.
+$K_{D}\left(\rho_{|_{D_{L}}}\circ\phi(0);(\rho_{|_{D_{L}}}\circ\phi)^{\prime}(0)\right)=K_{D}\left(0,v/h_{D_{L}}(v)\right)=K_{D}\left(\phi(0);\phi^{\prime}(0)\right).$
+Hence $\rho_{|_{D_{L}}}\circ\phi$ is a $K_{D}$-geodesic for $(0,v/h_{D_{L}}(v))$.
 Because of our hypothesis, we have that $v/h_{D_{L}}(v)$ is a holomorphic extreme point of $D$.
-By uniqueness of complex geodesics in $D$ (exposited as proposition 11.3.5 in [Jrncki_invrnt_dst]), it follows that $\rho|_{D_{L}}\circ\phi=\phi\circ\psi$, where $\psi\in{\rm Aut}(\Delta)$.
-Substituting $\psi^{-1}(h_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho|_{D_{L}}$.
+By uniqueness of complex geodesics in $D$ (exposited as proposition 11.3.5 in [Jrncki_invrnt_dst]), it follows that $\rho_{|_{D_{L}}}\circ\phi=\phi\circ\psi$, where $\psi\in{\rm Aut}(\Delta)$.
+Substituting $\psi^{-1}(h_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho_{|_{D_{L}}}$.
 Hence $D_{L}\subset Z$.
@@
 ∎
+An alternative short proof of this proposition is obtained simply by noting that it is a direct corollary of theorem 3.6; specifically, by applying the second part of that theorem to any one of the holomorphic retraction maps $\rho$ restricted to $D_{L}$.
+∎
 Some remarks about the foregoing pair of results are in order.
+Firstly, a remark about the last proof: note that the arguments up-till the conclusion that ${\rm image}(\tilde{\varphi})\subset\partial D$, goes through in the setting of above theorem (same as theorem 1.15), whose statement we may therefore refine as follows.
+If we call the largest complex analytic variety within $\partial D$ containing a point $p\in\partial D$ as the ‘holomorphic face’ of $\partial D$ at $p$, then we may say that the mapping (denote it by $F$ say), whose graph is the retract $Z$, as in theorem 1.15, takes values within the union of such holomorphic faces at boundary points $p$, for $p$ varying in $\partial D_{L}=\partial D\cap L$.
+As we already know that the range of the map $F$ is contained in $\ker(D\rho(0))$, the above helps refine this information about ${\rm range}(F)$.
+We thereby get another proof of proposition 1.3 as an immediate corollary of the foregoing theorem.
+This owes to the assumption that the holomorphic face on $\partial D_{L}$ becoming a singleton in the situation of that proposition.
 Remark 3.11.
-To indicate just a few concrete examples to illustrate where proposition 1.4 may be applied (but not the other results in this subject directly at-least), we first deal with domains where the holomorphic extremality condition is neither violated maximally nor minimally; examples for ‘maximal’ violation is also considered but in section 8.
-Among the simplest of domains satisfying the hypotheses of proposition 1.4 in this intermediate fashion, we consider the domain $\Omega_{1}:=\Delta^{2}\cap D_{q}$, where $D_{q}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<2^{q}\}$ for $10<q<$ 1 is a dilated copy of the $\ell_{q}$-ball.
-Note that $\Omega_{1}$ is a bounded balanced pseudoconvex domain and $\partial\Omega_{1}$ fails to be holomorphically extreme (in its entirety); though of-course, there exists an open piece of $\partial\Omega_{1}$, which is holomorphically extreme.
+To indicate just a few concrete examples to illustrate where proposition 1.3 may be applied (but not the other results in this context directly at-least), we first deal with domains where the holomorphic extremality condition is neither violated maximally nor minimally; examples for ‘maximal’ violation is also considered but in section 8.
+Among the simplest of domains satisfying the hypotheses of proposition 1.3 in this intermediate fashion, we consider the domain $\Omega_{1}:=\Delta^{2}\cap D_{q}$, where $D_{q}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<2^{q}\}$ for $0<q<1$ is a dilated copy of the $\ell_{q}$-ball.
+Note that $\Omega_{1}$ is a bounded balanced pseudoconvex domain and $\partial\Omega_{1}$ fails to be holomorphically extreme; though of-course, there exists an open piece of $\partial\Omega_{1}$, which is holomorphically extreme.
 A combined application of theorem 1.12, theorem 1.15 and corollary 2.38, proves that every retract passing through origin is either a linear subspace of $\Omega_{1}$ or a graph of a non-linear holomorphic function over some linear subspace of $\Omega_{1}$.
@@
 Further such examples of irreducible smoothly bounded balanced convex domains in any dimension $N\geq 3$, can be constructed by considering the domain
-:=$\mathcal{E}:=\{z\in\mathbb{C}^{N}\;:\;|z_{1}|^{p_{1}}+\ldots+|z_{N}|^{p_{N}}<1\},$
+$\mathcal{E}:=\{z\in\mathbb{C}^{N}\;:\;|z_{1}|^{p_{1}}+\ldots+|z_{N}|^{p_{N}}<1\},$
 where the exponents $p_{j}$’s are taken to be bigger than one and distinct (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$).
-Then, for every $1<k<N$, there exist a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent.
+Then, for every $1<k<N$, there exists a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent.
 This is easily seen for instance, by projecting $\mathcal{E}$ onto the span of different subsets of $k$-many of the standard basis vectors of $\mathbb{C}^{N}$ following the foregoing concrete examples.
-Remark 3.13.
-We shall assume until further notice that $D$ is convex.
-The existence in abundance of one dimensional retracts of strictly convex domains (in $\mathbb{C}^{N}$) is well-known due to the work of Lempert.
-Another well-known work which obtains retracts of higher dimensions under certain conditions is [Ian_Grhm] of Graham; as noted therein, the condition in the above theorem that $D_{L}=D\cap L$ be a linear retract of $D$ which is equivalent to the existence of a norm one projection onto $L$ (denoted by $\pi_{L}$ in the proof), is always met if $L$ is one dimensional – this follows by the Hahn–Banach theorem which may be used to first get a linear functional $f$ on $(\mathbb{C}^{N},\mu)$ of norm one which takes the value $1$ at a boundary point of $D_{L}$ (one may view $f$ as mapping $D$ onto $D_{L}\subset L\simeq\mathbb{C}$); the desired projection is then obtained by taking the linear projection of $\mathbb{C}^{N}$ onto $L$ along $\ker(f)$.
-On the other hand (also noted therein), if there exists a norm one projection of the Banach space $(\mathbb{C}^{N},\mu)$ onto a every linear subspace, then $(\mathbb{C}^{N},\mu)$ is actually a Hilbert space i.e., $\mu$ is actually induced by an inner product owing to classical results mentioned in the introduction namely that of Kakutani in [Kakutni] for the real case and Bohnenblust in [Bhnblst] for the complex case.
-Thus it must be noted that not every linear subspace intersected with $D$ need be a retract of $D$.
-But then again, we shall show in the next theorem that the retracts through the origin are all linear.
 The converse of above theorem 3.10 is also true.
@@
 Proof.
-Since $D_{L}$ is a linear retract of $D$, there exists a linear projection map $P$ such that $image(P)=D_{L}$.
-By idempotency of $P$, we can write $\mathbb{C}^{N}=L\oplus M$, where $M=\ker(P)$ and $L={\rm image}(I-P)$.
-Consider the map $\psi:D\to D$ defined by $\psi(z)=(T(z),\varphi(T(z)))$; this can be verified to be idempotent by a direct computation, thereby showing that $\psi$ is a holomorphic retraction from $D$ onto ${\rm Graph}(\varphi)$ over $D_{L}$, finishing the proof.
+Since $D_{L}$ is a linear retract of $D$, there exists a linear projection map $P$ such that $P(D)=D_{L}$.
+By idempotency of $P$, we can write $\mathbb{C}^{N}=L\oplus M$, where $M=\ker(P)$ and $L={\rm image}(P)$.
+Consider the map $\psi:D\to D$ defined by $\psi(z)=(P(z),\varphi(P(z)))$; this can be verified to be idempotent by a direct computation, thereby showing that $\psi$ is a holomorphic retraction from $D$ onto ${\rm Graph}(\varphi)$ over $D_{L}$, finishing the proof.
 ∎
@@
 However, the obstruction is substantial: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well.
-The proof of this, atleast as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true for every bounded balanced convex domain with $\mathbb{C}$-extremal boundary.
-This is of-course the result of Vesentini which we extend and give a more elementary proof as already mentioned in the introduction i.e., we now lay down the proof of theorem 1.3.
-Proof of theorem 1.3: Let $\rho:D\longrightarrow D$ be a retraction and $\rho(0)=0$.
-By lemma 3.7, $D\rho|_{0}$ is a projection of norm $1$ from $D$ onto $D_{L}=D\cap L$, where $L=T_{0}Z$.
-Note that $(D\rho|_{0})_{|_{D_{L}}}:D_{L}\longrightarrow D_{L}$ is the identity map on $D_{L}$ and that $D_{L}$ is a bounded balanced convex domain in $\mathbb{C}^{N}$.
-It follows from the proof of Krein-Milman theorem (the reader is referred to the book [Meggn] for the standard proof we have in mind here), that $D_{L}$ must have atleast one $\mathbb{R}$-extremal boundary point, say $z_{0}$.
-As $D\rho(0)$ acts as the identity map on $D_{L}$ (thereby in particular $D\rho(0)(z_{0})=z_{0}$), it follows that there is atleast one $\mathbb{R}$-extremal point in $D\rho(0)(D_{L})$ namely $z_{0}$ itself!
-To be precise $z_{0}$ is an $\mathbb{R}$-extremal point of $D$ not just $D_{L}$.
-Applying theorem 3.5 to the map $\rho_{|_{D_{L}}}$ viewed as a map from $D_{L}$ into the (larger) domain $D$, finishes the proof that $Z$ is linear.
+The proof of this, at least as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true in much greater generality as stated in proposition 1.3 .
 To not be amiss to discuss concrete examples and the simplest ones at that, recall that the first standard choices for $D$, namely the $l^{p}$-balls was already mentioned in the introduction, specifically theorem 1.6.
@@
 We now finish off the discussion of this $p=\infty$ by proving the characterization of its linear retracts as stated in proposition 1.7.
-[…]
+Proof of proposition 1.7: Assume $L$ is a linear retract of $\Delta^{N}$.
+We can write $L=Y\cap\Delta^{N}$, for some linear subspace $Y\subset\mathbb{C}^{N}$.
+By theorem-3 in [Heath_Suff], there exists $J=\{j_{1},j_{2},\ldots,j_{r}\}$ and holomorphic functions $F_{i}:M\to\Delta,~\ 1\leq i\leq N$ where $M=\{z\in\Delta^{N}:z_{j}=0~\text{if}~j\notin J\}$ such that $F_{i}(z)=z_{i}$ if $i\in J$ and
 $L=\Delta^{N}\cap Y=\{(F_{1}(z),\ldots,F_{N}(z)):z\in M\}.$
@@
 So, we get
-$F_{i}(z)=\begin{cases}z_{i},\quad\qquad\qquad\qquad\qquad\text{if}~{}i\in J\\ c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}},\quad\text{if}~{}i\notin J,\\ \end{cases}$
+[…]
 for some scalars $c_{i,j},~j\in J$.
 Fix $m\notin J$ and for $1\leq k\leq N$, define
-$w_{k}:=\begin{cases}\frac{|c_{m,j_{r}}|}{c_{m,j_{r}}},\ \text{if}\ k=j_{r}~{}\text{for some}~{}r\ \text{and}\ c_{m,j_{r}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}.$
+$w_{k}:=\begin{cases}\frac{|c_{m,j_{i}}|}{c_{m,j_{i}}},\ \text{if}\ k=j_{i}~\text{for some}~i\ \text{and}\ c_{m,j_{i}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}.$
 We know that $|F_{m}(z)|\leq 1$ for all $z\in\Delta^{N}$. If $w:=(w_{1},\ldots,w_{N})$ then
@@
 By definition of $F_{i}$, we get
-$F_{i}(e_{j_{k}})=\begin{cases}\delta_{i,j_{k}},\quad\qquad\text{if}~{}i\in J\\ c_{i,j_{k}},\qquad\text{if}~{}i\notin J,\\ \end{cases}$
+[…]
 where $\delta_{i,j}$ denotes the Kronecker delta function.
@@
 Define $J:=\{j_{1},\ldots,j_{r}\}$ and
-$F_{i}(z)=\begin{cases}z_{i},\quad\qquad\qquad\qquad\qquad\text{if}~{}i\in J\\ c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}},\quad\text{if}~{}i\notin J,\\ \end{cases}$
+[…]
 Note that $F_{i}(z)$ is a holomorphic function from $\Delta^{N}$ to $\Delta$, because by condition (ii) of the theorem.
@@
 As noted above, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors.
-The converse implication however is true and can now be derived as a corollary.
+The converse implication however, is true and can now be derived as a corollary.
 Corollary 3.15.
@@
 Without loss of generality assume that $\|v_{i}\|_{\infty}=1$ for $i=1,\dots,r$.
-There exists $J=\{j_{1},\ldots,j_{r}\}$ such that $\pi_{j_{i}}(v_{i})=1$, where $\pi_{i}$ denotes the projection to the $i^{th}$ component.
-We can write $v_{k}=e_{j_{k}}+w_{k}$.
+There exists $J=\{j_{1},\ldots,j_{r}\}$ such that $|\pi_{j_{i}}(v_{i})|=1$, where $\pi_{i}$ denotes the projection to the $i^{th}$ component.
+After multiplying an unimodular constant by $v_{i}$, we may assume that $\pi_{j_{i}}(v_{i})=1$.
+So, we can write $v_{k}=e_{j_{k}}+w_{k}$.
 Note that $w_{k}$’s are disjointly supported vectors.
 If $m\in\{1,2,\ldots,N\}\setminus J$, then we define $c_{m,{j_{k}}}:=\pi_{m}(w_{k})$.
-Since $\text{supp}(w_{i})$’s are disjoint, we get $\pi_{m}(w_{k})\pi_{m}(w_{l})=0$ for $k\neq l$.
+Since $\text{supp}(w_{i})$’s are disjoint, we get that $\pi_{m}(w_{k})\pi_{m}(w_{l})=0$ for $k\neq l$.
 Hence $\sum_{j\in J}|c_{m,j}|\leq 1$.
-By theorem 1.8, $Y=span\{v_{1},\ldots,v_{k}\}$ is a linear retract of $\Delta^{N}$.
+By proposition 1.7, $Y=span\{v_{1},\ldots,v_{k}\}$ is a linear retract of $\Delta^{N}$.
 ∎
-Turning now to the proof of theorem 1.15 and proceeding forth with it, as already remarked, the boundary of our product domain $D=B_{1}\times B_{2}$ does not enjoy the $\mathbb{R}$-extremality property.
-Nevertheless, we shall show that $Z$ is linear (indeed, $\rho(D_{L})=D_{L}=Z$) in several cases to be dealt with in the sequel and in all cases can be realized as holomorphic graphs over some subspace contained within the coordinate spaces $\mathbb{C}^{N_{1}}$ or $\mathbb{C}^{N_{2}}$ intercepted by $B_{1}$ or $B_{2}$ respectively.
-That is to say, while we already know that $\rho(D_{L})=Z$ is posited in $D$ as a holomorphic graph over $D_{L}$, what we shall be more precisely showing is that $Z$ can always be realized a holomorphic graph over some retract of $B_{1}$ or $B_{2}$ which in view of the last theorem is a linear subspace of the respective ball ($B_{1}$ or $B_{2}$).
-Indeed, we have finished our general observations and we shall now progress into the analysis of our specific $D$.
-First, a few words about notations about retracts $\rho$ of our specific product domain $D\subset\mathbb{C}^{N_{1}+N_{2}}$.
+Proposition 3.16.
+Let $W$ be a linear subspace of $\mathbb{C}^{N}$. Then $W\cap\Delta^{N}$ is a $k$-dimensional linear retract of $\Delta^{N}$ if and only if $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{\infty})$.
+Proof.
+One part of the implication will need tools that will be discussed only later below.
+Namely: if $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{\infty})$, then by proposition 5.4 below, $W\cap\Delta^{N}$ is a $k$-dimensional linear retract of $\Delta^{N}$.
+We shall therefore now deal with the converse here; to do this, assume that $W$ is a $k$-dimensional linear retract of $\Delta^{N}$.
+By proposition 1.7, there exists $J=\{j_{1},\ldots,j_{k}\}\subset\{1,2,\ldots,N\}$ such that $L$ is spanned by $v_{1},\ldots,v_{k}$, where each $v_{r}$ is of the form
+$v_{r}=\left(\sum_{i\notin J}c_{i,j_{r}}e_{i}\right)+e_{j_{r}}$
+and for each $m\in\{1,2,\ldots,N\}\setminus J$ satisfies $\sum_{j\in J}|c_{m,j}|\leq 1$.
+Without loss of generality assume that $J=\{j_{1},\ldots,j_{k}\}=\{1,2,\ldots,k\}$.
+Consider the linear map $S:W\to\mathbb{C}^{k}$ defined by $S(v_{i})=e_{i}$ for $i=1,2,\ldots,k$ and extended to other vectors of $W$ by linearity.
+Now we claim that $S$ is an isometry.
+To prove this claim, write each vector $v\in W$ as $v=\lambda_{1}v_{1}+\ldots+\lambda_{k}v_{k}$ for some $\lambda_{i}\in\mathbb{C}$.
+Note that $S(v)=(\lambda_{1},\ldots,\lambda_{k})$.
+Choose $i_{0}\in\{1,2,\ldots,k\}$ such that $\|(\lambda_{1},\ldots,\lambda_{k})\|_{\infty}=|\lambda_{i_{0}}|$.
+Hence $\|S(v)\|=|\lambda_{i_{0}}|$.
+Using the expression of $v_{r}$, we can expand $v$ as
+$v=\sum_{l=1}^{k}\sum_{i=k+1}^{N}\lambda_{l}(c_{i,l}e_{i}+e_{l})$
+Observe that the first $k$ components of $v$ are $\lambda_{1},\ldots,\lambda_{k}$ and for each $i=k+1~\text{to}~N$ , $i^{\text{th}}$ component of $v$ is $\sum_{l=1}^{k}\lambda_{l}c_{i,l}$.
+Note that for each $i=k+1$ to $N$, we have
+$\left|\sum_{l=1}^{k}\lambda_{l}c_{i,l}\right|\leq|\lambda_{i_{0}}|\left(\sum_{l=1}^{k}|c_{i,l}|\right)\leq|\lambda_{i_{0}}|$
+Hence we obtain that $\|S(v)\|_{\infty}=\|v\|_{\infty}$. This finishes the proof that $S$ is an isometry. ∎
+We turn now to the proof of theorem 1.17.
+First, a few words about notations about retractions $\rho$ specific to this theorem.
 When written out explicitly, $\rho$ has $N_{1}+N_{2}$ components: $(\rho_{1},\rho_{2},\ldots,\rho_{N_{1}},\rho_{N_{1}+1},\ldots,\rho_{N_{1}+N_{2}})$ where each of the $\rho_{j}^{\prime}s$ for $j=1,\ldots,N_{1}+N_{2}$ can apriori depend on all of the variables $w_{1},\ldots,w_{N_{1}},z_{1},\ldots,z_{N_{2}}$; so, we use $w=(w_{1},\ldots,w_{N_{1}})$ as our notation for the coordinates on $\mathbb{C}^{N_{1}}\supset B_{1}$ and $z=(z_{1},\ldots,z_{N_{2}})$ for that on $\mathbb{C}^{N_{2}}\supset B_{2}$.
+We denote the Minkowski functionals of $B_{1},B_{2}$ by $\mu_{1},\mu_{2}$ respectively.
+We spend some words for the relevant convex case i.e., the special case when $\mu_{1},\mu_{2}$ norms such that their unit balls are strictly convex – we refer to this as ‘the convex case’ for brevity, till the proof of theorem 1.17 is completed.
+This explicit spelling out is for the convenience of future use of this result in conjunction with results on norm one projections in Banach spaces from the functional analysis literature.
 Next, note that as our non-trivial retract $Z$ is necessarily a non-compact subset of $B_{1}\times B_{2}$, it will therefore meet the boundary of $D$ i.e., $\overline{Z}\cap\partial(B_{1}\times B_{2})\neq\emptyset$ where it may noted that
@@
 So $g_{1},g_{2}$ are holomorphic maps from the unit disc into the balls $B_{1},B_{2}$ respectively, mapping $0\in\Delta$ to the origin in the respective balls.
-We are thus essentially in the setting of the Schwarz lemma of Mazet recalled above as theorem 3.4, provided only that we verify that the derivatives of these maps at the origin satisfy the norm condition as in theorem 3.4.
+We are thus essentially in the setting of theorem 3.6 above (in-case $B_{1},B_{2}$ are strictly convex, the Schwarz lemma of Mazet recalled above as theorem 3.4 may also be applied), provided only that we verify that the derivatives of these maps at the origin satisfy the extremality condition as required.
 To do this, let us begin with the map $g_{1}$ whose derivative we compute using the chain-rule as:
 $dg_{1}(0)=d\pi_{1}\circ d\rho_{|_{\varphi(0)}}\circ d\varphi(0).$
-Then, using the facts that $d\pi_{1}\equiv\pi_{1}$ (as the projection $\pi_{1}$ is just linear), $\varphi$ maps $0\in\Delta$ to the origin in $B_{1}$ which inturn is fixed by $\rho$, and that $d\rho|_{0}$ is just the identity map on $T_{0}Z$ and that $(w^{0},z^{0})=d\varphi(0)\in T_{0}Z$, we get $dg_{1}(0)=w^{0}\in\partial B_{1}$, verifying the norm condition $\mu(w^{0})=1$ (and the $\mathbb{R}$-extremality of $dg_{1}(0)$, which is now trivial here owing to the hyporthesis that all boundary points of $B_{1}$ are $\mathbb{R}$-extremal).
-Hence, the aforementioned Schwarz lemma of Mazet, applied here to $g_{1}$ shows that that $\pi_{1}\circ\rho(tw_{0},tz^{0})=tw_{0}$, which means that $\rho^{1}:=\pi_{1}\circ\rho$ acts like the projection onto the $w$-axes on $D_{L}$:
+Then, using the facts that $d\pi_{1}\equiv\pi_{1}$ (as the projection $\pi_{1}$ is just linear), $\varphi$ maps $0\in\Delta$ to the origin in $B_{1}$ which inturn is fixed by $\rho$, and that $d\rho|_{0}$ is just the identity map on $T_{0}Z$ and that $(w^{0},z^{0})=d\varphi(0)\in T_{0}Z$, we get $dg_{1}(0)=w^{0}\in\partial B_{1}$ which is a holomorphically extreme boundary point by hypothesis.
+The aforementioned theorem applied to $g_{1}$ shows that that $\pi_{1}\circ\rho(tw_{0},tz^{0})=tw_{0}$, which means that $\rho^{1}:=\pi_{1}\circ\rho$ acts like the projection onto the $w$-axes on $D_{L}$:
 ${\rho^{1}}_{|_{D_{L}}}(w,z)=w.$
@@
 Finally, what the fact that the boundary of $Z=D_{L}$ (what we have denoted by $\partial D_{L}$), lies in $\partial B_{1}\times\partial B_{2}$ means for the functionals $\beta_{j}$’s appearing in the above description of $Z$, is that the linear map represented by their collective (whose graph is $Z$), namely
-$w\longmapsto\big{(}\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}$
-is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$.
+(3.4) $w\longmapsto\left(\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\right)$
+maps $B_{1}^{L}:=\pi_{1}(D_{L})$ into $B_{2}$.
+In this paragraph, we remark about the special case when $B_{1},B_{2}$ are both strictly convex.
+We may phrase the last observation as saying that the map (3.4) is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}$ which is the unit ball in $\pi_{1}(L)$ with respect to the norm $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$.
 Hence, we see that this places strong restrictions about when this case-(I) can happen.
 For instance, this case cannot occur if $N_{1}=N_{2}=N>1$ and $B_{1}=\Delta^{N}$ is the standard unit polydisc and $B_{2}=\mathbb{B}$, the standard ball with respect to the $l^{2}$-norm in $\mathbb{C}^{N}$ and ${\rm dim}(D_{L})=N$; or more generally, case-(I) cannot arise as soon as $\pi_{1}(D_{L})$ is not biholomorphic to any of the linear sections of $B_{2}$.
-But then again, one must not be amiss to note that case-(I) does happen atleast once regardless of what particular domains $B_{1},B_{2}$ are, namely when ${\rm dim}(D_{L})=1$; indeed, the complex line segment spanned by every point in $\partial B_{1}\times\partial B_{2}$ is a retract (as is more generally true of the $\mathbb{C}$-line segment spanned by any boundary point).
+But then again, one must not be amiss to note that case-(I) does happen at least once regardless of what particular balls $B_{1},B_{2}$ are, namely when ${\rm dim}(D_{L})=1$; indeed, the complex line segment spanned by every point in $\partial B_{1}\times\partial B_{2}$ is a retract (as is more generally true of the $\mathbb{C}$-line segment spanned by any boundary point of $D$, as we are dealing only with the convex case in this para).
 Case-(II):
-$\partial L\subset\partial B_{1}\times\overline{B}_{2}$.
+$\partial D_{L}\subset\partial B_{1}\times\overline{B}_{2}$.
 In this case, note that the projection $\pi_{1}:D_{L}\to B_{1}$ is proper and again the general fact that proper holomorphic maps cannot decrease dimension gives ${\rm dim}D_{L}\leq{N_{1}}$.
@@
 $D_{L}=\{\big{(}w,\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}\;:\;w\in\pi_{1}(D_{L})\},$
-for some $\mathbb{C}$-linear functionals $\beta_{1},\ldots,\beta_{N_{2}}$ on the complex linear subspace $\pi(L)\subset\mathbb{C}^{N_{1}}$; again, the fact that the boundary of $D_{L}$ lies in $B_{1}\times\partial B_{2}$ implies that the linear transformation given by
+for some $\mathbb{C}$-linear functionals $\beta_{1},\ldots,\beta_{N_{2}}$ on the complex linear subspace $\pi_{1}(L)\subset\mathbb{C}^{N_{1}}$; again, the fact that the boundary of $D_{L}$ lies in $\partial B_{1}\times B_{2}$ implies that the linear transformation given by
 $T_{\beta}(w)=\big{(}\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}$
-is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$.
-Picking a point $(w^{0},z^{0})\in\partial D_{L}$, considering the complex line segment joining the origin to this point and the restriction of the retraction of $\rho$ to this segment, we get by the arguments as in Case-(I) by applying the Schwarz lemma of Mazet, that ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ where $\rho^{1}=\pi_{1}\circ\rho$.
+maps $B_{1}^{L}:=\pi_{1}(D_{L})$ into $B_{2}$; this may be phrased in the convex case as: $T$ is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into ($\mathbb{C}^{N_{2}}$, $\mu_{2}$).
+The mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$.
+Picking a point $(w^{0},z^{0})\in\partial D_{L}$, considering the complex line segment joining the origin to this point and the restriction of the retraction of $\rho$ to this segment, we get by the arguments as in Case-(I) by applying theorem 3.6 (or in the convex case, the Schwarz lemma of Mazet), that ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ where $\rho^{1}=\pi_{1}\circ\rho$.
 As points of $D_{L}$ are of the form $(w,T_{\beta}(w))$ for $w\in\pi_{1}(D_{L})$, we have for all $w\in\pi_{1}(D_{L})$ that
@@
 $w\to\big{(}\rho_{N_{1}+1}(w,T_{\beta}(w)),\ldots,\rho_{N_{1}+N_{2}}(w,T_{\beta}(w))\big{)}.$
-for $w$ varying through $B_{1}^{L}$, the unit ball of $\pi_{1}(L)$ with respect to the norm $\mu_{1}$.
-Recalling that theorem 3.9 gives $Z=\rho(D_{L})$ whose form we just determined, finishes this case.
+for $w$ varying through $B_{1}^{L}$, the unit ‘ball’ of $\pi_{1}(L)$ with respect to the functional $\mu_{1}$.
+Recalling that theorem 3.10 gives $Z=\rho(D_{L})$ whose form we just determined, finishes this case, yielding that $Z$ is a graph of a $B_{2}$-valued holomorphic map on the linear subspace $B_{1}^{L}$, which we also denote by $(B_{1})_{L_{1}}$.
 Case-(III)
 $\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$.
-In this case, the projection onto the second-factor namely, $\pi_{2}:D_{L}\to B_{2}$ is proper and we deduce as usual at the outset, that the general fact that proper holomorphic maps do not decrease dimension implies that ${\rm dim}D_{L}\leq N_{2}$.
+In this case, the projection onto the second-factor namely, $\pi_{2}:D_{L}\to B_{2}$ is proper and we deduce as usual at the outset, that: the general fact that proper holomorphic maps do not decrease dimension implies that ${\rm dim}D_{L}\leq N_{2}$.
 We then start as in Case-(I) by picking a boundary point $(w^{0},z^{0})\in\partial D_{L}$, which we may assume to not be from the part of $D_{L}$ that lies (if it does at all) in $\partial B_{1}\times\partial B_{2}$ for this has already been dealt with in case-(I); the arguments that follow do not depend on such an assumption and we therefore do not make any assumption about which part of the boundary of $D_{L}$ the point $(w^{0},z^{0})$ comes from.
@@
 Indeed, compose the standard parametrization of $l$ namely $\varphi(t)=t(w^{0},z^{0})$ with $\rho$ followed by $\pi_{2}$ to get the holomorphic curve in $B_{2}$ given by the parametrization $g_{2}(t):=(\pi_{2}\circ\rho\circ\varphi)(t)$ for $t\in\Delta$.
-Noting that $g_{2}$ sends $0\in\Delta$ to the origin in $B_{2}$, computing by the chain-rule that $dg_{2}(0)=z^{0}\in\partial B_{2}$ and recalling that all boundary points of $B_{2}$ are $\mathbb{R}$-extremal, we see that all conditions required by theorem 3.4 are met, which when applied to this map $g_{2}$ yields that $g_{2}$ is a linear map: $g_{2}(t)=z^{0}t=tz^{0}$.
+Noting that $g_{2}$ sends $0\in\Delta$ to the origin in $B_{2}$, computing by the chain-rule that $dg_{2}(0)=z^{0}\in\partial B_{2}$ and recalling that all boundary points of $B_{2}$ are holomorphically extreme, we see that all conditions required by theorem 3.6 (or if one so wishes, the Mazet Schwarz lemma for the convex case) are met; which when applied to $g_{2}$ yields that this map is a linear map: $g_{2}(t)=z^{0}t=tz^{0}$.
 To see the implications of this, let us write this more elaborately using the original definition of $g_{2}$ as $\pi_{2}(\rho\circ\varphi(t))=tz^{0}$, which unravels explicitly as
@@
 As this verifies in particular that $\rho(D_{L})$ is a submanifold of $Z$, we may now apply the identity principle for analytic sets or just recall theorem 3.10, which allows us to say that the whole of $Z$ is obtained by applying $\rho$ just once to $D_{L}$ i.e., $\rho(D_{L})=Z$; as we just deciphered the form of $\rho(D_{L})$, this finishes determining $Z$.
-To conclude this case, note that equation (3.6) gives $(\pi_{2}\circ\rho)(0,z)=z$ for all $z\in\pi_{2}(D_{L})$.
-Moreover, $(\pi_{2}\circ\rho)(D)=\pi_{2}(Z)=\pi_{2}(D_{L})$, since $Z$ is a graph over $\pi_{2}(D_{L})$ i.e., $\pi_{2}\circ\rho$ is a retraction of $D$ with $\pi_{2}(D_{L})$ being the corresponding retract.
-Restricting $\pi_{2}\circ\rho$ to $B_{2}$ also shows that $\pi_{2}(D_{L})$ is a retract of $B_{2}$ itself.
-Similar arguments apply to the previous and forthcoming cases as well, which we skip for some conciseness.
 Remark 3.17.
@@
 As there are lines in $D_{L}$ which intersect $\Gamma_{1}$ as well as (other) lines in $D_{L}$ which intersect $\Gamma_{2}$, we have both the conclusions of these cases about the major pair of components of $\rho$ namely, $\rho^{j}:=\pi_{j}\circ\rho$ for $j=1,2$, which are that $\rho^{1}$ just returns the first set of variables on $D_{L}$: ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ and likewise for the second component ${\rho^{2}}_{|_{D_{L}}}(w,z)=z$.
-Putting together, in short we see that $\rho$ acts just as the identity map, first atleast on $U_{L}$, and thereby on $L$ as well, by the identity principle applied to both the holomorphic maps $\rho^{1},\rho^{2}$.
+Putting together, we therefore see that $\rho$ acts just as the identity map, first atleast on $U_{L}$, and thereby on $L$ as well, by the identity principle applied to both the holomorphic maps $\rho^{1},\rho^{2}$.
 As ${\rm dim}(D_{L})={\rm dim}(Z)$ obviously, this means that $D_{L}=Z$.
@@
 However, we note that though this theorem seems to provide a complete characterization of the retracts of $D$, we cannot pinpoint which of the linear subspaces of $D$ can happen to be its retract much as we cannot say which linear subspaces of $B_{1}$ or $B_{2}$ are their (linear) retracts as well – this depends on the specific nature of the balls $B_{j}$ i.e., the geometric features of their boundaries, for it is known for certain that not every linear subspace of $D$ can be its retract in general, for instance even in the simplest case $B_{1}=B_{2}=\Delta$, in which case however we can indeed so pinpoint, as was precisely the purpose of our theorem 1.7 above.
-Finally, we note that the maximum dimension of the nonlinear retracts of $D$ is $\max\{N_{1},N_{2}\}$ far less, in general, than the dimension of $D$ namely, $N_{1}+N_{2}$.
 More importantly perhaps is the following observation.
@@
 All the automorphisms of the polydisc are linear fractional transformations, thereby each one of them maps a complex linear subspaces to at-most a quadratic complex surface.
-Hence, as soon as $Z$ is ‘sufficiently’ non-linear (there are plenty of retracts in the polydisc which are graphs of monomial maps of degree $>2$ for instance, which are herein regarded as ‘sufficiently’ non-linear), it is certainly clear that there does not exist any automorphism of the polydisc mapping $Z$ onto $D_{L_{1}}$.
+Hence, as soon as $Z$ is ‘sufficiently’ non-linear (there are plenty of retracts in the polydisc which are graphs of monomial maps of degree >2>2 for instance, which are herein regarded as ‘sufficiently’ non-linear), it is certainly clear that there does not exist any automorphism of the polydisc mapping $Z$ onto $D_{L_{1}}$.
 Theorem 1.17 may therefore be viewed as providing an answer to the problem of recasting retracts in a domain into a simple ‘normal’ form (which cannot be reduced to a further simple one in general), by applying some appropriate automorphism of the domain.
 Remark 3.20.
-The condition of strict convexity is equivalent to the notion of $\mathbb{R}$-extremality (as will be discussed in the next section), which is obtained just by changing the ground field from $\mathbb{C}$ to the reals in that of $\mathbb{C}$-extremality already used in Vesentini’s theorem.
-The proof of the foregoing theorem will be phrased in a manner to make it clear that as a corollary to its proof it will follow that: we may relax the strict convexity condition on the pair of balls $B_{1},B_{2}$ to $\mathbb{C}$-extremality, to get the same conclusions for those linear retracts $D_{L}$ of $D=B_{1}\times B_{2}$ with the property that some point of $\partial D_{L}$ is an $\mathbb{R}$-extremal point of $\partial D$.
-This relaxation allows the applicability of this result to the case of balls which are $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal, for which simple examples exist: indeed, the ball in $\mathbb{C}^{N}$ with respect to the $l^{1}$-norm is one such (though balls in all other $l^{p}$-norms for $1<p<\infty$ satisfy the stronger condition of $\mathbb{R}$-extremality.
-As mentioned in the introduction, we shall skip the details of the extension of the arguments above to the case of the product of an arbitrary number of balls and only lay down the end result.
+It may very well happen that neither $\pi_{1}(Z),\pi_{2}(Z)$ need be linear subspaces of the domain, let alone retracts (neither of the respective factors $B_{1},B_{2}$ nor of $D$).
+For a simple example, consider the polydisc in $\mathbb{C}^{4}$ which we view as a product of $2$ copies of the bidisc i.e., let $D:=\Delta^{2}\times\Delta^{2}$; and, consider the linear subspace $L$ of $\mathbb{C}^{4}$ defined by $L:=span\{v_{1},v_{2}\}$, where $v_{1}=(1,1/2,1/4,0)\in\partial\Delta^{2}\times\Delta^{2},~v_{2}=(0,1/2,3/4,1)\in\Delta^{2}\times\partial\Delta^{2}$.
+We can identify first and second factor of $D$ as $\Delta^{2}\times\{0\}$ and $\{0\}\times\Delta^{2}$.
+Note that the holomorphic map $\rho$ on $D$ defined by $\rho(z_{1},z_{2},z_{3},z_{4})=(z_{1},\frac{z_{1}+z_{4}}{2},\frac{z_{1}+3z_{4}}{4},z_{4})$ gives a retraction map from $D$ onto $D_{L}$.
+Note that $D_{L}$ is a retract of $D$ and $\partial D_{L}$ intersects $\partial\Delta^{2}\times\Delta^{2}$ and $\Delta^{2}\times\partial\Delta^{2}$.
+Now we consider the projections restricted to $D_{L}$, which is given as follows:
+$\pi_{1}:L\to\pi_{1}(L)$ defined by
+$\pi_{1}(z_{1},\frac{z_{1}+z_{4}}{2},\frac{z_{1}+3z_{4}}{4},z_{4})=(z_{1},\frac{z_{1}+z_{4}}{2},0,0).$
+and $\pi_{2}:L\to\pi_{2}(L)$ defined by
+$\pi_{2}(z_{1},\frac{z_{1}+z_{4}}{2},\frac{z_{1}+3z_{4}}{4},z_{4})=(0,0,\frac{z_{1}+3z_{4}}{4},z_{4}).$
+Note that $\pi_{1}(D_{L})$ and $\pi_{2}(D_{L})$ are not linear subspaces of neither the respective factors ($\Delta^{2}\times\{0\}$, $\{0\}\times\Delta^{2}$) nor of $D$; indeed, this can be confirmed for instance by noting that that $(0,1/2)=\pi_{1}(v_{2})\in\pi_{1}(D_{L})$ and $(1/4,0)=\pi_{2}(v_{1})\in\pi_{2}(D_{L})$, but their scalings: $(0,3/4)\notin\pi_{1}(D_{L})$ and $(1/2,0)\notin\pi_{2}(D_{L})$ despite $(0,3/4,0,0)$ and $(1/2,0,0,0)$ are within $D$.
+As mentioned in the introduction, we shall skip the details of the extension of the arguments above to the case of an extended product of an arbitrary number of ‘balls’ and only lay down the end result; in-fact we only do this for the case of convex factors with $\mathbb{C}$-extremal boundaries which is of special importance as it leads to an associated result about norm-one projections with respect to those norms whose unit balls are such products.
+The corresponding statement for bounded balanced pseudoconvex domains with holomorphically extreme boundaries is also true and can be written down as in theorem 1.17.
 Theorem 3.21.
-Let $B_{1},B_{2},\ldots,B_{N}$ be a collection of bounded balanced strictly convex domains possibly in different complex Euclidean spaces (say, $B_{j}\subset\mathbb{C}^{n_{j}}$ for $j=1,\ldots,N$).
+Let $B_{1},B_{2},\ldots,B_{N}$ be a collection of bounded balanced convex domains with $\mathbb{C}$-extremal boundaries, possibly in different complex Euclidean spaces (say, $B_{j}\subset\mathbb{C}^{n_{j}}$ for $j=1,\ldots,N$).
 Then, every non-trivial retract through the origin of $D=B_{1}\times B_{2}\times\ldots\times B_{N}$ is a holomorphic graph over a linear subspace which itself is a retract of $D$; if the retract is not linear, this linear subspace is infact a subspace of the product of a subcollection of these balls $B_{j}$’s with values in the product of the remaining balls in this collection.
-More specifically and precisely, let $L=T_{0}Z$, $D_{L}=L\cap D$ and $\partial D_{L}=L\cap\partial D$.
-For each $p\in\{1,2,\ldots,N\}$, denote by $\Sigma_{p}$, the set of all $p$-tuples $(j_{1},j_{2},\ldots,j_{p})$ where $\{j_{1},j_{2},\ldots,j_{p}\}\subset\{1,2,\ldots,N\}$ with $j_{1}<j_{2}<\ldots<j_{N}$ (the cardinality of $\Sigma_{p}$ is ($N\choose p$) and note that there exists any given $Z,D_{L}$ as above, $p\in\{1,2,\ldots,N\}$ such that $\partial D_{L}$ is contained in that part of the boundary of the convex body $\overline{D}=\overline{B}_{1}\times\overline{B}_{2}\times\ldots\times\overline{B}_{N}$, obtained by replacing the factor $\overline{B}_{l}$ herein by $\partial B_{l}$ for each $l=j_{1},j_{2},\ldots,j_{p}$ where $\sigma=(j_{1},j_{2},\ldots,j_{p})\in\Sigma_{p}$.
-Then, $Z$ is the graph of a $B^{\sigma_{c}}$-valued holomorphic map over a $q$-dimensional linear retract of $B_{j_{1}}\times B_{j_{2}}\times\ldots\times B_{j_{p}}$ (which is included in the canonical manner into our product domain $D=B_{1}\times B_{2}\times\ldots\times B_{N}$), where:
+[…]
   • (i)
@@
   • (ii)
-    $B^{\sigma_{c}}$ is the product of balls given by $B_{k_{1}}\times B_{k_{2}}\times\ldots\times B_{k_{N-p}}$ where $\sigma_{c}\in\Sigma_{N-p}$ is the $(N-p)$-tuple complementary to $\sigma$ in the sense that $\sigma^{c}=(k_{1},\ldots,k_{N-p})\subset\{1,\ldots,N\}\setminus\{j_{1},\ldots,j_{p}\}$ and $k_{1}<k_{2}<\ldots<k_{N-p}$.
+    $B^{\sigma_{c}}$ is the product of balls given by $B_{k_{1}}\times B_{k_{2}}\times\ldots\times B_{k_{N-p}}$ where $\sigma_{c}\in\Sigma_{N-p}$ is the (Np)(N-p)-tuple complementary to $\sigma$ in the sense that $\sigma^{c}=(k_{1},\ldots,k_{N-p})\subset\{1,\ldots,N\}\setminus\{j_{1},\ldots,j_{p}\}$ and $k_{1}<k_{2}<\ldots<k_{N-p}$.
 Let $m=\min\{n_{1},n_{2},\ldots,n_{N}\}$, $\hat{\mu}_{m}=\max\{\mu_{1},\mu_{2},\ldots,\mu_{m-1},\mu_{m+1},\ldots,\mu_{N}\}$, where for all $j=1,2,\ldots,N$ we denote by $\mu_{j}$, the Minkowski functional of $B_{j}$.
@@
 In the pending case when $\partial D_{L}$ intersects more than one connected component of $\partial D\setminus\partial B_{e}$, the retract $Z$ is again a linear subspace of $D$ of complex dimension equal to the number of connected components that it intersects.
-The simplest case for $D$ as in the above theorem is obtained by taking all the $B_{j}$’s therein to be just the unit disc, in which case ofcourse $D$ is the standard unit polydisc $\Delta^{N}$, whose retracts were determined originally by Heath and Suffridge [Heath_Suff].
-To recover their result from the above theorem which only characterizes the retracts through the origin, we only have to note that the automorphism group of $\Delta^{N}$ acts transitively thereby implying that all retracts of $\Delta^{N}$ are obtained as biholomorphic images of those passing through the origin under some automorphism of $\Delta^{N}$; recalling that all automorphisms of $\Delta^{N}$ are actually fractional linear maps, we also note that the main essential feature of retracts through the origin described in the above theorem namely, as holomorphic graphs over some linear subspace is inherited by all retracts of $\Delta^{N}$, as fractional linear maps carry linear subspaces to affine-linear subspaces.
-Similar considerations apply whenever the factors $B_{j}$ are homogeneous domains; recall that a domain is said to be homogeneous if its automorphism group acts transitively and examples of bounded balanced convex domains which are also homogeneous include apart from $\Delta^{N}$, the standard Euclidean ball $\mathbb{B}^{N}$ and more generally, products of Cartan domains of various types (better known as bounded balanced symmetric domains).
-In such a case their product $D=B_{1}\times\ldots B_{N}$ is homogeneous as well and all its retracts are obtained as images under some element of ${\rm Aut}(D)$, of those retracts of $D$ passing through the origin which have been determined by the above theorem.
-However, we may not be able to say that every such retract (unless it passes through the origin), is a holomorphic graph over a linear subspace since elements of ${\rm Aut}(D)$ may well not be so simple (as linear fractional transformations) to preserve linearity as in the aforementioned instance of the polydisc.
-To give another concrete instance, where the main features of the retracts through the origin are indeed preserved, let us take each $B_{j}$ in the above theorem to be the standard unit ball with respect to the $l^{2}$-norm; so, let $\mathbb{B}_{1},\ldots,\mathbb{B}_{N}$ be standard Euclidean balls in complex Euclidean spaces of possibly different dimensions.
-Then as ${\rm Aut}(D)={\rm Aut}(\mathbb{B}_{1})\times\ldots\times{\rm Aut}(\mathbb{B}_{N})$, we see that the product domain $D=\mathbb{B}_{1}\times\ldots\times\mathbb{B}_{N}$ is also homogeneous and more importantly, all of its automorphisms are fractional-linear maps.
-However, despite all retracts of each of these $l^{2}$-balls being affine-linear, this does not mean that every retract of $D$ is affine-linear because it must be kept in mind that the operation of taking Cartesian product leads to new retracts in the form of holomorphic graphs (even through the origin), which are not linear.
-The foregoing remarks for a product of Euclidean balls apply as well to the case of products of balls with polydiscs such as $\Delta^{k}\times\mathbb{B}$ (or a larger collection of Euclidean balls and polydiscs in any order) which are also homogeneous domains with all of is automorphisms being fractional-linear maps.
-We conclude by noting that in these cases the characterization of retracts provided by theorems 1.15 and 3.20, is complete.
+Needless to say, when the product domain happens to be homogeneous such as a bounded balanced symmetric domain whose automorphisms are known, the above theorem suffices to give all its retracts.
 4. Proof of theorem 1.18
-The proof follows ideas from the artice [Fost] of Fornaess–Sibony.
+The proof follows ideas from the article [Fost] of Fornæss–Sibony.
 Let us begin by noting that it follows from theorem 1.15 proved in the foregoing section that every retract of a bounded balanced homogeneous domain is biholomorphic to the graph of some holomorphic function over a linear subspace of $D$.
@@
 Remark 4.1.
-We recall for one last time, that the class of homogeneous bounded balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein).
-It follows that a domain $\Omega$ as in the above theorem is a polyball (and convex), in-particular.
-5. Proof of Theorem 1.9
-Before discussing our result as in the title of this section on the retracts of $\ell_{q}$ balls for $10<q<$ 1 we first review the case $q\geq 1$.
-The main theorem about this case was already mentioned in theorem 1.7.
-However that formulation is highly coordinate dependent.
-Here we first re-formulate it in-terms of commutavity with standard projections; the proof is based on an idea from the work [Baronti-Pappini] of Baronti – Pappini.
+We recall for one last time, that the class of homogeneous bounded it balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein), a polyball.
+5. Proof of Theorems 1.8 and 1.10
+Before discussing our result as in the title of this section on the retracts of $\ell_{q}$ balls for $0<q<1$ we first review the case $q\geq 1$.
+This is not to repeat the details available in the classical treatise of Lindenstrauss – Tzafiri [Lindstraus], we note that this is still perhaps the only book where the the main theorem about this case namely, the one stated as theorem 1.6 is dealt with.
+To indicate our purpose here in reviewing this case, notice that the formulation of the determining criteria for the linear retracts of $\ell^{p}$-balls as stated in the just-mentioned theorem is highly coordinate dependent, leaving a a more geometric understanding to be desired.
+While geometric reformulations of the same have been recorded for instance in the survey [Beata-survey], unfortunately this article neither outlines a proof nor a reference for where it could be found; the book of Lindenstrauss – Tzafiri referred therein does not give one side of the result that we need, which to the best of our knowledge is not recorded elsewhere as well.
+We therefore proceed to unraveling in some detail, the geometric meaning of theorem 1.6 needed for a better understanding of the linear retracts of $l^{p}$-balls.
+To begin with, we observe that a linear slice of an $\ell^{p}$-ball need not be isometric to any lower dimensional $\ell^{p}$-ball.
+In fact, this happens only in very special circumstances, indeed precisely when the slice is a retract.
+As already mentioned, this result is not really new, being contained in essence in theorem 6.3 of [Beata-survey]; our purpose here is to provide the missing details, indeed a self-contained proof of this result, not to be found elsewhere.
+To attain this geometric result, we now recall a generalization of the notion of orthogonality in Euclidean spaces to normed spaces wherein the norm fails to be induced by any inner product; one that is most relevant for our aforementioned purposes first.
+Definition 5.1.
+Let $(X,\|\cdot\|)$ be a complex Banach space. For $x,y\in X$, we say that $x$ is $p$-orthogonal to $y$ and denoted $x\perp_{p}y$ if
+$\begin{cases}\|x+\lambda y\|^{p}=\|x\|^{p}+|\lambda|^{p}\|y\|^{p},~\forall\lambda\in\mathbb{C},~\text{when}~p<+\infty;\\ \|x+\lambda y\|=\max\{\|x\|,|\lambda|\|y\|\},~\forall\lambda\in\mathbb{C},~\text{when}~p=+\infty.\end{cases}$
+There are many notions of orthogonality and we shall not spend space in comparing the above with others except perhaps for the most popular known as Birkhoff-James orthogonality, which can be found in [Giles] and [Faulk].
+Denoted suggestively by ‘BJ’, this is defined as
+$x\perp_{BJ}y\Longleftrightarrow\|x+\lambda y\|\geq\|x\|~~\text{for every}~\lambda\in\mathbb{C}.$
+Note that if $x\perp_{p}y$ then $x\perp_{BJ}y$ and $y\perp_{BJ}x$.
+Therefore, $p$-orthogonality is a stronger condition than Birkhoff-James orthogonality.
+This generalized orthogonality relation is, in general not symmetric.
+If it is indeed the case that $x\perp_{BJ}y$ and $y\perp_{BJ}x$, then we say that $x$ and $y$ are bi-orthogonal (or symmetrically orthogonal).
+Note that if $u,v\in\mathbb{C}^{N}$ are mutually disjointly supported vectors then $u$ and $v$ are bi-orthogonal vectors in $(\mathbb{C}^{N},\|\cdot\|_{p})$ for $1\leq p\leq\infty$.
+But for each $p\neq 2$, there exists a linear subspace spanned by bi-orthogonal vectors which is not be spanned by mutually disjoint supported vectors, as illustrated in the following.
+Example: Consider the linear subspace $L=f^{-1}(0)$, where the linear functional $f:\mathbb{C}^{3}\to\mathbb{C}$ defined by $f(z_{1},z_{2},z_{3})=2z_{1}+z_{2}-z_{3}$.
+Next we claim that $L$ is spanned by the bi-orthogonal vectors (1,1,1)(1,-1,1) and $(0,1,1)$.
+To prove this claim, let $v_{1}=(1,-1,1)$ and $v_{2}=(0,1,1)$.
+Note that
+$\|v_{1}+\lambda v_{2}\|^{p}=\|(1,-1+\lambda,1+\lambda)\|^{p}=1+|1-\lambda|^{p}+|1+\lambda|^{p}$
+Since $|a+b|^{p}\leq 2^{p-1}(|a|^{p}+|b|^{p})$, we obtain that $\|v_{1}+\lambda v_{2}\|^{p}\geq 1+\frac{2^{p}}{2^{p-1}}=3=\|v_{1}\|^{p}$.
+This shows that $v_{1}$ is orthogonal to $v_{2}$.
+Using similar arguments, we obtain that $v_{2}$ is orthogonal to $v_{1}$.
+This finishes the proof of the claim.
+If $L$ is spanned by mutually disjoint supported vectors then by lemma 3.2 in [Baronti-Pappini], $f$ has atmost two non-null components.
+Hence, we obtain that $L$ cannot be realized as a subspace spanned by mutually disjoint supported vectors.
+Lemma 5.2.
+Let $P$ be a linear projection map on $\mathbb{C}^{N}$. Then $\|P\|=1$ if and only if ${\rm image}(P)$ is BJ-orthogonal to $\ker(P)$.
+Proof.
+Assume that $P$ is a projection of norm one.
+We want to show that ${\rm image}(P)$ is BJ-orthogonal to $\text{ker}(P)$.
+Towards this, let $z\in\text{image}(P)$ and $w\in\ker(P)$; so, $P(z+\lambda w)=P(z)$ by linearity of $P$.
+Being a projection of norm-one $P$ is contractive, thereby for all $\lambda\in\mathbb{C}$, we have $\|z\|=\|P(z+\lambda w)\|\leq\|z+\lambda w\|$.
+By the notion of BJ-orthogonality recalled above, $z$ is BJ-orthogonal to $w$.
+Conversely assume that $\text{image}(P)$ is BJ-orthogonal to $\ker(P)$.
+Since $P$ is a linear projection map, we have $\mathbb{C}^{N}=\text{image}(P)\oplus\ker(P)$.
+So, we can write each $z\in\mathbb{C}^{N}$ as $z=w_{1}+w_{2}$, where $w_{1}\in\text{image}(P)$ and $w_{2}\in\ker(P)$.
+Note that $\|P(z)\|=\|P(w_{1})\|=\|w_{1}\|$.
+As $w_{1}\perp_{BJ}w_{2}$, we get: $\|w_{1}\|\leq\|w_{1}+w_{2}\|=\|z\|$ i.e., $\|P(z)\|\leq\|z\|$, finishing the proof.
+∎
+Next, we recall the following result due to Lamperti (Corollary 2.1 of [Lamp]), which is an analogue of the parallelogram law for $\ell^{p}$-norms, determining precisely for which pair of vectors such a law holds.
+Lemma 5.3.
+Let $1\leq p<\infty$ with $p\neq 2$. Two vectors $u,v\in\mathbb{C}^{N}$ are mutually disjoint supported if and only if $u$ and $v$ are $p$-orthogonal.
+Though the following result is not new.
+For instance, refer to theorem 3 of chapter 6 in [Lacey].
+We therefore give a self-contained and direct exposition.
+For clarity, let us introduce a notation $\|\cdot\|_{p,k}$ to denote the $\ell_{p}$ norm on $\mathbb{C}^{k}$.
+Proposition 5.4.
+Let $W$ be a linear subspace of $(\mathbb{C}^{N},\|\cdot\|_{p,N})$.
+If $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{p,k})$ for $1\leq p\leq\infty$ with $p\neq 2$, then there exists a norm-one projection from $\mathbb{C}^{N}$ onto $W$.
+Proof.
+First, we consider the case $p\neq\infty$.
+Assume that $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{p,k})$.
+Then there exist a basis $\{u_{1},\ldots,u_{k}\}$ of $W$ and a linear isometry $S$ such that $\|u_{j}\|_{p,N}=1$ and $S(u_{j})=e_{j}$, where $e_{j}$ denotes the standard basis of $\mathbb{C}^{k}$ for $j=1,2,\ldots,k$.
+Note that $T:=S^{-1}$ is an isometry from $\mathbb{C}^{k}$ onto $W$ and $T(e_{j})=u_{j}$ for $j=1,2,\ldots,k$.
+First, we prove the result for $1\leq p<\infty$ with $p\neq 2$.
+Since $e_{j}$’s are disjointly supported, we can apply the above lemma 5.3 to $e_{j}$.
+Hence for each $j\neq l$, we have
+$\|e_{j}+e_{l}\|_{p,k}^{p}+\|e_{j}-e_{l}\|_{p,k}^{p}=2(\|e_{j}\|_{p,k}^{p}+\|e_{l}\|_{p,k}^{p})$
+As $T$ is an isometry, we obtain that
+$\|T(e_{j})+T(e_{l})\|_{p,N}^{p}+\|T(e_{j})-T(e_{l})\|_{p,N}^{p}=2(\|T(e_{j})\|_{p,N}^{p}+\|T(e_{l})\|_{p,N}^{p})$
+Again applying the above lemma 5.3, we get that the $u_{j}$’s are mutually disjoint supported vectors in $\mathbb{C}^{N}$.
+Following the proof of proposition 2.a.1 in [Lindstraus], the details regarding the construction of norm-one projection are given as follows.
+For each $j=1,2,\ldots,k$, let $\sigma_{j}:=supp(u_{j})$.
+Consider the linear functional $f_{j}:W\to\mathbb{C}$ defined by $f_{j}(u_{l})=\delta_{jl}$ (where $\delta_{jl}$ denotes the Kronecker-delta function) and extended to other vectors of $W$ by linearity.
+First we want to prove that $\|f_{j}\|_{p,N}=1$ for each $j=1,2,\ldots,k$. To prove this, write each $z\in W$ as $z=\sum_{l=1}^{k}c_{l}u_{l}$ for some scalars $(c_{l})_{l=1}^{k}$. Note that
+(5.1) $|f_{j}(z)|=\left|f_{j}\left(\sum_{l=1}^{k}c_{l}u_{l}\right)\right|=\left|\sum_{l=1}^{k}c_{l}f_{j}(u_{l})\right|=|c_{j}|\leq\|(c_{1},\ldots,c_{k})\|_{p,k}=\|S(z)\|_{p,k}.$
+Since $S$ is an isometry, we obtain that $|f_{j}(z)|\leq\|z\|_{p,N}$.
+This implies that $\|f_{j}\|_{p,N}\leq 1$ for all $j=1,2,\ldots,k$.
+Since $f_{j}(u_{j})=1$, we obtain that $\|f_{j}\|_{p,N}=1$.
+By the Hahn – Banach extension theorem, $f_{j}$ can be extended to a linear functional on $\mathbb{C}^{N}$ of the same norm which we also denote by $f_{j}$.
+So, $f_{j}:\mathbb{C}^{N}\to\mathbb{C}$ is a linear functional with $\|f_{j}\|_{p,N}=1$.
+For each $j$ with $1\leq j\leq k$, write $f_{j}(z)=\sum_{i=1}^{N}c_{j,i}z_{i}$ for some scalars $(c_{j,i})_{i=1}^{N}$.
+Define another linear functional $u_{j}^{*}$ on $\mathbb{C}^{N}$ by $u_{j}^{*}(z)=\sum_{i\in\sigma_{j}}c_{j,i}z_{i}$.
+Note that
+$\|u_{j}^{*}\|_{p,N}^{q}=\sum_{i\in\sigma_{j}}|c_{j,i}|^{q}\leq\sum_{i=1}^{N}|c_{j,i}|^{q}=\|f_{j}\|_{p,N}^{q}=1$
+where $q$ is the dual exponent of $p$ (i.e.
+$1/p+1/q=1$.).
+This implies in particular that $|u_{j}^{*}(\tilde{e_{i}})|=|c_{j,i}|\leq 1$, where $\tilde{e_{i}}$ denote the standard basis of $\mathbb{C}^{N}$.
+As $u_{j}^{*}(u_{j})=1$, we obtain that $\|u_{j}^{*}\|_{p,N}=1$ for every $j$.
+Now we define the projection $P:\mathbb{C}^{N}\to W$ by $P(z)=\sum_{j=1}^{k}u_{j}^{*}(z)u_{j}$.
+To prove this proposition, it suffices to show that $\|P\|_{p,N}=1$.
+If $z=\sum_{i=1}^{N}a_{i}\tilde{e_{i}}$ then $u_{j}^{*}(z)=\sum_{i=1}^{N}a_{i}u_{j}^{*}(\tilde{e_{i}})=\sum_{i\in\sigma_{j}}a_{i}c_{j,i}$.
+Since $\|u_{j}^{*}\|_{p,N}=1$, $|u_{j}^{*}(z)|^{p}\leq\sum_{i\in\sigma_{j}}|a_{i}|^{p}$.
+Hence we obtain that $\|P(z)\|_{p,N}^{p}=\sum_{j=1}^{k}|u_{j}^{*}(z)|^{p}\leq\|z\|^{p}$.
+This proves that $P$ is a norm-one projection from $\mathbb{C}^{N}$ onto $W$.
+Next, we prove the result for $p=\infty$.
+Write $S(z)=(S_{1}(z),\ldots,S_{k}(z))$.
+Note that $S_{j}(z)$ is a linear functional on $W$ for each $j$ with $1\leq j\leq k$.
+By the Hahn – Banach theorem, $S_{j}$ can be extended to a linear functional on $\mathbb{C}^{N}$ of the same norm which we denote by $\tilde{S_{j}}$.
+So, $\tilde{S_{j}}:\mathbb{C}^{N}\to\mathbb{C}$ is a linear functional so that $\|\tilde{S_{j}}\|=\|S_{j}\|\leq 1$.
+Define the linear map $\tilde{S}:\mathbb{C}^{N}\to\mathbb{C}^{k}$ by $\tilde{S}(z)=(\tilde{S}_{1}(z),\ldots,\tilde{S}_{k}(z))$.
+Note that $\|\tilde{S}\|\leq 1$ because $\|\tilde{S}\|=\sup\limits_{\begin{subarray}{c}1\leq j\leq N,~\|z\|=1\end{subarray}}|S_{j}(z)|\leq 1$.
+Since $\tilde{S}(u_{j})=e_{j}$, we get that $\|\tilde{S}\|=1$.
+Consider the linear map $P:\mathbb{C}^{N}\to W$ defined by $P(z)=T\circ\tilde{S}(z)$.
+Note that $P$ maps $\mathbb{C}^{N}$ onto $W$ and $P_{|_{W}}={\rm id}_{W}$.
+Hence, we get that $P$ is a norm one projection from $\mathbb{C}^{N}$ onto $W$.
+∎
+Proposition 5.5.
+Let $\mathbb{B}_{p}$ be the $\ell_{p}$-ball in $\mathbb{C}^{N}$ where $1\leq p\leq\infty$.
+The linear subspaces of $\mathbb{C}^{N}$ such that $(\mathbb{B}_{p})_{L}:=\mathbb{B}_{p}\cap L$ is a retract of $\mathbb{B}_{p}$ are precisely those for which $(\mathbb{B}_{p})_{L}$ is biholomorphic to a lower dimensional $\ell_{p}$-ball $\mathbb{B}_{p}^{k}$ for some $k$ with $1\leq k\leq N$.
+Proof.
+Assume that $(\mathbb{B}_{p})_{L}$ is biholomorphic to a lower dimensional $\ell_{p}$-ball $\mathbb{B}_{p}^{k}\subset\mathbb{C}^{k}$.
+By the theorem of W.
+Kaup and H.Upmeier in [Kaup-Upm] (exposited as Remark 16.4.2 d) in [Jrncki_invrnt_dst]), after passing to a different biholomorphic mapping we may assume that $\varphi(0)=0$.
+By an application of Cartan’s theorem, we get that $\varphi$ is linear.
+Note that for each $v\in(\mathbb{B}_{p})_{L}$, we have
+$\|v\|_{p}=h_{(\mathbb{B}_{p})_{L}}(v)=K_{(\mathbb{B}_{p})_{L}}(0,v)=K_{\mathbb{B}_{p}^{k}}(0,\varphi(v))=\|\varphi(v)\|_{p,k}$
+This implies that $\varphi$ is a linear isometry from $L$ onto $(\mathbb{C}^{k},\|\cdot\|_{p,k})$.
+By the above proposition 5.4, we obtain that $(\mathbb{B}_{p})_{L}$ is a linear retract of $\mathbb{B}_{p}$.
+Conversely assume that $(\mathbb{B}_{p})_{L}$ is a linear retract of $\mathbb{B}_{p}^{N}$.
+First we consider the case $1\leq p<\infty$.
+By theorem 2.a.4 in [Lindstraus], there exists an isometry from $L$ to $(\mathbb{C}^{k},\|\cdot\|_{p,k})$.
+This implies in particular that $(\mathbb{B}_{p})_{L}$ is biholomorphic to $\mathbb{B}_{p}^{k}$ for some $k$ with $1\leq k\leq N$.
+Next we consider the case for $p=\infty$.
+By proposition 3.16, there exists an isometry from $L$ to $(\mathbb{C}^{k},\|\cdot\|_{\infty})$.
+This proves that $\Delta^{N}_{L}$ is biholomorphic to lower dimensional polydisk in $\mathbb{C}^{k}$.
+∎
+We conclude the characterizations of retracts of $l^{p}$-balls in the convex case $p\geq 1$.
 Theorem 5.6.
-$R$ is a $k$-dimensional retract of the $\ell^{p}$-ball in $\mathbb{C}^{N}$ if and only if one of the following holds:
-  1. (1)
-    There exists a norm-one projection which maps $\mathbb{B}_{p}$ onto $R$ which commutes with the standard orthogonal projection onto one of the $(N-k)$ or $(N-(k-1))$-dimensional subspace spanned by co-ordinate vectors.
-  2. (2)
-    There exist a finite sequence of projections $P_{k},P_{k-1},\ldots,P_{1}$ such that for each $r=1,2,\ldots,k$, $P_{r}$ is a projection of norm one which maps $\mathbb{C}^{N}$ onto an $r$-dimensional subspace and $P_{k}(\mathbb{B}_{p})=R$.
+A linear subspace $R$ of the $\ell_{p}$-ball $\mathbb{B}_{p}$ is a $k$-dimensional retract of the $\mathbb{B}_{p}$ in $\mathbb{C}^{N}$ if and only if there exist a finite sequence of retracts $R_{1}\subset R_{2}\subset\ldots\subset R_{k}\subset\mathbb{B}_{p}$ such that for each $j=1,2,\ldots,k$, $R_{j}$ is a $j$-dimensional linear retract of $\mathbb{B}_{p}$ and $R_{k}=R$.
 Proof.
-It is easy to check that one implication of the theorem holds by using the definition of norm one projection.
-So we have only to prove the other implication.
-Assume that $R$ is a $k$-dimensional linear retract of $\mathbb{B}_{p}$.
-By theorem 2.a.4 in [Lindstraus], we obtain that $R=L\cap\mathbb{B}_{p}$ where $L$ is a $k$-dimensional linear subspace of $\mathbb{C}^{N}$ spanned by mutually disjointly supported vectors.
-Hence $L$ can be written as the solution set of $(N-k)$ linear equations, where each linear equation depends on atmost two variables.
-After applying Gaussian elimination and renaming variables, there is no loss of generality assume that above mentioned system of equations can be written as $AZ=0$, where $A$ is a $(N-k)\times N$ matrix $[B~{}B^{\prime}~{}B^{\prime\prime}]^{T}$ and $Z=[z_{1}~{}\ldots~{}z_{N}]^{T}$.
-The submatrices $B,B^{\prime},B^{\prime\prime}$ are of the following form:
-  1. (i)
-    $B=[B_{0}~{}B_{1}~{}\ldots B_{r-1}]^{T}$, where for each $i=0,1,\ldots,r-1$, $B_{i}=[b_{k_{i}+1}\ldots b_{k_{i+1}}]^{T}$ and $0=k_{0}<k_{1}<\ldots<k_{r}$.
-Futhermore for each $i$, $B_{i}$ satisfies the following properties:
-    1. (a)
-      For $j=k_{i}+1$ to $k_{i+1}-1$, $supp(b_{j})\cap supp(b_{j+1})\neq\phi$ and $|supp(b_{j})\cap supp(b_{j+1})|=1$.
-    2. (b)
-      For $i=0,1,\ldots,r-1$, if we define $supp(B_{i}):=\bigcup_{j=k_{i}+1}^{k_{i+1}}supp(b_{j})$ then $supp(B_{i})\cap supp(B_{j})=\phi$ for $i\neq j$.
-  2. (ii)
-    $B^{\prime}$ is of the form $[b_{1}^{\prime}~{}\ldots~{}b_{l}^{\prime}]^{T}$, where $supp(b_{i}^{\prime})\cap supp(b_{j}^{\prime})=\phi$ for $i,j=1,2,\ldots,l$ with $i\neq j$.
-  3. (iii)
-    $B^{\prime\prime}$ is of the form $[e_{s_{0}+1}~{}\ldots e_{s_{1}}]^{T}$. where $s_{0}=k_{r}+r+2l,s_{1}=N-k+r+l$ and $e_{i}$ denotes the standard basis of $\mathbb{C}^{N}$.
-Note that $|supp(B)|=|\bigcup_{i=0}^{r-1}supp(B_{i})|=k_{r}+r$, $|supp(B^{\prime})|=2l$ and $|supp(B^{\prime\prime})|=s_{1}-s_{0}$.
-Since the number of equation is $(N-k)$, we have $k_{r}+l+(s_{1}-s_{0})=N-k$.
-Because of this equation, we take $s_{1}=N-k+r+l$.
-Hence, we obtain the total number of dependent variables in the system of equations $AZ=0$ is equal to
-$|supp(B)|+|supp(B^{\prime})|+|supp(B^{\prime\prime})|$ $=$ $k_{r}+r+2l+(s_{1}-s_{0})$
-$=$ $k_{r}+r+2l+(N-k)-(k_{r}+l)$
-$=$ $N-k+r+l.$
-First we want to construct mutually disjoint supported vectors $(v_{i})_{i=1}^{k}$ such that $(v_{i})$ forms a basis for $L$.
-For each $i=1,\ldots,r$, consider the system of equations $B_{i}Z_{i}=0$, where $Z_{i}=[z_{k_{i-1}+1}~{}\ldots~{}z_{k_{i}}]^{T}$.
-This solution space is spanned by a non-zero vector $w_{i}=(\mu_{k_{i-1}+1},\ldots\mu_{k_{i}})\in\mathbb{C}^{k_{i}-k_{i-1}}$.
-Since every two co-ordinates of $w_{i}$ satisfies a linear equation in two variables, every component of $w_{i}$ is non-zero.
-For each $i=1,\ldots,r$, define $v_{i}:=\mu_{k_{i-1}+1}e_{k_{i-1}+1}+\ldots+\mu_{k_{i}}e_{k_{i}}$, where $e_{i}$ denote the standard basis in $\mathbb{C}^{N}$.
-Note that $v_{i}\in L$ and $supp(v_{i})=k_{i}-k_{i-1}$.
-For $i=r+1$ to $r+l$, choose a non-zero vector $v_{i}$ which is orthogonal to $b^{\prime}_{i-r}$ and $supp(v_{i})=supp(b^{\prime}_{i-r})$.
-For $i=r+l+1$ to $k$, $v_{i}=e_{i+t}$, where $t=N-k$.
-Define $u_{i}:=\frac{v_{i}}{\|v_{i}\|_{p}}$ for $i=1,2,\ldots,k$.
-Note that $(u_{i})_{i=1}^{k}$ forms a mutually disjointly supported vectors and a basis for $L$.
-Observe that $|supp(u_{i})|=2$ for $i=r+1$ to $r+l$ and $|supp(u_{i})|=1$ for $i=r+l+1$ to $k$.
-Using the idea in the proof of theorem 2.a.4 in [Lindstraus], we can construct projection $P$ of norm $1$ from given vectors $u_{1},\ldots,u_{k}$.
-Let $\sigma_{j}:=supp(u_{j})$ for $j=1,2,\ldots,k$.
-If we write $u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i}$, then we define the linear functional $u_{j}^{*}=\sum_{i\in\sigma_{j}}\frac{|\lambda_{i}|^{p}}{\lambda_{i}}e_{i}^{*}$.
-Observe that for $i=1,2,\ldots,r$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{i-1}+1},\ldots,z_{k_{i}}$, for $i=r+1,\ldots,r+l$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{r}+2i-r-1}$, $z_{k_{r}+2i-r}$ and for each $i\neq j$, $u_{j}^{*}$ does not depends on the variables $z_{k_{r}+2i-r-1}$ and $z_{k_{r}+2i-r}$.
-For $i=r+l+1$ to $k$, $u_{i}^{*}(z)=z_{i+t}$.
-Define a projection $P:\mathbb{C}^{N}\to\mathbb{C}^{N}$ by
-$P(z)=\sum_{j=1}^{k}u_{j}^{*}(z)u_{j}.$
-Note that $P$ is a projection of norm $1$ onto $L$. Next we are looking for standard projections which commutes with $P$. For $i=1,\ldots,r$, consider the standard projection,
-$\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
-Note that
-$P\circ\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
-For each $i=1,\ldots,r$, note that $k_{i-1}+1,\ldots,k_{i}$ components of $P$ only depends on $z_{k_{i-1}+1},\ldots,z_{k_{i}}$.
-So, $\pi_{k_{i-1}+1,\ldots,k_{i}}\circ P(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$.
-Let $m_{i}:=k_{r}+2i-r-1$. For $i=r+1,\ldots,r+l$, consider another projection
-$\pi_{m_{i},m_{i}+1}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{m_{i}-1},0,0,z_{m_{i+2}},\ldots,z_{N})$
-Using similar arguments as mentioned above, we can conclude that $P\circ\pi_{m_{i},m_{i}+1}=\pi_{s_{i},s_{i}+1}\circ P$.
-As we noted before, the total number of dependent variables in the system of equation $AZ=0$ is $N-k+r+l$.
-So, we consider another projection $\pi$ defined by $\pi(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{s_{1}},0,\ldots,0)$.
-Note that $P$ commutes with $\pi$ and $rank(\pi)=N-k+r+l$.
-If $r+l\geq 1$, then rank of $\pi$ is atleast $N-(k-1)$.
-If $r+l=0$, then $rank(\pi)$ is atleast $(N-k)$.
-This finishes the proof of (i).
-To prove (ii), take $P_{k}:=P$, which is a projection of norm one defined in the proof of (i).
-From the proof of (i), we obtain a standard projection $\pi_{k-1}$ of rank $(N-1)$ which commute with $P$.
-Let $P_{k-1}:=\pi_{k-1}\circ P$.
-Observe that $P_{k-1}$ is a norm one projection and $\rm image(P_{k-1})$ is a $k$-dimensional is a $(k-1)$-dimensional subspace of $\mathbb{C}^{N}$.
-Again applying the same argument to $P_{k-1}$, we get another projection $\pi_{k-2}$ of rank $(N-1)$ which commutes with $P$.
-Define $P_{k-2}:=\pi_{k-2}\circ P_{k-1}=\pi_{k-2}\circ\pi_{k-1}\circ P$.
-Hence we get that $P_{k-2}$ is a projection of norm one and it’s image is a $(k-2)$ dimensional subspace.
-Continuing this process, until we will obtain a projection $P_{1}$ of norm one such that $\rm image(P_{1})$ is a one dimensional subspace of $\mathbb{C}^{N}$.
+As the sufficiency part is utterly trivial, we need assume that $R$ is a linear retract of the $\ell_{p}$-ball $\mathbb{B}_{p}$ and verify the reachability of $R$ via a monotonically increasing (finite) sequence of retracts as asserted.
+By theorem 2.a.4 in [Lindstraus], $R$ is isometric to $k$-dimensional $\ell_{p}$-ball $\mathbb{B}_{p}^{k}$.
+Recalling there always exist retracts of all possible dimensions in $\mathbb{B}_{p}$, we have in-particular, a (k1)(k-1)-dimensional retract $R_{k-1}$ of $\mathbb{B}_{p}^{k}$.
+By lemma 2.13, $R_{k-1}$ is a linear retract of $\mathbb{B}_{p}$.
+Again applying theorem 2.a.4 in [Lindstraus] and lemma 2.13 to $\mathbb{B}_{p}^{k-1}$, we obtain that there exists a (k2)(k-2)-dimensional retract $R_{k-2}$ of $\mathbb{B}_{p}$.
+Continuing this process, until we get a one-dimensional retract $R_{1}$ of $\mathbb{B}_{p}$, renders the required sequence as in the statement of the theorem.
 ∎
-An apparently even more geometric formulation of the retracts of $\ell_{p}$ balls for $1\leq p<\infty,p\neq 2$ is that the retracts are precisely isometric copies of lower dimensional $\ell_{p}$-balls; see theorem 6.3 of [Beata-survey].
+We now put the above result together with the foregoing more geometric formulation of the linear retracts of $\ell_{p}$-balls as being precisely those linear intercepts which are isometric copies of lower dimensional $\ell_{p}$-balls.
+Doing so, leads to the conclusion that for any given retract $R$, there is a decreasing chain of isometric copies of lower dimensional $\ell^{p}$-balls reaching $R$ as in the above theorem.
 This is a feature shared by the $\ell_{p}$-‘balls’ even for positive values of the parameter $p<1$.
 Indeed, we now move onto this case dealt with by theorem 1.8.
-Before going into its proof 1.9, we need to make a couple of observations about the geometry of the $\ell_{q}$ ball $D_{q}$ in $\mathbb{C}^{N}$ (for $q<1$).
-Firstly, although its absolute image of is concave (easily and vividly seen in the two-dimensional case), $D_{q}$ is actually not concave (nor convex) domain in $\mathbb{C}^{N}$; not even locally, near any of the smooth boundary points.
+Before going into its proof, we need to make a couple of observations about the geometry of the $\ell_{q}$-ball $D_{q}$ in $\mathbb{C}^{N}$ for $q<1$ – for a vivid distinction between the convex case $p\geq 1$ with the non-convex case $p<1$, we use the symbol $q$ rather than $p$, for the latter case.
+To begin with then, although the absolute image of $D_{q}$ is concave (easily and vividly seen in the two-dimensional case), $D_{q}$ is actually not concave (nor convex) domain in $\mathbb{C}^{N}$; not even locally, near any of the smooth boundary points.
 We may establish this by computing the real Hessian at any smooth boundary point with positive real coordinates as $D_{q}$ is multicircular; we show this in the two-dimensional setting for simplicity, while it holds in higher dimensions as well by straight-forward extension of the arguments and computations.
@@
 To compute the real Hessian, note the following.
-$H$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}\frac{q^{2}}{4}|z|^{q-2}&0\\ 0&\frac{q^{2}}{4}|w|^{q-2},\end{bmatrix}$
-$\tilde{H}$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z^{2}}&\frac{\partial^{2}\varphi}{\partial z\partial w}\\ \frac{\partial^{2}\varphi}{\partial w\partial z}&\frac{\partial^{2}\varphi}{\partial w^{2}}\end{bmatrix}=\begin{bmatrix}\frac{q}{2}(\frac{q}{2}-1)|z|^{q}z^{-2}&0\\ 0&\frac{q}{2}(\frac{q}{2}-1)|w|^{q}w^{-2}\end{bmatrix}$
+$H$ $:=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}\frac{q^{2}}{4}|z|^{q-2}&0\\ 0&\frac{q^{2}}{4}|w|^{q-2},\end{bmatrix}$
+$\tilde{H}$ $:=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z^{2}}&\frac{\partial^{2}\varphi}{\partial z\partial w}\\ \frac{\partial^{2}\varphi}{\partial w\partial z}&\frac{\partial^{2}\varphi}{\partial w^{2}}\end{bmatrix}=\begin{bmatrix}\frac{q}{2}(\frac{q}{2}-1)|z|^{q}z^{-2}&0\\ 0&\frac{q}{2}(\frac{q}{2}-1)|w|^{q}w^{-2}\end{bmatrix}$
 Before proceeding, we remark that we shall utilize the above to compute the real Hessian only at the (smooth) absolute boundary points of $D_{q}$, such points being precisely those points $(x,y)$ of $\partial D_{q}$ both of whose coordinates are positive; while this suffices as $D_{q}$ is Reinhardt, it is important to note that the (tangent) vectors along which the Hessian form is evaluated need to be allowed to have non-real (complex) coordinates.
@@
 Then the real Hessian at the point $(x,y)$ evaluated along the direction $v$ is
-(5.1) $vH\overline{v}^{T}|_{(x,y)}+Re(vHv^{T})_{(x,y)}\\ =\frac{q^{2}}{4}\left(x^{q-2}|v_{1}|^{2}+y^{q-2}|v_{2}|^{2}\right)+\frac{q}{2}(\frac{q}{2}-1)Re\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right).$
+(5.2) $vH\overline{v}^{T}_{|_{(x,y)}}+Re(vHv^{T})_{(x,y)}\\ =\frac{q^{2}}{4}\left(x^{q-2}|v_{1}|^{2}+y^{q-2}|v_{2}|^{2}\right)+\frac{q}{2}(\frac{q}{2}-1)Re\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right).$
 If $v_{1},v_{2}$ are purely imaginary, then the above expression (5.2) is strictly positive. If $v_{1}$ and $v_{2}$ are real, then the above expression (5.2) becomes
@@
 which is negative as $q<1$.
-We conclude therefore that $D_{q}$ is neither convex nor concave (though of-course pseudoconvex) at $(x,y)\in\partial D_{q}$.
+We conclude therefore that $D_{q}$ is neither convex nor concave (though of course pseudoconvex) at $(x,y)\in\partial D_{q}$.
 Owing to already made remarks this holds at each point of $\partial D_{q}$ other than the negligible set of non-smooth boundary points of $D_{q}$.
-Proceeding now towards the retracts of $D_{q}$ (where we no longer assume $N=2$ for that would be an oversimplification unlike the matter in the preceding paragraph), it must first be noted that by proposition 1.4, every holomorphic retract of $\ell_{q}$-ball through the origin is actually linear even when $q<1$; for $1$-dimensional retracts, this follows from theorem 1.3 but as highlighted in the introduction, the thrust here is on retracts of arbitrary dimension.
+Proposition 5.7.
+Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$.
+Suppose $L$ is a linear subspace of $\mathbb{C}^{N}$ such that ${\rm co}(\overline{D_{L}})\neq C_{L}$, where $C_{L}:=C\cap L$ with $C={\rm co}(\overline{D})$, denoting the convex hull of $\overline{D}$.
+Then $D_{L}:=D\cap L$ is not a linear retract of $D$.
+Proof.
+To prove this proposition by contradiction, assume that $D_{L}$ is a linear retract of $D$.
+Then there exists a linear projection $P:\mathbb{C}^{N}\to L$ such that $P$ maps $D$ onto $D_{L}$.
+Since ${\rm co}(\overline{D_{L}})\neq C_{L}$ by hypothesis, there exists a point $p$ lying within the convex set $C_{L}$ but not in ${\rm co}(\overline{D_{L}})$.
+Write $p\in C_{L}\subset C$ as the convex combination of extreme points of $C$ i.e.
+$p=\sum_{j=1}^{k}\alpha_{j}e_{j}$, (where $\alpha_{j}$’s are positive reals with $\sum_{j=1}^{k}\alpha_{j}=1$), where $e_{j}$’s are extreme points of $C$ (here of course $e_{j}$’s need not be the standard basis vectors, though we use this notation only in this proof for convenience and the extreme points of $C$ need not be finite in number).
+Such an expression is possible due to the Krein-Milman theorem applied to the compact convex set $C$.
+By the linearity of $P$, we have
+$p=P(p)=\sum_{j=1}^{k}\alpha_{j}P(e_{j}).$
+As $P(e_{j})\in\overline{D_{L}}$, this shows that $p$ can be written as the convex combination of points in $\overline{D_{L}}$.
+This implies that $p\in{\rm co}(\overline{D_{L}})$, which contradicts the hypothesis that ${\rm co}(\overline{D_{L}})\neq C_{L}$.
+∎
+We now proceed to prove theorem 1.10.
+Proof of theorem 1.10:.
+Assume that $L$ is not spanned by vectors only from $\partial D\cap\partial C$.
+We want to show that $D_{L}$ is not a linear retract of $D$.
+By the above proposition, it suffices to show that ${\rm co}(\overline{D_{L}})\neq C_{L}$.
+Now we claim that there exists a vector $v\notin\partial D\cap\partial C$ such that $v$ is an extreme point of ${\rm co}(\overline{D_{L}})$.
+To prove this claim by contradiction, assume that every extreme point of ${\rm co}(\overline{D_{L}})$ lies in $\partial D\cap\partial C$.
+We may write each $w\in{\rm co}(\overline{D_{L}})$, as a convex combination of the extreme points in ${\rm co}(\overline{D_{L}})$.
+Let $w^{\prime}$ be an arbitrary point in $L$.
+Note that there exists $\lambda\in\mathbb{C}^{*}$ such that $\lambda w^{\prime}\in D_{L}$.
+So, we may express $\lambda w^{\prime}$ as a convex combination of the extreme points of ${\rm co}(\overline{D_{L}})$.
+Since these extreme points all lie within $\partial D\cap\partial C$, we get that $w^{\prime}$ belongs to span of vectors in $\partial D\cap\partial C$.
+Since this holds for every $w^{\prime}\in L$, $L$ is spanned by vectors in $\partial D\cap\partial C$, which contradicts the hypotheses of the theorem.
+This finishes the proof of the claim.
+By theorem 2.10.15 in [Meggn], $v\in\partial D_{L}=L\cap\partial D$.
+As $v\notin\partial C\cap\partial D$, we have $v\notin\partial C_{L}$.
+To prove ${\rm co}(\overline{D_{L}})\neq C_{L}$, choose $t_{0}>1$ such that $t_{0}v\in\partial C_{L}$.
+We want to show that $t_{0}v\notin{\rm co}(\overline{D_{L}})$.
+To prove this by contradiction, assume that $t_{0}v\in{\rm co}(\overline{D_{L}})$.
+Write $t_{0}v=c_{1}v_{1}+\ldots+c_{k}v_{k}$ (where $c_{j}$’s are positive reals with $\sum_{j=1}^{k}c_{j}=1$), where $v_{j}$’s are extreme points of ${\rm co}(\overline{D_{L}})$.
+After dividing $t_{0}$ on both sides, we obtain that
+$v=c_{1}\left(\frac{v_{1}}{t_{0}}\right)+\ldots+c_{k}\left(\frac{v_{k}}{t_{0}}\right)$
+Since $v_{j}/t_{0}\in D_{L}$ for each $j$, $v$ can be expressed as a convex combination of points in $D_{L}$.
+This contradicts the fact that $v$ is an extreme point of ${\rm co}(\overline{D_{L}})$.
+This proves that ${\rm co}(\overline{D_{L}})\neq C_{L}$.
+∎
+Lemma 5.8.
+Let $D$ be a bounded pseudoconvex domain in $\mathbb{C}^{N}$ and $p$ a boundary point of $D$ near which $\partial D$ is smooth.
+Fix $p\in\partial D$ and suppose that the (real) affine line $l(p,v):=\{p+tv:t\in\mathbb{R}\}$ with $v\in T_{p}^{\mathbb{C}}(\partial D)$ does not intersect $D$.
+Then the real Hessian $\mathcal{H}r(p)$ along $v$ is non-negative, where $r$ is any smooth defining function near $p$.
+Proof.
+Let us begin with the standard parametrization of the (real) line $l$ passing through $p$ in the direction of $v$ with $v\in T_{p}^{\mathbb{C}}(\partial D)$, namely $\varphi(t)=p+tv$.
+Let $r$ be a local defining function of $D$ defined on an open neighbourhood $U$ of $p$.
+So, $D\cap U=\{z\in U:r(z)<0\}$.
+Define $V:=\varphi^{-1}(U)$ and consider the smooth function $\psi:V\to\mathbb{R}$ defined by $\psi(t)=r\circ\varphi(t)$.
+Since ${\rm image}(\varphi)=l$ never intersects $D$, $\psi(t)\geq 0$ for all $t\in V$.
+This implies that $\psi$ attains a local minimum at $t=0$.
+Hence $\psi^{\prime\prime}(0)\geq 0$.
+Next we want to relate $\psi^{\prime\prime}(0)$ and $\mathcal{H}r(p)$ evaluated along the direction $v$.
+By an application of the chain rule, we obtain that
+$\psi^{\prime}(t)$ $=$ $\frac{\partial r}{\partial z_{1}}(\varphi(t))~v_{1}+\frac{\partial r}{\partial z_{2}}(\varphi(t))~v_{2}+\ldots+\frac{\partial r}{\partial z_{N}}(\varphi(t))~v_{N}$
+$\psi^{\prime\prime}(t)$ $=$ $\frac{d}{dt}\left(\frac{\partial r}{\partial z_{1}}(\varphi(t))~v_{1}\right)+\frac{d}{dt}\left(\frac{\partial r}{\partial z_{2}}(\varphi(t))~v_{2}\right)+\ldots+\frac{d}{dt}\left(\frac{\partial r}{\partial z_{N}}(\varphi(t))~v_{N}\right)$
+For each $j=1,2,\ldots,N$, we have
+$\frac{d}{dt}\left(\frac{\partial r}{\partial z_{j}}(\varphi(t))~v_{j}\right)=\left[\frac{\partial^{2}r}{\partial z_{1}\partial z_{j}}(\varphi(t))~v_{1}+\ldots+\frac{\partial^{2}r}{\partial z_{j}^{2}}(\varphi(t))~v_{j}+\ldots+\frac{\partial^{2}r}{\partial z_{N}\partial z_{j}}(\varphi(t))~v_{N}\right]v_{j}$
+Hence, we obtain that $\psi^{\prime\prime}(0)=\displaystyle\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial z_{k}}(p)~v_{j}v_{k}$. The real Hessian $\mathcal{H}r(p)$ evaluated along $v$ is
+(5.3) $2\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial\overline{z}_{k}}(p)~v_{j}\overline{v}_{k}+2Re\left(\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial z_{k}}(p)~v_{j}v_{k}\right)$
+By the (Levi) pseudoconvexity of $D$ at $p$, we then have for every vector $v\in T_{p}^{\mathbb{C}}(\partial D)$, that:
+$\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial\overline{z}_{k}}(p)~v_{j}\overline{v}_{k}\geq 0.$
+The second sum in (5.3) is non-negative, and it equals $2\psi^{\prime\prime}(0)$, which was noted to be non-negative.
+It therefore follows from (5.3) that the real Hessian $\mathcal{H}r(p)$ along $v$ is non-negative.
+∎
+Definition 5.9.
+Let $D$ be a domain in $\mathbb{C}^{N}$.
+We say that $D$ admits retracts along an (or, for some, or corresponding to an) open piece of directions at $p\in D$ if there exists a (non-empty) open subset $\Gamma$ of $\partial\mathbb{B}$ such that for each $v\in\Gamma$, there is a retract $Z$ of $D$ passing through $p$ with $v\in T_{0}Z$.
+We now restate theorem 1.8 as follows and prove it.
+Corollary 5.10.
+Consider the balanced pseudoconvex domain given by
+$D:=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:|z_{1}|^{q_{1}}+\ldots+|z_{N}|^{q_{N}}<1\}$
+where the $q_{j}$’s are all positive numbers less than one.
+Let $L$ be a linear subspace of $\mathbb{C}^{N}$ which cannot be spanned precisely by vectors only from the standard basis vectors i.e.
+$L\neq\operatorname{span}(S)$ for any $S\subset B_{std}=\{e_{1},\ldots,e_{N}\}$.
+Then $D_{L}$ is not a retract of $D$; (thereby, the retracts of $D$ are precisely the orthogonal projections of $D$ onto a span of a finite subset of $B_{std}$).
+Consequently, $D$ does not admit any retract along any open piece of directions at the origin.
+Proof.
+Let $C$ denote the convex hull of $\overline{D}$.
+First we claim that $C=\overline{\mathbb{B}}_{1}$, where $\mathbb{B}_{1}:=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:|z_{1}|+\ldots+|z_{N}|<1\}$ , the $\ell_{1}$-ball in $\mathbb{C}^{N}$.
+As $q_{j}<1$ for each $j$, we have $|z_{j}|<|z_{j}|^{q_{j}}$.
+This implies that $\sum_{j=1}^{N}|z_{j}|<\sum_{j=1}^{N}|z_{j}|^{q_{j}}$ and thereby $D\subset\mathbb{B}_{1}$.
+As $\overline{\mathbb{B}}_{1}$ is convex, $C\subset\overline{\mathbb{B}}_{1}$.
+Now, $\overline{\mathbb{B}}_{1}$ is the convex hull of the extreme points of $\overline{\mathbb{B}}_{1}$, but these extreme points belong to $\partial D$.
+Therefore $C=\overline{\mathbb{B}}_{1}$.
+Now, we claim that $\partial C\cap\partial D=\{\lambda e_{i}:\lambda\in\mathbb{C},~|\lambda|=1\}$, where $e_{i}$’s are the standard basis vectors of $\mathbb{C}^{N}$.
+To prove this claim by contradiction, assume that there exists $p$ different from the scalar multiple of $e_{i}$ belonging to $\partial C\cap\partial D$.
+First, consider the case that every component of $p$ is non-zero.
+In that case, $p$ is a smooth boundary point of $D$.
+Since $C$ is convex at $p$, there exists a supporting hyperplane $H_{p}$ to $C$ at $p$.
+Note that the supporting hyperplane at $p$ coincides with the tangent plane at $p$ because $p$ is a smooth boundary point of $D$.
+As $D\subset C$ and $p\in\partial C\cap\partial D$, $H_{p}$ serves as a supporting hyperplane to $\partial D$ at $p$.
+Let $l$ be a (real) affine line passing through $p$ in the direction of $v$, which is contained in $H_{p}$.
+By the above lemma 5.8, the real Hessian $\mathcal{H}r(p)(v)\geq 0$.
+Since this holds for all vectors $v\in T_{p}(\partial D)$, $\mathcal{H}r(p)$ is positive semi-definite on $T_{p}(\partial D)$.
+Now, we compute the real Hessian of $D$ at $p$.
+Since $D$ is multicircular, after applying an automorphism on $D$, we may assume $p$ is of the form $(x_{1},\ldots,x_{N})$, where $x_{i}$’s are positive.
+Upto a non-negative scaling factor of $2$, the real Hessian at the point $(x_{1},\ldots,x_{N})$ along the vector $(v_{1},\ldots,v_{N})$ is given by
+$\frac{q_{1}^{2}}{4}x_{1}^{q_{1}-2}|v_{1}|^{2}+\frac{q_{2}^{2}}{4}x_{2}^{q_{2}-2}|v_{2}|^{2}+\ldots+\frac{q_{N}^{2}}{4}x_{N}^{q_{N}-2}|v_{N}|^{2}\\ +Re\left(\frac{q_{1}}{2}(\frac{q_{1}}{2}-1)x^{q_{1}-2}v_{1}^{2}+\frac{q_{2}}{2}(\frac{q_{2}}{2}-1)x_{2}^{q_{2}-2}v_{2}^{2}+\ldots+\frac{q_{N}}{2}(\frac{q_{N}}{2}-1)x_{N}^{q_{N}-2}v_{N}^{2}\right)$
+When $v_{1},\ldots,v_{N}$ are positive reals, this real Hessian may be written as
+$\frac{q_{1}^{2}-q_{1}}{2}x_{1}^{q_{1}-2}|v_{1}|^{2}+\frac{q_{2}^{2}-q_{2}}{2}x_{2}^{q_{2}-2}|v_{2}|^{2}+\ldots+\frac{q_{N}^{2}-q_{N}}{2}x_{N}^{q_{N}-2}|v_{N}|^{2}$
+As $q_{i}<1$, this is negative, contradicting the above conclusion that the real Hessian is positive semi-definite on $T_{p}(\partial D)$. This finishes the proof of the claim.
+Next, we consider the case that at least one of the components of $p$ is zero.
+After a permutation of co-ordinates, we may assume that $p$ is of the form $(p_{1},\ldots,p_{M},0,\ldots,0)$ for some $M<N$.
+Consider the domain $D^{\prime}:=\{(z_{1},\ldots,z_{M})\in\mathbb{C}^{M}:|z_{1}|^{q_{1}}+\ldots+|z_{M}|^{q_{M}}<1\}$.
+Note that $p^{\prime}:=(p_{1},\ldots,p_{M})$ is a smooth boundary point of $D^{\prime}$ and $D^{\prime}$ is convex at $p^{\prime}$.
+By using the same argument applied to $D$, we conclude that this case does not arise.
+By the above theorem 1.10, if the linear subspace $L$ is not spanned only by vectors $e_{i}$ then $D_{L}$ is not a retract of $D$.
+To conclude, let $p=(p_{1},\ldots,p_{N})\in\partial D$ with $p_{i}\neq 0$ for all $i$.
+Note that $p$ does not belong to a non-trivial linear subspace spanned by standard basis vectors.
+Hence, there does not exist a retract of $D$ passing through the origin in the direction of $p$.
+Let $S:=\{(z_{1},\ldots,z_{N})\in\partial D:z_{i}\neq 0~\text{for ~all}~i\}$.
+Note that $S$ is an open dense subset of $\partial D$ and for each $p\in S$, there does not exist a retract of $D$ passing through the origin in the direction of $p$.
+Since the complement of $S$ in $\partial D$ forms a nowhere dense subset of $\partial D$, there does not exist any retract of $D$ passing through the origin along an open piece of directions.
+∎
+Theorem 1.10 can be employed to determine retracts of some non-decoupled domains as well. We demonstrate this in the following.
+Corollary 5.11.
+Consider the non-decoupled (balanced pseudoconvex) domain given by $\Omega:=\{(z_{1},z_{2})\in\mathbb{C}^{2}:|z_{1}|^{2q_{1}}+|z_{2}|^{2q_{2}}+|z_{1}z_{2}|^{2q_{3}}<1\}$ where the $q_{j}$’s are all positive numbers less than $1/2$.
+Let $L$ be a linear subspace of $\mathbb{C}^{2}$ not spanned by vectors only from the standard basis vectors.
+Then $\Omega_{L}$ is not a retract of $\Omega$; (thereby, the retracts of $\Omega$ are precisely the orthogonal projections of $D$ onto a span of the standard basis vectors).
+Consequently, there are no retracts of $\Omega$ passing through the origin along an open piece of directions.
+Proof.
+First, we compute the real Hessian of the defining function of $\Omega$ at positive real coordinates $(x,y)$. Consider the function
+$\varphi(z_{1},z_{2})=|z_{1}|^{2q_{1}}+|z_{2}|^{2q_{2}}+|z_{1}z_{2}|^{2q_{3}}-1=z_{1}^{q_{1}}\overline{z_{1}}^{q_{1}}+z_{2}^{q_{2}}\overline{z_{2}}^{q_{2}}+(z_{1}z_{2})^{q_{3}}(\overline{z_{1}z_{2}})^{q_{3}}-1$
+Note that
+$H$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}q_{1}^{2}|z_{1}|^{2(q_{1}-1)}+q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}|z_{1}|^{-2}&q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}(z_{1}\overline{z_{2}})^{-1}\\ q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}(\overline{z_{1}}z_{2})^{-1}&q_{2}^{2}|z_{1}|^{2(q_{2}-1)}+q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}|z_{2}|^{-2}\end{bmatrix}$
+In convenience, introduce the notation
+$\tilde{H}=\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z_{1}^{2}}&\frac{\partial^{2}\varphi}{\partial z_{1}\partial z_{2}}\\ \frac{\partial^{2}\varphi}{\partial z_{2}\partial z_{1}}&\frac{\partial^{2}\varphi}{\partial z_{2}^{2}}\end{bmatrix}$
+where
+$\frac{\partial^{2}\varphi}{\partial z_{1}^{2}}=q_{1}(q_{1}-1)|z_{1}|^{2q_{1}}z_{1}^{-2}+q_{3}(q_{3}-1)|z_{1}|^{2q_{3}}z_{1}^{-2}|z_{2}|^{2q_{3}},$
+$\frac{\partial^{2}\varphi}{\partial z_{1}\partial z_{2}}=\frac{\partial^{2}\varphi}{\partial z_{2}\partial z_{1}}=q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}(z_{1}z_{2})^{-1},$
+$\frac{\partial^{2}\varphi}{\partial z_{2}^{2}}=q_{2}(q_{2}-1)|z_{2}|^{2q_{2}}z_{2}^{-2}+q_{3}(q_{3}-1)|z_{1}z_{2}|^{2q_{3}}z_{2}^{-2}.$
+Let $v=[v_{1}~v_{2}]$ and $\overline{v}=[\overline{v_{1}}~\overline{v_{2}}]$.
+Assume that $v_{1},v_{2}$ are real.
+Upto a non-negative scaling factor of $2$, the real Hessian at the point $(x,y)$ along the vector $v$ is
+$\overline{v}Hv^{T}|_{(x,y)}+Re(v\tilde{H}v^{T})_{(x,y)}=\left[q_{1}(2q_{1}-1)x^{2(q_{1}-1)}+q_{3}(2q_{3}-1)x^{2(q_{3}-1)}y^{2q_{3}}\right]|v_{1}|^{2}+\\ \left[q_{2}(2q_{2}-1)y^{2(q_{2}-1)}+q_{3}(2q_{3}-1)x^{2q_{3}}y^{2(q_{3}-1)}\right]|v_{2}|^{2}+4q_{3}^{2}(xy)^{2q_{3}-1}xy$
+The complex tangent space at $(x,y)$ is
+$\{(v_{1},v_{2}):\left(q_{1}|z_{1}|^{2(q_{1}-1)}\overline{z}_{1}+q_{3}|z_{1}|^{2(q_{3}-1)}|z_{2}|^{q_{3}}\right)v_{1}=-\left(q_{2}|z_{2}|^{2(q_{2}-1)}\overline{z}_{2}+q_{3}|z_{2}|^{2(q_{3}-1)}|z_{1}|^{q_{3}}\right)v_{2}\}$
+This shows that every real vector $v=(v_{1},v_{2})$ in $T_{p}(\partial\Omega)$ for $p\in\partial\Omega$ with positive real co-ordinates whose components $v_{1},v_{2}$ are necessarily opposite in sign.
+Hence we obtain that the real Hessian at the point $(x,y)$ along $v$ is negative.
+Next, we claim that $C=\overline{\mathbb{B}}_{1}$, the $\ell_{1}$-ball in $\mathbb{C}^{2}$.
+Note that for each $j$, $|z_{j}|<|z_{j}|^{q_{j}}$.
+This implies that $|z_{1}|+|z_{2}|<|z_{1}|^{q_{1}}+|z_{2}|^{q_{2}}<|z_{1}|^{q_{1}}+|z_{2}|^{q_{2}}+|z_{1}z_{2}|^{q_{3}}$ and thereby $\Omega\subset\mathbb{B}_{1}$.
+As $\overline{\mathbb{B}}_{1}$ is convex, $C\subset\overline{\mathbb{B}}_{1}$.
+Note that $\overline{\mathbb{B}}_{1}$ is the convex hull of extreme points of $\overline{\mathbb{B}}_{1}$ which belongs to $\partial\Omega$.
+Therefore $C=\overline{\mathbb{B}}_{1}$.
+By using the same arguments in the above corollary, we can prove that $\partial C\cap\partial\Omega=\{\lambda e_{i}:\lambda\in\mathbb{C},~|\lambda|=1\}$, where $e_{i}$’s are the standard basis vectors in $\mathbb{C}^{N}$.
+After applying the above theorem 1.10, we obtain that if the linear subspace $L$ is not spanned only by vectors $e_{i}$, then $\Omega_{L}$ is not a retract of $\Omega$.
+Hence, there does not exist retracts of $\Omega$ passing through the origin with an open piece of directions.
+∎
+We now give a direct and more elementary proof for the retracts of the standard $\ell^{q}$-ball $D_{q}$ (where we no longer assume $N=2$ for that would be an oversimplification unlike the matter in the preceding paragraph), it must first be noted that by proposition 1.3, every holomorphic retract of $\ell_{q}$-ball through the origin is actually linear even when $q<1$.
 So, if $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$.
@@
 where
-$t_{j}=\frac{1}{\left(1+\sum_{k=r+1}^{N}|a_{j,k}|^{q}\right)^{1/q}}.$
+$t_{j}=\frac{1}{\left(1+\sum_{\begin{subarray}{c}k=r+1\end{subarray}}^{N}|a_{j,k}|^{q}\right)^{1/q}}.$
 For $1\leq j\leq r$, write
@@
 $P(e_{j})=\frac{1}{t_{j}}P(v_{j})-\left(\sum_{l=r+1}^{N}a_{j,l}P(v_{l})\right).$
-Since $P$ maps $\mathbb{C}^{N}$ into $Y$, $P(v_{j})\in Y$. Write
+Since $P(v_{j})\in Y$, there exists scalars $u_{1,j},\ldots,u_{r,j}$ such that
 (5.5) $P(v_{j})=\left(u_{1,j},\ldots,u_{r,j},\sum_{k=1}^{r}a_{k,r+1}u_{k,j},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,j}\right)~\text{for}~r+1\leq j\leq N.$
@@
 $P(e_{j})=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}\left(u_{1,l},\ldots,u_{r,l},\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)\right)\\ =\left(-\sum_{l=r+1}^{N}a_{j,l}u_{1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{j-1,l},1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l},-\sum_{l=r+1}^{N}a_{j,l}u_{j+1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{r,l},\right.\\ \left.a_{j,r+1}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,a_{j,N}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,N}u_{k,l}\right).$
-Note that $P(v_{j})\in\overline{D_{q}}$ for all $j$ So, we get the following inequality for $1\leq j\leq r$,
+Note that $P(v_{j})\in\overline{D_{q}}$ for all $j$ . So, we get the following inequality for $1\leq j\leq r$,
 (5.6) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|^{q}+\sum_{1\leq k\leq r~~k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}\left|a_{j,s}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,s}u_{k,l}\right|^{q}\leq 1.$
@@
 (5.7) $\sum_{k=1}^{r}|u_{k,j}|^{q}+\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}~u_{k,j}\right|^{q}\leq 1.$
-By reverse triangle inequality for $q^{\text{th}}$ power where $10<q<$ 1: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.4) becomes
-$-\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
+By reverse triangle inequality for $q^{\text{th}}$ power where $0<q<1$: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.6) becomes
+$\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~~k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\left|\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}a_{j,l}~a_{k,s}u_{k,l}\right|^{q}\leq 0.$
 Combining the above inequality with (5.7), we obtain the following inequality
-$-\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,j}\right|^{q}\right)+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
+$\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,l}\right|^{q}\right)+\sum_{1\leq m\leq r~~m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\left|\sum_{k=1}^{r}a_{j,l}~a_{k,s}u_{k,l}\right|^{q}\leq 0.$
 Simplifying this inequality, we get
@@
 This is obviously a contradiction, which completes the proof. ∎
-To make some noteworthy remarks here, let us also record a special case about all non-trivial retracts through the origin of such a domain in $\mathbb{C}^{2}$ where this result is not new (though our proof is direct and more elementary).
-Proposition 5.2.
-Let $q$ be an arbitrary positive number less than one.
-Consider the $\ell^{q}$-‘ball’ given by $D_{q}=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<1\}$ and the subspace $Y=\{(z,w)\in\mathbb{C}^{2}:z-mw=0\}$ (where $0\neq m\in\mathbb{C}$) of $\mathbb{C}^{2}$.
-Then the one dimensional subspace $Y_{q}=D_{q}\cap Y$ is not a retract of $D_{q}$.
-In-fact, the holomorphic retracts through the origin are linear and more precisely, the intercepts of the coordinate axes with $D_{q}$.
-Owing to remarks made in the introduction (in greater detail, the work of section 8 below), it follows as an immediate corollary that $D_{q}$ is not biholomorphic to any product domain; it must also be noted here that retracts can be used to immediately show inequivalence with convex domains – for instance, the non-existence of retracts of $D_{q}$ in certain directions suffices to prove its biholomorphic distinctness from bounded convex domains (the unbounded case is ruled out simply by hyperbolicity of $D_{q}$).
+For certain reasons below, we explicitly mention about the special case in dimension $2$.
+Owing to remarks made in the introduction, it follows as an immediate corollary that $D_{q}=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<1\}$ is not biholomorphic to any product domain.
 A more important relevance of this example is noted next.
@@
 We recall from [Vigue] that fixed point sets of holomorphic self-maps of a any given bounded (not necessarily balanced) convex domain $D$ are precisely the same as its retracts.
-The example of the above proposition proves that this coincidence fails as soon as we drop convexity.
-Indeed, for the non-convex (but pseudoconvex) balanced domain $D_{q}$ in $\mathbb{C}^{2}$ as above, the fixed point set of the holomorphic self-map of $D_{q}$ as simple as the flip $(z,w)\to(w,z)$, namely the linear subspace $\{(z,w)\in\mathbb{C}^{2}\;:\;w=z\}$, is not realizable as a holomorphic retract of $D_{q}$, which is part of the content of the proposition 5.2 and its higher dimensional analogue in theorem 1.9 above.
+The example $D_{q}$ in $\mathbb{C}^{2}$ above suffices to show that this coincidence fails as soon as we drop convexity.
+Indeed, for the non-convex (but pseudoconvex) balanced domain $D_{q}$ in $\mathbb{C}^{2}$ as above, the fixed point set of the holomorphic self-map of $D_{q}$ as simple as the flip $(z,w)\to(w,z)$, namely the linear subspace $\{(z,w)\in\mathbb{C}^{2}\;:\;w=z\}$, is not realizable as a holomorphic retract of $D_{q}$.
 The $\ell_{q}$-balls dealt with above are bounded by holomorphically extreme real hypersurfaces.
@@
 We begin with an example where we can show existence of higher dimensional retracts; the purpose being: to record atleast one simple example to affirmatively answer the question as to whether there exists non-homogeneous pseudoconvex domains in $\mathbb{C}^{N}$ (i.e., when $N$ is not restricted to be $2$ or the domain being convex) with the property that there are retracts through every point of the domain and all possible dimensions; given Lempert’s apt remarks about the non-existence of retracts of dimension any bigger than one, in general.
-As our purpose is in recording the simplest cases first, the examples may well not be impressive.
+As our purpose is in recording the simplest cases first, the next couple of examples may well not be impressive.
 Example: Consider the unbounded balanced non-convex domain given by $H^{N}:=\{(z_{1},\cdots,z_{N})\in\mathbb{C}^{N}:|z_{1}\ldots z_{N}|<1\}$.
@@
 It suffices to prove that the $k$-dimensional affine subspace $H^{N}\cap A$ is a retract of $H$.
-First we do this for $k=N-1$ and the rest follows recursively as noted below.
-To this end, we let $L_{c_{N}}:=\{(z_{1},\ldots,z_{N-1},z_{N})\in\mathbb{C}^{N}:z_{N}=c_{N}\}$ and show that the $(N-1)$-dimensional affine subspace $H^{N}\cap L_{c_{N}}$ of $H^{N}$ is a retract of $H^{N}$.
+First, we do this for k=N1k=N-1 and the rest follows recursively as noted below.
+To this end, we let $L_{c_{N}}:=\{(z_{1},\ldots,z_{N-1},z_{N})\in\mathbb{C}^{N}:z_{N}=c_{N}\}$ and show that the (N1)(N-1)-dimensional affine subspace $H^{N}\cap L_{c_{N}}$ of $H^{N}$ is a retract of $H^{N}$.
 We split into $2$ cases depending on $c_{N}$ being zero or not.
@@
 Continuing this process, we establish in $k$ steps, that $H^{N}\cap A$ is a $k$-dimensional retract of $H^{N}$.∎
-At this juncture, we can settle a question about whether holomorphic retraction mappings are automatically submersions so as (for instance) to render the hypotheses of theorem 1.17 redundant; to answer this, it suffices to consider the two-dimensional version of the above domain.
+At this juncture, we can settle a question about whether holomorphic retraction mappings are automatically submersions so as (for instance) towards rendering hypotheses such as (ii) in theorem 1.19 redundant in general.
+We briefly address this by considering the two-dimensional version of the above domain (we note here that $D_{L}$ is not bounded as stipulated by theorem 1.19).
 Theorem 5.13.
@@
 As $f(1)=0$, that constant is zero, thereby $f(w)=\rho_{1}(0,w)=0$ for all $w\in\mathbb{C}$.
-As $H^{2}$ is a logarithmically convex complete Reinhardt domain, $\rho_{1}$ can be expressed as a single power series expansion on $H^{2}$.
+As $H^{2}$ is a (logarithmically convex) complete Reinhardt domain, $\rho_{1}$ can be expressed as a single power series expansion on $H^{2}$.
 This means that there exist scalars $c_{j,k}$ such that for all $(z,w)\in H^{2}$,
@@
 Similarly we can consider another entire function $g:\mathbb{C}\to\Delta$ defined by $g(z)=\rho_{1}(z,0)$.
-By the same arguments, we obtain that $g$ for all $z\in\mathbb{C}$ and hence $P_{1}\equiv 0$.
+By the same arguments, we obtain that $P_{1}\equiv 0$.
 Therefore from the above equation (5.10), we arrive at our aforementioned deduction about the form of the first component namely, $\rho_{1}(z,w)=zwh(z,w)$ for some holomorphic function $h$ on $H^{2}$.
@@
 A more important point is noted next.
-A direct consequence of the above theorem is that unless we are in very special settings even within the class of balanced pseudoconvex domains, the hypothesis about submersivity of retractions as in theorem 1.17 cannot be dropped.
+A direct consequence of the above theorem is that unless we are in very special settings even within the class of balanced pseudoconvex domains, the hypothesis (ii) along-with boundedness of $D_{L}$ as in theorem 1.19 cannot be dropped.
 For a quick (counter-)example, observe that the retract $Z$ in the foregoing theorem 5.13 cannot be realized as the graph over any linear retract of $H^{2}$ as in the conclusion of theorem 1.19.
 Indeed, the only (non-trivial) linear retracts of $H^{2}$ are the pair of coordinate axes, on neither of which $Z$ can be expressed as a graph, for then $Z$ would be biholomorphic to $\mathbb{C}$, manifestly false.
-It is also possible to give a non-affine example in this simple domain itself to show that there are retracts which are impossible to be realized as a graph over any affine linear subspace of $H$, thereby showing (again) that the conclusion of theorem 1.17 can fail as soon as its hypotheses, particularly about submersivity of the retractions fail which is the purpose of the following corollary.
+It is also possible to give a non-affine example in this simple domain itself to show that there are retracts which are impossible to be realized as a graph over any affine linear subspace of $H$, thereby showing (again) that the conclusion of theorem 1.19 can fail as soon as some of its hypotheses is dropped as mentioned above, which is the purpose of the following corollary.
 Corollary 5.15.
@@
 Note that $\rho$ is a retraction map on $H^{2}$ and ${\rm image}(\rho)=Z=\{(\zeta,1):\zeta\in\Delta\}$ and we may apply theorem 5.13 to $\rho$.
-But then the submersitivity of $\tilde{\rho}$ at the origin is equivalent to that of $\rho$ at the origin, as $\psi$ is not just a conjugating automorphism but also fixes at the origin; this contradicts theorem 5.4 finishing the proof.
+But then the submersivity of $\tilde{\rho}$ at the origin is equivalent to that of $\rho$ at the origin, as $\psi$ is not just a conjugating automorphism but also fixes at the origin; this contradicts theorem 5.13, finishing the proof.
 ∎
 We next record a pair of bounded modifications of $H^{2}$ in the above examples wherein the some of the results hold as is, but some arguments cease to.
+More importantly, these give examples of bounded balanced non-convex (but pseudoconvex) domains wherein (i) for each point, there is a (non-trivial) retract passing through it and (ii) they are equivalent to product domains but contain an open piece of directions (an open subset of the unit sphere $\partial\mathbb{B}$ representing directions) along which retract exist.
 Example: Consider the domain $\Omega:=\Delta^{2}\cap H^{2}_{1/4}$, where $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}:|zw|<1/4\}$ (and $\Delta^{2}$ the standard bidisc, as usual).
@@
 However, we must check that $Z$ is a retract of $\Omega$, not just of $\Delta^{2}$.
-To do this of-course we need only consider the restriction $\rho|_{\Omega}$, which maps $\Omega$ into $Z$ and that would finish the proof provided we verify the following.
+To do this of-course we need only consider the restriction $\rho_{|_{\Omega}}$, which maps $\Omega$ into $Z$ and that would finish the proof provided we verify the following.
 Claim-2: $Z=L\cap\Omega$.
@@
 This finishes the proof of the claim.
-It now follows immediately that $\rho$ maps $\Omega$ into itself and ${\rm image}(\rho|_{\Omega})\subset L\cap\Omega=Z$; indeed, as we have for every $z\in Z=L\cap\Omega$ that $\rho(z)=z$, we have ${\rm image}(\rho|_{\Omega})=L\cap\Omega$.
+It now follows immediately that $\rho$ maps $\Omega$ into itself and ${\rm image}(\rho_{|_{\Omega}})\subset L\cap\Omega=Z$; indeed, as we have for every $z\in Z=L\cap\Omega$ that $\rho(z)=z$, we have ${\rm image}(\rho_{|_{\Omega}})=L\cap\Omega$.
 This finishes the proof completely that $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$ in the above case.
@@
 It remains to realize $W$ as a retraction of our smaller bounded $\Omega$, rather than the unbounded $H^{2}_{1/4}$.
-Again we need only consider the restriction $\varphi|_{\Omega}$ (which maps $\Omega$ into $W$), provided we verify the following.
+Again we need only consider the restriction $\varphi_{|_{\Omega}}$ (which maps $\Omega$ into $W$), provided we verify the following.
 Claim-3: $W=L\cap\Omega$.
@@
 As another difference from the unbounded $H^{2}$, thereby a purpose of the above $\Omega$, is that the sets: $S_{1}$ of unit vectors representing directions along which there are linear retracts in $\Omega$, as well as, $S_{2}$ of those unit vectors along which there are no non-trivial retracts through the origin, are both non-trivial subsets of the unit sphere neither of which have measure zero (and with both having non-empty interior in the sphere).
+Finally here, we also mention that there are non-linear retracts along the directions given by $S_{1}$ through the origin.
+Thus there exist bounded balanced pseudoconvex domains which are not biholomorphic to product domains but still have non-linear retracts through the origin.
+Indeed, we may establish $D:=\Delta^{2}\cap H^{1/2}$ is irreducible as follows.
+Assuming the contrary, we observe that it would follow by the Riemann mapping theorem that $D$ is biholomorphic to the bidisc $\Delta^{2}$.
+The homogeneity of $\Delta^{2}$ together with an application of Cartan’s theorem then renders a linear equivalence between $\Delta^{2}$ and $D$ which is impossible as the former is convex whereas the latter is not.
 We may ask for variants of the above $H^{2}_{1/4}\cap\Delta^{2}$ and more; we shall only record the case when the polydisc $\Delta^{2}$ is replaced by the Euclidean ball $\mathbb{B}$.
-Example: Consider the domain $\Omega:=\mathbb{B}\cap H^{2}_{1/4}$, where $\mathbb{B}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{2}+|w|^{2}<1\}$ be the euclidean ball centred at 00 of radius $1$ and $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}:|zw|<1/4\}$
+Example: Consider the domain $\Omega:=\mathbb{B}\cap H^{2}_{1/4}$, where $\mathbb{B}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{2}+|w|^{2}<1\}$ be the euclidean ball centred at 0 of radius $1$ and $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}:|zw|<1/4\}$ .
 We want to prove that for each $(a,b)\in\Omega$, there exist a retract (of dimension 1) of $\Omega$ passing through $(a,b)$.
-Let $L$ denotes the hyperplane $\{(z,w)\in\mathbb{C}^{2}:w=b\}$ Firstly note that
+Let $L$ denotes the hyperplane $\{(z,w)\in\mathbb{C}^{2}:w=b\}$ . Firstly note that
 $\partial\Omega=(\partial\mathbb{B}\cap H^{2}_{1/4})\cup(\mathbb{B}\cap\partial H^{2}_{1/4})\cup(\partial\mathbb{B}\cap\partial H^{2}_{1/4})$
-Note that $\partial(L\cap\Omega)=L\cap\partial\Omega$ First we prove the following claim:
+Note that $\partial(L\cap\Omega)=L\cap\partial\Omega$ . First we prove the following claim:
 Claim-1: Either $\partial(L\cap\Omega)\subset\partial\mathbb{B}$ or $\partial(L\cap\Omega)\subset\partial H^{2}_{1/4}$ for $i=1,2$.
-Proof of the claim-1: If possible assume that $\partial(L\cap\Omega)\nsubseteq\partial\mathbb{B}$ and $\partial(L\cap\Omega)\nsubseteq\partial H^{2}_{1/4}$ Then there exist points $(a_{1},b)$ and $(b_{1},b)$ in $\partial(L\cap\Omega)$ such that $(a_{1},b)\notin\partial\mathbb{B}$ and $(b_{1},b)\notin\partial H^{2}_{1/4}$.
+Proof of the claim-1: If possible assume that $\partial(L\cap\Omega)\nsubseteq\partial\mathbb{B}$ and $\partial(L\cap\Omega)\nsubseteq\partial H^{2}_{1/4}$ .
+Then there exist points $(a_{1},b)$ and $(b_{1},b)$ in $\partial(L\cap\Omega)$ such that $(a_{1},b)\notin\partial\mathbb{B}$ and $(b_{1},b)\notin\partial H^{2}_{1/4}$.
 From the above decomposition of $\partial\Omega$ and $(a_{1},b),(b_{1},b)\in\partial\Omega$, we obtain that $(a_{1},b)\in\mathbb{B}\cap\partial H^{2}_{1/4}$ and $(b_{1},b)\in\partial\mathbb{B}\cap H^{2}_{1/4}$.
@@
 Dividing $|a_{1}b|$ (which is equal to $1/4$) on both sides of the equation $|b_{1}b|<1/4$, we get $|b_{1}|<|a_{1}|$.
-Squaring and adding $|b|^{2}$ on both sides of the equation $|b_{1}|<|a_{1}|$, we get that $11=|b_{1}|^{2}+|b|^{2}<|a_{1}|^{2}+|b|^{2}<$ 1, which is vividly a contradiction.
+Squaring and adding $|b|^{2}$ on both sides of the equation $|b_{1}|<|a_{1}|$, we get that $1=|b_{1}|^{2}+|b|^{2}<|a_{1}|^{2}+|b|^{2}<1$, which is vividly a contradiction.
 This finishes the proof of the claim.
@@
 Now we prove that $Z$ is a retract of $\Omega$.
-To prove this, consider the map $\rho|_{\Omega}$, which maps $\Omega$ into $Z$.
+To prove this, consider the map $\rho_{|_{\Omega}}$, which maps $\Omega$ into $Z$.
 Next we prove the following claim:
@@
 This finishes the proof of the claim.
-Using the above claim-2, $\rho$ maps $\Omega$ into itself and $image(\rho|_{\Omega})\subset L\cap\Omega=Z$.
+Using the above claim-2, $\rho$ maps $\Omega$ into itself and $image(\rho_{|_{\Omega}})\subset L\cap\Omega=Z$.
 If $z\in Z=L\cap\Omega$, then $\rho(z)=z$.
-So $image(\rho|_{\Omega})=L\cap\Omega$.
+So $image(\rho_{|_{\Omega}})=L\cap\Omega$.
 Hence $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$.
@@
 Firstly note that $b\neq 0$ because $(a^{\prime},b)\in\partial H^{2}_{1/4}$. Consider the holomorphic map $\varphi:H^{2}_{1/4}\to H^{2}_{1/4}$ defined by $\varphi(z,w)=(\frac{1}{b}zw,b)$.
-Note that $\varphi$ is a retraction map on $H^{2}_{1/4}$. Now we consider the map $\varphi|_{\Omega}$, which maps $\Omega$ into $W$. Next we prove the following claim:
+Note that $\varphi$ is a retraction map on $H^{2}_{1/4}$. Now we consider the map $\varphi_{|_{\Omega}}$, which maps $\Omega$ into $W$. Next we prove the following claim:
 Claim-3: $W=L\cap\Omega$.
 Proof of claim-3: First we prove that $|a^{\prime}|\geq\frac{1}{4|b|}$.
-If $|a^{\prime}|<\frac{1}{4|b|}$ then $|a^{\prime}b|<1/4$ and thus implies that $(a^{\prime},b)\in H^{2}_{1/4}$, which contradicts the fact that $(a^{\prime},b)\in\partial H^{2}_{1/4}$ Hence we proved that $|a^{\prime}|\geq\frac{1}{4|b|}$
+If $|a^{\prime}|<\frac{1}{4|b|}$ then $|a^{\prime}b|<1/4$ and thus implies that $(a^{\prime},b)\in H^{2}_{1/4}$, which contradicts the fact that $(a^{\prime},b)\in\partial H^{2}_{1/4}$ .
+Hence we proved that $|a^{\prime}|\geq\frac{1}{4|b|}$
 If $(z,b)\in W$, then $|z|<\frac{1}{4|b|}$. Note that
@@
 The converse is the famous theorem of Lempert (exposited as theorem 11.2.1 in [Jrncki_invrnt_dst]). ∎
-6. Proof of theorem 1.10
+6. Proof of theorems 1.12 and 1.14
 We begin by connecting the notions of convexity as used in the statement of the theorem, usual geometric convexity and continuity of the Minkowski functional – throughout this section $D$ will stand for a bounded balanced domain of holomorphy in $\mathbb{C}^{N}$ without further mention; we may only mention additional assumptions if made.
@@
 Proof.
-Let us firstly quickly note that if $D$ is convex near $p$ is the sense of definition 2.30, then $D$ is geometrically convex near $p$ i.e., there exist an open ball $U$ in $\mathbb{C}^{N}$ centered at $p$ such that $U\cap D$ is convex in the usual sense of being closed under the formation of line segments.
+Let us firstly quickly note that if $D$ is convex near $p$ in the sense of definition 2.34, then $D$ is geometrically convex near $p$ i.e., there exist an open ball $U$ in $\mathbb{C}^{N}$ centered at $p$ such that $U\cap D$ is convex in the usual sense of being closed under the formation of line segments.
 Indeed, this follows by considering the closed half-spaces (containing the domain) given by the linear functionals as per definition 2.34 due to the points of $\partial D\cap U$ being convex boundary points; supporting hyperplanes (and likewise half-spaces containing $D\cap U$) at other points of $\partial(D\cap U)$ exist just by convexity of the ball $U$.
 As the intersection of these closed half-spaces precisely equals $\overline{U\cap D}$, the convexity of $U\cap D$ follows.
-Next, we show the subadditivity of $h$.
-Let $z,w\in U\cap D$.
-By definition of the Minkowski functional, there exist positive reals $t,s$ such that $z/t,w/s\in U\cap D$ and $t<h(z)+\epsilon$, $s<h(w)+\epsilon$ for some $\epsilon>0$.
-Writing $(z+w)/(s+t)$ as the convex combination:
-$\frac{z+w}{s+t}=\frac{t}{s+t}\cdot\frac{z}{t}+\frac{s}{s+t}\cdot\frac{w}{s},$
-it follows that by the convexity of $U\cap D$, that $(z+w)/(s+t)\in U\cap D$.
-Hence $h(z+w)\leq s+t<h(z)+h(w)+2\epsilon$.
-As $\epsilon$ is arbitrary, we get $h(z+w)\leq h(z)+h(w)$ for all $z,w\in U\cap D$, finishing the verification of the required subadditivity.
-We now proceed to the original goal about continuity of $h$.
-Let $q$ be arbitrary point in $U\cap\partial D$ and $(z_{n})$ be a sequence converging to $q\in U\cap\partial D$.
-(without loss of generality, we assume that $z_{n}\in U\cap D$ for all $n$).
-By subadditivity of $h$ and the well-known basic fact that the Minkowski functional is dominated by some scaling of the Euclidean norm, we have $|h(z_{n})-h(q)|\leq h(z_{n}-q)\leq c\|z_{n}-q\|$ for some $c>0$.
-The convergence of $z_{n}$ to $q$, thereby readily implies that $h(z_{n})$ converges to $h(q)$.
-Hence $h$ is continuous near $p$, finishing the proof owing to what was noted earlier.
+Let $C:=\bigcup_{z\in\partial D\cap U}\{\lambda z:\lambda\geq 0\}$ and $C_{D}:=C\cap D$.
+As $U\cap D$ is geometrically convex, so is the cone generated by it, thereby assuring that $C_{D}$ is convex.
+Next, we show the subadditivity of $h$ on $C$.
+Let $z,w\in C_{D}$.
+By definition of the Minkowski functional, for each $\epsilon>0$ there exist positive reals $t,s$ such that $z/t,w/s\in D$ and $t<h(z)+\epsilon$, $s<h(w)+\epsilon$.
+Since $C_{D}$ is open, there exist $t^{\prime},s^{\prime}$ such that $t^{\prime}<t,~s^{\prime}<s$ and $z/t^{\prime},w/s^{\prime}\in C_{D}$.
+Writing $(z+w)/(s^{\prime}+t^{\prime})$ as the convex combination:
+$\frac{z+w}{s^{\prime}+t^{\prime}}=\frac{t^{\prime}}{s^{\prime}+t^{\prime}}\cdot\frac{z}{t^{\prime}}+\frac{s^{\prime}}{s^{\prime}+t^{\prime}}\cdot\frac{w}{s^{\prime}},$
+it follows that by the convexity of $C_{D}$, that $(z+w)/(s+t)\in C_{D}$.
+Hence $h(z+w)\leq s^{\prime}+t^{\prime}<h(z)+h(w)+2\epsilon$.
+As $\epsilon$ is arbitrary, we get $h(z+w)\leq h(z)+h(w)$ for all $z,w\in C_{D}$.
+This shows that $h$ is sub-additive on $C_{D}$.
+To prove the subadditivity of $h$ on $C$, let $z,w$ be any two points in $C$.
+Choose $k$ so that $k>\max\{h(z),h(w)\}$.
+Note that $z/k,w/k\in C_{D}$.
+Applying the subadditivity of $h$ on $C_{D}$ and homogenity of $h$, we obtain that $h(z/k+w/k)\leq 1/k~h(z)+1/k~h(w)$.
+This shows that $h$ is subadditive on $C$.
+The subadditivity of $h$ on $C$ and together with homogenity of $h$ yields that $h$ is a convex function on $C$.
+Hence $h$ is continuous on $C$ and in particular $h$ is continuous near $p$.
 ∎
@@
 Consider the holomorphic map $\varphi:\Delta\to\mathbb{C}^{N}$ defined by $\varphi(t)=(p+tu)t$.
-As $p+tu\in\partial D$ for all $t\in\Delta$ and $h$ is continuous at $p+tu$ for all $t\in\Delta$, we have by a well-known fundamental fact (see for instance Remark 2.2.1 (h) of [Jrncki_invrnt_dst]) that $h(p+tu)=1$ for all $t\in\Delta$.
+As $p+tu\in\partial D$ for all $t\in\Delta$ and $h$ is continuous at $p+tu$ for all $t\in\Delta$, we have by a well-known fundamental fact (see proposition 2.9 (d)) that $h(p+tu)=1$ for all $t\in\Delta$.
 Hence
@@
 So, $\varphi$ maps $\Delta$ into $D$.
-Next we want to prove that $\varphi$ is a complex geodesic in $D$.
+Next we want to prove that $\varphi$ is a complex $C$-geodesic in $D$.
 First we claim that there exist $t_{0}\in\Delta^{*}:=\Delta\setminus\{0\}$ such that $D$ is convex at $p+t_{0}u$.
@@
 (6.1) $|L(z)|\leq h(z)~~\text{and}~~|L(p+t_{0}u)|=h(p+t_{0}u)=1.$
-Since $\mathbb{C}^{N}=Ker(L)\oplus span\{p\}$, we can write $u=u_{1}+u_{2}$, where $u_{1}\in Ker(L)$ and $u_{2}\in span\{p\}$. Write $u_{2}=\lambda p$ for some $\lambda\in\mathbb{C}$. So
+Since $\mathbb{C}^{N}=\ker(L)\oplus span\{p\}$, we can write $u=u_{1}+u_{2}$, where $u_{1}\in\ker(L)$ and $u_{2}\in span\{p\}$. Write $u_{2}=\lambda p$ for some $\lambda\in\mathbb{C}$. So
 $|L(p+tu)|=|L(p)+tL(u)|=|L(p)+t\left(L(u_{1})+L(u_{2})\right)|=|1+\lambda t||L(p)|=|1+\lambda t|.$
 We need only prove that there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$, which we do now.
-If $\lambda=0$, then $u\in Ker(L)$.
+If $\lambda=0$, then $u\in\ker(L)$.
 Hence (6.1) holds for any $t_{0}\in\Delta^{*}$.
@@
 $|re^{i\theta}+1|=1~~\text{ and }~~~r<|\lambda|.$
-Geometrically above equation represents the intersection of circle centred at $-1$ with radius $1$ and the disc centred at 00 of radius $|\lambda|$.
+Geometrically above equation represents the intersection of circle centred at 1-1 with radius $1$ and the disc centred at 0 of radius $|\lambda|$.
 So there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$.
@@
 The arguments above can actually be used to show that $\partial D$ is convex at all points near $p$ in the complex line segment through $p$ in the direction of $u$.
-To proceed further with the proof that $\varphi$ is a complex $C$-geodesic, let us now note that as $D$ is convex at the boundary point $p+t_{0}u$ by the foregoing, we have by proposition 2.3.1(b) in [Jrncki_invrnt_dst], with the same notations therein, that
+To proceed further with the proof that $\varphi$ is a complex $C$-geodesic, let us now note that as $D$ is convex at the boundary point $p+t_{0}u$ by the foregoing; thereby also at $(t_{0}/|t_{0}|)(p+t_{0}u)$.
+We then have by proposition 2.3.1(b) in [Jrncki_invrnt_dst], with the same notations therein, that
 $c_{D}^{*}(0,(p+t_{0}u)t_{0})=h((p+t_{0}u)t_{0})=|t_{0}|h(p+t_{0}u)=|t_{0}|;$
@@
 $h(p)=|\lambda|h(u).$
-By the hypotheses that $D$ is bounded and from remark 2.2.1 (l) in [Jrncki_invrnt_dst], we obtain that $h(u)\neq 0$.
-Hence we get $|\lambda|=1/h(u)$.
-We now split into two cases depend upon whether $h(u)>1$ or not.
+By the hypotheses that $D$ is bounded and from proposition 2.9 (f), we obtain that $h(u)\neq 0$. Hence we get $|\lambda|=1/h(u)$. We now split into two cases depend upon whether $h(u)>1$ or not.
 If $h(u)\leq 1$, then $p+\frac{1}{2\overline{\lambda}}u\in\partial D$. Since $p=\lambda u$, we have
@@
 This establishes the linear independence of $p$ and $u$.
-Now we get back to finishing the proof that $Z$ is linear, to which end, we get back to the equation (6.3).
+Now we get back to finishing the proof that $Z$ is non-linear, to which end, we get back to the equation (6.3).
 Since $p,u$ are linearly independent, we obtain in-particular, $e^{i2\theta}/2=1$.
 As such a $\theta$ does not exist, this concludes our proof by contradiction to establish that $Z$ is a non-linear retract of $D$.
-∎
+An alternative proof of this theorem is provided below without the use of complex geodesics.
+First, we consider the case that $\partial D$ is convex near $p$.
+As $\partial D$ is not $\mathbb{C}$-extremal at $p$, there exists a non-zero holomorphic map $\psi:\Delta\to F_{\mathbb{C}}(p,\partial D)$.
+After rescaling $\Delta$, if needed, we may assume that $\psi(t)$ is a non-vanishing holomorphic map on $\Delta^{*}$.
+We want to show that there exists $\delta>0$ such that $p+t\psi(t)\in\partial D$ for all $t\in\Delta_{\delta}$, where $\Delta_{\delta}$ denotes the disk centred at the origin of radius $\delta$.
+Now, we claim that there exists $\delta>0$ such that, for every $u\in F_{\mathbb{C}}(p,\partial D)\cap S$, we have $p+su\in\partial D$ for all $s\in\Delta_{\delta}$, where $S=\partial\mathbb{B}_{\ell_{1}}$ denote the unit sphere in the $\ell_{1}$ norm.
+To prove this, let $\text{co}(D)$ denote the convex hull of $D$.
+Note that $F_{\mathbb{C}}(p,\partial D)=F_{\mathbb{C}}(p,\partial(\text{co}(D)))$ because $D$ is convex near $p$.
+For any $u\in F_{\mathbb{C}}(p,\partial(\text{co}(D)))\cap S$ can be written as a convex combination of vectors $u_{j}\in F_{\mathbb{C}}(p,\partial(\text{co}(D)))\cap S$.
+i.e.
+$u=c_{1}u_{1}+\ldots+c_{k}u_{k}$, where $c_{j}$’s are non-negative reals with $c_{1}+c_{2}+\ldots+c_{k}=1$.
+Since each $u_{j}\in F_{\mathbb{C}}(p,\partial(\text{co}(D)))$, there exists $\epsilon_{j}>0$ such that $p+su_{j}\in\partial\text{co}(D)$ for all $s\in\Delta_{\epsilon_{j}}$.
+Let $\delta:=\min\{\epsilon_{1},\ldots,\epsilon_{k}\}$.
+Then for each $s\in\Delta_{\delta}$,
+$p+su=p+s(c_{1}u_{1}+\ldots+c_{k}u_{k})=c_{1}(p+su_{1})+\ldots+c_{k}(p+su_{k})$
+Each term $p+su_{j}\in\overline{\text{co}(D)}$, and since $\overline{\text{co}(D)}$ is convex, it follows that $p+su\in\overline{\text{co}(D)}$ for all $s\in\Delta_{\delta}$.
+Since $D$ is convex near $p$, we may assume that $D$ is convex at $p+su$ for all $s\in\Delta_{\delta}$.
+By proposition 2.3.1(b) in [Jrncki_invrnt_dst], we have $\gamma_{D}(0;p+su)=h(p+su)$.
+Applying proposition 2.3.1(d) in [Jrncki_invrnt_dst], we obtain that $h(p+su)=\widehat{h}(p+su)\leq 1$, where $\widehat{h}$ is the Minkowski function of $\text{co}(D)$.
+This implies that $p+su\in\overline{D}$ for all $s\in\Delta_{\delta}$.
+From the proof of proposition 2.17, we obtain that $p+su\in\partial D$ for all $s\in\Delta_{\delta}$, which finishes the proof of the claim.
+Since $F_{\mathbb{C}}(p,\partial D)$ is closed under scalar multiplication, for each $t\in\Delta^{*}$, we have
+$\frac{\psi(t)}{\|\psi(t)\|}\in F_{\mathbb{C}}(p,\partial D)\cap S.$
+Since $\phi(t)=t\|\psi(t)\|$ is continuous on $\Delta^{*}$, there exists $\delta_{0}>0$ such that $\|t\psi(t)\|\leq\delta$ for all $t\in\Delta_{\delta_{0}}$.
+Hence by the above claim, for all $t\in\Delta_{\delta_{0}}^{*}$, we have
+$p+t\psi(t)=p+t\|\psi(t)\|\left(\frac{\psi(t)}{\|\psi(t)\|}\right)\in\partial D$
+Define the holomorphic map $\tilde{\varphi}:\Delta\to\mathbb{C}^{N}$ by $\tilde{\varphi}(t)=p+\delta t\psi(\delta t)$.
+Now, we consider the map $\varphi:\Delta\to D$ defined by $\varphi(t)=t\tilde{\varphi}(t)$.
+Hence $Z:=\text{image}(\varphi)$ is the graph of the holomorphic map $\phi(t)=\delta t^{2}\psi(\delta t)$ over the linear subspace $\operatorname{span}(p)\cap D$ of $D$.
+By proposition 3.14, it follows that $Z$ is a retract of $D$.
+Next, we assume the weaker and general assumption that $\partial D$ is convex at $p$, together with the Minkowski functional of $D$ being continuous near $p$.
+Choose any holomorphic map $\tilde{\varphi}:\Delta\to(p+F_{\mathbb{C}}(p,\partial D))\cap\partial D$ , since $\partial D$ is not $\mathbb{C}$-extremal at $p$, there exists atleast one such holomorphic map $\tilde{\varphi}$.
+So, we write $\tilde{\varphi}(t)=p+\tilde{\psi}(t)$, where $\tilde{\psi}\in\mathcal{O}(\Delta)$ with $\tilde{\psi}(0)=0$ and $\tilde{\psi}(t)\in F_{\mathbb{C}}(p,\partial D)$ for all $t\in\Delta$.
+Now, we consider the holomorphic map $\varphi:\Delta\to\mathbb{C}^{N}$ defined by $\varphi(t)=t\tilde{\varphi}(t)$.
+In the above proof, we showed that $\varphi$ maps $\Delta$ into $D$.
+Note that $\text{span}(p)\cap\text{image}(\tilde{\psi})=\{0\}$ because if $u$ is a non-zero vector in $F_{\mathbb{C}}(p,\partial D)$ then $p$ and $u$ are linearly independent (as shown in the above proof; look at the paragraph following equation(6.3)).
+This implies that $Z=\text{image}(\varphi)$ is a graph of the holomorphic map $\phi(t)=t\tilde{\psi}(t)$ over the linear retract $\text{span}(p)\cap D$ of $D$.
+By theorem 3.14, it follows that $Z$ is a retract of $D$.
 Remark 6.3.
@@
 The next proposition establishes that this exhausts all possible examples of non-linear (one-dimensional) retracts through the origin in the direction of a boundary point that fails to be $\mathbb{C}$-extremal.
-Indeed, this proposition is an expanded version of theorem 1.12 whose proof will thereby get finished here.
+Indeed, this proposition is an important special case of theorem 1.14 whose proof will be based on this and we therefore deal with it first.
+Moreover, this gives finer information about the one-dimensional retracts than theorem 1.14, which is an assertion about retracts of all dimensions.
 Before we get to the details, it is very helpful to introduce a notation for brevity: if $a=(a_{1},\ldots,a_{k})$ is any point in $\mathbb{C}^{k}$ (for some $k\in\mathbb{N}$) and each of $v_{1},v_{2},\ldots,v_{k}$ are vectors in $\mathbb{C}^{N}$, then we define $a\cdot v=a_{1}v_{1}+a_{2}v_{2}+\ldots+a_{k}v_{k}$; indeed this is the ‘$a$-linear’ combination of the vectors $v_{1},\ldots,v_{k}$.
@@
 Suppose that $\partial D$ is convex at $p\in\partial D$ but not $\mathbb{C}$-extremal at $p$.
-Let $Z$ be an arbitrary one dimensional retract of $D$ passing through 00 in the direction of $p$; let $\rho:D\to Z$ be a retraction map.
-Then, $Z$ can be be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\tilde{\psi}(t)\cdot v$ where $\tilde{\psi}:\Delta\to\mathbb{C}^{N-1}$ is an analytic disc with $\tilde{\psi}(0)=0$ and where $v$ stands for a list of basis vectors $(v_{1},\ldots,v_{N})$ of $\ker\left(D\rho(0)\right)$.
+Let $Z$ be an arbitrary one-dimensional retract of $D$ passing through 0 in the direction of $p$; let $\rho:D\to Z$ be a retraction map.
+Then, $Z$ can be be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\tilde{\psi}(t)\cdot v$ where $\tilde{\psi}:\Delta\to\mathbb{C}^{N-1}$ is an analytic disc with $\tilde{\psi}(0)=0$ and where $v$ stands for a list of basis vectors $(v_{1},\ldots,v_{N-1})$ of $\ker\left(D\rho(0)\right)$.
 Also, after factoring out $t$, the resulting map $t\mapsto p+\tilde{\psi}(t)\cdot v$ gives an analytic disc in the boundary $\partial D$; indeed, this analytic disc is the radial projection of $Z$ onto the boundary of our balanced domain $D$.
@@
 $\psi(t)\cdot v\;=\;\psi_{1}(t)v_{1}+\ldots+\psi_{N-1}(t)v_{N-1}$
-with $v_{1},\ldots,v_{N-1}$ being a basis for the $(N-1)$-dimensional subspace $M:=\ker\left(D\rho(0)\right)$.
-Note that $\psi(0)=0$ and $\psi^{\prime}(0)=0$ because $Z$ passes through 00 in the direction of $p$ and $p,v_{1},v_{2},...,v_{N-1}$ are linearly independent.
+with $v_{1},\ldots,v_{N-1}$ being a basis for the (N1)(N-1)-dimensional subspace $M:=\ker\left(D\rho(0)\right)$.
+Note that $\psi(0)=0$ and $\psi^{\prime}(0)=0$ because $Z$ passes through 0 in the direction of $p$ and $p,v_{1},v_{2},...,v_{N-1}$ are linearly independent.
 Note that we can write $\psi(t)=t\widetilde{\psi}(t)$, for some holomorphic map $\widetilde{\psi}$ with $\widetilde{\psi}(0)=0$.
@@
 Since $\partial D$ is convex near $p$ and $\widetilde{\varphi}(0)=p$, there exist a smaller subdisc $\Delta^{\prime}$ of $\Delta$ such that $\partial D$ is convex near $\widetilde{\varphi}(t)$ for all $t\in\Delta^{\prime}$.
-Observe that that $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}$ is a harmonic function on $\Delta^{\prime}$ and $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}\leq 0$ on $\Delta^{\prime}$.
-Since $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}(0)=0$, by the maximum principle we obtain $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}\equiv 0$ on $\Delta^{\prime}$ Consider the holomorphic function $F:\Delta\to\mathbb{C}$ defined by
+Observe that that $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}$ is a harmonic function on $\Delta^{\prime}$ and $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}\leq 0$ on $\Delta^{\prime}$.
+Since $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}(0)=0$, by the maximum principle we obtain $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}\equiv 0$ on $\Delta^{\prime}$ .
+Consider the holomorphic function $F:\Delta\to\mathbb{C}$ defined by
 $F(t)=\sum_{j=1}^{N}\left((a_{j}-ib_{j})\widetilde{\varphi}_{j}(t)-\alpha\right),$
 where $\widetilde{\varphi}_{j}$ denotes the $j^{\text{th}}$ component of $\widetilde{\varphi}$.
-Since $\Phi\circ\widetilde{\varphi}|_{\Delta^{\prime}}$ vanishes on $\Delta^{\prime}$, ${\rm Re}(F)$ vanishes on $\Delta^{\prime}$.
+Since $\Phi\circ\widetilde{\varphi}_{|_{\Delta^{\prime}}}$ vanishes on $\Delta^{\prime}$, ${\rm Re}(F)$ vanishes on $\Delta^{\prime}$.
 By the open mapping theorem applied to $F$, we deduce that $F$ vanishes on $\Delta^{\prime}$.
@@
 Remark 6.5.
-In this claim, we obtain that $F_{\mathbb{C}}(p,\partial D)+p$ is contained in every complex supporting hyperplane to $\partial D$ at $p$ i.e., every complex supporting hyperplane to the boundary at the point $p$ has got to contain the copy of the face $F_{\mathbb{C}}(p,\partial D)$ translated to $p$.
-Continuing with the proof of the proposition 6.4, write $\widetilde{\varphi}(t)=p+\widetilde{\phi}(t)$, where $\widetilde{\phi}:\Delta\to\mathbb{C}^{N}$ is the holomorphic map given by $\widetilde{\phi}(t)=\widetilde{\psi}(t)\cdot v$.
-Let $L_{p}$ denote the smallest $\mathbb{R}$-linear subspace of $\mathbb{C}^{N}$ containing $image({\widetilde{\phi}})$ ($L_{p}$ is of-course uniquely determined as the intersection of all such $\mathbb{R}$-linear subspaces).
-We now prove that $L_{p}$ is a complex vector space.
-Note that $L_{p}$ is the solution of some (finite) system of linear equations with real coefficients, where each linear equation can be written in the following form
+As a consequence of this claim, it follows that $F_{\mathbb{C}}(p,\partial D)+p$ is contained in every complex supporting hyperplane to $\partial D$ at $p$ i.e., every complex supporting hyperplane to the boundary at the point $p$ has got to contain the copy of the face $F_{\mathbb{C}}(p,\partial D)$ translated to $p$.
+[…]
 $\sum_{j=1}^{N}a_{j}{\rm Re}(z_{j})+\sum_{j=1}^{N}b_{j}{\rm Im}(z_{j})=0.$
-Since $image(\widetilde{\phi})$ is contained in $L_{p}$, we have
+[…]
 $\sum_{j=1}^{N}a_{j}{\rm Re}(\widetilde{\phi}_{j})+\sum_{j=1}^{N}b_{j}{\rm Im}(\widetilde{\phi}_{j})=0,$
@@
 $\mathbb{C}^{N}\;=\;{\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)\oplus\tilde{V}$,
-where $\tilde{V}$ is a complementary subspace to ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$.
-If $\pi$ denotes the projection of $\mathbb{C}^{N}$ onto ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$, then note that $\pi(Z)=Z$.
-Indeed, this is just a re-statement of the fact that $Z$ is contained within ${\rm span}(p)\oplus F_{\mathbb{C}}(p,\partial D)$ i.e., has no $\tilde{V}$-component.
-From the fact that $Z=image(\widetilde{\varphi})$ and the expression of $\widetilde{\varphi}(t)$, it is now easy to see that $Z$ is the graph of the holomorphic map on ${\rm span}(p)$ given by $tp\to t\tilde{\psi}\cdot v$; where $\tilde{\psi}(t)\cdot v$ is a holomorphic map from $\Delta$ into the face $F_{\mathbb{C}}(p,\partial D)$ with the property that $p+\tilde{\psi}(t)\cdot v$ is contained within the connected component of
+[…]
 $\left(p+F_{\mathbb{C}}(p,\partial D)\right)\cap\partial D;$
@@
 Noting that this connected component is contained within the affine linear projection of $D$ into $p+F_{\mathbb{C}}(p,\partial D)$, we are led to the last statement of the proposition; translating this projection back to the origin, gives a bounded balanced pseudoconvex domain which is a projection of $D$ into the lower dimensional complex linear subspace given by the face $F_{\mathbb{C}}(p,\partial D)$.
-This finishes the proof in all, both of theorem 1.12 as well as the assertions of this proposition 6.4.
-∎
+This finishes the proof of the proposition (which was for one-dimensional retracts).
+Using the foregoing proposition and the arguments in its proof, it is possible to retracts of arbitrary dimension but we do this at the end of this section, after the following remarks and examples.
 Remark 6.6.
-When $N=2$, we may rephrase the conclusion above more precisely as follows: every non-linear retract $Z$ passing through the origin in the direction of $p$ can be realized as the image of an analytic disc of the form $\varphi(t)=tp+t\widetilde{\psi}(t)u$, where $u$ belongs to the face $F_{\mathbb{C}}(p,\partial D)$ and $\widetilde{\psi}$ maps $\Delta$ holomorphically onto bounded domain in $\mathbb{C}$ containing the origin, with $\psi(0)=0$ and $\psi^{\prime}(0)=0$.
-The proof of this is as follows.
-First observe that from the proof of above proposition, we obtain there exist non-zero $v\in\mathbb{C}^{N}$ such that $p+\widetilde{\psi}(t)v\in\partial D$ for all $t\in\Delta$.
-Since $Z$ is non-linear, $\widetilde{\psi}$ is not identically zero.
-By continuity of $\widetilde{\psi}$ and $\widetilde{\psi}(0)=0$, we get $v\in F_{\mathbb{C}}(p,\partial D)$.
-It follows from the connectedness of image of $\widetilde{\psi}$, that $\widetilde{\psi}$ maps $\Delta$ into $\widetilde{\Delta}$.
-By the open mapping theorem, $\widetilde{\psi}$ maps $\Delta$ onto an open subset $\widetilde{\psi}(\Delta)$ within ${\rm int}(\widetilde{\Delta})$.
+[…]
 Remark 6.7.
@@
 where $B$ is a strictly convex ball such as $l^{p}$ ball with $p\neq 1,\infty$ and $c=1$.
-In-fact, we can also let $B$ be a non-convex balanced domain of holomorphy such as the $\ell_{q}$-‘ball’ for $10<q<$ 1 (in which case we must choose $c$ appropriately for instance $c=1/2$) to obtain more concrete examples.
-Let us now conclude this section with a couple of remarks.
+In-fact, we can also let $B$ be a non-convex balanced domain of holomorphy such as the $\ell_{q}$-‘ball’ for $0<q<1$ (in which case we must choose $c$ appropriately for instance $c=1/2$) to obtain more concrete examples.
+Let us now conclude this chapter with a couple of remarks and more importantly, proof of theorem 1.14.
 We formulate the first one as the following corollary to highlight that: as a consequence of lemma 6.1 and our proof of theorem 1.12 immediately verifies the truth of that theorem in an important special case.
@@
     $D$ is neither convex at $p$ nor near $p$. In this case, there are no retracts of $D$ passing through origin in the direction of $p$.
+As mentioned above, we shall now use the foregoing proposition 6.4 and the arguments in its proof, to extend it to retracts of arbitrary dimension as stated in theorem 1.14, which we now prove.
+Proof of theorem 1.14: Observe that we may apply (the already proven) theorem 1.15 which applies to the setting of the theorem being proved.
+We can therefore express $Z$ as a graph of a holomorphic map $F$ from $D_{L}$ into $D_{M}$, where $M=\ker(D\rho_{|_{0}})$; notice that $\mathbb{C}^{N}=L\oplus M$.
+We then pick any point $p$ in $L\cap\partial D$ and consider the retract over the segment spanned by $p$ i.e., let $Z_{p}$ denote the graph of $F$ over $D_{p}:=D\cap{\rm span}(p)$ which is a linear retract precisely by virtue of the convexity hypothesis at $p$.
+This makes it clear that $Z_{p}$ can be expressed as $\{tp+F(tp):t\in\Delta\}$ and also that $Z_{p}$ is a retract.
+The foregoing proposition and the arguments in its proof are therefore applicable to this one-dimensional retract $Z_{p}$; we thereby get that $F_{|_{D_{p}}}$ takes values within the complex face $F_{\mathbb{C}}(p,\partial D)$.
+Viewing $L$ as the union of one-dimensional subspaces, we obtain that $F$ takes values in $\bigcup_{p\in\partial D}F_{\mathbb{C}}(p,\partial D)$.
+∎
 7. Proof of theorem 1.19
-Let $\pi:=D\rho|_{\rho(0)}$ and $L:=T_{p}Z$. Since $\rho$ is a submersion at the origin, $D\rho_{|_{0}}:\mathbb{C}^{N}\longrightarrow T_{p}Z$ is onto. Next, write
-$\mathbb{C}^{N}=Ker(D\rho_{|_{0}})\oplus V_{p},$
-where $V_{p}$ is some complemented linear subspace of $Ker(D\rho_{|_{0}})$.
-Now, the submersivity condition means that $D\rho_{|_{0}}$ when restricted to $V_{p}$ maps $V_{p}$ onto $T_{p}Z$.
-Since ${\rm dim}(V_{p})=\rm dim(T_{p}Z)$, it follows that $D\rho_{|_{0}}:V_{p}\longrightarrow T_{p}Z$ is a bijective linear map i.e., a linear isomorphism.
-Now, consider the map
-$G=(D\rho_{|_{0}})^{-1}\circ\pi\circ\rho:D_{L}\longrightarrow V_{p}.$
-Note then that
-$DG_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}\circ\pi\circ D\rho_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}\circ\pi(D\rho_{|_{0}}(v)).$
-Further, note that $\pi_{|_{T_{p}Z}}$ is just the identity map.
-So we get $DG_{|_{0}}(v)=(D\rho_{|_{0}})^{-1}(D\rho_{|_{0}}(v))=v$.
-We then get by Mazet’s Schwarz lemma that $G$ acts as just the identity map on $D_{L}$.
-Note also that
-$(D\rho_{|_{0}})^{-1}\circ\pi\circ\rho\equiv I_{D_{L}}.$
-Therefore, $Z={\rm image}(\rho)$ is the graph of holomorphic function over some linear subspace $D_{L}$ of $D$, finishing the proof.
+Consider the holomorphic map $G:D_{L}\to\mathbb{C}^{N}$ defined by the composition of the restricted mappings:
+$G(z)=D\rho_{|_{0}}\circ\rho_{|_{D_{L}}}(z).$
+Note that $G$ actually maps into $D_{L}$: as $\rho$ maps $D$ into itself and $D\rho_{|_{0}}$ is a linear projection mapping $D$ onto $D_{L}$, it follows that $G$ maps $D_{L}$ into $D_{L}$.
+Note that for all $z\in D_{L}$, we have
+$DG_{|_{0}}(z)=D\rho_{|_{0}}\circ D\rho_{|_{0}}(z)=D\rho_{|_{0}}(z)=z.$
+By the hypotheses that $D_{L}$ is convex, we can choose $v\in D_{L}$ such that $v$ is an $\mathbb{R}$-extremal boundary point in $D_{L}$.
+Note that $DG_{|_{0}}(v)=v$.
+By Mazet Schwarz lemma 3.5 applied to $G:D_{L}\to D_{L}$, we obtain that $G(z)=z$ for all $z\in D_{L}$.
+This means that $D\rho_{|_{0}}\circ\rho_{|_{D_{L}}}$ is the identity map on $D_{L}$.
+Hence we conclude that $Z$ is a graph of a holomorphic function over $D_{L}$, finishing the proof of theorem 1.19.
 We ought to make a few remarks to settle any first questions that may arise as to whether the result is vacuous(!) i.e., whether there exist retracts not passing through the origin at all.
-For this let us only recall Lempert’s results which guarantee the existence of one-dimensional retracts $Z$ through any given pair of points $p,q$ which we take to be linearly independent; if $Z$ passes through origin, then the strict convexity hypotheses here leads by Vesentini’s theorem to the conclusion that $Z$ is linear.
-But then the linear independence of $p,q$ means that ${\rm dim}(Z)>1$; this contradiction finishes the verification about the existence of many retracts not passing through the origin, the matter of the above theorem.
+For this let us only recall Lempert’s results which guarantee the existence of one-dimensional retracts $Z$ through any given pair of points $p,q$ which we take to be linearly independent.
+Such a $Z$ cannot pass through origin, for otherwise, the strict convexity hypotheses here leads by Vesentini’s theorem to the conclusion that $Z$ is linear.
+But then the linear independence of $p,q$ means that ${\rm dim}(Z)>1$, a contradiction; this finishes the verification about the existence of many retracts not passing through the origin, the matter of the above theorem.
 This discussion also answers another basic question: are there examples of a bounded balanced domain of holomorphy $D$, wherein there exist pairs of one-dimensional retracts which are not ${\rm Aut}(D)$-equivalent i.e., there is no automorphism carrying one onto the other?
 In light of the above discussion which actually only requires $\mathbb{C}$-extremality, we may readily cite $D$ to be the $\ell^{1}$-ball in which the origin is fixed by every of its automorphisms; a retract through any pair of linearly independent points in $D$ is therefore not ${\rm Aut}(D)$-equivalent to any of the retracts passing through the origin.
-To address other hypotheses of the foregoing theorem briefly, let us remark that a non-trivial holomorphic retraction map is always a submersion except perhaps on a set of measure zero (page $74$ of the book [Hrsch]).
-Nevertheless, we admit that it would be more significant to ascertain whether the origin falls outside such a measure zero set, for some retraction map $\rho$ as in the above theorem; or, if all retraction maps of such a class of domains is a submersion throughout the domain, making the submersive hypotheses redundant.
-Perhaps a more useful remark would be that as a corollary of the proof of this theorem, it will be seen that the strict convexity can be relaxed to $\mathbb{C}$-extremality to obtain the same description of $Z$, as soon as it ascertained that $L\cap\partial D$ has at-least one point which is $\mathbb{R}$-extremal in $\partial D$.
-8. Proof of proposition 1.6
-Before we begin to work out the proof, we need to make some general remarks; to indicate as to why there are enough instances of domains verifying the conditions stipulated in proposition 1.6.
-Indeed, even before this, let us indicate as to why it may perhaps be of mild interest atleast, to note in passing (as we did in the preliminaries section) that: one may deduce out of the general facts about retracts and the already mentioned well-known characterization of retracts of polydiscs, yet another proof of the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis which is, the biholomorphic inequivalence of two of the simplest bounded simply connected domains namely, the ball $\mathbb{B}$ and $\Delta^{N}$ (despite their diffeomeorphic equivalence), as soon as the dimension $N$ gets any bigger than one; and other such inequivalences using retracts.
-We have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (and so that proof is not repeated here).
-However here, we would indeed like to hasten to mention that this does not obtain for such $D$, results as good as those, which already well-known techniques in this realm can!
+Finally next here, we get to the question of at-least one example wherein all hypotheses of the theorem are satisfied.
+Example: Consider the $\ell_{1}$-ball
+$B:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<1\}.$
+Consider the holomorphic map $\varphi:\Delta\to B$ defined by $\varphi(\zeta)=\left((\frac{1+\zeta}{2})^{2},(\frac{1-\zeta}{2})^{2}\right)$.
+By proposition 16.2.2 in [Jrncki_invrnt_dst], we obtain that $\varphi$ is a complex geodesic in $B$.
+Hence
+$Z:={\rm image}(\varphi)=\Big{\{}\left((\frac{1+\zeta}{2})^{2},(\frac{1-\zeta}{2})^{2}\right)\;:\;\zeta\in\Delta\Big{\}}$
+is a retract of $B$ which we note: does not pass through the origin.
+Next we want to find a retraction map $\rho$ on $B$ such that $\text{image}(\rho)=Z$.
+Towards this, it suffices to determine a holomorphic left-inverse for the map $\varphi$ as discussed in proposition 2.32 of the preliminaries section.
+It is easy to see the map $\phi:B\to\Delta$ defined by $\phi(z,w)=z-w$ serves as the required left-inverse: $\phi\circ\varphi(\zeta)=\zeta$ for all $\zeta\in\Delta$.
+It follows that the map $\rho(z,w)=\varphi\circ\phi(z,w)$ gives the holomorphic retraction of $B$ onto $Z$.
+As $B$ is convex with $\mathbb{C}$-extremal boundary, the only hypotheses that needs to be checked is that about the derivative of $\rho$.
+For this we need its explicit expression which can be computed to be:
+$\rho(z,w)=\left(\left(\frac{1+z-w}{2}\right)^{2},\left(\frac{1-z+w}{2}\right)^{2}\right).$
+In-particular, $\rho$ maps the origin to $p=(1/4,1/4)$ at which point $L=T_{p}Z$ is given by $L=\{(z,w)\in\mathbb{C}^{2}:z+w=0\}$. Next, we compute the matrix of $D\rho_{|_{(z,w)}}$.
+$D\rho_{|_{(z,w)}}=\frac{1}{2}\begin{bmatrix}1+z-w&-(1+z-w)\\ -(1-z+w)&1-z+w\end{bmatrix}$
+In particular,
+$D\rho_{|_{(0,0)}}=\frac{1}{2}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$
+So, $D\rho_{|_{(0,0)}}(z,w)=(\frac{z-w}{2},\frac{-(z-w)}{2})$ which is apparently a projection operator and of norm-one with respect to the $\ell_{1}$-norm.
+This concludes the verification of all hypotheses of theorem 1.19.
+We discuss briefly the form of the conclusion as well.
+Now we want to show that $Z$ is a graph of a holomorphic function over $B_{L}=B\cap L$, where $L=\{(z,w)\in\mathbb{C}^{2}:z+w=0\}$ which we note is the tangent space to $Z$ at $p=(1/4,1/4)$.
+Firstly note that $B_{L}=\{(z,-z):|z|<1/2\}$.
+Consider the holomorphic map $\psi:B_{L}\to B$ defined by
+$\psi(z,-z)=\varphi(2z)=\left(\left(\frac{1+2z}{2}\right)^{2},\left(\frac{1-2z}{2}\right)^{2}\right).$
+Denote the projection operator $D\rho_{|_{0}}$ by $\pi$ for short. Note that
+$\pi\circ\psi_{|_{B_{L}}}(z,-z)=\pi\left(\left(\frac{1+2z}{2}\right)^{2},\left(\frac{1-2z}{2}\right)^{2}\right)=\frac{1}{2}(2z,-2z)=(z,-z)$
+The above computation shows that $\pi\circ\psi_{|_{B_{L}}}$ is the identity map on $B_{L}$.
+This implies that $Z$ is the graph of holomorphic function $\psi$ over $B_{L}$.
+The reason for the existence of non-linear retractions $\rho$ with $D\rho_{|_{0}}$ being a norm-one projection as in the above example, owes to the boundary failing to be $\mathbb{R}$-extremal (though it is $\mathbb{C}$-extremal).
+This is brought out by the next result.
+Proposition 7.1.
+Let $D$ be a bounded balanced strictly convex domain in $\mathbb{C}^{N}$.
+Let $\rho:D\to D$ be a retraction map on $D$ with $p=\rho(0)$ and $Z:={\rm image}(\rho)$.
+If $D\rho_{|_{0}}$ is a norm-one projection i.e., a projection which maps $D$ onto $D_{L}$ (where $L=T_{p}Z$ as usual), then $\rho$ is a linear map and hence $Z$ is a linear retract of $D$.
+Proof.
+Consider the holomorphic map $F:D_{L}\to D$ defined by $F(z)=\rho_{|_{D_{L}}}(z)$.
+Note that $DF_{|_{0}}(z)=D\rho_{|_{0}}(z)$, which is a linear projection onto $D_{L}$.
+In particular, $D\rho_{|_{0}}$ is the identity map on $D_{L}$.
+Applying Mazet’s Schwarz lemma to the map $F:D_{L}\to D$, we obtain that $F(z)=DF_{|_{0}}(z)$.
+Hence we get that ${\rm image}(F)=D_{L}\subset Z$.
+Note that $D_{L}$ and $Z$ have the same dimension.
+Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$.
+This finishes the proof of the proposition.
+∎
+8. Proof of theorems 1.51.9 and Corollary 1.11
+Before we begin to work out the proof of the pair of theorems as in the title of this section, we need to make some general remarks.
+Theorems 1.5 is about the irreducibility of balanced $\mathbb{C}$-extremal analytic polyhedra.
+It will be derived as a consequence of theorem 1.9 about retracts through the origin in such domains; this will therefore be established before theorem 1.5.
+Along the way, we shall attain useful propositions about biholomorphic inequivalences between co-analytic domains and analytic polyhedra (or more generally, bounded hyperconvex domains).
+Also included in this section is the proof of corollary 1.11 about the irreducibility of the anisotropic non-convex egg domains, whose boundaries are holomorphically extreme in good contrast to analytic polyhedra; nevertheless, this will be simultaneously attained alongwith the proof of theorem 1.5.
+The above exercise in showing irreducibility may seem to need justification, owing to the existence of well-known methods which yield better results for certain classes of domains even if they are different.
+To put this matter in perspective we recall that the famous inequivalence between the Euclidean ball $\mathbb{B}$ and the polydisc $\Delta^{N}$ for all $N\geq 2$, can be shown using retracts as well; while we believe that this proof will atleast be of mild interest to atleast note in passing, we have better justifications below.
+Now, we have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (so that proof is not repeated here).
+Here, we would indeed like to hasten to mention that this does not obtain for such $D$: results as good as can be obtained from already well-known techniques in this realm can!
 – so, this leads to the question as to why one may want to even consider this proof even if the problem of biholomorphic inequivalence and its special case of establishing irreducibility of a domain $D$ i.e., showing that $D$ is not biholomorphic to a product of lower dimensional domains, is undoubtedly fundamental.
@@
 We shall presently show that the route via retracts can very well dispense with such conditions, if $D$ is a balanced domain of holomorphy.
-To exposit this better by comparing and contrasting with the Remmert – Stein techniques, let us mention that the aforementioned proof by retracts for $\mathbb{B}\not\simeq\Delta^{N}$, shares (unlike a proof based on comparison of automorphism groups for instance), with the best possible techniques in this context, the essential idea about the difference in the lack or presence of analytic discs in the boundary of $D$ as being responsible for such inequivalences, as it has a bearing on the geometry of retracts through the origin in $D$.
-Moreover, this provides a shorter way to establish inequivalences without having to worry about the boundary extendibility (or boundary behaviour) of biholomorphic mappings since the boundary geometry is ‘internalized’ as features about the retracts through points in the interior; we believe this will be useful for establishing new inequivalences between pairs of bounded balanced domains in both of whose boundaries there reside non-trivial analytic discs.
-To illustrate this point with a concrete and simple example, consider the question of whether the analytic polyhedron
-$D_{h}=\{(z,w)\in\mathbb{C}^{2}:|w^{2}-z^{2}|<1\}\cap\{(z,w)\in\mathbb{C}^{2}:|zw|<2\}$
-is biholomorphic to a product domain  [^5] A motivation for the specific choices here, is that the simplest examples of bounded product domains namely polydiscs are defined as sublevel sets of a pair of (homogeneous) holomorphic linear polynomials and we are interested in bounded domains whose definition bears some semblance to that of polydiscs, but are nevertheless irreducible..
+To exposit this better by comparing with the Remmert – Stein techniques, let us mention that the aforementioned proof by retracts for $\mathbb{B}\not\simeq\Delta^{N}$, shares (unlike a proof based on comparison of automorphism groups for instance), with the best possible techniques in this context, the essential idea about the difference in the lack or presence of analytic discs in the boundary of $D$, as being responsible for such inequivalences; this difference has a bearing on the geometry of retracts through the origin in $D$.
+Moreover, this provides a shorter way to establish inequivalences without having to worry about the boundary extendibility (or boundary behaviour) of biholomorphic mappings since the boundary geometry is ‘internalized’ (atleast partly) as features about the retracts through points in the interior; for instance, holomorphic extremality of the boundary of a bounded balanced domain of holomorphy entails linearity of the retracts through its center and these linear slices have holomorphically extreme boundaries again.
+We believe this will be useful for establishing new inequivalences between pairs of bounded balanced domains in both of whose boundaries there reside non-trivial analytic discs.
+We now illustrate these points with a couple of progressively complex examples and theorems.
+We start with a concrete and simple example of a special analytic polyhedron; indeed, this is the simplest example of a bounded balanced analytic polyhedron with $\mathbb{C}$-extremal boundary and will serve as a running example for us.
+Specifically, consider the question of whether
+(8.1) $D_{h}:=\{(z,w)\in\mathbb{C}^{2}:|w^{2}-z^{2}|<1\}\cap\{(z,w)\in\mathbb{C}^{2}:|zw|<2\}$
+is biholomorphic to a product domain  [^11] A motivation for the specific choices here, is that the simplest examples of bounded product domains namely polydiscs are defined as sublevel sets of a pair of (homogeneous) holomorphic linear polynomials and we are interested in bounded domains whose definition bears some semblance to that of polydiscs, but are nevertheless irreducible.
+Indeed as will be seen, this example is the simplest special analytic polyhedron which is balanced and non-linear..
 To begin with analysing this example first, we note that while an open set cut-out by such polynomial inequalities in general may fail to be connected, such issues do not arise here: the homogeneity of the defining pair of quadratic polynomials leads to $D_{h}$ being balanced, thereby (immediately seen to be connected and) contractible.
@@
 $\left||z|-|w|\right|\leq|z+iw|$
-when squared and combined with 8.1, leads to
+when squared and combined with 8.2, leads to
 $|z|^{2}+|w|^{2}<5+2|zw|.$
@@
 While the removal of $A$ from $D_{h}$ never disconnects $D_{h}$, it always destroys the simply connectedness of $D_{h}$, as is always the case when proper analytic sets are subtracted from any domain.
-Though not required by the statement of proposition 1.6, the case $D=D_{h}$ also serves as a concrete example of a (hyperbolic) balanced domain of holomorphy with analytic discs in its boundary (without purposefully removing analytic sets!), which feature as already mentioned is the case of greatest interest in contributing to any new results or routes about establishing non-existence of a product structure in a domain.
 To establish the irreducibility of such $D_{h}^{A}$, let us start with the most natural attempt of assuming to obtain a contradiction, the existence of a biholomorphism $F$ from $D_{h}^{A}$ onto a product domain $D$; no less standard then, is to analyze the boundary behaviour of $F$.
 However, this is rather cumbersome owing to potential complexity of the boundary of the factors $B_{1},B_{2}$.
-It is here where consideration of retracts provides a far more convenient approach, as already alluded to in the above.
-However, before we discuss the proof of propositions about the irreducibility of analytic complements as in proposition 1.6, it is convenient to formulate part of its proof as the following more general lemma.
+It is here where consideration of retracts provides a far more convenient approach, as already alluded to in the above, particularly for theorem 1.5, one of the main theorems of this section.
+However, before we discuss the proof of theorem 1.5, it is convenient to formulate part of its proof as the following more general lemma.
 Lemma 8.1.
-Let $D$ be any bounded pseudoconvex domain in $\mathbb{C}^{N}$ either with a Stein neighbourhood basis or with a bounded plurisubharmonic defining function.
-Then any holomorphic retraction map on $D\setminus A$, where $A$ is a proper analytic subset of $D$ can be extended to a holomorphic retraction map on $D$.
+Let $D$ be any bounded hyperconvex domain in $\mathbb{C}^{N}$.
+Let $A$ be any non-trivial analytic subset of $D$.
+Then any holomorphic retraction map on $D\setminus A$, be extended to a holomorphic retraction map on $D$.
 Consequently, every retract $R^{A}$ of $D\setminus A$ is of the form $R^{A}=R\cap(D\setminus A)$ for some retract $R$ of $D$.
 Proof.
-Firstly recall that removing an analytic set from a domain does not disconnect it; by definition $A$ is closed in $D$, thereby $D\setminus A$ is a domain in $\mathbb{C}^{N}$.
+Firstly, recall that removing an analytic set from a domain does not disconnect it; by definition $A$ is closed in $D$, thereby $D\setminus A$ is a domain in $\mathbb{C}^{N}$.
 Let $\rho:D\setminus A\to D\setminus A$ be a holomorphic retraction map.
@@
 If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $a$.
-If $D$ admits a bounded plurisubharmonic defining function $\psi$, then we consider the composition $\phi:=\psi\circ\tilde{\rho}\circ\varphi$, which is a subharmonic function on $\Delta^{*}=\Delta\setminus\{0\}$; as it is also bounded (by the boundedness of $\psi$), it follows by that the origin is a removable singularity for this subharmonic function, thereby $\phi$ extends to a subharmonic function on $\Delta$ which we continue to denote by $\phi$.
+Since $D$ is hyperconvex, it follows from the theorem 2.50 and its remark that $\tilde{\rho}(a)$ is a local psh barrier point of $D$.
+So, there exists neighborhood $U$ of $b:=\tilde{\rho}(a)$ and plurisubharmonic function $\psi$ on $U\cap D$ such that $\psi<0$ on $U\cap D$ and $\lim_{U\cap D\ni z\to b}\psi(z)=0$.
+After shrinking the radius of the punctured disc $\Delta^{*}$ (if needed), we may assume that $image(\tilde{\rho}\circ\varphi)\subset U$.
+Consider the subharmonic function $\phi:\Delta^{*}\to\mathbb{R}$ defined by $\phi(z)=\psi\circ\tilde{\rho}\circ\varphi(z)$.
+By applying the removable singularity theorem for subharmonic functions, we obtain that 0 is a removable singularity for $\phi$.
+Hence $\phi$ extends to a subharmonic function on $\Delta$ which we continue to denote by $\phi$.
+By the limiting behaviour of $\psi$ at $b$, we obtain that $\phi(0)=0$.
 But then, the maximum principle for this subharmonic function is violated, under our assumption that $b\in\partial D$.
 We deduce therefore that $b$ cannot lie in $\partial D$ and it follows that $\tilde{\rho}(D)\subset D$.
-Even if the existence of such a $\psi$ is not known, it is indeed true that $\tilde{\rho}(D)$ lies within $D$ (the above argument argument was given to show a simple route under the aforementioned assumption about bounded plurisubharmonic defining functions).
-This follows from the last result in [Abe], owing to our hypothesis that $D$ has a Stein neighbourhood basis; as the work [Abe] has some non-major errors, we also cite the Masters thesis of Persson [Persson], wherein the errors are set right with a better exposition.
-Thus as $\tilde{\rho}(a)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
 We finish off by verifying that $\tilde{\rho}$ is a retraction on $D$.
@@
 Remark 8.2.
+We first recall that for the class of much interest in this article namely bounded balanced pseudoconvex domains, the hypothesis of the above is not satisfied automatically as hyperconvexity of such domains is equivalent to continuity of the Minkowski functional (see Remark 2.45 (iii) in the preliminaries section); nevertheless, it turns out that the conclusion of the above lemma is true regardless of such regularity, owing to certain results that we briefly touch upon now.
+For our applications however, we shall not require such more general observations.
+To remark on the extension then, the above lemma can be derived for bounded pseudoconvex domains $D$ even if the existence of such a $\psi$ as in its proof above is not known, provided $D$ has a Stein neighbourhood basis.
+The only point to be checked is $\tilde{\rho}(D)\subset D$.
+It is indeed true that $\tilde{\rho}(D)$ lies within $D$ and can be derived rather quickly from known results (the arguments above were given to show a simple route under the aforementioned assumption about bounded plurisubharmonic defining functions).
+This follows from the last result in [Abe], owing to our hypothesis that $D$ has a Stein neighbourhood basis; as the work [Abe] has some non-major errors, we also cite the Masters thesis of Persson [Persson], wherein the errors are set right with a better exposition.
+Thus as $\tilde{\rho}(a)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
+Remark 8.3.
 One may want to know to what extent we may weaken the hypotheses of the above lemma.
@@
 Notice then that $R^{A}=\{(z,1/z):z\in\mathbb{C}^{*}\}$ is a retract of $D$.
-We observe that there exists no retract $R$ of $\mathbb{C}^{2}$ such that $R^{A}=R\cap(D\setminus A)$.
-Indeed, this follows from the fact that the (smooth) analytic variety $R^{A}$ described by the equation $zw-1=0$, is irreducible, together with the principle of analytic continuation for complex analytic sets.
+We observe that there exists no retract $R$ of $\mathbb{C}^{2}$ such that $R^{A}=R\cap D$.
+Indeed, this follows from the fact that the (smooth) analytic variety $R^{A}$ described by the equation zw1=0zw-1=0, is irreducible, together with the principle of analytic continuation for complex analytic sets.
 Notwithstanding the example in the above remark, we may certainly remove analytic varieties of higher codimension to attain the analogue of the foregoing lemma for the case when the domain is the entire complex Euclidean space itself: to adapt the essential ideas of its proof to deal with the unboundedness now, we only need to employ (theorem 5B in [Whitney]) in place of the standard Riemann removable singularities theorem.
@@
 Proof of proposition 1.4 .
-Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ Since $D_{1}$ is bounded, there exist $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$.
+Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ .
+Since $D_{1}$ is bounded, there exist $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$.
 Since $D_{2}$ is open, there exist $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$.
@@
 Let $v_{1}\in T_{p_{1}}D_{1}$ be any non-zero vector.
-Then there exist $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, there is a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ such that $(w_{1},v_{2})\in T_{p}Z$, for some non-zero vector $w_{1}$ with $w_{1}\in span_{\mathbb{C}}\{v_{1}\}$.
+Then there exist $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, there is a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ with $T_{p}Z\ni(w_{1},v_{2})$, for some non-zero vector $w_{1}$ with $w_{1}\in\operatorname{span}_{\mathbb{C}}\{v_{1}\}$.
 Proof.
@@
 Again note that $F(p_{1},w)=(p_{1},p_{2})$ and $F^{\prime}(p)(w_{1},w_{2})=(w_{1},\phi^{\prime}(p_{1})(w_{1}))=(w_{1},v_{2})$.
-Hence $Z:=\{(z,\phi(z)):z\in D_{1}\}$ is a retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(w_{1},v_{2})\in T_{p}Z$, where $w_{1}\in span_{\mathbb{C}}<v_{1}>$.
+Hence $Z:=\{(z,\phi(z)):z\in D_{1}\}$ is a retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(w_{1},v_{2})\in T_{p}Z$, where $w_{1}\in\operatorname{span}_{\mathbb{C}}\{v_{1}\}$.
 ∎
-Next we shall prove proposition 1.6.
-Proof of proposition 1.6.
-Let us begin by considering the set of boundary points of the form $q/h(q)$ for $q$ varying within $D\setminus A$.
-Denote the subset of such boundary points satisfying the additional condition that $\partial D$ is non-convex at $q/h(q)$, by $S^{h}$.
-It follows from the hypotheses that $S^{h}$ is a dense subset of $\partial D$ (as is not difficult to see).
-Next, we observe that it follows from the above lemma 8.1 together with corollary 2.33, there is no one-dimensional retract through 00 in the direction of $p\in S^{h}$.
-Note that lemma 8.1 is applicable here owing to the fact that bounded balanced pseudoconvex domains are guaranteed to have a Stein neighbourhood basis by theorem 1.17.16 in [Jrnck_Frst_stps].
-With these observations, we are ready to prove the proposition by contradiction.
-Indeed, if $F$ were a biholomorphism mapping $D\setminus A$ onto a product domain $D$, then through $F(0)$ we have a retract along those directions given by proposition 8.4 – note that the hypothesis about Caratheodory hyperbolicity therein is verified owing to the fact that $D\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D\setminus A$ must be Caratheodory hyperbolic as well.
-In particular, there always exist a non-trivial retract (necessarily of dimension one here) of $D$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$.
-By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exist a (non-trivial) retract of $D$ through the origin with its tangent direction at 00 belonging to $S^{\prime}$.
-But this contradicts the hypotheses that $D$ is not convex at an open dense subset of $\partial D$ which implies that for each non-convex boundary point $a$ of $D$, there are no retracts of $D\setminus A$ passing through the origin in the direction of $a$.
-This finishes the proof.
+We shall now employ this proposition in applying retracts to establish various biholomorphic inequivalences amongst members of $3$ classes of domains namely, analytic polyhedra, product domains and co-analytic domains i.e., domains of the form $D\setminus A$ wherein $D$ is a domain of holomorphy, an assumption that we generally make unless explicitly stated otherwise.
+all such domains have the common feature that their boundaries are not entirely smooth and contain non-trivial analytic varieties.
+On the other hand (as also mentioned in the introductory section), there are differences in the location of the analytic boundary pieces: for instance, in a co-analytic domain there are interior analytic boundary points whereas analytic polyhedra are devoid of such ‘interior’ boundary points, leading them to be biholomorphically inequivalent.
+To transform this intuition into a rigorous proof as well as to attain a more result a bit more general, we note that co-analytic domains are not hyperconvex whereas analytic polyhedra are.
+The biholomorphic invariance of hyperconvexity immediately renders the desired inequivalence but take a moment to record this in the following proposition among other related statements in order to be able to refer and make use of it in the sequel.
+We first formalize the aforementioned notion of boundary points lying in the ‘interior’ of the domain, which we will be used later.
+Definition 8.6.
+Let $\Omega$ be a domain in $\mathbb{C}^{N}$ and $p\in\partial\Omega$.
+We say that $p$ is an interior boundary point of $\Omega$ if there exists a neighbourhood $U_{p}$ of $p$ such that $U_{p}\cap\overline{\Omega}=U_{p}$.
+Next, we prove a lemma regarding the product of hyperconvex domains in $\mathbb{C}^{N}$.
+Lemma 8.7.
+Let $D_{1},D_{2}$ be domains in $\mathbb{C}^{N},\mathbb{C}^{M}$ respectively. Then $D:=D_{1}\times D_{2}$ is hyperconvex if and only if $D_{1}$ and $D_{2}$ are hyperconvex.
+Proof.
+Assume that $D:=D_{1}\times D_{2}$ is hyperconvex and fix a point $(z_{0},w_{0})\in D$.
+Note that $D_{1}\times\{w_{0}\}$ and $\{z_{0}\}\times D_{2}$ are retracts of $D$.
+By proposition 2.49, $D_{1}$ and $D_{2}$ are hyperconvex.
+Conversely, assume that $D_{1},D_{2}$ are hyperconvex.
+Let $\varphi_{1},\varphi_{2}$ be psh exhaustion functions of $D_{1},D_{2}$ respectively.
+Consider the function $\varphi$ on $D_{1}\times D_{2}$ defined by $\varphi(z,w)=\max\{\varphi_{1}(z),\varphi_{2}(w)\}$.
+Then $\varphi$ is plurisubharmonic and negative on $D_{1}\times D_{2}$.
+For each $c<0$, let $S^{c}:=\{(z,w)\in D:\varphi(z,w)<c\}$, $S_{1}^{c}:=\{z\in D_{1}:\varphi_{1}(z)<c\}$ and $S_{2}^{c}:=\{w\in D_{2}:\varphi_{2}(w)<c\}$ .
+Note that
+$(z,w)\in S^{c}\Longleftrightarrow\varphi_{1}(z)<c~~\text{and}~\varphi_{2}(w)<c\Longleftrightarrow(z,w)\in S_{1}^{c}\times S_{2}^{c}.$
+This shows that $S^{c}=S_{1}^{c}\times S_{2}^{c}$.
+Since $S_{1}^{c}$ and $S_{2}^{c}$ are relatively compact subsets of $D_{1},D_{2}$ respectively, $S$ is a relatively compact subset of $D$.
+Hence, $D_{1}\times D_{2}$ is hyperconvex.
 ∎
+Proposition 8.8.
+  1. (i)
+    Let $D_{1}$ be a complex manifold, $A_{1}$ be a non-empty analytic subset of $D_{1}$ and let $D_{2}$ be a bounded hyperconvex domain in $\mathbb{C}^{N}$, $A_{2}$ be an (possibly empty) analytic subset of $D_{2}$.
+If $F:D_{1}\setminus A_{1}\to D_{2}\setminus A_{2}$ is a proper holomorphic map then $F$ can be extended as a holomorphic map $\tilde{F}$ from $D_{1}$ to $D_{2}$ with $\tilde{F}(A_{1})\subset A_{2}$.
+Consequently, there does not exist any proper holomorphic map from $D_{1}\setminus A_{1}$ to $D_{2}$ (as follows by considering the case $A_{2}=\phi$).
+In particular, there does not exist a biholomorphic map from $D_{1}\setminus A_{1}$ onto $D_{2}$.
+  2. (ii)
+    Let $D_{1},D_{2}$ be bounded hyperconvex domains in $\mathbb{C}^{N}$ and $A_{1},A_{2}$ be non-empty analytic subsets of $D_{1},D_{2}$ respectively.
+If $D_{1}\setminus A_{1}$ is biholomorphic to $D_{2}\setminus A_{2}$ then $D_{1}$ is biholomorphic to $D_{2}$.
+Proof.
+Assume that there exists a proper holomorphic map $F$ mapping $D_{1}\setminus A_{1}$ to $D_{2}\setminus A_{2}$.
+By applying the Riemann removable singularity theorem to each of the component functions of $F$, we get a holomorphic extension of $F$ which we denote by $\tilde{F}$.
+The fact that $F$ maps $D_{1}\setminus A_{1}$ into $D_{2}\setminus A_{2}$ implies that $\tilde{F}$ maps $D_{1}$ into $\overline{D_{2}}$.
+Using the same arguments in the proof of the lemma 8.1, we obtain that $\tilde{F}$ maps $D_{1}$ into $D_{2}$ and $\tilde{F}(A_{1})\subset A_{2}$.
+Considering the special case $A_{2}=\phi$, we get that there does not exist any proper map from $D_{1}\setminus A_{1}$ to $D_{2}$.
+This finishes the proof of (i).
+To prove (ii), let $F$ be a biholomorphic map from $D_{1}\setminus A_{1}$ onto $D_{2}\setminus A_{2}$ and $G$ be the inverse of $F$.
+By the arguments similar to those used in (i), we obtain that $\tilde{F}$ maps $D_{1}$ into $D_{2}$ and $\tilde{G}$ maps $D_{2}$ into $D_{1}$.
+It is easy to see that $\tilde{F}$ and $\tilde{G}$ are biholomorphic maps.
+To prove this, pick an arbitrary point $z_{0}\in A_{1}$ and a sequence $(z_{n})$ in $D_{1}\setminus A_{1}$ converges to $z_{0}\in A_{1}$.
+Since $\tilde{F}\equiv F$ on $D_{1}\setminus A_{1}$ and $\tilde{G}\equiv G$ on $D_{2}\setminus A_{2}$, $z_{n}=\tilde{G}\circ\tilde{F}(z_{n})\to\tilde{G}\circ\tilde{F}(z_{0})$.
+By the uniqueness of limits, $\tilde{G}\circ\tilde{F}(z_{0})=z_{0}$.
+This shows that $\tilde{G}\circ\tilde{F}$ is the identity map on $D_{1}$.
+Similarly, $\tilde{F}\circ\tilde{G}$ is the identity map on $D_{2}$.
+This proves that $\tilde{F}$ is a biholomorphism on $D_{1}$ and its inverse is $\tilde{G}$.
+Hence $D_{1}$ is biholomorphic to $D_{2}$.
+∎
+As a corollary of part (i) of the above proposition 8.8 and by above lemma 8.7, we obtain the following.
+Corollary 8.9.
+Let $D$ be a pseudoconvex domain in $\mathbb{C}^{N}$ and $A$ be a non-empty analytic subset of $D$. Then $D\setminus A$ is not biholomorphic to a product of bounded hyperconvex domains.
+In particular, it follows immediately from this corollary that an analytic polyhedron is never biholomorphically equivalent to a co-analytic domain.
 Remark 8.10.
-Though we had assumed in the statement of proposition 1.6 that the dimension $n=2$, the above proof goes through for arbitrary dimension to show atleast that if $D$ is a balanced domain of holomorphy in $\mathbb{C}^{N}$ with the same assumptions as in that proposition (and $A$ an analytic subset not passing through origin), then $D\setminus A$ is not biholomorphic to a product domain with a one-dimensional factor; in-particular, it is not completely reducible meaning that $D\setminus A$ is not biholomorphic to a product of one-dimensional domains.
-How about not complete reducibility of the domain but an intermediate one?
-Establishing such a complete non-existence of a product structure for a large class of domains is rather challenging.
-Let us only remark that as a direct application of proposition 8.4, we may derive the following corollary from theorem 1.9 proved in an earlier section.
-Corollary 8.6.
-Let $D_{q}$ denote the $\ell_{q}$-‘ball’ for $10<q<$ 1 and $A$ any analytic subset of $D_{q}$ not containing the origin.
-Then, neither $D_{q}$ nor $D_{q}\setminus A$, is biholomorphic to a product domain.
-We conclude this section with a complete determination of retracts of $D_{h}\setminus A$ through the origin, where $D_{h}$ is the polynomial polyhedron discussed around the beginning of this section.
-As above, $A$ is an analytic subset of $D_{h}$ not passing through origin, which is allowed to be empty but not $D_{h}$ itself.
+In general, it can very well happen that both $D$ and $D\setminus A$ are product (pseudoconvex) domains.
+Indeed the simplest example showing this is $D=\Delta^{2}$ and $D\setminus A=\Delta\times\Delta^{*}$ obtained by taking $A:=\{(z,w)\in\Delta^{2}:w=0\}$; while there are plenty of other examples, the next proposition shows that within the class of bounded balanced domains of holomorphy in dimension two, this is the only one.
+This proposition also serves as one of the simplest applications of retracts to establishing biholomorphic inequivalences.
+Proposition 8.11.
+Let $D$ be any bounded balanced domain of holomorphy in $\mathbb{C}^{2}$ (whose Minkowski functional we denote by $h=h_{D}$), which is not biholomorphic to $\Delta^{2}$.
+Let $A$ be any proper analytic subset of $D$.
+Then neither $D$ nor $D\setminus A$ is biholomorphic to a product domain.
+Proof.
+To prove this proposition by contradiction, assume that $D\setminus A$ is biholomorphic to a product domain $\Omega_{1}\times\Omega_{2}$, where $\Omega_{1},\Omega_{2}$ are domains in $\mathbb{C}$.
+Let $F$ be a biholomorphism from $D\setminus A$ onto $\Omega_{1}\times\Omega_{2}$ and $G$ be the inverse of $F$.
+Let $(p_{1},p_{2})\in\Omega_{1}\times\Omega_{2}$.
+Note that $\Omega_{1}\times\{p_{2}\}$ and $\{p_{1}\}\times\Omega_{2}$ are retracts of $\Omega_{1}\times\Omega_{2}$.
+Let $G$ be a biholomorphism from $\Omega_{1}\times\Omega_{2}$ onto $D\setminus A$.
+By proposition 2.4, $\tilde{R_{1}}:=G(\Omega_{1}\times\{p_{2}\})$ and $\tilde{R_{2}}:=G(\{p_{1}\}\times\Omega_{2})$ are retracts of $D\setminus A$.
+This implies that $\tilde{R_{1}},\tilde{R_{2}}$ are biholomorphic to $\Omega_{1},\Omega_{2}$ respectively.
+Since $D\setminus A$ is biholomorphic to $\Omega_{1}\times\Omega_{2}$, we obtain that $D\setminus A$ is biholomorphic to $\tilde{R_{1}}\times\tilde{R_{2}}$.
+By lemma 8.1, there exist one-dimensional retracts $R_{1},R_{2}$ of $D$ such that $\tilde{R_{1}}=R_{1}\cap(D\setminus A)$ and $\tilde{R_{2}}=R_{2}\cap(D\setminus A)$.
+If $R_{i}\subset D\setminus A$ for $i=1,2$ then $\tilde{R_{i}}=R_{i}$.
+Since $R_{i}$ is a simply connected Riemann surface, its biholomorphic image $\Omega_{i}$ is a simply connected domain in $\mathbb{C}$.
+By Riemann mapping theorem, $R_{i}$ is biholomorphic to $\Delta$.
+Hence we obtain that $D\setminus A$ is biholomorphic to $\Delta^{2}$, which contradicts part (i) of the above proposition 8.8.
+Next, we consider the case that $R_{i}\cap A\neq\phi$ for at least one $i$.
+Let $S$ be the set of all points of intersection of $R_{i}$ with $A$.
+As $R_{i}\cap A$ does not have any interior point and they are of dimension one, $S$ is a discrete set and thereby can be enumerated as $(p_{j})_{j=1}^{\infty}$.
+Fix one of the $p_{j}$’s say $p_{1}$ for simplicity.
+There exists an open set $U_{1}\subset R_{i}$, which is contained within at least one chart about $p_{1}$ for $R_{i}$.
+Let $\phi$ be the biholomorphism from $\tilde{R_{i}}$ onto $\Omega_{i}$.
+By Casorati-Weiestrass theorem, $p_{1}$ is not an essential singularity for $\phi_{|_{U_{1}^{*}}}$, where $U_{1}^{*}=U_{1}\setminus\{p_{1}\}$.
+So, the isolated singularity for $\phi$ at $p_{1}$ is removable when $\phi$ is viewed as a map into $\mathbb{C}_{\infty}$.
+Let $V_{1}:=\phi(U_{1})$ and $V_{1}^{*}=V_{1}\setminus\{q_{1}\}$, where $q_{1}=\phi(p_{1})$.
+By local-normal form theorem, after passing to suitable charts, we obtain that $\partial(\phi(U_{1}^{*}))=\phi(\partial U_{1})\cup\{q_{1}\}$ with $q_{1}\notin\phi(\partial U_{1})$.
+Hence $q_{1}$ is an interior boundary point of $V_{1}$.
+This argument holds for all $p_{j}$.
+Since $\phi$ is a biholomorphism on $\tilde{R_{i}}$, $\phi_{|_{S}}$ is one-one.
+Hence $R_{i}$ is biholomorphic to $\Omega_{i}^{\prime}:=\Omega_{i}\cup\phi(S)$.
+Since $R_{i}$ is simply connected, $\Omega_{i}^{\prime}$ is simply connected and thereby $R_{i}$ is biholomorphic to $\Delta$ by Riemann mapping theorem.
+Hence there exist discrete subsets $L_{1},L_{2}\subset\Delta$ (at least one of $L_{i}$ is non-empty) such that $\tilde{R_{1}}$ and $\tilde{R_{2}}$ are biholomorphic to $\Delta\setminus L_{1}$ and $\Delta\setminus L_{2}$ respectively.
+This implies that $D\setminus A$ is biholomorphic to $(\Delta\setminus L_{1})\times(\Delta\setminus L_{2})$.
+Write $(\Delta\setminus L_{1})\times(\Delta\setminus L_{2})=\Delta^{2}\setminus S^{\prime}$, where $S^{\prime}=(\Delta\times L_{2})\cup(L_{1}\times\Delta)$.
+Note that $S^{\prime}$ is a non-empty analytic subset of $\Delta^{2}$.
+By applying the Riemann removable singularity theorem to each of the component functions of $F$, we get a holomorphic extension of $F$ which we denote by $\tilde{F}$.
+The fact that $F$ maps $D\setminus A$ to $\Delta^{2}\setminus S^{\prime}$ implies that $\tilde{F}$ maps $D$ into $\overline{\Delta}^{2}$.
+By theorem 1.17.16 in [Jrnck_Frst_stps], both $D$ and $\Delta^{2}$ admit Stein neighborhood bases.
+Applying the same argument in the proof of the lemma 8.1 and its remark, we conclude that $\tilde{F}$ maps $D$ into $\Delta^{2}$ and $\tilde{G}$ maps $\Delta^{2}$ into $D$.
+Using the methods similar to the proof of part (ii) of the above proposition 8.8, we obtain that $\tilde{F}$ and $\tilde{G}$ are biholomorphic maps.
+This proves that $D$ is biholomorphic to the polydisc $\Delta^{2}$.
+∎
+Next,we prove the corollary of above proposition 8.5 and theorem 1.8.
+Proof of Corollary 1.11: If $F$ is a biholomorphic mapping from $D_{q}\setminus A$ onto a product domain $G=G_{1}\times G_{2}$, then through $F(0)$ we have a retract along those directions given by proposition 8.5 – note that the hypothesis about Caratheodory hyperbolicity therein (for $G$ here) is verified owing to the fact that $D_{q}\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D_{q}\setminus A$ must be Caratheodory hyperbolic as well (refer to theorem 18.2.1 of [Jrncki_invrnt_dst]).
+In particular, there always exists a non-trivial retract (necessarily of dimension one here) of $D_{q}$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$.
+By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exist a (non-trivial) retract of $D_{q}$ through the origin with its tangent direction at 0 belonging to $S^{\prime}$.
+But this contradicts theorem 1.8, which implies that $D_{q}$ does not admit any retract along an open piece of directions at the origin.
+Hence, $D_{q}\setminus A$ cannot be biholomorphic to a product domain.
+By the same reasoning, it follows that $D_{q}$ itself is not biholomorphic to a product domain.
+∎
+We conclude this section with a complete determination of retracts of $D_{h}\setminus A$ through the origin, where $D_{h}$ is the polynomial polyhedron (8.1).
+As above, $A$ is an analytic subset of $D_{h}$ not passing through the origin, which is allowed to be empty but not $D_{h}$ itself.
 This of-course is tantamount to determining retracts of $D_{h}$ by lemma 8.1 and we may therefore forget about $A$.
 Example 8.12.
-The point here is to record at-least one case of a (non-convex but pseudoconvex) domain with $\mathbb{C}$-extremal but not holomorphically extreme boundary; thereby an example where proposition 1.4 does not apply but the modified form of Vesentini’s result as in theorem 1.3 does.
-Indeed, $D_{h}$ serves this purpose as we now discuss.
+We are yet to determine the linear retracts of $D_{h}$ (the motivation for this has already been mentioned in the prequel).
 First, note that almost every point of $\partial D_{h}$ fails to be holomorphically extreme; specifically, every point of $\partial D_{h}$ which is not in the intersection of the real analytic varieties $H_{1}$ given by $|z^{2}-w^{2}|=1$ and $H_{2}$ given by $|zw|=2$.
@@
 Let us begin by settling at the outset, the non-existence of retracts along directions given by $D_{h}\setminus(H_{1}\cap H_{2})$: this follows by noting that every point of this open piece of the boundary is a non-convex boundary point of $D_{h}$ and invoking corollary 2.38.
-With this we may restrict ourselves to points of $H_{1}\cap H_{2}\subset\partial D_{h}$ and proceed to discussing retracts through such points alone; it also being borne in mind that by retracts we mean the linear ones for, by theorem 1.3 there are no non-linear holomorphic retracts through the origin in $D_{h}$.
+With this we may restrict ourselves to points of $H_{1}\cap H_{2}\subset\partial D_{h}$ and proceed to discussing retracts through such points alone; it also being borne in mind that by retracts we mean the linear ones for, by proposition 1.3 there are no non-linear holomorphic retracts through the origin in $D_{h}$.
 To proceed forth indeed then, pick any point $p=(p_{1},p_{2})$ in the intersection $H_{1}\cap H_{2}\subset\partial D_{h}$; so, $|p_{2}^{2}-p_{1}^{2}|=1$ and $|p_{1}p_{2}|=2$.
@@
 Indeed, let just first use the holomorphicity of $\varphi$ to deduce that the aforementioned maximum – call it $m$ – cannot be attained in the open piece $X_{1}\cup X_{2}$ within the boundary $\partial D_{h}$, where
-$X_{1}$ =$=\{|z^{2}-w^{2}|=1\}\cap\{|zw|<2\},$
-$X_{2}$ =$=\{|z^{2}-w^{2}|<1\}\cap\{|zw|=2\}.$
+$X_{1}$ $=\{|z^{2}-w^{2}|=1\}\cap\{|zw|<2\},$
+$X_{2}$ $=\{|z^{2}-w^{2}|<1\}\cap\{|zw|=2\}.$
 To see this, suppose $q=(q_{1},q_{2})$ is a point in $\partial D_{h}$ where $m$ is attained by $|\varphi|$.
@@
 subject to the simultaneous constraints:
-[…]
+(8.4) $\begin{cases}g_{1}(z,w):=|zw|^{2}\;=\;4\\ g_{2}(z,w):=|z^{2}-w^{2}|^{2}\;=\;1.\end{cases}$
 Despite the squares, this is of-course equivalent to the previous maximization problem.
@@
 First we define the middle one, namely the interval:
-$I_{0}=\text{ connected component containing the origin in $\mathbb{R}$ of its subset }\Big{\{}\theta\in\mathbb{R}\;:\;\sin^{2}\theta\leq\frac{1}{16}\Big{\}}.$
+[…]
 More concisely put $I_{0}=[-\epsilon_{0},\epsilon_{0}]$ where $\epsilon_{0}$ stands for the smallest positive number satisfying $\sin^{2}\epsilon_{0}=1/16$.
@@
 to not lose sight of our main goal in the midst of computations, we keep this aside for a moment, and observe that our main goal translates here to checking only whether: $\theta_{0}$ is a point of global maximum for $h$ or not.
-This is due to the following observation: if at all $p$ is a a point of maximum for $f(z,w)=\tilde{f}(r,\theta)$, then $\theta_{0}$ must be a maximum point for $h$.
+This is due to the following observation: if at all $p$ is a point of maximum for $f(z,w)=\tilde{f}(r,\theta)$, then $\theta_{0}$ must be a maximum point for $h$.
 Now, owing to its smoothness on the open interval $I^{0}$ (the interior of $I$), a point of global maximum for $h$ on $I^{0}$ if any, is not only a point of local maximum but also a point about which $h$ is smooth and so, the standard derivative tests of calculus are applicable.
@@
 But then we infer from (8.15) that: $h^{\prime}(\theta_{0})=0$ implies that $\theta_{0}=0,\pm\pi/2\text{ or }\pi$.
-Thus as soon as $\theta_{0}$ is not one among the very few possibilities: 00, ±$\pm\pi/2$, $\pi$, or the remaining end-points of the intervals occurring in the definition of $I$, it cannot even be a point of local extremum for $h$.
+Thus as soon as $\theta_{0}$ is not one among the very few possibilities: 0, $\pm\pi/2$, $\pi$, or the remaining end-points of the intervals occurring in the definition of $I$, it cannot even be a point of local extremum for $h$.
 Moreover, the pair of possibilities $\theta_{0}=\pm\pi/2$ are ruled out simply because $h(\theta_{0})$ is not real and therefore not of relevance for our consideration of the function $h$.
-We presently proceed to dispose off the cases $\theta_{0}=\pm\epsilon_{0}$ where $\epsilon_{0}$ be the smallest positive number such that $\sin^{2}\epsilon_{0}=1/16$ which means that ±$\pm\epsilon_{0}$ are the end-points of the interval $I$; by showing that $h_{\theta_{0}}(\theta)$ does not attain its (global) maximum on $I$ at $\theta=\theta_{0}=\epsilon_{0}$ – the argument for the other case $-\theta_{0}$ being similar is skipped.
+We presently proceed to dispose off the cases $\theta_{0}=\pm\epsilon_{0}$ where $\epsilon_{0}$ be the smallest positive number such that $\sin^{2}\epsilon_{0}=1/16$ which means that $\pm\epsilon_{0}$ are the end-points of the interval $I$; by showing that $h_{\theta_{0}}(\theta)$ does not attain its (global) maximum on $I$ at $\theta=\theta_{0}=\epsilon_{0}$ – the argument for the other case $\theta_{0}$ being similar is skipped.
 Now when $\theta_{0}=\epsilon_{0}$, the expression of $h(\theta)$ simplifies a bit and reads:
@@
 More specifically and precisely, there are no retracts in the directions of $p\in\partial D_{h}\setminus H$, where $H$ is the real analytic manifold cut out by $|z^{2}-w^{2}|=1$ and $|zw|=2$.
-As for points $p\in H$, there is a holomorphic retract in the direction of $p$ if and only if it is of the form $(e^{it_{1}}r_{1},e^{it_{2}}r_{2})$ where $r_{1},r_{2}$ are positive real numbers satisfying $r_{1}^{2}-r_{2}^{2}=\pm 1$ and $r_{1}r_{2}=2$ while $t_{1},t_{2}\in\mathbb{R}$ with $|t_{2}-t_{1}|$ being either 00 or $\pi$ (i.e., the arguments of $p_{1},p_{2}$ are either same or differ by $\pi$); and every such retract is linear.
+As for points $p\in H$, there is a holomorphic retract in the direction of $p$ if and only if it is of the form $(e^{it_{1}}r_{1},e^{it_{2}}r_{2})$ where $r_{1},r_{2}$ are positive real numbers satisfying $r_{1}^{2}-r_{2}^{2}=\pm 1$ and $r_{1}r_{2}=2$ while $t_{1},t_{2}\in\mathbb{R}$ with $|t_{2}-t_{1}|$ being either 0 or $\pi$ (i.e., the arguments of $p_{1},p_{2}$ are either same or differ by $\pi$); and every such retract is linear.
 Let us remark that there are of-course points $p$ in $H$ other than those of the form mentioned in the theorem and note more importantly, that those points are precisely the ones in the boundary of the domain $D_{h}$ which are farthest from the origin.
 Let us briefly justify these claims.
-For the former, note that for any $\alpha_{0}$ satisfying $160\leq\alpha_{0}\leq 1/$ 16, we can find $p=(p_{1},p_{2})=\left(re^{i\theta_{1}},(2/r)e^{i\theta_{2}}\right)\in H_{1}\cap H_{2}$ such that $\sin^{2}\theta_{0}=\alpha_{0}$, where $\theta_{0}=\theta_{1}-\theta_{2}$.
+For the former, note that for any $\alpha_{0}$ satisfying $0\leq\alpha_{0}\leq 1/16$, we can find $p=(p_{1},p_{2})=\left(re^{i\theta_{1}},(2/r)e^{i\theta_{2}}\right)\in H_{1}\cap H_{2}$ such that $\sin^{2}\theta_{0}=\alpha_{0}$, where $\theta_{0}=\theta_{1}-\theta_{2}$.
 Now choose $\theta_{0}\in I$ such that $\sin^{2}\theta_{0}=\alpha_{0}$.
@@
 Remark 8.14.
-The foregoing arguments – again using theorem 1.3 at the outset to reduce considerations from the generality of holomorphic retracts to linear ones – can be adapted to determine precisely the retracts through the origin of other ‘simpler’ examples of non-convex (but pseudoconvex) bounded Reinhardt domains such as $\mathbb{B}\cap H^{2}_{1/4}$ or $B_{p}\cap H^{2}_{1/4}$, where $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/4\}$ and where $B_{p}$ denotes the standard $\ell^{p}$-ball with $1\leq p<\infty$.
+The foregoing arguments – again using proposition 1.3 at the outset to reduce considerations from the generality of holomorphic retracts to linear ones – can be adapted to determine precisely the retracts through the origin of other ‘simpler’ examples of non-convex (but pseudoconvex) bounded Reinhardt domains such as $\mathbb{B}\cap H^{2}_{1/4}$ or $B_{p}\cap H^{2}_{1/4}$, where $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/4\}$ and where $B_{p}$ denotes the standard $\ell^{p}$-ball with $1\leq p<\infty$.
 The case $p=\infty$ is different as in that case (i.e., for $\Delta^{2}\cap H^{2}_{1/4}$) there are non-linear retracts through the origin and therefore, discussed separately below.
@@
 Both of these conditions are fulfilled by $D_{h}$.
-Despite its simplicity, one can still ask natural questions about the choice of $D_{h}$; without elaborating on them all, let us just mention one: why not consider $D_{q}\cap H^{2}_{1/4}$ (where $D_{q}$ denotes the ‘$\ell_{q}$-ball’ for $10<q<$ 1 as before), as this is also a maximally non-convex domain?
+Despite its simplicity, one can still ask natural questions about the choice of $D_{h}$; without elaborating on them all, let us just mention one: why not consider $D_{q}\cap H^{2}_{1/4}$ (where $D_{q}$ denotes the $\ell_{q}$-ball’ for $0<q<1$ as before), as this is also a maximally non-convex domain?
 Our answer to this is that though this is easier to analyse owing to it being Reinhardt, it turns out that through every non-smooth point of $D_{q}\cap H^{2}_{1/4}$ there is a linear retract and is a convex boundary point whereas this is not the case for $D_{h}$: not every non-smooth point of $\partial D_{h}$ is a point of convexity, thereby is an even better example of a maximally non-convex domain.
@@
 Consider the map $(z,w)\mapsto(z^{2},w)$; it effects a proper map from the ellipsoid $E_{4}=\{(z,w)\;:\;|z|^{4}+|w|^{2}<1\}$ to the Euclidean ball $\mathbb{B}$.
-Note that the linear subspace given by $w=z$ is retract of $E_{4}$ (as every one-dimensional linear subspace is).
+Note that the linear subspace given by $w=z$ is a retract of $E_{4}$ (as every one-dimensional linear subspace is).
 However, its image $\{z=w^{2}\}$ being the graph of a non-linear map is not a retract of $\mathbb{B}^{2}$.
+We now recall some of the basic definitions associated with analytic polyhedra; based on section 2.3 of [Forst_Stein], [petrosyan] and [Zim_generic].
+Analytic polyhedra are obtained by looking at subsets of a complex manifold $X$ occurring as the intersection of sublevel sets of moduli of holomorphic functions on $X$.
+Several issues arise here: firstly if $X$ is compact, the interior of such an intersection is empty; even if $X$ is non-compact and a Stein manifold, the intersection can have several connected components.
+We get around such issues by phrasing the definition as follows; we note that there are some variations in the definitions followed by different authors and we are laying the one that we shall adopt, for this reason as well.
+A relatively compact domain in a complex manifold $X$ of the form
+(8.18) $D=\{x\in U:|f_{j}(x)|<1~\text{for}~j=1,2,\ldots,m\}$
+where the $f_{j}$’s are holomorphic functions on some open subset $U\supset\overline{D}$, is said to be an analytic polyhedron.
+By definition then, $\overline{D}$ is compact and connected.
+We shall refer to the $f_{j}$’s as defining functions, though it must be noted that such a system of functions are not at all uniquely determined (as is seen immediately by taking their powers for instance); nevertheless for convenience, we shall refer to them as ‘the’ defining functions.
+By adding more functions to the list of defining functions (the $f_{j}$’s above), we can ensure that $F=(f_{1},\ldots,f_{m}):U\to\mathbb{C}^{m}$ embeds $D$ properly into the polydisc $\Delta^{m}\subset\mathbb{C}^{m}$.
+If $D$ is a domain in $\mathbb{C}^{N}$, this follows by considering the map $G:U\to\mathbb{C}^{m+N}$ defined by $G(z)=(f_{1}(z),\ldots,f_{m}(z),z_{1}/R_{1},\ldots,z_{N}/R_{N})$, where $R_{1},\ldots,R_{N}$ are positive reals such that $D\subset\Delta^{m}(0;R_{1},\ldots,R_{N})$, which exist by virtue of the boundedness of $D$ here.
+Extending this result to the setting of complex manifolds requires additional work but the result is well-known (see section 4.5.3 of [Noguchi] and section 2.3 in [Forst_Stein]).
+If the defining (analytic) functions can be chosen so that their number equals the dimension of the manifold i.e., $m=N$, then $D$ is called a special analytic polyhedron.
+Our case of first interest is $U$ being an open set in $\mathbb{C}^{N}$.
+Note then that as per the above stated definition of analytic polyhedron (which is the one that we shall adopt), $D$ is bounded; though a couple of observations in the sequel do not require boundedness, most of our main results do.
+More importantly, it may very happen that the (topological) closure $\bar{D}$ may fall short of being equal to the ‘algebraic/order’-closure:
+$\overline{D}^{alg}:=\{x\in U\;:\;|f_{j}(x)|\leq 1\text{ for }j=1,2,\ldots,m\}$
+– for concrete examples, see [closure_anal_poly].
+We shall see that when $D$ is balanced, the closures in the topological and algebraic senses coincide.
+Towards this and other main results, we require a couple of more definitions.
+If all the defining functions $f_{j}$’s can be chosen to be polynomials, then $D$ is called a polynomial polyhedron.
+Next, in all cases, we have the following.
+The boundary $\partial D$ of $D$ consists of the closed faces $\sigma_{j}:=\{z\in\partial D:|f_{j}(z)|=1\}$ intersecting along the level-$r$ ribs $\sigma_{j_{1},\ldots,j_{r}}:=\sigma_{j_{1}}\cap\ldots\cap\sigma_{j_{r}}$.
+That is, the boundary $\partial D$ can be expressed as the union of the faces; this is based on the fact that $\overline{D}\subset\{z\in U:|f_{j}(z)|\leq 1~\text{for}~j=1,2,\ldots,m\}$ which implies that $\partial D\subset\cup_{j=1}^{m}\{z\in U:|f_{j}(z)|=1\text{ and }|f_{k}(z)|\leq 1,\text{ for all }k\neq j\}=:{\partial D}^{alg}$.
+Associated with the faces $\sigma_{j}$, we need to introduce the following sets:
+$\hat{\sigma}_{j}:=\{z\in\mathbb{C}^{N}:|f_{j}(z)|=1\}=Z(|f_{j}(z)|^{2}-1),$
+where $Z(|f_{j}(z)|^{2}-1)$ denotes the zero set of $|f_{j}(z)|^{2}-1$.
+$\tilde{\sigma}_{j}:=\{z\in\partial D:|f_{j}(z)|=1,~|f_{l}(z)|<1~\text{for}~1\leq l\leq m~\text{with}~l\neq j\}$
+Note that $\sigma_{j}=\hat{\sigma}_{j}\cap\partial D$.
+The reader may find the articles [Singularity], [Real_Var],[Strat] to provide helpful background information concerning the structure of real analytic varieties and semi-varieties, though we do not make explicit use of their general structure results here.
+Following the definition from [Zim_generic], the analytic polyhedron $D$ is said to be generic (non-degenerate) if we can choose the functions $f_{1},\ldots,f_{m}$ so that whenever $\zeta\in\sigma_{j_{1},\ldots,j_{r}}$, the vectors
+$\nabla f_{j_{1}}(\zeta),\ldots,\nabla f_{j_{r}}(\zeta)$
+are $\mathbb{C}$-linearly independent.
+We shall refer to this as the genericity condition on the $f_{j}$’s.
+The symbol $\nabla$ denotes the complex gradient i.e., the complex derivative.
+Though this notation is more common for real differentiable functions.
+We use this here follows Zimmer [Zim_generic].
+The context will make it clear whether we use complex or real gradient.
+Remark 8.15.
+  1. (1)
+    Every linear retract of an analytic polyhedron is again an analytic polyhedron because if $D_{L}$ is a linear retract of $D$ for some linear subspace $L$ of $\mathbb{C}^{N}$ then $D_{L}=\{z\in U\cap L:|f_{j}(z)|<1~\text{for}~1\leq j\leq m\}$ is an analytic polyhedron on $L$.
+  2. (2)
+    Note that every analytic polyhedron $D$ in any complex manifold $X$ is Stein. Needless to say when $X=\mathbb{C}^{N}$, $D$ is a pseudoconvex domain.
+  3. (3)
+    Every (bounded) analytic polyhedon $D$ is hyperconvex because if $|f_{j}(p)|=1$ for some $j$ then $\varphi(z):=|f_{j}(z)|^{2}-1$ is a plurisubharmonic function on $D$ and $\lim_{D\ni z\to p}\varphi(z)=0$ (this suffices to verify hyperconvexity by the remark 2.51 in the preliminaries section 2.7).
+  4. (4)
+    Product of two analytic polyhedra is again an analytic polyhedron because if $D_{1}:=\{z\in U_{1}:|f_{j}(z)|<1~\text{for}~j=1,\ldots,m_{1}\},~~D_{2}:=\{w\in U_{2}:|g_{k}(z)|<1~\text{for}~k=1,\ldots,m_{2}\}$, where $f_{j}$’s are holomorphic on an open set $U_{1}$ containing $\overline{D}_{1}$ and $g_{k}$’s are holomorphic on an open set $U_{2}$ containing $\overline{D}_{2}$ then $D_{1}\times D_{2}:=\{(z,w)\in U_{1}\times U_{2}:|f_{j}(z)|<1,~~~~|g_{k}(w)|<1~\text{for}~j=1,\ldots,m_{1},~k=1,2,\ldots,m_{2}\}$.
+Lemma 8.16.
+Let $D$ be a (bounded) balanced analytic polyhedron in $\mathbb{C}^{N}$ given by
+$D:=\{z\in U:|f_{j}(z)|<1~\text{for}~j=1,2,\ldots,m\}$
+(where the $f_{j}$’s are holomorphic functions on some domain $U\subset\mathbb{C}^{N}$ containing $\overline{D}$).
+Suppose also that the system of defining functions $(f_{j})$ is minimal, i.e.
+no function $f_{j}$ can be removed without changing $D$.For each $j~\text{with}~1\leq j\leq m$, consider the analytic variety $\hat{\sigma}_{j}=Z(|f_{j}|^{2}-1)$.
+If $p\in\tilde{\sigma}_{j}$ is a smooth point of $\hat{\sigma}_{j}$ (recall that smooth points of $\hat{\sigma}_{j}$ form a dense subset of $\hat{\sigma}_{j}$.) for some $j$, then $p$ is a regular point of $\hat{\sigma}_{j}$ of dimension 2N12N-1.
+Consequently, for each $j$, $\text{dim}(\hat{\sigma}_{j})=2N-1$.
+Proof.
+[…]
+(8.19) $h(q_{n})=h\big{(}\frac{1}{\alpha_{n}}p_{n}\big{)}=\frac{1}{\alpha_{n}}=\frac{1}{\|p_{n}\|}$
+[…]
+Proposition 8.17.
+Let $D$ be a (bounded) balanced analytic polyhedron in $\mathbb{C}^{N}$ given by
+$D:=\{z\in U:|f_{j}(z)|<1~\text{for}~j=1,2,\ldots,m\}$
+(where the $f_{j}$’s are holomorphic functions on some domain $U\subset\mathbb{C}^{N}$ containing $\overline{D}$.) Suppose also that the system of defining functions $(f_{j})$ is minimal.
+Then we can choose the $f_{j}$’s to be homogeneous polynomials (so $D$ is a balanced polynomial polyhedron).
+Proof.
+First, we prove this proposition under the assumption that $D$ is a generic analytic polyhedron because the proof is more elementary in this case.
+We can express $D$ as $D=F^{-1}(\Delta^{m})$, where $F:U\to\mathbb{C}^{m}$ defined by $F(z)=(f_{1}(z),\ldots,f_{m}(z))$.
+Let $F(0)=(a_{1},\ldots,a_{m})$.
+As $a$ need not be zero, we compose $F$ with the automorphism $\varphi$ of $\Delta^{m}$ given by
+$\varphi(z)=\left(\frac{z_{1}-a_{1}}{1-\overline{a_{1}}z_{1}},\ldots,\frac{z_{m}-a_{m}}{1-\overline{a_{m}}z_{m}}\right).$
+Then $\varphi\circ F(0)=0$. After possibly shrinking $U$, to avoid $1/\overline{a_{j}}$ for all $j$ with $1\leq j\leq m$, define a holomorphic function $g_{j}:U\to\mathbb{C}$ by
+$g_{j}(z)=\frac{f_{j}(z)-a_{j}}{1-\overline{a_{j}}f_{j}(z)}.$
+Since $D=(\varphi\circ F)^{-1}(\Delta^{m})$, it follows that
+$D=\{z\in U:|g_{j}(z)|<1~\text{for}~j=1,2,\ldots,m\}.$
+Moreover, $D\varphi_{|_{0}}$ is an invertible linear map and therefore preserves linear independence.
+It follows that $g_{j}$’s satisfy the genericity condition.
+In the rest of the proof, we relabel these functions as $(f_{j})$ and regard them as the defining functions for $D$.
+Fix $z_{0}\in\tilde{\sigma}_{1}$. We can divide the proof into the following steps.
+Step-1: There exists positive integer $d(z_{0})$ such that for each $\lambda\in\Delta$, we have $f_{1}(\lambda z_{0})=\lambda^{d(z_{0})}B(\lambda,z_{0})$, where $B(\lambda,z_{0})$ is some Blaschke product on $\Delta$ with $B(0,z_{0})\neq 0$.
+Consider the holomorphic function $h$ on $\Delta$ defined by $h^{z_{0}}(\lambda)=f_{1}(\lambda z_{0})$.
+Note that $h^{z_{0}}$ maps $\Delta$ into $\Delta$ because $D$ is a balanced hyperconvex domain.
+We want to show that $h^{z_{0}}$ is a proper map.
+To prove this, for each $k$ with $1\leq k\leq m$, define
+$A_{k}:=\{\theta\in[-\pi,\pi]:|f_{k}(e^{i\theta}z_{0})|=1\}.$
+As $D$ is balanced, it follows that
+$\pi,\pi]=\bigcup_{k=1}^{m}A_{k}.$
+Now, we claim that the sets $A_{k}$ are pairwise disjoint.
+To prove this claim, consider the invertible $\mathbb{C}$-linear map $T:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $T(z)=e^{i\theta}z$ for some $\theta\in[-\pi,\pi]$.
+Note that $T$ maps $D$ into $D$, as $D$ is balanced.
+Since $T$ is an invertible linear transformation, $T$ maps $\partial D$ into $\partial D$.
+Note that $T$ preserves smooth boundary points of $D$ because if $p$ is a smooth boundary point of $D$ and $r$ be its smooth defining function near $p$ then $\tilde{r}(z)=r(e^{-i\theta}z)$ serves as a smooth defining function near $T(p)$.
+This finishes the proof of the claim that the sets $A_{k}$ are pairwise disjoint.
+Hence $A_{k}^{c}=\cup_{j=1,j\neq k}^{m}A_{j}$.
+Since the $A_{j}$’s are closed subsets of $\pi,\pi]$, $A_{k}^{c}$ is closed.
+This implies that $A_{k}$ is both open and closed subset of $\pi,\pi]$.
+As $\pi,\pi]$ is connected, either $A_{k}=\phi$ or $\pi,\pi]$.
+Thus there is atleast one $j$ such that $A_{j}=[-\pi,\pi]$.
+Without loss of generality we may assume that $j=1$, i.e.
+$|f_{1}(\lambda z_{0})|=1$ whenever $|\lambda|=1$.
+Hence, we obtain that $h^{z_{0}}$ is a proper map on $\Delta$.
+Therefore $h^{z_{0}}$ is of the form $h^{z_{0}}(\lambda)=B(\lambda,z_{0})$, where $B(\lambda,z_{0})$ is some finite Blaschke product.
+Since $h^{z_{0}}(0)=0$, there exists positive integer $d(z_{0})$ such that
+(8.20) $f_{1}(\lambda z_{0})=h^{z_{0}}(\lambda)=B(\lambda,z_{0})=\lambda^{d(z_{0})}\frac{P(\lambda,z_{0})}{Q(\lambda,z_{0})}$
+where $P(\lambda,z_{0})$ and $Q(\lambda,z_{0})$ are polynomials in $\lambda$, which does not vanishes at $\lambda=0$, and ${\rm deg}(P(\lambda,z_{0}))={\rm deg}(Q(\lambda,z_{0}))$.
+Since $|f_{1}(z_{0})|=1$, after multiplying $f_{1}$ by an unimodular constant, we may assume that $f_{1}(\lambda z_{0})=B(\lambda,z_{0})f_{1}(z_{0})$.
+Step-2: There exists $z^{\prime}\in\tilde{\sigma}_{1}$ such that $z^{\prime}$ is a growth vector of $f_{1}$ at 0 and $f_{1}$ does not vanish on $\Delta_{z^{\prime}}^{*}:=\{\lambda z^{\prime}:|\lambda|\leq 1,~\lambda\neq 0\}$.
+Let $Z_{f_{1}}$ denotes the zero set of $f_{1}$ and let $C^{t}:=C(Z_{f_{1}},0)$ denotes the tangent cone to $Z_{f_{1}}$ at the origin.
+By theorem 2.58, the set of all non-growth vectors for $f_{1}$ at 0 is precisely equal to the tangent cone $C^{t}$.
+Let $A:=C^{t}\cap\tilde{\sigma}_{1}$.
+First, we want to show that $A$ is a nowhere dense subset of $\tilde{\sigma}_{1}$.
+By proposition 2.55, $C^{t}$ is a closed nowhere dense subset in $\mathbb{C}^{N}$.
+Consider the radial projection $\pi^{r}:\mathbb{C}^{N}\setminus\{0\}\to\partial\mathbb{B}$ defined by $\pi^{r}(z)=z/\|z\|$, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{C}^{N}$.
+Note that $\pi^{r}(C^{t})=C^{t}\cap\partial\mathbb{B}$ because the cone generated by $C^{t}\cap\partial\mathbb{B}$ is precisely equal to $C^{t}$.
+Since $dim(C^{t})=2N-2$ is less than $dim(\partial\mathbb{B})$, $C^{t}\cap\partial\mathbb{B}$ is a nowhere dense subset of $\partial\mathbb{B}$.
+Owing to the hyperconvexity of $D$, $\pi^{r}_{|_{\partial D}}:\partial D\to\partial\mathbb{B}$ is a homeomorphism.
+Hence $(\pi^{r})^{-1}(C^{t}\cap\partial\mathbb{B})$ is a nowhere dense subset of $\partial D$.
+This shows that $A$ is a nowhere dense subset of $\tilde{\sigma}_{1}$.
+Let $\tau_{1}:=\tilde{\sigma}_{1}\setminus A$.
+To prove the claim in step-2, it suffices to show that there exists $z^{\prime}\in\tau_{1}$ such that $f_{1}$ does not vanish on $\Delta_{z^{\prime}}^{*}$.
+To prove this by contradiction, assume that for every $z\in\tau_{1}$, $f_{1}$ vanishes at some point in $\Delta_{z}^{*}$.
+This implies that for every $z\in\tau_{1}$, $f_{1}$ vanishes at some point in $\Delta_{z}^{*}$.
+It therefore follows that $\pi^{r}(Z_{f_{1}}^{*})\supseteq\pi^{r}(\tau_{1})$, where $Z_{f_{1}}^{*}:=Z_{f_{1}}\setminus\{0\}$.
+Denote by $Z_{f_{1}}^{\text{reg}},Z_{f_{1}}^{\text{s}}$ respectively, the regular and singular points of $Z_{f_{1}}^{*}$.
+Hence $\pi^{r}(Z_{f_{1}}^{*}\setminus Z_{f_{1}}^{s})\supseteq\pi^{r}(\tau_{1})\setminus\pi^{r}(Z_{f_{1}}^{s})$.
+By the principle of invariance of domain, it follows that $\pi^{r}(\tau_{1})$ is an open subset of $\partial\mathbb{B}$.
+Thereby, in particular, its Hausdorff dimension is 2N12N-1.
+By Corollary 1 of section 2.6 in [Chirka], the Hausdorff dimension of $\pi^{r}(Z_{f_{1}}^{s})$ is at most 2N42N-4.
+It follows that $f_{1}$ vanishes on a set of positive (2N1)(2N-1)-dimensional Hausdorff measure.
+Hence $f_{1}$ vanishes on $U$.
+This contradicts the fact that $f_{1}$ is a non-constant function, which finishes the proof of Step-2.
+Step-3: $f_{1}$ is a homogeneous polynomial on $\mathbb{C}^{N}$.
+By the above Step-2, there exists a growth vector $z^{\prime}\in\tilde{\sigma}_{1}$ of $f_{1}$ at 0 and $f_{1}$ does not vanish on $\Delta_{z^{\prime}}^{*}$.
+By proposition 2.55, the tangent cone $C^{t}$ is a closed nowhere dense subset of $\mathbb{C}^{N}$.
+Choose an open set $V$ around $z^{\prime}$ of $\tilde{\sigma}_{1}$ such that $\overline{V}$ is contained in the complement of $C(Z_{f_{1}},0)$.
+Consider the punctured (complex) cone $C_{V}^{\prime}$ generated by $V$.
+Note that $\overline{C^{\prime}}_{V}$ is contained in the complement of $C^{t}$.
+Now, we want to prove that there exists an open set $U^{\prime}$ of $D$ containing 0 such that $Z_{f_{1}}^{*}$ does not intersect $U^{\prime}\cap C_{V}^{\prime}$.
+To prove this by contradiction, assume that it so happens that for every neighbourhood $U^{\prime}$ of the origin, $Z_{f_{1}}^{*}\cap(U^{\prime}\cap C_{V}^{\prime})$ is non-empty.
+This gives rise to the existence of a sequence $q_{j}\subset Z_{f_{1}}^{*}\cap(U^{\prime}\cap C_{V}^{\prime})$ such that $q_{j}\to 0$.
+Let $p_{j}:=q_{j}/h(q_{j})$, where $h$ is the Minkowski functional of $D$.
+Since $(p_{j})$ is bounded, it has a convergent subsequence $(p_{j_{k}})$, which converges to some (growth vector) $v$ lying in $\overline{V}$.
+By definition of tangent cones, this implies that $v\in C(Z_{f_{1}},0)$, which contradicts $\overline{V}$ is disjoint from $C^{t}$.
+Therefore, there exists an open neighbourhood $U^{\prime}\subset D$ of 0 such that $Z_{f_{1}}^{*}\cap(U^{\prime}\cap C_{V}^{\prime})=\phi$.
+We shall now show that $f_{1}$ does not vanish in outside of $U^{\prime}$ if only we pass to a possibly smaller cone.
+To work this out rigorously, first choose $r>0$ such that $\mathbb{B}(0,r)\subset U^{\prime}$.
+Denote the closed annulus $\mathbb{A}[0;r,1]$ by $B$ for short.
+Consider the function $\tilde{f}_{1}(\lambda,z)=f_{1}(\lambda z)$, which is in particular continuous on the compact set $B\times\overline{V}$.
+Since $f_{1}$ is a non-vanishing function on $\Delta_{z^{\prime}}^{*}$, there exists $\epsilon>0$ such that
+$|f_{1}(\lambda z^{\prime})|>\epsilon~\text{for all}~\lambda\in\mathbb{A}[0;r,1].$
+By the uniform continuity of $\tilde{f}_{1}$ on $B\times\overline{V}$, there exists $\delta>0$ such that
+$\|\tilde{f}_{1}(\mu,z^{\prime})-\tilde{f}_{1}(\lambda,z)\|<\epsilon/2~\text{whenever}~\|(\lambda,z)-(\mu,z^{\prime})\|<\delta.$
+In particular, if $\|z-z^{\prime}\|<\delta$, then $\|\tilde{f}_{1}(\lambda,z^{\prime})-\tilde{f}_{1}(\lambda,z)\|<\epsilon/2$ for all $\lambda\in B$. By triangle inequality, we have
+$\|\tilde{f}_{1}(\lambda,z)\|\geq\|\tilde{f}_{1}(\lambda,z^{\prime})\|-\|\tilde{f}_{1}(\lambda,z^{\prime})-\tilde{f}_{1}(\lambda,z)\|>\epsilon/2$
+whenever $\|z-z^{\prime}\|<\delta$ and $z\in\overline{V}$. Define the punctured cone
+$C_{\delta}^{\prime}(z^{\prime}):=\{\lambda z:\|z-z^{\prime}\|<\delta,~z\in\partial C_{V}^{\prime}\cap\partial D,~\lambda\in\Delta^{*}\}.$
+Then note that $f_{1}$ is non-vanishing on $K^{\prime}:=C_{\delta}^{\prime}(z^{\prime})\setminus U^{\prime}$ and hence on all of $C_{\delta}^{\prime}$ as $Z_{f_{1}}$ does not intersect $U^{\prime}$ within $C_{\delta}^{\prime}(z^{\prime})$.
+Since every vector in $C_{\delta}^{\prime}$ is a growth vector of $f_{1}$ at 0, it follows from 8.20 now that $f_{1}(\lambda z)=\lambda^{d(z^{\prime})}f_{1}(z)$ for all $\lambda\in\Delta$.
+This shows that $d(z)=d(z^{\prime})$ for all $z$ near $z^{\prime}$, i.e.
+$d(z)$ is constant on $C_{\delta}^{\prime}$.
+By the identity principle, $f_{1}(\lambda z)=\lambda^{d}f_{1}(z)$ for all $z\in U$, $\lambda\in\Delta$.
+Using this functional equation, we can extend $f_{1}$ to all of $\mathbb{C}^{N}$ by defining $f_{1}(\lambda z)=\lambda^{d}f_{1}(z)$.
+It follows, therefore that $f_{1}$ is a homogeneous entire function, thereby a homogeneous polynomial.
+So, $f_{j}$ is a homogeneous polynomial for each $j=1,2,\ldots,m$.
+We now consider the case that $D$ is not necessarily a generic analytic polyhedron.
+As in the above case, we may assume that $f_{j}(0)=0$ and that the collection of defining functions $(f_{j})$ is minimal in the sense that no proper subset of $(f_{j})$ suffices to define $D$.
+The proof for this case is similar to the proof for the above case except for Step-2.
+Note that in this case, $\tilde{\sigma}_{1}$ need not be a manifold.
+So, we consider the analytic variety $\hat{\sigma}_{1}=Z(|f_{1}|^{2}-1)$.
+This analytic variety $\hat{\sigma}_{1}$ has a well-defined dimension $d$, which can at most be 2N12N-1, where $d$ is maximal among the dimensions of smooth points, as laid down precisely by definition 2.59 in the preliminaries section.
+Note that $\tilde{\sigma}_{1}^{reg}:=\hat{\sigma}_{1}^{reg}\cap\tilde{\sigma}_{1}$ — recall the notation and basic facts about such points in analytic variety from the preliminaries section 2.9 — is a smooth submanifold of real dimension $d$ (for further details regarding the structure of analytic varieties and semi-varieties, we refer the reader to [Strat]).
+By the above lemma 8.16, $\text{dim}(\tilde{\sigma}_{1}^{reg})=2N-1$.
+Now, we apply a similar argument in Step-2 to $\tilde{\sigma}_{1}^{reg}$ replacing $\tilde{\sigma}_{1}$ therein.
+Hence, we get the desired result of Step-2.
+As noted earlier, this was all that was required to finish the proof in the general case.
+∎
+Remark 8.18.
+Let $D$ be any balanced analytic polyhedron.
+So, by the above proposition, $D:=\{z\in\mathbb{C}^{N}:|f_{j}(z)|<1~\text{for}~j=1,2,\ldots,m\}$, where each $f_{j}$ is a homogeneous polynomial of degree $d_{j}:={\rm deg}(f_{j})$.
+Define $h:\mathbb{C}^{N}\to\mathbb{R}$ by $h(z):=\max\{|f_{j}(z)|^{1/d_{j}}:1\leq j\leq m\}$.
+Note that $h$ is a continuous function on $\mathbb{C}^{N}$ and $h(\lambda z)=|\lambda|h(z)$.
+Appealing to proposition 2.9 (c), we conclude that $h$ is the Minkowski functional on $D$.
+Furthermore, by proposition 2.9 (d), we therefore deduce that the topological closure equals the algebraic closure: $\overline{D}=\{z\in\mathbb{C}^{N}:|f_{j}(z)|\leq 1~\text{for}~j=1,2,\ldots,m\}$.
+Lemma 8.19.
+Let $f\in\mathbb{C}[z_{1},\ldots,z_{N}]$ be a homogeneous polynomial. If $|f|$ is a non-zero constant on a complex hyperplane, then $f$ is a power of some linear polynomial.
+Proof.
+Assume that $|f|$ is a non-zero constant, say $c$, on some complex hyperplane $l$.
+Observe that this, in particular, means $l$ does not pass through the origin.
+Parametrize $l$ in the standard manner as $\varphi(\zeta_{1},\ldots,\zeta_{N-1})=p+\zeta_{1}v_{1}+\ldots+\zeta_{N-1}v_{N-1}$ for some choice of a point $p$ on $l$ and non-zero vectors $v_{j}\in\mathbb{C}^{N}$.
+The maximum modulus principle (for holomorphic functions of several complex variables) applied to the holomorphic function $f\circ\varphi$ confirms that $f$ is constant (=c=c) on $l$.
+This implies that $f\circ\varphi(\zeta)$ is a non-zero constant, where $\varphi(\zeta_{1},\ldots,\zeta_{N-1})=p+\zeta_{1}v_{1}+\ldots+\zeta_{N-1}v_{N-1}$ is the parametrization of $l$ passing through $p$ in the direction of vectors $v_{j}$.
+After multiplying by $1/c$, we may assume that $f$ is identically equal to $1$ on $l$.
+We can also express $l$ as the zero set Z(g1)Z(g-1), where $g$ is a homogeneous linear polynomial in $\mathbb{C}[z_{1},\ldots,z_{N}]$.
+Let $\tilde{f}(z_{1},\ldots,z_{N}):=f(z_{1},\ldots,z_{N})-1$.
+Note that $\tilde{f}$ vanishes on Z(g1)Z(g-1).
+By an application of the Hilbert Nullstellensatz, there exists $k\in\mathbb{N}$ such that g1g-1 divides $\tilde{f}^{k}$.
+Being a linear polynomial g1g-1 is irreducible in (the unique factorization domain) $\mathbb{C}[z_{1},\ldots,z_{N}]$, thereby a prime element therein.
+This implies that g1g-1 divides $\tilde{f}$.
+So, we write $\tilde{f}=(g-1)\tilde{h}$ for some non-zero polynomial $\tilde{h}$.
+Substituting $\tilde{f}$ by (f1)(f-1), we obtain that
+$f=(g-1)\tilde{h}+1.$
+Note that $\tilde{h}(0)=1$ because $f(0)=0$.
+So, we can write $\tilde{h}(z_{1},\ldots,z_{N})=h(z_{1},\ldots,z_{N})+1$, where $h\in\mathbb{C}[z_{1},\ldots,z_{N}]$ with $h(0)=0$.
+Hence, the above equation reduces to f=gh+ghf=gh+g-h.
+Write $h$ in its standard homogeneous decomposition: $h=h_{1}+\ldots+h_{r}$, where $h_{i}$ is a homogeneous polynomial of degree $i$ (so the integer $r$ is one less than the degree of $f$).
+Note that $gh_{i}$ is of degree $i$ for each $i$.
+Thus, the above equation may be rewritten as
+$f=g(h_{1}+\ldots+h_{r})+g-(h_{1}+\ldots+h_{r})=(g-h_{1})+(gh_{1}-h_{2})+\ldots+(gh_{r-1}-h_{r})+gh_{r}$
+As $f$ is a homogeneous polynomial of degree $r+1$, we infer that the first $r$ homogeneous components (i.e., all except the last) are identically zero. We therefore conclude
+$h_{1}=g,~h_{2}=gh_{1},\ldots,\;h_{r}=gh_{r-1},\;\;gh_{r}=f.$
+This implies that $h=g+g^{2}+\ldots+g^{r}$ and $f=g^{r+1}$. ∎
+Remark 8.20.
+Let us note a special case for future use.
+If $f$ is a homogeneous polynomial in $z$ and $w$ (i.e.
+$N=2$) then the level set of $|f|$ does not contain any complex line unless $f$ is a power of a linear polynomial.
+Consequently, a homogeneous analytic polyhedron in $\mathbb{C}^{2}$ is $\mathbb{C}$-extremal i.e.
+$\partial D$ does not contain any (affine) $\mathbb{C}$-line segments iff for each $j$, $f_{j}$ is not a power of a linear polynomial.
+Proof of theorem 1.9
+.
+To prove this theorem by contradiction, assume that there exists a non-trivial retract $Z$ of $D$ passing through the origin in the direction of $p\in L:=T_{0}Z$.
+Let $\rho$ be a retraction map from $D$ onto $Z$.
+By lemma 3.8, $T:=D\rho_{|_{0}}$ is a linear projection map from $\mathbb{C}^{N}$ onto $L$ and $T$ maps $D$ onto $D_{L}:=D\cap L$.
+Let $H_{p}:=p+H_{0}$, where $H_{0}:={\rm\ker}(T)$.
+Note that the affine $\mathbb{C}$-linear subspace $H_{p}$ does not intersect $D$ because if $q\in H_{p}\cap D$ then $T(q)=p$, which contradicts the fact that $T$ maps $D$ into $D$.
+Since $\text{dim}(\ker(T))\geq 1$, there exist a non-zero vector $v\in\ker(T)$ such that the complex line $\ell_{p}:=\{p+\zeta v:\zeta\in\mathbb{C}\}$ does not intersect $D$.
+By the hypothesis, $p$ belongs to one of the open faces, i.e., $p\in\tilde{\sigma}_{j}$ for some $j$.
+This implies that $|f_{j}(p)|=1$ and $|f_{l}(p)|<1$ for $l\neq j$.
+Consider the holomorphic function $\psi:\mathbb{C}\to\mathbb{C}$ defined by $\psi(\zeta)=f_{j}(p+\zeta v)$.
+Since $f_{l}$’s are continuous, there exists $r>0$ such that $|f_{l}(p+\zeta v)|<1$ for all $\zeta\in\Delta_{r}$ and all $l$ with $l\neq j$.
+Hence, $|f_{j}(p+\zeta v)|\geq 1$ for all $\zeta\in\Delta_{r}$ because $\ell_{p}$ does not intersect $D$.
+By the minimum principle available for the non-vanishing holomorphic function $\psi$, it follows that $|\psi(\zeta)|=1$ for all $\zeta\in\Delta_{r}$.
+This contradicts our assumption that $\partial D$ is $\mathbb{C}$-extremal.∎
+Lemma 8.21.
+  1. (1)
+    Let $f$ be a non-constant homogeneous polynomial in $\mathbb{C}^{N}$.
+Consider the real algebraic variety $A:=Z(r(z))$, where $r(z)=|f(z)|^{2}-c^{2}$ for some $c\in\mathbb{R}^{+}$.
+Then $\nabla r$ is nowhere vanishing on an open dense subset of $A$.
+  2. (2)
+    Let $D$ be a balanced analytic polyhedron in $\mathbb{C}^{N}$ and let $(f_{j})_{j=1}^{m}$ be minimal defining functions for $D$.
+If $r_{j}(z):=|f_{j}(z)|^{2}-1$ then $\nabla r_{j}$ is nowhere vanishing on an open dense subset of $\sigma_{j}$.
+Proof.
+  1. (1)
+    Firstly note that $\nabla r(z)=|f(z)|\overline{\nabla f(z)}$.
+To prove (1), it suffices to show that $\nabla f$ never vanishes on an open dense subset of $A$, since $|f(z)|=1$ for all $z\in A$.
+Fix $\theta\in[-\pi,\pi]$.
+Consider the complex analytic hypersurface
+    $L_{\theta}:=\{z\in\mathbb{C}^{N}:f(z)=ce^{i\theta}\}.$
+    Suppose $\nabla f(z)$ vanishes on some open subset $W:=B\cap A$ of $A$.
+In particular, $\nabla f$ vanishes on an open subset $V:=W\cap L_{\theta}$ of $L_{\theta}$.
+Since $\nabla f$ is homogeneous, $\nabla f$ vanishes on the complex cone $C_{V}$ generated by $V$.
+It follows from the Weierstrass preparation theorem that $V$ is biholomorphic to an open subset of $\mathbb{C}^{2N-2}$.
+Hence, $C_{V}$ is an open subset of $\mathbb{C}^{N}$.
+By the identity principle applied to $\nabla f$, we obtain that $\nabla f$ vanishes on $\mathbb{C}^{N}$.
+Hence, $f$ is a constant function, which is a contradiction.
+  2. (2)
+    By the above proposition 8.17, for each $j$, $f_{j}$ is a homogenous polynomial in $\mathbb{C}^{N}$.
+Hence by (1), $\nabla f_{j}$ never vanishes on an open dense subset of $Z(|f_{j}|^{2}-1)$.
+This implies that $\nabla r_{j}$ never vanishes on an open dense subset of $\sigma_{j}$.
+Theorem 8.22.
+Let $D$ be a (bounded) balanced $\mathbb{C}$-extremal analytic polyhedron in $\mathbb{C}^{N}$. Then $D$ does not admit any retract along any open piece of directions at the origin.
+Proof.
+First, we assume that $D$ is a generic analytic polyhedron.
+To prove this theorem by contradiction, assume that there exists a non-empty open subset $\Gamma$ of $\partial\mathbb{B}$ such that for each $v\in\Gamma$ there is a retract $Z$ of $D$ passing through the origin with $v\in T_{0}Z$.
+Consider the map $\pi^{r}:\mathbb{C}^{N}\setminus\{0\}\to\partial\mathbb{B}$ defined by $\pi^{r}(z)=z/\|z\|$, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{C}^{N}$.
+Note that $\pi^{r}$ is a smooth map on $\mathbb{C}^{N}\setminus\{0\}$.
+As $D$ is bounded, it follows from the continuity of $\pi^{r}$ that $V:=(\pi^{r})^{-1}(\Gamma)\cap\partial D$ is a non-empty open subset of $\partial D$.
+By the theorem 1.9, $V$ is contained in the union of all ribs $\sigma_{j,k}$.
+Note that $\sigma_{j,k}$ is a smooth manifold of dimension 2N22N-2 , due to the genericity condition of $D$.
+As ${\rm dim}(\sigma_{j,k})<{\rm dim}(\partial\mathbb{B})$, it follows from Sards theorem (see specifically corollary 6.11 in [J.Lee]) applied to the restriction $\pi^{r}_{|_{\sigma_{j,k}}}$, that $\pi^{r}(\sigma_{j,k})$ is a measure zero subset of $\partial\mathbb{B}$.
+Since $V$ is contained in the union of finitely many ribs $\sigma_{j,k}$, $\pi^{r}(V)$ is a measure zero subset of $\partial\mathbb{B}$.
+Hence, $\pi^{r}(V)$ does not have an interior point in $\partial\mathbb{B}$.
+Next note that $\pi^{r}(V)=\Gamma$.
+While $\pi^{r}(V)\subset\Gamma$ comes just from definition of these sets, the reverse inclusion follows from the hypothesis that $D$ is bounded – indeed the fibers $(\pi^{r})^{-1}(z)$ are rays (from the origin), every one of which have got to intersect $\partial D$ as the origin lies within $D$.
+Thus $\pi^{r}(V)=\Gamma$; but then, this contradicts the conclusion of the previous para that $\pi^{r}(V)$ does not have an interior point in $\partial\mathbb{B}$.
+We now move to the case where $D$ is not necessarily a generic analytic polyhedron.
+By proposition 8.17, we may assume that $f_{j}$’s are homogeneous polynomials and form a minimal set of defining functions for the domain $D$.
+To prove this theorem, owing to theorem 1.9 (as argued above), it suffices to show that all ribs $(\sigma_{j,k})$ form a nowhere dense subset of $\partial D$.
+For simplicity of presentation, we write the details for the case $j=1,k=2$, other cases being similar.
+To prove this by contradiction, assume that there exists $p\in\sigma_{1,2}$ and an open subset $B$ of $\mathbb{C}^{N}$ containing $p$ such that $V:=B\cap\partial D\subset\sigma_{1,2}$.
+This means that for all $z\in V$, we have $|f_{1}(z)|=|f_{2}(z)|=1$.
+Note that the radial projection $\pi^{r}_{|_{\partial D}}:\partial D\to\partial\mathbb{B}$ is a homeomorphism.
+Indeed, $\pi^{r}_{|_{\partial D}}$ is a continuous bijection from $\partial D$ to $\partial\mathbb{B}$ with inverse that can be explicitly written down with ease: $g(z)=z/h(z)$, whose continuity is due to that of the Minkowski functional $h$ of our domain.
+As $V$ is an open subset of $\partial D$, it follows that $\pi^{r}(V)$ is an open subset of $\partial\mathbb{B}$.
+Though cones in general may have empty interior, this cannot happen for that spanned by $V$.
+To establish this, it suffices to prove that the the punctured cone generated by $V$ namely, $C_{V}^{\prime}:=\{tv:t\in\mathbb{R}^{+},~v\in V\}$ is an open subset of $\mathbb{C}^{N}$.
+To do this, first observe that $C_{V}^{\prime}$ is the same as the cone generated by $\pi^{r}(V)$, as follows by noting that the ray spanned by any $z\in V$ coincides with that spanned by $\pi^{r}(z)$.
+Now, note that the punctured cone spanned by $\pi^{r}(V)$ is the image of the open set $\mathbb{R}^{+}\times\pi^{r}(V)\subset\mathbb{R}^{+}\times\partial\mathbb{B}$ under the (backward) polar coordinates homeomorphism $\Psi:\mathbb{R}^{+}\times\partial\mathbb{B}\to\mathbb{C}^{N}\setminus\{0\}$ given by $\Psi(r,\omega)=r\omega$.
+It follows thereby that the cone $C_{V}^{\prime}$ is an open subset of $\mathbb{C}^{N}$, as was to be established.
+Let $z$ be an arbitrary point in $C^{\prime}_{V}$.
+Then there exists $z_{0}\in V$ and $t\in\mathbb{R}^{+}$ such that $z=tz_{0}$.
+Note that $t=h(z)$ and $z_{0}=z/h(z)$ because the Minkowski function $h$ is continuous.
+Let $d_{1},d_{2}$ be the degrees of homogeneous polynomials $f_{1},f_{2}$ respectively.
+As $|f_{1}(z_{0})|=|f_{2}(z_{0})|=1$, we have
+$|f_{2}(z)|=|f_{2}(tz_{0})|=t^{d_{2}}|f_{2}(z_{0})|=t^{d_{2}}~\text{and}~|f_{1}(z)|=|f_{1}(tz_{0})|=t^{d_{1}}|f_{1}(z_{0})|=t^{d_{1}}.$
+Thereby $|f_{2}(z)|^{d_{1}}=|f_{1}(z)|^{d_{2}}$.
+Since this equation holds for all $z\in V$, by the identity principle applied to the real analytic functions $|f_{1}|^{d_{2}},~|f_{2}|^{d_{1}}$ gives $|f_{2}(z)|^{d_{1}}=|f_{1}(z)|^{d_{2}}$ for all $z\in\mathbb{C}^{N}$.
+Thereby, the condition $|f_{2}(z)|<1$ holds whenever $|f_{1}(z)|<1$ holds i.e., we may as well drop one of these functions in defining the polyhedron.
+This contradicts the minimality of the defining functions for $D$, thereby completing the proof of this theorem.
+∎
+Proof of theorem 1.5 (by retracts)
+Following the proof of corollary 1.11 together with the above theorem and proposition 8.5, we obtain theorem 1.5.
+As we have enunciated theorems involving balanced analytic polyhedra with $\mathbb{C}$-extremal boundaries, we now deal with the issue of substantiating such theorems by confirming existence of ample examples; indeed, we now discuss techniques for constructing plenty of examples of bounded balanced $\mathbb{C}$-extremal analytic polyhedra.
+As noted in remark 8.18, for every such $D$, its Minkowski functional is given by $h(z)=\max\{|f_{j}(z)|^{1/d_{j}}:1\leq j\leq m\}$, where each $f_{j}$ is a homogeneous polynomial (whose degree we denote by $d_{j}$).
+Furthermore, according to proposition 2.9 (f), $D$ is bounded if and only if $h^{-1}(0)=\{0\}$.
+Therefore, to construct such examples, it suffices to show the existence of homogeneous polynomials $f_{j}$ satisfying the following conditions:
+  1. (A)
+    for each $j$, $f_{j}$ does not reduce to a non-zero constant on any line $l$ not passing through the origin. This will give the $\mathbb{C}$-extremal condition of the theorem 1.9.
+  2. (B)
+    $Z(f_{1},\ldots,f_{m})=\{0\}$, where $Z(f_{1},\ldots,f_{m})$ denotes the common zero set of $f_{1},\ldots,f_{m}$. This corresponds to the boundedness of the domain.
+First, we provide concrete examples and methods for constructing examples of homogeneous polynomials satisfying condition (A).
+Example-1: According to lemma 8.19, every homogeneous polynomial in $\mathbb{C}[z,w]$ that is not a power of a linear polynomial satisfies condition (A).
+Example-2: $f(z_{1},z_{2},z_{3})=(z_{1}^{2}+z_{2}^{2})(z_{2}^{2}+z_{3}^{2})$ satisfies condition (A), which we now prove by contradiction below.
+More generally, $(az_{1}^{2}+bz_{2}^{2})(cz_{2}^{2}+dz_{3}^{2})$ with $a,b,c,d\neq 0$ satisfies condition (A).
+Suppose that $f$ does not satisfy condition (A).
+Then there exists a line $\ell$ that does not pass through the origin and also fails to satisfy condition (A).
+We can write $f(p+tv)=g_{1}(t)g_{2}(t)$, where $g_{j}(t)=(p_{j}+tv_{j})^{2}+(p_{j+1}+tv_{j+1})^{2}$ for $j=1,2$.
+If $f$ is a non-zero constant on $l$ then for each $j$, $g_{j}$ is a non-zero constant polynomial because if $g_{j}$ is non-constant for some $j$ then their product is also a non-constant polynomial of degree >1>1.
+Now, we consider
+$g_{1}(p+tv)=(p_{1}^{2}+p_{2}^{2})+2(p_{1}v_{1}+p_{2}v_{2})t+(v_{1}^{2}+v_{2}^{2})t^{2}.$
+If $g_{1}$ is constant on $\ell$ then $v_{1}^{2}+v_{2}^{2}=0$ and $p_{1}v_{1}+p_{2}v_{2}=0$.
+Multiplying $p_{1}^{2}$ on both sides of the equation $v_{1}^{2}+v_{2}^{2}=0$, we obtain that $p_{1}^{2}v_{1}^{2}+p_{1}^{2}v_{2}^{2}=0$.
+Since $p_{1}v_{1}=-p_{2}v_{2}$, we have $p_{2}^{2}v_{2}^{2}+p_{1}^{2}v_{2}^{2}=(p_{1}^{2}+p_{2}^{2})v_{2}^{2}$.
+As $p_{1}^{2}+p_{2}^{2}$ is not a non-zero constant, we conclude that $v_{2}=0$.
+Furthermore, as $v_{1}^{2}=-v_{2}^{2}$, it follows that $v_{1}=0$.
+Next, we consider the function $g_{2}(t)$.
+Using similar arguments, we can prove that $v_{2}=v_{3}=0$.
+This contradicts the assumption that $\ell$ is a one-dimensional line that does not pass through the origin.
+This shows that $f$ satisfies condition (A).
+While the above example was a concrete one in three variables, we now formulate a general proposition to construct a fairly large collection of polynomials in any (finite) number of variables, satisfying condition (A).
+Proposition 8.23.
+Let $g$ be homogeneous polynomial in $\mathbb{C}[z_{1},\ldots,z_{N}]$.
+  1. (i)
+    If $f$ is a homogeneous polynomial in $\mathbb{C}[z_{1},\ldots,z_{N}]$ satisfies condition (A) then $h_{1}=f\cdot g$ is a homogeneous polynomial satisfying condition (A).
+  2. (ii)
+    Let $f_{1},\ldots,f_{k}$ be homogeneous polynomials in $\mathbb{C}[z_{1},\ldots,z_{N}]$ such that its common zero set $Z(f_{1},\ldots,f_{k})=\{0\}$.
+Then $h_{2}:=f_{1}\cdots f_{k}\cdot g$ is a homogeneous polynomial satisfying condition (A).
+Proof.
+To prove (i) by contradiction, assume that $h_{1}$ does not satisfy condition (A).
+Then there exists a line $\ell:=\{p+tv:t\in\mathbb{R}\}$ such that $h_{1}$ is a non-zero constant on $\ell$.
+In particular, if $h_{1}(p+tv)$ is a non-zero constant function, then $f(p+tv)$ must also be a non-zero constant function.
+This contradicts our assumption that $f$ satisfies condition (A).
+For proving (ii),consider the homogeneous polynomial $h_{2}(z)=f_{1}(z)\cdots f_{k}(z)\cdot g(z)$.
+If $h_{2}$ does not satisfy condition (A), then there exists a line $\ell:=\{p+tv:t\in\mathbb{R}\}$, where $v\neq 0$, such that $h_{2}$ is a non-zero constant on $\ell$.
+Then for all $t\in\mathbb{R}$, we can write
+$h_{2}(p+tv)=f_{1}(p+tv)\cdots f_{k}(p+tv)\cdot g(p+tv).$
+Since $h_{2}$ is a non-zero constant on $\ell$, $f_{j}(p+tv)$ and $g(p+tv)$ reduces to a non-zero constant function.
+Note that the coefficient of the highest power of $t$ in each $f_{j}(p+tv)$ is $f_{j}(v)$ due to the homogeneity of the polynomials.
+Therefore, we must have $f_{j}(v)=0$ for each $j=1,2,\ldots,m$.
+This implies that $v\in Z(f_{1},\ldots,f_{m})$, contradicting our assumption that $Z(f_{1},\ldots,f_{m})=\{0\}$.
+∎
+We shall actually be constructing examples which shall satisfy the stronger condition of non-degeneracy of the defining functions at every point of all the pairwise intersections of faces, briefly referred to as condition-(C).
+Lemma 8.24.
+Let $g_{1},g_{2},f$ be homogeneous polynomials in $\mathbb{C}[z_{1},\ldots,z_{N}]$ with degrees $d_{1},d_{2},d$ respectively.
+If $v$ is a common zero of $g_{1}(z),g_{2}(z),f(z)-e^{i\theta}$ for some $\theta\in(-\pi,\pi]$ then $e^{-i\theta/d}v$ is a common zero of $g_{1}(z),g_{2}(z),f(z)-1$.
+Proof.
+Assume that $v$ is a common zero of $g_{1}(z),g_{2}(z),f(z)-e^{i\theta}$.
+This implies that $g_{1}(v)=g_{2}(v)=0$ and $f(v)=e^{i\theta}$ for some $\theta\in(-\pi,\pi]$.
+Note that $f(e^{-i\theta/d}v)=e^{-i\theta}f(v)=1$ because by the homogeneity of $f$.
+Since the zero set of a homogeneous polynomial is closed under scalar multiplication, $e^{-i\theta/d}v$ is a zero of $g_{1}$ and $g_{2}$.
+Hence, we obtain that $e^{-i\theta/d}v$ is a common zero of $g_{1}(z),g_{2}(z),f(z)-1$.
+∎
+We shall now show that the functions $f_{j}$ in the foregoing examples satisfy the above mentioned condition-(C).
+Towards this, we prove that for each $j\neq k$, $\nabla f_{j}(z)$ and $\nabla f_{k}(z)$ are $\mathbb{C}$-linearly independent for each $z\in\sigma_{j,k}$.
+Consider the homogeneous polynomials
+$f_{1}(z_{1},z_{2},z_{3})=z_{1}z_{2}z_{3},~~~~~~~~f_{2}(z_{1},z_{2},z_{3})=(z_{1}^{2}+z_{2}^{2})(z_{2}^{2}+z_{3}^{2})(z_{3}^{2}+z_{1}^{2})$
+$f_{3}(z_{1},z_{2},z_{3})=(z_{1}+z_{2}-z_{3})(z_{1}-z_{2}+z_{3})(-z_{1}+z_{2}+z_{3})$
+Their gradients are given as follows:
+$\nabla f_{1}(z)=(z_{2}z_{3},z_{1}z_{3},z_{1}z_{2})$
+$\nabla f_{2}(z)=\left(2z_{1}(z_{2}^{2}+z_{3}^{2})(2z_{1}^{2}+z_{2}^{2}+z_{3}^{2}),2z_{2}(z_{1}^{2}+z_{3}^{2})(z_{1}^{2}+2z_{2}^{2}+z_{3}^{2}),\right.\\ \left.2z_{3}(z_{1}^{2}+z_{2}^{2})(z_{1}^{2}+z_{2}^{2}+2z_{3}^{2})\right)$
+$\nabla f_{3}(z)=\left(-3z_{1}^{2}+z_{2}^{2}+z_{3}^{2}+2z_{1}z_{3}-2z_{2}z_{3}+2z_{1}z_{2},z_{1}^{2}-3z_{2}^{2}+z_{3}^{2}+\right.\\ \left.2z_{2}z_{3}-2z_{1}z_{3}+2z_{1}z_{2},z_{1}^{2}+z_{2}^{2}-3z_{3}^{2}-2z_{1}z_{2}+2z_{1}z_{3}+2z_{2}z_{3}\right)$
+First, we prove that $\nabla f_{1}(z)$ and $\nabla f_{2}(z)$ are linearly independent for each $z\in\sigma_{1,2}$.
+It suffices to show that for every $\theta_{1},\theta_{2}\in(-\pi,\pi]$, $\nabla f_{1}(z)$ and $\nabla f_{2}(z)$ are linearly independent for each $z\in\sigma_{1,2}$ with $f_{1}(z)=e^{i\theta_{1}}$ and $f_{2}(z)=e^{i\theta_{2}}$.
+Initially, we shall show this for $\theta_{1}=\theta_{2}=0$.
+Let $z$ be a point in $\sigma_{1,2}$ such that $f_{1}(z)=1$ and $f_{2}(z)=1$.
+Suppose $\nabla f_{1}(z)$ is a scalar multiple of $\nabla f_{2}(z)$.
+Then
+$\frac{\partial f_{2}(z)/\partial z_{1}}{\partial f_{1}(z)/\partial z_{1}}=\frac{\partial f_{2}(z)/\partial z_{2}}{\partial f_{1}(z)/\partial z_{2}}=\frac{\partial f_{2}(z)/\partial z_{3}}{\partial f_{1}(z)/\partial z_{3}}$
+The above equation reduces to the following equations:
+$g_{1}(z_{1},z_{2},z_{3})=0,~~g_{2}(z_{1},z_{2},z_{3})=0$
+where
+$g_{1}(z_{1},z_{2},z_{3})$ $=$ $\frac{\partial f_{2}}{\partial z_{1}}\frac{\partial f_{1}}{\partial z_{2}}-\frac{\partial f_{2}}{\partial z_{2}}\frac{\partial f_{1}}{\partial z_{1}}$
+$=$ $z_{3}(z_{1}^{4}z_{2}^{2}+2z_{1}^{4}z_{3}^{2}-z_{1}^{2}z_{2}^{4}+z_{1}^{2}z_{2}^{4}-2z_{2}^{4}z_{3}^{2}-z_{2}^{2}z_{3}^{4}),$
+$g_{2}(z_{1},z_{2},z_{3})$ $=$ $\frac{\partial f_{2}}{\partial z_{2}}\frac{\partial f_{1}}{\partial z_{3}}-\frac{\partial f_{2}}{\partial z_{3}}\frac{\partial f_{1}}{\partial z_{2}}$
+$=$ $z_{1}(-z_{1}^{4}z_{2}^{2}+z_{1}^{4}z_{3}^{2}-2z_{1}^{2}z_{2}^{4}+2z_{1}^{2}z_{3}^{4}-z_{2}^{4}z_{3}^{2}+z_{2}^{2}z_{3}^{4}).$
+Now, we consider the common zeros of the equations $g_{1}=0,g_{2}=0,f_{2}=1$.
+Note that the common zero set of these equations is exactly the same as the variety of the ideal $I=\langle g_{1},g_{2},f_{2}-1\rangle$.
+Note that $\sqrt{I}=\mathbb{C}[z_{1},z_{2},z_{3}]$ [^13] We can compute radical of $I$ using the Macaulay2 software..
+Therefore, the functions $g_{1},g_{2},f_{2}-1$ have no common zero in $\mathbb{C}^{3}$.
+Fix $\theta\in(-\pi,\pi]$.
+We want to show that there is no common zero for the functions $g_{1},g_{2},f_{2}-e^{i\theta}$.
+Let $v$ be a common zero of these functions.
+By the above lemma 8.24, $e^{-i\theta/6}v$ is a common zero of $g_{1},g_{2},f_{2}-1$.
+This contradicts the above fact that there is no common zero for $g_{1},g_{2},f_{2}-1$.
+This shows that $\nabla f_{1}(z)$ and $\nabla f_{2}(z)$ are linearly independent for each $z\in\sigma_{1,2}$.
+Next, we prove that $\nabla f_{2}(z)$ and $\nabla f_{3}(z)$ are linearly independent for each $z\in\sigma_{2,3}$.
+Let $z$ be a point in $\sigma_{1,2}$ such that $f_{2}(z)=1$ and $f_{3}(z)=1$.
+If $\nabla f_{2}(z)$ is a scalar multiple of $\nabla f_{3}(z)$, then the ratios of each component of $\nabla f_{2}$ and $\nabla f_{3}$ are the same.
+Hence, we obtain the following equations.
+$h_{1}(z_{1},z_{2},z_{3})=0,~~h_{2}(z_{1},z_{2},z_{3})=0$
+where
+$h_{1}(z_{1},z_{2},z_{3})=\frac{\partial f_{2}}{\partial z_{1}}\frac{\partial f_{3}}{\partial z_{2}}-\frac{\partial f_{2}}{\partial z_{2}}\frac{\partial f_{3}}{\partial z_{1}}\\ =3z_{1}^{6}z_{2}-2z_{1}^{5}z_{2}z_{3}+2z_{1}^{5}z_{3}^{2}+9z_{1}^{4}z_{2}^{3}-2z_{1}^{4}z_{2}^{2}z_{3}+9z_{1}^{4}z_{2}z_{3}^{2}-4z_{1}^{4}z_{3}^{3}-9z_{1}^{3}z_{2}^{4}-6z_{1}^{3}z_{2}^{2}z_{3}^{2}+3z_{1}^{3}z_{3}^{4}+\\ 2z_{1}^{2}z_{2}^{4}z_{3}+6z_{1}^{2}z_{2}^{3}z_{3}^{2}+3z_{1}^{2}z_{2}z_{3}^{4}-2z_{1}^{2}z_{3}^{5}-3z_{1}z_{2}^{6}+2z_{1}z_{2}^{5}z_{3}-9z_{1}z_{2}^{4}z_{3}^{2}-3z_{1}z_{2}^{2}z_{3}^{4}+z_{1}z_{3}^{6}\\ -2z_{2}^{5}z_{3}^{2}+4z_{2}^{4}z_{3}^{3}-3z_{2}^{3}z_{3}^{4}+2z_{2}^{2}z_{3}^{5}-z_{2}z_{3}^{6}$
+$h_{2}(z_{1},z_{2},z_{3})=\frac{\partial f_{2}}{\partial z_{2}}\frac{\partial f_{3}}{\partial z_{3}}-\frac{\partial f_{2}}{\partial z_{3}}\frac{\partial f_{3}}{\partial z_{2}}\\ =z_{1}^{6}z_{2}-z_{1}^{6}z_{3}-2z_{1}^{5}z_{2}^{2}+2z_{1}^{5}z_{3}^{2}+3z_{1}^{4}z_{2}^{3}+3z_{1}^{4}z_{2}^{2}z_{3}-3z_{1}^{4}z_{2}z_{3}^{2}-3z_{1}^{4}z_{3}^{3}-4z_{1}^{3}z_{2}^{4}+4z_{1}^{3}z_{3}^{4}\\ +2z_{1}^{2}z_{2}^{5}+9z_{1}^{2}z_{2}^{4}z_{3}-6z_{1}^{2}z_{2}^{3}z_{3}^{2}+6z_{1}^{2}z_{2}^{2}z_{3}^{3}-9z_{1}^{2}z_{2}z_{3}^{4}-2z_{1}^{2}z_{3}^{5}-2z_{1}z_{2}^{5}z_{3}-2z_{1}z_{2}^{4}z_{3}^{2}\\ +2z_{1}z_{2}^{2}z_{3}^{4}+2z_{1}z_{2}z_{3}^{5}+3z_{2}^{6}z_{3}+9z_{2}^{4}z_{3}^{3}-9z_{2}^{3}z_{3}^{4}-3z_{2}z_{3}^{6}$
+We want to find the common zeros of $h_{1}=0,h_{2}=0,f_{2}=1$.
+Consider the ideal $J=\langle h_{1},h_{2},f_{2}-1\rangle$.
+Note that $\sqrt{J}=\mathbb{C}[z_{1},z_{2},z_{3}]$.
+So, we conclude that there is no common zero for the functions $h_{1},h_{2},f_{2}-1$.
+Using similar arguments to those employed in the proof of the linear independence of $\nabla f_{1}(z),\nabla f_{2}(z)$ that there is no common zero for the functions $h_{1},h_{2},f_{2}=e^{i\theta}$, where $\theta$ is some fixed real number in $(-\pi,\pi]$.
+This shows that $\nabla f_{2}(z)$ and $\nabla f_{3}(z)$ are linearly independent for each $z\in\sigma_{2,3}$.
+Finally, we prove that $\nabla f_{1}(z)$ and $\nabla f_{3}(z)$ are linearly independent for each $z\in\sigma_{1,3}$.
+Let $z$ be a point in $\sigma_{1,2}$ such that $f_{1}(z)=1$ and $f_{3}(z)=1$.
+If $\nabla f_{1}(z)$ and $\nabla f_{3}(z)$ are linearly dependent, then the ratios of each component of $\nabla f_{1}$ and $\nabla f_{3}$ are the same.
+Hence, we obtain the following equations.
+$k_{1}(z_{1},z_{2},z_{3})=0,~~k_{2}(z_{1},z_{2},z_{3})=0$
+where
+$k_{1}(z_{1},z_{2},z_{3})$ $=$ $\frac{\partial f_{1}}{\partial z_{1}}\frac{\partial f_{3}}{\partial z_{2}}-\frac{\partial f_{1}}{\partial z_{2}}\frac{\partial f_{3}}{\partial z_{1}}$
+$=$ $z_{3}(-3z_{1}^{3}+z_{1}^{2}z_{2}+2z_{1}^{2}z_{3}-z_{1}z_{2}^{2}+z_{1}z_{3}^{2}+3z_{2}^{3}-2z_{2}^{2}z_{3}-z_{2}z_{3}^{2})$
+$k_{2}(z_{1},z_{2},z_{3})$ $=$ $\frac{\partial f_{1}}{\partial z_{2}}\frac{\partial f_{3}}{\partial z_{3}}-\frac{\partial f_{1}}{\partial z_{3}}\frac{\partial f_{3}}{\partial z_{2}}$
+$=$ $z_{1}(z_{1}^{2}z_{2}-z_{1}^{2}z_{3}+2z_{1}z_{2}^{2}-2z_{1}z_{3}^{2}-3z_{2}^{3}+z_{2}^{2}z_{3}-z_{2}z_{3}^{2}+3z_{3}^{3})$
+Now, we look at the common zeros of $k_{1}=0,k_{2}=0,f_{1}-1$.
+Consider the ideal $K=\langle k_{1},k_{2},f_{1}-1\rangle$.
+Note that $\sqrt{K}=\mathbb{C}[z_{1},z_{2},z_{3}]$.
+So, we conclude that there is no common zero for the functions $k_{1},k_{2},f_{1}-1$.
+Using similar arguments to those employed in the proof of the linear independence of $\nabla f_{1}(z),\nabla f_{2}(z)$ that there is no common zero for the functions $h_{1},h_{2},f_{2}=e^{i\theta}$, where $\theta$ is some fixed real number in $(-\pi,\pi]$.
+This shows that $\nabla f_{1}(z)$ and $\nabla f_{3}(z)$ are linearly independent for each $z\in\sigma_{1,3}$.
+We would like to conclude this section by laying down at least one concrete example of a balanced analytic polyhedron whose boundary is not only $\mathbb{C}$-extremal but also has the feature of non-degeneracy of the defining functions throughout (unlike the example $D_{h}$ in $\mathbb{C}^{2}$ that we dealt with earlier) the various open faces and all pairwise intersections of the closed faces.
+Namely, the analytic polyhedron
+$D=\{(z_{1},z_{2},z_{3})\in\mathbb{C}^{3}:|f_{i}(z_{1},z_{2},z_{3})|<1~\text{for}~i=1,2,3\}$
+where $f_{1}(z)=z_{1}z_{2}z_{3},f_{2}(z)=(z_{1}^{2}+z_{2}^{2})(z_{2}^{2}+z_{3}^{2})(z_{3}^{2}+z_{1}^{2}),f_{3}(z)=(z_{1}-z_{2}+z_{3})(z_{1}+z_{2}-z_{3})(-z_{1}+z_{2}-z_{3})$.
 9. Retracts of Unbounded and topologically non-trivial domains – Proofs of theorems 1.201.24
-As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allows topological non-triviality, as we considered ‘analytic complements’ $D\setminus A$ at the beginning therein, the boundedness assumption prevailed.
+As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allow topological non-triviality, as we considered ‘analytic complements’ $D\setminus A$ at the beginning therein, the boundedness assumption prevailed.
 To go beyond, to more general settings, we require the following lemma:
@@
 Proof of theorem 1.21.
-Let $Z$ be a retract of ×$\mathbb{C}\times D$ passing through $p:=(a,c)$.
+Let $Z$ be a retract of $\mathbb{C}\times D$ passing through $p:=(a,c)$.
 Let $F:\mathbb{C}\times D\to\mathbb{C}\times D$ be a holomorphic retraction map such that ${\rm image}(F)=Z$.
@@
 By comparing the first component of $F\circ F=F$, we get $f_{1}(f_{1}(z,w),c)=f_{1}(z,w)$.
-Note that $F$ is a holomorphic retraction map on ×$\mathbb{C}\times D$ onto $Z_{1}\times\{c\}$, where $Z_{1}:={\rm image}(f_{1})$ on ×$\mathbb{C}\times D$.
+Note that $F$ is a holomorphic retraction map on $\mathbb{C}\times D$ onto $Z_{1}\times\{c\}$, where $Z_{1}:={\rm image}(f_{1})$ on $\mathbb{C}\times D$.
 By lemma 9.1, $Z_{1}$ is a retract of $\mathbb{C}$ passing through $a$.
@@
 ∎
-Next we shall detail the complete determination of the retraction maps of ×$\mathbb{C}\times\Delta$.
+Next we shall detail the complete determination of the retraction maps of $\mathbb{C}\times\Delta$.
 Proof of theorem 1.23:.
@@
 $f(z,w)=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots,$
-where $P_{i}(z,w)$ is a homogenous polynomial in $z,w$ of degree $i$.
+where $P_{i}(z,w)$ is a homogeneous polynomial in $z,w$ of degree $i$.
 Since $f(f(z,w),0)=f(z,w)$, we get
@@
 Continuing in this way, we get $P_{i}(z,w)=w(Q_{i-1}(z,w))$, where $Q_{i-1}$ is a polynomial in $z$ and $w$.
-Finally $f(z,w)=z+w(h(z,w))$, where $h$ is some holomorphic map on ×$\mathbb{C}\times\Delta$.
+Finally $f(z,w)=z+w(h(z,w))$, where $h$ is some holomorphic map on $\mathbb{C}\times\Delta$.
 In this subcase, $F(z,w)=(z+w(h(z,w)),0)$.
@@
 $f(z,w)=P_{1}(z,w)+P_{2}(z,w)+P_{3}(z,w)+\ldots,$
-where $P_{i}(z,w)$ is a homogenous polynomial in $z,w$ of degree $i$
+where $P_{i}(z,w)$ is a homogeneous polynomial in $z,w$ of degree $i$
 Since $f(f(z,w),w)=f(z,w)$, we get
@@
 In this subcase, $F(z,w)=\left(h(w),w\right)$. Hence we get $Z=\{(h(w),w):w\in\Delta\}$.
-Since ×$\mathbb{C}\times\Delta$ is homogenous, upto a conjugation of automorphism, every non-trivial retraction map is of the form $F(z,w)=\left(z+wh(z,w),0\right)$, where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$ or $F(z,w)=(h(w),w)$, where $h\in\mathcal{O}(\Delta)$
+Since $\mathbb{C}\times\Delta$ is homogeneous, upto a conjugation of automorphism, every non-trivial retraction map is of the form $F(z,w)=\left(z+wh(z,w),0\right)$, where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$ or $F(z,w)=(h(w),w)$, where $h\in\mathcal{O}(\Delta)$
 Next, we turn to the proofs of theorems dealing with domains or manifolds with much greater topological complexity.
@@
 For each fixed $w\in D$, $g_{w}:(\mathbb{C}\setminus A)\to D$ defined by $g_{w}(z)=g(z,w)$ is a bounded holomorphic map on $(\mathbb{C}\setminus A)$.
-Applying Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each component of $g_{w}$, $g_{w}$ can extend holomorphically as a map $g_{w}^{\prime}:\mathbb{C}\to\overline{D}$.
+Applying Riemann removable singularities theorem to each component of $g_{w}$, $g_{w}$ can extend holomorphically as a map $g_{w}^{\prime}:\mathbb{C}\to\overline{D}$.
 By Liouville’s theorem, $g_{w}^{\prime}$ is a constant function.
@@
 This shows that $f_{w}$ is a retraction map on $(\mathbb{C}\setminus A)$.
-Since every domain in $\mathbb{C}$, the only retraction maps are identity and constant function, either $f_{w}\equiv I$, the identity map on $(\mathbb{C}\setminus A)$ or $f_{w}\equiv$ constant function (depends on $w$), hence imples that $f_{w}$ is independent of $z$.
+Since every domain in $\mathbb{C}$, the only retraction maps are identity and constant function, either $f_{w}\equiv I$, the identity map on $(\mathbb{C}\setminus A)$ or $f_{w}\equiv$ constant function (depends on $w$), hence implies that $f_{w}$ is independent of $z$.
 Hence for all $z,w\in(\mathbb{C}\setminus A)\times D$, we have
@@
 (9.5) $f(z,\tilde{g}(w))=z.$
-Define $Z:={\rm image}(\tilde{g})$. According to (9.4) and (9.5), we divide into $2$ cases.
+[…]
 Case-1: $f(z,\tilde{g}(w))=H(\tilde{g}(w))$ for all $(z,w)\in(\mathbb{C}\setminus A)\times D$
 We claim that $\widetilde{Z}=\{(H(w),w):w\in Z\}$.
-To prove this claim, observe that for each $w\in Z$, we have $\tilde{g}(w)=w$ and $f(H(w),w)=f(H(w),\tilde{g}(w))=H(w)$ This implies that $\{(H(w),w):w\in Z\}\subset\widetilde{Z}$.
+To prove this claim, observe that for each $w\in Z$, we have $\tilde{g}(w)=w$ and $f(H(w),w)=f(H(w),\tilde{g}(w))=H(w)$ .
+This implies that $\{(H(w),w):w\in Z\}\subset\widetilde{Z}$.
 For attaining converse implication, let $(z_{0},w_{0})$ be an arbitrarily point in $\widetilde{Z}$.
-Since $\tilde{g}(w_{0})=w_{0}$ and $f(z_{0},w_{0})=z_{0}$ Hence we get $z_{0}=f(z_{0},w_{0})=f(z_{0},\tilde{g}(w_{0}))=H(\tilde{g}(w_{0}))=H(w_{0})$.
+Since $\tilde{g}(w_{0})=w_{0}$ and $f(z_{0},w_{0})=z_{0}$ .
+Hence we get $z_{0}=f(z_{0},w_{0})=f(z_{0},\tilde{g}(w_{0}))=H(\tilde{g}(w_{0}))=H(w_{0})$.
 This implies that $(z_{0},w_{0})=\left(H(w_{0}),w_{0}\right)$.
@@
 We shall next record the proof of theorem 1.24, whose verifying details will be shortened as most of its arguments are contained in the prequel.
-To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of ×$\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$.
+To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of $\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$.
 Indeed as mentioned earlier, this class contains domains with a wide variety of fundamental (or for that matter higher homotopy and homology) groups.
@@
 We shall now detail the proof of theorem 1.20 about retracts of the Hartogs triangle and its analytic complements.
-Making the identification $\Omega_{H}\simeq\Delta\times\Delta^{*}$, retracts of the Hartogs triangle can be deduced as a special corollary of lemma 8.1 taking $D$ to be just the bidisc $\Delta^{2}$.
-However, for analytic complements $\Omega_{H}\setminus A$ this cannot be done merely by citing lemma 8.1 (at the very least, directly!), as the Hartogs triangle $\Omega_{H}$ does not admit a Stein neighbourhood basis.
-We therefore prefer to give a direct proof as follows.
+Firstly, a very quick proof is as follows.
+Making the identification $\Omega_{H}\simeq\Delta\times\Delta^{*}$, retracts of the Hartogs triangle can be deduced as an immediate corollary of lemma 8.1 taking $D$ to be just the bidisc $\Delta^{2}$.
+An alternative direct albeit longer proof is as follows.
 Proof of theorem 1.20.
@@
 Applying the same argument, $\widetilde{F}_{1},\widetilde{F}_{2}$ can be extended to a holomorphic map from $\Delta^{2}$ to $\overline{\Delta}$.
-By abuse of notation, $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorphic maps from $\Delta^{2}$ to $\Delta$.
+By a little abuse of notation, we view $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorphic maps from $\Delta^{2}$ to $\Delta$.
 We now observe that the images of these maps actually does not contain any point from $\partial\Delta$ i.e., they map into $\Delta$; this follows by the open mapping theorem (non-constant holomorphic functions from domains in $\mathbb{C}^{N}$ are open maps into $\mathbb{C}$;).
@@
 This proves that $\widetilde{F}$ is a holomorphic retraction on $\Delta^{2}$.
-Let $\widetilde{Z}:={\rm image}(\widetilde{F})$.
-By theorem 3 in [Heath_Suff], $\widetilde{Z}$ is either of the form $\{(z,f(z)):z\in\Delta\}$ for some holomorphic self map $f$ on $\Delta$ or $\{(g(w),w):w\in\Delta\}$ for some holomorphic self map $g$ on $\Delta$.
-Since $Z\subset(\Delta\times\Delta^{*})\setminus A$ and $Z\subset\widetilde{Z}$ ($\widetilde{F}$ is a holomorphic extension of $F$), we get $Z$ is either of the form $\{(z,f(z)):z\in\Delta\}\cap((\Delta\times\Delta^{*})\setminus A)$ for some holomorphic map $f:\Delta\to\Delta$ or $\{(g(w),w):w\in\Delta^{*}\}\cap((\Delta\times\Delta^{*})\setminus A)$ for some holomorphic map $g:\Delta\to\Delta$.
-∎
+[…]
 10. Retracts of $\mathbb{C}^{2}$; proof of theorem 1.25
@@
 $Z:=\{(e^{z},ze^{-z}):z\in\mathbb{C}\}$
-is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},\;zwe^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
+is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},zw~e^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
 Proof.
@@
 As $\mathbb{C}\setminus L_{1}$ is simply connected, there exists atleast one (indeed infinitely many) branch of the (complex) logarithm on $\mathbb{C}\setminus L_{1}$; pick one of them and call it $\ell_{1}$.
-Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ Note that $L_{2}$ divides $D$ into two regions which shall be referred to by $D^{+}~{}\text{and}~{}D^{-}$.
+Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ .
+Note that $L_{2}$ divides $D$ into two regions which shall be referred to by $D^{+}~\text{and}~D^{-}$.
 As with $\ell_{1}$, we likewise pick and denote by $\ell_{2}$, a branch of logarithm on $\mathbb{C}\setminus L_{2}$.
@@
 Getting examples of retracts which are not conjugate to graphs is more non-trivial but it is not difficult to derive it from the work of Forstneric – Globevnik – Rosay [Forstn].
-To briefly recall their main result, let us recall as they do in the introduction of [Forstn] that Abhyankar and Moh in [Abynkr] and M.Suzuki in [Suzuki] proved that if $P$ is a polynomial embedding of $\mathbb{C}$ into $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ of $\mathbb{C}^{2}$ such that $\psi\circ P(\mathbb{C})=\mathbb{C}\times\{0\}$.
-This result is not true for holomorphic embeddings of $\mathbb{C}$.
-By proposition $1$ in [Forstn], there exists a proper holomorphic embedding $H:\mathbb{C}\to\mathbb{C}^{N}$ such that for no automorphism $\psi$ of $\mathbb{C}^{N}$, $\psi\circ H(\mathbb{C})=\mathbb{C}\times\{0\}\subset\mathbb{C}^{N}$.
-We want to prove that ${\rm image}(H)$ is a retract of $\mathbb{C}^{N}$.
-Towards this the manner in which this proper holomorphic embedding $H$ is defined and attained in the proof of proposition $2$ of their paper [Forstn] is now very briefly recalled.
-Start with the embedding $H_{0}:\mathbb{C}\to\mathbb{C}^{N}$ defined by $H_{0}(\zeta)=(\zeta,0)$.
-For $j\geq 1$, $H_{j}:=\psi_{j}\circ H_{j-1}$, where $\psi_{j}$ is some automorphism of $\mathbb{C}^{N}$.
-These $(H_{j})$’s consitute a sequence of proper embeddings of $\mathbb{C}$ satisfying some properties.
-Taking their uniform limit on compacts $H:=limH_{j}$ yields the desired embedding.
-Consider now, the holomorphic retraction $\rho_{0}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{0}(z,w)=(z,0)$, where $z\in\mathbb{C}$ and $w\in\mathbb{C}^{N-1}$.
+To briefly recall their main result, let us recall as they do in the introduction of [Forstn], that Abhyankar and Moh in [Abynkr] and M.Suzuki in [Suzuki] proved the following: if $P$ is a polynomial embedding of $\mathbb{C}$ into $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ of $\mathbb{C}^{2}$ such that $\psi\circ P(\mathbb{C})=\mathbb{C}\times\{0\}$.
+This result is not true for holomorphic embeddings of $\mathbb{C}$ in $\mathbb{C}^{2}$ or for that matter in $\mathbb{C}^{N}$ for any $N\geq 2$.
+By the first proposition in [Forstn], there exists a proper holomorphic embedding $H:\mathbb{C}\to\mathbb{C}^{N}$ such that for no automorphism $\psi$ of $\mathbb{C}^{N}$, we can attain: $\psi\circ H(\mathbb{C})=\mathbb{C}\times\{0\}\subset\mathbb{C}^{N})$.
+For the embedding provided by that proposition, we prove here that ${\rm image}(H)$ is indeed a retract of $\mathbb{C}^{N}$.
+Towards this, the manner in which this proper holomorphic embedding $H$ is defined and attained in the proof of proposition $2$ of their paper [Forstn], is now very briefly recalled.
+Start with the standard embedding $H_{0}:\mathbb{C}\to\mathbb{C}^{N}$ defined by $H_{0}(\zeta)=(\zeta,0)$.
+For $j\geq 1$, let $H_{j}:=\psi_{j}\circ H_{j-1}$, where $\psi_{j}$ is some automorphism of $\mathbb{C}^{N}$.
+These $(H_{j})$’s constitute a sequence of proper embeddings of $\mathbb{C}$ satisfying some properties which we do not need to repeat here.
+Taking their uniform limit on compacts $H:=\lim H_{j}$ yields the desired embedding; for further details, we refer the reader to [Forstn] but this recap will suffice for our aforementioned purpose which we finish off in the next paragraph.
+Denote by $\rho_{0}:\mathbb{C}^{N}\to\mathbb{C}^{N}$, the simplest of standard one-dimensional retraction maps namely, $\rho_{0}(z,w)=(z,0)$, where $z\in\mathbb{C}$ and $w\in\mathbb{C}^{N-1}$.
 Note that ${\rm image}(H_{0})={\rm image}(\rho_{0})$.
-For $j\geq 1$, let $\rho_{j}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{j}:=\psi_{j}\circ H_{j}\circ\psi_{j}^{-1}$, where $\psi_{j}$ is the automorphism mentioned in the above construction.
+For $j\geq 1$, let $\rho_{j}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ be defined by $\rho_{j}:=\psi_{j}\circ H_{j}\circ\psi_{j}^{-1}$, where $\psi_{j}$ is the automorphism mentioned in the above construction.
 Observe that each $\rho_{j}$ is a holomorphic retraction with ${\rm image}(\rho_{j})={\rm image}(H_{j})$.
 Note that the $\rho_{j}$’s converge uniformly on compacts.
-If we denote the limit map by $\rho$ then, $\rho$ is also a retraction – this follows by passing to the limit in the defining equation of a retraction wherein taking a double limit is justified owing to the uniform convergence.
+If we denote the limit map by $\rho$ then, $\rho$ is also a retraction – this follows by passing to the limit in the defining equation of a retraction for our specific map $\rho_{j}$ (i.e., writing $\rho_{j}\circ\rho_{j}=\rho_{j}$ elaborately), wherein taking a ‘triple limit’ is justified owing to the uniform convergence.
 Finally, we note that the limit of this retraction map $\rho$ satisfies ${\rm image}(H)={\rm image}(\rho)$.
 Hence we have verfied the existence of a one-dimensional holomorphic retract $Z$ of $\mathbb{C}^{N}$ which is not straightenable, thereby not conjugate to a graph as well.
-We conclude by expositing on results about polynomial retracts of $\mathbb{C}^{2}$, the simplest of domains admitting non-trivial retracts.
-As mentioned in the foregoing, holomorphic retracts in $\mathbb{C}^{2}$ are not in general rectifiable.
-Also, it seems difficult to get a complete characterization of holomorphic retracts of $\mathbb{C}^{2}$.
-However it is possible to deduce a complete characterization of polynomial retracts on $\mathbb{C}^{2}$ from known results about the polynomial ring in two complex variables and this yileds in-particular, that all polynomial retracts are rectifiable.
-Even this deduction is not entirely trivial and needs some work which was done by Shpilrain and Yu in [Shpilrain--Yu], which we exposit upon.
-In order to lay this work out properly, we begin with some definitions.
+Owing to the foregoing impossibility of attaining a rectifiability result even for one-dimensional holomorphic retracts of $\mathbb{C}^{N}$, it seems too difficult to obtain a complete characterization of holomorphic retracts, indeed even for $\mathbb{C}^{2}$ which is unknown.
+However it is possible to deduce a complete characterization of polynomial retracts on $\mathbb{C}^{2}$ from known results about the polynomial ring in two complex variables and this yields in-particular, that all polynomial retracts are rectifiable – this result is not at all new.
+However, even this deduction is not entirely trivial and needs some work which was done by Shpilrain and Yu in [Shpilrain--Yu] with the key ingredients being the theorem 3.5 in [Cost] and the Abhyankar – Moh theorem (main theorem of [Abynkr]).
+We shall show an alternative route without these ingredients as discussed below.
+An important point to note here is that though the Abhyankar – Moh theorem is a result that renders the rectifiability of polynomial embeddings of $\mathbb{C}$, it does not follow as an immediate corollary that polynomial retracts can be straightened.
+Firstly, though retracts are embedded complex submanifolds wherein ‘embedded’ here is in the sense of differential geometry, it is far from clear that a (non-trivial) holomorphic retract of $\mathbb{C}^{2}$ is the image of an embedding of $\mathbb{C}$ – this is true but seems to require the uniformization theorem.
+Furthermore, if the retract is a polynomial retract, it is not at all immediate that the embedding given by the uniformization theorem can be taken to be a polynomial one.
+In-fact, the converse namely, that the image of an embedding in the sense of the Abhyankar – Moh theorem is indeed a (polynomial) retract is not clear without invoking that theorem.
+Notwithstanding all this, it is possible to give a direct, self-contained and elementary proof for this (known) result concerning the rectifiability of polynomial retracts of $\mathbb{C}^{2}$; in-particular, neither using the result of Costa and nor the Abhyankar – Moh theorem, as we do in the next subsection.
+10.1. Proof of theorem 1.25
+Since $Z$ is a polynomial retract of $\mathbb{C}^{2}$, there exists an idempotent map $F:\mathbb{C}^{2}\to\mathbb{C}^{2}$ whose (both) components are (holomorphic) polynomials and ${\rm image}(F)=Z$; we may assume from the outset via a translation, that $Z$ passes through the origin, so that $F(0)=0$ in-particular.
+Again, by a linear change of variables, we may further assume $DF_{|_{0}}(z,w)=(z,0)$.
+So, we can write $F(z,w)=(z,0)+(P_{1}(z,w),P_{2}(z,w))$ for some $P_{1},P_{2}\in\mathbb{C}[z,w]$ each of which is either zero or have the property that all its monomials are of degree at-least two.
+We shall now proceed to proving that $P_{1}\equiv 0$ or $P_{2}\equiv 0$.
+To prove this by contradiction, suppose that neither $P_{1}$ nor $P_{2}$ are identically zero.
+Let $P_{1}^{(r)},P_{2}^{(k)}$ be the homogeneous constituents of highest degree in the homogeneous decomposition of $P_{1}$ and $P_{2}$ respectively.
+Indeed, let us introduce the notations for the homogeneous decompositions that we will need to work with in the sequel:
+$P_{1}(z,w)$ $=P_{1}^{(2)}(z,w)+\ldots+P_{1}^{(r)}(z,w)$
+$P_{2}(z,w)$ $=P_{2}^{(2)}(z,w)+\ldots+P_{2}^{(k)}(z,w),$
+where $P_{1}^{(i)},P_{2}^{(j)}$ satisfy the following conditions:
+  • Either $P_{1}^{(i)}\equiv 0$ or $P_{1}^{(i)}$ is a homogeneous polynomial of degree $i$ for $i=2,\ldots,r$.
+  • Either $P_{2}^{(j)}\equiv 0$ or $P_{2}^{(j)}$ is a homogeneous polynomial of degree $j$ for $j=2,\ldots,k$.
+To prove that $Z$ can be straightened, it suffices to show that by a change of variables, we can reduce the arbitrary polynomial retraction map $F$ to a corresponding ‘normal form’.
+To do so, we need to solve the equation of idempotence: $F(F(z,w))=F(z,w)$; indeed, in the notations for the components of $F$ already introduced, this leads to the following identities for the afore-mentioned polynomials:
+(10.2) $P_{1}^{(2)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},\;P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\\ \;\Large\text{+}\;P_{1}^{(3)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\\ \hskip-144.54pt\Large\Large\text{+}\;\cdots\cdots\;\\ \Large\text{+}\;P_{1}^{(r)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},\;P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\;\;\Large\text{=}\;\;\Large\text{0}$
+and
+(10.3) $P_{2}^{(2)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},\;P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\\ \;\Large\text{+}\;P_{2}^{(3)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\\ \hskip-144.54pt\Large\Large\text{+}\;\cdots\cdots\;\\ \hskip 144.54pt\Large\text{+}\;P_{2}^{(k)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},\;P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\\ \Large{\Large\text{=}}\;\;P_{2}^{(2)}(z,w)+\ldots+P_{2}^{(k)}(z,w)$
+We now divide the arguments into two cases depending on whether $r=k$ or not.
+Case-1: $r\neq k$.
+In this case, note that all the highest degree terms of left hand side of (10.2) are contained within $P_{1}^{(r)}\left(P_{1}^{(r)},P_{2}^{(k)}\right)$.
+Write $P_{1}^{(r)}(z,w)=\sum_{i+j=r}a_{ij}z^{i}w^{j}$.
+Hence
+$P_{1}^{(r)}\left(P_{1}^{(r)},P_{2}^{(k)}\right)=\sum_{i+j=r}a_{ij}(P_{1}^{(r)})^{i}(P_{2}^{(k)})^{j}.$
+Each summand in the above sum is a homogeneous polynomial of degree ri+jk=ri+(ri)kri+jk=ri+(r-i)k, which we may write as: rk+i(rk)rk+i(r-k).
+Let N(i):=rk+i(rk)N(i):=rk+i(r-k).
+Note that on the ordered finite set of integers: $\left(0,1,\ldots,r\right)$, the function $N(i)$ is strictly monotonic, as $r\neq k$.
+Thus $N(i)$ attains maximum at a unique point $i_{0}$.
+So, the highest degree monomials of $P_{1}^{(r)}\left(P_{1}^{(r)},P_{2}^{(k)}\right)$ are all contained within the homogeneous polynomial $a_{i_{0}j_{0}}\left(P_{1}^{(r)}\right)^{i_{0}}\left(P_{2}^{(k)}\right)^{j_{0}}$.
+Equating the highest degree terms of the left hand and right hand sides of (10.2), we obtain: $a_{i_{0}j_{0}}\left(P_{1}^{(r)}\right)^{i_{0}}\left(P_{2}^{(k)}\right)^{j_{0}}\equiv 0$.
+Since $a_{i_{0}j_{0}}\neq 0$, we have $\left(P_{1}^{(r)}\right)^{i_{0}}\left(P_{2}^{(k)}\right)^{j_{0}}\equiv 0$.
+This implies by the identity principle that $P_{1}^{(r)}\equiv 0$ or $P_{2}^{(k)}\equiv 0$, which contradicts our assumption that $P_{1}\cdot P_{2}\not\equiv 0$.
+Thus, either $P_{1}\equiv 0$ or $P_{2}\equiv 0$, contradicting the assumption for the case being dealt with here.
+Case-2: $r=k$.
+In this case, note that $P_{1}^{(k)}\left(P_{1}^{(k)},P_{2}^{(k)}\right)\equiv 0$ as right hand side of (10.2) is identically zero.
+We also have $P_{2}^{(k)}\left(P_{1}^{(k)},P_{2}^{(k)}\right)\equiv 0$, this time owing to the fact that the highest degree of right hand side of (10.3) is at-most $k$ whereas the degree of every monomial in the polynomial $P_{2}^{(k)}\left(P_{1}^{(k)},P_{2}^{(k)}\right)$ is $k^{2}$ unless this polynomial vanishes identically.
+Consider then the holomorphic map $F_{h}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $F_{h}(z,w)=\left(P(z,w),Q(z,w)\right)$, where $P(z,w)=P_{1}^{(k)}(z,w)$ and $Q(z,w)=P_{2}^{(k)}(z,w)$.
+First we consider the sub-case that ${\rm image}(F_{h})$ has non-empty interior.
+Since $P(P,Q)\equiv 0,Q(P,Q)\equiv 0$, it follows by the identity principle that $P$ and $Q$ are identically zero, which contradicts our assumption that $P_{1},P_{2}\not\equiv 0$.
+Next we deal with the other sub-case that ${\rm image}(F_{h})$ has empty interior.
+First we want to prove that there exists $\lambda_{0}\in\mathbb{C}$ such that $Q(z,w)=\lambda_{0}P(z,w)$.
+Let $S$ denote the zero set of $F_{h}$.
+If $(\alpha,\beta)\in S$, then for each $\lambda\in\mathbb{C}$, $\lambda(\alpha,\beta)\in S$ – this follows from the fact that $S$ is the zero set of a map whose components are both homogeneous polynomials of the same degree and here is where the hypothesis of ‘case-2’ is used.
+Thus, $S$ is an union of complex lines in $\mathbb{C}^{2}$.
+Let $S^{\prime}:=\pi(S)$, where $\pi:(\mathbb{C}^{2}\setminus\{0\})\to\mathbb{CP}^{1}$ is the canonical quotient map.
+We observe that $S^{\prime}$ is a finite set in $\mathbb{CP}^{1}$.
+To verify this, let $z:w]\in S^{\prime}$.
+Without loss of generality, assume that $z\neq 0$.
+Dividing by $z^{k}$ throughout each of the equations $P(z,w)=0$ and $Q(z,w)=0$, we get one variable polynomials $f,g$ both of degree at-most $k$, such that: $f(w/z)=0$ and $g(w/z)=0$.
+It follows that the values of the ratio $w/z$ such that it is a root of $f$ (likewise of $g$) are at-most $k$ many; consequently, that $S^{\prime}$ is a finite set in $\mathbb{CP}^{1}$ and in-particular that $S^{\prime}$ is a closed subset of $\mathbb{CP}^{1}$.
+As $F_{h}(\lambda z,\lambda w)=\lambda^{k}F_{h}(z,w)$, the map $F_{h}$ on $\mathbb{C}^{2}$ induces a holomorphic map $\tilde{F_{h}}:\mathbb{CP}^{1}\setminus S^{\prime}\to\mathbb{CP}^{1}$ defined by $\tilde{F_{h}}([z:w])=[P(z,w):Q(z,w)]$.
+Note that $\tilde{F_{h}}$ is a rational map from $\mathbb{CP}^{1}$ into $\mathbb{CP}^{1}$.
+Hence $\tilde{F_{h}}$ induces a regular map from $\mathbb{CP}^{1}$ into $\mathbb{CP}^{1}$.
+Since $\tilde{F_{h}}$ is not surjective, $\tilde{F_{h}}$ is a constant map.
+This implies that $\rm{image}(F_{h})$ is a one-dimensional complex line passing through origin.
+This shows that there exists $\lambda_{0}\in\mathbb{C}$ such that $Q(z,w)=\lambda_{0}P(z,w)$.
+We shall now use this to reduce the present case to the former.
+To this end, we apply an affine-linear transformation of coordinates given by $\psi(z,w)=(z,w-\lambda_{0}z)$.
+Its inverse has a similar form; indeed, $\psi^{-1}(z,w)=(z,w+\lambda_{0}z)$.
+Under this change of variables, our retraction map $F$ is transformed to its conjugate by $\psi$, namely $G:=\psi\circ F\circ\psi^{-1}$, again a polynomial retraction map on $\mathbb{C}^{2}$.
+As we shall need it below to make a careful observation, we compute its explicit expression:
+$\psi\circ F\circ\psi^{-1}(z,w)=\psi\circ F(z,w+\lambda_{0}z)$
+$\begin{split}=\psi\left(z+P_{1}^{(2)}(z,w+\lambda_{0}z)+\ldots+P_{1}^{(k)}(z,w+\lambda_{0}z),P_{2}^{(2)}(z,w+\lambda_{0}z)+\ldots\right.\\ \left.+P_{2}^{(k)}(z,w+\lambda_{0}z)\right)\end{split}$
+$\begin{split}=\left(z+P_{1}^{(2)}(z,w+\lambda_{0}z)+\ldots+P_{1}^{(k)}(z,w+\lambda_{0}z),(P_{2}^{(2)}(z,w+\lambda_{0}z)+\ldots\right.\\ \left.+P_{2}^{(k)}(z,w+\lambda_{0}z))-\lambda_{0}(z+P_{1}^{(2)}(z,w+\lambda_{0}z)+\ldots+P_{1}^{(k)}(z,w+\lambda_{0}z))\right)\end{split}$
+Denote the analogues of the component polynomials (as with the case of $F$) in the above by $P_{1}^{\prime}$ and $P_{2}^{\prime}$; or in other words, precisely so that
+$G(z,w)\;=\;\left(z+P^{\prime}_{1}(z,w),\;P^{\prime}_{2}(z,w)\right).$
+Then notice that unlike $F$, the degrees of the components of $G$ are not both equal.
+Indeed, observe that $\deg(P^{\prime}_{1})=\deg(P_{1})$ but $\deg(P^{\prime}_{2})<\deg(P_{2})$, as follows by observing the highest degree homogeneous constituents in the expressions occurring in the last equation (particularly the second component of $G$), and recalling that $P_{2}^{(k)}=\lambda_{0}P_{1}^{(k)}$.
+It follows therefore that all the hypotheses required for case-1 above, is satisfied by the components of $G$ allowing us to appeal to the arguments therein to conclude that either $P^{\prime}_{1}\equiv 0$ or $P^{\prime}_{2}\equiv 0$.
+As $P^{\prime}_{1}=P_{1}\circ\psi^{-1}$ and $\psi$ is an automorphism of $\mathbb{C}^{2}$, we conclude that the former case $P^{\prime}_{1}\equiv 0$ implies that $P_{1}\equiv 0$ as well.
+The latter case $P^{\prime}_{2}\equiv 0$, however does not lead to a similar conclusion; in this case, recall that relation between $P_{2}^{\prime}$ and $P_{1},P_{2}$ is given by
+$P_{2}^{\prime}(z,w)\;=\;P_{2}\circ\psi^{-1}(z,w)\;-\;\lambda_{0}\left(z+P^{\prime}_{1}(z,w)\right)$
+Thus, the latter case $P^{\prime}_{2}\equiv 0$ means:
+$P_{2}\left(\psi^{-1}(z,w)\right)\;-\;\lambda_{0}P_{1}\left(\psi^{-1}(z,w)\right)\;=\;\lambda_{0}z$
+It is then clear that this equation cannot hold thereby giving the required contradiction to conclude that: the case $P_{2}^{\prime}\equiv 0$ cannot happen; indeed, the right side of the above equation has no non-linear terms whereas the left hand side of the above has no linear terms, so that such an equation can at-best hold only when both sides are zero which in-turn can happen only when $\lambda_{0}=0$, a contradiction (recall that $\lambda_{0}\neq 0$).
+Thus, the only conclusion of case-2 is that $P_{1}\equiv 0$.
+The common conclusion of the foregoing cases $1$ and $2$, is that either $P_{1}\equiv 0$ or $P_{2}\equiv 0$. Now, we unravel what each of these possibilities means for our retraction map $F$.
+If $P_{1}\equiv 0$, then the equation (10.3) reduces to
+(10.4) $P_{2}^{(2)}\left(z,P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)+\cdots+P_{2}^{(k)}\left(z,P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\;=\;P_{2}^{(2)}(z,w)+\cdots+P_{2}^{(k)}(z,w)$
+Examine the left hand side for all monomials of the lowest degree and observe that it may altogether expressed as: $P_{2}(z,0)$, provided of-course that this is non-zero, else the aforementioned lowest degree is higher than two.
+Therefore, equating the lowest degree term on both sides of the above equation (10.4), we get that $P_{2}^{(2)}(z,0)=P_{2}^{(2)}(z,w)$, unless it is the case that $P_{2}\equiv 0$.
+In both cases, we conclude that $P_{2}^{(2)}(z,w)$ is independent of $w$ i.e., a polynomial function of the single variable $z$ (possibly zero).
+Further in any case, by canceling-off these entities from both sides, equation (10.4) reduces to
+(10.5) $P_{2}^{(3)}\left(z,P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)+\cdots+P_{2}^{(k)}\left(z,P_{2}^{(2)}+\ldots+P_{2}^{(k)}\right)\;=\;P_{2}^{(3)}(z,w)+\cdots+P_{2}^{(k)}(z,w).$
+Again comparing the lowest degree terms on both sides of the above equation, we obtain that $P_{2}^{(3)}$ is at-most a function of $z$ alone i.e., either a cubic polynomial in $z$ or the zero polynomial.
+Continuing this process, we finally get that $P_{2}^{(2)},P_{2}^{(3)},\ldots,P_{2}^{(k)}$ are at-most functions of $z$ alone.
+This shows that $P_{2}(z,w)$ is a function of $z$ alone.
+Hence, we have proven that $F$ is of the form $F(z,w)=(z,\varphi(z))$ for some $\varphi\in\mathbb{C}[z]$; specifically, $\varphi(z)=P_{2}(z,0)$.
+Finally here, if we conjugate our retraction map $F$ by another automorphism – which we call $\psi$ again – $\psi(z,w)=(z,w-\varphi(z))$ and denote the resulting retraction again by $F$, we obtain that $F$ is the standard orthogonal projection onto the $z$-axis with ${\rm image}(F)=\{(\zeta,0):\zeta\in\mathbb{C}\}$, rendering the required rectification result for both the retraction map as well as the retract.
+If $P_{2}\equiv 0$, then the equation (10.2) becomes
+(10.6) $P_{1}^{(2)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},0\right)+\cdots+P_{1}^{(r)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},0\right)\;=\;0.$
+Equating the smallest degree terms of both sides of the above equation (10.6), we get that $P_{1}^{(2)}(z,0)=0$.
+This means that all the monomials in $P_{1}^{(2)}$ are divisible by $w$, thereby $P_{1}^{(2)}(z,w)=wQ^{(1)}(z,w)$, for some polynomial $Q^{(1)}(z,w)$ of degree one.
+Using this back in the above equation (10.6), we may reduce it to:
+$P_{1}^{(3)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},0\right)+\cdots+P_{1}^{(r)}\left(z+P_{1}^{(2)}+\ldots+P_{1}^{(r)},0\right)\;=\;0.$
+Again comparing lowest degree terms on both sides, we get that $P_{1}^{(3)}(z,w)\equiv wQ^{(2)}(z,w)$ for some polynomial $Q^{(2)}(z,w)$ of degree two.
+Continuing this process, finally we obtain that for each $i=2,\ldots,r$, there exists a polynomial $Q^{(i-1)}(z,w)$ of degree (i1)(i-1) such that $P_{1}^{(i)}(z,w)=wQ^{(i-1)}(z,w)$.
+Therefore, we obtain this time that without any further change of variables, our retraction map is precisely of the form
+$F(z,w)=(z+wQ(z,w),0).$
+for some $Q\in\mathbb{C}[z,w]$ which vanishes at the origin; specifically in the above notations, $Q(z,w)=Q^{(1)}(z,w)+\ldots+Q^{(r-1)}(z,w)$.
+Hence, ${\rm image}(F)=\{(\zeta,0):\zeta\in\mathbb{C}\}$.
+∎
+We can also derive the precise form of all polynomial retraction mappings of $\mathbb{C}^{2}$, as a corollary of proof of the foregoing theorem 1.25.
+Corollary 10.2.
+If $F:\mathbb{C}^{2}\to\mathbb{C}^{2}$ be a polynomial retraction map, then after a translation and a (isotropic) scaling, $F$ has one of the following forms:
+  1. (1)
+    $F(z,w)=\left(d(az+bw)-b\varphi(az+bw),-c(az+bw)+a\varphi(az+bw)\right)$,
+    where $a,b,c,d\in\mathbb{C}$ with adbc=1ad-bc=1 and $\varphi$ is a polynomial of one variable with $\varphi(0)=\varphi^{\prime}(0)=0$.
+  2. (2)
+    $F(z,w)=\left(dL(z,w),-cL(z,w)\right)$,
+    where $a,b,c,d\in\mathbb{C}$ with adbc=1ad-bc=1, $L(z,w)=az+bw+(cz+dw)Q(az+bw,cz+dw)$ and $Q(z,w)$ is a polynomial of degree greater than one.
+Finally, as mentioned above, we may use the above analysis to give an elementary proof of Costa’s result on the form of the polynomial ring/algebra in two variables.
+To this end, we begin by recalling some definitions; a good exposition of associated matter can be found in [[Shplrn]].
 Definition 10.3.
-A holomorphic map $\beta:\mathbb{C}^{2}\to\mathbb{C}^{2}$ is said to be a polynomial retraction if both components of $\beta$ are polynomials and satisfy $\beta\circ\beta=\beta$.
-Definition 10.3 ([Shplrn]).
-Let $K[x,y]$ be the polynomial algebra in two variables over a field $K$ of characteristic 00.
+Let $K[x,y]$ be the polynomial algebra in two variables over a field $K$ of characteristic 0.
 An idempotent $K$-algebra homomorphism $\varphi$ from $K[x,y]$ into itself is called a retraction (or a projection) map on $K[x,y]$.
@@
 Thus, a subalgebra $R$ of $K[x,y]$ is called a retract if there is an idempotent homomorphism $\varphi:K[x,y]\rightarrow$ $K[x,y]$ such that $\varphi(K[x,y])=R$.
-To provide just one concrete example to illustrate the above definitions, take $p=x+x^{2}y$; then $K[p]$ is a retract of $K[x,y]$ (the retraction map here, is given by $x\mapsto p$ , $y\mapsto 0$).
+To provide just one concrete example to illustrate the above definition, take $p=x+x^{2}y$; then $K[p]$ is a retract of $K[x,y]$ (the retraction map here, is given by $x\mapsto p$ , $y\mapsto 0$).
 Of-course as always, before proceeding to such examples, one can cite the trivial ones, namely $K[x,y]$ or $(0)$; a retract of $K[x,y]$ which is neither of these two extremes, is referred to as a proper retract in this context.
@@
 But this just means that $\rho\circ\rho=\rho$ i.e., $\rho$ is a polynomial retraction mapping on $\mathbb{C}^{2}$.
+Now we can prove theorem 3.5 in [Cost] as a corollary of the above theorem 1.25.
+Corollary 10.5.
+If $R$ is a non-trivial retract of $\mathbb{C}[z,w]$, then $R=\mathbb{C}[r]$ for some $r\in\mathbb{C}[z,w]$.
+Proof.
+Since $R$ is a non-trivial retract of $\mathbb{C}[z,w]$, there exists an idempotent homomorphism $\tilde{\rho}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ such that $\text{image}(\tilde{\rho})=R$.
+Let $p(z,w):=\tilde{\rho}(z)$ and $q(z,w):=\tilde{\rho}(w)$.
+By lemma 10.4, $\tilde{\rho}$ induces a polynomial retraction map $\rho:\mathbb{C}^{2}\to\mathbb{C}^{2}$ defined by $\rho(z,w)=(p(z,w),q(z,w))$.
+If we denote $\text{image}(\rho)$ by $Z$ then by above theorem 1.25, there exists a polynomial automorphism $\psi$ on $\mathbb{C}^{2}$ such that $\psi(Z)=\{(\zeta,0):\zeta\in\mathbb{C}\}$.
+Let $\hat{\rho}:=\psi\circ\rho\circ\psi^{-1}$ and $\hat{Z}:=\{(\zeta,0):\zeta\in\mathbb{C}\}$.
+Note that $\hat{\rho}$ is a polynomial retraction map on $\mathbb{C}^{2}$ and $\text{image}(\hat{\rho})=\hat{Z}$.
+Consider the induced homomorphism $\psi^{*}$ on $\mathbb{C}[z,w]$ defined by $\psi^{*}(z)=\psi_{1}(z,w)$ and $\psi^{*}(w)=\psi_{2}(z,w)$, where $\psi_{1},\psi_{2}$ are the components of $\psi$.
+It is easy to see that the surjectivity of $\psi$ implies the injectivity of $\psi^{*}$.
+While it is true that the injectivity of $\psi$ implies the surjectivity of $\psi$ is true, it is far more non-trivial.
+As our purpose here is to give a self-contained proof, we proceed without splitting it to directly verify that $\psi^{*}$ is an automorphism on $\mathbb{C}[z,w]$.
+Let $\phi(z,w)=(\phi_{1}(z,w),\phi_{2}(z,w))$ denote the inverse of $\psi(z,w)$.
+Consider its induced homomorphism, $\phi^{*}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ defined by $\phi^{*}(z)=\phi_{1}(z,w)$ and $\phi^{*}(w)=\phi_{2}(z,w)$.
+To prove $\psi^{*}$ is an automorphism, it suffices to show that $\psi^{*}$ and $\phi^{*}$ are inverses to each other.
+Note that
+$\phi^{*}\circ\psi^{*}(z)=\phi^{*}(\psi_{1}(z,w))=\psi_{1}(\phi_{1}(z,w),\phi_{2}(z,w))$
+and
+$\phi^{*}\circ\psi^{*}(w)=\phi^{*}(\psi_{2}(z,w))=\psi_{2}(\phi_{1}(z,w),\phi_{2}(z,w))$
+Since $\psi\circ\phi(z,w)=(z,w)$, we obtain that $\phi^{*}\circ\psi^{*}(z)=z$ and $\phi^{*}\circ\psi^{*}(w)=w$.
+Similarly, we get that $\psi^{*}\circ\phi^{*}(z)=z$ and $\psi^{*}\circ\phi^{*}(w)=w$.
+This shows that $\psi^{*}\circ\phi^{*}$ and $\psi^{*}\circ\phi^{*}$ fixes the generators $z,w$.
+As $\psi^{*}$ and $\phi^{*}$ are ring homomorphisms, we obtain that $\phi^{*}\circ\psi^{*}$ and $\psi^{*}\circ\phi^{*}$ are the identity maps on $\mathbb{C}[z,w]$, finishing the verification that $\psi^{*}$ is an automorphism.
+Getting back to the proof proper (of this corollary), since $\text{image}(\hat{\rho})=\hat{Z}$, there exists $s(z,w)\in\mathbb{C}[z,w]$ such that $\hat{\rho}(z,w)=(s(z,w),0)$.
+To conclude, we only need to consider the idempotent homomorphism $\rho^{*}:\mathbb{C}[z,w]\to\mathbb{C}[z,w]$ defined by $\rho^{*}(z,w):=\psi^{*}\circ\tilde{\rho}\circ\phi^{*}$; and, Note that $\text{image}(\rho^{*})=\mathbb{C}[s]$.
+This implies that $\psi^{*}(R)=\mathbb{C}[s]$ and hence we obtain that $R=\mathbb{C}[r]$, where $r(z,w)=\phi^{*}(s(z,w))$.
+∎
+It is not clear if the above results can be extended to the case when the number of variables is greater than two, even if we continue restricting to the polynomial case.
+We hope to be able to take this up in a future article soon.
 10.2. Proof of theorem 1.25
+We now deal with the other already known proof.
 This is essentially an exposition of the proof of theorem $1.1$ of [Shpilrain--Yu] with greater details.
@@
 $=$ $(\tilde{q_{1}}(\tilde{r}(z,w)),\tilde{q_{2}}(\tilde{r}(z,w)))$
-where $\tilde{q_{1}}=q_{1}-cq_{2}^{k},~{}\tilde{q_{2}}=q_{2},~{}\tilde{r}(z,w)=r(z+cw^{k},w)$. Note that $\tilde{q_{1}}(\tilde{r}),~{}\tilde{q_{2}}(\tilde{r})$ generates $\mathbb{C}[r]$.
+[…]
 Observe that the degree of the single variable polynomial $\tilde{q}_{1}$ is strictly less than that of $q_{1}$; we shall call the polynomials $\tilde{q}_{1},\tilde{q}_{2}$ now simply by $q_{1},q_{2}$ and repeat the above process of passing from one retraction to another through conjugation by the (new and appropriate) automorphism $\psi$.